A Complete Bibliography of Publications in Aquaculture International

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

13 April 2024 Version 1.12

Title word cross-reference

$39.95 [563]. 3 ± 1 [2011]. 4n − 6 [1224]. 4 ± 2 [2056]. > [2570]. ^{+} [92]. ^{−} [1131]. ^{1} [1504]. ^{15} [1594]. ^{2+} [1023, 1540]. ^{6} [1158]. ^{9} [1158]. ^{vp} [2605]. _{1} [1818, 2244, 2407, 2105, 1250]. _{2} [875]. _{2α} [1828]. _{3} [2190]. _{50} [1910]. α [784, 2378, 1816, 2032, 767, 116, 2281, 2368, 1729]. β [1995, 1993, 291, 2363, 2251, 1769, 2708, 1211, 1806, 1388, 2066, 2611, 594, 1185, 249, 1021, 1317]. β1 [2276, 2276]. G_{4} [2559]. γ [1851]. κ [1816]. n − 3 [934, 1117, 1241, 994, 1454, 1524]. ω [67]. Ω3 [46]. p [2502]. × [1750, 1635, 574, 551, 2618, 2297, 1921, 2733, 2748, 1188, 2371, 852, 1636, 2717, 1453, 585, 1508, 1097, 2482, 1631, 1732, 1155, 246, 1751, 1404, 2076, 2652, 2004, 2265, 1561].

-1 [1993, 2363, 1806, 1185, 1021]. -3 [67]. -ATPase [134]. -carotene [2066]. -Carotene-enriched [1995]. -coumaric [2502]. -estradiol [291, 1211]. -farmed [2638]. -gal [249]. -glucan [2251, 2708, 1388]. -glucans [1769]. -hydroxy- [594]. -induced [1250]. -infected [2110]. -K [134, 92].

-ketoglutarate [2032]. -lipoic [2281]. -mannanase [2611]. -methylbutyrate [594]. -sensitized [2280]. -tocopherol [784, 767, 116]. -tocopheryl [767].

/HIF [1816]. /HIF-1 [1816]. /keap1 [2732]. /macro [1465]. /vacuum [1576].

0 [1008, 1975]. 0-632-06406-4 [554]. 0-643-06865-1 [565]. 0-85199-596-9 [553]. 0-85238-276-6 [564]. 0-group [40]. 0D [997].

1 [1344, 2378, 678, 1556, 1154, 1287, 2441, 2440, 2559]. 1-56022-104-6 [563]. 1-57808-443-1 [678]. 1-84376-601-9 [552]. 1-chloro-7-methyl-2-propyl-1h-benzo [2387]. 10 [2334]. 10th [812]. 11-ketotesterone [2212]. 13 [574]. 15 [2216]. 15-year [2086]. 17-3 [2642]. 17^{o} [1297]. 18-year [1407]. 19 [2599, 2507]. 1A [2592]. 1b [2164]. 1h [2387].

2 [2280, 1556, 2679, 2401, 2278]. 2-methylisoborneol [1492]. 2-phenoxyethanol [578]. 20^{o} [1297]. 20th [1518]. 222 [2539]. 24L [997]. 281 [574]. 2nd [231, 692, 716, 860, 862, 992, 1285, 680].

3 [2148, 2276, 2642]. 3-glucan [1806, 1185, 1021]. 3-month [1530]. 3/1 [1993, 2363]. 3102 [2151, 2076]. 3D [1246]. 3rd [736].

4 [2148, 2680]. 4-3 [2148]. 4-endoglucanase [896]. 48-h [1472].

5 [1374]. 5-HT [1374]. 5.8S [895]. 53 [2389]. 53-kDa [2389]. 5M [1975].

6 [1592]. 6-glucan [2363]. 6-glucans [1993]. 6w [165]. 6w-3 [165].

8 [926]. 8-cineole [2321].

978 [1008, 1975]. 978-1-78918-048-0 [1975]. 978-1-84971-259-0 [1008]. 9th [861].

Îles [138]. Îles-de-la-Madeleine [138].

A. [1826, 700, 957, 1283]. Aarset [708]. Abalone [514, 2029, 1085, 1673, 2344, 792, 1670, 370, 1813, 1701, 1115, 459, 1589, 2623, 1681, 2220, 2405, 1131, 896, 710, 1034, 1504, 185, 1367, 2465, 2652, 1875, 1871, 1327, 851, 1545]. abalones [1172]. abce1 [2534]. abdominalis [544, 556]. abilities [1680]. ability [1218, 2297, 1860, 1083, 2641, 2747, 355, 2742, 2536]. abiotic [1498]. Abnormal [104, 2547]. Abnormalities [519, 1642, 1651]. Abramis [489, 585]. abrogate [1639]. absorption [712, 773, 2076]. Abundance [271, 957]. acai [2569, 2753, 2100]. Acanthocephala [2613].

Acanthopagrus [2180]. Acanthus [2680]. Acartia [1240, 1310, 1472, 1624, 1771, 1632]. accelerate [1489]. acceleration [1544]. acceptance [1571]. access [299]. acclimated [1118]. acclimation [1635, 1315, 1794]. according [1037]. accretion [2651]. Accumulation [1120, 1913, 1784, 1629, 2742, 1035, 2091]. acei [1077]. acetate [1356, 767, 2180]. acetic [2219]. acetylcholinesterase [2053]. achievement [1407]. achieving [2504]. Acid [250, 671, 2173, 1907, 562, 2502, 1255, 1696, 1621, 2414, 1750, 1098, 2062, 1133, 444, 1121, 2190, 910, 1353, 266, 2651, 784, 2128, 899, 1173, 1860, 1028, 1784, 1813, 1141, 1188, 1999, 1109, 1841, 1552, 912, 994, 915, 1311, 821, 1643, 1014, 1224, 1263, 2719, 2526, 1803, 2156, 37, 2097, 1684, 2311, 1147, 2473, 428, 501, 2177, 165, 983, 767, 1886, 1773, 982, 373, 634, 464, 1767, 2697, 17, 1096, 1751, 1432, 331, 1296, 620, 2180, 1163, 900, 2412, 202, 1505, 2594, 2281, 2562, 1653, 959, 1348, 930, 1316, 2383, 204, 2300, 1694, 164, 1029]. acid [1662, 432, 914, 2087, 1425, 1106, 1581, 2219, 1405, 2143, 425, 1545]. acid-binding [2097]. acid-fermented [2473]. acid-supplemented [1886]. acidic [2565]. acidification [1913, 2196, 1633]. acidifier [2335, 1736, 2005]. acidifiers [1888]. Acids [239, 2560, 867, 934, 2461, 818, 1117, 1241, 824, 2335, 994, 1231, 2218, 1097, 2035, 46, 667, 1454, 2588, 100, 1524, 1900, 135, 669, 1128, 2219]. Acipenser [2702, 2618, 1227, 1328, 1903, 2024, 994, 1094, 1848, 1247, 2212, 875, 776, 1232, 750, 483, 1857, 2219, 2265]. Acipenseriformes [454]. acoustic [2603, 2188]. Acromitus [2636]. across [2651, 2547, 1279, 2353]. Acrossocheilus [2534]. Actinidia [2390]. actinomycetes [823]. activated [1319]. activating [1223, 1444]. activation [869]. active [1087]. activities [2063, 1957, 1708, 2513, 2003, 410, 2676, 463, 1560, 2322, 361, 1267, 2621, 2518, 2035, 1063, 925, 913, 1145, 92, 2000, 1108, 1735, 905, 1649]. Activity [506, 484, 2215, 4, 255, 1696, 1619, 549, 2313, 160, 1626, 1715, 2471, 924, 134, 258, 2609, 987, 2169, 1563, 1399, 2110, 1501, 2452, 1580, 2229, 2387, 2328, 1048, 1486, 1882, 1916, 2428, 932, 2473, 2390, 2675, 2619, 1254, 380, 2457, 2679, 2591, 2632, 1091, 840, 197, 2462, 1138, 1653, 2073, 2677, 1348, 85, 950, 386, 2720, 2565, 735, 782, 1727, 1592, 2318, 2711, 2424]. actors [2068]. actual [1276]. Acute [570, 1037, 2053, 1259, 1228, 2690, 2153, 1599, 2000, 513, 2297, 2605, 2227, 1510, 2218, 1441, 2576, 2463, 400, 2010, 24, 360, 1299, 1438, 2655, 2163, 2112]. acyl [1780]. Adams [448, 439]. adaptability [1174]. Adaptation [500, 1997, 1046]. adaptive [1261]. Add1 [2373]. added [1465, 298]. Addition [193, 341, 2516, 2370, 2304, 2355, 2608, 1685, 732, 1376, 1221, 1222]. additive [2246, 2455, 2399, 2346, 898, 2082]. additives [2385, 1766, 2372]. address [2239, 1375]. addressed [1747]. adenylate [1444]. adequate [1271]. adhesion [2641, 2105, 2375]. adjusting [2359]. adjuvant [1638, 1603]. administered [1901, 1809]. Administration [507, 2612, 2315, 2736, 2169, 1048, 1301, 2526, 2377, 1063, 1363, 1185, 1069, 1867, 2361, 1105, 1387, 1021]. adopted [1452]. Adoption [2141, 2423, 2134, 2445]. Adriatic [1574, 815, 1602, 534, 806]. adsorbents [1250]. Adult [568, 2310, 2090, 804, 2395, 2332, 202, 2548, 1443]. adults [721, 707, 2081]. Advancements [1605]. Advances [347, 929, 2443, 309, 1330]. Advancing [2517, 2570]. Aegle [2736]. aeglefinus [430]. aeneus [1868]. aequans [38, 441]. Aequipecten [332]. Aerated [1483, 294]. aeration [1960, 2233, 2537, 2236, 2422, 2069, 2489, 2065]. aerators [1598]. aerial [1437]. Aerobic [2196, 2267, 1970, 96]. aerolysin [2149]. Aeromonas [1986, 1033, 1638, 1790, 2567, 2215, 2502, 2666, 1835, 2336, 2579, 2014, 885, 2025, 2732, 2232, 1912, 1474, 987, 1999, 2071, 86, 1169, 2075, 1137, 2587, 2009, 2341, 87, 2034, 1388, 2675, 2482, 1998, 2200, 2151, 2261, 2353, 2290, 2058, 2037, 395, 1333, 1805, 1010, 2677, 1663, 2333, 2470, 1437, 2402, 2749, 2149, 2247, 2582, 2500, 1922, 2639, 2061]. aeruginosa [1182, 792, 2342, 2161]. AFB [1818]. aff. [164]. affect [2673, 2085, 63, 381, 1779, 590, 2291, 2106, 1952, 2270, 1476, 469, 2697]. affected [673, 532]. affecting [2494, 2317, 2233, 1074, 2333, 2283, 2149]. affects [1309, 291, 1474, 2354, 1090, 1927, 2431, 2113]. affine [2087]. affinis [1797]. Aflatoxin [1818, 2407, 1250]. AFLP [528, 1506, 874]. afra [2034]. Africa [2269, 683, 2721, 1368]. African [1888, 128, 492, 2142, 1670, 545, 546, 2626, 2752, 200, 813, 205, 206, 66, 65, 1589, 877, 993, 2118, 2405, 2498, 2200, 346, 1714, 2281, 185]. after [146, 1140, 1569, 1817, 1213, 2331, 2572, 1239, 1640, 29, 426, 1288, 2200, 797, 47, 473, 1717, 2531, 1258, 2100, 1105, 1922, 2629]. against [1957, 1033, 1638, 2192, 2215, 2109, 1801, 2604, 1619, 2618, 2095, 1498, 2437, 2643, 2471, 1676, 2609, 2736, 1474, 2460, 987, 1563, 2545, 2553, 1467, 2641, 915, 1169, 1417, 2229, 2387, 2428, 1469, 855, 794, 1914, 1925, 2341, 1757, 2308, 1388, 2329, 1872, 2443, 2694, 2482, 925, 1998, 1732, 1704, 2457, 465, 1315, 2592, 739, 2647, 2150, 2632, 1087, 2058, 2037, 2105, 1333, 2577, 2539, 2677, 1059, 2333, 2470, 1711, 1132, 1962, 82, 950, 1361, 1437, 1182, 2480, 1438, 1403, 2152, 1727, 1592, 2658, 2639, 1731]. agalactiae [2653, 1919, 2592, 2181, 1809, 1815, 2234, 1731]. agar [1673, 879]. agar-bound [1673]. Agardh [717]. agarophyte [2417]. Agarophyton [2295]. Age [248, 515, 126, 1439, 659, 462, 1002, 133, 1466]. age-structured [1002]. agent [2560, 1915, 1323, 2209, 2257, 1260]. agent-based [1915]. agents [1460, 2150, 407, 1161]. ages [1722]. agglutinating [85]. aggregative [2105]. aggression [313]. aggressiveness [1948, 1947]. agonistic [1266, 65]. agonists [255]. agricultural [2366]. Agriculture [678, 2286]. Agrimos(R) [1687]. agro [2052, 1385, 1386]. agro-ecological [1385]. agro-ecosystem [2052]. agro-industrial [1386]. Ah17 [2058, 2037, 2402]. ahead [1416]. AHPND [2604, 1510, 2162, 2209, 2576, 2010, 2163]. AHPND-positive [2162]. AI [2750]. AI-driven [2750]. Ailian [1102, 1272]. aiming [2558]. air [439, 279, 1576, 280, 177, 1112, 672, 2069]. air- [1576]. air-/vacuum-packaged [1576]. air-breathing [1112]. akaara [1364]. Akazaki [1503]. AKG [2032]. Akvakultur [708]. alba [1252]. Albendazole [2254]. albiflora [1350, 2130]. albino [2619]. albumin [1217]. albus [1267].

Alcalase [1927]. alevins [2303, 2420]. alexandri [1884, 2568]. alga [1041, 2388, 2400]. algae [1356, 1913, 2509, 396, 2331, 652, 1676, 1501, 199, 2165, 1471, 1084, 2013, 718, 2588, 2191, 892]. Algal [232, 2664, 1310, 2170, 1347, 1798, 289, 2683, 1820, 327, 36]. algal-based [36]. Algarve [391]. Algeria [2155]. Algerian [1696]. alginate [2604, 1261, 2515, 2171, 209, 2682]. alginate-chitosan [2515]. alginate-coated [1261]. Alginic [1773]. alginolyticus [1493, 2031, 1479, 1757, 1429, 1732, 1704, 941]. algorithm [1915, 1349]. alien [714]. alimentation [1012]. alkaline [2463, 971]. alkalinity [1381]. all-female [455]. alleviate [1250]. alleviated [2031]. alleviates [2662]. Allium [2593, 950, 1727]. allocation [80]. allochthonous [1420]. allometric [1209]. almond [2516]. almyra [358]. Aloe [1638]. alone [150, 1652, 2045, 767]. along [1565, 1756, 1755, 2417, 1433, 1695, 2528, 2120]. alpha [544]. Alpinia [1384]. alpinus [481, 1882, 1233, 667]. alter [1883]. Alteranthera [1252]. Alteration [1708]. Alterations [233, 2225, 2009, 2159, 983]. altered [1420]. altering [899]. Alternanthera [2601, 1616, 2720]. Alternative [1754, 1673, 1859, 2293, 2427, 712, 842, 2558, 956, 417, 1846, 2481, 2507, 2726, 407, 2616, 1294, 789, 1650]. alternatives [1295]. alters [2156]. altivelis [866]. alvarezii [2619, 893]. Alzheimer [2635]. amaranth [1989, 2525]. Amaranthus [1989]. Amazon [1971, 1709, 2460, 1964]. Amazonian [45, 2312]. amazonicum [1971]. Amberjack [247, 40, 1084, 694]. ambient [1035, 1430]. amblycephala [1530]. amboinensis [1025]. Ameliorative [2075, 2428]. amended [2631]. America [2364]. American [1789, 1844, 1421, 283, 2250, 2629]. americanum [1823]. americanus [1844, 1421, 509, 2211]. Amino [1128, 671, 562, 2190, 824, 1231, 2218, 2719, 2035, 1886, 1432, 1900, 1316, 135, 425]. Ammonia [1536, 1493, 664, 2278, 1037, 2610, 924, 12, 2367, 1228, 1351, 1198, 2633, 2314, 935, 2365, 2322, 2642, 1143, 1447, 1315, 400, 2537, 704, 972, 2397, 1430, 443, 592, 745, 2112]. ammonia-N [1351]. Ammonium [2195]. Among [487, 2093, 2014, 1558, 1564, 768, 86, 675, 1742, 1984, 644, 884, 103, 1539, 831]. Amphilophus [1636]. Amphimictic [487, 586]. Amphioctopus [1694]. amphipods [2172]. Amphiprion [2436, 1124, 2136, 2362, 1536, 1554]. amplification [2401, 2149]. amplification-lateral [2401]. amplified [86]. AMPNT [1592]. AMPNT-6 [1592]. Amur [1857, 2324, 2368]. amurensis [2359]. Amusium [283]. amylase [2169]. amyloliquefaciens [1372]. Anabas [2394]. anadromous [134]. anaerobic [2267, 840, 2167]. Anaesthetic [1601, 1599, 785]. analogue [588, 1160]. analogues [1215]. analyses [2269, 1459, 2726, 2688]. Analysing [947]. Analysis [1838, 86, 1681, 835, 2068, 665, 542, 2710, 656, 866, 2687, 2754, 2389, 2135, 2701, 2579, 1853, 1670, 849, 1202, 2499, 2325, 2221, 302, 2227, 1634, 2733, 2086, 23, 1491, 954, 2712, 1298, 1564, 2378, 2564, 66, 1080, 2442, 889, 1329, 2165, 1692, 2657, 2644, 2055, 1585, 2721, 765, 576, 2498, 2527, 1026, 2272, 2080, 2332, 2650, 1667, 1007, 2349, 2271, 2674, 1933, 1273, 1389, 2309, 2044, 1707, 1693, 2275, 2397, 2620, 2665, 2546, 607, 2222, 2168, 2174, 2368, 1660, 2217, 2077, 2692, 1373, 1728, 2234, 851, 2208, 1512, 380]. analyze [2073]. Anarhichas [1826, 408, 715]. Andaman [1113, 1112, 858]. Andhra [2668]. androgen [1000]. androgen-treated [1000]. androgenesis [1359]. androgenic [822, 293]. Andrographis [472]. anemia [1993]. anemone [2471]. anemone-isolated [2471]. anemonefish [1124]. anesthesia [1314]. Anesthetic [2321, 513, 629, 2419, 2257, 2357]. anesthetics [2118, 578, 1884]. anesthetized [2224]. ANFIS [2682]. angelfish [1375, 2712, 767]. Angelica [985]. Angilla [387]. angiogenic [2664]. Anguilla [1049, 2147, 2050, 701, 2256, 1093, 1177, 758, 1507, 1296, 387]. anguillarum [42, 1780, 1833, 1132, 1438]. anguillicaudatus [1157, 1156, 2244, 724, 741, 814, 831]. angulata [2735]. anhydrase [92]. Animal [700, 1671, 94, 365, 1283, 114, 277, 1718]. animals [872, 1071, 1390, 785]. Annelida [2528]. anniversary [1518]. Announcement [130, 187, 214, 224, 231, 308]. annua [1989]. annual [1710]. annulatus [435]. anomalies [302]. Anostraca [1696, 1754]. antagonisms [107]. Antagonistic [2677]. antagonists [1215]. Anthocidaris [1318]. anthocyanins [1991]. Anti [2387, 2663, 2633, 1916, 2220, 2636, 2679, 2216, 2420]. anti-apoptotic [1916]. Anti-bacterial [2663]. anti-infective [2216]. anti-oxidant [2420]. anti-oxidative [2220]. anti-proliferative [2636]. anti-stress [2633]. anti-VHSV [2679]. Anti-viral [2387]. Antibacterial [915, 2675, 925, 1132, 950, 2215, 1619, 472, 2471, 2341]. antibiofilm [2663]. antibiotic [2342, 2380, 2306, 343]. antibiotics [1294]. antibodies [1387]. antibody [1321, 1835, 1277]. antibody-based [1321, 1835]. anticipation [890]. anticoagulants [2348, 2224]. Antifungal [2457, 2621, 2390]. Antigen [1261, 1140]. antigens [2111]. antillarum [2503]. Antimicrobial [1957, 1998, 1696, 2466, 2185, 407, 1592]. Antioomycete [2565]. Antioxidant [2215, 1626, 2008, 331, 2370, 2560, 2552, 2147, 2618, 2437, 1680, 1860, 1552, 1417, 2328, 1882, 2428, 1914, 2672, 2032, 2395, 2463, 361, 359, 2329, 2013, 2514, 1254, 2491, 1478, 2595, 2262, 2479, 2372, 1978, 2281, 844, 2562, 1059, 436, 2361, 1735, 2324, 2435, 2495, 2382, 2742, 2745, 1657, 2536, 2500, 2646, 1649, 2658, 2569, 2753, 2686]. antioxidant-related [2370, 2372]. antioxidants [2035, 1221, 837]. antioxidation [1139]. antioxidative [1681, 2634]. Antiparasitic [2315, 2063]. Antiviral [2110, 2229, 2632, 1361]. Antofagasta [1306, 605]. any [2516]. aPDT [2185]. apiculata [622]. apidaecin [746]. apoptosis [1268, 2365, 2278]. apoptotic [1916]. Apostichopus [788, 1708, 2607, 1239, 1606, 839, 1423, 1546, 1547, 1243, 1003, 1192, 1299, 1483, 1607, 1517, 1318, 1373, 1494, 870, 1343, 2662]. Apparent [178, 200, 243, 2719, 762, 2318]. Application [970, 1655, 1289, 191, 2684, 1461, 1470, 2171, 166, 875, 2391, 2715, 2728, 1060, 1223, 2143, 1906, 895, 2240, 1793, 1878, 1427, 1944, 2487, 1497, 1197, 2443, 1639, 2257, 2412, 2044, 1991, 1114, 2339, 2628, 2082]. Applications [2422, 2215, 129, 258, 2400, 955, 1837]. applied [995, 2338, 2744, 2355, 294, 330]. appraisal [2052, 2144]. approach [1058, 2331, 682, 1427, 1549, 1747, 2055, 536, 1454, 2506, 2561, 2507, 978, 1746, 2337]. approaches [1637, 1544, 2163, 553]. approaching [2582]. appropriate [2707, 1924]. APRIL [2137]. AQUA [1416, 1407, 1413, 1408, 2117, 1412]. AQUA-TNET [1407, 1408, 2117]. Aquacult [574]. aquacultural [2475]. Aquaculture [217, 264, 498, 490, 238, 1518, 411, 506, 1600, 571, 2172, 347, 563, 887, 503, 2052, 274, 1551, 413, 1524, 1489, 2215, 54, 1917, 2413, 1075, 2067, 2001, 2360, 2036, 2269, 2451, 53, 819, 2007, 2423, 424, 1520, 1286, 706, 1463, 1611, 795, 1859, 1001, 1906, 2474, 1557, 1529, 2519, 1960, 541, 836, 2385, 2258, 2509, 2398, 2512, 2366, 1709, 1740, 2670, 129, 2673, 1829, 560, 2494, 1407, 1410, 531, 2558, 1525, 1865, 529, 467, 1061, 652, 1575, 2638, 1323, 1862, 1304, 2367, 1637, 1456, 1380, 2198, 1176, 1648, 1115, 1347, 1688, 1845, 2134, 2376]. aquaculture [2, 2627, 698, 1878, 415, 2027, 1241, 2256, 1198, 1564, 56, 1861, 1985, 1068, 575, 2493, 2184, 1521, 2392, 3, 365, 956, 974, 1284, 1837, 1756, 1755, 1391, 1606, 1580, 2705, 414, 683, 1647, 2064, 1608, 1747, 1738, 1337, 2574, 715, 926, 2671, 984, 1408, 889, 1308, 2403, 1546, 2092, 2411, 1349, 1533, 1953, 1460, 1808, 1492, 2657, 1079, 1593, 809, 2570, 747, 2022, 1850, 624, 1519, 1325, 2443, 2696, 2597, 2541, 2466, 74, 1631, 2252, 517, 518, 1787, 1279, 1605, 2124, 2041, 1984, 1386, 1893, 1686, 1078, 1031, 1825, 1027, 1400, 2551, 2292, 710, 634, 1144]. aquaculture [1411, 1911, 1973, 2506, 1639, 1378, 2561, 790, 1842, 2236, 2422, 2584, 1937, 1834, 1365, 41, 1426, 2259, 1414, 1415, 2412, 2715, 1877, 973, 2677, 418, 2384, 1968, 1935, 1683, 284, 2730, 2640, 1659, 491, 1095, 1710, 1225, 1294, 2012, 2678, 542, 1462, 1746, 2326, 1051, 686, 2692, 543, 2337, 1650, 2186, 1424, 1988, 2453, 2201, 691, 678, 692, 700, 862, 1070, 1071, 1151, 1152, 1436, 689, 725, 992, 1975, 329, 296, 353, 354, 774, 446, 625, 709, 1009, 1204, 1282, 1905, 553]. aquaculture-algae [652]. Aquaculture-based [2052]. aquafarming [15]. aquafeed [2202, 1497, 2074, 2082]. AquaMats [1198]. Aquaponic [627, 2421, 2131, 1514, 2446, 1989, 1383, 2452, 520, 521, 1497, 1495, 1677, 2364, 2504, 2374]. aquaponically [1951]. Aquaponics [2479, 2740, 2433, 2573, 2239, 1836, 1672, 1963, 2468, 2158, 2351, 2386, 2345, 2570, 2041, 1811, 1760, 2250, 2167]. aquaporin [1315]. aquarium [2516, 1924, 847]. aquarium-scale [847]. aquariums [1375]. Aquatic [688, 700, 785, 448, 1631, 128, 1671, 2601, 1416, 1413, 1071, 624, 2317, 2270, 1414, 1616, 1390, 1412, 2669, 2077]. aquatica [2740, 1495]. aqueous [2009, 2153]. Arabia [2530, 133, 2217]. Arabian [2530, 1695]. Arachidonic [432, 1224]. Arapaima [2348, 2545]. Arca [1144]. archipelago [1032]. archived [1969]. Arctic [25, 299, 481, 3, 50, 188, 7, 1127, 1882, 322, 91, 1233, 667]. arcticus [2530]. ardeolae [2379]. area [832, 2555, 1380, 2002, 1485]. areas [1917, 1906, 370, 94, 417, 829, 1183, 1112]. Arenas [1306]. areolata [827, 910, 909, 946, 979, 1001, 631, 756, 2477]. areolate [548]. Arg [1158]. argenteus [2228]. Argentina [978]. Argentine [681]. arginine [1116, 1081].

Argopecten [728, 605, 278, 402, 259, 348, 303, 2429, 626, 284, 397, 967]. argulosis [1608, 2600]. Argulus [2600]. argus [2579, 1914, 1561]. Argyrosomus [1017, 1270, 759, 445, 2157]. arid [2597]. Aristichthys [2426, 1816, 1815]. Ark [1144]. aromatase [1262]. Arosa [103]. array [2652]. arrow [2654]. Artemia [2370, 2088, 350, 166, 210, 2531, 1995, 68, 1696, 2617, 687, 796, 39, 1780, 546, 13, 1326, 1498, 435, 833, 1901, 122, 781, 757, 515, 647, 915, 1263, 1560, 341, 1823, 78, 433, 1063, 524, 770, 798, 465, 434, 67, 2462, 900, 116, 1199, 1132, 2300, 1025, 1795, 1466, 89, 2211]. Artemisia [1989, 2034, 2357]. Arthrospira [2506, 2286, 2388, 2436]. Article [1467, 1385, 2252]. articles [859, 1285]. Artificial [568, 2050, 902, 948, 1418, 9, 1817, 712, 966, 1532, 314, 1399, 2614, 830, 1856, 2580, 1941, 2057, 2203, 2405, 875, 1359, 1777, 964, 1296, 1365, 2289, 1955, 741, 1537, 2432, 2174]. Artificially [479, 2585]. arvensis [1677]. Asche [992]. ascidian [1606]. Ascorbate [204]. Ascorbic [1353, 1999, 1750, 1121, 266, 428, 767, 1751, 331, 204]. ash [1595]. Asia [2317]. Asian [1645, 988, 2583, 2442, 2343, 2329, 1691, 906, 1881, 2674]. asiatica [2041]. asinina [2029, 1673, 1612, 370, 2623, 1367, 2465, 1545]. asinina-fed [2465]. Aspects [610, 819, 1208, 2531, 1577]. assay [1835, 1441, 1476, 2401, 1697, 2149, 2374]. assays [301, 1910]. assemblage [2509, 657]. assemblages [1584, 1167]. assess [829, 2603, 1107]. assessed [954]. Assessing [2580, 1497, 758, 1031, 1752, 1876, 542]. Assessment [2413, 562, 1531, 1619, 1463, 301, 769, 998, 1198, 56, 1401, 1856, 2452, 2327, 2244, 1787, 427, 1183, 514, 892, 2646, 344, 2287, 1520, 1748, 1565, 59, 2351, 1854, 902, 50, 716, 2055, 2417, 96, 1073, 2407, 20, 1866, 2264, 2396, 1185, 2454, 2012, 1051, 2204]. asset [2466, 2349]. assignment [1187, 2244]. assignments [256]. assimilation [712, 773, 1862, 2584, 704]. assisted [656, 1273]. Associated [2285, 2666, 792, 1669, 1344, 1703, 1061, 2690, 1201, 429, 1429, 2194, 2028, 1987, 2528, 1881, 2102, 2383, 2116, 580, 1190, 686, 2393]. Association [945, 1444, 1670, 1949, 1707]. assurance [2240]. assured [284]. Astacidae [638, 833, 1117]. Astacus [304, 273, 406, 723, 752, 876, 1117, 1266, 917, 1426]. Astaxanthin [1914, 1446, 2437, 2733, 1124, 975, 2220, 2210, 2040, 2658]. astaxanthins [1139]. Asterias [1024, 2359]. astragalus [2114, 1680]. Astronotus [1446]. Aswan [2379]. ATCC [2045]. ater [1306]. Atlantic [313, 1621, 743, 109, 174, 635, 890, 16, 1140, 1309, 942, 227, 989, 639, 81, 155, 2651, 1566, 63, 228, 779, 1715, 1993, 12, 1865, 1013, 31, 49, 1314, 1110, 838, 1439, 195, 2637, 768, 1313, 2544, 22, 171, 29, 123, 2583, 1046, 960, 1395, 976, 361, 359, 193, 2657, 1073, 226, 46, 181, 1431, 1434, 1967, 2695, 92, 547, 394, 137, 1591, 431, 1042, 2724, 425, 2112]. ATPase [134, 92]. atrazine [916]. attacks [656]. Attempt [479, 314, 2110]. attenuated [87, 1872]. attitudes [945, 499, 1865, 2117]. attractants [2235, 2738]. attributes [405]. atypical [395]. AU$89.95 [565]. auction [1057]. augmented [69]. Aulacomya [1306]. Aulonocara [2118]. aurata [882, 535, 2133, 1626, 364, 1622, 390, 1565, 405, 2367, 2197, 1486, 1349, 770, 798, 1332, 339, 246, 464, 1535, 898, 510, 2496, 844, 557, 1089, 789, 1029]. auratus [738, 2463, 1500, 2469, 1018, 2587, 1925, 2526, 2511, 1495, 1155, 2622, 2105, 2333, 1991, 2339]. aureolum [143]. aureus [1956, 1188, 1453, 2152, 1592]. Austin [2385, 1254]. Australia [747]. Australian [1018, 115, 114, 388, 295, 2283, 565]. australiense [512]. australis [2552, 654]. Austropotamobius [345, 369]. Author [170, 215, 262, 315, 375, 419, 475]. Authors [225, 317]. autochthonous [1879, 1069]. autoclaved [1141, 1200]. autoethanogenum [2707]. Autogenous [1990]. autointoxication [664]. autolysis [2662]. Automated [494, 973, 1968]. Automatic [2253, 2352]. autotrophic [1686]. Availability [182, 1310]. Available [565, 1740, 1339]. avenue [1671]. aversive [890]. avian [1218]. avidus [1806]. avoids [2195]. awatschensis [1035]. axis [2749]. Azotobacter [191, 384].

B [411, 307, 552, 330, 366, 563, 690, 785, 860, 1039, 1816, 1818, 2407, 1250, 185, 2275]. B-like [2275]. B. [1253, 708, 625, 2694]. babylon [548, 827, 910, 909, 946, 979, 631, 756]. Babylonia [548, 827, 910, 909, 946, 979, 1001, 631, 756, 2477]. babylons [1001]. Bacillus [2131, 1668, 1652, 1256, 2319, 2148, 2169, 1563, 2045, 2641, 2184, 1048, 1137, 1126, 1449, 2314, 1757, 2463, 2608, 2316, 2675, 2694, 1063, 2151, 1184, 1372, 1873, 2105, 1646, 2076, 1182, 1966, 1735, 1517, 735, 782, 1883, 2245, 2375, 1731, 2176]. background [1924]. Bacteria [232, 348, 580, 1621, 2043, 2427, 1992, 1862, 1878, 321, 1263, 1420, 2526, 1684, 341, 429, 1814, 925, 2177, 2647, 306, 107, 2412, 1390, 1705, 82, 2089, 1494]. Bacterial [1515, 102, 14, 241, 386, 1883, 2540, 1085, 385, 2358, 1140, 472, 2385, 1377, 2019, 258, 1766, 1701, 1564, 1861, 889, 2411, 1808, 298, 1630, 1686, 1078, 1902, 1994, 2525, 1717, 1639, 2645, 1697, 2663, 1258, 107, 289, 2577, 2539, 362, 2102, 121, 2119, 972, 1705, 2300, 1105, 2575]. bacterial-based [972]. bactericidal [259]. bacterin [2572]. bacterins [1990]. bacteriocin [2105]. bacteriocin-producing [2105]. Bacteriophage [2557, 1078, 1323]. Bacteriophages [2124, 2448]. bacterium [2684, 897, 1970, 1805]. baculo [140]. baculo-like [140]. baerii [2618, 2024, 1247, 2219, 2265]. BAFF [2137]. bag [1601]. bagasse [2513, 2631]. bagrid [268]. Bahamas [1031]. Bahia [169, 566]. baicalensis [2311]. Baikal [2311]. bailiniae [1253]. Baird [1797]. Bait [1844]. Bait-subsidized [1844]. baits [1355]. Baja [169, 566]. Baker [625]. bakery [1033]. balance [1586, 922]. balances [1020]. Balearic [247]. Balss [731]. Baltic [1656, 163, 148, 211]. Ban [2231, 1786]. Ban-V [2231]. Banana [1274]. banded [1286]. Bangiales [895]. Bangladesh [2266, 2199, 848, 1015, 2413, 1721, 673, 1058, 2517, 2360, 2138, 1005, 2376, 682, 1549, 2602, 2476, 1647, 2574, 963, 1898, 2141, 2080, 2350, 2391]. bank [1534]. bannerfish [1375]. barb [2253, 2296, 2068]. barbata [2175]. barbatus [1144]. barbel [1215, 1162]. Barbier [552]. Barbonymus [2296]. Barbus [1215, 1162, 1484]. barcoo [2461, 1466, 1744]. barfin [1876]. barley [1658, 1783]. barnacle [995]. barramundi [997]. barren [2410]. barrier [2336, 2749]. barriers [2239, 1911]. based [1196, 953, 1801, 2740, 1915, 1269, 577, 2131, 799, 2083, 1321, 1835, 2573, 2716, 2007, 795, 2385, 2128, 1465, 390, 1793, 31, 1457, 1110, 281, 1487, 830, 1257, 388, 1919, 2052, 2327, 1395, 2521, 2505, 2533, 2649, 1681, 2244, 1497, 2482, 2072, 1154, 2514, 346, 380, 1964, 2615, 1026, 1400, 2551, 534, 538, 2663, 2537, 2703, 2584, 211, 2611, 2020, 2223, 834, 973, 972, 1979, 1125, 2454, 2728, 1392, 2038, 1038, 1191, 1398, 1509, 2464, 1644, 2130, 36, 1849, 682]. baseline [1505]. basic [687, 1461]. basidiobolomycosis [2746]. Basidiobolus [2746]. basil [2131]. Basilewsky [745]. basins [1574]. basis [417, 1276]. bass [2031, 519, 62, 901, 2054, 38, 161, 39, 340, 970, 1780, 2107, 660, 1774, 2335, 912, 1102, 1975, 852, 1171, 1308, 2580, 578, 606, 269, 165, 309, 367, 136, 90, 355, 1428, 294, 2157, 1092, 2020, 204, 2056, 2441, 2440, 1106]. batatas [2479]. Batch [621, 1450, 1930, 2374]. batch-specific [1930]. Bate [1146]. Bath [2427, 473, 1939]. baths [2404]. batrachus [846, 267, 1112]. Bavarian [481]. BAX/ [2278]. Bay [1838, 1748, 129, 1103, 416, 1176, 1102, 1786, 126, 790, 1597, 1272, 873, 159, 278, 158, 397, 967, 2155, 106, 2132, 158]. Bays [809, 808]. BCL [2278]. BCL-2 [2278]. BCR [2148]. Bdellovibrio [1701]. be [829, 1686, 2074]. beads [209, 1493]. bean [2386, 1496]. bearded [1144]. becomes [787]. bed [793, 520, 521, 627, 810, 1968]. beds [1749, 1726]. beet [2011]. before [1319, 405, 426, 1717]. beginning [1216]. Behavior [1941, 1770, 786, 1174, 1083, 1351, 1266, 2744, 2583, 2285, 2393, 2188, 2057]. behavioral [371, 1243]. behaviour [62, 113, 595, 459, 65, 1093, 280, 2445, 2073, 175]. Behavioural [1589, 918, 234, 637, 779, 1572, 1208, 1948, 1947]. Beijing [823]. being [1717]. Bekaert [690]. belcheri [1786]. Belgium [945, 646]. beliefs [2625]. bellied [2225, 2221]. Belonging [621]. Belt [1276]. beluga [1076, 1303, 2455, 2399, 1590, 1496]. Benedenia [794]. Beneficial [1775, 1607, 1994, 1646, 1483, 2089]. benefits [1480, 1519, 2005]. Bengal [1904, 1724]. benghalensis [1437]. Benin [2398]. Benthic [795, 984, 1099, 106, 144, 1612, 1793, 1120, 1786, 1749, 1726, 1275, 2465, 1607, 1424]. benzo [2387]. Benzocaine [2118, 1884, 2224]. berries [2569, 2753]. Best [768, 774, 872]. beta [896]. beta-1 [896]. betaine [2738]. Betanodavirus [2229, 1881, 1974, 2022, 2206, 1695, 1797]. Betanodavirus-associated [1881]. betle [1998]. Betta [2516, 1601, 1599, 1924, 2073]. Better [1788, 1439, 817, 1059]. Between [481, 772, 1836, 2638, 919, 2535, 273, 1320, 1572, 1852, 2442, 1171, 2625, 373, 937, 2645, 1050, 2073, 1316, 2600, 1938, 2477, 823, 1922, 2228, 272]. BFT [2716, 2610, 2532, 981, 1532, 1393, 1289, 1381, 1379, 1555, 2709, 2622, 2305, 2289, 1389, 2438]. BG11 [2294]. bi [607]. bi-culture [607]. bialatus [733]. bibliometric [1491]. bigelovii [1200]. Bighead [240, 2426, 1816, 1815]. Bight [1067]. Billard [2383]. bimaculatus [2445]. binder [732]. binding [2097]. Bio [714, 1901]. Bio-economical [714]. bio-encapsulation [1901]. bioaccessibility [2407]. bioaccumulation [473]. bioactive [2635, 2331, 2189]. bioactivities [2046]. Bioaugmentation [729]. bioavailability [891]. biochar [2242, 2631]. biochar-amended [2631]. Biochemical [48, 671, 1356, 1907, 903, 535, 1216, 979, 2297, 1173, 1291, 1642, 2589, 558, 1736, 2449, 2210, 2321, 2329, 2072, 101, 294, 1863, 1252, 2123, 2100, 2588, 1496, 2098, 2383, 1795, 1537, 1840, 2634, 363, 1360, 920, 735, 1038, 2318]. biochemistry [772, 2452, 2197, 2511, 2595, 2629]. Biocontrol [972, 2643]. Biodeposits [1870, 1272]. biodiesel [1690, 2581]. Biodiversity [374, 1535, 1203]. Bioeconomic [1533, 1034, 2208, 2498, 2507, 1327, 1237]. Bioencapsulated [1122, 511, 647, 298, 341]. bioencapsulation [39, 429]. bioenergetic [1980]. Biofilm [1532, 241, 1804, 1919, 1725, 1953, 1902, 2216, 941, 1516, 1968]. biofilms [1955, 1280]. biofilter [2446, 607]. biofilters [1001, 1564, 1825]. biofiltration [1686]. Biofloc [1671, 2083, 1555, 2355, 2486, 2551, 2622, 2391, 2337, 2242, 2370, 2716, 1334, 2026, 2532, 1558, 1740, 2394, 2304, 981, 1202, 1173, 1567, 2094, 2434, 1532, 1289, 1259, 1381, 1379, 1889, 2140, 2183, 1725, 1598, 2533, 2598, 1822, 1768, 1123, 2203, 2570, 1796, 2121, 2233, 2514, 2079, 1994, 2525, 2166, 2537, 1880, 1620, 2295, 1452, 2305, 1877, 2018, 2208, 2711, 2176, 2606]. Biofloc-based [2083, 2551, 2533, 2537]. bioflocs [2242, 1492, 1570]. biofoulers [2213]. Biofouling [1051, 1634]. biofuel [2223]. Biogen [1763, 1764]. biogenic [2636]. Biogeochemical [258]. Biolog [2186]. Biological [571, 494, 1402, 2676, 1323, 2614, 1681, 448]. Biology [726, 650, 861, 374, 713, 531, 533, 716, 860, 974, 1039, 1837, 926, 28, 736]. biomarkers [2492, 829, 2569, 2753]. Biomass [2383, 242, 2235, 577, 1134, 2286, 815, 2078, 2085, 1912, 191, 2493, 1006, 2202, 1014, 2013, 524, 2506, 2076, 1618, 823]. Biomaterials [688]. biomedicine [647]. biomedicines [836, 739]. biometrics [2050]. biometry [1873, 2629]. Biophysical [790, 2386]. Biopotentiality [1668]. biopsy [695]. bioreactors [793, 1953, 1808]. Bioremediation [2178, 1471, 1631, 2295]. biosecure [1689]. Biosecurity [700]. Biosensors [294]. biostimulant [2681, 2223]. Biotechnology [347, 678, 553, 1075, 887, 2259, 1070]. biotic [2104, 1498]. biotin [267]. Bioturbation [1099, 1279, 1424]. birdiae [1334, 1887, 610]. Biscay [1838]. bisexual [1696, 455]. bivalent [2329, 2482]. Bivalve [474, 2237, 2154, 811, 1834, 803, 143]. bivalves [1526]. Bivalvia [1306, 636, 283]. Bjørndal [992]. Black [329, 2523, 2538, 249, 1281, 2095, 396, 2734, 2325, 2638, 2738, 1135, 1480, 1032, 1145, 1952, 1315, 2591, 124, 2706, 952, 950, 2021, 2120, 1644, 500, 2710]. black-lip [396, 1032, 1952]. black-lipped [1281, 1480]. blackish [208]. blacklip [459, 469]. Blackslip [235]. Blackwell [564, 554]. bladder [1932, 1882, 1084, 1208, 1928]. blade [2187]. blades [2727]. Blanc [574]. Bleeding [228]. Bleeker [2742]. blend [1888, 2335, 789]. Bloch [671, 1116, 1081, 2206, 817, 2263, 2391, 1881, 2445, 1516, 2674, 1038]. blochii [2595]. Blood [639, 2497, 97, 2280, 2107, 2363, 2352, 838, 944, 2335, 1642, 1636, 2516, 2511, 2329, 2595, 2281, 2160, 2301, 2698, 1739, 1727, 1398, 2318, 2568]. bloom [1249, 199]. bloom-tanks [199]. Blue [411, 1924, 848, 1041, 1067, 1375, 2501, 804, 935, 26, 737, 371, 718, 963, 2191, 2283, 1618]. blue-green [1041, 963]. bluefin [1121, 1111, 1190]. bluegill [874]. blunt [184, 2035]. bocourti [988]. body [1448, 2699, 953, 1907, 1645, 1255, 1076, 799, 1049, 1679, 2253, 1716, 2147, 2394, 603, 780, 203, 919, 2299, 1028, 1784, 404, 1766, 1340, 1896, 944, 824, 1552, 1257, 912, 1135, 1364, 1162, 720, 2197, 817, 1473, 2291, 1040, 1651, 163, 148, 1636, 1699, 1550, 550, 1267, 2737, 338, 342, 2121, 2072, 268, 1357, 1369, 445, 884, 1372, 2180, 630, 742, 750, 1653, 1496, 959, 1595, 611, 1739, 1702, 1190, 685, 2222, 1148, 782, 1731]. Bohai [129, 1737, 2038]. Bohemia [2475]. bombardment [893]. Bone [736, 1339, 1125]. bones [1847]. Book [21, 34, 43, 72, 117, 141, 152, 213, 179, 194, 223, 244, 253, 261, 474, 466, 498, 527, 98, 564, 329, 411, 307, 552, 726, 27, 51, 296, 297, 330, 365, 366, 353, 354, 449, 447, 565, 554, 563, 608, 609, 650, 690, 691, 688, 679, 678, 692, 708, 700, 727, 716, 709, 736, 774, 785, 860, 861, 859, 863, 862, 1009, 957, 1008, 992, 956, 991, 955, 974, 1070, 1071, 1039, 1056, 1072, 1151, 1152, 1204, 1284, 1285, 1283, 1282, 1436, 1837, 1905, 1975, 448, 274, 446, 625, 374, 689, 553, 616, 725]. Book [680, 864]. Books [564]. Boone [1290, 1904, 1248, 1334, 1463, 1678, 1173, 2434, 1992, 1052, 2631, 1654, 2739, 1324, 1822, 2203, 1019, 373, 2295, 2100, 2102, 522, 526, 1711, 1021, 1397]. boost [2337]. borgei [2092]. Botechnology [498]. both [1261, 2623, 767]. Botryococcus [2085]. bottlenecks [1946, 1945]. Bottom [323, 299, 2015, 832, 1786, 401, 612, 1318]. Botucatu [2122]. boulardii [465]. boulengeri [2221]. bound [1673]. bovine [1603, 1105]. Bowen [860]. Brachionus [2370, 430, 1923, 1944, 621, 429, 1814, 775, 1798, 2262]. brachydactyla [731, 1291]. brachypomus [2225, 2460]. brachyuran [2550]. Brackish [510, 1380, 2289, 619]. Brackishwater [94, 2178, 2295, 1112]. brama [489, 585]. bran [764, 2340, 2435]. branchial [1980, 134]. Branchinella [1754]. Branchiopoda [1754]. Brandt [2219]. Branson [727]. brasiliensis [629]. braunii [2085]. Brazil [2135, 2122, 1251, 2468, 610, 1197, 1964, 1620, 1762, 2384, 1512]. Brazilian [629, 2458, 1836, 1899, 2590, 1964, 2076]. breach [2749]. breaking [528]. Bream [489, 507, 510, 562, 93, 882, 160, 1622, 390, 1565, 888, 886, 919, 405, 1083, 1304, 2367, 974, 2197, 1486, 167, 1481, 58, 585, 186, 2035, 983, 380, 1369, 1363, 309, 1332, 339, 136, 156, 898, 844, 1062, 557, 204, 1660]. breast [2664]. breathing [1112]. bred [2136, 805]. breeder [638]. breeders [1747]. Breeding [236, 1335, 1098, 1220, 1557, 1743, 598, 132, 1487]. Brest [106]. brewer [1169, 1082]. brewers [946]. Brewery [1393]. Bright [1975]. brill [1090]. brine [1780]. bristle [287]. British [271]. brooders [2555]. Broodstock [2654, 500, 511, 753, 1843, 368, 286, 1730, 910, 2138, 423, 2338, 923, 598, 1207, 1848, 1941, 2057, 1553, 771, 2022, 737, 1431, 667, 1939, 718, 442, 2102, 930, 1809, 1795, 2182]. Broodstocks [239, 1229, 1565, 1435, 204]. Brook [481, 1233, 884, 2160, 1158, 1581]. brown [1178, 2552, 1251, 1463, 1869, 871, 2331, 1207, 1501, 2524, 2175, 959, 611, 2096, 1727]. brown-marbled [2096]. brownbow [574, 551]. browned [1956]. Bruno [449]. Brycon [392, 2028]. bubble [212, 1960, 2516]. Buccinum [517, 518].

Buchmann [1056]. budget [1709, 705, 1584, 1023, 2002, 1741, 156, 1130, 1666]. building [1409]. buitendijki [2553]. Bulgaria [32]. Bulgarian [661]. bulgaricus [2515]. bulguricus [1484]. bullfrog [2629]. bullseye [435]. bungeanum [1595]. Bunting [1204]. burbot [1210, 1209, 1208, 1225]. Burchell [492, 2142, 801]. burrowing [1958]. burst [1981]. Butcher [2383]. butter [2293]. buttnerae [2315, 2254, 2613, 2409]. butyrate [1763, 1764, 2472, 2511]. butyricum [1925]. by-catch [846]. by-product [961, 844]. By-products [243, 1813, 1447, 1386]. byplant [291]. byssus [916].

C [297, 366, 353, 354, 563, 650, 690, 679, 678, 700, 709, 974, 1071, 1039, 1151, 1852, 1301, 601, 78, 1363, 1287, 1134, 2389, 2140, 1792, 1642, 2452, 720, 2355, 1685, 1297, 2151, 2537, 2056, 2076, 2652, 1443, 2011, 2718, 2702]. C-3102 [2076]. C-3102-supplemented [2151]. C-enriched [78]. C-reactive [1852]. c-type [1287, 2389]. C. [717, 650, 774, 1009, 1151, 689]. C.-S [700, 709]. C/ [2140, 2355]. C/N [1792, 2537]. C/N-CP [1792]. C1HD [2339]. C20015 [2115]. C5 [2288]. Ca [1023, 2381]. CABI [553]. cachara [1779, 1097]. cadmium [2593, 593, 175]. cadmium-contaminated [175]. Cage [567, 449, 2648, 866, 2155, 1748, 1906, 769, 962, 1427, 171, 977, 765, 528, 1714, 2194, 1597, 858, 1512, 1284]. cage-farmed [2194]. cage-reared [171]. cages [153, 904, 749, 535, 1100, 45, 370, 1210, 888, 886, 1004, 988, 305, 2343, 2022, 383, 1881, 2710]. cAgNPs [1619]. cainii [2385, 1254]. cajuputi [2553]. cakes [200]. cakes/meals [200]. calanoid [1534, 1632]. Calanopia [2370]. calcarifer [1645, 2442, 2343, 2329, 1691, 1881, 2674, 997]. calcification [1720, 2241]. calcitrans [1823, 1831]. Calcium [1587, 245, 1635, 827, 2434, 1289, 182, 2072, 1847]. calcium-supplemented [1635]. Calico [169, 626]. California [169, 566, 2404]. Call [231]. Callinectes [2126]. Campo [2050]. Can [1789, 2602, 1686, 2270, 313, 1256, 1534, 1404, 2749, 2074]. Canada [138, 2036, 1865, 1395, 2527]. Canadian [2527]. canadum [1037, 2279, 1511, 962, 2378, 765, 913, 1113, 961, 1512]. canal [657]. canaliculatus [717, 990]. canaliculus [615, 986, 2101, 2547, 2630, 2264, 1051]. canals [1167]. canarium [2485]. Candidate [503, 2635, 424, 1670, 560, 698, 2691, 1263, 1872, 517, 518, 2332, 2412, 2425, 892]. candidates [1621, 1776, 2068, 1866]. Cannibalism [1539, 1310, 2459, 1499, 1929]. canola [2651, 1543]. canthaxanthin [1636]. capacities [1681]. capacity [2618, 1992, 2132, 1587, 2428, 2672, 1816, 2395, 2220, 1471, 1478, 2479, 922, 2562, 2495, 2382, 2745, 1657, 2634, 2420, 2500, 2658]. capital [1416]. capitata [2267]. caprylic [915]. capsulatus [2427]. capsules [910]. captive [1098, 1826, 1992, 2136, 295]. captive-bred [2136]. Captivity [509, 508, 698, 781, 2524]. capture [2626, 2752]. Captured [500]. Caraguatatuba [1251]. Carangidae [508]. carapace [1370]. Carassius [738, 1382, 2587, 1925, 2526, 2463, 2511, 1495, 1155, 2622, 2105, 2333, 1991, 2339, 999, 1223, 1500, 2469]. carbo [656, 2475, 918, 1050]. carbohydrate [492, 1954, 1779, 1278, 361, 387, 611, 1438, 1613]. carbohydrate-related [1613]. carbohydrates [1640, 2091]. Carbon [1913, 1176, 1130, 1020, 2083, 1740, 1393, 1417, 2183, 2355, 2449, 1953, 1593, 2121, 2233, 2537, 1376, 1452, 1271, 1199, 542]. carbon/nitrogen [2233]. carbonic [92]. carbonylation [2282]. carboxylate [2387]. carcass [2513, 1081, 846, 994, 115, 114, 409, 1719, 1727, 1191, 1398, 2661]. Carcinus [938]. card [2358]. cardinalis [2601]. Cargill [859, 1285]. Cariaco [323, 401]. Carica [2647]. Caridea [403]. carinicauda [2097, 2271]. carnitine [2748, 820]. carnivorous [1884, 2074]. Carotene [1995, 2066]. carotenoid [1995, 2713, 64, 2085, 196, 1077, 1124, 1329, 2255]. carotenoids [1446, 64, 525, 1795, 327]. Carp [570, 243, 240, 1580, 1647, 241, 1638, 1578, 2475, 2567, 461, 2502, 150, 2104, 2207, 1980, 2137, 2513, 869, 2003, 1218, 1220, 588, 149, 1216, 198, 2336, 1238, 1213, 1293, 2274, 398, 849, 603, 1891, 1187, 1459, 94, 456, 987, 2140, 2535, 1999, 1792, 1235, 1584, 755, 1066, 1382, 633, 2276, 1231, 1002, 455, 2587, 1040, 2009, 918, 1217, 1201, 1362, 1699, 1925, 2426, 2672, 1816, 2032, 1550, 2707, 1576, 1684, 585, 1585, 666, 1583, 2121, 2072, 1495, 2041, 1714, 2033, 371, 400, 473, 451, 1574, 1234, 1863, 963, 1898, 2357, 2129, 654, 2650, 124, 1667, 2189, 1805, 1677, 1050]. carp [1723, 668, 1577, 1900, 969, 1666, 664, 670, 453, 826, 1030, 1595, 669, 665, 1167, 1735, 2179, 2339, 2435, 2749, 1203, 2288, 2278, 2565, 1219, 2152, 2146, 999, 1223, 1815, 1883, 2115, 2245, 1451, 1500, 2120, 2469, 97]. carpet [1843, 541, 1630]. carpio [1383, 2345, 1495, 2041, 2634, 1638, 1578, 2104, 570, 869, 2003, 1218, 1220, 1216, 336, 456, 987, 2140, 1584, 1383, 1642, 1417, 2276, 1231, 1002, 455, 2009, 918, 1217, 1201, 1576, 663, 585, 1583, 2121, 1155, 400, 473, 1574, 1234, 1863, 2357, 2650, 1805, 1677, 1050, 1723, 668, 1577, 1900, 1595, 669, 2435, 2278, 1219, 2245]. Carps [579, 953, 2660, 1242, 845, 2659, 2350]. carrageenan [1867]. carrageenophytes [2444]. carrier [1953]. carrying [2132, 2690, 1816]. CASA [1217]. case [2475, 1015, 2423, 2458, 541, 2509, 1793, 2154, 2590, 278, 2627, 1298, 2172, 2583, 1651, 176, 1571, 1982, 1325, 2106, 1787, 963, 1620, 1112, 1930, 803, 1979, 284, 664, 542, 911, 1203, 1660]. casei [1879]. casein [1217]. cases [1415]. Caspian [1782, 1109, 2321, 2747, 959]. Caspius [959, 2321, 1727]. catappa [2516, 2153]. Catarina [2429]. catch [846, 843]. catch-up [843]. Catfish [2214, 237, 1888, 128, 671, 1116, 1477, 1972, 492, 255, 2142, 1789, 1750, 2360, 145, 2190, 2559, 801, 2277, 2293, 2297, 200, 1081, 1226, 813, 205, 846, 1818, 206, 66, 1781, 988, 2476, 65, 877, 993, 2291, 2492, 1211, 1615, 325, 617, 881, 2554, 1802, 2263, 267, 2498, 2514, 2200, 268, 346, 2346, 1714, 1884, 1661, 1751, 1376, 2281, 1273, 2330, 2301, 360, 1246, 2698, 1488, 1702, 1733, 1682, 1665, 1405, 1509, 1922, 1981, 1988, 2568]. catfishes [2007]. Catla [2045, 2513, 969, 1191, 1269, 2137, 2513, 2045, 1449, 969, 1191, 1398]. caught [1344, 2555, 61]. Caulerpa [1001]. causative [2209, 1260]. caused [1986, 2483, 1589, 2361, 2163]. causes [1956, 2086, 2619, 1591]. causing [2605, 1441, 1615, 2645, 2216, 936]. caviae [2215, 2677]. cavitation [1249, 1738]. CD [297]. CD-ROM [297]. cDNA [2325, 2288]. cell [1212, 2280, 531, 2202, 1629, 2386, 2016, 594, 2160, 1981]. cell-mediated [594]. cells [1293, 1912, 2679, 2658, 746]. cellular [985]. cellularity [2638]. Cellulase [197]. Centella [2041]. central [2476, 1803, 2499]. Centropomus [1317]. Centropyge [2712]. cepa [2593]. cephalopod [1016]. cephalopods [1128]. cephalus [1888, 2235, 2248, 549, 392, 2023, 471, 555, 2028]. Ceratomyxosis [1730]. Cereal [243, 1585, 669]. cereals [1584]. cereus [1668, 1517]. cerevisiae [1812, 1033, 2084, 2262, 1731]. certifications [2240]. Cetti [537]. cf [2393]. cf. [2543, 2285]. CgGH [2559]. CGM [939]. cGnRH [255]. cGnRH-II [255]. Chabahar [2173]. Chaetoceros [1823, 1831]. Chaetomorpha [2588, 847]. chain [2266, 2560, 1531, 2651, 2590, 1774, 915, 2068, 2195, 2594, 972]. chains [183]. challenge [1033, 2031, 2284, 635, 42, 1999, 2111, 1169, 2008, 1805, 1185, 2402, 1438, 1922, 2234]. challenged [2653, 1887, 2148, 2075, 2218, 2034, 2200, 2151, 2018, 2723, 1966, 2061]. Challenges [1899, 1522, 2730, 2504, 2512, 1410, 1115, 2597, 2129]. Chandra [1205]. Chang [895]. Change [1301, 2036, 2258, 1427, 157, 1363, 1394, 1190, 1613]. Changes [1721, 146, 1213, 2331, 621, 596, 1450, 798, 1357, 250, 655, 597, 2119, 557, 135, 504, 2361, 1353, 2450, 2538, 883, 1239, 1513, 2637, 1230, 1895, 409, 1850, 770, 591, 371, 2648, 400, 445, 1562, 312, 2496, 1415, 2117, 1050, 672, 1979, 387, 293, 1108, 360, 2011, 2424]. Changhua [1513]. changing [1013, 414]. Channa [2579, 2741, 1914, 2058, 2037, 1437, 2402, 1038, 1561]. Channel [2528, 2501, 1615, 1802, 1273, 1702, 1682, 1405, 1981]. channels [2664]. Chanos [623, 1660]. char [1233, 667, 1158, 1581]. Characidae [1098, 392]. Characiformes [1098, 2225, 2053, 1828, 2028]. character [155, 302]. Characterisation [2364, 2309, 2542, 2567, 1395, 1155]. characteristic [183]. Characteristics [1220, 227, 1458, 229, 1763, 1018, 2523, 885, 1764, 2050, 1724, 879, 1292, 585, 1506, 1057, 1243, 1880, 1183, 61, 2529, 2160, 2179, 2375]. Characterization [792, 2161, 1996, 484, 2644, 576, 896, 1707, 1075, 2635, 2389, 2122, 2060, 2676, 2380, 1323, 1970, 2276, 2442, 980, 2448, 2362, 1908, 2727, 48, 2159, 943, 2400, 2563, 2139, 2478, 2746, 1069, 1361, 936, 2441, 2440, 2368, 2379, 2082]. characterizations [2587]. Characterizing [2030]. charcoal [1319]. Charentais [416]. Charente [416]. Charr [481, 25, 299, 3, 50, 188, 7, 1127, 1882, 322, 91]. chelate [1500]. Cheltenham [552]. Chemical [615, 1395, 2390, 82, 242, 167, 879, 932, 95, 2175, 47, 2462, 1857, 1425, 2011]. chemicals [20, 60]. chemism [1979]. chemistry [2352, 838, 944, 2335, 2516, 1357, 2281, 2698, 1739, 1857, 2568]. chemotherapeutics [1294]. Chennai [2178]. Cherax [2083, 2385, 1254, 822, 2372, 865, 1794, 2536]. Chiayi [1298]. chicken [2702, 846, 2473, 2728]. Child [218, 219]. Chile [1306, 728, 605, 303, 284, 851]. Chilean [236, 1967]. chilensis [560, 567]. Chilka [523]. chilled [2056]. China [1612, 926, 792, 129, 571, 769, 1176, 1102, 1864, 777, 889, 140, 2317, 2033, 1243, 1272, 1909, 936, 599, 1693, 1746, 2620, 2669, 2149, 823, 933, 686, 873, 2692, 1674, 1318, 1815, 2004, 1195]. chinensis [1743, 980, 1737, 2008, 1742, 734, 2038, 761, 905]. Chinese [1238, 2051, 1680, 1956, 2227, 1142, 980, 929, 932, 1156, 1737, 2672, 2395, 80, 2049, 2659, 931, 753, 930, 2373, 2361, 2382, 2548, 734, 1799, 2038, 2500, 1405, 2323, 1649, 1752, 2639]. Chinook [2156]. chitooligosaccharides [1675]. chitosan [1957, 2031, 1261, 2299, 2515, 2708, 2058, 2037, 2189, 2677, 2470, 2152, 1500]. chitosan-selenium [2677]. Chlamys [1176, 1597, 1438, 873]. Chloramine [2282]. chloramphenicol [755, 407]. chlorella [724, 2078, 1490, 2294, 1954, 2304, 196, 1014, 1045, 2223, 2383, 724]. chloride [2545, 1847, 1022]. chlorides [2434]. chlorine [1236]. chloro [2387]. chlorophyll [826]. chloroticus [697]. choice [676]. cholerae [2483]. cholesterol [604, 400, 1200, 1976]. Choluteca [95]. Chondrus [717, 1471]. chorionic [1160]. chromaticity [2100]. chromatophores [924]. chromide [2206]. chromium [604, 2220, 820]. Chromosome [256, 124]. chronic [2642]. chronicle [1407]. chrysops [852]. chu [1552]. chuatsi [856, 2672, 2323]. chub [2235]. chuii [175]. chum [1987]. chymotrypsin [1873]. Cichlasoma [645, 703]. Cichlid [645, 1894, 1821, 1077, 1636, 2118]. cichlids [1596, 703]. ciliate [1240, 2691]. cineole [2321]. cinnamaldehyde [2261]. cinnamon [2301, 2698]. Ciparay [2650]. circle [787]. circulans [1966]. circulated [644]. circulating [1302]. circulation [660]. Cirrhinus [2567, 461, 953, 1736, 2672]. cirrosa [1020]. cisplatin [2619]. Cissus [1252]. citric [2335, 1348]. citrinellus [1636]. Citrus [2554]. Cittarium [1955]. cladoceran [1481]. Cladophora [2408]. Clam [490, 534, 1843, 772, 439, 541, 1669, 871, 705, 1485, 1630, 1749, 1726, 829, 1787, 2642, 1279, 177, 1660, 2108]. clams [1196, 1958, 805, 536]. clandestinum [1307]. Clarias [1888, 492, 2142, 2190, 2559, 801, 813, 846, 877, 993, 2034, 267, 382, 1998, 2200, 2346, 1112, 1866, 2281]. clarki [842]. clarkii [1468, 2598, 2365, 1554, 1625, 2628, 2745, 2720, 1148, 1540]. class [1450, 17]. classes [165, 2188, 1016, 1159, 1149]. classification [682]. Claus [2026]. clava [1391, 1606, 1546]. claw [1794, 272]. clawed [345, 369]. clay [1940]. clay-turbid [1940]. cleaner [1025, 1837]. cleaning [2637, 2704]. Cleanser^{T M} [2385]. clear [1822, 2018, 2176, 2606]. clearance [127, 873]. cliftonii [879]. climat [453]. climate [1997, 2258, 2366, 1632, 2541, 2080, 1394]. climbing [2394]. clinical [885, 2409]. cloacae [1821]. clone [164]. clones [189]. Cloning [2378, 980, 2362, 2097, 2099, 1189, 2222, 2168, 2288, 2389, 2325, 2276, 2563, 2469]. closed [1953, 1814, 2106, 41, 1302, 2149]. closed-tube [2149]. closer [2381]. Clostridium [1925, 2707]. closure [1746]. Clove [513, 1601, 1599]. clown [1124]. clownfish [2436, 2136, 2362, 1536, 1554]. Cluster [2627]. co [1772, 2083, 1248, 2280, 2751, 1376, 900, 1302, 1858, 847, 2480, 1318]. co-culture [1772, 2083, 1248, 1376, 847, 1318]. co-cultured [2280, 1302]. co-fed [900]. co-infection [2480]. co-infections [2751]. co-pelletization [1858]. coagulans [1563, 1184]. coagulant [2728]. Coast [1904, 81, 1748, 1565, 1756, 1755, 558, 2417, 380, 2178, 2528, 2349, 858, 1237]. Coastal [238, 413, 1917, 1721, 1793, 2555, 1380, 575, 1549, 1970, 417, 2052, 2343, 984, 1593, 1485, 1787, 2141, 2080, 1183, 2383, 1693, 416]. coasts [2173, 1865]. coated [1261, 1470, 1601]. cobia [1037, 2279, 1511, 962, 2378, 2564, 765, 913, 1113, 961, 2168, 1512]. cockle [810]. coconut [444]. Cod [821, 69, 577, 1140, 1309, 639, 1874, 1110, 1439, 22, 1046, 960, 437, 1431, 2695, 620, 431, 2112]. codling [629]. coefficient [762, 2374]. cohnii [1771]. Coho [2303, 2420]. cohort [11]. cohorts [2345]. coibor [1552]. coin [843]. coinfection [2159, 2520]. coioides [2687, 2754, 550, 2022, 1184, 921, 1069, 1404, 1455, 2082]. Cold [564, 290, 299, 1086, 1472, 1641, 715, 1300, 2296, 670, 2441, 2440, 2655, 1794, 674]. cold-shock [1300]. cold-stored [1472]. cold-stress [2655]. Cold-water [564, 715]. coldwater [2746]. Cole [365]. collagen [557]. collagenase [1847]. collected [2173, 2410, 1485]. collection [77, 106, 771, 776]. Collette [860]. Collingwood [565]. colloidal [1619]. Colonization [298, 1994, 2412, 2238]. Color [2436, 1734, 1124, 1329, 1800, 1928, 1718, 1455, 1613, 2186, 2569, 2753]. coloration [2712, 1596]. colors [1971]. Colossoma [2315, 2053, 2254, 2419, 2352, 2308, 2737, 2153, 2613, 2409, 2224, 1731, 2569, 2753, 2201, 2176]. Colossomamacropomum [1964]. colour [1679, 1281, 64, 254, 1370, 1422, 703, 1636, 1550, 41, 1617, 2222]. colouration [975, 2255]. columbina [2066]. columnare [2380, 855]. colymbus [1703]. Comacchio [2050]. combat [2403, 2600]. combating [2466]. combination [150, 1652, 1358, 533, 2045, 2623, 925, 2747, 767, 963, 2219]. combinations [1735]. Combined [181, 1299, 2021, 1807, 1005, 652, 1168, 595, 2733, 2392, 1590, 2330]. combining [1163]. comes [1358]. Commercial [227, 866, 1085, 713, 1838, 947, 145, 1679, 2239, 1569, 1836, 1215, 138, 1558, 1923, 2558, 31, 2024, 415, 2637, 962, 1198, 2492, 1335, 1822, 1460, 613, 2345, 2417, 765, 75, 524, 868, 737, 1811, 710, 445, 2531, 2262, 2243, 362, 2444, 1161, 1988, 1731]. commercial-scale [31, 2345, 710]. commercial-size [445]. commercialization [1745]. commercially [856, 927, 926, 596, 2528, 597, 2182]. Common [489, 570, 243, 241, 504, 1638, 2502, 150, 2003, 1218, 149, 1216, 2336, 456, 987, 2436, 2140, 1584, 2276, 1002, 2386, 455, 2009, 1217, 948, 1201, 176, 1576, 1585, 666, 1583, 190, 338, 2121, 517, 518, 1143, 1359, 14, 400, 473, 1574, 1234, 1863, 2650, 1805, 1677, 1723, 2677, 668, 1577, 44, 1900, 664, 669, 1735, 2435, 2278, 2245, 97, 503]. Communal [487, 1487]. communally [586]. Communication [212, 161, 178, 200, 151, 140, 193, 186, 177, 1259, 1174, 1774, 2042, 414, 2327, 2487, 2657, 2076]. communities [2421, 144, 830, 1564, 1002, 889, 2411, 2525, 1104, 775, 1107, 2119, 2575]. community [2242, 1889, 1701, 1786, 1546, 1460, 1808, 429, 747, 2233, 1325, 1275, 362, 971, 386]. compact [1832]. Comparative [1638, 760, 1445, 2045, 1066, 1329, 2735, 1585, 578, 1508, 1368, 394, 1933, 2224, 2396, 436, 2164, 2397, 2152, 831, 2234, 2142, 1242, 2732, 183, 1307, 1827, 455, 1065, 452, 2546, 607, 1922]. compare [439, 1097]. compared [305, 1177, 2479]. Comparing [767, 1254]. Comparison [2414, 923, 2638, 1849, 1440, 567, 1399, 627, 1692, 1742, 722, 524, 1359, 246, 1579, 211, 2724, 1387, 1195, 36, 1067, 1558, 1634, 752, 633, 558, 520, 521, 2625, 1264, 809, 28, 906, 1948, 1947, 1694, 1938, 697, 2489, 272]. compartment [2389]. Compensatory [1768, 734, 2699, 22, 843, 1338, 1012, 2711]. competence [2005]. competent [561, 2453]. competition [1290, 1913, 1305, 1798]. Competitiveness [2051]. complement [2288]. Complementary [1316, 1902]. complements [2235]. complete [1142]. Completed [515]. completion [63]. Complex [621]. complexes [1540, 1883]. complications [1956]. component [203, 1123]. composite [2274]. Composition [471, 242, 1448, 2699, 953, 1356, 1907, 1645, 1255, 1696, 1076, 1133, 799, 1049, 444, 146, 2513, 903, 1319, 1334, 1716, 2190, 2046, 910, 979, 2147, 2394, 2455, 2399, 603, 981, 1173, 780, 1081, 1889, 203, 907, 2299, 1028, 1784, 846, 1766, 1340, 818, 1109, 1841, 944, 824, 1552, 1609, 1135, 1291, 994, 1311, 2589, 1364, 821, 115, 114, 1643, 975, 1162, 2197, 817, 1473, 2291, 2492, 1014, 558, 1040, 163, 148, 879, 409, 1598, 932, 1699, 157, 454, 1803, 550, 37, 1267, 2311, 2737, 2210, 338, 342, 2121, 2390, 95, 2072, 1450, 165, 268, 1369]. composition [2508, 982, 2175, 373, 634, 101, 1767, 2697, 17, 1372, 18, 393, 394, 2462, 2180, 1163, 630, 900, 742, 61, 1978, 1505, 431, 2562, 2588, 1496, 959, 1719, 930, 1666, 2098, 2383, 2301, 1595, 2300, 2520, 1795, 1128, 611, 1537, 1739, 1857, 685, 1618, 914, 1148, 782, 1727, 1581, 1931, 2219, 1038, 1191, 1398, 1883, 2130, 1731, 2374]. composition-odor [1319]. Compositions [250, 1188, 1124, 2508, 1296, 202, 363, 1425, 1106]. composted [699]. Compound [2639, 2502, 2580, 1585, 1234, 1007, 2634]. compounds [2331, 1615, 1802, 1361, 823]. Comprehensive [1854, 1557, 2236, 1530]. compromised [40]. computational [2573, 2716, 1544]. Computer [656, 393]. Computer-assisted [656]. concentrate [942, 1678, 1486, 1019, 1044, 1036, 2318]. concentrated [322, 2103]. concentrates [427, 2667]. concentration [1129, 1354, 693, 1690, 1310, 1023, 2212, 2261, 1910, 2420, 744]. concentrations [2054, 2147, 2450, 2084, 1124, 2291, 2516, 797, 2709, 312, 137, 750, 1723, 826, 823, 1425, 1219, 1561, 1317]. Concepcion [169]. concepts [955]. concerns [1985, 978]. conch [1720, 2241, 2485, 1892]. concrete [633, 1508, 133]. concurrent [845]. Condition [439, 568, 511, 749, 2656, 2230, 1385, 2055, 1485, 644, 702, 1539, 892]. conditioning [890, 260, 1553, 771, 737, 937, 718]. Conditions [500, 515, 493, 494, 513, 1178, 1245, 93, 2286, 2241, 109, 687, 796, 910, 198, 1490, 1913, 970, 1997, 1568, 1173, 1567, 1064, 595, 813, 598, 2352, 1118, 818, 1347, 757, 897, 661, 1958, 1572, 1827, 1606, 754, 1753, 879, 2165, 1681, 2449, 1920, 2463, 1576, 2405, 91, 1143, 1357, 1208, 1928, 586, 294, 1300, 1287, 906, 369, 1830, 1834, 1036, 2330, 641, 1225, 741, 589, 1799, 1527, 999, 1161, 1592, 1831]. conducted [977]. Conference [803, 812]. conferring [1059]. confers [987]. confinements [120]. confirm [2212]. conflicting [2260]. confusion [2705]. Congeneric [490]. conjoint [2526]. conjugated [2232]. Connecticut [278]. consecutive [391]. consensus [706]. conservation [2007, 284]. conserves [2256]. Consideration [227]. Considerations [238, 310]. considering [2731, 2529]. consistent [2516]. consisting [1391, 1606]. consortia [1639]. consortium [1684]. conspecific [804]. constant [1174, 1023, 148, 520, 521, 1312]. constitution [1362]. Constraints [238, 2198, 2, 2199, 706, 2626, 2752, 1551]. construction [1273]. Consumer [502, 646, 499, 2705, 2583, 2625, 1571, 2527, 1062, 1175]. Consumption [495, 1193, 2384, 1015, 1172, 2413, 945, 825, 924, 31, 49, 2367, 1351, 1647, 328, 1369, 2527, 156, 704, 443, 592, 1149, 745]. contact [137]. contactors [2614]. contain [1813]. containing [1133, 1215, 1226, 2748, 2738, 1231, 1447, 982, 797, 2503, 1234, 1863, 959, 1503]. contaminated [175]. contamination [2138, 1064]. content [2713, 1076, 2190, 1690, 2294, 196, 1587, 1329, 37, 2311, 666, 1583, 1450, 46, 437, 1713, 547, 1296, 1873, 41, 2305, 557, 670, 164, 669, 1662, 1350, 2087, 1649, 1352]. Contents [265, 318, 377, 421, 477, 1244, 2496, 2542]. context [2132]. contingent [1985]. Continuous [495, 838, 1629, 1092, 997, 1149]. contribute [1439]. contributing [1409, 1000, 2582]. Contribution [228, 1570, 1520, 751, 2064, 2533, 2163, 2718]. Contributions [447]. Control [700, 788, 2451, 16, 2315, 441, 2474, 1990, 2025, 2254, 1323, 2071, 1861, 1822, 269, 2252, 1605, 2124, 2185, 2613, 1078, 2006, 1528, 2129, 111, 2409, 1302, 973, 2038, 2089]. Controlled [236, 1177, 687, 796, 717, 1572, 2249, 1753, 1143, 1357, 1928, 2139, 906, 641, 589, 999, 1161]. Controlling [485, 1377, 1705]. conventional [2104, 979, 1849, 1880]. Conversion [510, 549, 749, 1081, 2331, 817, 1038, 1398]. Convict [645, 218, 219]. copepod [1354, 1770, 1240, 2170, 1771, 1632, 1614, 1842]. Copepoda [2530]. copepods [1534, 2530, 2211]. coping [1083]. copper [2232, 473, 1059]. coral [781]. Corallina [1859]. coreana [880]. Coregonus [180, 1154]. corkwing [635]. cormorant [656, 2475, 918]. cormorants [1050]. Corn [1541, 908, 2748, 1691, 1863]. cornea [1535]. Correction [2754, 2455, 1764, 2752, 1756, 2057, 1749, 1783, 1946, 2393, 2058, 2577, 1948, 2441, 2456, 2578, 2753]. Correlation [1852, 1996]. correlations [2535, 1341]. corrianus [857]. Corrigendum [118]. corruscans [1750, 1097, 1751, 2076]. Corsican [2524]. corteziensis [1118]. corticata [2589]. Cortisol [629, 1227, 1562, 750]. coruscus [2004]. corylifolia [472]. cost [2421, 2586, 1858, 1489, 2217]. cost-effective [1858]. Costa [411, 625, 2306]. costae [314]. Costatum [250]. costs [66, 868, 380]. Cotentin [558]. cottonii [2257]. cottonseed [1776, 1142, 379, 1824, 1900]. Could [381, 2079]. coumaric [2502]. Council [1055]. count [2160]. counterparts [2386]. Counties [1298]. countries [1950, 74]. County [1513, 1660]. coupled [1298]. courses [1409, 1911]. Cove [1306]. coverage [2381]. COVID [2599, 2507]. COVID-19 [2507]. COVID-19-induced [2599]. cow [1066]. COX [1224]. Coyle [1975]. CP [1792]. CpG [1511, 2689]. CpG-oligodeoxynucleotides [2689]. CPH [1213]. Crab [571, 2361, 2654, 934, 1377, 545, 423, 1680, 1956, 2227, 938, 2555, 1291, 1142, 2387, 929, 932, 2672, 935, 2395, 2377, 1325, 1368, 1704, 1983, 931, 928, 2141, 2080, 2632, 753, 930, 2098, 2373, 1646, 1060, 2382, 612, 1618, 2431, 1799, 933, 2500, 1260, 2639]. crabs [731, 2635, 1569, 927, 1745, 926, 937, 2548, 1649]. crassispina [1318]. Crassostrea [760, 460, 2714, 1344, 786, 807, 1118, 1298, 1102, 1786, 2735, 1692, 1778, 809, 1197, 1433, 2213, 427, 1965, 806, 1712, 808, 1482, 672, 1707, 1820, 2021, 1761, 1352, 2453]. crassus [2388]. Crayfish [1426, 2083, 345, 638, 2385, 842, 1637, 833, 271, 304, 273, 723, 752, 876, 1117, 1266, 115, 114, 388, 1370, 1468, 2598, 722, 917, 369, 822, 2372, 865, 1625, 2745, 1794, 2720, 1148, 2536, 1540]. created [809]. creation [2590]. credentials [1985]. credit [835]. Crimea [2523]. crispus [1471]. cristaefolium [1501]. criteria [2155, 682, 58, 2068]. criterion [1965]. Critical [1111, 2421, 1101, 911]. croaker [1552, 2691, 1949, 674, 745]. Croatia [305, 1144, 1574, 1597]. crocea [2691, 1908, 1949, 674]. Crocus [2452]. Cromileptes [866]. cropping [952]. crops [388]. cross [2246, 1990, 2327, 1315]. cross-protection [1990]. cross-sectional [2327]. cross-tolerance [1315]. crosses [1385]. crossing [456]. crosstalk [2732]. crowding [49]. crucian [1382, 2587, 1550, 1666, 2339, 2749, 999, 1223]. crude [2484, 1045]. cruentum [2679]. Crustacea [358, 1696, 732]. crustacean [2635, 1688, 524]. crustaceans [1738, 2550]. crustin [2039]. crustin-like [2039]. cryoconservation [715]. Cryopreservation [783, 2481, 2096, 1219, 743, 2701, 1212, 1464]. Cryopreserved [483, 1218]. cryoprotectant [1817, 1219]. cryoprotectants [802, 2356, 1758]. cryoprotectors [1222]. Crypthecodinium [1771]. Cryptocaryon [2691]. cryptophyte [2503]. Cryptosporidium [1064]. crystalline [1432, 1875]. CSIRO [565]. csiro.au [565]. Ctenolabrus [635, 1572, 395]. Ctenopharyngodon [1980, 1238, 1293, 849, 603, 1187, 1459, 1999, 1040, 1699, 2032, 2707, 451, 1030, 2179, 2288, 2146, 1451]. CTL [2389]. CTP4 [2046]. CTX [2216]. CTX-M-15 [2216]. cucumber [788, 713, 1708, 2310, 2607, 1606, 839, 889, 1423, 1193, 1243, 1003, 1192, 1299, 1403, 1483, 1607, 1517, 1318, 1373, 1494, 870, 2662]. Cucurbita [1989]. cues [615, 167]. culling [872]. cultch [605]. cultivable [1992]. cultivate [1513]. cultivated [2414, 1251, 2523, 632, 1103, 943, 1749, 1726, 2728, 2606, 272]. Cultivation [717, 1676, 567, 188, 79, 1067, 2026, 1754, 1490, 2294, 1954, 2684, 1461, 1786, 199, 2165, 1335, 1778, 2417, 287, 2498, 2119, 1858, 2444, 2040, 36]. culturable [2125]. cultural [1203, 2673]. Culture [474, 466, 500, 904, 567, 1347, 515, 366, 650, 1802, 566, 2462, 2305, 1069, 628, 572, 2235, 866, 1772, 699, 1789, 77, 995, 2286, 2360, 2083, 1838, 212, 349, 1932, 276, 1248, 764, 282, 2015, 1253, 638, 827, 40, 5, 2274, 681, 791, 358, 438, 832, 816, 842, 2731, 751, 2170, 1202, 2626, 2752, 169, 370, 1525, 811, 323, 1703, 2251, 1889, 828, 938, 1555, 94, 1120, 598, 1724, 2609, 1052, 769, 1345, 818, 1792, 1487, 998, 897, 830, 962, 1298, 1810, 66, 2751, 1461, 1102, 679, 348, 977, 2183, 176, 1725, 1598, 401, 158, 2203, 2210, 1197]. culture [624, 765, 55, 320, 333, 326, 280, 357, 1368, 845, 1450, 1902, 2213, 800, 1034, 2270, 469, 1579, 1946, 1945, 2695, 775, 1287, 1074, 1898, 306, 1597, 1376, 211, 1937, 1130, 1834, 1452, 2391, 1493, 2445, 362, 2264, 952, 2289, 1199, 1570, 1666, 733, 468, 1646, 1089, 847, 714, 1955, 2382, 607, 612, 1483, 933, 1330, 1527, 1318, 1700, 1530, 1237, 2108, 2438, 1355, 1512, 2711, 1831, 36, 449, 564]. Culture-independent [1069]. Cultured [502, 621, 510, 1196, 2666, 1838, 159, 1730, 1904, 1281, 1286, 707, 970, 2280, 391, 2734, 1923, 381, 1028, 738, 1854, 986, 988, 754, 1480, 1420, 2449, 140, 1941, 2057, 1508, 528, 1506, 2347, 2177, 1254, 1065, 2213, 2033, 2596, 2659, 1165, 1579, 1939, 126, 1547, 2645, 394, 1881, 2191, 1302, 1877, 1524, 2070, 2550, 1744, 556, 2149, 1451, 2120, 2182]. cultures [1673, 1240, 1632, 259, 303]. culturing [285, 13, 889, 1785]. cum [2131]. cumin [2095]. cumingii [1734, 1043, 1080, 777, 1329, 2164, 1195]. cuneata [698, 841]. cupped [1712]. curcumin [1556, 2518]. Current [6, 810, 2129, 2458, 1856, 2347, 1243, 686]. custard [69]. cut [1548, 1390]. cut-off [1548, 1390]. Cutaneous [2133]. Cuttlefish [568, 628, 825, 350, 391, 438, 458, 712, 773, 832, 357, 1918, 468, 2575]. Cuvier [903, 816, 842, 907, 404, 515, 1291]. cyanobacteria [2542, 2087]. cyanobacterial [1249]. cyanobacterium [2506]. cyclase [1444]. cyclase-activating [1444]. Cycle [515, 1306, 1011, 2651, 350, 391, 458, 2351, 184, 2210, 101, 283, 1096]. cycles [1326, 1498, 306]. cyclical [883]. cyclin [2665]. cyclin-dependent [2665]. Cyclina [2642, 1279]. cycling [258]. cycloancistrium [2545]. cyclooxygenase [1224]. Cyclopina [1842]. cyclopoid [1842]. Cyclopterus [1655, 1572]. cygnus [426]. CyHV [2280, 2276, 2401]. CyHV-2 [2280, 2401]. CyHV-2-infected [2280]. CyHV-3 [2276]. cylindrica [2589]. Cylindrotheca [1423]. Cynodon [2110, 1664]. Cynoglossus [1320]. Cypermethrin [570, 336]. Cyprinid [479, 573, 2280, 257, 1651, 1420, 2401]. Cyprinidae [1228]. Cyprinus [1638, 1578, 2104, 570, 869, 2003, 1218, 1220, 1216, 336, 456, 987, 2140, 1584, 1383, 1642, 1417, 2276, 1231, 1002, 455, 2009, 918, 1217, 1201, 1576, 663, 2345, 585, 1583, 2121, 1495, 2041, 1155, 400, 473, 1574, 1234, 1863, 2357, 2650, 1805, 1677, 1050, 1723, 668, 1577, 1900, 1595, 669, 2435, 2634, 2278, 1219, 2245]. cyst [1199]. Cystidicola [1982]. Cystoseira [2552, 2331]. cysts [1326, 1498, 435, 210, 116]. cytochalasin [185]. Cytogenetic [1233]. cytometric [1897]. cytometry [2713, 818]. cytoskeleton [2373]. cytosolic [1649]. cytotoxic [143]. cytotoxicity [746]. Czech [2475, 457, 666, 1582]. Czechia [2012]. CZM [416].

D [329, 353, 354, 449, 563, 650, 700, 860, 863, 864, 957, 956, 625, 150, 2387, 887, 1158]. d-Arg [1158]. d-Lys6 [150]. D. [1008, 1436, 680]. DAB1 [1182].

D’Abramo [1039]. dabryanus [2244, 711]. dactylon [2110, 1664]. Daily [1093, 749, 813, 919, 322, 30, 89]. damage [649, 1591]. Damariscotta [159]. damselae [2415, 2102, 2415, 2102]. Dana [638, 1472, 1624, 752, 1771, 1632]. danae [2126]. dangers [55]. Danio [2656, 1147, 1184]. Danish [229]. Danjiangkou [1746]. Danube [1574]. DAP1 [1182]. daphnia [724, 2026]. Dar [911]. darkness [876, 7]. Data [515, 2584, 1906, 1548, 528, 1357, 790, 2044]. data-scarce [1906]. date [1907]. davidi [2742]. Dawestrema [2545]. days [1681, 1208]. de-oiled [2340]. dead [1061]. dead-end [1061]. Dealing [1240]. DEB [2002]. Debaryomyces [2251]. debelius [403]. Decapoda [934, 512, 2566, 865]. decapsulated [435, 210]. decision [1079]. decline [2272]. Decomposing [622]. decoupled [2446]. decrease [1268]. decreased [2479]. decreases [1628, 24]. decussatus [1843, 772, 541, 771, 1429, 1630]. dedicated [1409]. deep [320, 326, 873]. defatted [2706]. defence [2514, 436, 985]. defences [2595, 2396]. defense [2032, 2329, 2491, 2018, 2076, 2639]. deficiency [1353]. defining [2421]. definition [1637, 537]. deformities [1507]. degradation [1179]. degrading [896, 1494, 1503]. Dehydration [441, 1326, 1498]. Delaware [809, 808]. Delayed [221, 1887]. delbrueckii [1484]. deletion [1872]. deliciosa [2390]. Delivery [647]. delousing [2724]. Delphi [706]. Delta [2213, 1394, 952]. Demaine [354]. DeMan [845]. Demand [62, 2434, 4]. Demand-feeding [62]. demissa [1174]. Demographic [1798, 2349]. denitrification [1970, 1968]. Denmark [1276]. Densities [235, 1915, 1251, 2559, 70, 392, 548, 816, 1792, 1513, 352, 1741, 2514, 2596, 1562, 644, 1712, 2382, 1660, 2646, 2438, 2453, 2211]. Density [493, 566, 1178, 1172, 721, 825, 299, 1569, 2015, 1253, 684, 687, 878, 1100, 340, 2394, 1186, 438, 832, 1310, 2170, 2094, 257, 811, 1637, 833, 281, 752, 1090, 459, 988, 408, 7, 65, 1422, 2739, 1054, 11, 839, 2598, 1977, 1045, 2429, 2233, 1368, 2359, 1825, 759, 1165, 2127, 1559, 1547, 1798, 2630, 702, 2391, 133, 968, 1539, 468, 641, 967, 1820, 1190, 1840, 2021, 1360, 1330, 999, 1698, 2201]. density-dependent [7]. Densu [2213]. dentatus [432]. dentex [305, 135]. Dependency [1648, 2673]. dependent [2741, 786, 7, 2665]. depending [1935]. deployment [607]. deposition [1281, 180]. depress [1127]. Depression [219]. deprivation [1054, 761, 905]. Depth [2062, 1834, 728, 281, 1597]. Depths [500]. derivatives [1294]. derived [2109, 155, 1813, 525]. Desbonnet [625]. descaling [1576]. descriptive [1180]. Design [2614, 2704, 1832, 2586, 2065, 2529, 2640, 2652]. designer [2191]. Designing [1409]. designs [1811]. desk [35, 73, 110, 162, 119]. Destratification [320]. destruction [14]. destructor [2536]. Detect [569, 33, 1441, 1651]. detected [730]. Detection [2136, 2703, 763, 1479, 1835, 2519, 1655, 2019, 2649, 2484, 2022, 1476, 1717, 2401, 1697, 2195, 2307, 2464]. deteriorates [1326]. Determinants [2445, 2266, 2476, 1898, 2384]. Determination [2192, 695, 384, 964, 797, 1225, 160, 2212, 2719, 2534, 2262, 1232, 1892, 1392, 711, 1509, 1644]. determine [2356, 1965, 1910, 2718, 2374]. determined [1930, 1390, 1570]. Determining [825, 755]. deterrent [632]. detoxification [1668, 639]. detritivorous [720]. develop [417]. developed [2292]. Developing [236, 467, 2, 463]. Development [1835, 507, 807, 938, 571, 1959, 2487, 1073, 1270, 15, 2312, 76, 2659, 2503, 1476, 965, 941, 2674, 1968, 1389, 2048, 2339, 2120, 1178, 1843, 2421, 2199, 54, 1354, 1720, 2207, 1838, 1520, 706, 441, 801, 1262, 541, 934, 160, 1407, 1227, 529, 291, 416, 2126, 1380, 2627, 757, 1958, 2447, 2485, 1231, 993, 2681, 1209, 1032, 1408, 1550, 1079, 585, 1097, 1519, 2405, 2524, 2597, 74, 2200, 2185, 1952, 1357, 850, 181, 2563, 1027, 28, 1918, 2080, 2332, 379, 654, 2650, 311, 750, 1489, 1108, 1412, 1710, 2652, 2669, 1875, 2431, 2582, 1633, 1698]. development [2186, 425]. Developmental [293, 531, 2249, 1400, 1830, 2222]. developments [2398, 416, 1683]. device [2704, 2065]. DGGE [889]. DHA [1628, 2311, 165, 164]. DHA-to-EPA [1628]. dhufish [295]. diabetes [2683]. diabetic [2680]. Diacronema [784]. Diadema [2503]. diagnoses [2540]. Diagnosis [2162, 2010, 2358, 2576, 1969, 1605, 1065]. Diagnostic [2163, 559, 1868]. diallel [2246]. diallyl [2404]. Diatom [250, 2388, 1052, 2230, 958, 1423, 965, 2465, 1831]. diatoms [1612, 1959, 1607]. Diazinon [663]. dicalcium [2205]. Dicentrarchus [2031, 519, 901, 535, 340, 970, 1780, 2673, 1766, 1774, 912, 1308, 1193, 578, 606, 367, 1428, 2157, 1092, 2020, 1868, 1132, 2056, 1543, 2005]. Dicologoglossa [698, 841]. Dictionary [297]. did [63]. Diel [136, 220, 788, 62, 1149]. Diet [222, 245, 676, 332, 1713, 1907, 1255, 1295, 286, 1129, 2604, 2388, 1673, 903, 1839, 1418, 796, 910, 979, 1610, 1186, 2418, 404, 2367, 833, 2126, 1538, 2748, 1307, 463, 1776, 696, 408, 720, 335, 1040, 1736, 2623, 157, 2032, 1885, 1796, 2035, 2486, 1886, 2151, 1019, 47, 452, 1658, 1783, 464, 2697, 822, 1830, 2706, 900, 431, 2100, 1991, 1366, 697, 1665, 1807, 1698, 1650, 1752, 2130, 1759, 1545]. Dietary [1812, 1888, 2031, 2560, 461, 671, 1116, 1477, 248, 1076, 1750, 1627, 1887, 2147, 2430, 1256, 2303, 2437, 1993, 1912, 1474, 1779, 1640, 1769, 1124, 2335, 1135, 1654, 2291, 1736, 239, 2314, 2533, 2426, 1757, 881, 2608, 428, 2072, 2281, 1824, 985, 1657, 2549, 1425, 1322, 1727, 1038, 1191, 1398, 2265, 1561, 2606, 1178, 2063, 1033, 953, 1020, 562, 492, 1645, 1763, 1049, 444, 1708, 2003, 1121, 882, 1286, 2653, 2054, 2315, 270, 801, 161, 1672, 827, 909, 946, 2277, 1017, 2313, 934, 996, 2025, 899, 1764, 780, 2319, 1081, 2607, 1028, 2743, 196, 2436, 2748, 1188, 406, 1117, 1266]. dietary [1109, 2738, 182, 944, 824, 1552, 1642, 1311, 623, 1169, 1364, 115, 975, 2197, 2328, 1791, 2371, 2693, 1486, 180, 1442, 1499, 604, 1126, 1449, 409, 1244, 1636, 1699, 1324, 1765, 2472, 2717, 2463, 550, 361, 359, 1267, 2220, 193, 601, 2580, 1684, 2737, 2377, 606, 2034, 1388, 1556, 267, 1687, 328, 1590, 165, 767, 1254, 2079, 46, 1184, 367, 880, 1454, 1535, 379, 1091, 547, 922, 331, 2611, 2667, 1138, 820, 742, 1271, 2305, 2372, 2594, 1653, 1496, 1719, 1348, 387, 753, 930, 969, 1374, 1666, 2098, 2330, 2470, 2301, 1030, 1595, 85, 1890, 1867, 2698, 2361, 1189, 1739, 1857, 2397, 2435]. dietary [2495, 2652, 2742, 2745, 1029, 2005, 1022, 1662, 685, 1875, 2634, 1350, 363, 914, 1148, 2420, 1517, 2152, 735, 782, 2219, 1139, 1405, 1976, 2115, 2245, 1021, 2569, 2753, 1751]. diethylstilbestrol [1211]. Diets [232, 568, 243, 573, 1843, 731, 2617, 2410, 1133, 799, 760, 145, 71, 642, 2310, 1463, 9, 942, 149, 1754, 1678, 423, 712, 2725, 178, 200, 1465, 1226, 1396, 435, 1766, 1844, 2230, 1340, 1440, 1813, 1563, 2738, 122, 1541, 2416, 1142, 388, 2708, 1231, 1473, 1882, 2205, 1181, 1420, 2707, 550, 555, 1453, 917, 2405, 517, 518, 346, 389, 1447, 982, 1026, 1332, 797, 1431, 667, 2503, 1863, 1767, 1258, 211, 961, 1007, 1365, 630, 2594, 1044, 2188, 1036, 959, 44, 1125, 1316, 611, 1966, 724, 762, 1537, 1733, 789, 1543, 589, 1146, 990, 1503, 1317]. diets [2568, 36]. differ [2568]. difference [1171]. Differences [772, 910, 1320, 286, 2107, 48, 937, 1965, 2477]. Different [500, 235, 573, 621, 627, 220, 237, 510, 2496, 516, 2179, 242, 1843, 519, 2729, 2059, 731, 788, 25, 1133, 2083, 947, 1619, 174, 642, 749, 1218, 2310, 2559, 2054, 1463, 70, 392, 430, 1215, 45, 970, 2532, 1612, 423, 358, 1923, 1567, 1874, 370, 1226, 390, 1289, 1381, 595, 643, 1555, 813, 1440, 1118, 1624, 2140, 752, 1164, 944, 1564, 659, 1958, 462, 768, 575, 321, 2452, 633, 1370, 2328, 1048, 2183, 2355, 168, 148, 756, 2343, 1385, 1725, 1598, 1420, 2362, 2449, 2408, 325, 1576, 555, 2516, 2417, 2210, 606, 1084, 1749, 1726, 333, 576]. different [2121, 1368, 2263, 352, 1063, 2547, 2514, 517, 518, 165, 1475, 1447, 2508, 1983, 982, 367, 76, 2709, 2622, 2591, 2497, 1234, 538, 718, 1562, 2537, 644, 30, 442, 806, 1712, 19, 1138, 1452, 1107, 1677, 2585, 1723, 1496, 952, 959, 2224, 1758, 733, 1030, 2300, 164, 1466, 865, 1735, 2382, 2620, 589, 685, 2222, 1146, 1148, 1682, 2288, 1219, 1660, 857, 1159, 2646, 1161, 1223, 905, 1722, 1592, 1976, 1149, 1317, 2211, 2510]. Differential [2732, 1487, 1002, 1331, 1026]. differentially [2579]. differentiate [970]. differentiation [522, 526, 1948, 1947, 1728]. differently [799, 1579, 1594]. differing [1453]. diffusing [1960]. diffusion [1548]. digest [2030]. Digestibility [243, 222, 1447, 766, 64, 2081, 178, 200, 196, 824, 1706, 2371, 2314, 2205, 2719, 606, 2611, 733, 1030, 1625, 762, 939, 2318]. digestion [174, 755, 704, 1824, 2683, 2167]. Digestive [2518, 1138, 733, 1477, 344, 772, 385, 1708, 2513, 2003, 2427, 2313, 301, 907, 2743, 1563, 1399, 2335, 2452, 2485, 932, 2032, 1560, 1093, 1267, 943, 2473, 1556, 2316, 1063, 1270, 770, 798, 913, 896, 1145, 850, 591, 2595, 1091, 922, 1504, 1653, 1348, 537, 2742, 1146, 735, 782, 892, 2646, 2219, 905, 2711]. Digestrom [2455, 2399]. digitata [1610]. dihydrate [1921]. diluted [2294]. dimension [212, 2043]. dinoflagellate [143]. diodes [1819]. dioxide [1176]. dipeptide [2313]. dipeptides [1231]. Diplazium [2522]. Diplectanidae [441]. Diplectanum [38, 441]. Diplodus [424, 1286, 506, 1465, 888, 886, 919, 372, 1422, 983, 505, 537]. Diploid [487, 1179, 1157, 993, 584, 1434, 586, 674, 748, 831]. diploid/tetraploid [1157]. dipotassium [1869]. dipstick [2401]. Direct [2258]. directions [2067]. disadoption [2423]. disc [1548]. discarded [847]. discards [1366]. discharge [1235, 1832, 1236, 1246]. discoloration [2734]. discoveries [2302]. discus [1172, 410, 1701, 2220, 896, 850, 710, 2652, 851, 1172, 896]. disease [1812, 1972, 1915, 673, 1049, 2635, 212, 2430, 40, 1346, 2670, 1344, 871, 2605, 2154, 1956, 1474, 1818, 1538, 1439, 1510, 2392, 2218, 1441, 2136, 1914, 2526, 1757, 359, 1974, 1684, 2750, 1796, 1388, 2316, 2252, 1145, 2124, 2557, 2033, 2272, 1082, 2561, 2478, 2195, 763, 2010, 2259, 1723, 1348, 2470, 2520, 1437, 985, 2575, 2163, 1405, 2115, 2245, 1731, 700]. disease-affected [673]. diseased [2567, 792, 2014, 140, 950, 2477, 2089]. diseases [2540, 2125, 2358, 2474, 2670, 2466, 295, 1775, 2550, 1705, 449]. Dish [690]. Disinfection [364, 440, 129, 351, 1470, 2282, 60, 137]. disk [896, 710]. dismutase [1649]. Disorders [449]. dispersal [2501]. Dispersion [227]. disruption [1780]. disruptive [632]. dissimilation [1862]. dissolved [1448, 1458, 2092, 1593, 1101, 2584]. distal [1140]. distinct [1564, 1965]. Distolasterias [2359]. distortions [104]. Distribution [2014, 251, 1711, 2284, 2036, 2519, 807, 723, 957, 125, 1593, 2591, 126, 898, 2469]. District [2650, 977]. disturbance [1107]. disturbances [1589]. diurnal [1101]. divergence [895]. diverse [1564]. diversicolor [1870]. Diversification [963]. Diversified [2236]. diversifolia [1998]. Diversity [1413, 503, 1902, 989, 2342, 1457, 1854, 860, 957, 777, 1953, 1964, 2312, 2648, 1502, 1114, 2120, 1728, 2182]. diverting [775]. diving [632]. division [2376]. dl [767, 1486]. dl- [767]. DL-methionine-supplemented [1486]. dmrt3 [2534]. DMTIOLA [1079]. DNA [649, 481, 954, 86, 894, 1060, 2339, 1392, 2038, 1398, 1509, 1728, 1644]. Do [1439, 1985, 469]. docosahexaenoic [165, 164]. doctoral [1409]. Dodoma [2597]. Does [1245, 2298, 590, 2106, 2079, 649, 681, 193]. dog [2485]. dojo [831]. Dolichospermum [2087]. Dombes [665]. domestic [948]. domesticated [2338]. Domestication [928, 1213, 1207]. dominance [272]. domingensis [1334]. domperidone [150, 2747]. Don [1786]. donkey [1367]. donor [1281]. Donuzlav [2523]. dopamine [1215]. DORB [2340, 332]. d’Orbigny [681, 1694]. dorsalis [2404]. dosage [384]. dose [953, 1629, 1773, 620, 2480]. doses [1289]. dosing [1921, 1131]. Dosinia [1485]. dotted [2530]. double [2030, 1705]. double-stranded [1705]. dover [122]. Downstream [2506, 2590, 2378]. DP [409]. DP/DE [409]. dredging [1749, 1726]. dried [2388, 1611, 396, 846, 524, 210, 990]. drifted [2493]. driven [2750]. driver [2648]. Drivers [2423, 2411]. driving [2376]. droebachiensis [968]. drop [2195]. droplet [1208]. drug [693, 2635, 2451]. drum [1020, 1906, 675, 269, 2640, 2130]. dry [749, 9, 796, 833, 964, 694, 211, 44, 670]. dryer [2586]. dual [326, 1969]. dubius [1989]. ducks [632, 292]. duckweed [2298, 1396, 451]. due [1353, 1083, 1304, 2272]. dumerili [495, 1084]. dumerilii [694]. Dunaliella [2713, 1446, 2678]. dung [398]. Dunham [553]. Dunker [954]. duorarum [1202]. dura [2349]. duration [2320]. During [228, 150, 1635, 507, 161, 2090, 867, 1377, 2532, 398, 302, 2419, 49, 1889, 203, 1052, 1228, 2126, 1487, 1399, 2378, 2447, 696, 2428, 409, 2521, 1681, 525, 1553, 1097, 2316, 1073, 566, 1582, 1150, 1143, 1357, 1208, 2563, 1601, 14, 951, 469, 1454, 2537, 92, 1096, 67, 1712, 1296, 1221, 1452, 2577, 2539, 750, 952, 1273, 557, 135, 669, 1003, 2575, 2662, 425, 1759, 2711]. dusky [1270, 1830]. Dutch [807, 23]. duvaucelii [2676]. dwelling [287]. Dybowski [2375]. dying [2602]. Dynamic [995, 2002, 704, 851]. Dynamics [1150, 1143, 1208, 2289, 1915, 1134, 795, 1672, 2385, 640, 1136, 769, 1278, 2345, 657, 854, 204, 1746, 2669, 1424].

e-Learning [1415, 1414]. e-mail [565]. E. [727, 1404, 2379]. E2 [1211]. E20 [1757]. each [2438]. ear [1367]. Early [447, 123, 1209, 2200, 1332, 1178, 1085, 1720, 2241, 2088, 1140, 981, 1227, 1459, 2538, 59, 205, 314, 1766, 1228, 1640, 351, 2543, 1080, 696, 2249, 585, 1097, 2316, 913, 1983, 28, 403, 750, 407, 580, 135, 2465, 1483, 1633, 425, 2424]. Early-weaning [1332]. earth [1500]. earth-chitosan [1500]. Earthen [510, 1888, 2673, 633, 388, 754, 756, 602, 619, 2615, 1104, 2718]. earthen-based [388]. Earthscan [1008]. Earthworm [2293]. earthworms [2346, 47]. East [1904, 2626, 2752, 1368]. Eastern [1275, 1748, 2274, 1210, 2555, 2417, 1130, 1778, 1027]. eating [945]. ebulus [2390]. EC [663]. echinata [1952]. Echinoculture [1365]. Echinodermata [713, 561, 923]. Echinoidea [561, 923]. Eclipta [1252]. eco [2337]. eco-friendly [2337]. Ecological [571, 765, 2620, 2509, 2132, 1549, 1385, 411]. Ecology [474, 863, 864, 955, 1749, 1726, 860]. Econometric [886, 1004, 888]. Economic [866, 673, 2360, 2135, 212, 2279, 1836, 2366, 751, 2499, 1349, 1042, 2489, 1674, 1689, 1512, 2699, 2421, 2601, 1520, 2239, 1724, 1456, 2376, 575, 2343, 176, 1335, 1497, 765, 1325, 1027, 1400, 1973, 1378, 2645, 2531, 2243, 292, 952, 2620, 911, 1660, 2710, 1355, 1008]. economical [714]. economically [1917, 2504, 2044]. Economics [679, 1298, 992, 32]. economies [963]. economy [962, 2012]. EcoPlates [2186]. Ecosystem [1853, 448, 2132, 2052, 1451]. ecosystems [1099, 1793, 984, 2174, 1530]. Ecotoxicological [1910]. ectoparasites [295]. Ecuador [1237]. ECVET [2117]. ed [329, 411, 726, 678, 727, 957, 991, 1072, 1151, 1152, 448, 725]. edema [2033, 2129]. edible [1610, 829]. Edited [552]. Edition [785, 692, 1285, 1283, 680]. Editor [296, 35, 73, 110, 1837, 119, 1518]. Editorial [450, 1, 131, 319, 412, 52, 99, 275]. Editors [307, 274, 162, 365, 366, 353]. edn. [716, 736, 860, 861, 862, 992]. edodes [1846, 1192]. eds [564, 354, 449, 447, 554, 563, 650, 690, 688, 679, 708, 700, 709, 774, 863, 864, 887, 956, 974, 1070, 1071, 1039, 1056, 1436, 1975, 446, 625]. edspubs.co.uk [565]. EDTA [1869, 2348, 2101]. Education [1412, 1407, 600, 1911, 1973, 1414]. Educational [1973, 1911, 1415]. edulis [681, 1245, 1870, 1067, 1610, 1402, 2501, 804, 828, 1643, 558, 737, 718, 1276, 744]. Edward [552]. Edwards [354, 625, 2532, 982, 2596, 753, 930]. Edwardsiella [1548, 1866, 2273, 2238, 1922]. edwardsii [434]. eel [1049, 2147, 701, 2256, 256, 1093, 1267, 1177, 1357, 758, 1507, 387]. eels [203, 2149]. Effect [953, 1172, 1041, 1356, 1645, 150, 2617, 1085, 2404, 605, 901, 2664, 549, 2003, 1218, 2310, 1690, 1215, 1672, 946, 1262, 878, 1017, 1384, 1005, 266, 1293, 495, 2394, 2095, 398, 568, 2081, 260, 784, 2025, 899, 981, 1173, 2094, 2319, 257, 1103, 1457, 1860, 949, 919, 1328, 1784, 2590, 1766, 1052, 2348, 402, 2126, 1141, 1351, 1467, 1117, 986, 824, 1609, 1383, 1311, 2589, 623, 1706, 1629, 1364, 1664, 1855, 975, 2197, 877, 2328, 1137, 1014, 163, 1217, 2459, 1211, 1499, 1929, 1265, 879, 1126, 1449, 2282, 239, 894, 1308, 1268, 935, 1542, 1816, 1765, 2472, 1576]. Effect [550, 37, 601, 10, 1471, 1147, 771, 1084, 2034, 624, 1388, 1685, 338, 333, 326, 280, 1687, 2035, 493, 382, 566, 173, 1554, 165, 108, 1495, 2041, 1507, 2213, 2175, 1928, 1082, 1918, 1428, 464, 2595, 1767, 931, 17, 2537, 30, 442, 1091, 1432, 1296, 1873, 1830, 2181, 1010, 1516, 648, 1758, 753, 930, 641, 453, 2470, 826, 653, 1595, 164, 2056, 2683, 1466, 1867, 2652, 1662, 685, 1840, 1012, 1350, 1148, 940, 920, 1159, 1931, 1161, 905, 1405, 870, 1021, 1649, 2186, 939, 1731, 2686, 2711, 2091, 2453, 519, 1196, 286, 2131, 2062, 444, 1652, 642, 749]. effect [430, 1817, 801, 1869, 1213, 1610, 867, 1780, 2455, 2399, 336, 1566, 254, 301, 2725, 2434, 1993, 1314, 1381, 2299, 802, 872, 846, 2743, 196, 2522, 705, 406, 876, 1587, 1841, 1207, 915, 1856, 1169, 988, 115, 65, 1370, 683, 1230, 2075, 703, 1602, 1224, 1736, 2428, 976, 2623, 794, 1925, 1934, 1940, 359, 1833, 2203, 2473, 2721, 1796, 1823, 1368, 2675, 501, 328, 1998, 1131, 1704, 1475, 1994, 1599, 28, 1034, 1658, 1783, 759, 1165, 2127, 1096, 343, 311, 331, 822, 2706, 1163, 1847, 41, 1107, 672, 1539, 154, 1711, 1132, 2300, 891]. effect [865, 2402, 1366, 1543, 697, 1875, 2720, 2021, 1219, 1527, 592, 1343, 1530, 2510, 768]. effected [1941, 2057]. Effective [2529, 2063, 441, 1919, 2236, 107, 1858]. effectively [2223]. Effectiveness [2330, 1220, 578, 1928, 1161]. effectivity [655]. Effects [2242, 2370, 1448, 2722, 1020, 2740, 248, 731, 1763, 1446, 1255, 368, 1134, 1354, 1720, 728, 1049, 300, 1971, 1716, 2015, 638, 684, 827, 909, 1100, 1238, 279, 2385, 2114, 2313, 2297, 934, 996, 1186, 358, 438, 458, 773, 832, 816, 2240, 779, 1229, 1874, 1081, 1419, 1396, 1891, 1680, 1379, 2607, 1028, 2230, 1340, 1813, 1538, 1047, 1472, 1624, 1188, 351, 1792, 1266, 2024, 1109, 1641, 2738, 1235, 1023, 944, 1552, 1642, 554, 408, 114, 1162, 1596, 916, 1422, 2183, 1473, 2492, 471, 2693, 1486, 1481, 1442, 631, 1786, 604, 1278, 409, 1244, 1806, 839]. Effects [932, 1423, 1636, 2526, 2032, 2449, 1977, 2717, 1560, 2395, 2463, 2322, 1267, 1453, 2220, 2554, 536, 2377, 1850, 1726, 186, 342, 2121, 2316, 1879, 1741, 221, 2514, 268, 622, 913, 2642, 1800, 2151, 1893, 226, 732, 367, 1478, 1431, 1661, 775, 1243, 898, 379, 434, 67, 922, 369, 1785, 19, 702, 820, 742, 431, 133, 2585, 292, 513, 1719, 24, 1900, 2425, 854, 1374, 2098, 172, 628, 1367, 1030, 1890, 611, 2698, 1625, 724, 762, 814, 1003, 1735, 1739, 1702, 1617, 1857, 1820, 2495, 2382, 2432, 2556, 2742, 1029, 589, 1618, 1794, 2634, 363, 674, 914, 1682, 2420, 1035, 1517, 2536, 89, 735, 782]. Effects [1106, 1633, 1139, 1280, 1455, 1613, 1729, 1604, 1698, 2115, 2245, 1149, 745, 1503, 2318, 2569, 2201, 1178, 1033, 2031, 1638, 2612, 2654, 2088, 2248, 825, 212, 1319, 299, 2653, 1511, 9, 1253, 161, 2246, 340, 2258, 1377, 2366, 1612, 350, 391, 1310, 2418, 2593, 2221, 924, 12, 1013, 1793, 1064, 2363, 2733, 1844, 769, 883, 2748, 2553, 701, 1541, 1584, 1124, 1513, 897, 830, 459, 912, 1291, 994, 1313, 1068, 1417, 8, 22, 2708, 720, 2355, 1094, 2371, 1002, 1040, 2739, 1747, 1480, 1725, 1598, 11, 1699, 2672, 2408, 2047, 303, 1460, 2365, 2580, 2516, 190, 1497, 2511, 2321]. effects [1297, 1484, 2694, 2013, 1590, 517, 518, 1279, 2636, 1254, 1773, 46, 437, 181, 1184, 937, 1454, 1559, 1287, 1880, 156, 2285, 2393, 2496, 1978, 362, 837, 2562, 968, 143, 1775, 387, 1666, 1646, 1876, 1795, 1182, 1189, 1299, 2435, 997, 1657, 1607, 1360, 2152, 1727, 1540, 1592, 2469, 2424, 1764, 1749, 2753, 554]. Efficacy [2419, 2251, 2747, 2315, 1740, 1314, 2460, 2045, 2321, 767, 1601, 395, 2409, 1250, 1809, 2333, 2724, 2069, 1752]. efficiencies [2158, 817, 963, 952]. Efficiency [2468, 1989, 226, 2699, 1058, 1789, 549, 1627, 1672, 979, 1743, 2668, 2673, 712, 773, 603, 1427, 1265, 2327, 2487, 855, 1808, 190, 338, 1475, 1886, 1579, 2178, 2105, 1659, 1030, 1693, 1146, 363, 1191, 2167, 1942, 2456]. efficient [2601, 2584, 2044]. effluent [1253, 2589, 2448, 1485, 2292]. effluents [1253, 1395]. Egg [1310, 1227, 511, 239, 1178, 69, 1303, 1133, 444, 1218, 910, 351, 273, 406, 876, 1927, 382, 14, 776, 2157, 1930, 100, 1710, 1876]. Eggs [226, 230, 1098, 38, 345, 129, 364, 1472, 1624, 1903, 1534, 2544, 1470, 2282, 1156, 1934, 440, 2457, 373, 76, 667, 79, 1987, 60, 137, 1936, 82, 1223]. Egypt [2093, 2248, 1551, 2116, 2379]. Egyptian [2133, 2060]. Egyptian-farmed [2133]. Ehrenberg [825, 357]. either [767, 395]. Elacatinus [2696]. Elastomer [1053]. electrochemical [2715]. electrochemically [1461]. electrolyte [1785]. electrolytes [136]. electrolytic [1831]. electrolyzed [2565]. electron [2380]. elegans [732]. elemental [2374]. elevate [1271]. elevated [1035]. elevates [1105]. elevation [1472]. Elgar [552]. elicits [1962]. eliminates [7]. Elimination [872, 429, 1832]. ELISA [33, 855]. elliptica [2370]. Elodea [2720]. embryo [531, 659, 2544, 2282]. embryogenesis [1108]. Embryonic [1178, 2524, 2207, 441, 1952, 2563, 1918, 1108]. Embryos [570, 618, 778, 2356, 1147, 337, 1623, 1758, 204, 1219]. Emergence [2116]. emerging [1671, 2670, 2416, 1522, 2557, 32]. emersion [279]. emissions [1530]. emitting [1819]. emlusion [2290]. emodin [1293, 1474]. emphasis [203, 1768, 46, 667]. empirical [2266, 2476, 1571, 1693]. employment [1411, 1898]. empurau [2433]. EMs [2063]. emulsified [157, 1158]. emulsions [430, 2633, 166]. encapsulated [1261]. encapsulation [2722, 1901, 2561]. encephalopathy [1797]. Enclosure [1666]. encoding [2288]. Encyclopedia [296]. encysted [2248]. end [1061]. Endangered [479, 1753, 715]. endocrine [11]. endogenous [2059]. endoglucanase [896]. endoparasitic [2530]. energetic [2501, 1054]. Energy [1827, 882, 801, 909, 2673, 802, 705, 1023, 2002, 1181, 156, 922, 2397, 1451, 1942, 2456, 2568]. Enfield [678]. engagement [2476]. engineering [2001, 691]. Engle [679, 689]. English [2528]. enhance [1342, 2618, 2232, 531, 647, 1886]. Enhanced [218, 1819, 1437, 219, 1888, 1995, 2437, 2436, 1999, 2503, 2462, 1785, 985]. Enhancement [307, 958, 1145, 1658, 1801, 2029, 2410, 138, 278, 927, 2231, 1469, 2038, 1280, 1783]. enhances [2430, 134, 1912, 191, 2680, 1676, 2736, 1914, 1757, 2156, 2346, 1713, 2223, 1403]. Enhancing [1791, 241, 2040, 2705, 1087, 1646]. Enlarging [503]. enough [2578, 2571]. Enriched [514, 1995, 2088, 1467, 1823, 78, 428, 770, 798, 2551, 1104, 67, 900, 468, 1025]. Enrichment [2260, 1241, 433, 465, 2036, 430, 1084, 166, 501, 17, 539]. enrofloxacin [693, 1293, 1921]. Enterobacter [1821, 2682]. Enterococcus [1894, 987, 1069]. Enterocytozoon [1904, 1909, 2546]. Enteromorpha [1361, 1403]. enterotoxin [1592]. enterprises [2012]. Entirely [232]. entrainment [2069]. envelopment [2044]. Environment [512, 1734, 2714, 786, 1459, 2351, 2684, 50, 1046, 2249, 889, 1802, 624, 2682, 103, 2145]. environment-dependent [786]. Environmental [1790, 1578, 2104, 1748, 1064, 20, 1236, 1095, 607, 1871, 54, 62, 1971, 106, 600, 1703, 1575, 2351, 182, 1985, 1068, 3, 365, 1850, 978, 542, 1746, 774, 329]. environmentally [529]. environments [1256, 717, 768, 289, 468, 2145]. enzymatic [2558, 2328, 2717, 1560, 1994, 2711]. Enzyme [385, 1708, 2513, 549, 2003, 2313, 1780, 1626, 1028, 1563, 2452, 1827, 1486, 1882, 932, 1927, 2322, 361, 1267, 2473, 2518, 2035, 1063, 913, 896, 1145, 1254, 2563, 400, 1091, 2462, 1138, 2000, 1653, 1348, 1108, 2324, 2720, 735, 782, 905, 2318]. enzymes [2560, 1477, 344, 2427, 2558, 2743, 1239, 1399, 2335, 2485, 2428, 2032, 1556, 2316, 770, 798, 1873, 1824, 733, 537, 2361, 2646, 2219, 1613, 1503, 2424]. enzymology [2595]. EPA [1628]. epibionts [303]. epibiotic [82]. Epidemiological [1548, 1608, 1390]. epidermal [1242]. epilimnion [326]. epilistictus [2530]. Epinephelus [866, 695, 1635, 2687, 2754, 2733, 314, 2748, 2530, 1364, 2717, 550, 2022, 1872, 1732, 1184, 739, 921, 1830, 1069, 1404, 1868, 985, 2096, 1455, 2082]. epinephrine [108]. episquamalis [2023]. epithelial [746]. Epsom [2740]. EQF [2117]. eques [2363]. Eradication [700, 1738]. ERASMUS [1413]. ERGIC [2389]. ERGIC-53 [2389]. Eriocheir [2114, 1680, 1956, 2227, 1142, 929, 932, 2672, 2395, 931, 753, 930, 2373, 2361, 2382, 2548, 2688, 1799, 2500, 1649, 2639]. Eriwe [2214]. erizo [79]. Erratum [142, 497, 574, 888, 864, 1751]. erythrinus [503, 504]. erythrogram [2629]. Escape [1083]. Escape-related [1083]. escapees [528]. Eschscholtz [406, 723, 752, 876, 1117, 1266, 917]. esculentum [2522, 686]. Esox [659, 2459]. Essential [239, 2656, 867, 2651, 2419, 2299, 2460, 1241, 1855, 2633, 1833, 2621, 2035, 925, 2457, 2290, 135, 1890]. essentiality [428]. EST [1707]. Establishing [2438, 1553]. establishment [2092, 343]. Estampador [933]. esterified [1124]. esthetic [2068]. Estimate [818]. estimated [671, 437]. Estimates [1734, 2246, 1468, 1246, 1459]. Estimating [1456, 2507]. Estimation [2713, 1622, 2292, 577, 1903, 1273]. Estonia [482]. estradiol [2741, 291, 1211, 1317]. estradiol-17 [1317]. estuarine [1433]. Estuary [610, 1804, 278, 158]. ethanol [2357]. ethanol-free [2357]. ethanolic [2619]. ethical [714]. Ethiopia [2198]. ethoxyquin [1882]. etiologies [2734]. Etroplus [2232, 2206, 1516]. EU [1409, 1520, 2239, 1456, 2317]. Eucheuma [2257]. Euclinostomum [2379]. Eugenol [1884, 1599, 2224]. Eupallasella [479, 573]. Euplotes [1240]. Eurasian [2287, 1929, 1171, 676, 486, 1928, 1943, 363, 443]. Europe [1803, 1519, 1411]. European [1519, 2031, 904, 88, 9, 70, 39, 1557, 340, 2258, 1402, 2673, 1407, 1226, 529, 203, 1766, 1774, 1115, 1347, 701, 2256, 912, 1643, 414, 1211, 2193, 1408, 1308, 2625, 1093, 1193, 606, 624, 1519, 1177, 1357, 309, 1428, 250, 1946, 1945, 2157, 1937, 2020, 1868, 1948, 1947, 714, 1175, 1543, 2005, 607, 1988, 2108, 272]. Eurospan [565]. euryhaline [1404]. euryhalinus [1354]. Eurytemora [2467]. Euterpe [2100, 2569, 2753]. euthanasia [1718]. eutrophic [655, 1834]. eutrophication [144, 1528, 2186]. evacuation [93, 813, 206]. evaluate [1740, 1568, 2226, 2069]. evaluated [1180, 1276]. Evaluating [2243, 2375, 1321, 740]. Evaluation [145, 1851, 574, 1286, 1490, 2427, 908, 1377, 2651, 1829, 396, 1202, 2338, 1043, 2140, 719, 1861, 2231, 696, 1614, 2408, 2707, 555, 2737, 2210, 78, 1506, 389, 2592, 1639, 2257, 1372, 2409, 1414, 1677, 2594, 1053, 2301, 1025, 544, 990, 656, 699, 1290, 2660, 551, 764, 1836, 2155, 1743, 129, 2050, 1465, 456, 2161, 1821, 2351, 1380, 372, 2614, 1214, 206, 682, 1066, 58, 2327, 2487, 2343, 855, 1335, 2429, 80, 346, 1027, 2312, 1994, 2586, 1234, 79, 1200, 60, 137, 1866, 292, 1042, 1128, 1693, 2038, 857]. Evaluations [417]. Evechinus [697]. event [528]. events [2268]. Evidence [161, 2493, 2266, 1950, 2240, 786, 2476, 2412, 1788, 537, 646, 2692]. Evolution [693, 411, 203, 860]. ex [1616, 1668]. examination [586, 2465, 1105]. Examining [2237]. example [2719, 1488]. ExbB [2333]. excavatus [2293]. Excel [2700]. exchange [523, 2610, 764, 1253, 1186, 1567, 2094, 631, 756, 2449, 442, 1935, 1462, 2432]. excretion [71, 2367, 1351, 1143, 1447, 704, 1625, 762, 940, 443, 592, 745]. exemplified [50]. exercise [8]. exercised [29]. exerted [2047]. Exiguobacterium [757]. exogenous [2059, 1028, 833, 1542, 2143, 1503, 2424]. Exopalaemon [2097, 2271]. exopolymeric [1490]. exopolysaccharides [2679]. expanded [37]. expansion [960]. expected [617]. expensive [1040]. experience [1439, 1249]. experienced [1969]. experiences [1336]. experiment [356, 1666]. Experimental [699, 106, 1740, 571, 1797, 1371, 2402, 933, 1397, 635, 1463, 1514, 1136, 2680, 2206, 2159, 400, 464, 1112, 2069]. experimentally [2123]. Experiments [660, 652, 2493, 177, 1365]. Exploitation [803, 2067, 812, 726]. exploited [596, 597]. Exploration [2681, 2598, 1615]. Exploratory [2527, 646]. Exploring [706, 2239, 2741, 2583, 1062]. export [2574, 1582, 2669]. export-oriented [2574]. exporting [1950]. exports [2051, 2317]. Exposed [240, 233, 237, 2450, 907, 148, 325, 2463, 2329, 360, 2397, 1430, 2144, 905]. exposure [2137, 1326, 1626, 1629, 2428, 1816, 2365, 2322, 280, 177, 672, 734, 593, 2368, 2112]. expressed [2579]. expressing [2608, 1705]. Expression [1140, 2372, 249, 2039, 2502, 2389, 1851, 2702, 2559, 1000, 1384, 856, 2741, 1293, 2114, 2313, 2280, 2232, 2434, 1715, 2325, 1419, 1396, 2638, 2712, 2637, 2378, 2276, 2328, 2442, 1914, 2362, 1816, 2209, 2334, 2511, 2316, 1879, 2200, 2563, 2591, 1287, 1562, 2216, 2706, 1163, 1805, 1338, 2309, 2099, 2520, 1331, 2678, 1966, 1189, 1299, 2275, 2441, 2435, 2440, 2222, 2168, 2288, 2368, 1729, 2082, 1649, 2228, 2130]. expressions [2482]. extend [1920]. Extended [2544]. extensive [673, 2517, 523, 1290, 1525, 1168, 1786, 1807]. extensive/semi [1290]. extensive/semi-intensive [1290]. extent [1050]. extra [1371, 1010]. extracellular [943]. extract [2552, 699, 2612, 2404, 645, 1384, 1715, 2522, 2553, 1135, 1169, 1664, 2681, 2009, 1920, 2484, 2554, 2516, 2153, 2619, 2257, 2647, 822, 2273, 2011, 2639]. extracted [2331, 1867, 1192]. Extraction [2037, 2396, 2058]. extracts [2294, 2736, 2683]. extruded [2725, 846, 1453, 2035, 1234, 2188, 611, 1682]. extrusions [2106]. Exudates [232]. eye [758]. eyed [748]. eyes [165].

F [2244]. F. [992, 1437, 1370]. Fab. [352]. faba [1496, 1496]. FABP [1729]. fabrication [2586]. Fabricius [569, 1773, 2706, 847]. Facey [860]. facilitates [2549]. Facilities [506, 1534, 835, 1712, 2236, 2364]. facing [1115]. factor [1385, 2368, 1373, 1444]. factor-1 [2368]. factor-2 [2368]. factor-3 [2368]. Factors [485, 617, 2317, 1074, 2283, 1790, 2421, 2059, 2666, 62, 106, 2114, 2313, 2014, 2494, 1703, 1996, 1642, 525, 173, 46, 1693, 911]. faecal [1458]. faecium [1894, 987, 1069]. failure [1932]. fairy [1754]. fallax [1370]. fallowing [769]. Fallu [565]. false [1536, 2195]. families [1281, 1457, 2052, 2334, 766]. family [2135, 2460, 1487, 1480, 1273, 2108]. family-based [1487]. fangsiao [1694]. fantail [1991]. Farfantepenaeus [1804, 981, 1202, 2106]. fario [1207]. farionis [1982]. farm [1915, 470, 795, 1997, 2240, 751, 640, 203, 281, 2624, 2690, 2589, 1002, 2052, 2448, 1615, 1485, 538, 1620, 906, 1236, 2454, 539, 386, 2326, 1840]. farm-based [281]. farm-raised [1615]. Farmed [506, 690, 945, 2142, 1531, 2248, 901, 2122, 2660, 1281, 535, 1869, 989, 2225, 1568, 2133, 2023, 2483, 779, 2154, 2638, 2348, 372, 1985, 1589, 2583, 1442, 960, 1681, 1765, 2625, 48, 2159, 1684, 2644, 1571, 1982, 2405, 1073, 1964, 46, 2369, 2194, 2028, 1762, 840, 1930, 202, 2588, 1933, 1062, 557, 1175, 646, 2048, 936, 1443, 1657, 1875, 2144]. farmer [1660, 565]. farmers [1513, 1984, 2615, 835, 2141, 2349, 2445, 1788]. Farming [551, 552, 1071, 2176, 1721, 673, 1058, 2517, 523, 2135, 1679, 439, 2155, 1748, 1005, 1899, 1853, 2366, 144, 1963, 2214, 1891, 1289, 2555, 416, 2624, 2376, 1584, 682, 2110, 1298, 1427, 1549, 366, 2476, 2327, 2487, 2381, 1522, 1079, 1264, 2721, 75, 352, 74, 1275, 1787, 2006, 2551, 2272, 2648, 534, 1967, 20, 963, 1898, 1183, 2295, 1112, 978, 2529, 1788, 1246, 1394, 32, 599, 2620, 911, 2710, 2217, 1674, 574, 680, 1039, 466]. farms [1917, 848, 693, 658, 2499, 600, 1793, 31, 1393, 66, 125, 1395, 984, 6, 617, 2306, 1717, 1979, 858, 2044, 1051]. farreri [1176, 1438, 873]. fasciata [2356]. fasciatus [1757, 1369, 1363, 2534, 939]. fascicularis [781]. fast [146, 1373]. fast-growing [1373]. faster [681]. fasting [2469, 2711]. Fat [2074, 1076, 2418, 180, 1583, 983, 1234, 670, 1662, 363, 1317]. Fate [2284, 1930]. Fattened [250]. fattening [2737, 2141]. Fatty [1696, 2461, 239, 2311, 1097, 634, 250, 100, 2560, 2173, 1907, 1255, 2414, 1098, 2062, 1133, 444, 910, 867, 934, 2651, 2418, 784, 2128, 899, 1173, 1028, 1784, 1141, 818, 1188, 1117, 1109, 1841, 1241, 1552, 912, 994, 915, 1311, 821, 1643, 1014, 1803, 37, 2097, 501, 983, 46, 982, 373, 667, 1454, 464, 1767, 2697, 17, 1096, 1296, 1163, 900, 202, 1505, 2594, 2562, 2588, 1653, 959, 1524, 930, 2383, 135, 2300, 1694, 669, 1029, 1662, 914, 2087, 1425, 1106, 1581, 2219, 1405, 1545]. fauna [2369]. Fe [1540]. feasibility [2360, 2279, 1810, 1521, 2672, 1031]. feasible [380, 2074]. feather [1316]. feces [2006]. Fecundity [489, 248, 2237, 659]. Fed [232, 235, 245, 510, 514, 2552, 2410, 1133, 799, 1708, 1445, 882, 1358, 1277, 430, 270, 910, 979, 423, 712, 2725, 1465, 1226, 2084, 1538, 1563, 2045, 1399, 2738, 1224, 1882, 322, 550, 359, 1571, 606, 1814, 2405, 2518, 524, 770, 798, 1886, 1447, 982, 47, 1234, 427, 1863, 30, 921, 2262, 1339, 630, 900, 1271, 2372, 1044, 2188, 959, 204, 436, 1694, 2520, 2723, 1955, 2465, 1966, 1537, 1022, 1500, 1317, 2211]. Feed [25, 299, 251, 579, 148, 2482, 437, 510, 1172, 1801, 2617, 1076, 799, 549, 642, 749, 882, 1716, 2190, 1611, 909, 2293, 2467, 908, 1743, 867, 545, 2455, 2399, 681, 996, 603, 780, 2558, 1081, 1226, 1766, 1340, 1563, 1467, 1399, 1841, 2738, 701, 1552, 2543, 2231, 1771, 1632, 1364, 821, 1919, 2075, 1048, 2492, 2693, 1895, 163, 180, 1265, 1614, 2623, 2719, 1699, 1560, 550, 37, 1768, 1123, 2580, 1585, 624, 338, 342, 1687, 845, 2013, 268, 1475, 2346, 1886, 1369, 964, 2551, 1846, 634, 2340, 312, 766, 1880, 2603, 18, 2611, 1007, 2372, 1978, 2588, 2191]. feed [2700, 44, 1030, 2616, 1128, 1466, 2678, 1437, 556, 685, 363, 914, 940, 1807, 1106, 999, 1038, 1398, 2204]. Feed-based [2482, 1801, 1919]. Feed-growth [25]. feed-mixing [2700]. feedback [2487]. feeder [540, 834]. feeders [4, 2603]. Feeding [681, 2725, 2299, 2352, 1118, 781, 3, 353, 7, 29, 511, 621, 352, 221, 1369, 494, 509, 403, 508, 2720, 2211, 1578, 4, 1789, 93, 2612, 62, 146, 299, 1770, 1716, 149, 2277, 45, 458, 603, 1874, 1013, 1174, 643, 813, 1784, 2367, 833, 1515, 662, 2744, 701, 1235, 1584, 994, 1632, 171, 1580, 2226, 1230, 1231, 1596, 958, 1481, 1265, 1725, 976, 426, 1765, 2395, 1312, 10, 298, 334, 1796, 624, 338, 342, 2233, 1297, 1590, 268, 1475, 226, 83, 797, 355, 775, 644, 884, 921, 210, 67, 1007, 19, 704, 742, 1591, 837]. feeding [2188, 2073, 44, 1225, 2300, 669, 1016, 2465, 741, 814, 1794, 1682, 89, 1581, 592, 761, 1149, 2228, 2406, 2568]. feedings [1886]. Feeds [227, 2298, 1923, 846, 1347, 1164, 2071, 2416, 696, 1370, 2371, 2521, 853, 2541, 524, 1691, 737, 2586, 1234, 2531, 2020, 2180, 2726, 2562, 1682, 956]. fem [2164]. fem-1b [2164]. female [1194, 2618, 2741, 1213, 1229, 291, 2733, 2748, 273, 455, 1747, 2212, 1908, 2717, 2405, 1096, 2164, 2652, 2265]. females [270, 1214, 1828, 822, 1029, 1195]. Feminization [1688, 1317]. feminized [758]. fenbendazole [2409]. Fenneropenaeus [1743, 980, 1737, 2008, 1742, 1796, 982, 1087, 1185, 1059, 950, 1361, 1182, 734, 2038, 761, 905]. feral [869, 787]. fermentans [2082]. fermentation [1258]. Fermented [2346, 1445, 1396, 2533, 2473, 1846, 2435, 1807]. fern [2522]. fertility [588]. fertilization [2287, 1218, 1103, 321, 1151, 993, 1094, 715, 587, 2747, 826, 1219, 1223, 2510]. fertilized [764, 2157]. fertilizer [1041, 258, 384, 1066, 653]. fertilizers [94, 2541]. FIA [1835]. fibropapilloma [2133]. fiddlehead [2522]. Field [549, 2155, 256, 623, 1267, 326, 1036]. fields [848]. figaro [2696]. fighting [168, 2516, 1601, 1599, 1924]. filamentous [1858, 847]. fillet [2660, 1319, 1109, 1841, 1311, 2492, 1615, 445, 393, 1978, 2562, 2698, 1718, 1657]. fillets [2128, 1920, 1576, 394, 2056, 2011]. film [2740, 1514, 627]. filter [540, 2367, 1470, 2640]. filtration [1061, 1861]. fimbriatus [2513]. Fin [1212, 644, 424, 2280, 1591]. fin-fish [424]. final [1577]. Financial [1620]. findings [2129]. fine [1960]. fine-bubble [1960]. Finfish [353, 449, 2155, 1212, 714]. finger [2514]. fingerling [461, 671, 1116, 300, 1081, 817, 555, 2072, 1038, 1191, 1398]. Fingerlings [493, 237, 953, 1763, 908, 1764, 1028, 846, 1637, 2024, 206, 1066, 2183, 2009, 2428, 342, 2121, 2329, 2514, 1258, 197, 1053, 969, 360, 2646]. Fingerman [688]. fingerprinting [84]. Finishing [1581]. Finland [1264]. fire [403]. First [248, 149, 487, 698, 515, 479, 357, 1007, 1892, 2746, 44, 801, 545, 636, 314, 1421, 462, 298, 1208, 921, 2160, 468, 1466, 874, 1875]. Fish [53, 1839, 485, 222, 567, 297, 330, 565, 554, 690, 1523, 484, 2649, 75, 602, 619, 1019, 532, 978, 1448, 2540, 2266, 457, 656, 1099, 2059, 1915, 344, 945, 2414, 77, 499, 693, 385, 300, 649, 2660, 1261, 658, 2253, 71, 1166, 424, 795, 2458, 910, 979, 2277, 2293, 1748, 1342, 1557, 1678, 2519, 1212, 2467, 2509, 1346, 1894, 751, 2170, 1626, 2499, 2019, 257, 600, 2214, 291, 1793, 1891, 2363, 1912, 203, 191, 384, 2590, 132, 1813, 533, 1345, 1538, 1047, 2748, 1188, 2624, 1587, 1109, 2744, 2071, 1841, 2416, 830, 1985]. fish [66, 1427, 677, 321, 1142, 8, 991, 1055, 1837, 2229, 977, 720, 1647, 125, 1473, 2371, 168, 1895, 1651, 1395, 676, 2672, 1803, 1757, 2047, 6, 2625, 2156, 37, 2345, 1264, 2516, 1101, 1571, 87, 2721, 1982, 2153, 55, 320, 326, 2390, 845, 2443, 2347, 2696, 2013, 2541, 657, 1691, 2252, 346, 1969, 983, 2407, 2557, 1886, 1984, 2615, 2312, 337, 1332, 797, 2648, 1601, 1599, 1478, 1760, 2250, 294, 445, 1165, 1946, 1945, 538, 655, 1797, 20, 1378, 1594, 2257, 835, 2697, 1074, 2645, 2478, 2369, 2340, 2353, 1880, 379, 2028, 393, 1258, 1924]. fish [1007, 60, 1112, 2255, 41, 704, 2412, 2746, 2445, 2562, 1044, 2529, 292, 972, 2396, 668, 1524, 1775, 2160, 1316, 2098, 1544, 2470, 826, 1910, 858, 653, 436, 539, 85, 1132, 1175, 32, 646, 386, 2402, 1462, 1488, 2275, 2326, 1543, 327, 911, 1203, 990, 940, 1392, 920, 2464, 1752, 2578, 2571, 2686, 2182, 727, 956, 1056, 1072, 564, 690]. Fish-farming [978]. fish-free [1880]. Fish-management [75]. fish-meal [2371]. Fisheries [498, 553, 2599, 2626, 2752, 1113, 1411, 1911, 2174, 716]. fishermen [1746]. fishery [1582, 2080, 447]. Fishes [221, 885, 2060, 183, 729, 2743, 860, 269, 1695, 736]. Fishing [564]. Fishmeal [2113, 942, 435, 846, 1776, 2521, 2473, 2035, 961, 2020, 2588, 2318]. fishmeal-based [2020]. fishponds [457, 452, 655, 1586]. five [1696, 777, 1288, 91, 105, 2465]. fixation [1176, 2608]. fixed [45]. fjords [1915]. flagellatus [2636]. flaked [2623]. flat [1245, 88, 1402, 560, 1643]. Flatfishes [726]. flathead [2023]. flatworms [2027]. Flavobacterium [2380, 855, 2644]. flavones [2418]. flavor [1615, 1274, 665]. fleet [2529]. Flesh [132, 579, 482, 2638, 2084, 1658, 1783, 840, 2549, 1139]. flesus [1265]. Fletcher [1070]. flexibility [607]. flexuosa [1361]. floating [627, 6, 550, 809, 528, 2525]. floc [981, 2391]. flocculants [1006]. flocculation [1831]. flocks [918]. FLOCponics [2570]. flood [2615]. flooded [1074]. flora [385, 1122, 1994, 121, 1604]. florfenicol [1517]. Florida [158]. flounder [1418, 996, 1301, 2171, 2693, 1265, 1806, 1542, 1876, 432, 748, 920]. flour [880]. Flow [285, 2713, 1129, 1835, 548, 1555, 818, 1383, 1897, 520, 521, 2165, 1940, 6, 613, 1495, 1893, 2401, 2537, 311, 626]. Flow-through [285, 1129, 548, 1940, 613, 626]. fluctuation [788]. Fluctuations [2344, 1023, 1101, 2317, 2537]. fluid [1655, 454, 2747]. flumequine [328, 395]. fluorescence [2713]. fluorescens [2095, 2333]. fluted [1196]. fluvial [1777]. fluviatilis [1166, 2287, 1929, 1171, 1753, 486, 1928, 1943, 1937, 1221, 1222, 1938, 363, 443]. flux [1120]. fluxes [984, 1279, 1424, 1451]. fly [2706]. focus [2311]. focusing [168]. follicles [659]. Following [249, 146, 1301, 157, 1833, 1872, 1363, 395, 1805, 1069, 2441, 2440, 734, 2288, 2234]. fontinalis [481, 1233, 2160, 1158, 1581]. Food [495, 1210, 563, 1312, 274, 628, 1578, 2235, 4, 1354, 825, 192, 1673, 1870, 642, 1416, 1358, 764, 9, 687, 979, 198, 2673, 358, 350, 773, 2170, 1173, 435, 813, 919, 404, 1083, 1345, 2354, 662, 719, 1241, 1102, 2172, 388, 2492, 2574, 1054, 1420, 328, 1554, 1902, 964, 2622, 2127, 1594, 1842, 369, 41, 972, 1524, 1203, 2204, 761, 905, 1280, 1455, 2323, 1149, 2061]. footprint [2610]. forage [388]. forces [6]. Foreword [1406]. Forez [665]. formalin [441, 2171, 1910]. formalin-inactivated [2171]. formation [23, 916, 2516, 2216]. formic [1813]. forms [1921, 2585]. formulated [2617, 642, 910, 979, 2298, 696, 2708, 2623, 737, 1767, 1258, 2706, 630, 44, 1545]. formulation [2558, 1123]. formulations [1128]. Forskål [545, 546]. forskali [1193]. Forsskål [623]. fortuitum [2284, 2490]. fossil [2673]. fossilis [671, 1116, 1081, 817, 2263, 2391]. Fouling [469, 1304]. four [1220, 548, 2081, 305, 2408, 524, 393, 60]. Fractal [2043]. framework [1915]. France [416, 665, 144, 558, 2528]. franciscana [2370, 2617, 1780, 915, 1560, 2462, 1132, 2300, 1795]. Francisella [1762]. Free [2218, 1534, 659, 1231, 770, 798, 209, 2401, 2357, 1880, 921]. free-living [921]. freezer [2481]. frenatus [2362]. French [1032]. frequencies [644, 2384]. frequency [299, 2342, 2725, 1013, 1596, 1265, 338, 1369, 1773, 1296, 369, 742, 89]. Fresh [229, 1173, 384, 696, 1686, 427, 547, 211, 1366]. freshness [405, 445]. Freshwater [466, 221, 2726, 1917, 2109, 68, 69, 25, 1635, 2388, 1734, 2389, 2137, 1839, 1627, 1342, 2430, 2385, 2626, 2752, 1043, 1724, 1637, 1563, 2522, 304, 723, 752, 876, 1117, 1266, 2071, 2614, 1257, 912, 1827, 115, 114, 388, 29, 2387, 125, 1048, 1137, 1481, 2249, 754, 1126, 1292, 1385, 777, 1803, 917, 2721, 78, 1388, 1685, 2263, 845, 2675, 1832, 139, 1082, 1884, 2659, 1797, 2689, 1252, 2369, 2632, 442, 67, 393, 1785, 1007, 763, 822, 1112, 2396, 2425, 293, 733, 2542, 865, 2402, 1794, 823, 2247, 2087, 857, 1195, 2711, 366, 1039, 1282, 602]. Freund [1638]. friendly [2184, 2337]. fries [2433, 2450]. fright [890]. Frimodt [297]. fringe [2513]. fringe-lipped [2513]. frisii [964, 2747]. frog [2221]. frontier [1058]. frozen [458, 1481, 294, 559, 547]. fructooligosaccharide [1645, 2361]. fructooligosaccharides [1774]. Fry [240, 471, 1033, 128, 1133, 39, 878, 2025, 899, 257, 662, 661, 1642, 2291, 24, 2005]. FTA(R) [2358]. fucata [954, 1506, 2246, 719, 1505]. Fucoidan [1665, 1501]. fucose [2238]. FucP [2238]. Fucus [1715, 1882]. fuels [2673]. Fuiman [447]. Fujian [2735]. fulgens [2652]. full [1457, 2288]. full-length [2288]. full-sibling [1457]. fulvic [1860]. fulvidraco [2297, 881, 1488, 1733, 1665, 1509]. Function [506, 1326, 1254, 2265, 870, 1343]. Functional [2189, 2071, 2097, 2648, 2332, 2441, 2440, 612, 2089, 2061]. functions [2736, 2665]. fund [2499]. fungal [1470, 2608, 1846]. fungi [1858]. fungus [2390]. Funka [790]. Fusarium [2621, 2694]. fuscoguttatus [866, 1635, 2733, 2748, 2717, 1872, 1732, 2096]. fusiformis [1423]. fusome [2373]. fusome-related [2373]. Future [274, 2067, 1720, 2302, 1523, 1911, 2129, 1683, 2600, 599, 2326]. Fuzzy [834, 973]. Fuzzy-logic-based [834].

G [2190, 447, 708, 2054]. G. [860, 956, 991, 1070, 1284]. g6 [897]. GABA [1374]. gaditana [1814]. Gadus [1140, 1309, 639, 1110, 1439, 1046, 1431, 2695, 620, 431, 2112]. gain [1018, 2076, 1398]. gainful [1898]. Gaining [1951]. gal [249]. galanga [1384]. galangal [1384]. Galaxea [781]. galbana [2653, 1841, 1609, 164]. Galicia [103]. Galician [636]. galloprovincialis [1772, 2062, 1838, 2523, 1305, 916, 558, 1602, 1553, 2004]. Gambier [1032]. Gambusia [1797]. gamete [587]. Gametogenesis [786, 509, 1419]. gametogenic [101]. gametophytes [1103]. gamma [1851, 1629]. Gammarus [1845, 2172, 904]. gaping [1305]. gaps [2274]. gardening [808]. gargal [1713]. gariepinus [1888, 492, 2142, 2190, 2559, 801, 813, 877, 993, 2034, 382, 1998, 2200, 2346, 1866, 2281]. garlic [2031, 2404, 1727]. garnish [1760]. Garrett [365]. garvieae [1323, 1852, 2123]. gas [212, 1530]. Gastric [93, 206]. Gastrointestinal [701, 1621, 2437, 2746]. gastropod [946]. Gastropoda [1892]. Gate [1412]. gavage [2612, 2689]. gayi [842]. Gayral [2383]. gear [281, 809]. gel [1638]. gelatinized [685]. Gelatinolytic [484]. gelatinosus [1912, 2084]. Gelidium [2215]. Gene [856, 249, 1479, 2579, 2114, 2313, 2280, 2232, 2434, 1715, 1396, 2637, 2328, 980, 1914, 1816, 2209, 2097, 2511, 2316, 1154, 2200, 2139, 2591, 2216, 2706, 1338, 1962, 2520, 2678, 1966, 2441, 2435, 2440, 2168, 2149, 2288, 2307, 1649, 2469]. generated [1548]. generation [852, 357, 1544, 2163]. generations [391, 1288]. Generic [1411]. genes [2370, 2502, 1140, 1384, 2342, 2638, 2712, 256, 2378, 2690, 2334, 1556, 1879, 2035, 2482, 2306, 2563, 1562, 2332, 1163, 2372, 2271, 1877, 2309, 2099, 2018, 2275, 2222, 2039, 2582, 2130, 1444]. Genetic [490, 310, 1119, 2714, 155, 1743, 1670, 84, 1459, 1565, 481, 954, 2535, 1080, 2447, 168, 777, 48, 1796, 1154, 1964, 553, 1502, 522, 526, 893, 2145, 2038, 1722, 147, 2182, 1734, 2246, 1212, 1622, 1457, 1854, 2162, 1382, 1756, 1755, 1848, 1468, 1737, 1908, 528, 1506, 2030, 2312, 2674, 1933, 787, 1114, 2048, 874, 1373, 1761, 2120, 1728, 1327]. genetically [1257, 1442, 1765, 87, 1443, 1657, 1840]. Genetics [498, 1113, 453, 1373, 330]. Genome [725, 2674, 1949, 2546]. genome-wide [2674, 1949]. Genomic [2302, 1743, 1362, 1544, 2546, 2163]. genotype [1734, 2714, 1459, 2145]. genotypic [2342]. genotyping [1429]. gentian [2733]. geoduck [1853, 1958]. geographic [771, 790, 1948, 1947]. Geographical [54]. geosmin [1492]. German [1067, 1600, 1362]. Germany [1656]. germinal [1328]. germplasm [893]. Geukensia [1174]. GFP [1257]. GGIV [2070]. GH [1556]. Ghana [1917, 1531, 2423, 1850, 2213]. Ghanaian [1386]. gher [1898, 963]. ghost [2161]. GHR [2099]. GI [2256]. Giant [466, 1196, 2109, 995, 1839, 2430, 2090, 786, 2522, 1257, 2490, 1126, 1278, 1292, 1385, 805, 78, 1388, 1685, 845, 1082, 2689, 1785, 19, 2425, 2070, 847]. Giardia [1064]. gibbesii [2400]. gibbus [626]. gibel [1925, 1500, 2469]. gibelio [738, 2463, 1500, 2469]. Gibson [726]. GIFT [1442, 1287, 1562, 1189, 1657, 2234]. giga [2714]. gigantea [1172]. Gigartinales [717]. gigas [1720, 2241, 760, 460, 1344, 786, 807, 2348, 2545, 1298, 1102, 805, 1692, 1778, 427, 1965, 806, 1712, 1892, 1482, 672, 1707, 2021, 1761, 1352, 2453]. Gill [500, 1940, 92, 2637, 1816, 1073, 950]. gills [1626, 2460, 1331, 2688]. gilthead [93, 882, 2133, 1622, 390, 1565, 405, 2367, 974, 1486, 58, 528, 770, 798, 309, 1332, 464, 156, 898, 2496, 844, 1089, 789, 1029]. Girard [1797]. GIS [54, 682]. GIS-based [682]. given [2298, 2611, 2180]. glacialis [1024]. Gladioferens [2170]. gland [2142, 907, 822, 1504, 293]. glanis [1226, 1164, 1211, 1988]. glassy [2478]. Global [2573, 2716, 1745, 563]. globin [183]. glucan [2363, 2251, 2708, 1806, 1388, 1185, 1021]. glucans [1993, 1769]. glucomannan [1729, 1604]. glucose [2363, 604]. GLUT [2680]. GLUT-4 [2680]. glutamine [2313, 980, 2742]. glutaraldehyde [364, 440, 137]. glutaredoxin [2325]. glutathione [2495]. Gluten [579, 2748, 1541, 1231, 1691, 1863]. glycerides [2594]. glycogen [1352]. Glycyrrhiza [2418]. Gmelin [986, 283]. gnotobiotic [915]. GnRH [255, 1215, 1160]. goby [2572, 2696]. gold [2636]. goldblotch [314]. golden [1854, 2074, 1392]. goldfish [2280, 2511, 1495, 2401, 1991]. goldsinny [635, 1572, 395]. Golfo [323, 401]. Golgi [2389]. Gonad [248, 2410, 2405, 801, 1419, 1032, 1274, 1767, 379, 1296, 968, 1366, 2652, 1029, 697, 1875, 1280, 1149, 1341]. Gonadal [2212, 1843, 695, 2523, 291, 1214, 1828, 1542, 2332, 2534]. gonadoliberins [1934]. Gonadosomatic [504]. gonadotropin [255, 1160, 1159]. gonadotropins [1934]. gonads [2164]. gonionotus [2296]. good [2351, 1776]. Goods [1525]. goramy [2284, 2490, 2070]. Gotukola [2041]. Gould [566]. gourami [2490, 2356, 1473, 1800, 2070, 2048]. governance [1264]. Governorate [2379]. Gracilaria [1772, 1334, 1887, 1001, 567, 1176, 2589, 879, 610, 1535, 2349, 1719, 990, 1503]. Gracilariales [610]. Gracilariopsis [1253, 2623]. grade [1281]. graded [942, 2492, 1339, 1500]. gradient [1433]. grading [300, 768, 2459, 154, 172]. graduate [1411]. grafter [1337]. Grain [243, 2533, 1749, 1726]. grains [1585, 1846]. Gram [1805]. Gram-negative [1805]. granulata [2566]. Grass [451, 628, 1980, 1238, 1293, 849, 603, 1187, 1459, 1999, 1040, 1699, 2032, 2707, 1030, 2179, 2288, 2565, 2146, 1883, 1451]. Grateloupia [2400]. gravel [520, 521, 627]. gravity [2422]. grayling [9, 70, 2311]. Grazing [2465, 451]. Great [2475, 349, 285, 286, 106, 260, 1170, 113, 1028, 1109, 2515, 918, 104, 613, 102, 356, 108, 101, 311, 112, 100, 580, 1276]. greater [694]. Greece [800, 1659]. Greek [499, 502, 252]. Green [2272, 327, 1041, 1356, 615, 2088, 1248, 1463, 1992, 1676, 1135, 2386, 2013, 339, 2506, 963, 2223, 1662, 2337, 2091]. green-lipped [615]. Greenhouse [1530, 1700]. greenhouses [2366, 2208]. Greenshell^{T M} [986, 2101, 2630, 2264, 1051, 2547]. Greville [1334, 717]. grey [549, 2023, 555]. Grifola [1713]. Griseb [1616]. groEL [1479]. ground [2739]. grounds [2410]. groundwater [623, 1183]. group [40, 675, 1948, 1947]. Grouper [739, 866, 695, 1635, 2687, 2754, 2733, 314, 2748, 2530, 1364, 2717, 2047, 550, 1974, 2022, 1732, 1184, 921, 1830, 1069, 1404, 1868, 985, 2096, 1455, 2082]. groups [890, 1325, 1400]. Grow [370, 614, 2521, 2596, 1547, 1452, 2585, 1404]. Grow-out [370, 2596, 1547, 1452, 2585]. growing [549, 455, 756, 409, 1522, 193, 80, 1373, 1530]. growing-out [756]. grown [282, 2386, 1614, 2345, 1450, 209, 1994]. Growout [548]. Growth [1801, 2552, 248, 232, 2029, 1133, 799, 159, 2513, 1628, 882, 299, 2190, 88, 282, 70, 979, 45, 2532, 568, 712, 235, 218, 1226, 222, 1703, 240, 251, 404, 2084, 1563, 838, 515, 462, 1102, 305, 579, 756, 1882, 1469, 1032, 805, 322, 401, 334, 610, 606, 2329, 493, 566, 91, 983, 226, 2508, 1714, 982, 482, 2596, 509, 371, 710, 2709, 47, 237, 1863, 312, 921, 654, 906, 1712, 510, 2180, 1339, 630, 1426, 900, 2188, 1723, 1496, 959, 154, 514, 847, 508, 241, 1966, 1537, 2669, 556, 2087, 1500, 242, 2568, 1178]. growth [1812, 2370, 2063, 1033, 1448, 519, 2729, 2699, 313, 1196, 2560, 671, 953, 1172, 1041, 1907, 1020, 1477, 2740, 731, 1645, 1446, 1255, 368, 2617, 1076, 788, 1134, 1354, 2286, 2241, 2062, 25, 2088, 728, 1838, 760, 1049, 947, 1734, 145, 1708, 300, 549, 825, 174, 903, 2003, 1627, 642, 749, 1319, 1251, 460, 2610, 1067, 1248, 1334, 2310, 1716, 2559, 1119, 1690, 9, 684, 827, 909, 946, 2277, 198, 1000, 1100, 1017, 2147, 2246, 2430, 2427, 279, 2298, 472, 2313, 1122, 1669, 2651, 1558, 2394, 2455, 2399, 996, 1186, 2095, 1612, 350, 438, 458, 773]. growth [832, 816, 1256, 1566, 254, 603, 2085, 2304, 981, 1202, 780, 966, 2725, 2434, 1874, 2319, 2221, 1081, 302, 257, 1419, 1622, 1396, 1043, 1459, 2363, 643, 1103, 1457, 1860, 2638, 2607, 949, 203, 919, 384, 1676, 2299, 1328, 1028, 729, 1784, 2736, 872, 846, 1474, 1766, 1849, 2230, 1818, 1779, 1340, 402, 2743, 281, 2352, 883, 1538, 1701, 1047, 1077, 1141, 2436, 1118, 2522, 1896, 1188, 1999, 1439, 752, 1266, 2024, 1109, 1399, 986, 1841, 2738, 701, 698, 841, 1090, 1878, 1541, 1769, 1023, 897, 944, 1320, 824, 1552, 2335, 459, 1383, 1135, 1291, 1427, 994, 1642]. growth [647, 677, 768, 1313, 2231, 2589, 623, 1706, 2631, 2452, 1629, 633, 1364, 2002, 1080, 988, 8, 3, 22, 50, 843, 171, 408, 115, 114, 7, 29, 65, 1370, 975, 1162, 2197, 1596, 1422, 2681, 399, 817, 2183, 1473, 2291, 2492, 2371, 1959, 2693, 1486, 558, 1040, 163, 148, 180, 2459, 1442, 631, 1211, 1499, 1929, 1265, 2017, 879, 1126, 1278, 1736, 2739, 1209, 1201, 409, 1725, 1598, 976, 1244, 1819, 2623, 2314, 2533, 839, 932, 1423, 1468, 1636, 1699, 2362, 2526, 2672, 426, 2032, 2449, 1757, 1765, 1822, 2472, 2707, 2717, 325, 303, 1288, 550, 1742, 37, 1267, 1453, 2220, 2554, 1768]. growth [1045, 2580, 2345, 1684, 10, 2516, 2203, 2737, 2210, 1084, 722, 2034, 2473, 190, 1630, 1497, 1796, 2511, 1823, 1675, 1433, 1685, 338, 342, 2121, 2316, 1687, 845, 1741, 501, 328, 2072, 524, 1590, 324, 1554, 517, 518, 1131, 2486, 268, 913, 1145, 1800, 1475, 767, 1254, 1886, 2151, 2041, 1893, 1369, 1773, 1983, 2213, 2175, 732, 2272, 1713, 1478, 1431, 452, 880, 1661, 1428, 759, 469, 464, 126, 2595, 1559, 1243, 1252, 2647, 2531, 1842, 644, 766, 884, 1597, 1372, 30, 2603, 18, 1965, 898, 1276, 379, 2285, 2393, 1200, 434, 331, 2216, 369, 2611, 289]. growth [822, 2667, 1830, 702, 2706, 2223, 742, 2391, 1010, 1978, 431, 1516, 2281, 2726, 2562, 968, 1653, 1273, 1824, 1719, 1338, 648, 1935, 1348, 1900, 2425, 854, 1374, 2098, 172, 2330, 2383, 641, 826, 1367, 2454, 2301, 2099, 1030, 1595, 2300, 2616, 1795, 1938, 611, 865, 1867, 2698, 1625, 2361, 724, 762, 814, 1707, 1739, 1617, 1857, 1820, 2495, 2382, 2652, 2556, 2742, 2745, 1543, 734, 1662, 2283, 997, 685, 1657, 1840, 1875, 2549, 2720, 1012, 1350, 1607, 363, 914, 1148, 1682, 2021, 2420, 2247, 940, 1035, 1425, 1517, 1799, 2536, 920, 89, 2152, 735, 782, 873, 2146, 1106, 892, 1727, 2219, 1038, 1191, 1398, 905]. growth [1280, 1405, 1373, 1540, 1722, 1592, 1976, 2115, 2245, 1021, 1195, 1149, 1561, 1649, 939, 831, 2113, 2130, 2318, 1988, 1444, 2718, 1731, 2569, 2753, 1317, 2711, 2091, 2201]. growth-promoting [1959]. growth-related [2638, 1468, 1824, 1707, 2146]. grypus [1484, 1879]. guajava [1998]. gueldenstaedtii [1848]. guide [863, 864]. Gulf [1275, 1336, 2530, 1695, 1113, 1482]. Gulfweed [1650]. gulio [2291]. Gut [2090, 1434, 2242, 2370, 2109, 1477, 1140, 1122, 2494, 813, 1774, 2169, 1701, 2111, 1436, 1126, 1420, 298, 2518, 1063, 1258, 2105, 1866, 1069, 2723, 2465, 2749, 2005]. gut-derived [2109]. gut-liver [2749]. guttatus [1353, 2697]. Gynogenesis [711, 902, 1247, 201, 1359, 2296, 1232, 581, 674, 748]. gynogenetic [1156]. Gynogenic [487, 586]. gynogens [902, 2212]. Gyrodinium [143].

H [354, 449, 1152, 274, 1472, 1504]. H-NMR [1504]. H. [1172, 2532, 2596, 753, 930, 2652]. habitat [1580, 809]. habits [1420, 2527, 2384]. haddock [430, 254]. haemato [1442, 1736, 2428, 2646]. haemato-biochemical [1736]. haemato-immune [2646]. haemato-immunological [1442, 2428]. haematocheila [720, 1425]. Haematococcus [2733, 2210, 2432, 2040]. haematocrit [1642, 136]. Haematological [535, 1972, 336, 1449, 1244, 2263, 2514, 1185]. haematology [1541, 898]. haematopterus [2634]. haematozoans [2263]. haemoglobins [183]. haemolytic [85]. haemotological [2003]. hairtail [1540]. haitanensis [895, 1913, 2187, 2727]. Haizhou [873]. hake [842, 1710]. Half [1480, 1320]. Half-sib [1480]. half-smooth [1320]. halibut [743, 838, 462, 768, 902, 976, 341, 181, 394, 137, 1042, 425]. Haliotis [1172, 2029, 1085, 1673, 1612, 370, 1701, 1115, 459, 1589, 2623, 2220, 2405, 1131, 896, 710, 1034, 1367, 514, 2465, 2652, 1875, 851, 1545]. halophyte [1462]. Ham. [908]. Hamilton [461, 953, 878, 1449, 1123, 1684, 2072, 2478, 921, 1258, 1667, 1191, 1398, 1455]. HAMP [1287]. HAMP-1 [1287]. Handbook [689]. Handling [228, 153, 1309, 49, 96, 759, 1527]. hanging [1067]. hannai [1701, 2220, 710, 2652, 851]. hansenii [2251]. hapas [2343]. Haptophyceae [784]. harbor [2525]. Harbour [81]. Hard [490, 1660]. Hardback [678, 1975]. hardened [2282]. hardness [1147, 382, 1928, 360]. Hardy [680]. Hargreaves [774]. harm [2619]. harvest [2266, 2714, 2535, 723, 1749, 1726]. harvested [723]. harvesting [1319, 2731, 845, 2243, 1858, 2728]. harveyi [2109, 1619, 1384, 2437, 1563, 915, 1806, 2448, 1548, 2717, 2329, 1872, 2049, 739, 1087, 950, 2480, 2039, 2307, 1260]. hassle [2401]. hassle-free [2401]. hasta [2572]. hatch [210]. hatchability [2142, 382]. hatched [1903, 1694, 1618]. hatcheries [2093, 730, 2690, 614, 1553, 737, 1332, 1987, 2004]. Hatchery [2458, 481, 1127, 199, 737, 718, 516, 2108, 1085, 548, 910, 909, 1001, 279, 155, 1622, 872, 954, 2614, 399, 631, 793, 1864, 805, 1737, 2429, 2101, 2316, 868, 91, 937, 2531, 2650, 111, 522, 526, 787, 2116, 397, 1955, 2283]. hatchery-bred [805]. hatchery-produced [1955, 2283]. hatchery-reared [548, 909, 399, 937, 787, 2116, 397]. hatchery-released [1737]. Hatching [1326, 38, 441, 1342, 340, 1472, 1624, 186, 2563, 1758, 1219]. hatchlings [350, 924]. Havbruk [708]. having [2516]. Hawaii [140]. Haworth [563]. HB [564]. hbk [552, 554]. hCG [2338, 1160]. head [2042, 763, 1805, 1665, 1981]. head-kidney [1981]. Health [1067, 2370, 2517, 945, 1319, 2427, 2670, 871, 2319, 2607, 1047, 365, 1436, 2492, 2693, 1765, 2707, 2750, 1073, 277, 2622, 586, 2697, 2647, 2028, 2667, 2119, 1185, 1595, 1543, 2578, 2571]. healthcare [1671]. healthy [892]. heart [2344]. Heat [1157, 1293, 2220, 2334, 1315, 1866, 1003, 2634, 1373]. heat-inactivated [1866]. heat-killed [2634]. heat-resistant [1373]. heat-shock [1293, 1003]. Heat-shock-induced [1157]. heated [799]. heavily [2225]. Heavy [2687, 2413, 2144, 2754]. Heckel [703]. Hediste [1870]. hedonic [1180]. Held [220, 2513, 2352]. Helfman [860]. helgolandica [1676]. Hellenic [1332]. hellerii [2014]. help [1400]. hemato [2297, 2299, 2140, 1890, 1360]. hemato-biochemical [2297]. hemato-immunological [2299, 2140, 1890, 1360]. hematobiochemical [2737]. hematologic [1869]. Hematological [2538, 2123, 2063, 2729, 2059, 1645, 1763, 629, 2455, 2399, 1764, 2348, 2024, 2709, 1661, 1428, 1863, 1496, 2224, 2160, 593]. hematology [1750, 1328, 2452, 2371, 1852, 1914, 1372, 1751, 2180, 1719, 1900, 2324, 1931, 1731]. hematopoietic [2685]. Hemicentrotus [1633]. Hemingway [554]. hemocyte [1810, 1268]. hemocytes [2114, 2689]. hemolymph [2434]. hemorrhagic [987, 1301, 1806]. Henderson [955]. Heniochus [1375]. heparin [1869, 2348]. hepatic [1293, 1882, 2619, 2742, 2646]. Hepato [2619].

Hepatobacter [2043]. hepatopancreas [2043, 1956, 2227, 2576, 2689, 2271, 1389, 2435, 2720, 2278, 2536]. hepatopancreatic [2605, 1956, 1510, 2218, 1441, 943, 931, 2010, 2102, 2163]. hepatopenaei [1904, 1909, 2546]. Hepatoprotective [1417]. hepatus [1375]. Herbal [836, 1810, 739, 1087, 2357, 2639]. herbivore [2470]. herbivorous [1307]. herbs [472, 2403, 2049, 1252, 735, 782]. Heritability [1903, 1341]. Herklots [2249]. Hermetia [2706]. herpesvirus [2280, 1925, 2401, 1482]. herpesvirus-2 [2280]. herring [211]. Hertlein [1118]. Hesperozygis [2419]. heteroclada [2623]. heterogametic [2212]. heterogeneity [313]. heterogenic [581]. heterologous [2678]. Heterophyes [2248]. Heterophyid [2159]. Heteropneustes [671, 1116, 1081, 817, 2263, 2391]. heterosis [2244]. heterostomum [2379]. heterotrophic [1052, 1970, 2183, 972]. heterozygosity [155]. heterozygous [189]. HIB D [2684]. Hibiscus [1991]. Hierarchical [811]. hif [2378]. HIF-1 [1816, 2378]. High [1986, 989, 1045, 1717, 1271, 2082, 1728, 788, 1870, 2701, 149, 1256, 2418, 2094, 2748, 1351, 1513, 1609, 1224, 1054, 2727, 37, 2127, 644, 547, 2000, 826, 967, 2477, 1330, 2143]. high- [2477]. High-density [1045, 967, 1330]. high-fat [2418]. High-level [2082]. high-performance [788]. high-quality [1870]. High-resolution [1717]. high-salinity [2000]. high-temperature [1256, 2727, 644]. high-throughput [2701]. higher [1911, 1973, 1199]. highly [1342, 934, 994]. hilarii [1098]. Hills [1027]. Himalayan [2644]. hindering [2198]. Hippocampus [1358, 2042, 544, 556]. Hippoglossus [743, 838, 976, 394, 1042, 425]. Hippolytidae [403]. histamine [1035]. Histological [1465, 233, 850, 1763, 2656, 2225, 1764, 2538, 1211, 983, 1233, 2496, 1050]. histology [2437, 2363, 2738, 2371, 2212, 2473, 2518, 2698, 2005, 1875]. histomorphological [2726]. histomorphology [1874, 1543]. Histopathological [2450, 400, 2324, 2063, 2248, 1353, 2009, 1940, 2463, 2159, 1073, 1065, 559, 2123, 2144, 2379]. histopathologicalstudies [1986]. Histopathology [2028, 2706, 2688]. historical [1574]. histories [1749, 1726]. history [723, 22, 1009]. HMB [594]. Hoffmann [2087]. Hokkaido [790]. holding [1066]. holes [2069]. Holmberg [1769]. Holothuria [713, 2310, 1193]. Holothuroidea [713]. Holt [991]. holy [2131]. Homarus [904, 1844, 1421]. homeostasis [1980]. homestead [1647]. homogenate [1213]. Homologous [2678]. homologs [2389, 2159]. homoserine [1780]. homozygosity [902]. homozygous [189]. Honduras [95]. honeycomb [695]. Hong [176]. Honshu [1103]. horizontal [1344, 1298]. Horizontally [706]. Hormonal [1679, 1160, 1229, 1934, 1508, 1661, 1159]. hormonal-sex-reversed [1508]. hormone [255, 2142, 2190, 2362, 875, 1296, 1338]. Hormones [220, 996, 1542, 770, 798, 136, 379, 1892, 1824]. Hormozgan [1380]. hornemannii [2683]. horneri [1650]. hornorum [1508]. horrens [2543, 2285, 2393]. horse [1144]. Host [2109, 1534, 1602, 1815]. hosts [2724]. hot [2153, 670]. household [1787, 1746]. housekeeping [2309]. Howell [307]. Hsp60 [2334]. Hsp60/10 [2334]. HSP70 [1860, 1299]. hsp90 [1299]. HT [1374]. Huangguan [1335]. Huanghai [1743]. Huangyou [1922]. Huangyou-1 [1922]. hub [2271]. Hucho [1739, 1617, 2495]. HUFA [430, 647, 67, 2496, 930]. HUFA-enriched [67]. Huffmanela [2687]. Hufmanela [2754]. human [2413, 1416, 1066, 1160, 1880]. humans [2019]. humoral [1309, 2572, 855]. humpback [866]. Hungarian [1220]. Hungary [658, 1575, 1207, 676]. hunting [918]. husbandry [1170, 1589, 2405]. Huso [1076, 1303, 2455, 2399, 1170, 1028, 1109, 2515, 1590, 1496]. Hybrid [574, 1269, 1750, 1635, 551, 2653, 2618, 2297, 660, 2748, 1188, 2111, 2371, 852, 2717, 1097, 269, 2482, 1631, 1732, 2151, 1751, 133, 1404, 2076, 2652, 1922, 2265, 1561]. Hybridization [483, 1908]. Hybrids [246, 1172, 290, 2244, 585, 1508, 1233, 1155, 1871, 2004]. hydration [1326, 1498]. Hydration/ [1498]. hydration/dehydration [1326]. hydraulic [1249, 2041]. hydrochloride [1921, 395]. hydrodromous [2387, 2632]. Hydrodynamic [2165]. hydrodynamics [5]. Hydrogen [1342, 16, 1259, 2282]. hydrogeochemical [1183]. hydrolases [2642]. hydrolysate [2354, 853, 2717, 1138, 1540]. hydrolysates [149, 2277]. hydrolysed [1440]. hydrolysis [344, 2558]. hydrolyzed [1824, 1316]. hydrophila [1033, 1638, 1790, 2502, 1835, 2336, 885, 2025, 2232, 1912, 1474, 987, 1999, 1137, 2587, 2341, 2034, 1388, 2675, 2482, 1998, 2200, 2151, 2261, 2058, 2037, 1333, 1805, 1010, 2333, 2470, 1437, 2402, 2749, 2247, 2500, 1922, 2061]. Hydroponic [627, 2267, 1849, 1677, 1462]. hydrostatic [2320]. hydroxide [1289, 1131]. hydroxy [594]. hyperosmotic [2181]. hyperoxia [779]. hyperoxic [294]. Hyphessobrycon [2363]. Hypnea [1867]. hypo [1299]. hypo-osmotic [1299]. Hypophthalmichthys [2535, 1792, 2152]. hypophthalmus [1477, 1972, 2277, 1818, 1781, 2492, 2554, 2514, 1150, 2659, 1661, 1653, 2330, 2301, 1394, 2698, 2646]. hyposaline [560]. hypoxia [2297, 931, 2703, 2271, 2168, 2368]. hypoxia-inducible [2368]. hypoxia-related [2168]. hypoxic [2378, 2564, 734]. Hyriopsis [733, 1734, 1043, 1080, 754, 777, 1329, 733, 2164, 1195].

I. [691, 863]. ibericus [1798]. ichthyofauna [2648]. icing [2011]. ICSR [812]. ictaluri [1922]. Ictalurus [1615, 1802, 1273, 1702, 1682, 1405, 1981]. ide [1213, 1228, 662, 1651]. idella [1980, 849, 603, 1187, 1459, 1999, 1040, 1699, 451, 1030, 2179, 2288]. idellus [1238, 1293, 2032, 2707, 2146, 1451]. Identification [2635, 2122, 2478, 1092, 1937, 1185, 2070, 2275, 1986, 2567, 2248, 2687, 2754, 2253, 1375, 2579, 1669, 738, 897, 1079, 1197, 2306, 1987, 965, 1053, 2628, 2089, 2307, 1494, 2004]. identify [528, 2332]. Identifying [2444]. idus [1213, 1228, 662, 1651]. IFN [1851]. IFN- [1851]. IGF [1556, 2099, 2559]. IGF-1 [1556, 2559]. IGF-2 [1556]. IGF-I [2099]. IGF-II [2099]. IgY [1387]. II [678, 255, 2429, 619, 737, 343, 2099]. III [193]. IL-17-3 [2642]. ilicifolium [2552, 2229]. ilicifolius [2680]. illucens [2706]. illuminated [1210].

Illustrated [297]. Image [502, 656, 1634, 2055, 1175]. images [2464]. imaging [2661]. imbricata [401]. Imelda [1284]. imidazole [2387]. imidazole-5-carboxylate [2387]. immersion [2284, 2404, 605, 2653, 2015, 645, 2572, 1704]. immersion-challenged [2653]. immobilized [209, 1493, 1302]. Immobilizing [488]. Immune [2280, 1901, 2691, 359, 1684, 1333, 2333, 2520, 2370, 1995, 1763, 1446, 1049, 1708, 1261, 2003, 1140, 1309, 1603, 1384, 2430, 2618, 2579, 1238, 989, 2114, 2297, 2532, 2025, 1764, 2437, 2232, 2434, 1680, 2538, 2148, 2161, 2736, 1474, 1538, 1640, 1351, 1467, 2024, 2738, 1501, 2231, 2328, 1126, 855, 1914, 2526, 2032, 1977, 2463, 2365, 2322, 1833, 2580, 2203, 2621, 2511, 1675, 1388, 2316, 2329, 1484, 1879, 2035, 2694, 2482, 2514, 2200, 2346, 1254, 1773, 2049, 1082, 1562, 1087, 2181, 1010, 2000, 1723, 1877, 2396, 648, 1185, 1059, 2470, 2723, 1867, 1966, 1105, 1438, 2179, 2435, 2495, 2749, 2005, 1403]. immune [1794, 1430, 1517, 1871, 2646, 1540, 1922, 870, 1021, 1343, 2685, 1500, 2130, 2318, 2686]. immune-modulating [1105]. immune-physiological [1723]. immune-related [1384, 2114, 2434, 2328, 1914, 2511, 1484, 1879, 2482, 2000, 1877, 2435, 2130]. immunity [1448, 2031, 2439, 2612, 1277, 2336, 1242, 2095, 2607, 1239, 1563, 2748, 2017, 1806, 2314, 2717, 2156, 2554, 2721, 2511, 1475, 2151, 2596, 739, 2645, 594, 1809, 2480, 1192, 1735, 2745, 2536, 2152, 2500, 2115]. immunization [1261, 1469, 1387]. immunocompetence [2560, 2689, 85]. immunodot [1321]. Immunogene [1805]. immunogenic [2192]. immunoglobulin [1321]. immunogold [1835]. immunohistochemical [1065]. Immunological [2144, 2729, 1972, 1645, 2248, 2687, 2754, 535, 2455, 2399, 2299, 2140, 1996, 1442, 1449, 2428, 1914, 2672, 2034, 1687, 931, 1890, 2556, 1360, 2219]. immunology [1852, 898]. immunomodulator [1851]. Immunomodulatory [2609, 1169, 1184, 2612, 2047, 1711]. immunophysiological [1822]. immunoregulation [2556]. Immunoserology [1868]. immunostimulant [2251, 1998, 1773]. immunostimulants [1810, 2177, 1294]. Immunostimulating [739]. immunostimulation [1887, 941]. Immunostimulatory [1511, 1610, 1704]. immunosuppressive [2593]. IMNV [1397]. IMP3 [2046]. Impact [144, 1549, 579, 1337, 845, 2615, 1983, 1027, 1842, 1272, 2616, 2011, 1424, 1988, 815, 1748, 541, 804, 1575, 2351, 723, 1878, 2027, 1967, 655, 20, 884, 2507, 971, 2724, 580, 1394, 542, 1167, 607, 543, 1761, 718]. impacted [1485]. Impacts [2729, 457, 2599, 329, 1630, 2630, 289, 2655, 1151, 1275, 2141, 2119, 714, 1746, 2745]. impairs [1980]. imparipes [2170]. Implant [1053]. implants [368, 1747]. Implementation [1514, 1899, 1950, 2584, 2117]. implementing [2350]. implication [1099]. Implications [1129, 227, 1679, 1305, 2501, 791, 1120, 2446, 1487, 1756, 1755, 2226, 1002, 2574, 2345, 2721, 280, 1325, 1113, 1597, 612, 1051, 2174]. Implicit [1057]. import [2086, 1582]. Importance [1845, 667, 795, 198]. important [856, 927, 926, 1803, 1937, 2528, 1616, 2182]. imposed [1498]. Improve [507, 1245, 1789, 2298, 1256, 1993, 757, 1769, 2314, 2429, 2611, 2281, 2264, 121]. Improved [1358, 1995, 2560, 1997, 1122, 1257, 1442, 2649, 1765, 2516, 2141, 2350, 2703, 2479, 1858, 1443, 1657, 1840]. Improvement [2336, 2304, 1258, 2458, 132, 2535, 1080, 48, 2359, 1348, 893, 1327]. improvements [2350]. improves [1812, 2109, 1627, 2656, 2593, 2515, 1135, 2631, 2218, 2486, 2622, 2195, 2223, 2588, 2076, 1192, 2639]. Improving [2158, 677, 1274, 1601, 1718, 2143]. IMTA [2505, 1834]. in-pond [660]. in-vitro [733]. inactivated [2171, 1866, 2582]. inaquosorum [2105]. inbreeding [1419, 1457, 1742]. inchi [2371]. Incidence [306, 1312, 2272]. Incisor [304]. including [2708]. Inclusion [2363, 1018, 1792, 1535, 2706, 2100, 1496, 1824, 2569, 2753]. inclusions [1795]. income [751]. incomes [1746]. incorporated [2486]. incorporating [2171, 1658, 1783]. Incorporation [579]. increase [2277, 1000, 193, 2611]. Increased [2207, 1000, 63, 1199, 1122, 1326, 2516]. increases [2148, 2169, 2544, 2608]. increasing [882, 270, 67, 822]. Incubation [337, 1874, 351, 1927, 186, 14, 1876]. independency [1930]. independent [1069]. Index [170, 215, 216, 315, 316, 375, 376, 419, 420, 475, 476, 1489, 504, 262, 263, 439, 2055, 702, 968, 1366, 1427]. indexes [2318]. India [713, 523, 1904, 2274, 2555, 2086, 1724, 2624, 2052, 1385, 2417, 2006, 2272, 2194, 1183, 1130, 1112, 2349, 2746, 2445, 2444, 1284]. Indian [1906, 2567, 461, 671, 953, 1116, 255, 2207, 2660, 2599, 2532, 1242, 2676, 1684, 2516, 2644, 1796, 2072, 1605, 982, 2551, 2596, 2350, 2478, 1087, 1667, 2189, 969, 1955, 1361, 1437]. indica [2552, 2331, 1467, 2273]. indicator [1107]. Indicators [448, 234, 2036, 535, 2140, 1351, 1940, 2349, 1273, 539, 85, 2435, 543]. indices [2370, 2455, 2399, 1793, 2086, 2024, 50, 2291, 2400, 1863, 1653, 1050, 1890, 920, 1727, 1398]. indicus [2532, 1467, 1796, 982, 2551, 2596, 1087, 2058, 2037, 1185, 1059, 950, 1361, 1182]. indigenous [1305, 807, 2589, 2178, 2444]. indirect [1480]. Individual [460, 2107, 1935, 1799, 4, 2253, 2226, 2425]. Indonesia [866, 2136, 2650, 1489]. Indonesian [2051]. indoor [1567, 897, 1881, 1159]. indoor-reared [1159]. induce [649, 2637]. Induced [2016, 201, 581, 496, 2699, 2142, 2599, 2418, 1157, 1417, 1247, 1806, 324, 1359, 2491, 1250, 2278, 2658]. induces [1059]. inducible [2368]. inducing [972]. Induction [195, 1156, 525, 2296, 185, 748, 480, 923, 1160, 2735, 2320, 1300, 674]. inductions [1086]. industrial [2400, 1386, 1683]. industrialized [74]. Industry [229, 1295, 1557, 836, 2398, 23, 2602, 1031, 2650, 2730, 32, 2692]. inedible [2461]. inert [437, 900]. infected [1980, 2043, 2336, 2225, 2280, 885, 2232, 1993, 1912, 2110, 1664, 1855, 1852, 1916, 2159, 1065, 1184, 2123, 2058, 2037, 1331, 2685]. infecting [2448]. infection [1033, 2031, 2284, 2687, 2754, 730, 1730, 2618, 2579, 2095, 2023, 2025, 2732, 2437, 2538, 1676, 2206, 2229, 1137, 1833, 2621, 2675, 2694, 1695, 1732, 2049, 1797, 2592, 739, 1562, 559, 1087, 2353, 343, 2028, 1987, 395, 1333, 1482, 1010, 1059, 2470, 1962, 2480, 936, 1949, 1105, 1744, 2324, 1403, 2288, 2039, 2152, 1727, 2639, 1397]. Infections [252, 2071, 2751, 2347, 1969, 2577, 2539, 2144]. infectious [1993, 1065, 2070, 2685, 1397]. infective [2216]. Infectivity [26]. inference [1965]. Inferred [481, 84]. infestation [2093, 81, 560, 2645, 2528]. infested [2530, 287]. infiltration [2181]. inflammatory [2656]. inflation [1932, 1084, 1208, 1928, 1190]. Influence [1843, 2699, 1972, 232, 2502, 2286, 490, 1804, 903, 1281, 1018, 1251, 2610, 1770, 38, 198, 472, 512, 2128, 780, 390, 2367, 281, 180, 1054, 1920, 361, 1675, 1825, 101, 90, 1252, 1597, 18, 1535, 547, 2667, 668, 969, 669, 208, 219, 2219, 999, 1178, 2059, 721, 1076, 1098, 1520, 2190, 2147, 2294, 603, 1498, 1703, 205, 404, 2624, 841, 171, 696, 1231, 2210, 177, 2079, 739, 1087, 840, 594, 1824, 1979, 557, 2160, 468, 1361, 1425, 1373, 1561]. influenced [463]. influences [1477, 1750, 62, 2335, 50, 821, 1046, 2263, 1369, 1751, 2130]. influencing [617, 1693]. Information [508, 54, 1743, 790]. infreshwater [273]. ingestion [2298, 167, 1312, 1455]. ingredient [2013, 2726, 990, 2204]. ingredients [1813, 2202, 2719, 1846, 2611, 2061]. inhabiting [1154]. inheritance [1679, 1550]. Inhibition [410, 2664, 2053, 2261]. inhibitors [1262, 301, 1224]. Inhibitory [987, 614, 1182, 693, 2641, 2408, 2261, 1592]. inhibits [1715, 1105]. iniae [2618, 2111, 2329, 2482, 1969, 1562]. Initial [493, 221, 281, 841, 884, 814, 2453]. initiative [1973]. initiatives [2705]. Injection [249, 1180, 1328, 1296, 395]. Injections [496]. injury [2418, 1417]. Inland [809, 808, 549, 623, 2631, 879, 2739, 740, 714]. innate [1448, 1261, 1242, 2095, 1563, 1640, 1806, 2314, 1977, 2034, 2721, 2076, 1867, 2480, 2686]. inner [1329, 1938]. inner-shell [1329]. Innovation [1415, 812]. innovations [2001, 2259]. innovative [1963]. Ino [1701]. inocula [2103]. Inoculated [567]. inoculation [1640]. Inolens [1920]. inopinatus [2028]. Inorganic [1822, 2205, 2541]. inositol [1857]. input [2286, 1151]. Insect [2235, 1571]. insects [128]. Insemination [587, 1817]. insensibilis [2172]. inserting [1620]. Insight [2012, 1080, 1274, 2600]. insights [1865, 2305, 646, 2688, 1951]. instar [426]. Instructions [225, 317]. instrument [1581]. Int [574]. Intake [251, 996, 1226, 1912, 813, 919, 404, 2738, 701, 163, 148, 180, 845, 1106, 2323]. Integrated [2227, 567, 2624, 1760, 2166, 2295, 1053, 413, 1462, 1245, 819, 1334, 706, 1529, 1899, 2509, 1709, 1176, 1102, 1391, 1606, 2381, 520, 521, 2411, 1593, 2570, 15, 2292, 1898, 1834, 292]. integration [1134, 2366]. integrin [2276]. Intelligent [2326]. intended [1016]. intensification [911]. intensify [1127]. intensities [1624]. intensity [2699, 2294, 254, 381, 2436, 1014, 932, 1918, 1050]. Intensive [506, 493, 494, 577, 93, 1290, 470, 2610, 1248, 1253, 795, 2519, 40, 2731, 2094, 1168, 833, 1198, 66, 1771, 1580, 2533, 1460, 1585, 333, 352, 2498, 1143, 1902, 339, 1034, 2695, 2243, 2507, 763, 952, 834, 2289, 1666, 2454, 1225, 1689]. intensively [9, 2734, 431]. inter [777, 1948, 1947]. inter-populational [1948, 1947]. inter-simple [777]. Interaction [1699, 1459, 919, 1661, 2600, 2145]. interactions [1734, 2714, 1304, 1572, 11, 1941, 2057, 637]. Interactive [1033, 2739, 1264]. interests [2260]. interface [1279, 1424]. interference [2665]. interferon [1851]. interleukin [2642]. intermediate [2389]. intermedius [1419, 1864, 1274, 1559, 1280, 1650, 1149, 1341]. intermittent [1149]. internal [1211, 715]. International [217, 264, 1518, 926, 803, 1409, 2240, 2086, 2077]. interrupts [2207]. Interspecific [1908, 1913]. intertidal [772, 1569, 2015, 1174]. intervention [1608]. interventions [2615]. Intestinal [2225, 1895, 2066, 2063, 1750, 1708, 549, 2336, 2313, 2319, 1992, 134, 2363, 2607, 2733, 1766, 2743, 1538, 2354, 2738, 1769, 2371, 1486, 2032, 2449, 2472, 2707, 2322, 2580, 2518, 2151, 2622, 2127, 1372, 1535, 1751, 2412, 2119, 2330, 2520, 1890, 2076, 2742, 2745, 2005, 2634, 2247, 1517, 2238, 2646, 2219, 1604, 2245, 746, 2130]. intestine [2147, 1465, 1230, 2473, 1879, 2496, 1543, 2375]. Intramuscular [249]. intraperitoneal [1833, 395, 1387]. Intraspecific [884, 1290]. introduced [636, 271, 2030]. introducing [292]. Introduction [1519, 1895, 861, 1008]. introductions [310, 277]. introductory [862]. Introgression [481]. intrusion [1394]. inulin [1465, 2708, 601, 2745]. Invasive [230, 1303, 2253, 1305, 970, 776, 393, 1892, 2661]. invertebrate [1002, 1786]. investigate [2612, 890, 1793]. investigating [2185, 2655]. Investigation [659, 1572, 1916, 234, 2013, 746, 813, 2522, 2356, 1608, 794, 1454, 2194, 2640, 2163, 2692]. Investigations [670, 1796]. Investing [1416]. investment [786]. involved [1140, 1239, 306]. Involvement [2373, 945, 1315]. involving [2397]. iodide [1733]. iodine [1342, 2577, 2539, 2500]. Ionic [485, 454, 879, 2508, 92]. IPNGS16 [2148]. IPNV [33]. Ipomoea [2740, 1495, 2479]. Iran [2173, 1380, 2478, 2369, 1502]. Iranian [994]. Ireland [81, 871, 2154, 1115, 1985]. iridovirus [1363, 2070]. iron [2428, 193, 898]. irradiance [1471]. irradians [278, 397, 967, 278, 397, 967]. irradiated [1156, 581]. Irradiation [219, 218, 1861, 1629, 1247, 674]. irrigation [55, 326]. irritans [2691]. ISABMC [926]. ISABMC-2009 [926]. ISBN [564, 552, 565, 554, 563, 678, 1008, 1975, 553]. ISBN-10 [678]. ISKNV [2070]. Island [2002, 356]. islands [2657, 2134]. ISO [164, 1450]. Isochrysis [2653, 1841, 1609, 1450, 2383, 164]. isoelectric [168]. isoeugenol [1314]. isoflavones [1875]. isolate [2105]. isolated [1621, 693, 1619, 385, 635, 2046, 2014, 2342, 1829, 885, 2676, 2471, 2680, 1323, 259, 2386, 140, 1879, 2675, 2682, 1258, 1866, 2412, 1390, 971, 2383, 950, 1182, 2477, 2546, 892, 2375]. isolates [1085, 2380, 2202]. Isolation [2567, 2060, 897, 1970, 1897, 2448, 1429, 936, 2628, 2089, 2307, 738, 583, 2139]. isoleucine [1324, 1191]. isonitrogenous [550]. isothermal [2149]. isotopes [1570, 542, 2718]. isozymes [174]. Israel [320, 333]. Israeli [75]. Israelitrial [269]. ISSR [777, 1964]. issue [1519]. issues [529, 414, 365]. Italian [1571]. Italy [815, 2050, 1793, 1336, 951, 543, 538, 1482]. items [358]. ITG [2276]. ITS-1 [1154]. ITS-5.8S [895]. ivory [2477]. Iwagaki [1820].

J [574, 565, 554, 690, 700, 727, 709, 774, 861, 991, 1039, 1072, 446]. J.

[690, 861, 1071, 1152, 1837, 1905, 1975]. Jacobsen [690]. jade [2461, 1466, 1744]. Jahncke [365]. Jana [563]. Japan [126, 1103, 927, 2317, 790, 1987, 1330]. Japanese [109, 996, 2607, 1351, 2335, 1224, 1506, 758, 1507, 126, 790, 914, 920, 851]. japonica [1035, 2687, 2754, 1049, 1958, 2717, 758, 1507, 1296]. japonicus [788, 1708, 1256, 2607, 1239, 2335, 1102, 1391, 1606, 839, 1423, 1546, 1270, 2491, 759, 1547, 1243, 2119, 1003, 1192, 1299, 1146, 1483, 1607, 914, 1517, 1318, 1373, 1494, 870, 1343, 2662]. Jasus [434]. Java [2650]. javelin [2572]. Jayachandran [374]. Jellyfish [2636, 1345, 686, 2228]. Jersey [356]. jet [1249, 2704]. Jewel [2363]. Jian [1595]. Jiangsu [1909]. John [725]. Joseph [1284]. journey [878]. juice [2543]. Juvenile [251, 515, 573, 234, 233, 1404, 514, 1812, 519, 2729, 1196, 1477, 562, 731, 1076, 77, 1635, 1049, 146, 1708, 1037, 904, 1839, 1121, 642, 749, 460, 1716, 1309, 684, 796, 548, 827, 909, 979, 1001, 2277, 2293, 1017, 1353, 2430, 2090, 2298, 867, 2455, 2399, 438, 773, 603, 780, 2419, 113, 1226, 2468, 1259, 1110, 1779, 2348, 1340, 402, 833, 1701, 2545, 2748, 1999, 1109, 1769, 2416, 1023, 944, 824, 1552, 1776, 2335, 2564, 459, 1958, 1135, 2543, 994, 677, 768, 1781, 1364, 1654, 115, 114, 29, 2197, 1422, 877, 1048, 1137, 2355, 2492]. juvenile [1046, 1040, 1651, 163, 148, 2459, 1442, 631, 756, 1211, 1882, 1265, 1736, 976, 2533, 2205, 2719, 932, 1699, 2426, 1324, 1765, 881, 1267, 1024, 334, 578, 917, 2473, 190, 1429, 2153, 1368, 1297, 2035, 356, 1590, 1732, 1475, 1143, 1886, 2151, 1369, 437, 732, 1713, 1846, 1884, 452, 759, 2127, 1454, 739, 931, 2697, 1091, 1092, 2066, 1111, 369, 2611, 2020, 822, 1138, 702, 2706, 742, 1271, 127, 2562, 968, 143, 1824, 959, 1719, 1900, 1374, 2098, 172, 2330, 453, 1595, 1225, 611, 2361, 1003, 1537, 1488, 1739, 2495, 589, 612, 997, 685, 1012, 1483, 1607, 914, 1148]. juvenile [1665, 1425, 1799, 920, 1322, 892, 443, 592, 1931, 2204, 1650, 1700, 2265, 870, 1561, 1649, 745, 2130, 2318, 2718, 2569, 2753, 1317, 2201, 2568]. juveniles [2560, 1972, 1134, 1652, 946, 707, 2297, 2607, 949, 907, 2045, 661, 1169, 2708, 590, 199, 2449, 2463, 550, 1885, 606, 1687, 767, 591, 694, 2709, 2166, 1165, 1579, 79, 2595, 1535, 2285, 2393, 1830, 1677, 1653, 1496, 1539, 24, 1042, 1694, 1955, 865, 2465, 741, 2435, 761, 905, 1545].

K. [374]. kappa [1867]. Kappaphycus [2619, 893]. Karaman [1879]. kasignete [1842]. Kaup [1091, 1089]. kDa [2389]. Keap1 [2491, 2732]. keep [405]. keeping [1166]. Kellogg [1795]. kelp [2410, 1611, 1103]. Kenya [1963, 977, 1984, 835, 1951]. Kenyi [1821, 1596]. kept [62, 509, 2018]. keratin [2218]. kernel [1467]. kessleri [2728]. ketoglutarate [2032]. ketotesterone [2212]. key [2269, 2599, 2332, 2534, 2584, 1373]. Keyword [216, 263, 316, 376, 420, 476]. kidney [2619, 1805, 2070, 1981]. Killary [81]. killed [2572, 2634]. killing [794]. kinase [2665]. kinds [439]. kinetics [1045]. King [716, 282]. kingfish [1244, 794, 1312]. Kishinouye [686]. kisutch [2303, 2420]. Kjørsvik [564]. knockout [1962]. knowledge [2705]. kob [1270]. koi [633, 1495, 2041, 2033, 2357, 2115, 1383, 2345, 1495, 2041]. Komárek [2087]. Kong [176]. Konjac [1729, 1604]. Korea [2583, 2317].

Korean [1297]. Kraft [879]. KT29 [2480]. kuda [2042]. Kumar [1205, 1284]. Kurokura [488]. kuruma [1256, 2491]. Kutty [692, 1039]. kutum [1399, 2747, 964, 1399, 964]. Kützing [1334, 847].

L [2267, 2450, 365, 449, 447, 554, 861, 1070, 1039, 1975, 1138, 671, 1266, 1499, 820, 1374, 2748]. l-carnitine [820, 2748]. l-threonine [671]. l-tryptophan [1266, 1499]. l-tryptophan-supplemented [1374]. L. [1178, 1033, 1448, 519, 1638, 457, 489, 656, 1020, 2029, 901, 570, 799, 743, 500, 904, 1166, 882, 1251, 1067, 1220, 588, 1140, 942, 1215, 1216, 684, 687, 796, 856, 1869, 1213, 989, 970, 2267, 2673, 350, 396, 2023, 1715, 1622, 600, 487, 1013, 1226, 1179, 323, 1210, 595, 828, 938, 1565, 2680, 456, 1328, 872, 1340, 838, 662, 1584, 1383, 1944, 785, 1643, 1589, 1230, 1162, 590, 455, 1486, 558, 918, 1217, 2459, 1160, 1211, 1499, 1929, 1171, 1753, 1209, 1054, 676, 976, 1032, 853, 587, 584, 1908, 1934, 1576, 361, 359, 1833, 596]. L. [585, 578, 1583, 2121, 1484, 356, 517, 518, 770, 798, 913, 1143, 2041, 1233, 591, 1208, 1928, 2196, 1332, 2272, 1431, 1434, 452, 586, 1943, 1234, 1165, 1579, 1939, 464, 1562, 442, 2650, 488, 395, 620, 1221, 1222, 1805, 594, 597, 641, 1367, 1089, 1108, 1225, 2056, 2616, 669, 504, 1437, 581, 593, 589, 443, 480, 592, 1159, 1931, 999, 1161, 1223, 2245, 1942, 2456, 744, 2112]. labeling [1043]. Labeo [953, 1269, 2207, 1321, 2513, 1851, 1262, 878, 1100, 908, 1474, 2075, 2428, 1123, 1684, 2072, 1258, 1667, 2611, 2562, 1338, 2470, 2600]. Laboratory [813, 517, 518, 1178, 2241, 549, 2287, 1011, 934, 623, 754, 879, 536, 1208, 369, 2339, 1799]. laboratory-raised [1011]. laboratory-reared [934]. labrax [2031, 519, 901, 535, 340, 970, 1780, 2673, 1766, 1774, 912, 1308, 1193, 578, 606, 367, 1428, 2157, 1092, 2020, 1868, 1132, 2056, 1543, 2005]. Lac. [1171]. Lack [867]. lactic [1621, 2335, 1263, 2526, 1684, 2177, 2412]. lactis [1560, 1132, 1560, 1132]. Lactobacillus [2722, 2109, 1708, 1652, 2169, 2515, 1484, 1879, 2177, 1428, 1723, 1705, 2115]. lactococcosis [1323]. Lactococcus [1323, 1852, 1560, 2123, 1132]. lactoferrin [1603, 1105]. lactonase [1780]. lactones [1780]. Lactuca [2267, 1849, 1248, 1913, 2505]. laevigata [514]. Laevistrombus [2485]. Lagoon [2050, 951, 523, 541, 144, 1064, 1243, 1804]. lagoons [534]. lagowskii [2324, 2368, 2375]. Lake [479, 573, 1154, 769, 2194, 2523, 1210, 1850]. lakes [777, 596, 597]. lalandi [1244, 794, 1312]. Lam. [2523]. Lamarck [1129, 728, 605, 636, 2230, 402, 348, 1602, 1553, 800]. lamellar [2446]. Lamellidens [857]. Lamiaceae [2460]. Laminaria [1610]. LAMP [1441, 2149]. Lampetra [1753]. lamprey [1753]. lanceolatu [2733, 2717]. lanceolatus [1635, 2748, 1732, 1404]. Land [1380, 1196, 795, 390, 1793, 31, 1110, 682, 1395, 2505, 380, 538, 2584, 1979]. land-based [1196, 795, 390, 1793, 31, 1110, 1395, 2505, 380, 538, 2584, 1979]. land-suitability [682]. Landlinks [565]. landlocked [2544, 91]. landscape [1203]. Language [1413]. Lanka [147]. lanthanid [437]. Lantibiotics [2466]. Large [1864, 482, 1673, 890, 1771, 2691, 756, 2417, 547, 1339, 1949, 674, 711, 1512]. large- [1512]. Large-scale [1864, 1673, 1771, 756, 2417, 711]. largemouth [2054, 1171, 2580, 1106, 1975]. larger [300]. largest [777]. Larimichthys [2691, 1908, 1949]. larva [1534, 1429]. Larvae [232, 570, 235, 221, 622, 494, 516, 572, 1178, 68, 615, 577, 285, 2654, 1098, 1673, 1932, 561, 1418, 149, 687, 42, 2467, 472, 1122, 934, 1377, 2532, 545, 546, 1186, 804, 2170, 1874, 1210, 813, 59, 872, 1766, 2230, 1421, 1903, 1901, 122, 463, 1944, 2690, 2485, 1162, 2218, 1094, 958, 167, 1499, 1929, 1753, 1614, 2282, 2598, 1934, 1940, 325, 1312, 613, 341, 2429, 1435, 2101, 1630, 1097, 1823, 78, 602, 619, 428, 102, 501, 1270, 332, 1554, 165, 770, 798, 913, 177, 389, 1507, 1777, 850, 181, 964, 1208, 1928, 337, 339]. larvae [2622, 90, 355, 427, 464, 2695, 2531, 18, 921, 210, 434, 111, 2496, 1007, 1830, 112, 431, 2588, 121, 1877, 2073, 1389, 1570, 641, 204, 1132, 741, 814, 1617, 1820, 2283, 1618, 1350, 89, 1527, 999, 1698, 1759, 2711, 2406, 2453, 2112]. Larval [2241, 2501, 1958, 1046, 226, 505, 626, 1129, 77, 2207, 424, 1770, 430, 441, 867, 160, 1612, 396, 254, 828, 938, 2126, 1487, 698, 58, 2459, 839, 1927, 771, 1084, 2524, 382, 501, 1952, 373, 1431, 452, 1287, 1930, 403, 2191, 1646, 580, 135, 1876, 1938, 724, 432, 208, 1190, 1146, 2021, 2431, 1807, 1455, 991]. larviculture [1971, 1375, 1481, 2503, 67, 289]. larvicultures [577]. laser [1043]. lasting [669]. Late [566, 150]. Lateolabrax [2335, 1102, 2441, 2440, 914]. lateral [2401]. Lates [1645, 2442, 2343, 2329, 1691, 1881, 2674, 997]. Latin [2364]. latipes [1224]. Latreille [1839]. Latris [618, 778, 440]. latus [2180, 685]. launcher [2157]. laurenti [283]. lavaretus [180, 1154]. Lawrence [554]. layer [2247]. LC [1910]. LCDV [2136, 2478]. LDPE [1601]. Lea [857]. leachate [1795]. lead [2722]. leaf [2025, 2680, 2386, 2516, 83]. Learning [1413, 1415, 1410, 1414, 2117]. Leaves [622, 2009, 2153]. lecithin [2088, 1133, 944, 753]. lectin [2389, 2232, 2275]. LED [2175]. LEDs [1819]. Lee [700, 709]. left [748]. left-eyed [748]. leg [2631, 1614, 2739, 2295, 2102, 2480]. legal [2239, 607]. legislation [2193]. legume [178]. Lekang [691]. lemaneiformis [1176, 990, 1503]. Lemmermann [2087]. Lemna [2298, 1396, 2292, 2616]. lemon [2299]. Length [1385, 878, 2107, 90, 2288]. leniusculus [638, 833, 752]. lens [2239, 1347, 2503]. lentillifera [1001]. Lentinan [1192]. Lentinus [1846, 1192]. Lepeophtheirus [1655, 2724]. Lepomis [874]. Leporinus [1623]. leptodactylus [304, 273, 406, 723, 752, 876, 1117, 1266, 917]. Lereboullet [345, 369]. less [1040]. less-expensive [1040]. lessons [747, 2615]. lethal [1553, 559, 1910, 1868]. lettuce [2267, 2468, 2166, 2167]. lettuce-juvenile [2468]. leucaena [83]. leucine [1398]. Leuciscus [1213, 1228, 662, 1651]. leucocephala [1437]. Leuconostoc [1866]. leukocyte [2160]. leukocytes [2280]. Leung [679, 709]. levamisole [2409]. Level [482, 1121, 1409, 2559, 1784, 1340, 1345, 1364, 817, 1852, 180, 1224, 2314, 2707, 550, 96, 2079, 884, 379, 1091, 742, 1107, 44, 1059, 1666, 1035, 2082, 1500, 2113].

Levels [248, 506, 2729, 1020, 1133, 1321, 882, 1463, 942, 149, 270, 161, 909, 1017, 996, 899, 780, 1081, 1226, 291, 1381, 802, 2738, 1552, 623, 2452, 2492, 1442, 1828, 1542, 1765, 359, 1267, 1453, 606, 1063, 2541, 1590, 165, 982, 1661, 1234, 30, 1200, 922, 1138, 1339, 1107, 1496, 959, 1374, 1030, 436, 1189, 2324, 2477, 2652, 1875, 1350, 363, 914, 1148]. LFD [2401]. LHRH [588, 875, 496]. LHRH-A [875, 496]. LHRHa [195]. lice [16, 81]. licheniformis [1182, 2319, 2148, 1126, 2675, 2694, 1063, 2245]. Life [515, 447, 2404, 2651, 350, 391, 458, 2351, 351, 1920, 1576, 2210, 1208, 1639, 394, 407, 580, 2056, 861, 957]. lifelong [1410, 2117]. ligand [2137]. ligands [2587]. ligase [2442]. light [1971, 890, 2294, 2581, 254, 595, 381, 838, 1472, 1624, 1632, 1014, 1819, 932, 2362, 2175, 1918, 1092, 1236, 1225, 164, 1613]. light-emitting [1819]. light-substrate [595]. lightweight [2519, 2703]. lignite [1860]. Ligurian [386]. ligustica [847]. like [2741, 381, 1701, 140, 2097, 2275, 2039]. Lim [353, 449, 650]. limacinum [2388]. limbatu [1885]. lime [258]. liming [491]. Limit [251, 1498]. limitation [2643]. limitations [2067]. limiting [1637]. limits [960]. Limnoperna [1305]. Limnoscapha [754]. limon [2554]. line [2280]. Lineage [621]. lineata [618, 778, 440]. lineatus [2481]. lines [1220, 57]. Linguistic [1413]. link [548, 979, 631]. linkage [675, 874]. linkage-group [675]. linked [1896, 1107]. Linn. [549, 1384, 846, 2075]. Linnaeus [1037, 1730, 1673, 1248, 340, 1612, 391, 458, 568, 832, 2133, 370, 1854, 1792, 948, 2521, 601, 2524, 737, 718, 1597, 468, 628, 508, 1618, 1545]. Linne [471, 771]. linoleic [899]. linolenic [899]. linseed [1163]. lip [396, 1032, 1952]. lipase [1189, 2082]. Lipid [2046, 907, 471, 501, 250, 202, 1016, 425, 1020, 492, 1286, 1690, 909, 1017, 2147, 2581, 2085, 1626, 780, 1226, 203, 196, 2748, 2738, 1552, 1609, 1629, 1364, 956, 975, 1897, 2291, 2282, 1699, 550, 1267, 1590, 1450, 165, 982, 373, 1713, 1478, 17, 312, 18, 1296, 1873, 1163, 1505, 431, 1653, 1858, 668, 1025, 1189, 2742, 789, 1618, 2074, 914, 1148, 2087, 1425, 1106, 1729, 1561, 2113, 1545]. lipid-enriched [1025]. lipid-related [1163]. lipid-rich [1897]. lipidcomposition [270]. Lipidomic [1665]. lipids [2078, 1490, 2400, 17, 1271, 2091]. lipoic [2281]. Lipolytic [463]. lipopeptide [1592]. lipoprotein [1189]. liposomes [433]. lipped [615, 2513, 1281, 2023, 1480]. Lippia [2656]. liquid [2681, 2481]. liquor [908]. List [217, 264]. literature [1533]. lithium [1869]. Lithodes [1539]. Litopenaeus [1119, 1011, 522, 526, 2699, 2552, 2604, 1619, 1904, 1652, 1248, 1334, 2043, 764, 1463, 1817, 1384, 1678, 2427, 2437, 1173, 1567, 2094, 2725, 2434, 2319, 1396, 1992, 1393, 1289, 1259, 1381, 1379, 2251, 1860, 1555, 2148, 1676, 1784, 1052, 883, 1538, 2684, 1023, 944, 824, 1198, 2602, 2631, 1827, 1654, 1664, 1855, 2218, 2355, 1614, 1598, 2314, 2533, 2205, 2719, 894, 1268, 525, 1324, 1822, 2092, 2472, 1460, 2322, 2608, 2316, 2518, 2694, 524, 1886, 389, 2079, 1019, 2006, 2049, 1880, 1620, 1502, 331, 2243, 763, 2295, 1452, 2000, 2594, 2100, 2102, 2188, 1036, 1877, 2289, 1389]. Litopenaeus [1570, 1374, 1711, 1962, 2018, 1331, 2723, 1966, 2480, 2397, 2432, 2655, 2546, 2549, 2039, 735, 782, 2089, 1613, 1722, 1976, 1021, 1424, 2318, 1397, 1355]. Little [354]. littoral [596]. Liu [725]. Live [511, 628, 1033, 642, 2467, 867, 438, 458, 396, 2170, 435, 662, 1347, 1399, 1241, 1771, 2202, 1632, 696, 1614, 1560, 2580, 1583, 1982, 1872, 1554, 964, 1842, 18, 1007, 1866, 2582, 1355]. livebearing [729]. livelihood [2052]. livelihoods [2421, 2624, 417]. Liver [2579, 69, 2147, 2418, 1841, 1417, 821, 2371, 361, 359, 2473, 983, 473, 1287, 1296, 2496, 1163, 61, 2281, 2301, 2698, 2749, 1029, 1543, 1561]. lividus [1859, 1610, 923, 2230, 2055, 800, 1767, 1365]. living [1008, 104, 921]. Liza [2023, 720, 1425]. Lmk. [1838, 558]. loach [1157, 1156, 324, 724, 741, 814, 711, 831]. load [472, 1766, 2282, 289]. loaded [2299]. Loading [538, 2041, 2462]. loads [1861]. Lobatus [1892]. Lobelia [2601]. lobster [904, 1421, 1458, 426, 434, 2706, 82, 272]. lobsters [1844]. Local [1386, 910, 2359, 1203]. Localisation [2111]. locally [1740]. located [2555]. location [1079]. locations [1915, 2529]. loci [1670, 675, 583, 852, 1848, 1201]. lockdown [2599]. locomotor [2073]. logarithmic [1609]. logic [834, 973]. lombardoi [1821, 1596]. London [1008]. Long [218, 586, 1559, 968, 1795, 2651, 2291, 2503, 17, 669, 1021]. long-chain [2651]. long-lasting [669]. long-spined [2503]. Long-term [586, 1559, 968, 1795, 17, 1021]. long-whisker [2291]. longitudinal [2086, 1073, 2349]. longline [2062, 984, 2264]. longlines [637]. look [2381]. loop [2158, 2149]. loop-mediated [2149]. Lophiosilurus [1884, 2568]. Loss [552, 2266, 1271]. losses [1569, 632, 2645, 211]. Lota [1210, 1209, 1208, 1225]. Low [942, 1253, 2187, 240, 1435, 2480, 2699, 2421, 2344, 541, 1566, 2094, 12, 2084, 1538, 1513, 1629, 1654, 163, 1324, 1796, 1675, 96, 1886, 2508, 2491, 1825, 2586, 210, 2585, 787, 44, 1489, 1702, 2477, 2549, 1752, 2686]. low- [2699]. low-cost [2421, 2586, 1489]. low-density [1825]. low-dose-rate [1629]. low-fish [1886]. low-hatch [210]. low-level [96]. low-oxygen [1702]. low-salinity [1654, 1324, 2508, 2585, 2549]. Low-salt [2187]. low-virulence [2477]. Lowe [1830]. lower [2033]. lowland [271]. LPL [1189]. LPS [1640]. LpxD [1962]. Lucas [1071]. lucioperca [1932, 642, 749, 2267, 1229, 1782, 2338, 1588, 1944, 1169, 1230, 1160, 1499, 2521, 1927, 1934, 1940, 1941, 2057, 2563, 2196, 452, 1165, 1579, 1939, 1936, 1930, 630, 1933, 1935, 641, 2099, 480, 1159, 1931, 1161, 1942, 2456]. lucius [659, 2459]. lumpfish [1655, 1401, 1572, 1856, 1756, 1755]. lumpus [1655, 1572]. lupus [1826]. luscus [1265]. Luten [690]. lutheri [427]. Lutjanus [1353, 2697, 1503]. Lutra [457, 676]. lyase [2682]. lycidas [1918]. lycopersicum [2681]. lymphocystis [2136, 2478]. lyophilized [2100, 2569, 2753]. Lyropecten [323]. Lys [2702]. Lys-c [2702]. Lys6 [150]. lysine [1699, 798, 2698, 2549]. Lysmata [403, 1025]. lysolecithin [1791]. Lysozyme [2702, 1287, 1727]. Lytechinus [561].

M [574, 1253, 692, 863, 864, 956, 1070, 1039, 1284, 448, 274, 2216, 1207]. M. [365, 366, 688, 716, 859, 955, 974, 1039, 1285, 1710, 2544]. Maas [2636]. maccoyii [1190]. macdonaldi [2708]. Machakos [977]. Macquarie [61].

Macro [415]. macro-morphological [1465]. macroalga [1772]. Macroalgae [514, 2552, 1859, 2193, 1547, 2264]. macroalgal [2173]. macrobenthos [362]. macrobrachion [2249]. Macrobrachium [2109, 1801, 2612, 2389, 1971, 1839, 1627, 2610, 270, 2430, 512, 2304, 2558, 1724, 1563, 2522, 1896, 2553, 1792, 2614, 1776, 1257, 2447, 366, 1048, 1137, 2249, 1126, 1292, 1385, 1288, 1823, 1388, 1685, 1687, 845, 2675, 1777, 1082, 2139, 2689, 1252, 1371, 2332, 1785, 1010, 2425, 293, 2665, 2247, 1807, 2113, 2711, 2510]. macrocephalus [1623]. macrochirus [874]. macrocopa [1481]. macroinvertebrates [1099]. macronutrient [742]. macrophages [1916, 2556, 1981]. macrophyte [144]. macrophytes [1891]. macropomum [2315, 2053, 2254, 2419, 2352, 2308, 2737, 2153, 2613, 2409, 2224, 1731, 2569, 2753, 2201, 2176]. macrozoobenthos [1167]. maculata [1561]. maculatum [1614]. maculatus [1995, 2007, 2206, 2441, 2440]. madaka [1172]. made [435, 1382, 2218, 1050, 147]. Madeleine [138]. maenas [938]. maeotica [500]. magellancius [281]. magellanicus [383]. magnesium [2740, 2434, 2265]. Magnifera [1467]. mail [565]. main [2625, 1280]. Maine [159]. maintained [2022]. Maintenance [1942, 378, 1365, 2456]. Maize [579, 192]. Maja [731, 1291]. Major [2125, 2567, 461, 953, 2207, 2660, 1242, 2072, 2272, 1476, 2350, 1667, 2189, 969, 562, 1481]. making [1079]. malabaricus [985]. Malaga [282]. Malaysia [2383, 2730]. Malaysian [1641, 1427]. malcolmsonii [1371]. male [2618, 1229, 2733, 2748, 2447, 1908, 2717, 2395, 2516, 2405, 1092, 1924, 2164, 2652, 2665, 2265, 272]. males [2548, 1029, 1195]. Mali [1597]. Maliakos [1275]. mallow [2003]. Malmquist [1427]. Malva [2003]. mammalian [255, 1160]. man [1382, 147, 1839, 1724, 1792, 1048, 1137, 1126, 1292, 1388, 1010]. man-made [1382, 147]. managed [1594]. Management [416, 661, 774, 868, 413, 2517, 2104, 2654, 2287, 198, 1826, 1740, 791, 1555, 2446, 598, 1862, 1588, 2027, 575, 1549, 2392, 633, 1008, 414, 1848, 2633, 1608, 2327, 926, 587, 1197, 75, 333, 576, 2359, 352, 1787, 2557, 534, 2622, 1113, 532, 20, 1378, 2350, 1880, 2584, 2259, 812, 1663, 1788, 2683, 1167, 2174, 2182, 2629, 716, 374]. managements [1565]. managing [2260]. Mandal [330]. mandarin [830, 2672, 1392]. mandibles [304]. manganese [1649]. Mangrove [552, 2381]. mangrove-shrimp [2381]. Manila [439, 871, 705, 177]. manipulate [2216]. manipulated [335]. manipulation [1826, 1229, 1588, 2233, 2397]. manipulations [124]. Manipur [2624]. mannan [1477, 1972, 2607, 2045, 917, 1254, 648]. mannan-oligosaccharides [2045]. mannanase [2611]. mannanoligosaccharides [1769]. Mantle [139, 907]. manual [466, 2106]. manure [1066, 1795]. many [1390]. map [874]. mapping [2573, 2353, 2146, 1352]. Mar [543]. marbled [2147, 1370, 2096]. Marennes [416]. margaritifera [1281, 396, 1337, 1480, 1032, 469]. marginatus [1830]. Mariculture [252, 2199, 815, 2415, 2002, 2359, 1693]. maricultured [1656]. Marine [564, 2509, 347, 861, 805, 446, 2414, 2388, 2635, 1628, 1887, 2268, 1490, 2260, 1954, 1899, 2081, 1865, 1241, 1564, 1629, 1008, 259, 305, 1897, 417, 1464, 1768, 624, 1073, 2675, 2696, 925, 1695, 1969, 209, 337, 1332, 2682, 1620, 107, 60, 289, 121, 1968, 436, 539, 2683, 1016, 1128, 2678, 1462, 2074, 2208, 605, 955]. marine-farmed [1073]. maritima [2383]. Maritime [416]. maritimum [693]. marker [738, 1728]. Markers [1346, 481, 1216, 1669, 849, 1506, 1964, 437, 2312, 1502, 2674, 1185, 1544, 1114, 874, 1060, 1707, 2339, 2628, 1392, 1509, 1761, 2120, 1644]. Market [502, 2512, 1557, 2051, 2202, 679, 1057, 1951]. marketable [548, 2731, 633, 1559]. Marketing [689, 1015, 499, 390]. markets [1967]. marmelos [2736]. marmorata [2147]. Maróz [1210]. Marphysa [949]. marron [1254]. marsh [2172]. Marsupenaeus [1256, 2491]. Mart. [1616]. Martens [865]. martensii [954, 1506, 1537]. Marthasterias [1024]. Martin [365]. Masculinization [645]. masculinized [195]. Masoten(R) [2315]. Mass [561, 1586, 2268, 1377, 1461, 1429, 2541, 2213, 2645, 2116, 1260, 1761]. Master [1408]. mat [1302]. material [632, 1110, 1856]. materials [1440, 491]. maternal [1439, 2544, 1809]. Mathematical [2744]. mating [876]. matrinxã [392]. matrix [1715, 2228]. Matsushima [1103]. matter [1107, 670]. matters [1834, 1948, 1947]. Maturation [248, 270, 482, 504, 1103, 195, 1313, 1370, 525, 1553, 1357, 1825, 1092, 840, 1296, 19, 1365, 1029]. Mature [2525, 712]. matured [121]. Maturity [248, 695, 801, 462, 123, 758, 2647, 1035]. Mau [2381]. maxima [500, 1255, 947, 2388, 1122, 1186, 601, 1271, 242]. maximizing [2340]. maximum [919, 1111]. maximus [1129, 349, 470, 1770, 707, 1179, 1996, 2416, 677, 1977, 613, 1024, 2320, 356, 394, 1824, 1108, 407, 516, 580, 748, 2307]. may [1127, 829, 121]. maya [943, 922]. Maylandia [1821, 1596]. Mayotte [1906]. MC13 [987]. MC4R [2469]. McIntyre [957]. McLarney [1282]. McVey [446]. meagre [1017, 2524, 445, 2157]. Meal [222, 1907, 799, 1445, 71, 1839, 1018, 1463, 2293, 1678, 2427, 780, 1465, 291, 846, 1538, 1047, 1141, 2748, 1541, 1706, 1142, 114, 1473, 2371, 1040, 2719, 1691, 2486, 346, 1886, 83, 1019, 880, 1863, 30, 379, 1200, 1258, 961, 2706, 41, 2726, 1044, 1496, 1900, 1125, 1316, 2098, 1625, 762, 1733, 940, 1807, 920, 1752]. meal-based [1465, 346]. meals [410, 178, 200, 1047, 1776, 322, 1571, 524, 1824, 1016, 1128]. means [818, 647, 1092, 806]. measure [1101]. measured [1217]. measurement [2716, 393, 1659, 744]. measurements [1634, 1257, 1583]. measures [1605, 1412]. measuring [758, 1236]. Meat [1125, 1018, 2336, 1541, 1257, 2301, 2728, 1592]. mechanical [778, 536]. mechanism [639, 2341, 2688, 2565]. mechanisms [1978, 1059, 436]. medaka [1224]. Media [288, 2740, 699, 1600, 1461, 1470, 2657, 2210, 1825, 1677, 2087]. media-based [2740]. median [1910]. mediated [2636, 594, 2149]. mediates [2549, 2238]. Medicinal [2403, 1294, 2571, 2502, 2064, 2721, 1252, 1775, 2724, 735, 782, 2578]. medicines [1238]. Mediterranean [1336, 247, 1275, 386, 1772, 2414, 2062, 88, 541, 40, 2509, 495, 2524, 380, 105, 542]. Medium [488, 915, 2594, 972, 1876]. medium-chain [915, 2594]. Megalobrama [1530]. megalopa [1260]. megalopae [1983]. meiotic [902, 2212, 201, 2296, 1232, 581]. Mekong [952, 1246, 1394]. Melaleuca [2553]. melanin [2676]. melanization [104, 2106]. melanocortin [2469]. melanocortin-4 [2469]. Melanogrammus [430]. melanosis [639]. melatonin [368]. melops [635]. membrane [1801, 1861, 2333]. Mendelian [675]. menhaden [944]. Menippe [2126]. Mentha [1677]. menthol [2118]. mercenaria [872, 490, 802, 872, 1454]. mercury [666]. Meretrix [1669, 1660]. Meristic [302, 155]. merluccius [1710]. Merlucius [842]. merra [695]. Merrifield [1436]. mesenteroides [1866]. mesh [370]. Mesocosm [505, 2241]. mesopotamicus [1912, 1541, 1769, 1747, 1828, 2709, 1760, 2497, 2127, 2066, 2305, 2686, 2182]. mesotrophic [1210]. meta [2721, 1007]. meta-analysis [2721, 1007]. Metabolic [387, 219, 1430, 1527, 701, 1090, 1827, 2428, 1681, 2595, 1432, 1824, 761, 2569, 2753]. metabolism [1779, 1239, 2748, 2738, 1054, 2395, 1093, 1478, 931, 2000, 2098, 1438, 1488, 2397, 2548, 1029, 1729, 2113, 425]. metabolism- [2000]. metabolite [1504, 2480]. Metabolites [220, 2664, 2747, 136, 1561]. metabolome [2395, 2156]. metabolomics [2653, 2331, 2748]. metacercariae [2248]. metacercarial [2159]. Metagenomic [1564]. metagenomics [2256]. metal [2144]. metalloproteinase [2228]. metalloproteinase-9 [2228]. metalloproteinases [1715]. metals [2413, 2091]. metamorphosis [1418, 1859, 1612, 958, 1554, 108, 311, 2283]. metanauplii [2300, 341]. Metazoan [247]. methanotroph [2427]. methionine [461, 1486, 770, 1752]. method [2601, 2433, 16, 2519, 5, 1061, 59, 1249, 1651, 2719, 2649, 1553, 559, 2181, 2073, 1042, 1876, 2149, 2464, 1831]. methodological [1637]. methodology [2078, 1205, 740, 1287]. Methods [567, 2713, 743, 2643, 1634, 1347, 1298, 2226, 1048, 2735, 1576, 2347, 670, 857]. methyl [2387]. methylbutyrate [594]. methylene [1342]. methylisoborneol [1492]. methylmercury [666]. Methylococcus [2427]. metrics [2492, 2371, 2516, 2569, 2753]. metschnikovii [936]. mexicanum [757]. Mexico [54, 169, 2627, 868, 2507, 522, 526, 787, 566]. mGnRHa [1623]. Miamiensis [1806]. MIC [1390]. mica [2131]. micelle [2693]. micro [1465, 1304, 718, 2640, 893, 2152]. micro- [1465, 2152]. micro-/macro-morphological [1465]. micro-algae [718]. micro-fouling [1304]. micro-particle [893]. micro-screen [2640]. Microalga [1841, 1954, 209, 2223, 2091]. Microalgae [235, 1885, 2406, 1178, 1075, 1628, 2046, 1629, 1391, 1897, 1014, 1614, 157, 1546, 1814, 1823, 2185, 389, 634, 2262, 2103, 289, 1858, 2300, 2728, 2678, 175]. Microalgal [1831, 1843, 699, 760, 1754, 2667, 1955]. Microarray [1026]. Microbacterium [757]. microbe [2184]. microbeads [1402]. microbes [2462]. Microbial [2202, 1546, 289, 2242, 1134, 1627, 1278, 1953, 1460, 429, 2233, 775, 2520, 1604, 1883, 678]. Microbially [121]. Microbiological [2194, 1577, 543, 885, 1150, 2011]. microbiology [887]. microbiome [2494, 2005, 2245]. microbiomes [2670]. microbiota [1477, 2090, 2733, 1774, 2169, 2354, 1895, 1126, 2472, 2707, 2322, 1630, 2518, 1063, 1535, 2066, 1069, 1890, 2723, 2745, 2634, 2247, 2219]. microcosm [602, 619]. microdiet [167]. microdiets [2496]. microecological [2408]. Microencapsulated [232, 1402, 1855, 1146]. Microencapsulation [2515, 1991]. Microeukaryote [1889]. microfeed [760]. microflora [2580]. micromorphological [1616]. microorganism [2282]. microorganisms [2063, 1957, 1202]. microparticle [724]. microparticles [2171]. microparticulate [1673, 435, 464, 1807]. microplate [1876]. micropropagation [2601]. Micropterus [2054, 1171, 2580, 1106]. Microsatellite [738, 583, 1187, 675, 852, 1848, 2244, 2312, 2674, 2048, 1060, 2339, 2628, 1392, 1509, 1761, 2120, 1728, 1644]. microsatellite-based [2244]. microsatellites [1854, 1692, 528, 1352]. microscopy [2380]. microspheres [1261]. microsporidian [1904, 1909]. microstructural [104]. microtiter [337]. micrura [68, 69, 724]. mid [2147]. mid-intestine [2147]. midae [1589, 2405, 1131, 1875]. middle [806]. midgut [2689]. Midnorth [2590]. Miedwie [1154]. migration [1328]. Miichthys [745]. miiuy [745, 745]. milestones [2272]. milkfish [1253, 418, 1660]. mill [2331]. millenium [276]. Miller [879]. millet [2514]. millet-based [2514]. Milne [2532, 753, 930, 982, 2596]. Milne-Edwards [2596]. milt [2544, 993]. miltiorrhiza [2618]. mimicus [2483]. mineral [346]. mineralization [2072, 1339, 969]. minerals [604, 2265]. mini [2150, 1276]. mini-mussel [1276]. Minimally [776, 230, 1303]. minimum [693, 2731]. Mining [1669]. Minnow [479, 573, 2324, 2368]. minor [2298, 1826, 408, 2292, 2616]. mint [1677]. miRNA [2227]. miRNA-mRNA [2227]. mirror [1362]. Mischke [1151]. Misgurnus [1157, 1156, 2244, 724, 741, 814, 831]. Mitchill [1233, 1158]. mitigate [2110]. mitigating [2154]. mitigation [607]. Mitochondrial [481, 954, 2038]. mitotic [1359, 2296]. Mitten [2361, 1680, 1956, 2227, 1142, 929, 932, 2672, 2395, 931, 753, 930, 2373, 2382, 2548, 1799, 2500, 1649, 2639]. mix [1066, 2218]. Mixed [506, 2230, 1109, 2345, 1508, 1714]. mixed-diatom [2230]. mixed-sex [1508]. mixed-size [2345]. mixing [2700]. mixotrophic [2040]. mixotrophically [2191]. mixture [2560, 2594]. mixtures [718, 2098]. Mizuhopecten [1351, 2359, 790]. mobile [2487]. mobility [1409, 1412]. Mobilization [2131]. mode [434]. model [2284, 1915, 577, 1668, 1851, 888, 886, 1004, 1345, 1549, 2002, 2092, 1797, 2703, 1965, 704, 2700, 1246, 2556, 1327, 1237]. model-based [577]. modeled [1276]. Modeling [1906, 227, 1960, 1963, 1257, 1488, 1592, 1853, 2744, 1643, 1400]. Modelling [815, 31, 1533, 320, 2682]. models [54, 995, 947, 966, 404, 1550, 1973, 790, 2620]. moderate [2699]. moderate- [2699]. Modern [1700]. modes [788, 2395]. modification [304, 1581]. modifications [255]. modified [2087]. Modiolus [1144]. modulate [1640, 2209, 2749]. modulated [2437, 1774, 1701]. modulates [2169, 2500]. Modulating [1063, 1105]. Modulation [2017, 1556, 1603, 1684, 2066]. modulatory [1833]. module [1832]. Mohit [1205]. Moina [68, 69, 1481, 724]. moist [979, 694, 211]. moisture [437]. Moksness [564, 307]. mola [1647]. molasses [2355, 1685]. molds [1987]. Molecular [1075, 2389, 895, 849, 2380, 2019, 533, 2276, 554, 2587, 2442, 2022, 2306, 2563, 2139, 1987, 1114, 2440, 2368, 2004, 2540, 2248, 1375, 885, 2227, 1675, 1155, 1667, 2339, 2688, 2379, 2441]. molecules [2137]. Molina [1306, 1539]. molitorella [2672]. molitrix [2535, 1792, 2152]. mollusc [614]. Mollusca [636, 1892]. Molluscan [2671]. molluscs [143]. Molluses [474]. mollusk [1464]. molt [1539]. molting [932]. Monastir [1748]. monitored [1236]. Monitoring [2605, 1762, 863, 864]. mono [371]. mono- [371]. monoammonium [2205]. monocalcium [2205]. Monoclonal [1321, 1835, 1277]. monoculture [722, 1714, 586]. monodon [2517, 1134, 730, 1277, 2138, 2090, 472, 1011, 640, 2325, 1086, 2736, 569, 2738, 2162, 647, 1791, 511, 958, 1278, 1435, 352, 1741, 622, 1145, 1475, 1773, 1315, 2591, 1639, 559, 1130, 2588, 952, 847, 1430, 892, 1644]. Monogenea [441, 2027]. monogenean [794]. monogeneans [2093, 2460, 2308]. Monopterus [1267]. monosex [2450]. monospecific [1612]. Monovalent [2490, 2329]. Montagu [949]. Montet [887]. month [1530]. monthly [1681]. Months [515]. Moore [736]. mooring [57]. morbidity [1939]. Moreau [841]. morhua [1140, 1309, 639, 1110, 1439, 1046, 1431, 2695, 620, 431, 2112]. Mori [1602]. moribund [2060]. Moringa [2025, 1307, 2009]. Moritella [1993]. Morogoro [911]. Morone [2107, 852]. Morpho [2248]. Morpho-molecular [2248]. Morphoecology [424]. morphogenesis [934, 186]. Morphological [2159, 591, 1234, 2424, 2687, 2754, 1465, 123, 1230, 1651, 2263, 2379]. Morphology [512, 2370, 1477, 1750, 2147, 1226, 2743, 1538, 1769, 1434, 1372, 1751, 969, 1890, 2634, 1517, 1604, 2245, 2130, 272]. Morphometric [2523, 58, 754, 1155, 1292, 1050]. morphometry [1566, 2151]. morphotypes [2068]. Morrissey [861]. mortalities [1986, 1429, 2194, 1260]. Mortality [1344, 287, 111, 1482, 159, 642, 282, 2268, 802, 2690, 1370, 1337, 1806, 303, 1969, 177, 1939, 1454, 2645, 1597, 1881, 2116, 2397]. mortem [1179, 975, 557]. MOS [1465, 917]. mosaic [1157]. Mosig [565]. mossambicus [1679, 1921, 1440, 2371, 2473, 1556, 2482, 1631, 2273]. most [2438]. motile [1127, 2290, 1663, 2009]. Motility [485, 1221, 1222, 588, 1217, 674, 1219]. Moult [1011]. moulting [935]. moving [1968]. Mozambique [1440, 2473, 241]. Mr [1371]. mrigal [953]. mrigala [2567, 461, 953, 1736]. mRNA [2227, 1287, 1649]. MrNV [1010]. MS [2577, 1314, 2539]. MS-222 [2577, 2539]. MS222 [649]. mucosa [2127]. Mucosal [2572, 2482, 2435]. mucus [1995, 1242, 794, 2396, 2495]. Mud [571, 2080, 934, 1377, 545, 423, 2555, 2672, 1325, 1368, 324, 1983, 937, 1243, 928, 2141, 1646, 1483, 933]. mudflats [1120]. Mugil [1888, 2248, 549, 2023, 471, 555]. mugwort [2357]. Müller [1923, 1814]. Mullet [471, 1888, 549, 1730, 2023, 555, 1425, 89]. mullets [2159]. mulloway [759]. Mullus [1730]. Multi [682, 1965, 1245, 819, 2279, 2155, 2509, 2014, 2158, 1176, 1521, 1391, 1606, 1546, 1288, 357, 2292, 1462, 2438].

Multi-criteria [682, 2155]. multi-generation [357]. multi-loop [2158]. Multi-model [1965]. multi-phase [2438]. multi-trait [1288]. multi-trophic [1245, 819, 2279, 2509, 1176, 1391, 1606, 1546, 2292, 1462]. multi-use [1521]. multi-virulence [2014]. multidrug [950]. multidrug-resistant [950]. Multienzymatic [1992]. Multifunctional [658]. Multiple [1886, 620, 1708, 120]. Multiplex [1187, 1197, 1697, 1476, 2048]. multispecies [1646, 2520]. multitrophic [2411, 1834]. multivalent [1993]. multivariate [576]. Murcia [88]. murrel [2058, 2037, 2402]. musciformis [1867]. muscle [901, 1000, 64, 856, 1179, 2638, 196, 755, 912, 1231, 976, 2608, 983, 1082, 473, 840, 1296, 1163, 202, 61, 1900, 557, 891, 1662, 1840, 2549, 2720, 1405, 2130]. mushroom [1713]. mushrooms [1192]. Mussel [502, 1275, 1178, 615, 2062, 1838, 2388, 1734, 2279, 1251, 1067, 1569, 1305, 681, 2501, 804, 1043, 828, 23, 986, 1080, 916, 754, 777, 1329, 2101, 737, 1057, 2272, 1144, 718, 2630, 1276, 637, 2726, 2264, 2191, 733, 1659, 2044, 2283, 1051, 1527, 857, 2004, 1195, 744]. mussels [1772, 2268, 632, 1174, 1553, 2547, 139, 1276]. mutant [1872]. mutiara [2190, 2559]. mycelia [1846]. mycobacteriosis [2490]. Mycobacterium [2284, 2490]. Mycoplasmas [2060]. mycotoxin [1250]. myersiana [754]. mykiss [2722, 1295, 1133, 444, 574, 551, 2358, 1603, 1997, 1568, 2299, 1047, 1141, 1311, 1470, 2226, 2328, 1852, 604, 1469, 409, 1920, 2334, 2644, 606, 1982, 1065, 2457, 482, 1713, 1846, 1658, 1783, 1476, 1096, 343, 1432, 1873, 1163, 1847, 820, 2746, 1978, 837, 2726, 1044, 1719, 648, 387, 1758, 1656, 762, 1662, 1387, 2222, 940, 1360, 2011, 2204, 1139, 2685]. Mylonas [974]. Mylopharyngodon [2120]. Mymensingh [2376]. myo [1857]. myo-inositol [1857]. myogenic [1444]. myonecrosis [2102, 1397]. MYP [1419]. mysid [1358, 207, 358]. Mysidopsis [358]. mysids [350]. mysis [1035]. Mystus [2291]. Mytilicola [2488, 1602]. Mytilidae [1306]. Mytilus [1772, 2062, 1838, 1870, 1067, 2523, 1610, 1305, 681, 2501, 791, 804, 828, 916, 558, 1602, 1553, 737, 718, 1276, 2004, 744]. Myxobolus [2023]. Myxosporean [252].

N [1351, 726, 692, 1039, 274, 2496, 1134, 1780, 2140, 720, 2355, 1685, 1594, 2537, 2718]. n-3 [2496]. N-acyl [1780]. N-CP [1792]. nacre [1734, 1281]. Nadu [713, 1183]. Nagabhushanam [688]. Nakazawa [1035]. Nannochloropsis [2192, 1814, 1798, 2262, 2103, 2143]. nano [2633, 2290, 2152]. nano-emlusion [2290]. nano-emulsions [2633]. nano-scale [2152]. nanocolloids [2216]. nanoemulsion [2128, 2357]. nanoemulsions [2457]. nanoflower [2677]. nanoformulations [2392]. nanomicelle [2518]. nanoparticle [1619]. nanoparticles [1957, 2232, 2299, 2428, 2636, 2663, 2470]. nanosilver [1470]. nanosilver-coated [1470]. nanotechnology [2252, 2259]. NaOH [1131]. Narrow [1305]. Nash [1009]. National [1055]. native [255, 1197, 126, 2076, 2182]. Natural [764, 9, 2061, 1995, 1578, 2235, 2284, 731, 903, 1358, 278, 1077, 321, 703, 2633, 2203, 2405, 1964, 2006, 400, 1594, 2255, 630, 844, 837, 1933, 292, 2070, 397]. naturally [1980, 885, 1065]. nauplii [1770, 687, 796, 39, 1498, 833, 781, 915, 1263, 1435, 1823, 78, 433, 770, 798, 465, 2462, 1132, 1466]. Nauru [418]. Navicula [1959]. ND [1154]. ND-1 [1154]. near [1720, 95]. near-future-predicted [1720]. nearshore [1067]. necessary [1532, 2020]. necrosis [2137, 1377, 2605, 1956, 1510, 2218, 1441, 2576, 1974, 1065, 2010, 2102, 2070, 2116, 1868, 2163, 2685]. Need [572]. needed [74]. needs [1411]. negative [2516, 1658, 1783, 1805, 1435]. negatively [2697]. negatives [2195]. nematocyst [2636]. Nematoda [2028]. nematode [1241, 921]. Neochinorhynchidae [2315]. Neoechinorhynchus [2315, 2254, 2613, 2409]. Neohelice [2566]. neomale [1320]. Neomysis [1035]. neon [2696]. Neoporphyra [2727]. Neotropical [2007, 2648, 2568, 1894]. nepa [1839]. nervous [1974, 2116, 1868]. nest [2338, 2516]. nest-spawning [2338]. nested [730, 1065]. nested-PCR [1065]. NEt [1158, 749, 1210, 384, 1304, 2637, 2343, 971, 120]. net-isolated [971]. netcages [550]. Netherlands [2268, 66]. Nets [500, 1634]. network [1980, 1407, 1413, 966, 2077]. neural [966]. Neutral [2237]. neutralizing [1587]. Nevada [621]. Newfoundland [2036]. newly [2046, 1903, 2292, 1694, 1618]. News [564]. next [276, 1544, 2163]. next-generation [1544, 2163]. NF [1816]. NF- [1816]. Ng [956]. NHPB [2043]. niacin [953]. Niantic [278]. Nibea [1552, 1350, 2130]. Nigeria [2499, 2214, 2615]. nigrofasciatum [645]. Nile [1812, 1957, 1033, 1448, 1888, 2729, 248, 1763, 1194, 2439, 2093, 2666, 2131, 799, 2083, 2122, 1319, 1000, 2298, 1709, 2095, 2450, 2734, 2025, 899, 301, 1764, 2732, 2380, 2483, 2593, 780, 178, 222, 291, 1889, 2161, 1515, 2169, 2354, 189, 2452, 1897, 2183, 2633, 2017, 855, 853, 1181, 48, 1508, 1675, 2030, 1447, 1893, 1714, 83, 2166, 2592, 2647, 2645, 1372, 442, 1200, 1762, 906, 2290, 1138, 1339, 2577, 2539, 2700, 1125, 1316, 1890, 1867, 1022, 2145, 2582, 1322, 1444, 1759, 2061, 2661, 2424]. niloticus [2242, 1812, 1986, 2063, 1957, 1033, 1448, 1888, 2729, 1763, 1194, 2439, 2666, 2131, 799, 2083, 1679, 2122, 1319, 2298, 2095, 2450, 1990, 2731, 2418, 2734, 2025, 899, 1764, 2380, 2483, 2593, 780, 1921, 2161, 2084, 2169, 1640, 1188, 1792, 1501, 2452, 2183, 2371, 2633, 1442, 2017, 855, 853, 1181, 1765, 1453, 1508, 1675, 2482, 2030, 1631, 1447, 1893, 2306, 1478, 2592, 1287, 1562, 2647, 2645, 2194, 1372, 442, 1200, 1762, 2290, 1138, 1339, 1250, 2577, 2539, 787, 834, 973, 2700, 1125, 1316, 854, 971, 2116, 1867, 1189, 1718, 1022, 2145, 1443, 1657, 1840, 2144, 2582, 1322]. niloticus [746, 1444, 2661, 2424]. nine [1385]. nipon [2359]. nippona [1820]. nipponense [1776, 2447, 2665, 1807, 2113]. nisin [2056]. Nitrate [1808, 1356, 1379, 1103, 2013]. Nitric [1718]. nitrification [1970, 793]. nitrifying [2367, 2684, 1639]. nitrite [2319, 924, 2227, 1171, 935, 1816, 1536, 1493, 1095, 1022]. Nitrogen [1709, 640, 2446, 1020, 2083, 1690, 1913, 2385, 1120, 769, 2684, 1458, 2449, 1822, 2092, 2121, 2233, 2642, 2481, 1594, 306, 1130, 1493, 1199, 542, 1625, 940, 1424]. nitrogenous [1668]. Nizimuddinia [2552]. nm [218, 219]. NMR [2653, 1504]. No [1166, 1743, 1780, 1335]. noae [1144]. Noah [1144]. noatunensis [1762]. nobilis [2426, 1816, 1815]. Noble [1426]. Nocardia [1969, 1744]. noctiluca [1626]. NOD [2741]. NOD-like [2741]. nodavirus [1371, 1010]. nodifrons [2126]. Nodipecten [323]. nodosus [323]. Nodulisporium [2480]. nom1 [2534]. nomenclature [1811]. Non [1501, 393, 2253, 2246, 970, 632, 807, 1680, 134, 2607, 2111, 2231, 2447, 793, 2282, 1553, 1197, 1675, 74, 2514, 2346, 2151, 46, 559, 1892, 1834, 121, 292, 2396, 1758, 1868, 2724, 1192, 1403, 1794, 2536, 2152, 1540, 1503, 2318, 2661, 1511]. non-additive [2246]. non-anadromous [134]. non-CpG [1511]. non-dietary [46]. non-disruptive [632]. non-eutrophic [1834]. non-hardened [2282]. non-indigenous [807]. non-industrialized [74]. non-integrated [292]. Non-invasive [393, 2253, 970, 1892, 2661]. non-lethal [1553, 559, 1868]. non-medicinal [2724]. non-native [1197]. non-opportunistic [121]. non-penaeid [793]. non-permeable [1758]. non-reproductive [2447]. Non-specific [1501, 1680, 2607, 2231, 1675, 2514, 2346, 2151, 2396, 1192, 1403, 1794, 2536, 2152, 1540, 2318]. non-starch [1503]. non-vaccinated [2111]. nonspecific [2672, 2717, 2049, 739]. normal [1320]. normality [631]. Normandy [558, 2528]. norsk [708]. north [1210, 2476, 92, 2499, 1844, 1521]. north-central [2476, 2499]. north-eastern [1210]. Northeast [2445, 610, 1964]. northeastern [596, 597]. northern [728, 1103, 1648, 1602, 889, 1749, 1726, 1378, 1243, 2349, 2154, 2306, 534, 284]. northwest [380, 2369, 522, 526]. northwestern [2523]. Norway [945, 1129, 349, 1985, 613, 1683]. Norwegian [1915, 1756, 1755]. notata [802, 872]. Note [1518]. Notes [314, 1782]. Notices [223, 244, 253, 21, 34, 43, 72, 117, 141, 152, 179, 194, 261, 98, 27, 51]. notochord [1507]. Nottinghamshire [489]. novel [2612, 1529, 2641, 1073, 2252, 2674, 2677, 2640, 1707, 2688, 2565]. November [926]. NPY [2323]. NRC [1055]. Nrf2 [2418, 2491, 2732]. Nrf2-Keap1 [2491]. Nuclear [481, 1212]. nucleotide [946, 2430, 1896, 1010, 2099]. nucleotides [1538]. nucleus [1119, 1337]. nudiventris [1232]. number [406, 633, 1127]. numerical [1246]. Nursery [566, 1955, 2370, 2433, 2610, 198, 2532, 13, 1567, 2094, 1052, 281, 1526, 2343, 805, 602, 619, 469, 1712, 1452, 1483, 933, 2510]. nursing [662]. nutraceutical [2127]. Nutrient [1055, 2374, 1772, 1020, 2740, 1514, 1525, 1168, 384, 1784, 846, 2158, 1047, 1141, 1421, 1584, 1706, 1151, 1040, 1265, 2017, 984, 627, 1453, 606, 2196, 1713, 1104, 538, 1200, 211, 2611, 1246, 1625, 762, 1682, 2087, 2318, 2167, 353]. nutrient-enriched [1104]. nutrients [1989, 433, 1279, 1146]. nutrition [1671, 1436, 1523, 102, 309, 367, 844, 668, 650, 991]. Nutritional [1923, 157, 1823, 1200, 2552, 430, 2581, 1326, 1889, 2638, 2202, 22, 2291, 2505, 1796, 1880, 537, 2616]. nutritive [1258]. nuttallii [2720]. NV [1371]. NW [103, 1305, 868]. nyassae [2118].

O. [1679, 1188, 691, 2371, 1453, 2482]. O.-I. [691]. oat [2525]. Oaxaca [787]. obliquus [2091]. O’Bryan [709]. O’Bryen [700]. Obscure [496]. obscurus [496]. Observation [218, 1550]. Observations [2268]. obtained [948, 1934, 140]. occurrence [2475, 2555, 918, 1435, 2106, 1507]. ocean [1720, 1838, 1913, 805, 2750, 1717, 1906]. oceanica [2143]. oceans [957]. ocellaris [2436, 1124, 1536]. ocellatus [1446, 1906, 675]. Ocimum [2131]. O’Connor [859, 1285]. Octopus [515, 1004, 903, 816, 842, 2081, 924, 907, 1004, 404, 515, 1291, 943, 922, 900, 2211]. oculata [1798, 2262, 2103]. Odisha [1400]. odor [1319]. odorous [1802, 823]. Oehlenschläger [690]. OER [1911]. Off [486, 2062, 2015, 535, 1548, 1615, 356, 1390, 665]. off-bottom [2015]. off-flavor [1615, 665]. Off-Season [486]. off-shore [535]. offer [737]. offers [2600]. officinale [2736]. officinalis [350, 391, 438, 458, 568, 712, 773, 832, 1882, 1920, 468, 628]. offshore [1067, 1521]. offspring [270, 423, 456, 1439, 1809]. Oil [513, 69, 1255, 1295, 444, 2656, 2385, 2651, 2095, 2128, 2419, 2299, 1047, 1188, 1109, 944, 821, 956, 1855, 1791, 157, 37, 1833, 606, 166, 983, 2457, 1208, 1601, 1599, 1478, 2697, 2357, 547, 2290, 1163, 844, 837, 2562, 436, 1890, 789, 1543, 2130, 36]. oil-based [2385, 2130]. oil-coated [1601]. oiled [2340]. oils [2460, 1047, 1109, 2633, 2621, 925, 367, 1029]. oilseed [200, 1047, 1447, 1258]. Oithona [1614]. Olafsen [408]. old [455]. oleaginous [1858]. oleifera [2025, 1307, 2009]. oleracea [2100, 2569, 2753]. Oléron [416]. oligo [2556]. oligo-porphyran [2556]. oligodeoxynucleotides [1511, 2689]. oligosaccharide [1477, 1972, 1254, 648]. oligosaccharides [2607, 2045, 601, 917]. olivaceus [996, 1301, 2693, 1806, 748, 920]. Olive [844, 2128, 1301, 2171, 2693, 1806, 837]. Olsen [564]. olseni [2547]. omega [2651]. omega-3 [2651]. Ompok [2293, 2445]. on-growing [1522]. Onchorynchus [837]. Oncorhynchus [2722, 1295, 1133, 444, 574, 551, 2358, 1603, 1997, 1568, 2303, 2299, 1047, 1141, 1311, 1470, 2226, 2328, 1852, 604, 1469, 409, 1920, 2156, 2334, 2644, 606, 1982, 1065, 2457, 482, 1713, 1846, 1658, 1783, 1476, 1096, 343, 1432, 1873, 1163, 1847, 820, 2746, 1978, 2726, 1044, 1719, 648, 387, 1758, 1656, 762, 1662, 1387, 2222, 2420, 940, 1360, 2011, 2204, 1139, 2685]. one [1269, 730, 123]. one-sea-winter [123]. one-step [730]. ongrowing [888, 886, 1004, 1579, 1942, 2456]. ongrown [13]. Onion [2593]. only [628]. onset [833]. Onsite [2167]. onto [696]. ontogenesis [741]. ontogenetic [1943]. Ontogenic [2485]. ontogeny [1228, 1399, 696, 28, 2424]. Oocystis [2092, 2541]. oocyte [2237, 993]. oocytes [61]. oomycetes [1987]. Oosterschelde [2268]. Open [1911, 1973, 1245, 1838, 1712, 2040]. open-sea [1245]. operated [1968]. Operating [2233, 868]. operation [652, 756]. operations [2700]. opinions [2527]. Oplegnathus [1757, 1369, 1363]. opportunistic [121]. Opportunities [2476, 706, 1410, 2626, 2752, 1911, 2730]. opportunity [836]. optical [1634]. Optimal [1915, 492, 2054, 2467, 545, 2249, 1300, 1121, 1490, 811, 2598, 1079, 880, 1225].

Optimisation [828, 2226, 1349]. optimising [1347]. Optimization [2731, 2558, 1606, 1247, 2320, 1287, 2682, 2700, 869, 5, 2704, 2541, 2065, 1689]. optimize [1915]. Optimized [2078, 1944, 1345]. Optimizing [1859, 2695]. Optimum [546, 817, 1123, 1653, 384, 964, 1831]. option [2627]. Oral [2736, 2169, 855, 2561, 1105, 2192, 2604, 1261, 42, 2490, 2171, 941, 1387]. orally [1723]. orange [2687, 2754, 2206, 2621, 2022, 1455]. orange-spotted [2687, 2754, 2022, 1455]. Oratosquilla [1839]. Orbetello [538]. orbital [166]. orders [565]. oregano [1833, 2621]. Oreochromis [2242, 1812, 1986, 2063, 1957, 1033, 1448, 1888, 2729, 1763, 1194, 2439, 2666, 2131, 799, 2083, 1679, 2122, 1319, 2653, 2298, 2095, 2450, 1990, 2731, 2418, 2734, 2025, 899, 1764, 2380, 2483, 2593, 780, 1921, 2161, 1440, 2084, 2169, 1640, 1188, 1792, 2111, 1501, 2452, 1919, 2183, 2371, 2633, 1442, 2017, 855, 853, 1181, 1765, 1453, 1508, 2473, 1675, 1556, 2482, 2030, 1631, 1447, 1893, 2306, 1478, 1760, 2592, 1287, 1562, 2647, 2645, 2194, 1372, 442, 379, 1200, 1762, 2290, 1138, 1339, 1250, 2273, 2577, 2539, 787, 834, 973, 2700, 1125, 1316, 854, 971, 2116, 1867]. Oreochromis [1189, 1718, 1022, 2145, 1443, 1657, 1840, 2144, 2582, 1322, 1815, 746, 1444, 2661, 2424]. Organic [1006, 1041, 2104, 2036, 2294, 1393, 384, 1110, 1593, 1104, 1376, 2180, 1107, 2396, 653, 539]. organically [2660]. organics [1494]. organics-degrading [1494]. organised [1325]. organism [545, 2270]. organisms [1701, 2202, 2750, 1686, 688]. organization [2373]. Organochlorine [506]. organs [1211]. oriental [2447, 2665, 1807]. orientalis [1121, 2488, 1602, 1762, 1111]. oriented [2574]. Origanum [1833]. origin [771, 829]. originating [204]. Orissa [523, 1027]. ornamental [1342, 1894, 2363, 729, 2696, 2033, 2312, 2369, 2255]. ornata [2632]. ornatus [696, 1458, 2706]. orthogonal [2652]. orthologues [2442]. Oryzias [1224]. OsHV [1344]. OsHV-1 [1344]. osmolality [454, 1299]. Osmoregulation [2548, 1635]. osmosis [2000]. osmosis- [2000]. Osmotic [485, 136, 92, 1299]. Osphronemus [2284, 2490, 2070]. Osteichthyes [537]. Ostrea [1245, 1402, 1643]. ostreatus [1846]. ostreid [1482]. other [1358, 106, 974, 305, 2547, 107, 1834, 2444]. otter [457]. otters [676]. Out-of-Season [480, 1159]. Outbreak [2604, 2033, 2746, 2070]. outbreaks [2483, 395]. outcomes [1788]. Outdoor [1534, 470, 199, 1814]. outer [1801, 2333]. outlook [2129]. Output [1110]. outputs [1906, 1693]. Ovalipes [2377]. Ovaprim [993]. ovarian [1262, 525, 19]. Ovary [1710, 1357, 101]. ovatus [1854, 1841, 508, 2074]. overfeeding [590]. overfished [284]. overlapping [2464]. overview [2751, 1408, 1533, 2047, 1663, 2406]. overwintering [802]. ovigerous [1844]. Ovopel [1213]. Ovulated [230]. Ovulation [496, 1160, 1542, 1623, 1158]. Oxford [564, 554]. oxidant [2420]. Oxidase [506]. oxidation [1308]. oxidative [2297, 2743, 1999, 2428, 2365, 2322, 2220, 2497, 2630, 2548, 1561, 2658]. oxide [2428, 193, 1718]. oxidised [436]. Oxidized [2697, 1729, 1604]. oxidizing [1970]. oxolinic [620]. Oxygen [495, 49, 1448, 31, 2367, 1351, 1816, 1101, 156, 2584, 704, 1702, 443, 592, 745]. oxygen-carrying [1816]. Oxygenated [2662]. oxygenation [1960]. oxytetracycline [2450, 395, 891]. Oyster [235, 236, 566, 808, 1245, 760, 1281, 2701, 88, 2246, 2714, 2509, 1402, 1186, 396, 560, 786, 807, 1703, 2132, 1491, 1118, 954, 719, 1513, 1298, 1102, 1643, 1786, 1337, 1480, 199, 1032, 984, 2016, 2735, 1692, 401, 1778, 809, 1197, 1506, 1952, 469, 427, 806, 2528, 1482, 1505, 1537, 1707, 1820, 2021, 2174, 2692, 1761, 1352, 2453]. Oysters [250, 460, 2015, 1513, 1197, 2213, 1712, 672]. ozonated [618]. ozone [1916].

P [574, 129, 354, 449, 700, 709, 859, 955, 1071, 1285, 1283, 446, 625, 720, 2568]. P. [1750, 1011, 2297, 679, 709, 1056, 1751, 2076]. P.E.P [2455, 2399]. p53 [2278]. p53-BAX [2278]. p53-BAX/ [2278]. Pabda [2360, 2445, 2293]. Pacifastacus [638, 833, 752]. Pacific [2306, 2699, 1907, 760, 1290, 1904, 238, 1121, 460, 1248, 1119, 1375, 1384, 2427, 2714, 786, 2319, 169, 807, 1865, 1396, 1379, 1860, 1901, 1487, 824, 1198, 1298, 1102, 1654, 2327, 2487, 199, 2314, 1681, 1324, 1692, 2472, 1778, 2322, 2608, 2518, 2079, 1019, 373, 2006, 2049, 427, 1502, 806, 1111, 2295, 1482, 2000, 2594, 2100, 2102, 1302, 672, 1036, 1389, 2520, 2018, 2480, 1707, 2655, 2549, 2021, 1761, 1976, 1352, 1397, 2453]. pacifica [2467]. packaged [1576]. packed [1920, 1808]. pacu [2225, 1541, 1769, 2709, 1760, 2497, 2127, 2305, 2686, 2182, 2066]. Paddlefish [230, 661, 454]. Padhi [330]. Padina [2552]. PAGE [160, 344]. Pagellus [503, 504]. pages [552, 554, 553]. Pagrus [562, 1018, 975, 1481, 1054, 494, 246, 1053]. pairing [2464]. Pakistan [2659, 2616]. Palaemon [732]. Palaemonetes [732, 628]. Palaemonid [374]. Palaemonidae [512, 732]. Pallas [479, 573, 1473, 1617]. pallipes [345, 369]. palm [1791, 2562]. Palmaria [1471]. palmata [1471]. Pampus [2228]. Panagrellus [921]. Panagrolaimus [1241]. pancreatic [1065]. pandemic [2507]. Pandora [504, 338, 503]. panel [2048]. Pangasianodon [1477, 1972, 1781, 2492, 2514, 1150, 1661, 1653, 2330, 2301, 1394, 2698, 2646]. pangasius [2413, 2277, 1818, 988, 2554, 2659, 1394]. paniculata [472]. Panopea [1958]. Pantoea [897]. Panulirus [696, 1458, 426, 2706]. papaya [2647]. Paperback [1008]. Papers [231]. Paracanthurus [1375]. Paracentrotus [1859, 1610, 923, 2230, 2055, 800, 1767, 1365]. Parachlorella [2728]. Paracoccus [1970]. paradigm [2557]. paraffin [559]. parahaemolyticus [2604, 2605, 2471, 2538, 2148, 1676, 2162, 915, 1441, 2209, 1962, 2723, 1182, 1966, 2163]. paralarvae [924, 1291, 900, 2211]. Paralarval [515]. Paralichthys [996, 1301, 2693, 1806, 432, 748, 920]. parallelus [1317]. paramagnetic [2428]. paramamosain [423, 1983, 1646, 933]. Parambassis [2478]. parameter [1459, 2704]. Parameters [218, 2729, 1196, 671, 2059, 1907, 1645, 1763, 1734, 1119, 535, 2078, 2714, 1213, 2297, 1242, 336, 1764, 2434, 2107, 1622, 2299, 2140, 1090, 1552, 8, 1224, 1217, 1442, 1449, 1736, 2739, 1725, 1468, 1914, 1681, 2449, 1977, 2463, 2417, 2737, 2034, 2320, 2233, 2263, 1687, 2329, 1484, 2514, 1400, 2709, 1658, 1783, 2497, 1252, 1221, 1222, 820, 2224, 2289, 2324, 2435, 2556, 2145, 1657, 1840, 2634, 2536, 1360, 735, 2646, 2219, 1722, 1988, 1731, 97]. Paramisgurnus [2244, 711]. Paraneetroplus [1636]. Parasite [1656, 2404, 1904, 38, 2027, 1909, 2724]. Parasites [247, 1056, 794, 1982, 2547, 2647]. Parasitic [2369, 2027, 2751, 2691, 2645, 2144]. parasitica [1980, 2593, 2390, 2275]. Parastacidae [822, 865]. Paratelphusa [2387, 2632]. parauberis [2171]. Parentage [1392, 1509, 1644, 1187, 2244, 1667, 1060, 2628]. Parental [246, 1922]. parenteral [855]. Parke [1609]. parr [313, 779]. parrot [1636, 1757]. parsley [1760]. part [1782, 737, 718]. parthenogenetic [1696]. Partial [291, 1142, 943, 2588, 396, 1047, 696, 1473, 2448, 1154, 983, 2243, 920]. Particle [125, 127, 1061, 1349, 893]. particle-associated [1061]. parts [2461]. Passing [1164]. Passive [1277, 2603, 1053, 1469, 1387]. past [1523]. paste [1831]. Patagonian [2377, 2412]. paternal [2544]. pathobiology [1056]. pathogen [2677, 1132, 2402, 2275]. pathogenesity [647]. pathogenic [1957, 2137, 2075, 1429, 2390, 465, 1639, 2546, 2089]. Pathogenicity [2273, 1986, 2502, 140, 2261, 2746, 2582]. pathogens [2019, 1061, 915, 2403, 2443, 1605, 2645, 1697, 1937, 2216]. pathological [2644, 1796]. pathology [1072]. pathway [1816, 2397, 2278]. Pathways [2421, 2418]. Patinopecten [2002, 1318]. Patos [1804]. pattern [562, 845, 136, 537, 1799]. patterns [62, 1838, 2253, 106, 1209, 976, 2411, 2563, 1965, 654]. paulensis [1804, 981, 2106]. Pavlidis [974]. Pavlova [427]. Pawpaw [2647]. Paz [566]. Pb [2688]. PB45 [1668]. Pbk [565, 563]. PBLs [2280]. PCR [1479, 1269, 730, 2702, 1655, 1187, 569, 889, 1435, 1197, 1065, 1476, 1697, 2195, 2309, 2048]. PCR-based [1269]. PCR-negative [1435]. Pdp11 [1089]. Pearl [235, 566, 1734, 1281, 2246, 1186, 396, 1043, 2733, 954, 719, 1080, 754, 1337, 1480, 1032, 777, 1329, 2016, 401, 1506, 139, 469, 1505, 733, 1537, 1195]. pearlscale [2712]. Pecten [1129, 349, 707, 613, 1024, 356, 407, 516, 580]. Pectinidae [283]. pectoralis [1800, 2048]. pedigree [1743, 1692]. Pediococcus [2247]. pediveliger [2241]. peel [1274, 2621, 2011]. Pelagia [1626]. pelagic [1003]. pelagicus [935, 1618]. pellet [694]. Pelleted [227, 2725, 388, 1453, 2188, 1682]. pelletization [1858]. pellets [2533]. Pelodiscus [1012, 1700, 1752]. Pelteobagrus [2297, 881, 1488, 1665, 1509]. pen [2637, 1074]. penaei [2043]. penaeid [2125, 793, 26, 140, 2106, 1114, 2406]. Penaeus [2243, 2370, 2560, 1907, 2517, 1134, 2088, 730, 1290, 1277, 1119, 2138, 2090, 472, 1011, 2313, 2532, 640, 2325, 1086, 381, 2736, 569, 1467, 2738, 2162, 647, 2690, 2231, 1706, 1791, 511, 958, 1278, 2739, 2327, 2487, 2576, 2203, 1435, 352, 1741, 2486, 622, 1145, 1475, 1773, 2508, 373, 1994, 2551, 2596, 1315, 2591, 1639, 559, 2531, 2507, 1130, 2058, 2037, 2667, 2585, 2588, 513, 1302, 952, 2119, 522, 526, 971, 2454, 847, 891, 1146, 1430, 892, 1644]. Penang [2383]. Pennisetum [1307]. pens [1804, 1401]. pentosaceus [2247]. pepo [1989]. peptidase [2424]. Peptide [2354, 2192]. peptides [2635]. peracetic [1147]. Perca [1166, 2287, 1929, 1171, 486, 1928, 1943, 1937, 1221, 1222, 1938, 363, 443].

Perceived [229, 2327]. perception [646]. perceptions [414, 747, 2527]. Perch [486, 300, 1166, 2287, 2394, 2461, 1230, 1929, 1171, 493, 1928, 1943, 1937, 1221, 1222, 630, 61, 1933, 1948, 1947, 1466, 1744, 363, 443, 2323]. percid [1946, 1945]. percula [2136]. perenurus [479, 573]. Performance [470, 1001, 1567, 222, 2205, 1953, 494, 509, 2640, 1033, 1448, 2031, 2729, 1578, 1772, 1907, 2740, 1446, 1194, 1255, 2617, 788, 799, 1049, 1445, 2716, 2003, 1067, 1358, 2015, 270, 827, 909, 2427, 1960, 2298, 340, 1997, 2313, 867, 2651, 2394, 2455, 2399, 423, 899, 2437, 981, 1202, 1173, 1923, 2094, 2338, 1226, 1396, 1676, 2299, 1328, 1474, 1849, 2367, 1818, 2084, 1456, 1538, 1563, 1047, 1077, 1141, 2436, 2522, 1188, 1999, 2024, 1841, 2738, 1769, 2614, 824, 1135, 768, 1364, 8, 50, 29, 975, 1596, 1422, 305, 2183, 1473, 2693, 1486]. performance [1481, 1442, 1278, 1736, 2739, 409, 2314, 1308, 2672, 2032, 1757, 2472, 2707, 550, 1312, 2220, 2554, 2580, 1684, 1508, 1796, 2511, 1675, 1433, 1685, 2316, 1368, 1687, 2329, 845, 2072, 1590, 1145, 983, 1886, 1773, 2079, 2508, 1714, 710, 2709, 1478, 1661, 1428, 2503, 427, 718, 1378, 1559, 2697, 2647, 1372, 1376, 822, 1930, 1339, 630, 1426, 820, 2391, 2726, 2562, 2100, 121, 2188, 1723, 1496, 648, 753, 930, 1374, 2098, 641, 1646, 1025, 1795, 2480, 2698, 1625, 2361, 1537, 2495, 2382, 2432, 2742, 2745, 1543, 208, 685, 1840, 1350, 1607, 914, 1148, 2247, 920, 2710, 2152, 2219, 1405, 2245, 1649, 939]. performance [831, 1500, 2113, 1731, 2091, 2510, 246]. performances [2452]. performing [1288]. Pericaridea [358]. period [1216, 662, 186, 2316, 2537, 1092, 67]. periodicity [2465]. periods [1703, 426, 905]. Perionyx [2293]. peripheral [2280]. periphery [523]. periphytic [2203]. periphyton [1134, 2513]. Perkinsus [2547]. permeabilizer [2560]. permeable [1758]. Perna [1178, 615, 2279, 1251, 986, 2101, 2547, 2272, 2630, 2264, 1051]. Pernambuco [1620]. peroxidation [1626, 2282, 1106]. peroxide [16, 1342, 1259, 2282]. Persian [2702, 1227, 1903, 1094, 875, 776]. persicus [2702, 1227, 1903, 994, 1094, 875, 776]. persistence [1236]. persistent [343]. perspective [2599, 74, 1605, 1095]. Perspectives [541, 1180, 53, 2302, 1519, 737, 1911, 1683, 563]. pertusa [1819]. Peru [2512, 2028]. Peruvian [2460]. pesticide [2053]. Petrie [1283]. PGF [1828]. pH [897, 1958, 325, 1131, 1367, 360, 2616, 870]. Ph.D. [1409]. Phaeodactylum [2581]. Phaeophyceae [1501]. phage [2592]. phage-1A [2592]. phagocytic [1912]. phagocytosis [2114, 1981]. Phalacrocorax [656, 2475, 918, 1050]. pharaoh [825]. pharaonis [825, 357, 2575]. pharmacokinetic [620]. Pharmacokinetics [1921, 891]. Pharmacological [1059]. phase [1628, 2741, 2532, 2094, 203, 1609, 409, 2521, 2737, 2585, 2383, 2438, 1759]. phase-dependent [2741]. phased [845]. phases [754]. phenol [2479]. phenolic [2502]. Phenoloxidases [2591]. Phenotypic [937, 1743, 2342, 1341]. phenoxyethanol [578]. Philipose [1284]. philippinarum [439, 871, 705, 536, 829, 951]. Philippines [417, 805, 2671]. Philipsson [2388]. philoxeroides [1616, 2720]. phosphate [897, 321, 2205, 826]. phosphate-solubilizing [897]. phosphates [2205]. phosphatidylcholine [166, 2496, 959]. phospholipids [270, 433]. phosphoric [2335]. phosphorous [2385]. phosphorus [71, 1672, 827, 769, 2426, 881, 2608, 1832, 1586, 1130, 1339, 969, 1125, 1666, 1625, 1488, 1739, 940]. phosphoryl [1980]. Photo [1588, 2253]. photo-identification [2253]. Photo-thermal [1588]. Photoautotrophic [1858]. Photobacterium [2415, 2102]. photodynamic [2185]. photolabile [1092]. Photoperiod [1826, 1313, 220, 2431, 2467, 254, 260, 408, 1014, 997]. Photoperiods [237, 1566, 462, 2362]. photosynthesis [1236, 2223]. phototrophic [1902]. Phoxinus [2324, 2368]. Phragmites [654]. Phycidae [629]. phycobiliprotein [2542]. Phylichthyid [2530]. phyllanthin [1417]. phyllosoma [426, 434]. phyllosomata [696]. Phylogenetic [889]. Physical [227, 1738, 61, 1395, 1053]. physicochemical [1961]. Physiological [2553, 2708, 1485, 837, 2713, 639, 1568, 779, 1314, 1818, 883, 2631, 1046, 2739, 1244, 1941, 2057, 2486, 371, 1234, 931, 1723, 2119, 2098, 2548, 120, 1607, 2368, 2265, 1021]. physiology [681, 1118, 1093]. phytase [1627, 942, 1672, 2608, 1258, 762]. phytase-producing [1258]. phytase-sprayed [762]. phytochemical [2552]. phytochemicals [2403]. phytoene [2678]. phytogenic [2560, 2455, 2399]. phytoplankton [1334, 1748, 719, 1006, 1460, 1236, 733, 823]. phytoremediated [2041]. Phytotherapy [2392]. Piaractus [2225, 1912, 2460, 1541, 1769, 1747, 1828, 2709, 1760, 2497, 2127, 2066, 2305, 2686, 2182]. pica [1955]. Piccolo [543]. piceus [2120]. Picochlorum [1614]. picolinate [604, 820]. Pierce [411, 625]. pig [398]. pigeons [2159]. pigment [2676, 1515, 1107, 1991]. pigmentation [1995, 196, 1077, 18, 327, 208, 1139, 1352]. pigments [1490, 784, 703, 2506]. Pike [493, 659, 1230, 2459, 630, 1933]. Pike-perch [493, 1230, 1933]. Pikeperch [1782, 480, 1932, 642, 749, 2267, 1229, 2338, 1588, 1944, 1160, 1499, 2521, 1927, 1934, 1940, 1941, 2057, 2563, 2196, 452, 1165, 1579, 1939, 1936, 1930, 1935, 641, 2099, 1159, 1931, 1161, 1942, 2456]. pili [2261]. pillars [2625]. Pillay [692]. Pilot [207, 977, 2185, 710, 2715]. pilot- [710]. Pilot-scale [207]. Pimelodidae [1097]. Pimelodus [2007]. pin [1847]. Pinctada [1281, 2246, 1186, 396, 954, 719, 1337, 1480, 1032, 401, 1506, 469, 1505, 1537]. pink [981, 1202, 305]. pinnatifida [1103]. Piper [1998]. PirA [2605, 2209]. PirAB [2690]. Pirarucu [643]. PirB [2605]. Pisces [629, 508]. piscicida [2238]. PIT [1273, 1053, 544, 1931]. PIT-tag-assisted [1273]. PIT-tagging [544]. pituitary [2142, 588, 1213, 1338, 1444]. Pixelar [2055]. pkmyt1 [2534]. Placopecten [281, 383]. plague [723]. plankton [854, 971, 1666, 653, 2718, 863, 864]. planktonic [1738]. Plant [2085, 222, 241, 2235, 2601, 2502, 410, 301, 1862, 1383, 1040, 1497, 2035, 1026, 2006, 1432, 2611, 1616, 1125, 1294]. plant-based [1497, 1026]. plant-protein-based [1125]. plantarum [2722, 2169, 1484, 1428, 1723, 1705, 2115]. plantings [967]. plants [2609, 2064, 2721, 1998, 1631, 1775, 2578, 2571, 1071]. Plasma [484, 1216, 161, 996, 291, 1996, 2328, 1828, 1542, 92, 379, 1432, 1857, 2634, 1931].

Plasmid [249]. platensis [1041, 2286, 1690, 681, 2084, 2436, 1311, 1473, 2506, 1978, 242]. plates [337]. Platichthys [1265, 1542]. platyfish [1995]. Platyhelminthes [2027]. Platymonas [1676]. play [2534]. players [2269]. plecoglossicida [1949]. pleopodal [406, 876]. Pleurotus [1846]. plicatilis [430, 1923, 1944, 621, 429, 1814, 2262]. ploidy [1096]. plot [1569]. plus [724, 36]. pluvialis [2733, 2210, 2432, 2040]. pod [1437]. Poecilia [663]. point [277, 2464]. Poland [1210, 596, 1154, 597]. policies [1520, 415, 1746]. Policy [2705, 2239, 2627, 414, 2615]. Pollution [554, 972]. Poly [1301, 1363]. polyactis [1908]. polyandry [1480]. polycaprolactone [1953, 1808]. polycaprolactone-packed [1808]. polychaete [1870, 540, 1336]. Polychaetes [539]. Polyculture [383, 2433, 1789, 2513, 2668, 398, 2672, 2570, 722, 845, 1495, 1714, 371, 2659, 1594, 2350, 1834, 971, 826, 1660]. polycultured [1888]. polycytidylic [2156]. Polydora [287]. Polydorid [560, 2528]. polyethylene [1825]. polygyny [1480]. polyherbal [2231]. polyhydroxyalkanoate [1953]. Polyinosinic [2156]. polymerase [2401, 2195]. polymorphic [86]. polymorphisms [1896, 2099]. Polynesia [1032]. Polyodon [454]. polypeptide [293, 1444]. Polyprion [509]. polyrhiza [451]. Polysaccharide [1403, 1878, 896, 2663, 985, 1517, 870, 1343, 1503]. polysaccharide-based [2663]. polysaccharide-degrading [896, 1503]. polysaccharides [1887, 2618, 2114, 1680, 2229, 2150, 2435]. Pomacanthus [1375]. Pompano [508, 1854, 2074]. Pond [176, 1378, 1099, 1578, 2235, 2475, 658, 1932, 2423, 2673, 1157, 660, 1168, 1575, 258, 94, 191, 598, 1792, 1587, 1584, 659, 56, 1151, 1580, 1002, 918, 754, 1941, 2057, 1585, 1802, 657, 2615, 1165, 1939, 2350, 2603, 1961, 19, 668, 854, 971, 826, 1167, 1424, 1451, 1530]. pond-based [2615]. pond-cultured [1941, 2057, 1165, 1939]. pond-raised [1802]. pond-reared [659, 19]. ponderosa [1485]. Ponds [250, 510, 1888, 457, 2475, 62, 2104, 1290, 795, 1960, 640, 1891, 191, 384, 1345, 1235, 830, 575, 1970, 321, 1382, 633, 2172, 1002, 756, 676, 1423, 2165, 2092, 1460, 666, 576, 602, 619, 95, 1279, 2596, 1104, 1594, 306, 654, 1130, 362, 292, 1199, 453, 653, 665, 2040, 1203, 823, 1494, 2718]. poor [2421]. Population [490, 1382, 554, 1755, 621, 2345, 1113, 1119, 155, 1743, 804, 2050, 1622, 1854, 455, 583, 1742, 845, 1154, 2462, 2021, 1761, 1756]. populational [1948, 1947]. populations [1826, 1136, 1565, 278, 1854, 723, 1382, 1753, 777, 1737, 596, 1506, 1177, 1964, 1933, 597, 397, 1938, 2179, 1922, 2120, 147, 2182]. Porgy [494, 220, 975, 1054, 334, 342, 1053]. pork [983]. porosity [1825]. Porphyra [895]. porphyran [2556]. Porphyridium [2679]. portable [2586]. Portieria [2683]. Portugal [391, 2062, 92]. portunid [927]. Portunidae [934]. Portunus [1457, 935, 1704, 2098, 936, 1060, 1618, 2431, 1260]. Positioning [1410]. positions [255]. Positive [2745, 2162]. possibilities [2]. Possible [143, 2635, 2741, 2107, 1064, 872, 1385, 634, 1987, 2412, 2683]. Possjet [126]. Post [2266, 707, 1179, 1939, 858, 1129, 1673, 472, 867, 2532, 31, 49, 1314, 1766, 2111, 1313, 975, 2218, 1499, 1929, 1614, 1435, 2531, 210, 2588, 1877, 1389, 1570, 1483, 1219, 1759, 2711, 557]. post-challenge [2111]. Post-harvest [2266]. post-larvae [1673, 1766, 1499, 1929, 1614, 1435, 2531, 210, 1877, 1389, 1570, 1759]. post-larval [1129, 867]. Post-mortem [1179, 975]. post-nursery [1483]. post-smolt [31, 49, 1313]. post-smolts [1314]. Post-spawning [1939]. post-thaw [1219]. Post-transport [707]. Post-tsunami [858]. postbiotics [2723]. postlarvae [69, 1085, 760, 106, 2090, 1567, 2094, 1421, 2522, 647, 524, 1145, 331, 1302, 1367]. postlarval [792, 981, 311]. postmetamorphic [79]. postprandial [1447, 1432]. potassium [2434, 1733]. potato [2011]. potency [2515]. potent [2013, 2105]. Potential [1603, 2643, 1168, 1068, 236, 683, 1024, 2350, 1276, 2255, 2199, 2722, 1801, 1075, 1295, 276, 1611, 1375, 638, 1490, 1529, 2467, 2512, 2532, 1894, 632, 1829, 2170, 1323, 872, 2416, 1241, 2493, 2202, 1521, 1080, 3, 1308, 2526, 1560, 2721, 55, 2400, 1145, 2185, 1386, 2596, 800, 1332, 2292, 1144, 1797, 2592, 2150, 1842, 210, 2028, 2010, 2223, 1125, 1705, 1938, 2444, 2678, 893, 1718, 2179, 2665, 1051, 2375, 1530]. potentialities [1899]. potentially [1870, 94, 1429]. potentials [1551]. potentiating [1501]. poultry [1018, 2218, 2521, 961]. poultry-based [2521]. poverty [683]. Povidone [2500]. Povidone-iodine [2500]. powder [1795, 2698, 2074, 1727]. Powersim [2454]. Powersim-simulated [2454]. pp [564, 565, 563, 678, 1008, 1975]. PPAR [1729]. Practical [2576, 1839, 1340, 1008, 1473, 1181, 1420, 555, 1267, 1832, 1019, 1036, 2149]. Practice [320]. Practices [774, 2517, 2376, 2327, 333, 2141, 2350, 1788, 1051, 692]. Pradesh [2668]. Prawn [366, 511, 848, 2109, 69, 2389, 1971, 1839, 1627, 2430, 1709, 1724, 1345, 1563, 2522, 1792, 2614, 1257, 2447, 1048, 1137, 2249, 1126, 1278, 1292, 1385, 2092, 2097, 1823, 78, 1388, 1685, 845, 2675, 1777, 1082, 2689, 1252, 963, 1898, 210, 1376, 1236, 1785, 19, 952, 293, 847, 2247, 1430, 1807, 2711, 466]. prawn-carp-rice [963]. Prawns [466, 374, 249, 1039, 67, 2425, 2665]. Praziquantel [2308, 2254]. Pre [226, 1313, 1934, 355, 532, 685]. pre- [1313]. Pre-feeding [226, 355]. pre-gelatinized [685]. pre-seasonally [1934]. pre-slaughter [532]. Prebiotic [1759, 1901, 2745]. prebiotics [2067, 1465, 2708, 2047, 601, 2330, 1436]. Precision [2001]. precocious [195]. precocity [929, 932]. Predation [128, 2359, 457, 1024, 469, 612]. predator [269]. predators [303]. predicted [1720]. Predicting [1984]. Prediction [966, 998, 2661]. Predictive [1643]. predictors [123]. predominant [2300]. Preface [651, 478, 530, 582, 1206, 1573, 1153, 1088, 422, 1926]. preference [4, 2465, 2211]. preferences [437, 2068, 1062]. pregrowing [1712]. prejudice [1980]. Preliminary [109, 1323, 2522, 2356, 794, 2407, 177, 246, 445, 1454, 508, 2204, 1262, 1529, 113, 2493, 917, 342, 1357, 800, 1051]. premium [1985]. premix [2552]. prenanti [1729, 1604]. preparation [1934, 2510]. preparations [1215, 2408, 1809]. prepared [731, 2410]. presence [1240, 1304, 2690, 1791, 1602]. present [2067, 723, 1523, 1802, 2306, 599]. presentation [434]. presented [803].

Preservation [335, 1212]. preserved [427]. Press [565, 563]. Pressure [1950, 960, 2320, 136]. pressures [1746]. presumptive [2576]. pretreatment [762]. Prevalence [730, 1904, 2342, 1909, 519, 2093]. prevent [2490, 2565]. Preventing [2154, 1809]. Prevention [700, 2290, 2009, 2633, 1796, 2006, 400, 2561, 395, 2520]. preventive [1974]. Prey [511, 1166, 438, 458, 842, 924, 1399, 1771, 1312, 2580, 434, 1455, 2211]. Price [564, 552, 565, 554, 563, 553, 23, 1513, 1985, 1057]. prickly [1595]. primarily [1640]. Primary [2146, 649, 384, 1810, 1107, 653]. primers [965]. Principles [692, 1087, 1204]. print [1600]. private [32]. Pro [1158]. Proales [1798]. probiont [1621]. probionts [647, 775]. Probiotic [1652, 507, 1048, 1484, 1705, 2063, 2722, 2109, 2029, 1894, 1821, 987, 2045, 1901, 2024, 1878, 2515, 2641, 2314, 1757, 1822, 1560, 1814, 465, 1873, 2105, 1866, 2412, 1069, 2330, 1089, 2089, 892, 2219, 2375, 1759]. probiotics [2439, 2067, 1708, 2474, 2336, 1122, 1377, 1829, 1068, 1436, 2017, 1263, 2526, 2047, 1460, 2377, 1879, 2694, 1145, 1646, 2361, 2634, 1988, 1731]. problem [2028]. problems [40]. Probopyrus [2553]. Procamallanus [2028]. Procambarus [842, 1370, 1468, 2598, 2365, 1625, 2628, 2745, 2720, 1148, 1540]. Procedure [569, 872, 965, 1718]. procedures [1589]. process [2078, 2558, 1825, 1858]. processed [2013]. Processing [690, 229, 2081, 1813, 998, 1150, 2506, 1577, 891, 2728, 1625]. prochilod [2481]. Prochilodus [2481]. produce [2558]. produced [2062, 2279, 633, 199, 2505, 1492, 1435, 2506, 1389, 1955, 2283, 1951]. producers [2364]. produces [1524]. producing [385, 1258, 107, 2105]. product [1838, 2317, 1579, 961, 844, 1577]. Production [2026, 129, 1298, 189, 575, 388, 259, 503, 1814, 2103, 952, 516, 32, 1888, 1015, 69, 1058, 1129, 1838, 1804, 2007, 2701, 1018, 1520, 561, 1690, 1253, 1836, 2078, 1100, 45, 1954, 1899, 1377, 2366, 2267, 2673, 207, 2581, 2085, 1923, 2626, 2752, 1782, 1419, 1532, 2468, 660, 1724, 271, 2198, 927, 876, 1117, 1587, 962, 1745, 623, 1066, 2202, 8, 1152, 1596, 2355, 2052, 1864, 1458, 1349, 1768, 1045, 2210, 1585, 2429, 1850, 1519, 352, 2541, 380, 2615, 2527, 2261, 1034, 2659, 1760, 2166, 1579, 1946, 1945, 655, 1378, 1559, 2682, 2603, 442, 1276, 1376, 2479]. production [2243, 2507, 763, 2295, 1107, 133, 292, 834, 973, 1710, 665, 2012, 2480, 1462, 2040, 1203, 1807, 2337, 1260, 1592, 1195, 1149, 1689, 851, 2208, 1355, 1512, 1831, 1151]. productive [2298, 1081, 2068, 1191, 2438]. Productivity [1427, 2454, 1772, 2260, 2258, 191, 384, 1647, 1045, 2141, 2350, 292, 653]. Products [563, 1295, 1950, 243, 1813, 2202, 2583, 2193, 1582, 1447, 1386, 2006, 362, 2301, 2669, 1203, 2077, 2061]. proficiency [2161]. profile [2062, 2605, 784, 1202, 1173, 1784, 1141, 2354, 2256, 912, 2505, 501, 983, 1717, 464, 1096, 2588, 1653, 959, 1185, 2160, 293, 2044, 2616, 1662, 2087, 1405]. profiles [2173, 2414, 1098, 1568, 2128, 899, 1028, 1421, 1552, 2631, 879, 1767, 1432, 2383, 1694, 2039]. profiling [856, 2380, 1872, 2349, 1504, 1665, 1527]. profit [2240, 1074]. Profitability [2668, 80, 888, 886, 1004, 1298, 2417, 2243, 2391, 1674]. profitable [1917]. progeny [1156]. program [1119, 2535, 808, 1327]. programmes [1408]. programming [2700]. programs [1557, 1487]. Progress [1510, 1946, 1945, 2575]. Progressive [134]. progressively [1326, 67]. project [138]. prolifera [1403]. proliferation [2016]. proliferative [2636]. prolonged [146, 907, 8]. promise [467]. Promising [503, 1212, 1771, 2392, 1710]. promote [313, 615, 415]. promoted [1701]. promoters [2525]. promotes [1887, 993, 2480, 2238]. promoting [1959, 96]. Promotion [1413, 145, 683]. prone [2052]. Propagated [567]. propagation [875, 1616, 1159]. proper [880]. properties [1821, 2641, 2386, 2013, 1998, 2663, 1961, 2105, 2189, 557, 1105, 1840]. property [2681]. Prophylactic [2653, 2093]. propionate [2511, 2180, 2005]. Propiscin [578]. propolis [1049, 2197]. proposal [534]. proposals [2239]. proposed [2374]. propyl [2387]. prospective [1608]. Prospects [238, 2272, 599]. protease [410, 160, 1872]. proteases [301, 2743]. protection [1990, 987, 1563, 2049, 1333, 1059, 1962, 1403, 1056]. Protective [632, 1925, 2127, 1801, 2577, 2539, 2333]. protects [1732]. Protein [248, 1781, 1181, 1420, 1033, 344, 2617, 2389, 174, 1627, 1286, 2190, 2054, 942, 149, 909, 2293, 1017, 1678, 2427, 1293, 996, 2085, 2558, 1081, 1226, 291, 1779, 1340, 1640, 2354, 1307, 2416, 824, 623, 1706, 115, 114, 2291, 1852, 1486, 1040, 1442, 2282, 2521, 853, 1765, 2707, 555, 1267, 1453, 2097, 2334, 1045, 1796, 2079, 1019, 2139, 880, 379, 1091, 922, 1873, 1138, 1271, 2305, 1044, 1036, 387, 1125, 2098, 2333, 2373, 1711, 1128, 611, 2652, 1012, 1148, 2420, 1038, 1191, 1540, 1752, 2113, 2318, 1942, 2456, 2606, 2568]. protein-fed [1271]. protein-like [2097]. protein-to-carbohydrate [1779]. protein-to-energy [2568]. protein/carbohydrate [611]. proteinases [943]. Proteins [222, 1801, 1440, 168, 1895, 2035, 1432, 1331, 2247, 2091]. proteolytic [1715, 1091]. Proteomic [2477, 2164, 2234]. protexin [2219]. protocol [1247, 1073, 389, 744]. protocols [198, 1548, 2503, 2695, 537]. Protozoan [247, 830]. provide [2688]. provides [1274, 2412]. Province [1786, 1787, 1909]. proving [829]. provokes [1623]. Proximate [1803, 146, 2455, 2399, 1202, 2508, 1767, 61, 2383]. proxy [1402]. Psetta [1255, 947, 500, 1122, 601, 1271]. pseudo [339]. pseudo-green [339]. Pseudoalteromonas [1814]. Pseudobagrus [1716]. Pseudodiaptomus [1354]. Pseudomonas [792, 2342, 2095, 2161, 2684, 2333, 1182, 1949, 1883]. Pseudoplatystoma [1750, 1779, 1097, 2151, 1751, 2123, 2076]. Pseudosciaena [674]. Pseudotropheus [1077]. Psidium [1998]. PSO [1915]. Psoralea [472]. PTCC [1879]. Pteria [1703, 566]. Pterophyllum [1894, 2312, 2309]. Pterophylum [767]. Public [1865, 1375, 2627, 414, 365]. Publisher [308]. Publishers [678]. Publishing [564, 552, 565, 1975, 553]. publishing.sales [565]. PUFA [433]. Puffer [245, 496, 435, 182, 939]. Puget [1853]. pulcherrimus [1633]. pullastra [772]. pumpkin [1989]. punctata [2741, 1437]. punctatus [1615, 1802, 1273, 1702, 1682, 1038, 1405, 1981]. Pungent [2216]. Punta [1306]. puntazzo [424, 506, 1465, 888, 886, 919, 372, 1422, 983, 505, 537]. Puntigrus [2253]. Puntius [2068]. purchasing [2583]. Purification [2676, 1529]. Purified [245]. purple [2410, 1734, 1329, 2055]. purpose [326]. purpuratus [2410, 728, 605, 402, 259, 348, 303, 284]. pusillum [2215]. putative [1684]. putrefaciens [1089]. puzzling [2734]. pygmaea [1719]. Pyocyanin [2451, 1639]. pyrenoidosa [2078, 724]. Pyrethroid [570]. pyriformis [2324]. Pyropia [1913, 2187, 2717, 2066].

Q. [736]. Qihe [2587]. qPCR [2484]. qRT [2309]. qRT-PCR [2309]. QTL [1201, 2146, 1352]. quadrangularis [1252]. quadricarinatus [2083, 822, 2372, 865, 1794]. Quagrainie [689]. quahog [1749, 1726]. qualifying [2157]. qualitatively [321]. Quality [2660, 66, 690, 579, 511, 239, 482, 229, 513, 1489, 491, 2370, 1578, 1196, 2517, 1870, 2287, 1248, 1334, 764, 1817, 270, 1215, 1216, 198, 1748, 2336, 1005, 970, 1740, 398, 423, 2240, 2581, 2304, 1326, 1202, 2725, 1419, 390, 2638, 59, 729, 132, 2084, 1996, 998, 1541, 1235, 1584, 56, 994, 1080, 2183, 58, 756, 1278, 1480, 1598, 1423, 2672, 2449, 1927, 1934, 2516, 1497, 1685, 333, 576, 2121, 2316, 2400, 1741, 2694, 2747, 1150, 1983, 1601, 534, 2270, 1658, 1783, 445, 532, 1579, 1376, 547, 840, 978, 1516, 292, 1302, 100, 2289, 668, 1577, 1979]. quality [1389, 1666, 2301, 1876, 1128, 1366, 2698, 2432, 697, 2549, 2720, 1988, 36, 863, 864]. qualityindicators [294]. Quantification [1634, 428, 2519, 5]. Quantifying [267]. Quantitative [1201, 1324, 1363, 2269, 2331, 2055, 1373]. Quantity [588, 1215]. Quasipaa [2221]. Québec [138]. queen [1720, 2241, 1892, 103]. quercetin [797]. quercetin-containing [797]. Quick [2401].

R [129, 296, 330, 365, 447, 565, 688, 692, 736, 862, 956, 1039, 446, 689, 553]. R. [726, 678, 862, 1072]. rabbitfish [702]. race [839]. raceway [2165, 1131, 44, 154, 2040]. Rachycentron [1037, 2279, 1511, 962, 2378, 765, 913, 1113, 961, 1512]. rack [1298]. RAD [2030]. radiation [381, 2658]. radiation-induced [2658]. radiography [206]. Radix [2612]. Rafinesque [677]. raft [1298, 1547]. raft-cultured [1547]. raft-string [1298]. Rainbow [1997, 566, 482, 229, 2722, 4, 1295, 1789, 1133, 444, 1445, 212, 2358, 1603, 64, 266, 1568, 2128, 134, 2299, 196, 1047, 1141, 1214, 1311, 1470, 2226, 2328, 604, 1916, 1469, 409, 11, 1920, 322, 37, 2334, 10, 2644, 606, 1982, 96, 328, 1065, 1026, 2457, 1713, 1846, 47, 1658, 1783, 1476, 1697, 312, 644, 766, 884, 30, 1096, 343, 1432, 211, 1873, 1163, 820, 2746, 1978, 837, 2726, 1044, 1719, 648, 1758, 1656, 762, 327, 1662, 1387, 2222, 940, 1360, 2011, 2204, 1139, 2685]. raised [153, 1011, 2367, 2533, 1615, 1802, 1902]. ramada [2023]. ranching [1024, 307]. randomly [86]. ranga [2478]. range [2405, 1404]. Rao [1284]. Rapa [1698]. Rapana [1698]. RAPD [84, 86]. rapeseed [1044, 1733]. Rapid [1583, 1835, 2358, 56, 2484, 559]. rare [1500]. RAS [2360, 2451, 2734, 2638, 1849, 2416, 2256, 1832, 2551, 1639, 2640, 2201].

RAS-based [1849]. Rate [482, 2617, 2142, 2604, 1838, 1281, 2344, 1342, 603, 924, 1958, 1629, 633, 993, 1094, 163, 148, 1337, 1201, 1819, 935, 190, 186, 342, 268, 1495, 2041, 1758, 826, 1225, 2683, 1938, 2465, 1794, 2431, 873, 2453]. Rates [500, 221, 509, 510, 443, 1716, 31, 1555, 281, 1383, 399, 958, 1725, 1765, 303, 1024, 1297, 1475, 1893, 2537, 289, 1935, 2454, 1682, 2211]. ratio [492, 1134, 1627, 1628, 2140, 2381, 409, 1181, 1685, 2233, 2537, 133, 387, 611, 1398, 2568]. Ration [1338, 817, 30, 378, 556, 89]. rations [749]. ratios [2083, 909, 899, 1779, 944, 2355, 2449, 2121, 685]. rats [2680, 2619]. RAU2C [2682]. raw [1440, 2738, 685]. RAW264.7 [2556]. Ray [678, 887]. RBR [2442]. RC [887]. rDNA [895]. re [146, 1960, 644, 1302, 1012]. re-alimentation [1012]. re-circulated [644]. re-circulating [1302]. re-feeding [146]. re-oxygenation [1960]. reaction [2195]. reactions [2687, 2754, 1236]. reactive [1852]. reactor [2374]. reactors [1968]. ready [1024]. real [1479, 2702, 1655]. real-time [1479, 2702, 1655]. rear [2101]. Reared [493, 482, 2242, 519, 93, 749, 299, 2559, 9, 70, 548, 909, 934, 1173, 1567, 2094, 302, 1992, 1210, 1259, 1381, 1379, 659, 462, 2631, 1654, 171, 305, 399, 2459, 631, 2739, 1324, 1940, 1297, 2514, 1695, 1893, 2079, 2508, 1983, 1928, 937, 2497, 294, 445, 644, 19, 61, 431, 1036, 787, 2116, 397, 589, 2145, 2549, 914, 1799, 592, 1159, 2646, 2201, 2510]. rearedgreat [279]. Rearing [796, 1336, 122, 515, 1753, 494, 505, 246, 516, 1245, 68, 109, 192, 687, 340, 867, 396, 1676, 314, 1052, 833, 1534, 698, 912, 1944, 677, 1771, 1422, 1094, 1046, 2343, 2533, 613, 341, 1630, 2524, 1143, 1777, 1983, 339, 371, 1658, 1783, 1562, 2697, 806, 1924, 1830, 626, 1365, 972, 2425, 135, 2300, 1876, 741, 2021, 1807, 999]. reason [1166]. Reassessing [2574]. reassessment [378]. Receptivity [1418]. receptor [1000, 2323, 2469]. receptors [2741, 2362]. recipient [1979, 1167]. Reciprocal [585]. reciprocating [520, 521]. recirculated [371]. recirculating [2360, 2451, 192, 1286, 1463, 548, 827, 1001, 640, 2638, 1862, 1989, 2367, 2614, 1564, 2459, 631, 1308, 1953, 1808, 1497, 2022, 1631, 1495, 1893, 1686, 1825, 2551, 710, 1639, 1937, 1365, 41, 1426, 1877, 1968, 1935, 2640, 592, 1689, 1988, 2201]. Recirculation [1526, 947, 470, 2298, 2256, 1861, 973, 1462, 1355, 2453]. Recombinant [1732, 2691, 2608, 1711]. recombinase [2401]. Recommended [1632]. reconstruction [1212]. record [2160]. recording [57]. recovered [2644, 2746]. recovery [2604, 707, 2446, 281, 1135, 2042, 2428, 1101, 96, 2196, 2058, 2037, 761, 2167, 2629]. recreational [1002]. Recruitment [951]. recrystallized [1921]. rectal [1387]. recycling [2715]. Red [2237, 1897, 494, 220, 1772, 562, 1085, 1730, 1679, 2653, 1906, 2225, 2111, 1641, 2589, 1364, 975, 675, 2371, 1481, 1054, 839, 1468, 2598, 1471, 334, 722, 186, 342, 269, 2400, 2482, 1631, 1363, 1034, 1760, 1300, 378, 906, 2295, 1719, 1053, 2683, 1625, 2339, 2749, 2745, 1794, 1148, 2011, 1540, 1728, 1237]. red-bellied [2225]. red-drum [269]. redclaw [2083]. redivivus [921]. redlip [1425]. redtail [80]. reduce [313, 632, 2684, 647, 868, 1939]. reduced [1367]. reduces [71, 2593, 1818]. reducing [2722, 1092, 2397]. reduction [381, 824, 683, 960, 1308]. reductions [2000]. reed [654]. reef [809, 2174].

Reeve [439]. refeeding [905]. referees [217, 264]. reference [562, 871, 2086, 2009, 2013, 2177, 1961, 1866]. reforms [2599]. refrigerated [1920, 1576, 294, 445]. refuge [854]. regardless [2525]. regeneration [2589, 720]. regime [1899, 1632, 171, 2226, 2233, 775, 1092, 1225, 1161]. regimen [431]. Regimes [621, 220, 45, 1874, 643, 752, 677, 575, 633, 631, 756, 334, 576, 76, 19, 164, 607, 1149]. Region [1306, 2475, 1793, 1549, 2597, 1964, 2306, 2038, 2644]. Regional [2657, 962, 978]. regions [1385, 1693, 911]. regius [1017, 445, 2157]. regression [966, 1550]. regulate [2238]. regulating [2323, 2113]. Regulation [2137, 2581, 2741, 2732, 2447, 976, 2247]. regulators [2085]. Regulatory [2418, 2647]. rehabilitation [858]. Reilly [365]. reineckii [2601]. Reinertsen [274]. rejections [2086]. related [2370, 1384, 2114, 2313, 2434, 2107, 2638, 1083, 662, 1351, 2712, 2637, 2328, 1468, 1914, 2511, 1484, 1879, 2482, 2106, 1411, 1163, 2372, 2000, 1824, 1877, 2373, 2018, 1707, 2435, 2556, 2222, 2168, 2146, 1613, 2130]. Relation [373, 588, 106, 769, 838, 1580, 766, 1798, 1653, 556]. Relationship [823, 2228, 460, 273, 754, 1385, 1050, 2073]. Relationships [487, 103, 25, 86, 675, 58, 1292, 75, 602, 619]. relative [198, 2386, 397]. relaxing [2615]. Release [288, 1547, 384, 195, 1146, 2374]. released [1737]. releasing [255]. relevance [1577, 2182]. Reliable [505]. remains [457]. remediation [1193, 2292]. remotely [790]. Removal [1492, 230, 1772, 1303, 1555, 1006, 1249, 1808, 1847, 1493, 2724]. remove [1061]. removing [1989, 1198]. rendalli [1307]. renoprotective [2619]. Reovirus [2555, 2179, 2288]. reoxygenation [2271]. repeat [777]. repeated [759]. repeats [1544]. repens [2601]. repetitive [173]. replaced [2492, 1040, 1686]. Replacement [68, 2388, 71, 222, 846, 2521, 1691, 961, 1180, 1839, 291, 1188, 1776, 956, 696, 114, 1473, 1420, 2473, 2035, 983, 1019, 1478, 2531, 2588, 1044, 2098, 920]. replacer [2371]. replaces [284]. replacing [1900, 940, 2318]. Repopulation [169]. report [636, 800, 2746, 116]. Reproduce [479]. Reproduction [397, 368, 1354, 1220, 2467, 2050, 1224, 1941, 2057, 1177, 1842, 2372, 2665, 1161]. Reproductive [1194, 713, 1306, 423, 2027, 28, 283, 2617, 444, 1358, 1216, 638, 2741, 1213, 1826, 1173, 786, 184, 1077, 1588, 2447, 1473, 2055, 1096, 753, 930, 1025, 1795, 208]. Republic [2475, 457, 1582]. required [1390]. Requirement [245, 461, 671, 1116, 1353, 2303, 1781, 2426, 1324, 881, 2737, 267, 2072, 1653, 1443, 1322, 1038, 1191, 1398, 2265, 1942, 2456, 2606]. Requirements [994, 353, 1286, 1654, 3, 1055, 2249, 112, 432, 1976]. rerio [2656, 1147, 1184]. Research [690, 1055, 1205, 74, 1693, 686, 2573, 2716, 349, 1491, 1510, 859, 1285, 1523, 929, 2353, 1544]. reseeded [113]. Reserve [605, 805]. reserves [802, 417]. Reservoir [1746, 2648, 452]. reservoirs [977, 55, 320, 326, 147]. residue [2450]. residues [1393, 755, 891]. resistance [1812, 2109, 1972, 693, 1049, 946, 1384, 2430, 2618, 2297, 1346, 2342, 2095, 2437, 2593, 2319, 1676, 1474, 1818, 1538, 1901, 1999, 1439, 1137, 1914, 2526, 1757, 2717, 2727, 359, 2220, 1684, 1388, 2316, 2694, 1145, 767, 2151, 1773, 2306, 1082, 465, 2157, 2647, 1010, 1723, 1348, 2470, 1437, 1949, 2361, 985, 2179, 2655, 2247, 2152, 2500, 2307, 1727, 761, 1405, 2143, 2115, 2245, 1731, 2201]. resistant [950, 1373, 2234]. resists [2261]. resolution [1717]. resource [2668, 80]. Resources [1973, 1212, 1008, 176, 1911, 803, 812, 789]. Respiratory [593, 1981]. respirometer [1101]. Response [1950, 573, 1112, 1543, 2370, 2729, 953, 1763, 1446, 1133, 2137, 2344, 629, 2559, 1140, 2656, 1603, 2078, 2430, 2579, 1238, 2025, 1764, 2732, 2232, 1227, 924, 1314, 258, 2148, 1818, 1538, 818, 1640, 1901, 1467, 2738, 2564, 1642, 2276, 2691, 22, 2226, 918, 1126, 1054, 855, 2032, 1822, 1977, 2365, 359, 2580, 1684, 2203, 1675, 1388, 2329, 2035, 173, 2346, 2491, 2049, 1082, 2689, 1287, 931, 2603, 2630, 1432, 2058, 2037, 2181, 704, 1504, 2000, 2577, 2539, 2271, 2100, 750, 672, 1877, 959, 2470, 2520, 2723, 1867, 1438, 2655, 2749, 120, 593, 612, 1665, 2368, 1871, 761, 1540, 1922]. response [2685]. Responses [2063, 321, 1995, 1972, 1049, 1261, 2003, 1309, 2297, 2437, 1170, 1229, 1680, 2538, 2572, 1474, 2231, 1827, 1589, 2587, 960, 2428, 1244, 1806, 2526, 2008, 1288, 334, 2694, 2486, 1800, 1773, 2497, 2123, 2262, 1333, 1010, 1496, 1316, 2333, 985, 2495, 2548, 1403, 1794, 1430, 1517, 2646, 1021, 1500, 2686, 2629, 554]. Responsible [446]. restaurant [2492, 2204]. resting [129, 1534]. restitution [1753]. restoration [830, 810, 808, 803, 967]. restores [284]. restricted [2352, 1591, 2044]. restriction [1768, 1338, 1012]. restrictive [334]. restrictively [882]. resulted [2516]. Resulting [511]. Results [487, 455, 246, 109, 698, 342, 445, 1199]. Retention [1146, 64, 196, 986, 1337, 1453, 1713, 2264, 1682, 1038, 1191]. reticulata [663]. reticulatum [1750, 1779, 1097, 1751, 2076]. retinopathy [1797]. Retracted [1467, 1385]. retrieved [2380, 2089]. Retrospective [1969]. return [575]. returns [673]. reuse [638]. reusing [1889]. reutilisation [1168]. revealed [2237, 1896, 2493, 777, 1761, 1728]. Revealing [2044]. reveals [2221, 2227, 1969, 2271, 2397, 2004]. reversal [1679, 822, 1759]. Reverse [2612, 2689]. Reverse-gavage [2612]. reversed [1214, 1508]. reversion [1320]. Review [244, 474, 466, 498, 329, 411, 307, 1115, 726, 296, 297, 330, 365, 366, 353, 354, 449, 447, 608, 609, 650, 690, 691, 688, 679, 678, 692, 708, 700, 727, 716, 709, 736, 774, 785, 860, 861, 859, 863, 862, 864, 1009, 957, 1008, 992, 956, 991, 955, 974, 1070, 1071, 1039, 1056, 1072, 1151, 1152, 1204, 1284, 1285, 1283, 1282, 1436, 1837, 1905, 1975, 448, 274, 446, 625, 374, 689, 725, 680, 2540, 1099, 2059, 2439, 2067, 2001, 2125, 53, 2573, 2716, 349, 98, 564, 2474, 1899, 2398, 1346, 2651, 2494, 2643, 2626, 2752, 598]. review [2609, 2743, 416, 927, 2071, 1878, 1235, 2202, 2392, 3, 565, 563, 2064, 2671, 1464, 1974, 2750, 2721, 2443, 2252, 1386, 1811, 616, 2250, 1574, 2150, 2236, 2422, 2010, 2255, 1577, 2550, 2504, 188]. Reviews [223, 253, 527, 552, 554, 553, 21, 34, 43, 72, 117, 141, 152, 213, 179, 194, 261, 27, 51]. revitalise [2602]. revival [2272]. revolution [848, 963, 411]. reward [890]. Rhabdosargus [1660]. rhabdovirus [26]. Rhizophora [622]. Rhodomonas [2503]. Rhodophyta [895, 717, 610]. rhombus [1090]. Rhopilema [686]. rhubarb [1517]. Rhynchocypris [2375]. rhythm [741]. Rhythms [220, 1093]. Ría [103]. ribbed [1174]. Rica [2306]. rice [1005, 256, 1267, 845, 1019, 963, 1898, 2340, 2382, 848]. ricefields [1074]. rich [1897]. Richardson [750, 2130]. Ricin [2275]. Riconada [605]. ridge [304]. ridgetail [2097]. rigida [1356, 1614, 1535, 514]. ring [871, 2442]. RING-between-RING [2442]. ringens [2419]. Ringø [1436]. ripening [2523]. Rise [1070, 156]. risk [2413, 2110, 2052, 2327, 2359, 1787, 2243, 1656, 665]. risk-prone [2052]. risks [2327]. Rissik [863, 864]. Risso [495, 694]. River [466, 159, 278, 1246, 271, 2447, 988, 1753, 722, 1823, 2665, 1807, 1971, 1709, 2120]. Rivers [489]. rivulatus [702]. RMCP [2639]. RNA [2332, 1705, 2665, 1398]. RNA/DNA [1398]. road [1416]. Roberts [1072]. Rock [612, 426, 1952, 1369, 1363, 434]. rockfish [2638, 2538, 1135, 1297, 742]. Röding [401]. rohita [953, 1269, 2207, 1321, 1851, 1262, 878, 1100, 908, 1474, 2075, 2428, 1123, 1684, 2072, 1258, 1667, 2611, 2562, 1338, 2470, 2600]. Rohu [221, 241, 953, 2207, 1321, 1262, 878, 1123, 428, 1258, 197, 2611, 1333, 2562, 2470]. Role [2439, 2009, 2491, 241, 2323, 1569, 2474, 2741, 1344, 2017, 984, 1982, 2577, 2539, 413, 2099, 2683, 2089]. Roles [2247, 2670, 2647, 2534]. ROM [297]. root [2386, 1437]. rooting [1616]. rope [791]. ROS [2262]. rose [1353, 2697]. roselle [1991]. rosemary [1882, 1920]. rosenbergii [845, 1010, 2109, 1801, 2612, 2389, 1839, 1627, 2610, 270, 2430, 2304, 2558, 1724, 1563, 2522, 1896, 2553, 1792, 2614, 1257, 1048, 1137, 1126, 1292, 1385, 1288, 1388, 1685, 1687, 2675, 1082, 2139, 2689, 1252, 1371, 2332, 1785, 1010, 2425, 293, 2247, 2711, 2510, 366]. Rosmarinus [1882, 1920]. Ross [785]. rotating [2614]. rotifer [129, 1923, 429, 1814, 775, 1330, 36]. Rotifers [572, 430, 546, 1944, 157, 298, 1084, 2551, 17, 2262]. rotten [1272]. rotundiformis [2370]. roundworms [2225]. RPA [2401]. RPA-LFD [2401]. RSIV [1363]. RSM [2682]. RT [1476]. RTG [2679]. RTG-2 [2679]. rubens [1024]. RUBISCO [895]. rubra [459]. Rubrivivax [1912, 2084]. ruckeri [2471, 1916, 343, 1727]. rude [1371]. Ruditapes [1843, 772, 871, 705, 771, 1429, 1630, 829]. rufescens [1085, 1034]. rupestis [635]. rupestris [1572, 395]. rural [640, 2134, 2476, 354]. Russia [126]. Russian [1848]. Rusten [708]. ruthenus [1328, 2212]. Rutilus [1399, 2747, 964]. Ryan [2385].

S [552, 365, 449, 679, 700, 709, 774, 1071, 1975, 448, 2714, 2733, 1885]. S. [1204, 720, 1233]. SA [1493]. Sabaki [2217]. sabdariffa [1991]. Sabella [540, 1336]. sac [1418, 1208]. Saccharina [1335, 2717]. Saccharomyces [1812, 1033, 2084, 465, 2262, 1731]. Saccostrea [1952]. Sacha [2371]. Sæbø [690]. safe [2451, 1639, 2578, 2571]. safensis [2641]. Safety [529, 690, 1150, 2409]. safflower [1255]. saffron [2452]. Sagay [805]. Saharan [683]. Salaam [911]. salar [2544, 1621, 635, 890, 942, 989, 779, 1715, 1013, 2637, 361, 359, 1073, 1434, 547, 1847, 2724]. Salicornia [1200, 1001]. salicylic [2143]. salina [2713, 1446, 1954, 781, 2383, 2678]. saline [1829, 623, 2631, 879, 2739, 2463, 971]. saline-alkaline [2463]. Salinities [240, 2532, 1567, 2343, 2596, 1404]. Salinity [205, 471, 2739, 1907, 2654, 1037, 549, 2344, 340, 934, 981, 2094, 949, 838, 1118, 1641, 1023, 897, 1654, 1094, 2249, 839, 1324, 2209, 186, 2233, 2013, 1952, 177, 1507, 2508, 181, 2491, 90, 1661, 1798, 1112, 2000, 2585, 1394, 1820, 2548, 2549, 2021, 1698, 1343, 745]. salinized [2585]. saliphilus [1970]. Salminus [1098]. Salmo [1621, 635, 574, 551, 890, 942, 1869, 989, 779, 1715, 1013, 2637, 1207, 2544, 2282, 361, 359, 2321, 1073, 1434, 547, 1847, 959, 2724, 611, 1727]. salmoides [2054, 1171, 2580, 1106]. Salmon [1621, 227, 228, 226, 313, 1180, 150, 255, 174, 635, 890, 16, 942, 989, 81, 155, 2651, 1566, 2303, 63, 779, 1715, 1993, 12, 1013, 31, 49, 1314, 195, 2637, 1985, 1313, 2544, 992, 171, 29, 123, 2583, 163, 148, 1522, 2156, 361, 359, 193, 1073, 46, 2527, 383, 28, 1434, 1967, 92, 547, 1987, 1847, 1591, 2724, 2420, 2710]. salmonicida [1169, 87, 395]. salmonid [145, 71, 87]. Salmonidae [2160]. salmonids [2302, 1895, 624]. salmonis [1655, 2724]. Salt [1180, 2740, 2187]. salting [1180]. salts [2180]. saltwater [25]. Salvelinus [481, 1882, 1233, 667, 2160, 1158, 1581]. Salvia [2618]. Sambucus [2390]. same [843]. samphire [1836]. Sample [755, 1273]. samples [1655]. sampling [2358, 11]. Sander [1932, 642, 749, 2267, 1229, 1782, 2338, 1588, 1944, 1169, 1230, 1160, 1499, 2521, 1927, 1934, 1940, 1941, 2057, 2563, 2196, 1165, 1579, 1939, 1936, 1930, 630, 1933, 1935, 641, 2099, 480, 1159, 1931, 1161, 1942, 2456]. sandwich [33]. Sanggou [2132, 1176]. sanguinea [949]. sanitary [1485]. sanitizer [2565]. santolla [1539]. Saponin [233, 1704]. Saprolegnia [1980, 2593, 2390, 2275]. Saprolegniasis [2457, 2593, 2565]. sarba [1660]. sarcoplasmic [168]. Sarcotaces [2530]. sardine [2354]. Sargassum [2552, 2543, 1501, 2229, 1650]. sargus [505, 505]. Sathirathai [552]. satiation [882]. sativa [2267, 1849]. sativum [950, 1727]. sativus [2452]. Saudi [2530, 133, 2217]. Sauvage [1972, 988, 2330, 711]. saxatilis [2107, 852]. Say [802]. SBSBR [793]. scabra [2310]. scalare [1894, 767, 2312, 2309]. scale [1917, 1129, 1673, 1001, 2366, 207, 640, 1923, 2499, 31, 2198, 1771, 1647, 756, 1864, 2345, 2417, 1325, 74, 710, 1378, 835, 2195, 2189, 2700, 2364, 1788, 847, 711, 2152, 1512]. scaled [371]. scales [2701, 1811]. scallion [1760]. Scallop [277, 516, 285, 286, 605, 109, 349, 276, 282, 106, 279, 707, 138, 260, 1136, 323, 278, 281, 1176, 1351, 2002, 259, 348, 104, 2704, 158, 613, 2429, 280, 102, 287, 332, 108, 101, 283, 105, 790, 111, 311, 103, 112, 100, 284, 967, 407, 580, 1438, 873, 1318, 680]. Scallops [234, 1129, 159, 310, 169, 113, 402, 1024, 2359, 356, 383, 126, 626, 127, 397, 612]. scaly [1219]. scanning [2380]. SCAR [849]. scarce [1906]. Scarfe [700]. scenarios [1549]. Scenedesmus [2091]. schedule [701, 10]. schedules [442, 2507]. scheme [788]. scherzeri [1392]. Schizochytrium [2388, 1545]. Schizothorax [1729, 1604]. Schlegel [759]. schlegeli [1135, 1297, 742]. schlegelii [2638, 2538]. schmitti [1393, 1355]. school [2519]. schrenckii [2618, 1857, 2265]. Schultze [1894]. Sciaena [2524]. Sciaenops [1906, 675].

Science [447, 678, 2398, 1283]. sciences [1413, 1414, 1412]. Scienti.c [554]. scientific [859, 1285, 646]. scientometric [2353]. scleractinian [781]. scope [402, 1118]. Scophthalmus [470, 1770, 1179, 1996, 1090, 2416, 677, 1977, 2320, 394, 1824, 1108, 748, 2307]. score [2417]. Scoring [1050, 2576, 1073]. Scortum [2461, 1466, 1744]. Scotland [791]. screen [1810, 2640]. Screening [2712, 1361, 1494, 1621, 2468, 1897, 2049, 2149]. scutellaria [2612]. Scylla [934, 1377, 545, 546, 423, 571, 1325, 1368, 1983, 937, 928, 1646, 933]. SDDV [2195]. SDS [160, 344]. SDS-page [344]. SE [88, 1838]. Sea [88, 507, 307, 1336, 1304, 247, 2637, 1102, 471, 126, 510, 1656, 2031, 519, 1245, 562, 788, 93, 2410, 62, 713, 901, 1708, 192, 882, 2310, 16, 561, 1611, 38, 161, 39, 1906, 1610, 138, 340, 970, 81, 1780, 160, 2471, 1419, 1622, 390, 2607, 1565, 888, 886, 919, 1004, 405, 1083, 2367, 2230, 281, 1774, 184, 1345, 2335, 912, 1401, 974, 1606, 123, 2197, 399, 1486, 167, 1481, 58, 1864, 984, 839, 889, 1308, 1423, 1681, 1274, 1193, 1024, 2055, 578, 606, 2022, 186, 2359, 165, 983, 380, 309, 1332, 339, 367, 383, 136, 1428, 2503]. sea [294, 1767, 2157, 1559, 1547, 1243, 156, 898, 1092, 1712, 2020, 1881, 844, 968, 1062, 557, 204, 2056, 386, 1366, 1003, 1192, 1299, 2441, 2440, 1403, 612, 697, 1483, 1607, 1517, 1660, 1633, 1280, 1318, 1373, 1494, 1650, 870, 1149, 1341, 2662, 2414, 500, 2523, 1612, 1782, 807, 1521, 1602, 1737, 2524, 1113, 806, 2038, 2710]. sea-based [1681]. sea-bass [2335]. Seabass [1193, 1645, 2673, 1766, 2442, 2343, 2329, 1691, 1881, 2674, 1868, 1175, 1543, 2005, 914]. seabirds [637]. Seabream [506, 815, 1286, 2133, 302, 1465, 372, 1422, 1349, 528, 770, 798, 1363, 464, 2496, 2180, 1089, 1175, 789, 1029, 685]. seafood [1950, 2583, 690]. seafood-exporting [1950]. seahorse [1358, 2042, 2449, 544]. seahorses [556]. searching [407]. Seas [815]. Season [486, 480, 150, 281, 2447, 1337, 840, 1159, 1530]. Seasonal [286, 2505, 1593, 55, 657, 1943, 504, 2174, 873, 62, 381, 1850, 105]. seasonal-like [381]. seasonally [1934]. seasons [25, 113, 2547, 952]. seawater [618, 470, 548, 1013, 134, 2614, 1461, 1827, 29, 631, 108, 2185, 1777, 311, 2715, 2018, 914, 1633]. Seaweed [567, 2623, 2150, 2223, 1683, 699, 1248, 1463, 1253, 1176, 1878, 2589, 2681, 2175, 2257, 2663, 1272, 2295, 1719, 847, 990]. Seaweed-based [2223]. seaweeds [1001, 2417, 2178]. sebago [2544]. Sebastes [2638, 2538, 1135, 1297, 742]. second [852]. secondary [2535, 1107]. secreted [2605]. sectional [2327]. sections [559]. sector [2599, 1416, 1456, 2705, 1408, 2616]. sectors [1411]. securis [1305, 636]. security [2673, 2499, 2052, 2574]. sedation [2419]. Sedative [785]. sedentary [2213]. sediment [1534, 1587, 56, 2631, 2411, 1749, 1726, 1741, 1279, 2119, 1494, 1424]. Sedimentation [227]. sediments [1099, 2675, 1272, 1961, 386]. seed [1907, 1557, 1377, 2095, 178, 927, 1596, 1864, 2647, 442, 1200, 2273, 133, 1437, 1807, 1260]. seeded [2359]. seeding [1569, 356, 2630]. seedling [2493]. seedlings [2386]. seeds [1595]. Seeking [1938]. segment [1140].

Segmentation [2464]. Segregation [852, 1269, 675]. Selected [1421, 730, 1708, 1216, 1896, 1742, 666, 2721, 1087, 766, 2462, 2330, 911]. Selection [2433, 1263, 2727, 1288, 1119, 2155, 2302, 1670, 1346, 1210, 2376, 709, 1127, 1034, 79, 790, 2068, 127, 121, 2542, 865, 1003, 1761]. Selective [595, 1249, 1557, 1487]. selectivity [1312]. selenium [2303, 1244, 1254, 2677]. Selenka [1239, 839, 1003, 1299, 1483, 1607, 1373, 870, 1343]. Self [1136, 2352, 334, 1400, 1591, 2568]. self-feeding [2352, 334, 1591, 2568]. self-help [1400]. Self-thinning [1136]. semen [743, 2481]. Semi [245, 93, 795, 1580, 1941, 2057, 1585, 352, 2597, 2243, 2507, 1968]. semi-arid [2597]. semi-artificial [1941, 2057]. semi-automated [1968]. semi-intensive [93, 1290, 795, 1580, 1585, 352, 2243, 2507]. semicircle [1375]. semicirculatus [1375]. semilaevis [1320]. Seminal [484, 1216, 1996, 2328, 454, 2747]. semisulcatus [2088, 513]. Semisulcospira [880]. Senegal [572]. senegalensis [867, 501, 1091, 1089]. Senegalese [721, 501, 1091, 1089]. sensed [790]. Sensitivity [778, 2553, 2195, 1003]. sensitized [2280]. sensor [57]. Sensory [372, 1180, 1880, 2011]. Sepia [825, 350, 391, 438, 458, 568, 712, 773, 832, 357, 1918, 468, 628, 2575]. septicaemia [1663]. septicemia [987, 1301, 1806, 2290, 2009]. sequence [2389, 2276, 777, 1544, 2222]. sequences [1154]. sequencing [2733, 2030, 2332, 1544, 2546, 2163]. Serbia [1979]. serial [1131]. serial-use [1131]. series [1117]. Seriola [2404, 495, 1244, 794, 1312, 1084, 694]. seriolae [2404, 794, 1969, 1744]. serious [2645]. Serovar [343]. Serrasalmidae [2315, 2225, 2053]. serrata [934, 1377, 545, 546, 571, 1325, 1368, 937, 928]. Serum [220, 85, 1321, 2299, 755, 2452, 2197, 604, 794, 2072, 136, 400, 820, 1496, 2435, 1840, 1360, 735, 1561]. service [2529]. services [1853, 1525]. sesame [1258]. sesamin [1163]. sessilis [1252]. set [124, 1390]. seticornis [2654]. sets [1187]. settlement [615, 795, 106, 1612, 640, 103, 2283]. settlements [105]. settler [2446]. settling [804]. seven [1938]. Several [218]. severe [2699, 2645]. severe-intensity [2699]. severum [703]. Sex [822, 695, 1679, 2523, 404, 1214, 1320, 2212, 1508, 1714, 2332, 2534, 379, 1232, 1892, 133, 750, 2160, 711, 1759]. sex-reversed [1214]. sexual [801, 161, 2447, 1941, 2057, 1553, 758, 2647, 1092, 1365, 2665, 1035]. SFE [2589]. sGnRH [150]. sGnRHa [1747, 2747, 1158]. SGSY [1027]. shade [1196]. shading [2285, 2393]. shallow [769, 104, 154, 1318]. Shanghai [926]. shape [721, 1979]. Sharma [1205]. sharp [888, 886, 983]. Sharpsnout [506, 302, 1465, 919, 372, 1422]. Sharptooth [801]. Sheffield [1975]. shelf [1920, 1576, 394, 2056]. shelf-life [394]. shell [1843, 1734, 1745, 1080, 631, 104, 1329, 1630, 14, 1144, 2528, 891, 1955, 2021, 1352]. shelled [1012, 1700, 1752]. Shellfish [416, 1479, 144, 1525, 1064, 1120, 1345, 1528, 1272, 803, 812, 297]. shellfisheries [2501]. shells [2528]. shelter [370, 1368, 369]. Shenzhen [792]. Shewanella [1089, 892]. shi [1020]. shifts [2090]. shiitake [1192]. ship [1232]. shock [1293, 1086, 1157, 778, 2334, 1359, 1300, 2296, 1003, 674]. shore [1569, 535, 938]. Short [212, 1309, 161, 178, 200, 1259, 1174, 1774, 2042, 151, 140, 193, 186, 177, 2076, 2560, 2001, 1173, 1239, 1472, 17, 1221, 1504, 2594, 593]. short- [17]. short-chain [2560, 1774, 2594]. Short-term [1309, 1173, 1239, 1472, 1221, 1504, 593]. shovelnose [201]. show [2679]. shows [1564]. Shrimp [232, 2232, 569, 622, 2259, 628, 1355, 2699, 2560, 1479, 2552, 1907, 1721, 68, 673, 2517, 2088, 730, 1619, 2451, 523, 1290, 1904, 1652, 1248, 1358, 1277, 2043, 1119, 1887, 1463, 1836, 2277, 2138, 1754, 1384, 1678, 2427, 1005, 2090, 1899, 1780, 2532, 13, 207, 358, 1256, 1950, 2240, 1829, 2494, 981, 1202, 640, 966, 2558, 2325, 2319, 1396, 1992, 1393, 1289, 1259, 1379, 1860, 381, 2086, 1784, 2609, 2736, 1538, 1901, 1487, 2738, 998, 2110, 824, 1198, 1810, 647, 915, 575, 1549, 2231, 1970, 2589, 2602, 2631, 1654, 1055, 2218, 1791, 1441, 1614, 2739, 2327, 2487]. shrimp [980, 2381, 1725, 1598, 853, 2314, 2448, 2205, 2719, 889, 1268, 1737, 1324, 2472, 2576, 26, 140, 1460, 2322, 1768, 2203, 2621, 1485, 2608, 1796, 80, 576, 2316, 2518, 1741, 2694, 95, 524, 868, 2177, 1145, 1773, 2079, 1902, 2491, 1019, 982, 373, 2006, 2049, 1994, 2551, 2596, 1315, 2586, 2591, 1104, 1717, 559, 2703, 306, 1880, 1620, 2603, 1183, 1502, 2243, 2507, 1961, 1130, 1138, 2295, 2105, 1493, 941, 1909, 2000, 2594, 362, 2100, 403, 2588, 2102, 1302, 2119, 1036, 522, 526, 1348, 1185, 1389, 971, 1788, 1711, 2454, 1705, 1962, 1025, 891, 2520, 2018, 1114, 1361, 2480]. shrimp [2620, 2655, 734, 2549, 1146, 2038, 735, 782, 1592, 1689, 1644, 1397, 2438, 2208, 2406, 679, 552]. Shrimps [233, 2125, 2342, 2347, 1660, 1976]. sHsp [2334]. siamensis [2600]. Siamese [168, 2516, 1601, 1599, 1924]. sib [1480]. Siberian [1247, 208, 2219]. sibling [1457]. sides [843]. sidoides [2656]. Siganus [702, 990]. sign [2618, 2733, 2748, 1908, 2717, 2652, 2265]. signal [638, 271, 998, 2278]. signaling [2418, 2491]. significant [2645]. significantly [1628]. silage [1839, 2473]. Siluriformes [1097]. Silurus [1226, 1164, 1211, 1988]. Silver [251, 237, 1619, 300, 2535, 1792, 325, 2296, 360, 1660, 2152]. silymarin [2693]. similar [2525]. similis [2026, 1798]. Simple [569, 50, 1544, 1835, 56, 1651, 777, 389]. simulate [1915]. simulated [2419, 321, 778, 1101, 528, 2454]. Simulation [2701, 2132, 2002, 2604, 851]. simulations [602, 619]. simultaneous [1476, 1697]. simultaneously [1013]. Sinaloa [54]. since [1519]. sinensis [656, 2475, 2114, 1680, 1956, 2227, 1142, 918, 929, 932, 2672, 2395, 2642, 1279, 931, 1050, 753, 930, 2373, 2361, 985, 2382, 2548, 2688, 1012, 1799, 2500, 1700, 1649, 1752, 2639]. Singhi [221]. Single [1896, 2230, 2526, 395, 2330, 2099]. Single-nucleotide [1896, 2099]. singly [1358]. singularly [946]. Siniperca [856, 2672, 1392, 2323]. sink [1593]. sinks [1586]. Sinuolinea [2122]. Site [2155, 790]. sites [1578, 2036, 1067]. situ [793, 1668]. Six [250, 289]. Size [313, 248, 1448, 4, 300, 548, 1670, 281, 273, 841, 659, 768, 720, 125, 754, 2459, 1929, 2345, 1368, 2359, 758, 1034, 1943, 445, 1165, 79, 126, 1547, 434, 702, 2188, 672, 1273, 172, 1876, 697, 1149]. sized [536, 1466]. sizes [1619, 2054, 2731, 113, 575, 1749, 1726, 576, 644]. skeletal [302, 1231]. skeleton [1642]. Skeletonema [250, 2046]. skill [1337]. skills [1411]. Skin [2685, 1995, 1446, 2670, 2712, 1124, 975, 1596, 1800, 1050, 2396, 2495, 2569, 2753]. skirt [2704]. slag [2260]. Slaughter [228, 171, 532]. slaughtered [1583]. slaughtering [456, 405]. SLE [2681]. slip [2479]. SLP1 [1814]. sludge [2294, 2267, 2196]. Small [977, 1917, 640, 2499, 2198, 1066, 1647, 2343, 1738, 1325, 74, 1378, 835, 1371, 1010, 2700, 2364, 1788, 1871, 2204, 1512]. small-scale [1917, 640, 2499, 1647, 1325, 74, 1378, 835, 2700, 2364, 1788, 1512]. smallholder [1984]. smallscale [1989]. smart [2541, 2715]. Smith [2126]. smoked [1180, 547, 670]. smolt [635, 1013, 31, 49, 1313]. smoltification [92]. smolting [63]. smolts [81, 12, 1314]. smooth [1320]. SN1 [2746]. snail [880, 2477]. snakehead [176, 2058, 2037, 1437, 2402, 1561]. snakeskin [1800, 2048]. snapper [1018, 1353, 2697, 1503]. snook [1317]. snout [888, 886, 2035, 983]. SNP [1669, 1707]. SNPs [1692, 1352]. Social [819, 1724, 1325, 2077, 768, 50]. Socio [2349, 911, 1027, 1400]. Socio-demographic [2349]. Socio-economic [911, 1027, 1400]. socking [632]. socks [1251]. sod [1299]. Sodium [2545, 2005, 1356, 1763, 1764, 2472, 2511, 1131, 2180, 1022]. soft [12, 1745, 1012, 1700, 1752]. soft-shell [1745]. soft-shelled [1012, 1700, 1752]. soil [1005, 2386, 2479]. soil-grown [2386]. soils [95]. solani [2621, 2694]. Solanum [472, 2681]. Solar [1611, 2586]. Solar-dried [1611]. soldier [2706]. Sole [572, 721, 687, 122, 698, 841, 1320, 501, 1091, 1089]. Solea [867, 501, 1091, 1089]. solids [1555, 2709]. Solomon [2134]. solubility [1706]. solubilizing [897, 321]. soluble [433]. Solution [488, 869, 1342, 176]. solutions [1817, 1522, 1223, 175, 2374]. solvent [2396]. somatic [968, 1280]. somatostatin [1469]. somatostatin-14 [1469]. Some [885, 1637, 247, 579, 484, 610, 2046, 336, 2060, 183, 2633, 585, 2034, 1484, 173, 2457, 820, 2089]. sorting [1929]. sorubim [1750, 1751, 2123]. Souahlia [2155]. Sound [1853, 529]. Sounds [416]. Source [628, 192, 1870, 2293, 2427, 2267, 350, 2558, 1393, 196, 1077, 115, 1953, 2233, 2400, 1091, 1376, 1524, 1125, 2074, 1545]. sources [2617, 2142, 1490, 64, 1740, 996, 1307, 2416, 1102, 956, 2328, 2183, 2355, 1040, 2521, 2449, 555, 1593, 2498, 2175, 880, 1586, 2537, 1987, 2255, 1452, 1106, 1752]. South [391, 2062, 1782, 2417, 283, 2475, 1789, 1853, 1670, 545, 546, 1612, 1589, 2583, 2136, 2405, 2317, 2250, 185]. south-eastern [2417]. Southeast [751, 2317]. southeastern [2173, 1251]. Southern [282, 251, 1786, 2198, 2172, 303, 158, 1144, 434, 1190, 2135, 2721, 1788]. Southgate [1071]. southwest [848, 1721, 673, 2214, 2615]. Sow [489]. Sowerby [2429]. Sox17 [256]. Sox9 [256]. soy [2088, 942, 1678, 1486, 1895, 1036]. soya [1295, 71, 166, 1875]. Soybean [1776, 880, 1133, 799, 1445, 1018, 1465, 944, 1706, 1142, 114, 1040, 2719, 606, 547, 1824, 959, 1900, 753, 1316, 1625, 762, 940, 1807, 920, 2130, 2318]. soybean-meal-based [799]. sp [2192, 2109, 2131, 2046, 1375, 1859, 1490, 350, 1240, 1396, 1901, 2684, 2111, 757, 1241, 2543, 2690, 1919, 1959, 703, 715, 1814, 287, 2541, 1450, 2151, 2306, 1760, 379, 1866, 2746, 2480, 1815, 1545]. spacecraft [2157]. spaced [322]. spacer [895]. Spain [88, 103, 945, 282, 1305, 2172, 2364]. spallanzanii [1336]. Spanish [1565]. Sparidae [424, 974, 537]. sparids [305]. sparing [1271, 2305]. Sparus [882, 535, 2133, 1626, 364, 1622, 390, 1565, 405, 2367, 2197, 1486, 1349, 770, 798, 1332, 339, 246, 464, 1535, 898, 510, 2496, 844, 557, 1089, 789, 1029, 685]. spat [728, 279, 707, 1344, 986, 287, 2264]. spatfall [605]. spathula [454]. spatial [2036, 2155, 2519]. spatio [106]. spatio-temporal [106]. spawners [1869]. Spawning [486, 480, 150, 255, 286, 1817, 786, 2338, 923, 314, 1534, 698, 948, 1828, 771, 269, 1939, 103, 19, 1161]. SPC [2318]. special [203, 2086, 2311, 1519, 2013, 2177, 46, 1866]. Species [1479, 238, 183, 503, 621, 246, 284, 483, 2059, 2173, 2601, 424, 1859, 1342, 106, 1913, 290, 2366, 1558, 398, 791, 84, 2081, 2170, 636, 2460, 2045, 752, 719, 2376, 698, 86, 1848, 715, 1420, 1803, 140, 2721, 1197, 2694, 517, 518, 105, 2250, 393, 2528, 2746, 127, 2384, 733, 670, 1710, 2300, 1016, 1128, 2465, 1488, 2149, 2379, 2108, 709, 974]. Species-characteristic [183]. Species-specific [1479, 1848, 1197]. Specific [2411, 1479, 1603, 871, 1655, 1680, 2607, 1501, 2231, 1848, 1197, 1675, 2514, 2346, 2151, 1930, 965, 2396, 1705, 1192, 2655, 1403, 1794, 2536, 2152, 1540, 2318]. spectra [2362]. spectrophotometric [744]. spectroscopic [2400]. speed [1856, 1111]. Speight [955]. Sperm [869, 1817, 1216, 715, 1464, 483, 744, 2287, 588, 1215, 266, 783, 1996, 1117, 1127, 1217, 1247, 1156, 584, 2747, 581, 674, 748, 2096, 1219]. spermatogenesis [161, 2373]. spermatophore [2106]. Spermatozoa [485, 488, 1218, 2747, 1221, 1222]. Spezia [1482]. Sphoeroides [435]. spider [731]. spinach [2740, 1383, 1495]. spined [2503]. spinifera [713]. spiny [2221, 1458]. spiny-bellied [2221]. Spionidae [2528]. Spirocamallanus [2028]. Spirodela [451]. Spirulina [2436, 2604, 1041, 1690, 2084, 1077, 1311, 1596, 1473, 2506, 1978, 1302, 242]. spleen [2653, 1805, 2070]. spleen-NMR-metabolomics [2653]. splendens [2516, 1601, 1599, 1924, 2073]. splendidus [635, 1403]. splenic [2234]. Spondylus [1885]. sponge [1031]. spontaneous [97]. Spore [567]. sporophytes [1103]. Spot [569, 730, 1277, 1887, 2609, 2736, 1467, 2110, 1664, 1855, 2387, 2218, 1473, 894, 2008, 2484, 2566, 1773, 2150, 2632, 559, 1961, 1185, 1059, 1711, 2018, 1331, 1361, 2480]. spotted [2687, 2754, 548, 827, 910, 909, 946, 979, 1001, 1353, 783, 351, 902, 1364, 408, 631, 756, 2022, 2697, 2441, 2440, 1455, 1503]. spp [490, 2653, 410, 2014, 2641, 2177, 2185, 850, 2123, 2089, 2375]. sprayed [762]. spring [97]. SPUM [1970]. Squalius [2235]. squamosa [1196]. squid [2676]. squilla [1839]. squirt [984]. SRAP [849]. Sri [147]. ssp. [1484]. SSR [1502]. SSRs [1544]. stability [2192, 2462]. stable [1570, 542, 2718]. Stackhouse [2215]. stage [519, 1418, 1119, 1954, 981, 2538, 369, 1539, 2425, 468, 2438, 2510]. stage-2 [369]. Stages [226, 1108, 504, 2404, 2088, 2090, 1227, 205, 314, 1640, 351, 2249, 1639, 1712, 1830, 403, 407, 580, 1003, 2222, 1799, 1976, 447]. staining [2237, 1897]. stakeholder [2599, 706, 2260]. stand [1212]. standard [1548]. Standardization [2287]. stands [654]. Staphylococcus [1956, 2152, 1592]. starch [1954, 1030, 2683, 685, 1503]. starry [1542]. stars [1024, 2359]. start [1874, 2537]. start-feeding [1874]. start-up [2537]. starter [9]. Starvation [405, 761, 2699, 907, 1504, 1618]. starved [426, 591, 1694]. State [54, 1950, 2552, 2713, 2458, 1620]. States [617]. static [207, 358, 2176]. station [598]. Statistical [1400, 576]. Statistics [1283]. Status [138, 2398, 571, 2217, 2692, 1319, 1748, 2147, 2427, 871, 2626, 2752, 883, 1999, 1745, 1765, 2707, 359, 193, 1941, 2057, 2750, 810, 2347, 2597, 1275, 2647, 1502, 2650, 2667, 2281, 1551, 648, 599, 2742, 686, 2144, 2646]. Staurogyne [2601]. steelmaking [2260]. steep [908]. Steindachner [1353, 314, 2697]. stellatus [1542, 1503]. Stenorhynchus [2654]. step [730, 2584]. steps [1285, 859]. Stepwise [976]. sterlet [1328, 2212]. sterna [566]. steroid [1828, 1542, 379]. steroids [161, 750]. Sterol [2173]. stewartii [897, 897]. Stichopus [2543, 1391, 2285, 2393]. Stickney [296, 862, 446]. still [191]. Stimpson [381, 2126]. stimulant [2621]. stimulate [1261, 167, 1981]. stimulated [1357]. stimulation [2137, 1213, 2689, 1159]. stimulations [2587]. stimuli [890]. stinging [2360, 1081]. Stizostedion [452]. Stochastic [2498, 1058, 2507]. stock [2235, 595, 927, 1320, 918, 1197, 284, 2038, 307]. Stocking [721, 833, 1090, 11, 2429, 493, 566, 825, 299, 2559, 2015, 70, 2458, 684, 548, 878, 1100, 2394, 1186, 2731, 1310, 2094, 811, 281, 1792, 988, 408, 65, 2739, 1054, 839, 2598, 1977, 2233, 1368, 352, 2514, 759, 1165, 2127, 1559, 1547, 1712, 702, 2391, 133, 968, 1539, 641, 1820, 2382, 1840, 1360, 1660, 2646, 999, 1698, 2438, 2453, 2201]. Stocks [481, 954, 1002, 948, 1385, 91, 1965, 1728]. stomach [1943, 2305]. stomachs [1655]. Ston [1597]. storage [345, 2544, 1576, 1936, 1221, 557, 669, 2662]. stored [1472, 1920, 2056, 2011]. Strain [507, 2559, 1997, 155, 2187, 738, 2641, 1970, 2136, 2727, 48, 87, 107, 1702, 2339, 2546, 2021, 1373]. Strains [482, 1194, 1696, 693, 1445, 1652, 2246, 298, 2030, 906, 1258, 1390, 522, 526, 2477]. stranded [1705]. Strategies [2007, 278, 2340, 349, 2731, 2338, 2468, 2227, 1688, 2027, 868, 1787, 913, 1378, 2243, 1834, 2724, 2510]. strategy [786, 2367, 2627, 859, 1285, 1974, 1078, 1034, 2650, 2117, 2600, 893, 2397, 1581]. stratification [55, 326]. stratified [1042]. streaked [2481]. streams [1979]. Streptococcosis [1990, 1919, 2633, 1809]. Streptococcus [2653, 2618, 2111, 1919, 2171, 2329, 2482, 1969, 2306, 2592, 1562, 2181, 1809, 1815, 2234, 1731]. Streptomyces [1652]. Stress [228, 1170, 234, 2629, 1812, 1907, 153, 2344, 1309, 392, 535, 2297, 981, 2094, 1227, 1229, 2319, 49, 2227, 2743, 2140, 2637, 2378, 2564, 1135, 1642, 1827, 2633, 2428, 1054, 1977, 2365, 2322, 2220, 2334, 536, 96, 2316, 173, 2013, 2642, 1800, 767, 371, 1315, 2497, 2630, 395, 1112, 1805, 2577, 2539, 750, 85, 2361, 2441, 2440, 2655, 2548, 2688, 2278, 1871, 870, 1343, 1561, 2658, 2686, 2201, 448]. stress-related [2637]. stresses [1299]. stressful [1568]. stressors [1498, 1170]. striata [2058, 2037, 2402]. string [1298]. Stringed [793]. striped [1888, 1477, 1972, 1730, 618, 2277, 2107, 660, 1781, 778, 2356, 2492, 852, 2554, 440, 2514, 90, 355, 2330, 2301, 1246, 1025, 2698]. Stripping [489]. Strombus [1720, 2241]. strong [2658]. Strongylocentrotus [2410, 1419, 1864, 1274, 1559, 968, 1280, 1650, 1149, 1341]. Structural [2386, 255]. Structure [490, 621, 901, 2523, 811, 1345, 1854, 1214, 1382, 1756, 1755, 1231, 1786, 1211, 1460, 845, 2030, 1154, 1275, 2648, 655, 293, 2301, 1948, 1947, 669, 386, 1761, 1451, 2182]. structured [1002]. structures [1701, 2547, 591, 2536]. Struthers [863]. Student [1412]. Studies [16, 571, 2530, 584, 2105, 549, 1793, 2351, 206, 623, 50, 26, 1233, 534, 1798, 1616, 803, 664, 85, 831]. Study [340, 1345, 962, 1362, 1737, 2092, 105, 473, 1104, 1232, 1976, 2658, 457, 2475, 953, 1015, 2067, 2142, 2458, 1262, 1529, 2509, 1242, 1740, 2732, 183, 113, 2214, 2154, 278, 1307, 1298, 1985, 1827, 977, 1651, 2327, 176, 2341, 2735, 2097, 1571, 917, 1073, 2106, 2185, 1065, 452, 963, 1087, 1620, 620, 1930, 1505, 594, 2224, 1979, 436, 537, 542, 1949, 2164, 1875, 2069, 1660, 1922]. stunted [1262]. sturgeon [1076, 2702, 2618, 2455, 2399, 1170, 1227, 1328, 1028, 1903, 1109, 2515, 994, 1094, 1848, 1247, 201, 1590, 875, 776, 1232, 750, 1857, 208, 2219, 2265]. stutzeri [1883]. Styela [1391, 1606, 1546]. stylirostris [381]. Sub [683, 627, 2310, 809, 2261, 2686]. sub-adult [2310]. sub-inhibitory [2261]. Sub-systems [627]. sub-tidal [809]. sub-tropical [2686]. subjected [629, 1170, 2622, 312, 1723]. Sublethal [12, 1661, 2397]. submerged [904, 2704, 2710]. submersion [1174]. submitted [1138]. subsequent [280, 1012]. subsidized [1844]. subsp [2105]. subsp. [2415, 897, 1560, 1762, 2102, 1132]. subspecies [183]. subspecies-characteristic [183]. substances [1490, 614, 259, 1959]. substantially [2479]. Substitute [1678]. Substituting [2562]. substitution [396, 1047, 1109, 1142]. Substrate [241, 865, 1134, 2513, 2385, 595, 830, 2079, 1243, 160]. Substrate-SDS-PAGE [160]. substrates [1532, 1401, 1856, 2203, 2525, 2270, 1104, 865, 2432]. subtidal [772]. subtilis [2045, 2169, 1048, 1137, 1449, 1757, 2463, 2608, 2316, 2694, 1063, 2151, 1873, 2105, 2076, 1966, 1883, 1731, 2176]. subtropical [2366]. Success [662, 2421, 286, 1472, 1624, 771, 1165, 1219, 1159]. Successful [345, 483, 2509, 1828]. Succession [2575, 1902]. successive [2106, 1761]. suckers [146, 145, 153]. suecica [818, 1461, 1823, 209, 2679]. sugar [2011]. Sugarcane [2631, 2513]. suggestions [2458]. Suitability [2267, 1380, 682, 1614, 85]. Suitable [743, 2170, 1973, 2433, 1401, 2250, 2444]. Sulawesi [2136]. sulfate [2740, 473, 2195]. sulfated [1887, 2229]. sulfide [2404, 2232]. sulphur [1970, 306]. sulphur-oxidizing [1970]. Sumatra [2253]. Sumich [861]. summer [398, 1681, 432, 1650]. sunfish [874]. supended [401]. super [2428, 1902, 2289, 1689]. super-intensive [1902, 2289, 1689]. super-paramagnetic [2428]. superfamily [2137]. supernatant [2267]. superoxide [1649]. superzist [2024]. supplement [796, 266, 833, 2231, 2551, 2020, 2594]. supplemental [2543, 2035, 669].

Supplementary [1584, 579, 908, 1235, 1580, 1280, 1649]. Supplementation [2130, 1812, 2063, 1477, 1972, 1645, 1255, 2029, 2604, 1049, 2003, 1627, 1319, 1887, 827, 946, 2430, 1256, 1173, 1052, 2743, 2748, 2738, 824, 2335, 2075, 1217, 604, 1126, 1278, 1449, 1736, 2032, 1757, 2472, 2463, 1687, 346, 1254, 1432, 2066, 2611, 2281, 1719, 1199, 1316, 2301, 1437, 2698, 2742, 1662, 1517, 2639, 2176]. supplementations [2314, 820]. supplemented [1635, 1766, 1563, 1231, 1882, 1814, 1886, 2151, 1713, 47, 1200, 2180, 1374, 1966, 1486]. supplements [2277, 204]. supplied [1991]. Supply [495, 1076, 687, 2590, 1083, 369]. Supplying [1733]. support [1414]. suppressed [2462]. suppression [7]. suratensis [2232, 1516]. Suratthani [1786]. Suresh [1284, 1205]. Surface [60, 137, 2078, 370, 1287, 2076, 2247, 2710]. surfacing [2703]. surgery [2357]. Surgical [230, 1303, 776]. surmuletus [1730]. Surubim [2076, 1779]. surveillance [2226, 2750]. survey [499, 1557, 871, 1575, 2327]. survey-based [2327]. Survival [232, 1635, 500, 568, 235, 240, 426, 325, 1433, 493, 566, 233, 324, 237, 510, 516, 360, 2723, 1196, 2552, 731, 2617, 2142, 2654, 1354, 2088, 728, 749, 439, 460, 1119, 9, 70, 684, 946, 1100, 279, 472, 1122, 934, 2532, 1186, 1612, 350, 438, 458, 832, 816, 1256, 531, 1874, 257, 1703, 1043, 949, 2148, 1766, 2230, 1340, 1701, 1439, 1351, 752, 1266, 1487, 841, 757, 1958, 1642, 647, 2589, 2631, 633, 988, 65, 1162, 2218, 1473, 958, 2249, 2459, 631, 1211, 1499, 1929, 805, 2623, 839, 935, 2449, 401, 1288, 1742, 1084, 722]. survival [1630, 1796, 1823, 280, 501, 524, 1554, 517, 518, 913, 1983, 2213, 181, 982, 373, 1928, 1601, 2622, 469, 464, 1559, 1842, 18, 921, 2285, 2393, 1200, 434, 1712, 331, 369, 2462, 2667, 1830, 702, 431, 968, 2588, 172, 641, 453, 1367, 580, 2300, 1795, 865, 1867, 724, 814, 1105, 1617, 1820, 208, 1190, 1618, 1794, 2431, 1035, 782, 1280, 1722, 831, 2453, 2201]. Survivorship [399]. Susceptibility [635, 2206, 2566, 1022, 2380, 1171]. susceptible [2234]. Suspended [1778, 1251, 282, 681, 370, 1703, 793, 383, 2709, 637, 873]. Sustainability [2650, 819, 2214, 2625, 740, 2252]. Sustainable [467, 2416, 2541, 2421, 2286, 1520, 2673, 529, 415, 1549, 709, 1204, 1031, 2506, 2561, 2080, 1111, 2259, 2504, 2616, 2669, 1051, 2174, 563, 274]. sustainably [1985]. sustained [195]. sustained-release [195]. sustaining [2730]. Suthers [864]. Svasand [307]. SVC [97]. Svennevig [274]. swamp [1468, 2598, 722, 1625, 2745, 1148, 1540]. swarm [1349]. swarms [1626]. Sweden [1382, 1264]. sweet [1989, 2011]. sweetness [1341]. sweetpotato [2479]. Swim [1932, 1084, 1208, 1928]. Swimbladder [1190, 960]. swimmer [1618]. Swimming [355, 1770, 1127, 935, 2377, 96, 1704, 1111, 2098, 1060, 2431, 1260]. switched [677]. SWOT [2269]. Syda [1284]. sylvestris [2003]. Symphodus [635]. Symphysodon [410, 850]. symposium [926]. symptoms [1377]. synbiotic [1901, 2508, 2585, 2520, 2361, 1759, 2510]. synbiotics [2067, 2723]. Synchronization [1158, 2749]. Syndrome [569, 2666, 730, 1277, 1887, 2609, 2736, 1467, 2110, 2387, 2218, 894, 2008, 2484, 2566, 1773, 2006, 2150, 2632, 559, 1961, 1185, 1059, 1711, 2018, 1331, 1361, 2480]. Synechogobius [2572]. Synergistic [2511]. synergistically [1952]. synergy [2645]. synspilus [1636]. synthase [2678]. Synthesis [812, 2311, 624, 2636, 611]. synthetase [980]. Synthetic [570, 824, 703, 875, 2292]. System [627, 494, 2242, 2370, 54, 2433, 577, 1129, 2131, 947, 2451, 192, 2279, 1248, 1334, 1286, 2026, 1463, 1836, 1514, 827, 1001, 1529, 1960, 2298, 2532, 1558, 2394, 2668, 207, 358, 2734, 1532, 2468, 652, 2638, 1555, 2446, 1862, 1989, 1052, 2140, 2256, 1383, 2452, 709, 1391, 1606, 2183, 2386, 631, 1725, 1598, 2505, 2533, 520, 521, 1546, 2032, 1822, 1940, 1953, 2576, 1808, 613, 1593, 2345, 341, 1101, 2203, 1497, 602, 619, 2121, 1270, 15, 2514, 1495, 1893, 2079, 2508, 2551, 371, 2709, 2622, 1760, 790, 963, 1898, 1087, 2537, 2422, 1232, 2178, 2295]. system [626, 1365, 41, 1452, 2715, 1677, 2585, 1302, 834, 973, 2289, 1968, 44, 154, 1389, 2640, 2504, 1462, 2382, 2069, 2489, 892, 592, 1318, 2438, 2208, 1355, 2453, 2201, 2176]. systematic [2573, 2716, 1899, 2721, 2422]. systemic [2482]. Systems [516, 1746, 2421, 1015, 2740, 285, 77, 2360, 2083, 523, 470, 439, 2610, 548, 1709, 681, 2304, 390, 811, 1061, 1393, 660, 1259, 1168, 1889, 2555, 258, 1849, 2158, 2367, 2351, 1198, 1564, 1861, 388, 1526, 2459, 793, 2052, 2381, 627, 1308, 2411, 722, 2022, 1814, 333, 1368, 740, 74, 1983, 1825, 1811, 710, 1639, 2236, 1937, 763, 1426, 1107, 292, 1877, 1053, 1935, 2432, 1807, 1689, 1988, 1512, 2606, 1152].

T [307, 449, 992, 1056, 1905, 1450, 2117, 164]. T-ISO [164, 1450]. T. [692, 720]. T3SS [2238]. tabaci [2682]. tablet [2331]. Tachysurus [1733]. tackling [2705, 1946, 1945]. tadpoles [2221]. tag [1273]. tagging [1053, 1931, 544]. tags [528]. tail [1358, 1077]. taimen [1739, 1617, 2495, 1739, 1617, 2495]. Taiwan [2260, 1513, 1298, 765, 1660]. Takifugu [496, 939]. Tamarindus [2273]. tambaqui [45, 2254, 2419, 2352, 2308, 2737, 2613, 2718, 1731, 2569, 2753, 2201, 2176, 2606]. tambatinga [2590]. tambroides [2433]. Tamil [713, 1183]. Tampa [158]. tangs [1375]. tank [254, 1422, 1800, 1928, 61, 1617]. tank-reared [61]. tanks [2513, 5, 1567, 1110, 1066, 633, 1738, 199, 1508, 1881, 133]. tannin [2728]. tannin-based [2728]. Tanzania [2030, 2597, 911]. Tapes [439, 541, 536, 951]. tapetis [635]. Taranto [1336, 543]. tarda [1548, 2273]. targeting [1402, 2691, 2149]. taurine [2020, 1350]. tauvina [739]. Taxonomy [374]. TCM [1238]. TCTP [2139]. tea [1135, 1662, 1405]. teaching [1414]. Technical [1058, 1659, 1782, 1115, 1427, 1522]. technique [2740, 1303, 1514, 627, 339, 2481, 776, 1892, 121]. technique-type [1514]. Techniques [785, 2716, 1375, 970, 2019, 923, 938, 998, 830, 56, 1630, 686]. Techno [2343]. Techno-economic [2343]. technologies [1212, 1984, 2163, 725]. Technology [505, 2391, 1671, 2716, 2398, 1740, 981, 1202, 2434, 1289, 1259, 1381, 1379, 2355, 2487, 2598, 1768, 2570, 2233, 2622, 2166, 1946, 1945, 1620, 2305, 911, 2337, 2208]. tehuelche [332]. tehuelchus [332]. telemetry [53]. Teleost [484, 856, 1556, 2074, 990]. Teleostei [392, 1828]. Temminck [759]. temperate [1891, 453]. Temperature [471, 2209, 226, 219, 788, 2654, 1354, 2207, 825, 1690, 38, 340, 391, 1256, 784, 63, 1874, 2221, 12, 919, 404, 402, 883, 1118, 1472, 705, 351, 1351, 752, 897, 1609, 677, 1313, 1162, 720, 163, 2249, 976, 839, 1268, 2727, 771, 190, 186, 2106, 1952, 1359, 177, 181, 2033, 1928, 28, 76, 90, 1661, 2595, 644, 156, 311, 822, 672, 24, 453, 2056, 2021, 2536, 1698, 2143, 745, 2686]. temperature-control [788]. Temperatures [242, 519, 1720, 1566, 1164, 148, 1312, 1297, 2497, 172]. Temporal [2036, 106, 2276]. temporally [322]. ten [1693]. Tenacibaculum [693]. Tench [687, 600, 487, 598, 488, 599, 588, 684, 796, 595, 1340, 590, 583, 948, 587, 584, 596, 585, 1143, 1359, 591, 586, 594, 597, 668, 581, 593, 589, 592]. tension [57]. tension-recording [57]. tensions [1101]. Tenth [803]. tenuiflorum [2131]. tenuistipitatum [2295]. Teod. [2713]. Term [218, 1309, 1173, 1239, 1472, 586, 1559, 17, 1221, 1504, 968, 1795, 593, 1021]. Terminalia [2516, 2153]. terms [2202]. Terrestrial [688]. terrestris [645]. Test [627, 4, 455, 583, 1263, 520, 521, 1548, 1832, 2157, 2145]. tested [586]. Testing [487, 1480]. testosterone [195]. Tests [675, 2468]. testudineus [2394]. tetra [2363]. tetrachloride [1417]. tetrachloride-induced [1417]. Tetrahymena [2324]. tetrameric [2628]. Tetraploid [1086, 1157, 1156, 2735, 324, 831]. tetraploidy [2320]. Tetraselmis [2046, 396, 818, 1461, 1823, 209, 2679, 175]. tetrasomic [1848]. tetrazona [2253, 2068]. text [862]. textural [557, 1657]. Texture [95, 1840]. Thaibinh [1787]. Thailand [640, 1786, 139, 1113, 19, 858, 491, 552]. thailandensis [1754]. Thalassiosira [958, 965]. Thau [144]. thaw [1219]. Their [251, 1033, 1172, 1269, 2666, 1696, 500, 2155, 2081, 183, 1064, 203, 1844, 1996, 1066, 863, 864, 1127, 2633, 2739, 2327, 984, 2362, 2672, 2334, 2345, 585, 1741, 1994, 1332, 1661, 1639, 2068, 1987, 1112, 2528, 2412, 2119, 1809, 1544, 1788, 2099, 85, 1114, 1707, 2179, 1203, 2219]. thelycum [2106]. them [2239]. thematic [1407]. therapeutic [2093, 2592, 2150, 2600]. therapy [1916, 2557, 2185, 1078]. thermal [901, 773, 1588, 1827, 326, 1805, 2728, 1003, 1299, 1161]. thermally [1584]. thermodynamics [1045]. thermotolerant [1045]. thiamine [1328]. thickness [1856]. thin [2187, 2023, 2727]. thin-blade [2187]. thin-lipped [2023]. thinning [1136]. Third [785, 1283]. thraustochytrid [1545]. threat [2645]. Three [627, 482, 2173, 2601, 1085, 947, 2701, 439, 2246, 410, 1558, 84, 816, 113, 1565, 2460, 719, 1164, 1473, 1385, 2735, 2395, 359, 2498, 524, 1250, 2412, 127, 2039, 1752, 1728]. three-spot [1473]. threonine [671, 1322]. throat [2069]. throughout [350, 1941, 2057, 1830]. throughput [2701]. Thunberg [460, 806]. Thunnus [1121, 1111, 1190]. Thymallus [2311]. thyme [1855, 2290, 1890]. thymol [2341]. Thyroid [220, 996, 136]. tidal [2501, 802, 809]. Tidwell [1039, 1152, 1975]. Tiger [245, 511, 233, 2068, 249, 866, 2088, 730, 1358, 1277, 2138, 2090, 13, 2325, 2738, 182, 1791, 1278, 1145, 1773, 1315, 2591, 559, 19, 952, 847, 1644]. Tilapia [248, 2670, 222, 1556, 269, 2306, 2194, 241, 2238, 1812, 1957, 1033, 1448, 1888, 2729, 1015, 2413, 1058, 1763, 1194, 2439, 2093, 2666, 1531, 2131, 799, 2083, 145, 1679, 2122, 1319, 2653, 2026, 1000, 2298, 290, 2051, 1709, 1558, 2095, 2450, 84, 2731, 1829, 2418, 2734, 2025, 899, 301, 1764, 2732, 2380, 2483, 2593, 780, 1921, 178, 2468, 291, 1889, 2161, 1440, 1515, 2084, 2169, 1640, 2354, 1188, 1792, 2134, 2111, 1641, 2627, 682, 189, 2751, 1501, 2452, 2183, 2371, 2633, 1442, 2064, 2017, 855, 853, 1181, 1765, 1453, 48, 1508, 2473, 1675, 333, 2482]. tilapia [2030, 1631, 1447, 1893, 1714, 83, 371, 1478, 1760, 2166, 1300, 2592, 1287, 1562, 2647, 2645, 1372, 378, 442, 379, 1200, 1762, 906, 2290, 1138, 2181, 1339, 2577, 2539, 133, 787, 834, 973, 2700, 1809, 1125, 1316, 854, 971, 2116, 1890, 1867, 1189, 1105, 1718, 1022, 2145, 1443, 1657, 1840, 2217, 2379, 2582, 1674, 1322, 1815, 746, 1728, 2234, 1444, 1237, 1759, 2061, 2661, 147, 2424, 1307, 650, 1833]. tilapias [48]. tilapiine [183]. Time [1591, 1479, 605, 2702, 439, 2015, 546, 1655, 2352, 876, 1590, 1554, 964, 30]. Time-restricted [1591, 2352]. times [1164, 414, 538, 137]. timing [771, 2320, 1359, 814]. Tinca [588, 684, 687, 796, 600, 487, 595, 598, 1340, 590, 948, 587, 584, 596, 585, 1143, 591, 586, 488, 594, 597, 668, 581, 593, 589, 592]. Tiostrea [560]. Tisbe [122]. Tissue [1402, 2284, 2687, 2754, 901, 2358, 1000, 2651, 2450, 899, 2732, 1179, 196, 2637, 912, 2386, 2072, 2547, 2697, 2332, 898, 1978, 930, 2655, 1035, 1106, 1139, 1649, 2469]. tissue-specific [2655]. tissues [856, 1329, 2311, 2482, 797, 2591, 1711, 2222, 2288]. titer [1301]. Tithonia [1998]. TLR [2418, 2587]. TLR1 [2587]. TLR2 [2587]. TNET [1416, 1407, 1413, 1408, 2117, 1412]. Tocher [956]. tocopherol [784, 767, 331, 116]. tocopheryl [767]. together [292]. Togo [2700]. Tolerance [618, 924, 219, 2112, 1812, 1907, 2610, 290, 2187, 1498, 981, 2094, 2460, 1641, 960, 2316, 1315, 1702, 1871]. Tomato [2681, 2386, 2362]. tomography [393]. tongue [1320]. tonsa [1240, 1310, 1472, 1624, 1771, 1632]. tool [2358, 600, 405, 1753, 1079, 1097, 1197, 829, 2252, 534, 20, 2603, 1667, 1868, 542, 2077]. tools [529, 1414]. Top [456, 1955]. Top-crossing [456]. Tor [2433, 1879]. Torr. [1200]. Total [228, 1044, 2418, 291, 1188, 1329, 881, 666, 2035, 2709, 1478, 17, 1795]. Totoaba [2708]. toxic [2031, 924]. Toxicity [570, 2428, 663, 513, 2722, 1037, 2053, 1259, 1379, 1228, 2162, 2356, 1147, 1536, 2153, 1599, 400, 1639, 1250, 1022, 2144]. toxin [2690]. toxins [2605]. Tra [1818, 1661]. trace [2091]. traceability [1950]. tracer [1594]. Trachinotus [1854, 1841, 2595, 508, 2074]. tract [1621, 2256, 850, 1146]. tracts [385]. trade [2086, 1982, 1967, 2669, 2077, 679]. trading [1525]. Traditional [418, 1238, 2049]. training [8]. trait [1201, 1288]. traits [1734, 444, 1119, 2246, 2714, 1669, 1622, 1459, 2535, 1257, 1080, 171, 1480, 1468, 1335, 1428, 445, 2068, 1930, 2425, 1707, 2146, 1373, 1722, 1444]. Transcription [2018]. Transcriptional [2733, 1877, 2370]. Transcriptome [2221, 2271, 2688, 2579, 1872, 2324, 2397]. Transcriptomic [2564, 2227, 2708]. transfer [1980, 1212, 1013, 134, 2637, 29, 964, 1809]. transferred [25]. transfers [310, 277]. Transformation [2357, 893]. transformed [2192]. transgene [2192]. transgenesis [2190]. transgenic [2190, 2559]. transient [2734]. transition [642, 1554, 431]. transitional [1793]. translate [1985]. translationally [2139]. translocation [2680, 1982]. transmission [1344, 1371]. transmitted [1815]. transmontanus [750]. transplantations [139]. transplanted [397]. Transponder [1053]. Transport [392, 345, 707, 2419, 778, 1042, 2662]. Transportation [153, 1601, 2539, 857, 2577]. transportations [1239]. Transpositional [741]. trash [910, 979]. trays [806]. Treasurer [1837]. treated [942, 1000, 1584, 1461]. treating [1808, 2675]. Treatment [407, 901, 629, 470, 1253, 773, 560, 2338, 195, 993, 1738, 1927, 2101, 2347, 2196, 47, 2296, 343, 2273, 82, 2167]. Treatments [2613, 295, 2093, 588, 2267, 1939, 2010, 1625]. T(R) [2282]. Trends [2086, 1582, 2573, 2716, 1898]. Trent [489]. Treptacantha [2175]. Trevally [251]. trial [1196, 1990]. Trials [806, 332, 517, 518]. triangle [1080]. Tribulus [645]. tributyrin [2472, 2130]. trichlorfon [2315, 2053]. trichlorphon [441]. Trichodinids [2093]. Trichogaster [2356, 1800]. Trichopodus [1473, 2048]. trichopterus [1473]. Trichosporon [2082]. trickling [2446, 2367]. tricornutum [2581]. Tridacna [1196, 805]. trilobatum [472]. trimaculatus [2377]. Triploid [574, 487, 2007, 551, 1179, 993, 584, 1434, 586, 831]. triploidization [877]. triploidy [1314, 1300, 185]. Tripura [2445]. trituberculatus [1457, 1704, 2098, 936, 1060, 2431, 1260]. Trophic [1451, 1245, 819, 2279, 2509, 1345, 1176, 1391, 1606, 1546, 1275, 2292, 1462]. Tropical [238, 1778, 2029, 1673, 2366, 370, 323, 1912, 1501, 977, 1786, 1032, 1458, 1850, 357, 268, 1952, 306, 379, 2470, 2686]. Trout [481, 482, 229, 2722, 4, 1295, 1789, 1133, 444, 2135, 1445, 212, 2358, 1603, 64, 1869, 266, 1997, 1568, 2128, 134, 2299, 196, 1047, 1141, 1214, 1207, 1311, 1470, 2226, 2328, 604, 1916, 1469, 409, 11, 1920, 322, 37, 2334, 10, 2644, 606, 1982, 96, 2321, 328, 1233, 1065, 1026, 2457, 1713, 1846, 47, 1658, 1783, 1476, 1697, 312, 644, 766, 884, 30, 1096, 343, 1432, 211, 1873, 1163, 1847, 41, 820, 2746, 1978, 837, 2726, 1044, 959, 1719, 648, 1979, 387, 1758, 2160, 1656, 611, 762, 327, 1662, 1387, 2222, 940, 1360, 2011]. trout [1727, 2204, 1139, 2685]. tructae [2644]. trumpeter [618, 778, 440]. Trushenski [1905]. trutta [1207, 959, 574, 551, 1869, 2282, 611]. trypsin [174, 1873]. tryptophan [1266, 1654, 1499, 2737, 1374]. tshawytscha [2156]. tsunami [858]. tube [287, 2149, 2065]. tube-dwelling [287]. tuberculata [1115]. tubiashii [2477]. Tucker [774]. tulipa [2213]. tumor [2137, 2139]. tumour [2664]. tuna [1121, 166, 1111, 1190]. Tunisia [1748]. turbid [1940]. Turbinaria [2632]. Turbot [500, 239, 1255, 368, 947, 470, 33, 1770, 42, 1122, 59, 1996, 1534, 463, 2416, 462, 677, 1977, 601, 298, 2320, 173, 18, 840, 394, 137, 202, 1271, 1824, 24, 154, 172, 1108, 120, 748]. turbots [1179]. turbulence [108]. Turchini [956]. Turkey [723]. turmeric [2075, 2518]. turtle [1700, 1752]. turtles [1012]. Tuscany [1793]. Tuticorin [713]. Two [77, 1954, 506, 953, 2414, 1619, 1445, 1067, 1286, 1119, 1215, 1913, 391, 780, 1921, 2558, 1793, 802, 2045, 752, 1164, 2738, 843, 1370, 455, 2442, 2362, 2411, 2377, 1749, 1726, 1177, 1760, 1454, 922, 1258, 1415, 1107, 2188, 172, 2145, 1976]. two-banded [1286]. Two-stage [1954, 1119]. two-year-old [455]. type [2389, 2702, 1514, 281, 732, 1287, 1842, 369, 999]. Types [510, 519, 1269, 1218, 1164, 2362, 1576, 2211]. typical [2735, 2411]. typing [86]. tyrp1a [2222]. tyrp2 [2222]. Tyrrhenian [815].

ubiquitin [2442]. UK [489, 552, 565, 553, 2239, 20]. ulceration [2042]. ulcerative [2666]. ultra [2481, 674]. ultra-freezer [2481]. ultra-violet [674]. ultrasonographic [2661]. Ultrasound [1214]. ultrastructural [2687, 2754]. ultrastructure [901]. Ultraviolet [218, 219]. Ultraviolet-A [218, 219]. Ulva [1356, 1248, 1913, 780, 1141, 2493, 2505, 1819, 1535, 514]. ulvan [1819]. Ulvella [1347]. umbra [2524]. Umbrina [1020]. Uncertainties [414]. uncertainty [2498, 2243]. uncontrolled [906]. Undaria [1103]. undatum [517, 518]. under-sized [536]. undergoing [1119, 674]. Understanding [2376, 747, 1746, 1905]. underwater [2519]. undesirable [1693]. unequals [768]. Uneven [2134]. unexercised [29]. unexplained [1969]. Unio [2388]. Union [2193]. Unique [447, 2256]. United [617]. unsaturated [934, 994]. unsuccessful [1828]. unveil [2733]. update [188]. upon [1626, 1895, 918, 1167]. upregulates [2482]. Uptake [1594, 2092, 1984]. upwelling [1703]. Urban [2696, 158, 625]. urchin [2410, 561, 1611, 1610, 1419, 2230, 184, 1864, 1274, 2055, 2503, 1767, 1559, 968, 1366, 697, 1633, 1280, 1318, 1149, 1341]. urchins [192, 399, 1650]. urea [400]. urine [1066]. urmiana [1063]. urolepis [1508]. Urophycis [629]. Uroteuthis [2676]. US$ [553]. USA [159, 1853, 278, 809, 2317, 808]. Usage [2006]. Use [69, 344, 1303, 192, 1673, 2277, 13, 196, 757, 853, 87, 829, 1777, 2611, 1452, 1036, 1089, 1366, 2661, 1545, 1915, 947, 16, 1603, 1672, 2366, 2668, 2673, 350, 842, 2158, 1307, 1521, 1002, 1725, 1598, 917, 96, 2013, 1131, 1386, 83, 1332, 494, 634, 2481, 1623, 1528, 400, 2270, 20, 835, 1892, 1414, 2264, 2330, 1466]. used [1075, 1068, 1084, 1902, 1825, 2525, 1717, 491]. useful [1197, 2349]. Using [1568, 531, 2633, 2177, 1107, 628, 483, 1203, 2718, 1479, 2713, 4, 2253, 2702, 1281, 890, 2026, 942, 1859, 796, 2155, 2078, 2336, 1743, 1402, 1740, 301, 966, 2331, 1086, 2572, 94, 833, 998, 2614, 1810, 56, 515, 1291, 388, 168, 1160, 199, 2521, 1819, 2719, 1156, 2598, 1349, 1953, 2244, 1585, 576, 1506, 433, 2030, 389, 2041, 337, 2292, 371, 2166, 2401, 1594, 1287, 790, 559, 2682, 1502, 1712, 961, 1232, 2290, 2700, 2224, 185, 1570, 1158, 1246, 581, 874, 893, 2432, 1146, 2087, 1352, 2208, 1831]. ussuriensis [1716]. Utilisation [226, 846, 1580, 312, 766, 1942, 2456]. Utilising [2191]. Utility [967]. Utilization [1907, 64, 540, 221, 1020, 2617, 799, 882, 1716, 909, 780, 1766, 1340, 1563, 1047, 1141, 2354, 1552, 1706, 1364, 2693, 1040, 1265, 2017, 1699, 550, 606, 342, 1687, 268, 1026, 1208, 90, 2340, 1200, 837, 2588, 2529, 685, 914, 940]. Utilizing [1375, 2260, 2735]. UV [1861, 1156]. UV-irradiated [1156].

V [2231, 692, 374, 2105]. V. [2483, 915]. vaccinate [87]. vaccinated [2111, 2181, 1333]. Vaccination [39, 1638, 2192, 1140, 2572, 2490, 2171, 359, 1974, 1872, 2561]. vaccine [1993, 2691, 1919, 894, 1872, 2482, 1732, 1809]. vaccines [855, 2329, 2443, 2582]. vachelli [2297]. vacuum [1920]. vacuum-packaged [1576]. vacuum-packed [1920]. Valenciennes [1792, 697, 1698]. Valenti [366, 1039]. Validation [1848, 2484, 2155, 2309]. Valle [2050]. valorise [1203]. valorization [2286]. valuable [2400, 1579, 2506, 1524]. valuation [1985]. value [2266, 2552, 1531, 430, 1529, 1923, 1081, 456, 1066, 1791, 809, 2101, 1258, 1390, 1191]. values [1869, 2348, 2202, 414, 1548, 1390]. valve [1305]. vannamei [1555, 2602, 891, 2370, 2699, 2560, 2552, 1907, 2604, 1619, 1290, 1904, 1652, 1248, 1334, 2043, 1119, 764, 1463, 1817, 1384, 1678, 2427, 1011, 2313, 2437, 1173, 1567, 2094, 2725, 2434, 2319, 1396, 1992, 1289, 1259, 1381, 1379, 2251, 1860, 2148, 1676, 1784, 1052, 883, 1538, 2684, 1023, 944, 824, 1198, 2690, 2231, 2602, 1706, 2631, 1827, 1654, 1664, 1855, 2218, 2355, 1614, 2739, 2327, 2487, 1598, 2314, 2533, 2205, 2719, 894, 1268, 525, 1324, 1822, 2092, 2472, 2576, 1460, 2322, 2203, 2608, 2316, 2518, 2694, 524, 2486, 1886, 389, 2079, 2508, 1019, 373, 2006, 2049, 1994, 2531]. vannamei [1880, 1620, 1502, 331, 2243, 2507, 763, 2667, 2295, 1452, 2000, 2594, 2100, 2585, 2102, 2188, 1302, 1036, 1877, 522, 526, 2289, 1389, 1570, 971, 1374, 1711, 2454, 1962, 891, 2018, 1331, 2723, 1966, 2480, 2397, 2432, 2655, 2546, 2549, 2039, 735, 782, 2089, 1613, 1722, 1976, 1021, 1424, 2318, 1397]. var [1595]. var. [2267, 1383, 2589, 2345, 1495, 2041]. varia [1597]. variabilis [2223]. variability [106, 2107, 583, 1737, 1943, 1107, 2674, 787]. variable [2713, 436]. variables [2104, 779, 1723]. variance [155, 617]. varians [732, 628]. Variation [81, 46, 1273, 300, 460, 84, 818, 954, 2544, 168, 2505, 1244, 1681, 884, 702, 2048, 147]. variations [2162, 92, 873]. variegatus [561, 902]. varieties [2538, 168, 1760]. variety [1269, 1335]. Various [243, 513, 64, 885, 178, 200, 2338, 2356, 1741, 1639]. Varunidae [2566]. varying [801, 1340, 623, 1312, 2596, 1200, 630]. Vegetable [789, 1109, 1366, 1029]. vegetables [1951]. vegetal [2525]. Vegetative [567]. veined [1698]. veliger [1178, 2485]. veligers [2241]. velocity [986, 1127]. Veneridae [2108]. Venerupis [772]. Venezuela [323, 401]. Venice [951]. venom [2636]. venom-mediated [2636]. venosa [1698]. ventricosus [2429]. venturi [2069, 2489]. Venus [1279]. vera [1638]. Verasper [902]. Verbenaceae [2656]. vermicompost [1795]. veronii [1986, 2567, 2666, 2579, 2075, 2582, 2639]. verrucosa [1772]. versatile [2165]. versus [1532, 520, 521, 2580, 524, 1880, 1007, 646, 2176]. vertebra [1566]. vertical [2079, 2640]. vesicle [1328]. vesiculosus [1715, 1882]. vessel [639]. vessels [2529]. veterinary [1283]. VHSV [33, 1301, 2679]. vi [1975, 544]. VI-alpha [544]. via [2418, 2261, 1809, 1544, 2749, 2113]. Viability [209, 1961, 2237, 1817, 266, 888, 886, 1004, 2544, 916, 2282, 771, 382, 1620, 1876, 1981, 1355]. viability/profitability [888, 886, 1004]. viable [1156, 1623]. vibration [2157]. vibrio [259, 2031, 1479, 2109, 2604, 1619, 635, 42, 1384, 1780, 2605, 2437, 2483, 2471, 2538, 2572, 2148, 1676, 1563, 2162, 2641, 915, 2690, 2042, 1441, 1806, 2448, 1548, 1757, 2717, 2209, 1833, 1429, 2329, 1872, 1732, 1704, 2185, 1184, 2049, 465, 739, 1087, 2462, 1493, 941, 1132, 1962, 950, 2723, 1182, 1966, 2480, 936, 1438, 2477, 1403, 2039, 2163, 2089, 2307, 1260]. vibrios [1639]. Vibriosis [2483, 2192, 1801, 2451, 946, 2643, 1872, 2216]. Vicia [1496]. vicious [787]. Video [2226]. VIE [1053]. Vietnam [751, 1648, 2381, 1787, 1378, 952, 1788, 1246, 1394, 2669]. Vietnamese [1950, 2240, 1150]. view [277]. vigour [935]. village [2214]. villages [911]. violet [674]. Viral [2138, 2347, 2125, 2474, 2387, 1301, 1974, 1605, 1797, 2116, 1868, 2550, 1705]. viremia [97]. virginalis [1370]. virginica [809, 1433, 808]. viridis [2272]. Virulence [2582, 1790, 2502, 1780, 2014, 2342, 2380, 2162, 2209, 2477]. virulent [2149]. Virus [569, 730, 1277, 1887, 1993, 2609, 2736, 1467, 2110, 1664, 1855, 2387, 2218, 1301, 2136, 1806, 894, 140, 2008, 2566, 1773, 2033, 2150, 2478, 2632, 559, 2194, 1371, 2129, 2195, 1961, 1010, 1185, 1059, 2070, 1711, 2018, 1331, 1361, 2480, 2685, 1397, 2484]. virus-infected [1664, 1855]. viruses [1476]. viscera [2558, 846]. viscosa [1993]. Visible [1053, 2253, 1624]. Visual [167, 2149, 2036]. vital [2237]. vitality [113]. Vitamin [112, 1443, 2729, 406, 1642, 2452, 2428, 1244, 525, 359, 78, 982, 436, 2652, 1657]. Vitamins [239, 359]. Vitexin [2680]. vitrification [1817]. vitro [1621, 174, 1851, 2287, 266, 2081, 301, 2471, 1821, 2460, 2544, 2621, 2390, 1936, 1866, 1616, 733, 1223, 746, 2341, 594]. vivo [255, 2471, 1821, 1263, 855, 2156, 2621, 2262, 1866, 1361, 594]. vlkianum [784]. VNN [1974]. vol [887]. volatile [1615, 1802]. Volta [1850]. Volume [170, 215, 216, 217, 262, 315, 265, 318, 263, 316, 264, 375, 377, 376, 419, 421, 420, 475, 477, 476, 347, 678, 2391]. Voluntary [960, 701]. Vp [2690]. VP28 [894, 1711]. VQV8 [2641]. vrolikii [2712]. vs [2104, 1934, 2447]. vulgare [1833]. vulgaris [903, 1286, 2294, 816, 842, 2081, 924, 907, 1004, 404, 196, 515, 1291, 1014, 1045, 2357, 900]. vulnerable [2696, 2080]. vulnificus [2572, 2042, 1184]. VZ5 [1493].

W [366, 449, 860, 956, 1039, 1204, 2261]. W-pili [2261]. W. [1603, 2303, 1282]. W.-K [956]. Wacklin [2087]. Wadden [807]. Walbaum [2722, 1568, 2226, 409, 2644, 343, 2746, 1044, 1656, 762, 2420]. wall [1928, 1455]. Wallingford [553]. walls [2386, 1981]. warmer [1997]. Waste [227, 1862, 1668, 1740, 2684, 2172, 2492, 1492, 1193, 352, 2204, 2167]. wastes [94, 125]. wastewater [2026, 2512, 1323, 1614, 2041, 2292, 2506, 2728]. wastewater-grown [1614]. Water [1248, 1334, 2385, 1952, 1376, 510, 513, 1302, 1489, 2370, 1196, 2740, 2612, 2207, 728, 523, 549, 2610, 299, 564, 764, 1253, 198, 1748, 1529, 1005, 1740, 2267, 1186, 398, 207, 358, 1202, 1567, 2094, 2725, 12, 1992, 1889, 191, 384, 1676, 729, 2158, 1380, 1587, 986, 998, 1235, 1584, 1383, 2530, 1856, 2631, 1654, 863, 864, 1470, 2183, 631, 756, 879, 1278, 2739, 1598, 104, 715, 2314, 2448, 1423, 2704, 2672, 2449, 1324, 1822, 1940, 6, 325, 1808, 298, 2516, 1147, 2101, 1497, 1685, 333, 326, 576, 619, 2121, 2316, 1297, 1741, 382, 2694, 433]. water [1832, 1131, 1279, 177, 1495, 1893, 1686, 2508, 1928, 339, 371, 1601, 2270, 1717, 2595, 1243, 644, 1597, 211, 1785, 1987, 978, 1847, 41, 1516, 2585, 121, 292, 672, 2119, 2289, 972, 1935, 1979, 44, 1666, 360, 2300, 2432, 2548, 2549, 2565, 873, 592, 1318, 1455, 1883, 2186, 1424, 1988, 2686, 2453, 2176, 2510, 2606]. water-exchange [756]. water-level [44]. water-reusing [1889]. water-soluble [433]. waterlogged [94]. waters [636, 1593, 2383, 714]. Watson [1283]. wavelengths [1624]. way [2295]. wealth [2590]. Weaning [435, 841, 694, 546, 1766, 696, 601, 913, 1332, 1165, 1466]. web [1345]. Webster [353, 563, 650]. wedge [698, 841]. Weeks [250]. weight [1018, 2107, 1634, 919, 404, 1896, 2535, 754, 1385, 1675, 884, 1702]. weissflogii [958, 965]. welfare [878, 1568, 533, 1977, 2345, 829, 1893, 532, 2724, 1718, 727]. well [1876]. wellbeing [2141]. Wellness [507]. wels [1164]. Werner [447]. West [81, 2650, 2349, 1904, 1724, 558, 1955]. Western [247, 386, 40, 426, 105, 542, 541, 295, 1656]. wet [1634]. wetland [740, 1107]. WFS [2006]. Wheat [1846, 764, 1231, 2525, 880, 1030, 2435]. whelk [517, 518, 1698]. whisker [2291]. White [232, 1904, 1487, 559, 1355, 2699, 1907, 730, 146, 145, 153, 1290, 1652, 1248, 1277, 2043, 1119, 1887, 345, 1384, 1678, 2427, 2532, 2319, 1396, 1393, 1259, 1379, 1860, 2609, 2736, 1901, 1467, 2110, 824, 1198, 2631, 1654, 1664, 1855, 2387, 2218, 1614, 2739, 2327, 2487, 2314, 894, 1329, 1268, 1737, 1324, 2472, 2008, 1460, 2322, 2097, 1485, 722, 2608, 1796, 2316, 2518, 2566, 524, 1773, 2079, 1019, 982, 373, 2006, 2049, 2551, 2596, 1082, 2150, 2632, 2703, 1502, 1961, 369, 2295, 1163, 2000, 2594, 2100, 750, 2102, 1302, 1036, 522, 526, 1348, 1185, 1059, 1389, 1711]. white [1868, 1962, 1025, 2520, 2018, 1331, 1361, 2480, 2655, 2549, 782, 1976, 1503, 1397, 569, 2484]. white-clawed [345, 369]. white-leg [2631, 2739]. white-spotted [1503]. white-striped [1025]. whitefish [180, 1154]. whiteleg [2560, 1463, 1836, 2218, 2621, 2694, 2006, 2243, 1711, 2454]. Whitmarsh [1008]. who [1513]. Whole [750, 2546, 1448, 912, 2202, 2291, 2386, 2737, 2072]. Whole-body [750, 1448, 2291, 2737]. whole-cell [2202]. Whole-genome [2546]. wholesome [1524]. wide [1171, 2674, 1404, 1949]. Wild [690, 1269, 945, 2142, 2414, 901, 155, 1344, 2555, 2638, 954, 1903, 1382, 948, 1737, 2625, 1742, 943, 269, 1368, 1506, 1177, 937, 394, 202, 61, 522, 526, 397, 1175, 1955, 646, 2021, 2120, 2182, 272]. wild-caught [1344, 2555]. wild-types [1269]. wildlife [1002]. winged [1703]. winning [533]. winter [1418, 398, 63, 123, 1454]. wintering [676]. Wistar [2619]. withdrawal [326]. Withell [879]. within [795, 2376, 607]. without [1822, 2516, 1658, 1783, 1104, 1271, 1877, 1316, 572, 1733, 2176]. Wofasteril(R) [2282]. wolffish [1826, 783, 351, 408, 190, 14, 28, 76, 44]. women [1027, 1400, 1898]. won’t [614]. Woo [449, 1056]. world [957, 2353]. worms [287]. wormwood [1989]. Would [2673]. wounding [656]. wrack [1882]. wrasse [635, 1572, 395]. Wreckfish [509]. Writing [859, 1285]. WSSV [569, 2609, 2736, 1467, 2110, 1435, 2566, 2150, 559, 1961, 1361].

WTP [1985].

xenic [757]. Xenocypris [2742]. Xenostrobus [636]. xi [553]. Xia [1253]. xii [565]. xiii [563]. Xiphophorus [1995, 2014]. XSV [1371, 1010]. xv [1008]. xvii [554]. xylooligosaccharides [1774].

Y3 [2082]. Yangtze [2120]. YB [2441, 2440]. YB-1 [2441, 2440]. year [1407, 2086, 455, 1159]. yearlings [1262]. Yeast [1981, 1812, 1033, 946, 2045, 1640, 1169, 335, 157, 2220, 1082, 465, 2262, 1517, 1807, 870, 1343, 36]. yellow [2297, 1077, 2691, 881, 2696, 763, 1949, 1488, 1733, 327, 674, 1665, 1509, 1922, 2130, 1737, 2038]. yellow-head [763]. yellowfin [2180, 685]. Yellowtail [495, 2404, 1244, 1312, 1554]. Yersinia [2471, 1916, 343, 1727]. Yesso [2002]. yessoensis [1351, 2002, 2359, 790, 1318]. yezoensis [2717]. Yield [2274, 1292, 1129, 815, 381, 1257, 1080, 988, 2681, 1819, 617, 722, 1074, 1858, 697, 2661]. Yields [613, 2381, 1199]. Yolk [221, 226, 1218, 1418, 1208, 90, 208]. yolk-sac [1418]. YOLOv5 [2649, 2703]. York [563]. Young [495, 408, 338, 1276, 468]. youth [2476]. YSP [1343]. Yuehu [1243]. yui [1777]. Yunlin [1298, 1660].

zanardini [2552]. Zanthoxylum [1595]. Zealand [810, 2547]. zearalenone [2031]. Zebrafish [2412, 1147, 1184, 2556]. zeolite [1470, 400]. Zero [523, 764, 1567, 2449, 1462]. zero-exchange [1462]. Zero-water [523, 1567, 2449]. Zeuxapta [2404, 794]. Zhang [1253]. Zhangzidao [2002]. Zhanjiang [725, 2692]. Zheng [895]. Zhikong [873]. zillii [1833]. zinc [2328]. Zingiber [2736]. zoea [935]. Zoeae [515, 1291, 2377, 1983, 2211]. Zone [416, 413]. zones [596]. zoobenthos [1584]. zoonotic [2019]. zooplankton [77, 1584, 602, 619, 428, 1104, 655, 854]. Zootechnical [2100, 1568, 2079, 2646, 2510].

References

Anonymous:1993:E

[1] Anonymous. Editorial. Aquaculture International, 1(1):1, September 1993. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic).

      URL https://link.springer.com/article/10.1007/BF00692660.

Hempel:1993:CPD

[2] E. Hempel. Constraints and possibilities for developing aquaculture. Aquaculture International, 1(1):2–19, September 1993. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

      //link.springer.com/article/10.1007/BF00692661.

Jobling:1993:FGE

[3] Malcolm Jobling, Even H. Jørgensen, Arne M. Arnesen, and Einar Ringø. Feeding, growth and environmental requirements of Arctic charr: a review of aquaculture potential. Aquaculture International, 1(1):20–46, September 1993. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/BF00692662.

Alanara:1993:TIF

[4] Anders Alanärä and Eva Brännäs. A test of the individual feeding activity and food size preference in rainbow trout using demand feeders. Aquaculture International, 1(1):47–54, September 1993. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00692663.

Cripps:1993:MQO

[5] S. J. Cripps and M. G. Poxton. A method for the quantification and optimization of hydrodynamics in culture tanks. Aquaculture International, 1(1):55–71, September 1993. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00692664.

Loland:1993:CFW

[6] Geir Løland. Current forces on, and water flow through and around, floating fish farms. Aquaculture International, 1(1):72–89, September 1993. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic).

URL https://link.springer.com/article/10.1007/BF00692665.

Jorgensen:1993:FDE

[7] Even H. Jørgensen and Malcolm Jobling. Feeding in darkness eliminates density-dependent growth suppression in Arctic charr. Aquaculture International, 1(1):90–93, September 1993. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00692666.

Jobling:1993:EPE

[8] M. Jobling, B. M. Baardvik, J. S. Christiansen, and E. H. Jørgensen. The effects of prolonged exercise training on growth performance and production parameters in fish. Aquaculture International, 1(2):95–111, December 1993. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/BF00692614.

Carlstein:1993:NFA

[9] Mikael Carlstein. Natural food and artificial, dry starter diets: effects on growth and survival in intensively reared European grayling. Aquaculture International, 1(2):112–123, December 1993. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00692615.

Makinen:1993:EFS

[10] Timo Mäkinen. Effect of feeding schedule on growth of rainbow trout. Aquaculture International, 1(2):124–136, December 1993. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00692616.

Leatherland:1993:SDC

[11] J. F. Leatherland. Stocking density and cohort sampling effects on endocrine interactions in rainbow trout. Aquaculture International, 1(2): 137–156, December 1993. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00692617.

Fivelstad:1993:SEA

[12] Sveinung Fivelstad, Harald Kallevik, Helge M. Iversen, Trond Møretrø, Klement Vøge, and Martin Binde. Sublethal effects of ammonia in soft water on Atlantic salmon smolts at a low temperature. Aquaculture International, 1(2):157–169, December 1993. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00692618.

Dhert:1993:UOA

[13] Ph. Dhert, R. B. Bombeo, and P. Sorgeloos. Use of ongrown Artemia in nursery culturing of the tiger shrimp. Aquaculture International, 1(2): 170–177, December 1993. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00692619.

Pavlov:1993:BDE

[14] D. A. Pavlov and E. Moksness. Bacterial destruction of the egg shell of common wolffish during incubation. Aquaculture International, 1(2): 178–186, December 1993. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00692620.

Naegel:1994:DIA

[15] Ludwig C. A. Naegel. Development of an integrated aquafarming system. Aquaculture International, 2(1):1–9, March 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00118528.

Bruno:1994:SUH

[16] D. W. Bruno and R. S. Raynard. Studies on the use of hydrogen peroxide as a method for the control of sea lice on Atlantic salmon. Aquaculture International, 2(1):10–18, March 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00118529.

Rainuzzo:1994:ESL

[17] Jose R. Rainuzzo, Kjell I. Reitan, and Yngvar Olsen. Effect of shortand long-term lipid enrichment on total lipids, lipid class and fatty acid composition in rotifers. Aquaculture International, 2(1):19–32, March 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic).

URL https://link.springer.com/article/10.1007/BF00118530.

Reitan:1994:ILC

[18] Kjell Inge Reitan, Jose R. Rainuzzo, and Yngvar Olsen. Influence of lipid composition of live feed on growth, survival and pigmentation of turbot larvae. Aquaculture International, 2(1):33–48, March 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/BF00118531.

Sangpradub:1994:EDF

[19] Sakon Sangpradub, Arlo W. Fast, Somkiat Piyatiratitivorakul, and Piamsak Menasveta. Effects of different feeding regimes on ovarian maturation and spawning of pond-reared giant tiger prawn in Thailand. Aquaculture International, 2(1):49–58, March 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00118532.

Priyan:1994:EIA

[20] Deva Priyan and Hance Smith. Environmental impact assessment as a management tool for the use of chemicals for fish farming in the UK. Aquaculture International, 2(1):59–64, March 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00118533.

Anonymous:1994:BRNa

[21] Anonymous. Book reviews and notices. Aquaculture International, 2 (1):65–74, March 1994. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00118534.

Jobling:1994:CGR

[22] M. Jobling, O. H. Meløy, J. dos Santos, and B. Christiansen. The compensatory growth response of the Atlantic cod: effects of nutritional history. Aquaculture International, 2(2):75–90, June 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/BF00128802.

Gibbs:1994:APF

[23] Juliette Gibbs, Susan A. Shaw, and Mark Gabbott. An analysis of price formation in the Dutch mussel industry. Aquaculture International, 2 (2):91–103, June 1994. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00128803.

Staurnes:1994:EAT

[24] Magne Staurnes. Effects of acute temperature decreases on turbot fry and juveniles. Aquaculture International, 2(2):104–113, June 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/BF00128804.

Arnesen:1994:FGR

[25] A. M. Arnesen, E. H. Jørgensen, and M. Jobling. Feed-growth relationships of Arctic charr transferred from freshwater to saltwater at different seasons. Aquaculture International, 2(2):114–122, June 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/BF00128805.

Lu:1994:ISR

[26] Y. Lu and P. C. Loh. Infectivity studies of rhabdovirus in the penaeid blue shrimp. Aquaculture International, 2(2):123–127, June 1994. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00128806.

Jobling:1994:BRN

[27] Malcolm Jobling. Book reviews and notices. Aquaculture International, 2(2):128–132, June 1994. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00128807.

Pavlov:1994:RBE

[28] D. A. Pavlov and E. Moksness. Reproductive biology, early ontogeny, and effect of temperature on development in wolffish: comparison with salmon. Aquaculture International, 2(3):133–153, September 1994. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00231511.

Jorgensen:1994:FGE

[29] Even H. Jørgensen and Malcolm Jobling. Feeding and growth of exercised and unexercised juvenile Atlantic salmon in freshwater, and performance after transfer to seawater. Aquaculture International, 2(3): 154–164, September 1994. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00231512.

Reddy:1994:EDM

[30] P. K. Reddy, J. F. Leatherland, M. N. Khan, and T. Boujard. Effect of the daily meal time on the growth of rainbow trout fed different ration levels. Aquaculture International, 2(3):165–179, September 1994. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00231513.

Forsberg:1994:MOC

[31] Odd Inge Forsberg. Modelling oxygen consumption rates of postsmolt Atlantic salmon in commercial-scale, land-based farms. Aquaculture International, 2(3):180–196, September 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00231514.

Varvarigos:1994:PEE

[32] Panagiotis Varvarigos. Production economics of the emerging private fish farming industry in Bulgaria. Aquaculture International, 2(3):197– 205, September 1994. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00231515.

Branas:1994:SED

[33] M. Vázquez Brañas, J. Coll Morales, and A. Estepa. A sandwich ELISA to detect VHSV and IPNV in turbot. Aquaculture International, 2(3): 206–217, September 1994. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00231516.

Anonymous:1994:BRNb

[34] Anonymous. Book reviews and notices. Aquaculture International, 2(3): 218–223, September 1994. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00231517.

Anonymous:1994:ED

[35] Anonymous. From the editor’s desk. Aquaculture International, 2(4): iii–iv, December 1994. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00123430.

oie:1994:CRC

[36] G. øie, K. I. Reitan, and Y. Olsen. Comparison of rotifer culture quality with yeast plus oil and algal-based cultivation diets. Aquaculture International, 2(4):225–238, December 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00123431.

Luzzana:1994:EEF

[37] Umberto Luzzana, Giovanni Serrini, Vittorio Maria Moretti, Corrado Gianesini, and Franco Valfrè. Effect of expanded feed with high fish oil content on growth and fatty acid composition of rainbow trout. Aquaculture International, 2(4):239–248, December 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00123432.

Cecchini:1994:ITH

[38] Stefano Cecchini. Influence of temperature on the hatching of eggs of Diplectanum aequans, a parasite of sea bass. Aquaculture International, 2 (4):249–253, December 1994. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00123433.

Chair:1994:VGE

[39] M. Chair, R. S. J. Gapasin, M. Dehasque, and P. Sorgeloos. Vaccination of European sea bass fry through bioencapsulation of Artemia nauplii. Aquaculture International, 2(4):254–261, December 1994. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00123434.

Crespo:1994:ICG

[40] S. Crespo, A. Grau, and F. Padrós. The intensive culture of 0-group amberjack in the western Mediterranean is compromised by disease problems. Aquaculture International, 2(4):262–265, December 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/BF00123435.

Schuster:1994:EFM

[41] C. Schuster. The effect of fish meal content in trout food on water colour in a closed recirculating aquaculture system. Aquaculture International, 2 (4):266–269, December 1994. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00123436.

Chair:1994:OCT

[42] M. Chair, M. Dehasque, S. Van Poucke, H. Nelis, P. Sorgeloos, and A. P. De Leenheer. An oral challenge for turbot larvae with Vibrio anguillarum. Aquaculture International, 2(4):270–272, December 1994. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00123437.

Anonymous:1994:BRNc

[43] Anonymous. Book reviews and notices. Aquaculture International, 2(4): 273–278, December 1994. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00123438.

Strand:1995:FFC

[44] Hans K. Strand, Tove K. Hansen, Arvid Pedersen, Inger Britt Falk-Petersen, and Victor Øiestad. First feeding of common wolffish on formulated dry feed diets in a low water-level raceway system. Aquaculture International, 3(1):1–10, March 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00240916.

Chellappa:1995:GPG

[45] S. Chellappa, N. T. Chellappa, W. B. Barbosa, F. A. Huntingford, and M. C. M. Beveridge. Growth and production of the Amazonian tambaqui in fixed cages under different feeding regimes. Aquaculture International, 3(1):11–21, March 1995. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00240917.

Olsen:1995:VCF

[46] Yngvar Olsen and Harald Skjervold. Variation in content of Ω3 fatty acids in farmed Atlantic salmon, with special emphasis on effects of non-dietary factors. Aquaculture International, 3(1):22–35, March 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/BF00240918.

Pereira:1995:GRT

[47] J. óscar Pereira and Emídio F. Gomes. Growth of rainbow trout fed a diet supplemented with earthworms, after chemical treatment. Aquaculture International, 3(1):36–42, March 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00240919.

Macaranas:1995:GIF

[48] Julie M. Macaranas, Liza Q. Agustin, Ma.Carmen A. Ablan, Ma.Josefa R. Pante, Ambekar A. Eknath, and Roger S. V. Pullin. Genetic improvement of farmed tilapias: Biochemical characterization of strain differences in Nile tilapia. Aquaculture International, 3(1):43–54, March 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00240920.

Forsberg:1995:OCP

[49] Odd Inge Forsberg. Oxygen consumption of post-smolt Atlantic salmon during crowding and handling stress. Aquaculture International, 3(1): 55–59, March 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/BF00240921.

Jobling:1995:SIA

[50] Malcolm Jobling. Simple indices for the assessment of the influences of social environment on growth performance, exemplified by studies on Arctic charr. Aquaculture International, 3(1):60–65, March 1995. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00240922.

Jobling:1995:BRN

[51] Malcolm Jobling, Victor øiestad, Clive Talbot, and Chris Carter. Book reviews and notices. Aquaculture International, 3(1):66–75, March 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00240923.

Poxton:1995:E

[52] Michael G. Poxton. Editorial. Aquaculture International, 3(2):iii–iv, June 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic).

URL https://link.springer.com/article/10.1007/BF00117875.

Baras:1995:FTA

[53] Etienne Baras and Jean-Paul Lagardère. Fish telemetry in aquaculture: review and perspectives. Aquaculture International, 3(2):77–102, June 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic).

URL https://link.springer.com/article/10.1007/BF00117876.

Aguilar-Manjarrez:1995:GIS

[54] José Aguilar-Manjarrez and Lindsay G. Ross. Geographical information system (GIS) environmental models for aquaculture development in Sinaloa State, Mexico. Aquaculture International, 3(2):103–115, June 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic).

URL https://link.springer.com/article/10.1007/BF00117877.

Milstein:1995:SSF

[55] A. Milstein, M. Zoran, and H. J. Krambeck. Seasonal stratification in fish culture and irrigation reservoirs: potential dangers for fish culture. Aquaculture International, 3(2):116–122, June 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00117878.

Hussenot:1995:AQP

[56] J. Hussenot and J.-L. M. Martin. Assessment of the quality of pond sediment in aquaculture using simple, rapid techniques. Aquaculture International, 3(2):123–133, June 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00117879.

Priour:1995:TRS

[57] Daniel Priour and Yves Degres. A tension-recording sensor for mooring lines. Aquaculture International, 3(2):134–142, June 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/BF00117880.

Koumoundouros:1995:MRC

[58] G. Koumoundouros, Z. Kiriakos, P. Divanach, and M. Kentouri. Morphometric relationships as criteria for the evaluation of larval quality of gilthead sea bream. Aquaculture International, 3(2):143–149, June 1995.

CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00117881.

Gatesoupe:1995:MEA

[59] F. J. Gatesoupe. A method for the early assessment of the quality of turbot larvae. Aquaculture International, 3(2):150–154, June 1995. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00117882.

Salvesen:1995:SDE

[60] I. Salvesen and O. Vadstein. Surface disinfection of eggs from marine fish: evaluation of four chemicals. Aquaculture International, 3(3): 155–171, September 1995. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00118098.

Sheikh-Eldin:1995:PCP

[61] Melika Sheikh-Eldin, Sena S. De Silva, Trevor A. Anderson, and Geoff Gooley. Physical characteristics, and proximate composition of oocytes, liver and muscle of wild caught and tank-reared Macquarie perch. Aquaculture International, 3(3):172–185, September 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00118099.

Anras:1995:DFB

[62] Marie-Laure Begout Anras. Demand-feeding behaviour of sea bass kept in ponds: diel and seasonal patterns, and influences of environmental factors. Aquaculture International, 3(3):186–195, September 1995. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00118100.

Duston:1995:IWT

[63] J. Duston and R. L. Saunders. Increased winter temperature did not affect completion of smolting in Atlantic salmon. Aquaculture International, 3(3):196–204, September 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00118101.

Choubert:1995:UCV

[64] Georges Choubert, Josø-Carlos G. Milicua, Ramon Gomez, Sophie Sancø, Høløne Petit, Geneviøve Nøgre-Sadargues, Renø Castillo, and Jean-Paul Trilles. Utilization of carotenoids from various sources by rainbow trout: muscle colour, carotenoid digestibility and retention. Aquaculture International, 3(3):205–216, September 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00118102.

Kaiser:1995:ESD

[65] H. Kaiser, O. Weyl, and T. Hecht. The effect of stocking density on growth, survival and agonistic behaviour of African catfish. Aquaculture International, 3(3):217–225, September 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00118103.

Ijzerman:1995:QCI

[66] H. C. A. Ijzerman, J. P. Hoogland, J. H. Boon, and W. de Wit. Quality costs in intensive fish culture: an analysis of African catfish farms in The Netherlands. Aquaculture International, 3(3):226–235, September 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00118104.

Romdhane:1995:EFH

[67] M. S. Romdhane, B. Devresse, Ph. Løger, and P. Sorgeloos. Effects of feeding (ω-3) HUFA-enriched Artemia during a progressively increasing period on the larviculture of freshwater prawns. Aquaculture International, 3(3):236–242, September 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00118105.

Alam:1995:RAM

[68] M. J. Alam, K. J. Ang, and M. Begum. Replacement of Artemia with Moina micrura in the rearing of freshwater shrimp larvae. Aquaculture International, 3(3):243–248, September 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00118106.

Alam:1995:UEC

[69] M. J. Alam, K. J. Ang, and M. Begum. Use of egg custard augmented with cod liver oil and Moina micrura on production of freshwater prawn postlarvae. Aquaculture International, 3(3):249–259, September 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00118107.

Carlstein:1995:GSG

[70] Mikael Carlstein. Growth and survival of European grayling reared at different stocking densities. Aquaculture International, 3(3):260– 264, September 1995. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00118108.

Bergheim:1995:RFM

[71] Asbjørn Bergheim and Harald Sveier. Replacement of fish meal in salmonid diets by soya meal reduces phosphorus excretion. Aquaculture International, 3(3):265–268, September 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00118109.

Anonymous:1995:BRN

[72] Anonymous. Book reviews and notices. Aquaculture International, 3(3): 269–274, September 1995. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00118110.

Anonymous:1995:ED

[73] Anonymous. From the editor’s desk. Aquaculture International, 3(4): 275–276, December 1995. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00121618.

Naegel:1995:RFS

[74] Ludwig C. A. Naegel. Research with a farming systems perspective needed for the development of small-scale aquaculture in nonindustrialized countries. Aquaculture International, 3(4):277–291, December 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
BF00121619.

Milstein:1995:FMR

[75] Ana Milstein. Fish-management relationships in Israeli commercial fish farming. Aquaculture International, 3(4):292–314, December 1995. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00121620.

Pavlov:1995:DWE

[76] D. A. Pavlov and E. Moksness. Development of wolffish eggs at different temperature regimes. Aquaculture International, 3(4):315–335, December 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/BF00121621.

Andersson:1995:TSC

[77] Jan Andersson. Two systems for collection of zooplankton for the culture of larval and juvenile fish. Aquaculture International, 3(4): 336–354, December 1995. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00121622.

Merchie:1995:EVC

[78] G. Merchie, P. Lavens, J. Radull, H. Nelis, A. De Leenheer, and P. Sorgeloos. Evaluation of vitamin c-enriched Artemia nauplii for larvae of the giant freshwater prawn. Aquaculture International, 3(4): 355–363, December 1995. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00121623.

Porez:1995:CEE

[79] M. C. Pørez, M. L. Gonzølez, D. A. Løpez, and J. Zøøiga. Cultivation of the erizo: an evaluation of eggs and postmetamorphic juveniles size selection. Aquaculture International, 3(4):364–369, December 1995. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00121624.

Miao:1995:PER

[80] Sha Miao and Shunchi Tu. Profitability evaluation of resource allocation for growing redtail and Chinese shrimp. Aquaculture International, 3(4): 370–378, December 1995. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00121625.

Costelloe:1995:VSL

[81] J. Costelloe, M. Costelloe, and N. Roche. Variation in sea lice infestation on Atlantic salmon smolts in Killary Harbour, West Coast of Ireland. Aquaculture International, 3(4):379–393, December 1995. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/BF00121626.

Uglem:1996:CTL

[82] Ingebrigt Uglem, Lars Elias Uksnøy, and øivind Bergh. Chemical treatment of lobster eggs against epibiotic bacteria. Aquaculture International, 4(1):1–8, March 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00175216.

Osman:1996:ULL

[83] M. F. Osman, A. Eglal Omar, and A. M. Nour. The use of leucaena leaf meal in feeding Nile tilapia. Aquaculture International, 4(1):9–18, March 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic).

URL https://link.springer.com/article/10.1007/BF00175217.

Dinesh:1996:GVI

[84] K. R. Dinesh, T. M. Lim, W. K. Chan, and V. P. E. Phang. Genetic variation inferred from RAPD fingerprinting in three species of tilapia. Aquaculture International, 4(1):19–30, March 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00175218.

Tort:1996:SHA

[85] Lluis Tort, Eladi Gømez, Daniel Montero, and J. Oriol Sunyer. Serum haemolytic and agglutinating activity as indicators of fish immunocompetence: their suitability in stress and dietary studies. Aquaculture International, 4(1):31–41, March 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00175219.

Inglis:1996:ADR

[86] V. Inglis, D. Colquhoun, M. D. Pearson, M. Miyata, and T. Aoki. Analysis of DNA relationships among Aeromonas species by RAPD (randomly amplified polymorphic DNA) typing. Aquaculture International, 4(1):43–53, March 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00175220.

Marsden:1996:UGA

[87] M. J. Marsden, A. Devoy, L. M. Vaughan, T. J. Foster, and C. J. Secombes. Use of a genetically attenuated strain of Aeromonas salmonicida to vaccinate salmonid fish. Aquaculture International, 4(1):55–66, March 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic).

URL https://link.springer.com/article/10.1007/BF00175221.

Cano:1996:GGE

[88] Juana Cano and José Rocamora. Growth of the European flat oyster in the Mediterranean Sea (Murcia, SE Spain). Aquaculture International, 4(1):67–84, March 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00175222.

Yoshimatsu:1996:EDR

[89] Takao Yoshimatsu and Chikara Kitajima. Effects of daily ration and feeding frequency of Artemia on the growth of mullet larvae. Aquaculture International, 4(1):85–88, March 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00175223.

Peterson:1996:ITS

[90] R. H. Peterson, D. J. Martin-Robichaud, and O. Berge. Influence of temperature and salinity on length and yolk utilization of striped bass larvae. Aquaculture International, 4(2):89–103, June 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00140591.

Naslund:1996:GFL

[91] Ingemar Näslund and Jan Henricson. Growth of five landlocked Arctic charr stocks under hatchery conditions. Aquaculture International, 4 (2):105–116, June 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00140592.

Reis-Henriques:1996:GKA

[92] M. A. Reis-Henriques, L. Silva, and J. Coimbra. Gill Na^{+}-K^{+} AT-Pase, carbonic anhydrase activities and plasma osmotic and ionic variations during smoltification of Atlantic salmon in the north of Portugal. Aquaculture International, 4(2):117–128, June 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00140593.

Andrade:1996:GEF

[93] José Pedro Andrade, Karim Erzini, and Jorge Palma. Gastric evacuation and feeding in the gilthead sea bream reared under semi-intensive conditions. Aquaculture International, 4(2):129–141, June 1996. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00140594.

Garg:1996:BCC

[94] S. K. Garg. Brackishwater carp culture in potentially waterlogged areas using animal wastes as pond fertilizers. Aquaculture International, 4 (2):143–155, June 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00140595.

Munsiri:1996:TCC

[95] Prasert Munsiri, Claude E. Boyd, David Teichert-Coddington, and Ben F. Hajek. Texture and chemical composition of soils from shrimp ponds near Choluteca, Honduras. Aquaculture International, 4(2):157– 168, June 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/BF00140596.

Meyer:1996:AUL

[96] Walter F. Meyer and Peter A. Cook. An assessment of the use of low-level aerobic swimming in promoting recovery from handling stress in rainbow trout. Aquaculture International, 4(2):169–174, June 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/BF00140597.

rehulka:1996:BPC

[97] J. řehulka. Blood parameters in common carp with spontaneous spring viremia (SVC). Aquaculture International, 4(2):175–182, June 1996. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00140598.

Beveridge:1996:BRN

[98] M. Beveridge, M. Jobling, Colin E. Adams, and R. Williamson. Book review and notices. Aquaculture International, 4(2):183–188, June 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00140599.

Poxton:1996:E

[99] Michael G. Poxton. Editorial. Aquaculture International, 4(3):189, September 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/BF00117380.

Soudant:1996:FAE

[100] Philippe Soudant, Yanic Marty, Jeanne Moal, and Jean François Samain. Fatty acids and egg quality in great scallop. Aquaculture International, 4 (3):191–200, September 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00117381.

Pazos:1996:IGC

[101] A. J. Pazos, G. Romøn, C. P. Acosta, M. Abad, and J. L. Sønchez. Influence of the gametogenic cycle on the biochemical composition of the ovary of the great scallop. Aquaculture International, 4(3):201–213, September 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic).

URL https://link.springer.com/article/10.1007/BF00117382.

Moal:1996:BNG

[102] J. Moal, J. F. Samain, S. Corre, J. L. Nicolas, and A. Glynn. Bacterial nutrition of great scallop larvae. Aquaculture International, 4(3): 215–223, September 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00117383.

Roman:1996:RAE

[103] G. Román, M. J. Campos, and C. P. Acosta. Relationships among environment, spawning and settlement of Queen scallop in the Ría de Arosa (Galicia, NW Spain). Aquaculture International, 4(3):225–236, September 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/BF00117384.

Larvor:1996:AMM

[104] H. Larvor, J. P. Cuif, and N. Devauchelle. Abnormal melanization and microstructural distortions of the shell of great scallop living in shallow water. Aquaculture International, 4(3):237–252, September 1996. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00117385.

Peoa:1996:SSS

[105] J. B. Peøa, J. Canales, J. M. Adsuara, and M. A. Sos. Study of seasonal settlements of five scallop species in the western Mediterranean. Aquaculture International, 4(3):253–261, September 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00117386.

Chauvaud:1996:ECG

[106] Laurent Chauvaud, Gérard Thouzeau, and Jacques Grall. Experimental collection of great scallop postlarvae and other benthic species in the Bay of Brest: settlement patterns in relation to spatio-temporal variability of environmental factors. Aquaculture International, 4(3): 263–288, September 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00117387.

Ruiz:1996:MBS

[107] C. M. Ruiz, G. Román, and J. L. Sánchez. A marine bacterial strain effective in producing antagonisms of other bacteria. Aquaculture International, 4(3):289–291, September 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00117388.

Nicolas:1996:EES

[108] L. Nicolas, R. Robert, and L. Chevolot. Effect of epinephrine and seawater turbulence on the metamorphosis of the great scallop. Aquaculture International, 4(3):293–297, September 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF00117389.

Barret:1996:PRR

[109] J. Barret, C. Mingant, and R. Robert. Preliminary results of rearing Japanese scallop in Atlantic conditions. Aquaculture International, 4(3): 299–302, September 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00117390.

Anonymous:1996:ED

[110] Anonymous. From the editor’s desk. Aquaculture International, 4 (4):303, December 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00120946.

Robert:1996:MCS

[111] R. Robert, P. Miner, and J. L. Nicolas. Mortality control of scallop larvae in the hatchery. Aquaculture International, 4(4):305–313, December 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic).

URL https://link.springer.com/article/10.1007/BF00120947.

Seguineau:1996:VRG

[112] C. Seguineau, A. Laschi-Loquerie, J. Moal, and J. F. Samain. Vitamin requirements in great scallop larvae. Aquaculture International, 4(4): 315–324, December 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00120948.

Fleury:1996:PSB

[113] P. G. Fleury, C. Mingant, and A. Castillo. A preliminary study of the behaviour and vitality of reseeded juvenile great scallops, of three sizes in three seasons. Aquaculture International, 4(4):325–337, December 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF00120949.

Jones:1996:ERA

[114] Paul L. Jones, Sena S. De Silva, and Brad D. Mitchell. Effects of replacement of animal protein by soybean meal on growth and carcass composition in juvenile Australian freshwater crayfish. Aquaculture International, 4(4):339–359, December 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF00120950.

Jones:1996:EDP

[115] Paul L. Jones, Sena S. De Silva, and Brad D. Mitchell. The effect of dietary protein source on growth and carcass composition in juvenile Australian freshwater crayfish. Aquaculture International, 4(4): 361–376, December 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00120951.

Shanmugasundaram:1996:TAC

[116] G. K. Shanmugasundaram, V. Ramasubramanian, and N. Munuswamy. α-Tocopherol in Artemia cysts: a report. Aquaculture International, 4 (4):377–378, December 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00120952.

Anonymous:1996:BRN

[117] Anonymous. Book reviews and notices. Aquaculture International, 4(4): 379–389, December 1996. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF00120953.

Anonymous:1996:C

[118] Anonymous. Corrigendum. Aquaculture International, 4(4):390, December 1996. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
BF00120954.

Poxton:1997:ED

[119] Michael Poxton. From the editor’s desk. Aquaculture International, 5(1):i, January 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF02764782.

Waring:1997:PRT

[120] Colin P. Waring, Michael G. Poxton, and Ronald M. Stagg. The physiological response of the turbot to multiple net confinements. Aquaculture International, 5(1):1–12, January 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/BF02764783.

Skjermo:1997:MMW

[121] J. Skjermo, I. Salvesen, G. Øie, Y. Olsen, and O. Vadstein. Microbially matured water: a technique for selection of a non-opportunistic bacterial flora in water that may improve performance of marine larvae. Aquaculture International, 5(1):13–28, January 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF02764784.

Heath:1997:RDS

[122] Philip L. Heath and Colin G. Moore. Rearing dover sole larvae on Tisbe and Artemia diets. Aquaculture International, 5(1):29–39, January 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF02764785.

Kadri:1997:EMP

[123] Sunil Kadri, Neil B. Metcalfe, Felicity A. Huntingford, John E. Thorpe, and David F. Mitchell. Early morphological predictors of maturity in one-sea-winter Atlantic salmon. Aquaculture International, 5(1):41– 50, January 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/BF02764786.

Rothbard:1997:CSM

[124] Shmuel Rothbard, William L. Shelton, Zeev Kulikovsky, Israel Rubinshtein, Yair Hagani, and Boaz Moav. Chromosome set manipulations in the black carp. Aquaculture International, 5(1):51–64, January 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/BF02764787.

Kelly:1997:PSD

[125] L. A. Kelly, A. Bergheim, and J. Stellwagen. Particle size distribution of wastes from freshwater fish farms. Aquaculture International, 5(1):65–78, January 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF02764788.

Pozdnyakova:1997:ASD

[126] Ludmila A. Pozdnyakova, Alla V. Silina, and George A. Evseev. Age, size distribution and growth of native and cultured Japanese scallops in Possjet Bay, Sea of Japan, Russia. Aquaculture International, 5 (1):79–88, January 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF02764789.

Shumway:1997:PCS

[127] S. E. Shumway, T. L. Cucci, M. P. Lesser, N. Bourne, and B. Bunting. Particle clearance and selection in three species of juvenile scallops. Aquaculture International, 5(1):89–99, January 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/BF02764790.

Adeyemo:1997:PAI

[128] A. A. Adeyemo, A. F. Yakubu, G. A. Oladosu, and O. A. Ayinla. Predation by aquatic insects on African catfish fry. Aquaculture International, 5(1):101–103, January 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/BF02764791.

Dhert:1997:PDE

[129] Ph. Dhert, K. Schoeters, P. Vermeulen, J. Sun, S. Gao, Z. Shang, X. Naihong, H. Van Duffel, and P. Sorgeloos. Production, disinfection and evaluation for aquaculture applications of rotifer resting eggs from Bohai Bay, P. R. of China. Aquaculture International, 5(2):0, March 1997. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1018331200146.

Anonymous:1997:A

[130] Anonymous. Announcement. Aquaculture International, 5(2):1, March 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1018327116076.

Anonymous:1997:Ea

[131] Anonymous. Editorial. Aquaculture International, 5(2):1, March 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1018313612383.

Gjedrem:1997:FQI

[132] Trygve Gjedrem. Flesh quality improvement in fish through breeding. Aquaculture International, 5(3):197–206, May 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1014546816984.

Siddiqui:1997:ESR

[133] A. Q. Siddiqui and A. H. Al-Harbi. Effects of sex ratio, stocking density and age of hybrid tilapia on seed production in concrete tanks in Saudi Arabia. Aquaculture International, 5(3):207–216, May 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1018383201054.

Fuentes:1997:PTS

[134] J. Fuentes, J. L. Soengas, P. Rey, and E. Rebolledo. Progressive transfer to seawater enhances intestinal and branchial Na+-K+-ATPase activity in non-anadromous rainbow trout. Aquaculture International, 5(3):217– 227, May 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1018387317893.

Tulli:1997:CAA

[135] F. Tulli and E. Tibaldi. Changes in amino acids and essential fatty acids during early larval rearing of dentex. Aquaculture International, 5(3):229–236, May 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1018339401963.

Pavlidis:1997:DPH

[136] M. Pavlidis, M. Berry, P. Divanach, and M. Kentouri. Diel pattern of haematocrit, serum metabolites, osmotic pressure, electrolytes and thyroid hormones in sea bass and sea bream. Aquaculture International, 5(3):237–247, May 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1018391418801.

Salvesen:1997:SDG

[137] I. Salvesen, G. Oie, and O. Vadstein. Surface disinfection of Atlantic halibut and turbot eggs with glutaraldehyde: evaluation of concentrations and contact times. Aquaculture International, 5(3):249–258, May 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1018343602872.

Cliche:1997:SCS

[138] G. Cliche, S. Vigneau, and M. Giguere. Status of a commercial sea scallop enhancement project in Îles-de-la-Madeleine (Québec, Canada). Aquaculture International, 5(3):259–266, May 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1018395619710.

Panha:1997:MTF

[139] Somsak Panha and Prasuk Kosavititkul. Mantle transplantations in freshwater pearl mussels in Thailand. Aquaculture International, 5 (3):267–276, May 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1018347703780.

Lu:1997:SCP

[140] Y. Lu, L. M. Tapay, P. C. Loh, R. B. Gose, and J. A. Brock. Short communication: The pathogenicity of a baculo-like virus isolated from diseased penaeid shrimp obtained from China for cultured penaeid species in Hawaii. Aquaculture International, 5(3):277–282, May 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1018399720619.

Anonymous:1997:BRNa

[141] Anonymous. Book reviews and notices. Aquaculture International, 5 (3):283–288, May 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1018351804689.

Anonymous:1997:Eb

[142] Anonymous. Erratum. Aquaculture International, 5(3):289, May 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1018303921527.

Smolowitz:1997:PCE

[143] Roxanna Smolowitz and Sandra E. Shumway. Possible cytotoxic effects of the dinoflagellate, Gyrodinium aureolum, on juvenile bivalve molluscs.

Aquaculture International, 5(4):291–300, July 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1018355905598.

DeCasabianca:1997:ISF

[144] M.-L. De Casabianca, T. Laugier, and D. Collart. Impact of shellfish farming eutrophication on benthic macrophyte communities in the Thau lagoon, France. Aquaculture International, 5(4):301–314, July 1997. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1018308022436.

Bandeen:1997:ECC

[145] Janice Bandeen and John F. Leatherland. Evaluation of commercial catfish, tilapia and salmonid diets for growth promotion of white suckers. Aquaculture International, 5(4):315–326, July 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1018360006506.

Bandeen:1997:CPC

[146] Janice Bandeen and John F. Leatherland. Changes in the proximate composition of juvenile white suckers following re-feeding after a prolonged fast. Aquaculture International, 5(4):327–337, July 1997. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1018312123344.

deSilva:1997:GVT

[147] C. D. de Silva. Genetic variation in tilapia populations in manmade reservoirs in Sri Lanka. Aquaculture International, 5(4):339–349, July 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1018364107415.

Koskela:1997:FIG

[148] J. Koskela, J. Pirhonen, and M. Jobling. Feed intake, growth rate and body composition of juvenile Baltic salmon exposed to different constant temperatures. Aquaculture International, 5(4):351–360, July 1997. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1018316224253.

Carvalho:1997:FFC

[149] A. P. Carvalho, A.-M. Escaffre, A. Oliva Teles, and P. Bergot. First feeding of common carp larvae on diets with high levels of protein hydrolysates. Aquaculture International, 5(4):361–367, July 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1018368208323.

Alok:1997:EDL

[150] Deoraj Alok, Devika Pillai, and Lalit C. Garg. Effect of d-lys6 salmon sGnRH alone and in combination with domperidone on the spawning of common carp during the late spawning season. Aquaculture International, 5(4):369–374, July 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1023/A:1018320325162.

Kouril:1997:SC

[151] J. Kouril, O. Linhart, and P. Relot. Short communication. Aquaculture International, 5(4):375–377, July 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1018372309232.

Anonymous:1997:BRNb

[152] Anonymous. Book reviews and notices. Aquaculture International, 5 (4):379–383, July 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1018324526070.

Bandeen:1997:THS

[153] Janice Bandeen and John F. Leatherland. Transportation and handling stress of white suckers raised in cages. Aquaculture International, 5(5): 385–396, September 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1018376510141.

Strand:1997:GEG

[154] Hans K. Strand and Victor Oiestad. Growth and the effect of grading, of turbot in a shallow raceway system. Aquaculture International, 5(5): 397–406, September 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1018328626979.

Crozier:1997:GHM

[155] W. W. Crozier. Genetic heterozygosity and meristic character variance in a wild Atlantic salmon population and a hatchery strain derived from it. Aquaculture International, 5(5):407–414, September 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1018380611049.

Requena:1997:ETR

[156] A. Requena, J. Fernandez-Borras, and J. Planas. The effects of a temperature rise on oxygen consumption and energy budget in gilthead sea bream. Aquaculture International, 5(5):415–426, September 1997. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1018332727888.

Lie:1997:NCR

[157] Oyvind Lie, Herborg Haaland, Gro-Ingunn Hemre, Amund Maage, Einar Lied, Grethe Rosenlund, Kjartan Sandnes, and Yngvar Olsen. Nutritional composition of rotifers following a change in diet from yeast and emulsified oil to microalgae. Aquaculture International, 5(5):427–438, September 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1018384711958.

Lu:1997:CSB

[158] Yantian Lu and Norman J. Blake. The culture of the southern bay scallop in Tampa Bay, an urban Florida estuary. Aquaculture International, 5(5): 439–450, September 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1018336828796.

Barber:1997:GMC

[159] Bruce J. Barber and Christopher V. Davis. Growth and mortality of cultured bay scallops in the Damariscotta River, Maine (USA). Aquaculture International, 5(5):451–460, September 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1018388812866.

Diaz:1997:SSP

[160] M. Diaz, F. J. Moyano, F. L. Garcia-Carreno, F. J. Alarcon, and M. C. Sarasquete. Substrate-SDS-PAGE determination of protease activity through larval development in sea bream. Aquaculture International, 5 (5):461–471, September 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1018340929705.

Cerda:1997:SCE

[161] Joan Cerda, Silvia Zanuy, and Manuel Carrillo. Short communication: Evidence for dietary effects on plasma levels of sexual steroids during spermatogenesis in the sea bass. Aquaculture International, 5(5):473–477, September 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1018392913775.

Anonymous:1997:ED

[162] Anonymous. From the editors desk. Aquaculture International, 5(6):1, November 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1018345030613.

Koskela:1997:ELT

[163] J. Koskela, J. Pirhonen, and M. Jobling. Effect of low temperature on feed intake, growth rate and body composition of juvenile Baltic salmon. Aquaculture International, 5(6):479–488, November 1997. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1018397014684.

Tzovenis:1997:EDL

[164] I. Tzovenis, N. De Pauw, and P. Sorgeloos. Effect of different light regimes on the docosahexaenoic acid (DHA) content of Isochrysis aff. galbana (clone T-ISO). Aquaculture International, 5(6):489–507, November 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1018349131522.

Navarro:1997:EDD

[165] J. C. Navarro, L. A. McEvoy, M. V. Bell, F. Amat, F. Hontoria, and J. R. Sargent. Effect of different dietary levels of docosahexaenoic acid (DHA, 22:6w-3) on the DHA composition of lipid classes in sea bass larvae eyes. Aquaculture International, 5(6):509–516, November 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1018301215592.

McEvoy:1997:ASP

[166] L. A. McEvoy, J. C. Navarro, F. Amat, and J. R. Sargent. Application of soya phosphatidylcholine in tuna orbital oil enrichment emulsions for artemia. Aquaculture International, 5(6):517–526, November 1997. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1018353232431.

Kolkovski:1997:VCC

[167] S. Kolkovski, A. Arieli, and A. Tandler. Visual and chemical cues stimulate microdiet ingestion in sea bream larvae. Aquaculture International, 5 (6):527–536, November 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1018305416501.

Khoo:1997:GVD

[168] Gideon Khoo, Eric Y. F. Loh, Tit Meng Lim, and Violet P. E. Phang. Genetic variation in different varieties of Siamese fighting fish using isoelectric focusing of sarcoplasmic proteins. Aquaculture International, 5 (6):537–549, November 1997. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1018357400571.

Felix-Pico:1997:RCG

[169] E. F. Felix-Pico, A. Tripp-Quezada, J. L. Castro-Ortiz, G. Serrano-Casillas, P. G. Gonzalez-Ramirez, M. Villalejo-Fuerte, R. Palomares-Garcia, F. A. Garcia-Dominguez, M. Mazon-Suastegui, G. Bojorquez-Verastica, and G. Lopez-Garcia. Repopulation and culture of the Pacific calico scallops in Bahia Concepcion, Baja California Sur, Mexico. Aquaculture International, 5(6):551–563, November 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1018309517410.

Anonymous:1997:AIV

[170] Anonymous. Author index to volume 5 (1997). Aquaculture International, 5(6):565–566, November 1997. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1018361501480.

Johansen:1998:IFR

[171] S. J. S. Johansen and Malcolm Jobling. The influence of feeding regime on growth and slaughter traits of cage-reared Atlantic salmon. Aquaculture International, 6(1):1–17, January 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009250118318.

Sunde:1998:ESG

[172] Leif M. Sunde, Albert K. Imsland, Arild Folkvord, and Sigurd O. Stefansson. Effects of size grading on growth and survival of juvenile turbot at two temperatures. Aquaculture International, 6(1):19–32, January 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009265602388.

Mugnier:1998:ESR

[173] Chantal Mugnier, Alexis Fostier, Sylvie Guezou, Jean-Louis Gaignon, and Loic Quemener. Effect of some repetitive factors on turbot stress response. Aquaculture International, 6(1):33–45, January 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009217719227.

Bassompierre:1998:VPD

[174] M. Bassompierre, T. H. Ostenfeld, E. McLean, and K. Rungruangsak Torrissen. In vitro protein digestion, and growth of Atlantic salmon with different trypsin isozymes. Aquaculture International, 6(1):47–56, January 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009269703297.

daCosta:1998:BMT

[175] Antonio C. A. da Costa and Francisca P. de Franca. The behaviour of the microalgae Tetraselmis chuii in cadmium-contaminated solutions. Aquaculture International, 6(1):57–66, January 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1009221820135.

Lai:1998:PCS

[176] Lawrence Lai and Ken Lam. Pond culture of snakehead in Hong Kong: a case study of an economic solution to common resources. Aquaculture International, 6(1):67–75, January 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009273804206.

Numaguchi:1998:SCP

[177] Katsuyuki Numaguchi. Short communication: Preliminary experiments on the influence of water temperature, salinity and air exposure on the mortality of Manila clam larvae. Aquaculture International, 6(1):77–81, January 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009225921044.

Fagbenro:1998:SCAa

[178] Oyedapa A. Fagbenro. Short communication: Apparent digestibility of various legume seed meals in Nile tilapia diets. Aquaculture International, 6(1):83–87, January 1998. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009277905114.

Anonymous:1998:BRNa

[179] Anonymous. Book reviews and notices. Aquaculture International, 6 (1):89–93, January 1998. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1017179121953.

Koskela:1998:IDF

[180] J. Koskela, M. Jobling, and R. Savolainen. Influence of dietary fat level on feed intake, growth and fat deposition in the whitefish Coregonus lavaretus. Aquaculture International, 6(2):95–102, March 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009230006023.

Ottesen:1998:CET

[181] Oddvar H. Ottesen and Sylvie Bolla. Combined effects of temperature and salinity on development and survival of Atlantic halibut larvae. Aquaculture International, 6(2):103–120, March 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1009234122861.

Hossain:1998:AED

[182] M. Amzad Hossain and Masayuki Furuichi. Availability of environmental and dietary calcium in tiger puffer. Aquaculture International, 6(2):121– 132, March 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009286206931.

Falk:1998:SCS

[183] T. M. Falk, E. K. Abban, W. Villwock, and L. Renwrantz. Speciescharacteristic and subspecies-characteristic haemoglobins in some tilapiine fishes and a comparative study on their globin chains. Aquaculture International, 6(2):133–145, March 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009238323770.

Guillou:1998:RCB

[184] M. Guillou and L. J. L. Lumingas. The reproductive cycle of the ‘blunt’ sea urchin. Aquaculture International, 6(2):147–160, March 1998. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009290307840.

Stepto:1998:ITG

[185] N. K. Stepto and P. A. Cook. Induction of triploidy in the South African abalone using cytochalasin B. Aquaculture International, 6(2):161–169, March 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009242424678.

Mihelakakis:1998:SCE

[186] Apostolos Mihelakakis and Takao Yoshimatsu. Short communication: Effects of salinity and temperature on incubation period, hatching rate and morphogenesis of the red sea bream. Aquaculture International, 6 (2):171–177, March 1998. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009294408749.

Anonymous:1998:Aa

[187] Anonymous. Announcement. Aquaculture International, 6(2):179, March 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1017182025587.

Jobling:1998:RCG

[188] Malcolm Jobling, Helge Tveiten, and Bjarne Hatlen. REVIEW cultivation of Arctic charr: an update. Aquaculture International, 6(3):181– 196, May 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009246509657.

Hussain:1998:PHH

[189] M. G. Hussain, D. J. Penman, and B. J. McAndrew. Production of heterozygous and homozygous clones in Nile tilapia. Aquaculture International, 6(3):197–205, May 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009250626496.

McCarthy:1998:ETG

[190] Ian McCarthy, Erlend Moksness, and Dimitri A. Pavlov. The effects of temperature on growth rate and growth efficiency of juvenile common wolffish. Aquaculture International, 6(3):207–218, May 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009202710566.

Garg:1998:AAE

[191] S. K. Garg, Anita Bhatnagar, and Neeru Narula. Application of azotobacter enhances pond productivity and fish biomass in still water ponds. Aquaculture International, 6(3):219–231, May 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1009254727404.

Basuyaux:1998:UMF

[192] Olivier Basuyaux and Jean-Louis Blin. Use of maize as a food source for sea urchins in a recirculating rearing system. Aquaculture International, 6(3):233–247, May 1998. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009206811475.

Maage:1998:SCA

[193] Amund Maage and Harald Sveier. Short communication: Addition of dietary iron(III) oxide does not increase iron status of growing Atlantic salmon. Aquaculture International, 6(3):249–252, May 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009258828313.

Anonymous:1998:BRNb

[194] Anonymous. Book reviews and notices. Aquaculture International, 6 (3):253–259, May 1998. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1017170932005.

Henry:1998:IPM

[195] J. C. Henry, E. McLean, I. Mayer, and E. M. Donaldson. Induction of precocious maturation in masculinized Atlantic salmon by treatment with sustained-release LHRHa and testosterone. Aquaculture International, 6 (4):261–268, August 1998. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009255815481.

Gouveia:1998:UCV

[196] Luisa Gouveia, Georges Choubert, Emídio Gomes, Paulo Rema, and José Empis. Use of Chlorella vulgaris as a carotenoid source for rainbow trout: effect of dietary lipid content on pigmentation, digestibility and retention in the muscle tissue. Aquaculture International, 6(4):269–279, August 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009251714573.

Saha:1998:CAR

[197] Amlan Krishna Saha and Arun Kumar Ray. Cellulase activity in rohu fingerlings. Aquaculture International, 6(4):281–291, August 1998. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009210929594.

Chakrabarti:1998:IMP

[198] Rina Chakrabarti and Jai Gopal Sharma. Influence of management protocols on carp growth under nursery conditions: relative importance of food and water quality. Aquaculture International, 6(4):293–301, August 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009203831411.

Laing:1998:HCG

[199] Ian Laing and Ren-Mou Chang. Hatchery cultivation of Pacific oyster juveniles using algae produced in outdoor bloom-tanks. Aquaculture International, 6(4):303–315, August 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009247613664.

Fagbenro:1998:SCAb

[200] Oyedapo A. Fagbenro. Short communication apparent digestibility of various oilseed cakes/meals in African catfish diets. Aquaculture International, 6(4):317–322, August 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009299630503.

Mims:1998:IMG

[201] Steven D. Mims and William L. Shelton. Induced meiotic gynogenesis in shovelnose sturgeon. Aquaculture International, 6(5):323–329, September 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009280804945.

Serot:1998:LFA

[202] T. Sérot, G. Gandemer, and M. Demaimay. Lipid and fatty acid compositions of muscle from farmed and wild adult turbot. Aquaculture International, 6(5):331–343, September 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009284905854.

Garcia-Gallego:1998:EBC

[203] M. García-Gallego and H. Akharbach. Evolution of body composition of European eels during their growth phase in a fish farm, with special emphasis on the lipid component. Aquaculture International, 6(5):345–356, September 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009207904037.

Terova:1998:ADE

[204] G. Terova, M. Saroglia, Z. Gy. Papp, and S. Cecchini. Ascorbate dynamics in embryos and larvae of sea bass and sea bream, originating from broodstocks fed supplements of ascorbic acid. Aquaculture International, 6(5):357–367, September 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009237022692.

Gbulubo:1998:SIE

[205] A. J. Gbulubo and E. S. Erondu. Salinity influence on the early stages of the African catfish. Aquaculture International, 6(5):369–379, September 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009228820875.

Hossain:1998:ERS

[206] Mostafa A. R. Hossain, Graham S. Haylor, and Malcolm C. M. Beveridge. An evaluation of radiography in studies of gastric evacuation in African catfish fingerlings. Aquaculture International, 6(5):379–385, September 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009289006763.

Domingues:1998:PSP

[207] Pedro M. Domingues, Philip E. Turk, Jose P. Andrade, and Phillip G. Lee. Pilot-scale production of mysid shrimp in a static water system. Aquaculture International, 6(5):387–402, September 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009232921784.

Williot:1998:IYB

[208] P. Williot. Influence of yolk blackish pigmentation of Siberian sturgeon on reproductive performance and larval survival. Aquaculture International, 6(6):403–410, December 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009241116526.

Pane:1998:VMM

[209] L. Pane, M. Feletti, C. Bertino, and A. Carli. Viability of the marine microalga Tetraselmis suecica grown free and immobilized in alginate beads. Aquaculture International, 6(6):411–420, December 1998. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009279300596.

Ribeiro:1998:PDL

[210] F. A. L. T. Ribeiro and D. A. Jones. The potential of dried, lowhatch, decapsulated artemia cysts for feeding prawn post-larvae. Aquaculture International, 6(6):421–440, December 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1009231417434.

Ruohonen:1998:CNL

[211] K. Ruohonen, J. Vielma, and D. J. Grove. Comparison of nutrient losses into the water from rainbow trout culture based on fresh Baltic herring, moist and dry diets. Aquaculture International, 6(6):441–450, December 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009283401505.

Batzios:1998:SCE

[212] C. Batzios, G. Fotis, and I. Gavriilidou. Short communication economic dimension of gas bubble disease effects on rainbow trout culture. Aquaculture International, 6(6):451–455, December 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1009235518343.

Anonymous:1998:BR

[213] Anonymous. Book reviews. Aquaculture International, 6(6):457–462, December 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1017114902413.

Anonymous:1998:Ab

[214] Anonymous. Announcement. Aquaculture International, 6(6):463–464, December 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1017139619252.

Anonymous:1998:AIV

[215] Anonymous. Author index to volume 6 (1998). Aquaculture International, 6(6):465–466, December 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1017191603322.

Anonymous:1998:KIV

[216] Anonymous. Keyword index to volume 6 (1998). Aquaculture International, 6(6):467–471, December 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1017143720160.

Anonymous:1998:LRV

[217] Anonymous. List of referees for volume 6 of aquaculture international. Aquaculture International, 6(6):473, December 1998. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1017195704230.

Fidhiany:1999:LTO

[218] Lucia Fidhiany and Klaus Winckler. Long term observation on several growth parameters of convict child under an enhanced ultraviolet-a (320– 400 nm) irradiation. Aquaculture International, 7(1):1–12, January 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009263901180.

Winckler:1999:TTM

[219] Klaus Winckler and Lucia Fidhiany. Temperature tolerance and metabolic depression of a convict child under the influence of enhanced ultraviolet-a (320–400 nm) irradiation. Aquaculture International, 7 (1):13–27, January 1999. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009236618927.

Pavlidis:1999:DRS

[220] M. Pavlidis, M. Paspatis, M. Koistinen, T. Paavola, P. Divanach, and M. Kentouri. Diel rhythms of serum metabolites and thyroid hormones in red porgy held in different photoperiod regimes. Aquaculture International, 7(1):29–44, January 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1023/A:1009292702997.

Mookerji:1999:RYU

[221] Nandita Mookerji and T. Ramakrishan Rao. Rates of yolk utilization and effects of delayed initial feeding in the larvae of the freshwater fishes rohu and singhi. Aquaculture International, 7(1):45–56, January 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009244819835.

Fontainhas-Fernandes:1999:RFM

[222] A. Fontaínhas-Fernandes, E. Gomes, M. A. Reis-Henriques, and J. Coimbra. Replacement of fish meal by plant proteins in the diet of Nile tilapia: Digestibility and growth performance. Aquaculture International, 7(1):57–67, January 1999. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009296818443.

Anonymous:1999:BRNa

[223] Anonymous. Book reviews and notices. Aquaculture International, 7(1): 69, January 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1017288117284.

Anonymous:1999:A

[224] Anonymous. Announcement. Aquaculture International, 7(1):71–72, January 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1017260728612.

Anonymous:1999:IA

[225] Anonymous. Instructions for authors. Aquaculture International, 7 (1):73–75, January 1999. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1017208701354.

Ojanguren:1999:ETG

[226] Alfredo F. Ojanguren, Felipe G. Reyes-Gavilán, and Rolando Rodríguez Muñoz. Effects of temperature on growth and efficiency of yolk utilisation in eggs and pre-feeding larval stages of Atlantic salmon. Aquaculture International, 7(2):81–87, March 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009214804949.

Chen:1999:PCC

[227] Yrong-Song Chen, Malcolm C. M. Beveridge, and Trevor C. Telfer. Physical characteristics of commercial pelleted Atlantic salmon feeds and consideration of implications for modeling of waste dispersion through sedimentation. Aquaculture International, 7(2):89–100, March 1999. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009249721787.

Erikson:1999:CBT

[228] U. Erikson, T. Sigholt, T. Rustad, I. E. Einarsdottir, and L. Jørgensen. Contribution of bleeding to total handling stress during slaughter of Atlantic salmon. Aquaculture International, 7(2):101–115, March 1999. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009236628690.

Ronsholdt:1999:QCF

[229] B. Rønsholdt and E. McLean. Quality characteristics of fresh rainbow trout as perceived by the Danish processing industry. Aquaculture International, 7(2):117–127, March 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009201805858.

Stech:1999:MIS

[230] Ludek Štěch, Otomar Linhart, William L. Shelton, and Steven D. Mims. Minimally invasive surgical removal of ovulated eggs from paddlefish. Aquaculture International, 7(2):129–133, March 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1009253806766.

Anonymous:1999:ACP

[231] Anonymous. 2nd announcement and call for papers. Aquaculture International, 7(2):135, March 1999. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1017227125422.

Alabi:1999:IAE

[232] A. O. Alabi, Z. Che Cob, D. A. Jones, and J. W. Latchford. Influence of algal exudates and bacteria on growth and survival of white shrimp larvae fed entirely on microencapsulated diets. Aquaculture International, 7(3):137–158, May 1999. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009257902630.

Nagesh:1999:SHA

[233] T. S. Nagesh, N. Jayabalan, C. V. Mohan, T. S. Annappaswamy, and T. M. Anil. Survival and histological alterations in juvenile tiger shrimps exposed to saponin. Aquaculture International, 7(3):159–167, May 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009239319468.

Maguire:1999:IBI

[234] Julie A. Maguire, David A. O’Connor, and Gavin M. Burnell. An investigation into behavioural indicators of stress in juvenile scallops. Aquaculture International, 7(3):169–177, May 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1009291323103.

Doroudi:1999:GSB

[235] Mehdi S. Doroudi, Paul C. Southgate, and Robert J. Mayer. Growth and survival of blackslip pearl oyster larvae fed different densities of microalgae. Aquaculture International, 7(3):179–187, May 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009207810808.

Jeffs:1999:PDC

[236] A. G. Jeffs. The potential for developing controlled breeding in the Chilean oyster. Aquaculture International, 7(3):189–199, May 1999. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009276329463.

Piaia:1999:GSF

[237] Rosamari Piaia, Clarice R. Townsend, and Bernardo Baldisserotto. Growth and survival of fingerlings of silver catfish exposed to different photoperiods. Aquaculture International, 7(3):201–205, May 1999. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009299830102.

Bell:1999:NSC

[238] J. D. Bell and M. Gervis. New species for coastal aquaculture in the tropical Pacific — constraints, prospects and considerations. Aquaculture International, 7(4):207–223, July 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009245612050.

Lavens:1999:EDE

[239] P. Lavens, E. Lebegue, H. Jaunet, A. Brunel, Ph. Dhert, and P. Sorgeloos. Effect of dietary essential fatty acids and vitamins on egg quality in turbot broodstocks. Aquaculture International, 7(4):225– 240, July 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009225028889.

Garcia:1999:SGB

[240] L. M. B. Garcia, C. M. H. Garcia, A. F. S. Pineda, E. A. Gammad, J. Canta, S. P. D. Simon, G. V. Hilomen-Garcia, A. C. Gonzal, and C. B. Santiago. Survival and growth of bighead carp fry exposed to low salinities. Aquaculture International, 7(4):241–250, July 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009277012959.

Umesh:1999:EGC

[241] N. R. Umesh, K. M. Shankar, and C. V. Mohan. Enhancing growth of common carp, rohu and Mozambique tilapia through plant substrate: The role of bacterial biofilm. Aquaculture International, 7(4):251–260, July 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009229113868.

deOliveira:1999:GCC

[242] M. A. C. L. de Oliveira, M. P. C. Monteiro, P. G. Robbs, and S. G. F. Leite. Growth and chemical composition of Spirulina maxima and Spirulina platensis biomass at different temperatures. Aquaculture International, 7(4):261–275, July 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1023/A:1009233230706.

Fagbenro:1999:ADV

[243] O. Fagbenro. Apparent digestibility of various cereal grain by-products in common carp diets. Aquaculture International, 7(4):277–281, July 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009285214776.

Anonymous:1999:BRNb

[244] Anonymous. Book review and notices. Aquaculture International, 7 (4):283–287, July 1999. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1017211008995.

Hossain:1999:CRT

[245] M. A. Hossain and M. Furuichi. Calcium requirement of tiger puffer fed a semi–purified diet. Aquaculture International, 7(5):287–293, September 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009237429590.

Paspatis:1999:PRR

[246] M. Paspatis, G. Markakis, G. Koumoundouros, and M. Kentouri. Preliminary results on rearing of Sparus aurata × Pagrus pagrus hybrids. Performance comparison with the parental species. Aquaculture International, 7(5):295–306, September 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009296000457.

Grau:1999:SPM

[247] A. Grau, F. Riera, and E. Carbonell. Some protozoan and metazoan parasites of the amberjack from the Balearic Sea (Western Mediterranean). Aquaculture International, 7(5):307–317, September 1999. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009212520021.

AlHafedh:1999:EDP

[248] Y. S. Al Hafedh, A. Q. Siddiqui, and M. Y. Al-Saiady. Effects of dietary protein levels on gonad maturation, size and age at first maturity, fecundity and growth of Nile tilapia. Aquaculture International, 7(5):319–332, September 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009276911360.

Sulaiman:1999:GEB

[249] Z. H. Sulaiman, R. H. Ming Chan, and P. M. Simanjuntak. Gene expression in black tiger prawns following intramuscular injection of β-gal plasmid. Aquaculture International, 7(5):333–340, September 1999. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009261913545.

Piveteau:1999:CLF

[250] F. Piveteau, G. Gandemer, J-P. Baud, and M. Demaimay. Changes in lipid and fatty acid compositions of European oysters fattened with skeletonema costatum diatom for six weeks in ponds. Aquaculture International, 7(5):341–355, September 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009244530383.

Gardner:1999:FIG

[251] D. Gardner and C. G. Carter. Feed intake and growth of juvenile silver trevally at the southern limit of their distribution. Aquaculture International, 7(5):357–360, September 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009296531292.

Rigos:1999:MIG

[252] G. Rigos, P. Christophilogiannis, M. Yiagnisi, A. Andriopoulou, M. Koutsodimou, I. Nengas, and M. Alexis. Myxosporean infections in Greek mariculture. Aquaculture International, 7(5):361–364, September 1999. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009200732200.

Anonymous:1999:BRNc

[253] Anonymous. Book reviews and notices. Aquaculture International, 7(5): 365–367, September 1999. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1017204222402.

Downing:2000:EPT

[254] G. Downing and M. K. Litvak. The effect of photoperiod, tank colour and light intensity on growth of larval haddock. Aquaculture International, 7 (6):369–382, November 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009204909992.

Alok:2000:VAS

[255] D. Alok, G. P. Talwar, and L. C. Garg. In vivo activity of salmon gonadotropin releasing hormone (GnRH), its agonists with structural modifications at positions 6 and 9, mammalian GnRH agonists and native cGnRH-II on the spawning of an Indian catfish. Aquaculture International, 7(6):383–392, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009262230465.

Huang:2000:CAR

[256] X. Huang, R. Zhou, L. Liu, H. Cheng, Y. Guo, and Q. Yu. Chromosome assignments of the rice field eel Sox9 and Sox17 genes. Aquaculture International, 7(6):393–397, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1009230814535.

Feldlite:2000:EDS

[257] M. Feldlite and A. Milstein. Effect of density on survival and growth of cyprinid fish fry. Aquaculture International, 7(6):399–411, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009282805391.

Ganguly:2000:BCB

[258] S. Ganguly, J. Chatterjee, and B. B. Jana. Biogeochemical cycling bacterial activity in response to lime and fertilizer applications in pond systems. Aquaculture International, 7(6):413–431, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009289823138.

Jorquera:2000:PBS

[259] M. A. Jorquera, C. E. Riquelme, L. A. Loyola, and L. F. Muñoz. Production of bactericidal substances by a marine vibrio isolated from cultures of the scallop Argopecten purpuratus. Aquaculture International, 7(6): 433–448, November 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009246007208.

Duinker:2000:EPC

[260] A. Duinker, C. Saout, and Y. M. Paulet. Effect of photoperiod on conditioning of the great scallop. Aquaculture International, 7(6):449–457, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009298024046.

Anonymous:2000:BRN

[261] Anonymous. Book reviews and notices. Aquaculture International, 7(6): 459–464, November 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1017245822085.

Anonymous:2000:AIVa

[262] Anonymous. Author index to volume 7 (1999). Aquaculture International, 7(6):465–466, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1017238325576.

Anonymous:2000:KIVa

[263] Anonymous. Keyword index to volume 7 (1999). Aquaculture International, 7(6):467–471, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1017230726485.

Anonymous:2000:LRV

[264] Anonymous. List of referees for volume 7 of aquaculture international. Aquaculture International, 7(6):473, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1017291008117.

Anonymous:2000:CVa

[265] Anonymous. Contents of volume 7. Aquaculture International, 7(6): 475–476, November 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1017278609647.

Ciereszko:2000:EAA

[266] A. Ciereszko and K. Dabrowski. Effect of ascorbic acid supplement in vitro on rainbow trout sperm viability. Aquaculture International, 8(1):1– 8, January 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009253330451.

Mohamed:2000:QDB

[267] J. Shaik Mohamed, B. Ravisankar, and A. Ibrahim. Quantifying the dietary biotin requirement of the catfish,Clarias batrachus. Aquaculture International, 8(1):9–18, January 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009264714521.

Ng:2000:EFR

[268] Wing-Keong Ng, Kim-Sun Lu, Roshada Hashim, and Ahyaudin Ali. Effects of feeding rate on growth, feed utilization and body composition of a tropical bagrid catfish. Aquaculture International, 8(1):19–29, January 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009216831360.

Milstein:2000:TWS

[269] A. Milstein, Y. Eran, E. Nitzan, M. Zoran, and D. Joseph. Tilapia wild spawning control through predator fishes: Israelitrial with red-drum and hybrid bass. Aquaculture International, 8(1):31–40, January 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009268815430.

Cavalli:2000:MPO

[270] Ronaldo O. Cavalli, Gerben Menschaert, Patrick Lavens, and Patrick Sorgeloos. Maturation performance, offspring quality and lipidcomposition of Macrobrachium rosenbergii females fed increasing levels of dietary phospholipids. Aquaculture International, 8(1):41–58, January 2000. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009220932268.

Guan:2000:API

[271] Rui-Zhang Guan. Abundance and production of the introduced signal crayfish in a British lowland river. Aquaculture International, 8(1):59– 76, January 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009272916339.

vanderMeeren:2000:CCM

[272] Gro I. van der Meeren and Lars E. Uksnøy. A comparison of claw morphology and dominance between wild and cultivated male European lobster. Aquaculture International, 8(1):77–94, January 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009225001318.

Harlioglu:2000:RBE

[273] M. Mustafa Harlioğlu and Ibrahim Türkgülü. The relationship between egg size and female size infreshwater crayfish, Astacus leptodactylus. Aquaculture International, 8(1):95–98, January 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1009281118156.

McLean:2000:BRS

[274] Ewen McLean. Book review: Sustainable Aquaculture: Food for the Future? N. Svennevig, H. Reinertsen and M. New (Editors). Aquaculture International, 8(1):99–102, January 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009224707890.

Strand:2000:E

[275] Øivind Strand and Gavin Burnell. Editorial. Aquaculture International, 8 (2–3):111–112, March 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009233211357.

Bourne:2000:PSC

[276] Neil F. Bourne. The potential for scallop culture — the next millenium. Aquaculture International, 8(2–3):113–122, March 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009212226803.

Mortensen:2000:SIT

[277] S. Mortensen. Scallop introductions and transfers, from an animal health point of view. Aquaculture International, 8(2–3):123–138, March 2000.

CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009282112691.

Goldberg:2000:SEN

[278] R. Goldberg, J. Pereira, and P. Clark. Strategies for enhancement of natural bay scallop, Argopecten irradians irradians, populations; a case study in the Niantic River estuary, Connecticut, USA. Aquaculture International, 8(2–3):139–158, March 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009242429529.

Christophersen:2000:EAE

[279] Gyda Christophersen. Effects of air emersion on survival and growth of hatchery rearedgreat scallop spat. Aquaculture International, 8(2–3):159– 168, March 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009294413599.

Minchin:2000:EAE

[280] Dan Minchin, Geir Haugum, Hanne Skjæggestad, and Øivind Strand. Effect of air exposure on scallop behaviour, and the implications for subsequent survival in culture. Aquaculture International, 8(2–3):169– 182, March 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009246530438.

Grecian:2000:ISI

[281] Lorelei A. Grecian, G. Jay Parsons, Patrick Dabinett, and Cyr Couturier. Influence of season, initial size, depth, gear type and stocking density on the growth rates and recovery of sea scallop, placopecten magellancius, on a farm-based nursery. Aquaculture International, 8(2– 3):183–206, March 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009298631346.

Cano:2000:GMK

[282] Juana Cano, Ma Jesús Campos, and Guillermo Román. Growth and mortality of the king scallop grown in suspended culture in Malaga, Southern Spain. Aquaculture International, 8(2–3):207–225, March 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009202801304.

Penchaszadeh:2000:RCS

[283] Pablo E. Penchaszadeh, Claudio Paredes, and Juan J. Salaya. Reproductive cycle of the south American scallop Amusium laurenti (Gmelin, 1791) (Bivalvia, Pectinidae). Aquaculture International, 8(2–3):227–235, March 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009239831230.

Stotz:2000:WAR

[284] Wolfgang Stotz. When aquaculture restores and replaces an overfished stock:is the conservation of the species assured? The case of the scallop Argopecten purpuratus in Northern Chile. Aquaculture International, 8 (2–3):237–247, March 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009215119051.

Andersen:2000:FTS

[285] Sissel Andersen, Gavin Burnell, and Øivind Bergh. Flow-through systems for culturing great scallop larvae. Aquaculture International, 8(2–3):249– 257, March 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009271220868.

Andersen:2000:SDE

[286] Sissel Andersen and Halldis Ringvold. Seasonal differences in effect of broodstock diet on spawning success in the great scallop. Aquaculture International, 8(2–3):259–265, March 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009279422686.

Mortensen:2000:MSS

[287] S. Mortensen, T. Meeren, A. Fosshagen, I. Hernar, L. Harkestad, L. Torkildsen, and Ø. Bergh. Mortality of scallop spat in cultivation, infested with tube-dwelling bristle worms, Polydora sp. Aquaculture International, 8(2–3):267–271, March 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009287706756.

Anonymous:2000:MR

[288] Anonymous. Media release. Aquaculture International, 8(2–3):273, March 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1017373327579.

Salvesen:2000:MEM

[289] Ingrid Salvesen, Kjell Inge Reitan, Jorunn Skjermo, and Gunvor Øie. Microbial environments in marine larviculture:impacts of algal growth rates on the bacterial load in six microalgae. Aquaculture International, 8(4):275–287, July 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009200926452.

Cnaani:2000:CTT

[290] A. Cnaani, G. A. E. Gall, and G. Hulata. Cold tolerance of tilapia species and hybrids. Aquaculture International, 8(4):289–298, July 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009299109614.

Fontainhas-Fernandes:2000:PTR

[291] A. Fontaínhas-Fernandes, M. Monteiro, A. Figueiredo, E. Gomes, J. Coimbra, and M. A. Reis-Henriques. Partial or total replacement of fish meal byplant protein affects gonadal development and plasma 17β-estradiol levels in female Nile tilapia. Aquaculture International, 8(4):299–313, July 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009217310464.

Soliman:2000:EID

[292] Adel K. Soliman, Abdou A. A. El-Horbeety, Mohamed A. R. Essa, Mohamed A. Kosba, and Ibrahim A. Kariony. Effects of introducing ducks into fish ponds on water quality, natural productivity and fish production together with the economic evaluation of the integrated and nonintegrated systems. Aquaculture International, 8(4):315–326, July 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009252910522.

Sun:2000:DCS

[293] Piera S. Sun, Tina M. Weatherby, Marilyn F. Dunlap, Kristi L. Arakaki, Donis T. Zacarias, and Spencer R. Malecha. Developmental changes in structure and polypeptide profile of the androgenic gland of the freshwater prawn Macrobrachium rosenbergii. Aquaculture International, 8 (4):327–334, July 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009205027360.

Poli:2000:BAB

[294] B. M. Poli, G. Zampacavallo, G. Parisi, A. Poli, and M. Mascini. Biosensors applied to biochemical fish qualityindicators in refrigerated and frozen sea bass reared in aerated or hyperoxic conditions. Aquaculture International, 8(4):335–348, July 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009209113248.

Pironet:2000:TED

[295] F. N. Pironet and J. B. Jones. Treatments for ectoparasites and diseases in captive Western Australian dhufish. Aquaculture International, 8(4):349–361, July 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009257011431.

Jobling:2000:BRE

[296] Malcolm Jobling. Book review: Encyclopedia of Aquaculture, R. R. Stickney (Editor). Aquaculture International, 8(4):363–364, July 2000. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009253212286.

Jobling:2000:BRI

[297] Malcolm Jobling. Book review: An Illustrated Dictionary of Fish and Shellfish (CD-ROM), C. Frimodt. Aquaculture International, 8(4):365– 366, July 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009219630942.

Makridis:2000:CGF

[298] Pavlos Makridis, Anders Jon Fjellheim, Jorunn Skjermo, and Olav Vadstein. Colonization of the gut in first feeding turbot by bacterial strains added to the water or bioencapsulated in rotifers. Aquaculture International, 8(5):367–380, September 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009251531832.

Brannas:2000:GEG

[299] Eva Brännäs and Johan Linnér. Growth effects in Arctic charr reared in cold water: Feed frequency, access to bottom feeding and stocking density. Aquaculture International, 8(5):381–389, September 2000. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009235130015.

Barki:2000:ELF

[300] Assaf Barki, Sheenan Harpaz, Gideon Hulata, and Ilan Karplus. Effects of larger fish and size grading on growth and size variation in fingerling silver perch. Aquaculture International, 8(5):391–401, September 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009274726380.

El-Sayed:2000:AEP

[301] A.-F. M. El-Sayed, I. Nmartínez, and F. J. Moyano. Assessment of the effect of plant inhibitors on digestive proteases of Nile tilapia using in vitro assays. Aquaculture International, 8(5):403–415, September 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009262407725.

Favaloro:2000:MCA

[302] Eugenia Favaloro and Antonio Mazzola. Meristic character analysis and skeletal anomalies during growth in reared sharpsnout seabream. Aquaculture International, 8(5):417–430, September 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1009284421354.

Lopez:2000:EEP

[303] D. A. López, V. A. Riquelme, and M. L. González. The effects of epibionts and predators on the growth and mortality rates of Argopecten purpuratus cultures in southern Chile. Aquaculture International, 8(5):431–442, September 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009240202273.

Harlioglu:2000:IRM

[304] Muzaffer Mustafa Harlıoğlu. Incisor ridge modification of the mandibles in freshwater crayfish, Astacus leptodactylus. Aquaculture International, 8(5):443–453, September 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1023/A:1009259702699.

Katavic:2000:GPP

[305] I. Katavic, L. Grubisic, and N. Skakelja. Growth performance of pink dentex as compared to four other sparids reared in marine cages in Croatia. Aquaculture International, 8(5):455–461, September 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009276204516.

Rao:2000:IBI

[306] P. S. Srnivasa Rao, Iddya Karunasagar, S. K. Otta, and Indrani Karunasagar. Incidence of bacteria involved in nitrogen and sulphur cycles in tropical shrimp culture ponds. Aquaculture International, 8(5): 463–472, September 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1009250004999.

Dutil:2000:BRS

[307] Jean-Denis Dutil. Book review: Stock Enhancement and Sea Ranching. B. Howell, E. Moksness and T. Svasand (Editors). Aquaculture International, 8(5):473–474, September 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1009279816830.

Bradshaw:2000:PA

[308] Alison J. Bradshaw. Publisher’s announcement. Aquaculture International, 8(6):475, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1023/A:1009263932544.

Oliva-Teles:2000:RAG

[309] Aires Oliva-Teles. Recent advances in European sea bass and gilthead sea bream nutrition. Aquaculture International, 8(6):477–492, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009236517555.

Beaumont:2000:GCT

[310] Andy Beaumont. Genetic considerations in transfers and introductions of scallops. Aquaculture International, 8(6):493–512, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1009210404351.

Robert:2000:ESF

[311] René Robert and Laurence Nicholas. The effect of seawater flow and temperature on metamorphosis and postlarval development in great scallop. Aquaculture International, 8(6):513–530, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009274707986.

Rasmussen:2000:GFU

[312] R. S. Rasmussen, T. H. Ostenfeld, and E. McLean. Growth and feed utilisation of rainbow trout subjected to changes in feed lipid concentrations. Aquaculture International, 8(6):531–542, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1009291227550.

Adams:2000:SHC

[313] Colin Adams, Felicity Huntingford, Jimmy Turnbull, Steve Arnott, and Aly Bell. Size heterogeneity can reduce aggression and promote growth in Atlantic salmon parr. Aquaculture International, 8(6):543–549, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1009255612529.

Glamuzina:2000:NFA

[314] B. Glamuzina, N. Glavić, P. Tutman, V. Kožul, and B. Skaramuca. Notes on first attempt at artificial spawning and rearing of early stages with goldblotch grouper, Epinephelus costae (Steindachner, 1875). Aquaculture International, 8(6):551–555, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1009211929367.

Anonymous:2000:AIVb

[315] Anonymous. Author index to volume 8 (2000). Aquaculture International, 8(6):557–558, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1017256526010.

Anonymous:2000:KIVb

[316] Anonymous. Keyword index to volume 8 (2000). Aquaculture International, 8(6):559–560, November 2000. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1017200408263.

Anonymous:2000:IA

[317] Anonymous. Instructions for authors. Aquaculture International, 8(6): 561–567, November 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1017324203411.

Anonymous:2000:CVb

[318] Anonymous. Contents of volume 8. Aquaculture International, 8(6): 569–573, November 2000. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1017292123284.

Burnell:2001:E

[319] Gavin Burnell. Editorial. Aquaculture International, 9(1):1–2, January 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1012546928955.

Milstein:2001:DDF

[320] Ana Milstein, Myron Zoran, and Hans Juergen Krambeck. Destratification in deep fish culture reservoirs in Israel: Practice and modelling. Aquaculture International, 9(1):3–15, January 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1012538711129.

Jana:2001:RPS

[321] B. B. Jana, J. Chatterjee, S. Ganguly, and T. Jana. Responses of phosphate solubilizing bacteria to qualitatively different fertilization in simulated and natural fish ponds. Aquaculture International, 9(1):17–34, January 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1012594828876.

Linner:2001:GGA

[322] Johan Linnér and Eva Brännäs. Growth in Arctic charr and rainbow trout fed temporally concentrated or spaced daily meals. Aquaculture International, 9(1):35–44, January 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1012534610220.

Freites:2001:BCT

[323] L. Freites, J. H. Himmelman, J. M. Babarro, C. J. Lodeiros, and A. Vélez. Bottom culture of the tropical scallop Lyropecten (Nodipecten) nodosus (L.) in the Golfo de Cariaco, Venezuela. Aquaculture International, 9 (1):45–60, January 2001. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1012526325242.

Nam:2001:SGI

[324] Yoon Kwon Nam, Geyong Cheol Choi, Dong Joon Park, and Dong Soo Kim. Survival and growth of induced tetraploid mud loach. Aquaculture International, 9(1):61–71, January 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1012540024333.

Lopes:2001:SGS

[325] J. M. Lopes, L. V. F. Silva, and B. Baldisserotto. Survival and growth of silver catfish larvae exposed to different water pH. Aquaculture International, 9(1):73–80, January 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1012512211898.

Milstein:2001:EWW

[326] A. Milstein and M. Zoran. Effect of water withdrawal from the epilimnion on thermal stratification in deep dual purpose reservoirs for fish culture and field irrigation. Aquaculture International, 9(1):81–86, January 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1012530409312.

Welker:2001:GAC

[327] C. Welker, P. De Negro, and M. Sarti. Green algal carotenoids and yellow pigmentation of rainbow trout fish. Aquaculture International, 9(1):87– 93, January 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1012582426150.

Moutou:2001:EDF

[328] K. A. Moutou, I. D. McCarty, and D. F. Houlihan. The effect of dietary flumequine on food consumption and growth in rainbow trout. Aquaculture International, 9(1):95–102, January 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1012586627059.

Carter:2001:BRE

[329] C. Carter. Book review: Environmental Impacts of Aquaculture, K. D. Black (ed.). Aquaculture International, 9(1):103–104, January 2001. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1012542812038.

Jobling:2001:BRA

[330] Malcolm Jobling. Book review: Applied Fish Genetics, B. K. Padhi and R. K. Mandal. Aquaculture International, 9(1):105–106, January 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1012590727967.

Ruff:2001:AED

[331] N. Ruff, P. Lavens, J.-Z. Huo, P. Sorgeloos, H. J. Nelis, and A. De Leenheer. Antioxidant effect of dietary tocopherol and ascorbic acid on growth and survival of Litopenaeus vannamei postlarvae. Aquaculture International, 9(2):115–126, March 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1014292223946.

Narvarte:2001:DTT

[332] Maite A. Narvarte and Marcela S. Pascual. Diet trials on tehuelche scallop Aequipecten tehuelchus (d’Orb) larvae. Aquaculture International, 9(2):127–131, March 2001. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1014257008017.

Milstein:2001:EDM

[333] A. Milstein, M. Zoran, M. Kochba, and Y. Avnimelech. Effect of different management practices on water quality of intensive tilapia culture systems in Israel. Aquaculture International, 9(2):133–152, March 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1014209124855.

Maragoudaki:2001:GFR

[334] D. Maragoudaki, M. Paspatis, and M. Kentouri. Growth and feeding responses of juvenile red porgy to restrictive self-feeding regimes. Aquaculture International, 9(2):153–170, March 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1014261104112.

Kim:2001:PMY

[335] Joong Kyun Kim and Hae-Yoon Chung. Preservation of manipulated yeast diet. Aquaculture International, 9(2):171–181, March 2001. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1014242526754.

Dorucu:2001:ECS

[336] Mustafa Dörücü and Asiye Girgin. The effect of cypermethrin on some haematological parameters of Cyprinus carpio. Aquaculture International, 9(2):183–187, March 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1023/A:1014297528028.

Panini:2001:IEL

[337] Elsa B. Panini, Constantinos C. Mylonas, Sylvia Zanuy, Manuel Carrillo, Jesus Ramos, and Michael P. Bruce. Incubation of embryos and larvae of marine fish using microtiter plates. Aquaculture International, 9(2):189– 196, March 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1014261830098.

Mihelakakis:2001:EFF

[338] A. Mihelakakis, T. Yoshimatsu, and C. Tsolkas. Effect of feeding frequency on growth, feed efficiency, and body composition in young common pandora. Aquaculture International, 9(2):197–204, March 2001. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1014203618871.

Papandroulakis:2001:PGW

[339] N. Papandroulakis, P. Divanach, P. Anastasiadis, and M. Kentouri. The pseudo-green water technique for intensive rearing of sea bream (Sparus aurata) larvae. Aquaculture International, 9(3):205–216, May 2001. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1016813623122.

Conides:2001:SER

[340] Alexis J. Conides and Branko Glamuzina. Study on the effects of rearing density, temperature and salinity on hatching performance of the European sea bass, Dicentrarchus labrax (Linnaeus, 1758). Aquaculture International, 9(3):217–224, May 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1015330608607.

Makridis:2001:ABB

[341] P. Makridis, Ø. Bergh, J. Skjermo, and O. Vadstein. Addition of bacteria bioencapsulated in Artemia metanauplii to a rearing system for halibut larvae. Aquaculture International, 9(3):225–235, May 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1016815929846.

Mihelakakis:2001:EFR

[342] Apostolos Mihelakakis, Takao Yoshimatsu, and Christos Tsolkas. Effects of feeding rate on growth, feed utilization and body composition of red porgy fingerlings: preliminary results. Aquaculture International, 9(3):237–245, May 2001. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1015345224537.

Rigos:2001:EAT

[343] G. Rigos and R. Stevenson. The effect of antibiotic treatment on the establishment of persistent infection with Yersinia ruckeri Serovar II in rainbow trout Oncorhynchus mykiss (Walbaum). Aquaculture International, 9(3):247–253, May 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1023/A:1015382625446.

Alarcon:2001:USP

[344] F. J. Alarcón, F. J. Moyano, and M. Díaz. Use of SDS-page in the assessment of protein hydrolysis by fish digestive enzymes. Aquaculture International, 9(3):255–267, May 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1016809014922.

Celada:2001:SST

[345] J. D. Celada, J. M. Carral, J. R. Pérez, M. Sáez-Royuela, and C. Muñoz. Successful storage and transport of eggs of the white-clawed crayfish (Austropotamobius pallipes Lereboullet). Aquaculture International, 9 (3):269–271, May 2001. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1015337211733.

Ng:2001:EMS

[346] Wing-Keong Ng, Lay-Pheng Ang, and Fuen-Lan Liew. An evaluation of mineral supplementation of fish meal-based diets for African catfish. Aquaculture International, 9(3):273–283, May 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1015359715815.

Jobling:2001:RAM

[347] Malcolm Jobling. Recent advances in marine biotechnology, volume 4: Aquaculture. Aquaculture International, 9(3):285–286, May 2001. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1015373601424.

Jorquera:2001:BCS

[348] M. A. Jorquera, F. R. Silva, and C. E. Riquelme. Bacteria in the culture of the scallop Argopecten purpuratus (Lamarck, 1819). Aquaculture International, 9(4):285–303, July 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1020449324456.

Bergh:2001:GSP

[349] Øivind Bergh and Øivind Strand. Great scallop, Pecten maximus, research and culture strategies in Norway: a review. Aquaculture International, 9(4):305–317, July 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1023/A:1020452715518.

Domingues:2001:UAS

[350] Pedro M. Domingues, António Sykes, and José P. Andrade. The use of artemia sp. or mysids as food source for hatchlings of the cuttlefish (Sepia officinalis L.); effects on growth and survival throughout the life cycle. Aquaculture International, 9(4):319–331, July 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1020416811568.

Hansen:2001:EED

[351] T. K. Hansen and I. B. Falk-Petersen. Effects of egg disinfection and incubation temperature on early life stages of spotted wolffish. Aquaculture International, 9(4):333–344, July 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1020437023395.

Mohanty:2001:FMW

[352] Rajeeb K. Mohanty. Feeding management and waste production in semi-intensive farming of Penaeus monodon (Fab.) at different stocking densities. Aquaculture International, 9(4):345–355, July 2001. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1020421309414.

Jobling:2001:BRN

[353] Malcolm Jobling. Book review: Nutrient Requirements and Feeding of Finfish for Aquaculture. C. D. Webster and C. E. Lim (editors). Aquaculture International, 9(4):367–368, July 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1020466312058.

Jobling:2001:BRR

[354] Malcolm Jobling. Book review: Rural Aquaculture. P. Edwards, D. C. Little and H. Demaine (eds.). Aquaculture International, 9(4):369– 370, July 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1020416013913.

Peterson:2001:SAP

[355] R. H. Peterson and P. Harmon. Swimming ability of pre-feeding striped bass larvae. Aquaculture International, 9(5):361–366, September 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1020533316106.

Morel:2001:SEJ

[356] G. M. Morel and S. F. Bossy. A seeding experiment of juvenile great scallops (Pecten maximus (L.)) off the Island of Jersey. Aquaculture International, 9(5):367–377, September 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1020561909788.

Minton:2001:FMG

[357] J. W. Minton, L. S. Walsh, P. G. Lee, and J. W. Forsythe. First multigeneration culture of the tropical cuttlefish Sepia pharaonis Ehrenberg, 1831. Aquaculture International, 9(5):379–392, September 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1020535609516.

Domingues:2001:EDF

[358] Pedro M. Domingues, Philip E. Turk, Jose P. Andrade, and Phillip G. Lee. Effects of different food items on the culture of the mysid shrimp mysidopsis almyra (crustacea: Pericaridea) in a static water system. Aquaculture International, 9(5):393–400, September 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1020569624151.

Lygren:2001:IRD

[359] Bjarte Lygren, Brit Hjeltnes, and Rune Waagbø. Immune response and disease resistance in Atlantic salmon (Salmo salar L.) fed three levels of dietary vitamin E and the effect of vaccination on the liver status of antioxidant vitamins. Aquaculture International, 9(5):401–411, September 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1020509308400.

Townsend:2001:SSC

[360] Clarice Ramalho Townsend and Bernardo Baldisserotto. Survival of silver catfish fingerlings exposed to acute changes of water pH and hardness. Aquaculture International, 9(5):413–419, September 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1020592226860.

Lygren:2001:IDC

[361] Bjarte Lygren and Gro-Ingunn Hemre. Influence of dietary carbohydrate on antioxidant enzyme activities in liver of Atlantic salmon (Salmo salar L.). Aquaculture International, 9(5):421–427, September 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1020530432508.

Shishehchian:2001:ECB

[362] F. Shishehchian, F. M. Yusoff, and M. Shariff. The effects of commercial bacterial products on macrobenthos community in shrimp culture ponds. Aquaculture International, 9(5):429–436, September 2001. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1020582417487.

Xu:2001:EDF

[363] X. L. Xu, P. Fontaine, C. Mélard, and P. Kestemont. Effects of dietary fat levels on growth, feed efficiency and biochemical compositions of Eurasian perch Perca fluviatilis. Aquaculture International, 9(5): 437–449, September 2001. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1020597415669.

Escaffre:2001:DSA

[364] A. M. Escaffre, D. Bazin, and P. Bergot. Disinfection of Sparus aurata eggs with glutaraldehyde. Aquaculture International, 9(5):451–458, September 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1020538701557.

Jobling:2001:BRMa

[365] Malcolm Jobling. Book review: M. L. Jahncke, E. S. Garrett, A. Reilly, R. E. Martin and E. Cole (editors), Public, animal, and environmental aquaculture health issues. Aquaculture International, 9(5):459–460, September 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1020510032722.

Jobling:2001:BRMb

[366] Malcolm Jobling. Book review: M. B. New and W. C. Valenti (editors), Freshwater Prawn Culture: The farming of Macrobrachium rosenbergii. Aquaculture International, 9(5):461–462, September 2001. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1020546715884.

Parpoura:2001:EDD

[367] Alkistis C. R. Parpoura and Maria N. Alexis. Effects of different dietary oils in sea bass (Dicentrarchus labrax) nutrition. Aquaculture International, 9(6):463–476, November 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1020590701325.

Alvarino:2001:EMI

[368] J. M. R. Alvariño, P. G. Rebollar, M. Olmedo, B. Alvarez-Blázquez, E. Ubilla, and J. B. Peleteiro. Effects of melatonin implants on reproduction and growth of turbot broodstock. Aquaculture International, 9(6): 477–487, November 2001. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1020590111031.

Saez-Royuela:2001:EST

[369] M. Sáez-Royuela, J. M. Carral, J. D. Celada, and J. R. Pérez. Effects of shelter type and food supply frequency on survival and growth of stage-2 juvenile white-clawed crayfish (Austropotamobius pallipes Lereboullet) under laboratory conditions. Aquaculture International, 9(6):489–497, November 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1020509627870.

Fermin:2001:GCT

[370] Armando C. Fermin and Shela Mae Buen. Grow-out culture of tropical abalone, Haliotis asinina (Linnaeus) in suspended mesh cages with different shelter surface areas. Aquaculture International, 9(6):499–508, November 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1020535301193.

Papoutsoglou:2001:GPC

[371] S. E. Papoutsoglou, H. Miliou, N. P. Karakatsouli, M. Tzitzinakis, and S. Chadio. Growth and physiological changes in scaled carp and blue tilapia under behavioral stress in mono- and polyculture rearing using a recirculated water system. Aquaculture International, 9(6):509–518, November 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1020506814609.

Hernandez:2001:SEF

[372] M. D. Hernández, F. J. Martínez, and B. García García. Sensory evaluation of farmed sharpsnout seabream (Diplodus puntazzo). Aquaculture International, 9(6):519–529, November 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1020513931447.

Palacios:2001:RBL

[373] Elena Palacios, Ilie S. Racotta, Horacio Heras, Yanic Marty, Jeanne Moal, and Jean-François Samain. Relation between lipid and fatty acid composition of eggs and larval survival in white Pacific shrimp (Penaeus vannamei, Boone, 1931). Aquaculture International, 9(6):531–543, November 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1020589924810.

New:2001:BRK

[374] Michael B. New. Book review: K. V. Jayachandran, Palaemonid Prawns: Biodiversity, Taxonomy, Biology and Management. Aquaculture International, 9(6):545–548, November 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1020581119018.

Anonymous:2001:AIV

[375] Anonymous. Author index to volume 9 (2001). Aquaculture International, 9(6):549–550, November 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1020574823256.

Anonymous:2001:KIV

[376] Anonymous. Keyword index to volume 9 (2001). Aquaculture International, 9(6):551–552, November 2001. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1020536306418.

Anonymous:2001:CV

[377] Anonymous. Contents of volume 9 (2001). Aquaculture International, 9 (6):553–556, November 2001. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1020599303266.

Richter:2002:RMR

[378] Hartmut Richter, George Francis, and Klaus Becker. A reassessment of the maintenance ration of red tilapia. Aquaculture International, 10(1):1– 9, January 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1021316926354.

Rinchard:2002:EDC

[379] Jacques Rinchard, Godfrey Mbahinzireki, Konrad Dabrowski, Kyeong-Jun Lee, Mary-Ann Garcia-Abiado, and Joseph Ottobre. Effects of dietary cottonseed meal protein level on growth, gonad development and plasma sex steroid hormones of tropical fish tilapia Oreochromis sp. Aquaculture International, 10(1):11–28, January 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1021379328778.

Oca:2002:LBS

[380] J. Oca, L. Reig, and R. Flos. Is land-based sea bream production a feasible activity on the northwest Mediterranean coast? Analysis of production costs. Aquaculture International, 10(1):29–41, January 2002. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1021395522585.

Garen:2002:CSL

[381] P. Garen and J.-L. M. Martin. Could a seasonal-like reduction in light radiation intensity affect cultured shrimp (Penaeus stylirostris Stimpson) yield? Aquaculture International, 10(1):43–55, January 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1021312702438.

Molokwu:2002:EWH

[382] C. N. Molokwu and G. C. Okpokwasili. Effect of water hardness on egg hatchability and larval viability of Clarias gariepinus. Aquaculture International, 10(1):57–64, January 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1021395122919.

Parsons:2002:PSS

[383] G. Jay Parsons, Sandra E. Shumway, Sue Kuenstner, and Alexander Gryska. Polyculture of sea scallops (Placopecten magellanicus) suspended from salmon cages. Aquaculture International, 10(1):65–77, January 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1021324610930.

Garg:2002:DDA

[384] S. K. Garg and Anita Bhatnagar. Determination of dosage of Azotobacter and organic fertilizer for optimum nutrient release, net primary productivity and fish growth in fresh water fish ponds. Aquaculture International, 10(2):87–107, March 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1021388328025.

Bairagi:2002:EPB

[385] Abhinanda Bairagi, Keka Sarkar Ghosh, Sukanta Kumar Sen, and Arun Kumar Ray. Enzyme producing bacterial flora isolated from fish digestive tracts. Aquaculture International, 10(2):109–121, March 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1021355406412.

Vezzulli:2002:BCS

[386] Luigi Vezzulli, Elisabetta Chelossi, Giovanna Riccardi, and Mauro Fabiano. Bacterial community structure and activity in fish farm sediments of the Ligurian sea (Western Mediterranean). Aquaculture International, 10(2):123–141, March 2002. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1021365829687.

Suarez:2002:MEC

[387] M. D. Suárez, A. Sanz, J. Bazoco, and M. García-Gallego. Metabolic effects of changes in the dietary protein: carbohydrate ratio in eel (Angilla anguilla) and trout (Oncorhynchus mykiss). Aquaculture International, 10(2):143–156, March 2002. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1021371104839.

Jones:2002:PGA

[388] Paul L. Jones, Jose R. Chavez, and Brad D. Mitchell. Production of Australian freshwater crayfish in earthen-based systems using pelleted diets and forage crops as food. Aquaculture International, 10(2):157– 175, March 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1021304308880.

Nunez:2002:EMD

[389] Maximiano Nuñez, César Lodeiros, Marcos de Donato, and César Graziani. Evaluation of microalgae diets for Litopenaeus vannamei larvae using a simple protocol. Aquaculture International, 10(3):177–187, May 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1022102032684.

Flos:2002:IMD

[390] R. Flos, L. Reig, J. Oca, and M. Ginovart. Influence of marketing and different land-based systems on gilthead sea bream (Sparus aurata) quality. Aquaculture International, 10(3):189–206, May 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1022100928523.

Domingues:2002:ETL

[391] Pedro M. Domingues, António Sykes, and José P. Andrade. The effects of temperature in the life cycle of two consecutive generations of the cuttlefish Sepia officinalis (Linnaeus, 1758), cultured in the Algarve (South Portugal). Aquaculture International, 10(3):207–220, May 2002. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1022148802078.

Carneiro:2002:TSM

[392] P. C. F. Carneiro and E. C. Urbinati. Transport stress in matrinxã, Brycon cephalus (Teleostei: Characidae), at different densities. Aquaculture International, 10(3):221–229, May 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1022166411089.

Romvari:2002:NIM

[393] R. Romvári, Cs. Hancz, Zs. Petrási, T. Molnár, and P. Horn. Noninvasive measurement of fillet composition of four freshwater fish species by computer tomography. Aquaculture International, 10(3):231–240, May 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1022196413528.

Ruff:2002:CCS

[394] N. Ruff, R. D. Fitzgerald, T. F. Cross, and J. P. Kerry. Comparative composition and shelf-life of fillets of wild and cultured turbot (Scophthalmus maximus) and Atlantic halibut (Hippoglossus hippoglossus). Aquaculture International, 10(3):241–256, May 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1022175200191.

Samuelsen:2002:ESI

[395] Ole Bent Samuelsen, Arne Ervik, Lise Torkildsen, and Øivind Bergh. The efficacy of a single intraperitoneal injection of either flumequine or oxytetracycline hydrochloride in prevention of outbreaks of atypical Aeromonas salmonicida infection in goldsinny wrasse, Ctenolabrus rupestris L., following stress. Aquaculture International, 10(3):257–264, May 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1022187627019.

Doroudi:2002:EPS

[396] Mehdi S. Doroudi, Paul C. Southgate, and Robert J. Mayer. Evaluation of partial substitution of live algae with dried Tetraselmis for larval rearing of black-lip pearl oyster, Pinctada margaritifera (L.). Aquaculture International, 10(4):265–277, July 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1022492416267.

Tettelbach:2002:RHR

[397] Stephen T. Tettelbach, Christopher F. Smith, Peter Wenczel, and Edward Decort. Reproduction of hatchery-reared and transplanted wild bay scallops, Argopecten irradians irradians, relative to natural populations. Aquaculture International, 10(4):279–296, July 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1022429500337.

Dhawan:2002:EPD

[398] Asha Dhawan and Satinder Kaur. Effect of pig dung on water quality and polyculture of carp species during winter and summer. Aquaculture International, 10(4):297–307, July 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1022406800631.

Kelly:2002:SGR

[399] Maeve S. Kelly. Survivorship and growth rates of hatchery-reared sea urchins. Aquaculture International, 10(4):309–316, July 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1022414516929.

Peyghan:2002:HSE

[400] R. Peyghan and G. Azary Takamy. Histopathological, serum enzyme, cholesterol and urea changes in experimental acute toxicity of ammonia in common carp Cyprinus carpio and use of natural zeolite for prevention. Aquaculture International, 10(4):317–325, July 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1022408529458.

Lodeiros:2002:GSP

[401] César Lodeiros, Damelys Pico, Antulio Prieto, Noelis Narváez, and Alejandro Guerra. Growth and survival of the pearl oyster Pinctada imbricata (Röding 1758) in supended and bottom culture in the Golfo de Cariaco, Venezuela. Aquaculture International, 10(4):327–338, July 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1022451608405.

Gonzalez:2002:ETS

[402] M. L. González, D. A. López, M. C. Pérez, and J. M. Castro. Effect of temperature on the scope for growth in juvenile scallops Argopecten purpuratus (Lamarck, 1819). Aquaculture International, 10 (4):339–348, July 2002. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1022429209469.

Simoes:2002:FEL

[403] Fernando Simoes, Fernando Ribeiro, and David A. Jones. Feeding early larval stages of fire shrimp Lysmata debelius (Caridea, Hippolytidae). Aquaculture International, 10(5):349–360, September 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1023366423144.

Gimenez:2002:GFI

[404] F. Aguado Giménez and B. García García. Growth and food intake models in Octopus vulgaris Cuvier (1797): influence of body weight, temperature, sex and diet. Aquaculture International, 10(5):361–377, September 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1023335024053.

Gines:2002:SBS

[405] R. Ginés. Starvation before slaughtering as a tool to keep freshness attributes in gilthead sea bream (Sparus aurata). Aquaculture International, 10(5):379–389, September 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1023365025292.

Harlioglu:2002:EDV

[406] Muzaffer Mustafa Harlioğlu, Kenan Köprücü, and Yaşar Özdemir. The effect of dietary vitamin E on the pleopodal egg number of Astacus leptodactylus (Eschscholtz, 1823). Aquaculture International, 10(5):391–397, September 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1023376705651.

Torkildsen:2002:TEL

[407] Lise Torkildsen. Treatment of the early life stages of scallop (Pecten maximus) with antimicrobial agents; searching for an alternative to chloramphenicol. Aquaculture International, 10(5):399–409, September 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1023347020675.

Jonassen:2002:EPS

[408] Thor M. Jonassen. Effects of photoperiod, stocking density and diet on growth in young spotted wolffish (Anarhichas minor Olafsen). Aquaculture International, 10(5):411–420, September 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1023374921581.

Lanari:2002:ECD

[409] D. Lanari and E. D’Agaro. Effects of changes in dietary DP/DE ratio during the growing phase on growth performance and carcass composition of rainbow trout (Oncorhynchus mykiss, Walbaum). Aquaculture International, 10(5):421–432, September 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1023396524383.

Chong:2002:IPA

[410] Alexander Chong, Roshada Hashim, and Ahyaudin Bin Ali. Inhibition of protease activities in discus Symphysodon spp. by three plant meals. Aquaculture International, 10(5):433–441, September 2002. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1023358216960.

Carter:2002:BRE

[411] C. Carter. Book review: Ecological Aquaculture: The Evolution of the Blue Revolution, B. A. Costa–Pierce (ed.). Aquaculture International, 10 (5):443–445, September 2002. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1023369110271.

Burnell:2002:E

[412] G. Burnell, P. Goulletquer, and S. Stead. Editorial. Aquaculture International, 10(6):3–4, November 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1023963402665.

Stead:2002:ARI

[413] S. M. Stead, G. Burnell, and P. Goulletquer. Aquaculture and its role in integrated coastal zone management. Aquaculture International, 10(6): 447–468, November 2002. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1023926421723.

Kaiser:2002:UVG

[414] M. Kaiser and S. M. Stead. Uncertainties and values in European aquaculture: communication, management and policy issues in times of “changing public perceptions”. Aquaculture International, 10(6): 469–490, November 2002. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1023963326201.

Hishamunda:2002:MPP

[415] Nathanael Hishamunda and Neil B. Ridler. Macro policies to promote sustainable commercial aquaculture. Aquaculture International, 10(6): 491–505, November 2002. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1023985430206.

Goulletquer:2002:SFC

[416] P. Goulletquer and O. Le Moine. Shellfish farming and Coastal Zone Management (CZM) development in the Marennes–Oléron Bay and Charentais Sounds (Charente Maritime, France): a review of recent developments. Aquaculture International, 10(6):507–525, November 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1023975418669.

Kuhlmann:2002:EMR

[417] K.-J. Kühlmann. Evaluations of marine reserves as basis to develop alternative livelihoods in coastal areas of the Philippines. Aquaculture International, 10(6):527–549, November 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1023955626357.

Spennemann:2002:TMA

[418] Dirk H. R. Spennemann. Traditional milkfish aquaculture in Nauru. Aquaculture International, 10(6):551–562, November 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1023900601000.

Anonymous:2002:AIV

[419] Anonymous. Author index to volume 10 (2002). Aquaculture International, 10(6):563–564, November 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1023987418268.

Anonymous:2002:KIV

[420] Anonymous. Keyword index to volume 10 (2002). Aquaculture International, 10(6):565–566, November 2002. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1023983302338.

Anonymous:2002:CV

[421] Anonymous. Contents of volume 10. Aquaculture International, 10(6): 567–570, November 2002. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1023935403247.

Wille:2003:P

[422] Mathieu Wille and Patrick Sorgeloos. Preface. Aquaculture International, 11(1–2):1–2, January 2003. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1024137403187.

Djunaidah:2003:RPO

[423] Iin S. Djunaidah. Reproductive performance and offspring quality in mud crab (Scylla paramamosain) broodstock fed different diets. Aquaculture International, 11(1–2):3–15, January 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1024188507215.

Boglione:2003:MLF

[424] C. Boglione. Morphoecology in larval fin-fish: a new candidate species for aquaculture, Diplodus puntazzo (Sparidae). Aquaculture International, 11(1–2):17–41, January 2003. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1024119032359.

Zhu:2003:LAA

[425] Peihong Zhu, Christopher C. Parrish, and Joseph A. Brown. Lipid and amino acid metabolism during early development of Atlantic halibut (Hippoglossus hippoglossus). Aquaculture International, 11(1–2):43–52, January 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1024128200224.

Liddy:2003:SGI

[426] G. C. Liddy, B. F. Phillips, and G. B. Maguire. Survival and growth of instar 1 phyllosoma of the western rock lobster, Panulirus cygnus, starved before or after periods of feeding. Aquaculture International, 11 (1–2):53–67, January 2003. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1024100110378.

Ponis:2003:APG

[427] Emanuele Ponis. Assessment of the performance of Pacific oyster (Crassostrea gigas) larvae fed with fresh and preserved Pavlova lutheri concentrates. Aquaculture International, 11(1–2):69–79, January 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1024176210427.

Mitra:2003:DEA

[428] Gopa Mitra and P. K. Mukhopadhyay. Dietary essentiality of ascorbic acid in rohu larvae: Quantification with ascorbic acid enriched zooplankton. Aquaculture International, 11(1–2):81–93, January 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1024158226110.

Martinez-Diaz:2003:EAM

[429] Sergio F. Martínez-Díaz. Elimination of the associated microbial community and bioencapsulation of bacteria in the rotifer Brachionus plicatilis. Aquaculture International, 11(1–2):95–108, January 2003. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1024117109362.

Castell:2003:EDH

[430] John Castell. The effect of different HUFA enrichment emulsions on the nutritional value of rotifers (Brachionus plicatilis) fed to larval haddock (Melanogrammus aeglefinus). Aquaculture International, 11(1–2): 109–117, January 2003. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1024154106656.

Shields:2003:EDT

[431] R. J. Shields. Effects of diet transition regimen on survival, growth and lipid composition of intensively reared Atlantic cod, Gadus morhua, larvae. Aquaculture International, 11(1–2):119–130, January 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1024120301031.

Willey:2003:AAR

[432] Stephen Willey, David A. Bengtson, and Moti Harel. Arachidonic acid requirements in larval summer flounder, Paralichthys dentatus. Aquaculture International, 11(1–2):131–149, January 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1024148625202.

Monroig:2003:EAN

[433] Óscar Monroig. Enrichment of Artemia nauplii in PUFA, phospholipids, and water-soluble nutrients using liposomes. Aquaculture International, 11(1–2):151–161, January 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1023/A:1024199305622.

Ritar:2003:APS

[434] Arthur J. Ritar. Artemia prey size and mode of presentation: Effects on the survival and growth of phyllosoma larvae of southern rock lobster (Jasus edwardsii). Aquaculture International, 11(1–2):163–182, January 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1024160908454.

Garcia-Ortega:2003:WBP

[435] A. García-Ortega, I. Abdo, and C. Hernández. Weaning of bullseye puffer (Sphoeroides annulatus) from live food to microparticulate diets made with decapsulated cysts of Artemia and fishmeal. Aquaculture International, 11(1–2):183–194, January 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1024100618838.

Tocher:2003:CSA

[436] D. R. Tocher. Comparative study of antioxidant defence mechanisms in marine fish fed variable levels of oxidised oil and vitamin E. Aquaculture International, 11(1–2):195–216, January 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1024127003997.

Otteraa:2003:FPJ

[437] Håkon Otterå. Feed preferences in juvenile cod estimated by inert lanthanid markers — effects of moisture content in the feed. Aquaculture International, 11(1–2):217–224, January 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1024144026308.

Domingues:2003:ECD

[438] P. Domingues, R. Poirier, L. Dickel, E. Almansa, A. Sykes, and J. P. Andrade. Effects of culture density and live prey on growth and survival of juvenile cuttlefish, Sepia officinalis. Aquaculture International, 11 (3):225–242, May 2003. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1024803802486.

Boscolo:2003:CIA

[439] Rossella Boscolo, Michele Cornello, and Otello Giovanardi. Condition index and air survival time to compare three kinds of Manila clam Tapes philippinarum (Adams & Reeve) farming systems. Aquaculture International, 11(3):243–254, May 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1023/A:1024888608791.

Morehead:2003:DST

[440] D. T. Morehead and P. R. Hart. Disinfection of striped trumpeter (Latris lineata) eggs with glutaraldehyde. Aquaculture International, 11(3):255– 260, May 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1024884620233.

Cecchini:2003:DME

[441] S. Cecchini and A. M. Cognetti-Varriale. Dehydration is more effective for the control of embryonic development and larval hatching of Diplectanum aequans (Monogenea, Diplectanidae) than formalin and trichlorphon. Aquaculture International, 11(3):261–265, May 2003. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1024892822958.

Ridha:2003:EDS

[442] M. T. Ridha and E. M. Cruz. Effect of different schedules for broodstock exchange on the seed production of Nile tilapia Oreochromis niloticus (L.) in freshwater. Aquaculture International, 11(3):267–276, May 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1024853107029.

Zakes:2003:ROC

[443] Zdzisław Zakeś, Krystyna Demska-zakeś, and Krzysztof Kata. Rates of oxygen consumption and ammonia excretion of juvenile Eurasian perch Perca fluviatilis L. Aquaculture International, 11(3):277–288, May 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1023/A:1024839903568.

Ballestrazzi:2003:EDC

[444] Rodolfo Ballestrazzi, Simona Rainis, Francesca Tulli, and Arianna Bracelli. The effect of dietary coconut oil on reproductive traits and egg fatty acid composition in rainbow trout (Oncorhynchus mykiss). Aquaculture International, 11(3):289–299, May 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/A:1024876024720.

Poli:2003:PRQ

[445] B. M. Poli, G. Parisi, G. Zampacavallo, F. Iurzan, M. Mecatti, P. Lupi, and A. Bonelli. Preliminary results on quality and quality changes in reared meagre (Argyrosomus regius): body and fillet traits and freshness changes in refrigerated commercial-size fish. Aquaculture International, 11(3):301–311, May 2003. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1024840804303.

McLean:2003:BRR

[446] Ewen McLean. Book review: Responsible Marine Aquaculture. R. R. Stickney and J. P. McVey (eds). Aquaculture International, 11(3):313– 316, May 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1024844120406.

Jobling:2003:BRF

[447] Malcolm Jobling. Book review: Fishery Science: The Unique Contributions of Early Life Stages. L. A. Fuiman and R. G. Werner (eds). Aquaculture International, 11(3):316–317, May 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1024861102659.

Jorgensen:2003:BRB

[448] Even H. Jørgensen. Book review: Biological Indicators of Aquatic Ecosystem Stress. S. M. Adams (ed.). Aquaculture International, 11(3):318– 319, May 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1024804006452.

Jobling:2003:BRD

[449] Malcolm Jobling. Book review: Diseases and Disorders of Finfish in Cage Culture, Woo, P. K. T., D. W. Bruno and L. H. S. Lim (eds). Aquaculture International, 11(3):320–322, May 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1024860421224.

Adamek:2003:E

[450] Zdenek Adámek. Editorial. Aquaculture International, 11(4):323, July 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1025767122381.

Pipalova:2003:GCC

[451] Irena Pípalová. Grass carp (Ctenopharyngodon idella) grazing on duckweed (Spirodela polyrhiza). Aquaculture International, 11(4):325–336, July 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1025703227216.

Peterka:2003:DGL

[452] J. Peterka, J. Matína, and J. Lipka. The diet and growth of larval and juvenile pikeperch (Stizostedion lucioperca (L.)): a comparative study of fishponds and a reservoir. Aquaculture International, 11(4):337– 348, July 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1025791208123.

Szumiec:2003:EGT

[453] Maria A. Szumiec and Henryk Bialowas. Effect of genetics and temperature on carp juvenile survival in ponds of the temperate climat. Aquaculture International, 11(4):349–356, July 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/A:1025765506531.

Linhart:2003:ICO

[454] Otomar Linhart, Steve D. Mims, Boris Gomelsky, Ana E. Hiott, William L. Shelton, Jacky Cosson, Marek Rodina, David Gela, and Jan Bastl. Ionic composition and osmolality of paddlefish (Polyodon spathula, acipenseriformes) seminal fluid. Aquaculture International, 11(4):357– 368, July 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1025773707439.

Kocour:2003:RCG

[455] Martin Kocour, Otomar Linhart, and David Gela. Results of comparative growing test of all-female and bisexual population in two-year-old common carp (Cyprinus carpio L.). Aquaculture International, 11(4):369– 378, July 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/A:
1025728921552.

Gela:2003:TCE

[456] David Gela, Marek Rodina, and Otomar Linhart. Top-crossing with evaluation of slaughtering value in common carp (Cyprinus carpio L.) offspring. Aquaculture International, 11(4):379–387, July 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1023/A:1025721723369.

Adamek:2003:IOL

[457] Zdeněk Adámek, David Kortan, Pavel Lepič, and Jaroslav Andreji. Impacts of otter (Lutra lutra L.) predation on fishponds: a study of fish remains at ponds in the Czech Republic. Aquaculture International, 11 (4):389–396, July 2003. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/A:1025787330366.

Domingues:2003:EFL

[458] Pedro Domingues, António Sykes, Anne Sommerfield, and José P. Andrade. Effects of feeding live or frozen prey on growth, survival and the life cycle of the cuttlefish, Sepia officinalis (Linnaeus, 1758). Aquaculture International, 11(5):397–410, September 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/B:AQUI.0000004195.92236.3a.

Huchette:2003:EDB

[459] Sylvain M. H. Huchette, C. S. Koh, and Rob W. Day. The effects of density on the behaviour and growth of juvenile blacklip abalone (Haliotis rubra). Aquaculture International, 11(5):411–428, September 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000004194.85299.38.

Boudry:2003:IGV

[460] P. Boudry, B. Collet, H. McCombie, B. Ernande, B. Morand, S. Heurtebise, and A. Gérard. Individual growth variation and its relationship with survival in juvenile Pacific oysters, Crassostrea gigas (Thunberg). Aquaculture International, 11(5):429–448, September 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/B:AQUI.0000004189.49517.c0.

Ahmed:2003:DMR

[461] Imtiaz Ahmed, Mukhtar A. Khan, and A. K. Jafri. Dietary methionine requirement of fingerling Indian major carp, Cirrhinus mrigala (Hamilton). Aquaculture International, 11(5):449–462, September 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000004181.89420.a2.

Imsland:2003:GAF

[462] Albert K. Imsland and Thor M. Jonassen. Growth and age at first maturity in turbot and halibut reared under different photoperiods. Aquaculture International, 11(5):463–475, September 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/B:AQUI.0000004191.43885.b2.

Hoehne-Reitan:2003:LAD

[463] K. Hoehne-Reitan, E. Kjørsvik, and K. I. Reitan. Lipolytic activities in developing turbot larvae as influenced by diet. Aquaculture International, 11(5):477–489, September 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/B:AQUI.0000004192.36467.0d.

Pousao-Ferreira:2003:EEM

[464] Pedro Pousão-Ferreira, Paulo Santos, António Paulo Carvalho, Sofia Morais, and Luís Narciso. Effect of an experimental microparticulate diet on the growth, survival and fatty acid profile of gilthead seabream (Sparus aurata L.) larvae. Aquaculture International, 11(5):491–504, September 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000004190.13871.f3.

Patra:2003:EAN

[465] S. K. Patra and K. S. Mohamed. Enrichment of Artemia nauplii with the probiotic yeast Saccharomyces boulardii and its resistance against a pathogenic Vibrio. Aquaculture International, 11(5):505–514, September 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000004193.40039.54.

Anonymous:2003:BRF

[466] Anonymous. Book review: Farming Freshwater Prawns; A manual for the Culture of the Giant River Prawn. Aquaculture International, 11(5): 515–516, September 2003. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/B:AQUI.0000004459.05434.bb.

Frankic:2003:SAD

[467] Anamarija Frankic and Carl Hershner. Sustainable aquaculture: developing the promise of aquaculture. Aquaculture International, 11(6): 517–530, November 2003. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/B:AQUI.0000013264.38692.91.

Sykes:2003:ICD

[468] António V. Sykes, Pedro M. Domingues, Maria Loyd, Anne Sommerfield, and José P. Andrade. The influence of culture density and enriched environments on the first stage culture of young cuttlefish, Sepia officinalis (Linnaeus, 1758). Aquaculture International, 11(6):531–544, November 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000013262.15437.e4.

Pit:2003:FPH

[469] Josiah H. Pit and Paul C. Southgate. Fouling and predation; how do they affect growth and survival of the blacklip pearl oyster, Pinctada margaritifera, during nursery culture? Aquaculture International, 11(6): 545–555, November 2003. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/B:AQUI.0000013310.17400.97.

Borges:2003:POS

[470] Maria-Teresa Borges, Aurora Morais, and Paula M. L. Castro. Performance of outdoor seawater treatment systems for recirculation in an intensive turbot (Scophthalmus maximus) farm. Aquaculture International, 11(6):557–570, November 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/B:AQUI.0000013308.33265.62.

Kheriji:2003:STE

[471] S. Khériji, M. El Cafsi, W. Masmoudi, J. D. Castell, and M. S. Romdhane. Salinity and temperature effects on the lipid composition of mullet sea fry (Mugil cephalus, Linne, 1758). Aquaculture International, 11(6): 571–582, November 2003. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/B:AQUI.0000013321.93743.6d.

Citarasu:2003:IAH

[472] T. Citarasu, K. Venkatramalingam, M. Micheal Babu, R. Raja Jeya Sekar, and M. Petermarian. Influence of the antibacterial herbs, Solanum trilobatum, Andrographis paniculata and Psoralea corylifolia on the survival, growth and bacterial load of Penaeus monodon post larvae. Aquaculture International, 11(6):581–595, November 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/B:AQUI.0000013322.53358.53.

Peyghan:2003:SBC

[473] R. Peyghan, M. Razijalaly, M. Baiat, and A. Rasekh. Study of bioaccumulation of copper in liver and muscle of common carp Cyprinus carpio after copper sulfate bath. Aquaculture International, 11(6):597–604, November 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000013323.69018.ff.

Anonymous:2003:BRB

[474] Anonymous. Book review: Bivalve molluses, ecology and culture. Aquaculture International, 11(6):605–607, November 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/B:AQUI.0000013398.37740.fb.

Anonymous:2003:AIV

[475] Anonymous. Author index to volume 11 (2003). Aquaculture International, 11(6):609–610, November 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/B:AQUI.0000013387.94128.01.

Anonymous:2003:KIV

[476] Anonymous. Keyword index to volume 11 (2003). Aquaculture International, 11(6):611–612, November 2003. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/B:AQUI.0000013388.88977.79.

Anonymous:2003:CV

[477] Anonymous. Contents of volume 11. Aquaculture International, 11(6): 613–617, November 2003. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/B:AQUI.0000013389.75996.8f.

Anonymous:2004:P

[478] Anonymous. Preface. Aquaculture International, 12(1):1, January 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000017212.43148.2a.

Kaminski:2004:FAA

[479] Rafał Kamiński, Jan Kusznierz, Leszek Myszkowski, and Jacek Wolnicki. The first attempt to artificially reproduce the endangered cyprinid lake minnow Eupallasella perenurus (Pallas). Aquaculture International, 12 (1):3–10, January 2004. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/B:AQUI.0000017182.17408.85.

Zakes:2004:ISS

[480] Zdzislaw Zakes and Miroslaw Szczepkowski. Induction of out-of-season spawning of pikeperch, Sander lucioperca (L.). Aquaculture International, 12(1):11–18, January 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1023/B:AQUI.0000017183.40691.7d.

Gross:2004:GIB

[481] Riho Gross, Bernhard Gum, Reinhard Reiter, and Ralph Kühn. Genetic introgression between Arctic charr (Salvelinus alpinus) and brook trout (Salvelinus fontinalis) in Bavarian hatchery stocks inferred from nuclear and mitochondrial DNA markers. Aquaculture International, 12(1):19– 32, January 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000017184.92614.e1.

Paaver:2004:GRM

[482] Tiit Paaver, Riho Gross, and Piret Ilves. Growth rate, maturation level and flesh quality of three strains of large rainbow trout (Oncorhynchus mykiss) reared in Estonia. Aquaculture International, 12(1):33–45, January 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000017185.10472.1d.

Urbanyi:2004:SHA

[483] Béla Urbányi, Ákos Horváth, and Balázs Kovács. Successful hybridization of Acipenser species using cryopreserved sperm. Aquaculture International, 12(1):47–56, January 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/B:AQUI.0000017186.59304.6d.

Kowalski:2004:CGA

[484] R. Kowalski, J. Glogowski, D. Kucharczyk, M. Mak, S. Dobosz, Z. Zakes, and A. Ciereszko. Characterization of gelatinolytic activity in seminal plasma of some teleost fish. Aquaculture International, 12(1):57–68, January 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000017188.36365.37.

Cosson:2004:IOF

[485] Jacky Cosson. The ionic and osmotic factors controlling motility of fish spermatozoa. Aquaculture International, 12(1):69–85, January 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000017189.44263.bc.

Migaud:2004:SSG

[486] Herve Migaud, Jean-Noel Gardeur, Patrick Kestemont, and Pascal Fontaine. Off-season spawning of Eurasian perch Perca fluviatilis. Aquaculture International, 12(1):87–102, January 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/B:AQUI.0000017190.15074.6c.

Flajshans:2004:FRR

[487] Martin Flajshans, Martin Kocour, David Gela, and Veronika Piackova. The first results on relationships among amphimictic diploid, diploid gynogenic and triploid tench, Tinca tinca L. under communal testing. Aquaculture International, 12(1):103–118, January 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:/

/link.springer.com/article/10.1023/B:AQUI.0000017191.52711.
53.

Rodina:2004:KSI

[488] M. Rodina, J. Cosson, D. Gela, and O. Linhart. Kurokura solution as immobilizing medium for spermatozoa of tench (Tinca tinca L.). Aquaculture International, 12(1):119–131, January 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/B:AQUI.0000017192.75993.e3.

Adamek:2004:SFC

[489] Zdeněk Adámek, Jaroslav Andreji, and Alan Henshaw. Stripping fecundity of common bream (Abramis brama L.) from the Rivers Trent and Sow (Nottinghamshire, UK). Aquaculture International, 12(1):133–137, January 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000017187.04566.99.

Arnold:2004:ICA

[490] William S. Arnold, Sarah L. Walters, Jonathan S. Fajans, Sarah C. Peters, and Theresa M. Bert. Influence of congeneric aquaculture on hard clam (Mercenaria spp.) population genetic structure. Aquaculture International, 12(2):139–160, March 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/B:AQUI.0000032078.90352.03.

Thunjai:2004:QLM

[491] Taworn Thunjai, Claude E. Boyd, and Mali Boonyaratapalin. Quality of liming materials used in aquaculture in Thailand. Aquaculture International, 12(2):161–168, March 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/B:AQUI.0000032062.89145.5b.

Ali:2004:ODC

[492] M. Z. Ali and K. Jauncey. Optimal dietary carbohydrate to lipid ratio in African catfish Clarias gariepinus (Burchell 1822). Aquaculture International, 12(2):169–180, March 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/B:AQUI.0000032065.28059.5b.

Molnar:2004:EIS

[493] T. Molnár, Cs. Hancz, M. Bódis, T. Müller, M. Bercsényi, and P. Horn. The effect of initial stocking density on growth and survival of pike-perch fingerlings reared under intensive conditions. Aquaculture International, 12(2):181–189, March 2004. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/B:AQUI.0000032079.62056.8c.

Papandroulakis:2004:BPR

[494] Nikos Papandroulakis, Maroudio Kentouri, and Pascal Divanach. Biological performance of red porgy (Pagrus pagrus) larvae under intensive rearing conditions with the use of an automated feeding system. Aquaculture International, 12(2):191–203, March 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/B:AQUI.0000032080.59789.5f.

DeLaGandara:2004:ECF

[495] F. De La Gándara, M. Jover, and A. García-Gómez. Effect of continuous food supply on the oxygen consumption of young Mediterranean yellowtail (Seriola dumerili Risso, 1810). Aquaculture International, 12 (2):205–213, March 2004. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/B:AQUI.0000032081.04020.51.

Yang:2004:IOO

[496] Zhou Yang and Ya-Fen Chen. Induced ovulation in obscure puffer Takifugu obscurus by injections of LHRH-a. Aquaculture International, 12 (2):215–223, March 2004. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/B:AQUI.0000032082.17825.f2.

Anonymous:2004:E

[497] Anonymous. Erratum. Aquaculture International, 12(2):225–236, March 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000032122.95518.f0.

Anonymous:2004:BRB

[498] Anonymous. Book review: Botechnology and genetics in fisheries and aquaculture. Aquaculture International, 12(2):237–238, March 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000032132.30077.54.

Arvanitoyannis:2004:MSG

[499] I. S. Arvanitoyannis, A. Krystallis, P. Panagiotaki, and A. J. Theodorou. A marketing survey on Greek consumers’ attitudes towards fish. Aquaculture International, 12(3):259–279, May 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/B:AQUI.0000036137.29397.12.

Basaran:2004:SRB

[500] Fatih Basaran and Necati Samsun. Survival rates of Black Sea turbot (Psetta maxima maeotica, L. 1758) broodstock captured by gill nets from different depths and their adaptation culture conditions. Aquaculture International, 12(3):321–331, May 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/B:AQUI.0000036183.39217.2a.

Morais:2004:LEG

[501] S. Morais, L. Narciso, E. Dores, and P. Pousão-Ferreira. Lipid enrichment for Senegalese sole (Solea senegalensis) larvae: effect on larval growth, survival and fatty acid profile. Aquaculture International, 12(3):281– 298, May 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000036184.13187.6b.

Batzios:2004:GGC

[502] Ch. Batzios, P. Angelidis, E. P. Papapanagiotou, D. K. Moutopoulos, Ch. Anastasiadou, and V. Chrisopolitou. Greek consumer’s image of the cultured mussel market. Aquaculture International, 12(3):239– 257, May 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000036185.15659.68.

Klaoudatos:2004:PEC

[503] S. D. Klaoudatos, G. Iakovopoulos, and D. S. Klaoudatos. Pagellus erythrinus (Common Pandora): a promising candidate species for enlarging the diversity of aquaculture production. Aquaculture International, 12(3):299–320, May 2004. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/B:AQUI.0000036186.31318.4a.

Valdes:2004:SCG

[504] P. Valdés, A. García-Alcázar, I. Abdel, M. Arizcun, C. Suárez, and E. Abellán. Seasonal changes on gonadosomatic index and maturation stages in common pandora Pagellus erythrinus (L.). Aquaculture International, 12(4–5):333–343, July 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/B:AQUI.0000042136.91952.9e.

Papandroulakis:2004:MRT

[505] N. Papandroulakis, M. Kentouri, E. Maingot, and P. Divanach. Mesocosm: a reliable technology for larval rearing of Diplodus puntazzo and Diplodus sargus sargus. Aquaculture International, 12(4–5):345– 355, July 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000042134.21211.ab.

Corsi:2004:MFO

[506] Ilaria Corsi, Antonio Mazzola, and Silvano Focardi. Mixed function oxidase activity and organochlorine levels in farmed sharpsnout seabream (Diplodus puntazzo) from two intensive aquaculture facilities. Aquaculture International, 12(4–5):357–375, July 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/B:AQUI.0000042132.39510.23.

Carnevali:2004:APS

[507] Oliana Carnevali, Maria Claudia Zamponi, Roberto Sulpizio, Arianna Rollo, Miria Nardi, Carla Orpianesi, Stefania Silvi, Massimo Caggiano, Alberta Maria Polzonetti, and Alberto Cresci. Administration of probiotic strain to improve sea bream wellness during development. Aquaculture International, 12(4–5):377–386, July 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/B:AQUI.0000042141.85977.bb.

Tutman:2004:PIF

[508] Pero Tutman, Nikša Glavić, Valter Kožul, Boško Skaramuca, and Branko Glamuzina. Preliminary information on feeding and growth of pompano, Trachinotus ovatus (Linnaeus, 1758) (Pisces; Carangidae) in captivity. Aquaculture International, 12(4–5):387–393, July 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:/

/link.springer.com/article/10.1023/B:AQUI.0000042135.88381.
f4.

Papandroulakis:2004:FRG

[509] N. Papandroulakis, M. Suquet, M. T. Spedicato, A. Machias, C. Fauvel, and P. Divanach. Feeding rates, growth performance and gametogenesis of wreckfish (Polyprion americanus) kept in captivity. Aquaculture International, 12(4–5):395–407, July 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1023/B:AQUI.0000042133.69455.95.

Sadek:2004:GSF

[510] S. Sadek, M. Fathy Osman, and M. Adel Mansour. Growth, survival and feed conversion rates of sea bream (Sparus aurata) cultured in earthen brackish water ponds fed different feed types. Aquaculture International, 12(4–5):409–421, July 2004. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/B:AQUI.0000042131.29346.93.

Kian:2004:BCE

[511] Annita Yong Seok Kian, Saleem Mustafa, and Ridzwan A. Rahman. Broodstock condition and egg quality in tiger prawn, Penaeus monodon, resulting from feeding bioencapsulated live prey. Aquaculture International, 12(4–5):423–433, July 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1023/B:AQUI.0000042139.26018.42.

Dimmock:2004:IEM

[512] A. Dimmock, I. Williamson, and P. B. Mather. The influence of environment on the morphology of Macrobrachium australiense (Decapoda: Palaemonidae). Aquaculture International, 12(4–5):435–456, July 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000042140.48340.c8.

Soltani:2004:ATA

[513] M. Soltani, G. H. Marmari, and M. R. Mehrabi. Acute toxicity and anesthetic effects of clove oil in Penaeus semisulcatus under various water quality conditions. Aquaculture International, 12(4–5):457–466, July 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1023/B:
AQUI.0000042137.26705.7b.

Taylor:2004:GAJ

[514] Matthew H. Taylor and Elena Tsvetnenko. A growth assessment of juvenile abalone Haliotis laevigata fed enriched macroalgae Ulva rigida. Aquaculture International, 12(4–5):467–480, July 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:/

/link.springer.com/article/10.1023/B:AQUI.0000042138.14961.
c2.

Iglesias:2004:CLC

[515] J. Iglesias, J. J. Otero, C. Moxica, L. Fuentes, and F. J. Sánchez. The completed life cycle of the octopus (Octopus vulgaris, Cuvier) under culture conditions: Paralarval rearing using Artemia and zoeae, and first data on juvenile growth up to 8 months of age. Aquaculture International, 12(4–5):481–487, July 2004. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1023/B:AQUI.0000042142.88449.bc.

Torkildsen:2004:HPS

[516] Lise Torkildsen and Thorolf Magnesen. Hatchery production of scallop larvae (Pecten maximus) — survival in different rearing systems. Aquaculture International, 12(4–5):489–507, July 2004. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1023/B:AQUI.0000042143.53903.21.

Nasution:2005:LTEa

[517] Syafruddin Nasution and David Roberts. Laboratory trials on the effects of different diets on growth and survival of the common whelk, Buccinum undatum L. 1758, as a candidate species for aquaculture. Aquaculture International, 12(6):509–521, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-005-5745-2.

Nasution:2005:LTEb

[518] Syafruddin Nasution and David Roberts. Laboratory trials on the effects of different diets on growth and survival of the common whelk, Buccinum undatum L. 1758, as a candidate species for aquaculture. Aquaculture International, 12(6):509–521, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-004-5745-7.

Abdel:2005:AJS

[519] I. Abdel, E. Abelløn, O. Løpez-Albors, P. Valdøs, M. J. Nortes, and A. Garcøa-Alcøzar. Abnormalities in the juvenile stage of sea bass (Dicentrarchus labrax L.) reared at different temperatures: types, prevalence and effect on growth. Aquaculture International, 12(6):523–538, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-004-0349-9.

Lennard:2005:CRFa

[520] Wilson A. Lennard and Brian V. Leonard. A comparison of reciprocating flow versus constant flow in an integrated, gravel bed, aquaponic test system. Aquaculture International, 12(6):539–553, January 2005. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-005-8528-x.

Lennard:2005:CRFb

[521] Wilson A. Lennard and Brian V. Leonard. A comparison of reciprocating flow versus constant flow in an integrated, gravel bed, aquaponic test system. Aquaculture International, 12(6):539–553, January 2005. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-8528-2.

Soto-Hernandez:2005:GDHa

[522] J. Soto-Hernández and J. M. Grijalva-Chon. Genetic differentiation in hatchery strains and wild white shrimp Penaeus (Litopenaeus) vannamei (Boone, 1931) from northwest Mexico. Aquaculture International, 12(6): 593–601, January 2005. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-5750-5.

Balasubramanian:2005:ZWE

[523] C. P. Balasubramanian, S. M. Pillai, and P. Ravichandran. Zerowater exchange shrimp farming systems (extensive) in the periphery of Chilka lagoon, Orissa, India. Aquaculture International, 12(6):555– 572, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-004-0350-3.

Naegel:2005:CGS

[524] Ludwig C. A. Naegel and Sonia Rodríguez-Astudillo. Comparison of growth and survival of white shrimp postlarvae (Litopenaeus vannamei) fed dried Artemia biomass versus four commercial feeds and three crustacean meals. Aquaculture International, 12(6):573–581, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-1024-x.

Linan-Cabello:2005:IFD

[525] Marco Agustín Liñán-Cabello and Jesús Paniagua-Michel. Induction factors derived from carotenoids and vitamin a during the ovarian maturation of Litopenaeus vannamei. Aquaculture International, 12(6): 583–592, January 2005. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-004-1088-7.

Soto-Hernandez:2005:GDHb

[526] J. Soto-Hernández and J. M. Grijalva-Chon. Genetic differentiation in hatchery strains and wild white shrimp Penaeus (Litopenaeus) vannamei (Boone, 1931) from northwest Mexico. Aquaculture International, 12(6): 593–601, January 2005. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-004-5750-x.

Anonymous:2005:BR

[527] Anonymous. Book reviews. Aquaculture International, 12(6):603–611, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-004-0399-z.

Miggiano:2005:AMG

[528] E. Miggiano, S. De Innocentiis, A. Ungaro, L. Sola, and D. Crosetti. AFLP and microsatellites as genetic tags to identify cultured gilthead seabream escapees: data from a simulated floating cage breaking event. Aquaculture International, 13(1–2):??, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-004-9024-4.

Focardi:2005:SIS

[529] S. Focardi, I. Corsi, and E. Franchi. Safety issues and sustainable development of European aquaculture: new tools for environmentally sound aquaculture. Aquaculture International, 13(1–2):??, January 2005. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-9036-0.

Focardi:2005:P

[530] S. Focardi. Preface. Aquaculture International, 13(1–2):1, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-9021-7.

Epel:2005:UCD

[531] D. Epel. Using cell and developmental biology to enhance embryo survival in aquaculture. Aquaculture International, 13(1–2):19–28, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-004-9037-z.

Poli:2005:FWQ

[532] B. M. Poli, G. Parisi, F. Scappini, and G. Zampacavallo. Fish welfare and quality as affected by pre-slaughter and slaughter management. Aquaculture International, 13(1–2):29–49, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-004-9035-1.

Gornati:2005:MBF

[533] R. Gornati, S. Gualdoni, R. Cavaliere, G. Terova, M. Saroglia, and G. Bernardini. Molecular biology and fish welfare: a winning combination. Aquaculture International, 13(1–2):51–55, January 2005. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-9034-2.

Pellizzato:2005:CFQ

[534] M. Pellizzato and L. Da Ros. Clam farming quality as a management tool: a proposal based on recent studies in Northern Adriatic lagoons. Aquaculture International, 13(1–2):57–66, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-004-9030-6.

Caruso:2005:HBI

[535] G. Caruso, L. Genovese, G. Maricchiolo, and A. Modica. Haematological, biochemical and immunological parameters as stress indicators in Dicentrarchus labrax and Sparus aurata farmed in off-shore cages. Aquaculture International, 13(1–2):67–73, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-004-9031-5.

Marin:2005:EMS

[536] M. G. Marin, V. Moschino, F. Meneghetti, and L. Da Ros. Effects of mechanical stress in under-sized clams, Tapes philippinarum: a laboratory approach. Aquaculture International, 13(1–2):75–88, January 2005. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-9029-z.

Tramati:2005:SPD

[537] C. Tramati, B. Savona, and A. Mazzola. A study of the pattern of digestive enzymes in Diplodus puntazzo (Cetti, 1777) (Osteichthyes, Sparidae): evidence for the definition of nutritional protocols. Aquaculture International, 13(1–2):89–95, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-004-9028-0.

Porrello:2005:LNL

[538] S. Porrello, M. Lenzi, G. Ferrari, E. Persia, F. Savelli, and P. Tomassetti. Loading of nutrient from a land-based fish farm (Orbetello, Italy) at different times. Aquaculture International, 13(1–2):97–108, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-9027-1.

Tomassetti:2005:PIM

[539] P. Tomassetti and S. Porrello. Polychaetes as indicators of marine fish farm organic enrichment. Aquaculture International, 13(1–2):109– 128, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-004-9026-2.

Giangrande:2005:UFF

[540] A. Giangrande, A. Cavallo, M. Licciano, E. Mola, C. Pierri, and L. Trianni. Utilization of the filter feeder polychaete sabella. Aquaculture International, 13(1–2):129–136, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-004-9025-3.

Chessa:2005:PDL

[541] L. A. Chessa, F. Paesanti, A. Pais, M. Scardi, S. Serra, and L. vitale. Perspectives for development of low impact aquaculture in a Western Mediterranean lagoon: the case of the carpet clam Tapes decussatus. Aquaculture International, 13(1–2):147–155, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-004-9022-6.

Vizzini:2005:ASC

[542] S. Vizzini, B. Savona, M. Caruso, A. Savona, and A. Mazzola. Analysis of stable carbon and nitrogen isotopes as a tool for assessing the environmental impact of aquaculture: a case study from the western Mediterranean. Aquaculture International, 13(1–2):157–165, January 2005. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-9023-5.

Zaccone:2005:MIA

[543] R. Zaccone, M. Mancuso, A. Modica, and D. Zampino. Microbiological indicators for aquaculture impact in Mar Piccolo (Taranto, Italy). Aquaculture International, 13(1–2):167–173, January 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-004-9038-y.

Woods:2005:EVA

[544] Chris M. C. Woods. Evaluation of VI-alpha and PIT-tagging of the seahorse Hippocampus abdominalis. Aquaculture International, 13(3): 175–186, May 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-004-1087-8.

Davis:2005:OFF

[545] Jerome A. Davis, Mathieu Wille, Thomas Hecht, and Patrick Sorgeloos. Optimal first feed organism for South African mud crab Scylla serrata (Forskål) larvae. Aquaculture International, 13(3):187–201, May 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-1496-8.

Davis:2005:OTW

[546] J. A. Davis, M. Wille, T. Hecht, and P. Sorgeloos. Optimum time for weaning South African Scylla serrata (Forskål) larvae from rotifers to Artemia. Aquaculture International, 13(3):203–216, May 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-004-1915-x.

Roraa:2005:IHC

[547] Anna Maria Bencze Rørå, Bente Ruyter, Jon Skorve, Rolf K. Berge, and Karl-Erik Slinning. Influence of high content of dietary soybean oil on quality of large fresh, smoked and frozen Atlantic salmon (Salmo salar). Aquaculture International, 13(3):217–231, May 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-004-1074-0.

Chaitanawisuti:2005:GHR

[548] N. Chaitanawisuti, S. Kritsanapuntu, and S. Y. Natsukari. Growout of hatchery-reared juvenile spotted babylon (Babylonia areolate link 1807) to marketable size at four stocking densities in flow-through and recirculating seawater systems. Aquaculture International, 13(3):233– 239, May 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-004-1585-8.

Barman:2005:EIW

[549] Utpal Kumar Barman, S. N. Jana, S. K. Garg, Anita Bhatnagar, and A. R. T. Arasu. Effect of inland water salinity on growth, feed conversion efficiency and intestinal enzyme activity in growing grey mullet, Mugil cephalus (Linn.): Field and laboratory studies. Aquaculture International, 13(3):241–256, May 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-004-2479-5.

Luo:2005:EDL

[550] Zhi Luo, Yong-Jian Liu, Kang-Sen Mai, Li-Xia Tian, Dong-Hui Liu, Xiao-Ying Tan, and Hei-Zhao Lin. Effect of dietary lipid level on growth performance, feed utilization and body composition of grouper Epinephelus coioides juveniles fed isonitrogenous diets in floating netcages. Aquaculture International, 13(3):257–269, May 2005. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-2478-6.

Blanc:2005:FEB

[551] J. M. Blanc and P. Maunas. Farming evaluation of the brownbow triploid hybrid (Oncorhynchus mykiss × Salmo trutta). Aquaculture International, 13(3):271–281, May 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-004-1086-9. See erratum [574].

Goddard:2005:BRS

[552] S. Goddard. Book reviews: Shrimp Farming and Mangrove Loss in Thailand. Edited by E. B. Barbier and S. Sathirathai. Edward Elgar Publishing, Cheltenham, UK, 2004, +288 pages, ISBN 1-84376-601-9 (hbk) (Price £69.95). Aquaculture International, 13(3):283–285, May 2005. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-1170-1.

Peruzzi:2005:BRA

[553] Stefano Peruzzi. Book reviews: Aquaculture and Fisheries Biotechnology: Genetic Approaches. By R. A. Dunham. Wallingford, UK: CABI Publishing, 2004, xi + 372 pages, ISBN 0-85199-596-9 (Price £75, US$ 140). Aquaculture International, 13(3):285–288, May 2005. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-004-1314-3.

Jobling:2005:BRE

[554] Malcolm Jobling. Book reviews: Effects of Pollution on Fish: Molecular Effects and Population Responses. By A. J. Lawrence and K. L. Hemingway (eds). Oxford: Blackwell Scienti.c, 2003, xvii + 342 pages, ISBN 0-632-06406-4 (hbk) (Price £79.50). Aquaculture International, 13(3):288–290, May 2005. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-004-1498-6.

Luzzana:2005:EDP

[555] Umberto Luzzana, Franco Valfrè, Marco Mangiarotti, Cinzia Domeneghini, Giuseppe Radaelli, Vittorio Maria Moretti, and Marco Scolari. Evaluation of different protein sources in fingerling grey mullet Mugil cephalus practical diets. Aquaculture International, 13(4):291–303, July 2005. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-3099-9.

Woods:2005:GCS

[556] Chris M. C. Woods. Growth of cultured seahorses (Hippocampus abdominalis) in relation to feed ration. Aquaculture International, 13 (4):305–314, July 2005. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-004-3100-7.

Suarez:2005:CMC

[557] M. D. Suárez, M. Abad, T. Ruiz-Cara, J. D. Estrada, and M. García-Gallego. Changes in muscle collagen content during post mortem storage of farmed sea bream (Sparus aurata): influence on textural properties. Aquaculture International, 13(4):315–325, July 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-004-3405-6.

Kopp:2005:CGB

[558] J. Kopp, F. Cornette, and C. Simonne. A comparison of growth and biochemical composition of Mytilus galloprovincialis (Lmk.) and Mytilus edulis (L.) on the West coast of Cotentin, Normandy, France 1999–2000. Aquaculture International, 13(4):327–340, July 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-004-6548-6.

Rajendran:2005:WSS

[559] K. V. Rajendran, K. K. Vijayan, T. C. Santiago, and J. J. S. Rajan. White spot syndrome virus (WSSV) infection in tiger shrimp Penaeus monodon: a non-lethal histopathological rapid diagnostic method using paraffin and frozen sections. Aquaculture International, 13(4):341– 349, July 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-004-6549-5.

Dunphy:2005:PIF

[560] B. J. Dunphy, R. M. G. Wells, and A. G. Jeffs. Polydorid infestation in the flat oyster, Tiostrea chilensis: hyposaline treatment for an aquaculture candidate. Aquaculture International, 13(4):351–358, July 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-6550-z.

Buitrago:2005:MPC

[561] Esperanza Buitrago, Cèsar Lodeiros, Karem Lunar, Daniela Alvarado, Federico Indorf, Khenia Frontado, Pulido Moreno, and Zoila Vasquez. Mass production of competent larvae of the sea urchin Lytechinus variegatus (Echinodermata: Echinoidea). Aquaculture International, 13 (4):359–367, July 2005. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-004-6551-y.

Alam:2005:ARD

[562] Md. Shah Alam, Shin ichi Teshima, Dedy Yaniharto, Ophirtus Sumule, Manabu Ishikawa, and Shunsuke Koshio. Assessment of reference dietary amino acid pattern for juvenile red sea bream, Pagrus major. Aquaculture International, 13(4):369–379, July 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-005-0614-6.

Jobling:2005:BRS

[563] Malcolm Jobling. Book review: Sustainable Aquaculture: Global Perspectives. B. B. Jana & C. D. Webster (eds). New York: Haworth Press (Food Products Press), 2003. xiii + 365 pp. ISBN 1-56022-104-6 (Pbk). Price $39.95. Aquaculture International, 13(4):381–382, July 2005. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-004-2879-6.

Brown:2005:BRC

[564] J. A. Brown. Book review: Culture of Cold-water Marine Fish. E. Moksness, E. Kjørsvik and Y. Olsen (eds). Fishing News Books (Blackwell Publishing), Oxford, 2004. 528 pp. ISBN 0-85238-276-6. Price £95 (HB). Aquaculture International, 13(4):383–385, July 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-004-4089-7.

Jobling:2005:BRA

[565] Malcolm Jobling. Book review: Australian Fish Farmer. J. Mosig & R. Fallu. Collingwood, Landlinks Press, 2004. xii + 444 pp. ISBN 0-643-06865-1 (Pbk). Price AU$89.95. Available from Eurospan, UK (email: orders@edspubs.co.uk) or CSIRO Publishing (e-mail: publishing.sales@csiro.au). Aquaculture International, 13(4):387–389, July 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-005-0269-3.

Monteforte:2005:ESD

[566] Mario Monteforte, Horacio Bervera, Juan José Ramírez, Pedro Saucedo, and César O. López. Effect of stocking density on growth and survival of the rainbow pearl oyster Pteria sterna (Gould 1852) during nursery and late culture in Bahía de La Paz, Baja California Sur, México. Aquaculture International, 13(5):391–407, October 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-005-1265-3.

Halling:2005:CSI

[567] C. Halling, G. Aroca, M. Cifuentes, A. H. Buschmann, and M. Troell. Comparison of spore inoculated and vegetative propagated cultivation methods of Gracilaria chilensis in an integrated seaweed and fish cage culture. Aquaculture International, 13(5):409–422, October 2005. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-005-6977-x.

Domingues:2005:EAD

[568] Pedro M. Domingues, Paul F. Dimarco, Jose P. Andrade, and Phillip G. Lee. Effect of artificial diets on growth, survival and condition of adult cuttlefish, Sepia officinalis Linnaeus, 1758. Aquaculture International, 13 (5):423–440, October 2005. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-6978-9.

Hameed:2005:SPP

[569] A. S. Sahul Hameed, V. Parameswaran, S. Syed Musthaq, R. Sudhakaran, G. Balasubramanian, and K. Yoganandhan. A simple PCR procedure to detect White Spot Syndrome Virus (WSSV) of shrimp, Penaeus monodon (Fabricius). Aquaculture International, 13(5):441– 450, October 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-005-7898-4.

Aydin:2005:ATS

[570] Rahmi Aydın, Kenan Köprücü, Mustafa Dörücü, Sibel Şimşek Köprücü, and Murat Pala. Acute toxicity of synthetic pyrethroid cypermethrin on the common carp (Cyprinus carpio L.) embryos and larvae. Aquaculture International, 13(5):451–458, October 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-005-0615-5.

Gui-zhong:2005:SBS

[571] Wang Gui-zhong, Li Shao-jing, Zeng Chao-shu, Lin Shu-jun, Kong Xiang-hui, Ai Chun-xiang, and Lin Qiong-wu. Status of biological studies and aquaculture development of the mud crab, Scylla serrata, in China: an experimental ecological studies. Aquaculture International, 13 (5):459–468, October 2005. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-1887-5.

Villalta:2005:CGS

[572] Mireia Villalta and Alicia Estévez. Culture of Senegal sole larvae without the need for rotifers. Aquaculture International, 13(5):469– 478, October 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-005-6981-1.

Kaminski:2005:RJC

[573] Rafał Kamiński, Michał Korwin-Kossakowski, Jan Kusznierz, Leszek Myszkowski, L. Andrzej Stanny, and Jacek Wolnicki. Response of a juvenile cyprinid, lake minnow Eupallasella perenurus (Pallas), to different diets. Aquaculture International, 13(5):479–486, October 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-005-7899-3.

Blanc:2005:EBJ

[574] J. M. Blanc and P. Maunas. Erratum: Blanc J. M. and Maunas P. 2005. Farming Evaluation of the ‘brownbow’ Triploid Hybrid (Oncorhynchus mykiss × Salmo trutta). Aquacult. Int. 13: 271–281. Aquaculture International, 13(5):487, October 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-005-0330-2. See [551].

Islam:2005:PER

[575] M. S. Islam, A. Milstein, M. A. Wahab, A. H. M. Kamal, and S. Dewan. Production and economic return of shrimp aquaculture in coastal ponds of different sizes and with different management regimes. Aquaculture International, 13(6):489–500, November 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-005-9000-7.

Milstein:2005:CWQ

[576] A. Milstein, M. S. Islam, M. A. Wahab, A. H. M. Kamal, and S. Dewan. Characterization of water quality in shrimp ponds of different sizes and with different management regimes using multivariate statistical analysis. Aquaculture International, 13(6):501–518, November 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-005-9001-6.

Alver:2005:SMB

[577] Morten O. Alver, Jo A. Alfredsen, and Gunvor Øie. A system for modelbased biomass estimation of larvae in intensive cod larvicultures. Aquaculture International, 13(6):519–541, November 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-005-9004-3.

Marsic-Lucic:2005:CEP

[578] Jasna Maršić-Lučić, Ivona Mladineo, and Mladen Tudor. Comparative effectiveness of 2-phenoxyethanol and propiscin as anesthetics for juvenile sea bass Dicentrarchus labrax L. Aquaculture International, 13(6): 543–553, November 2005. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-9005-2.

Kaur:2005:IMG

[579] V. I. Kaur and P. K. Saxena. Incorporation of maize gluten in supplementary feed and its impact on growth and flesh quality of some carps. Aquaculture International, 13(6):555–573, November 2005. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-005-7995-4.

Torkildsen:2005:BAE

[580] Lise Torkildsen, Christophe Lambert, Are Nylund, Thorolf Magnesen, and Øivind Bergh. Bacteria associated with early life stages of the great scallop, Pecten maximus: impact on larval survival. Aquaculture International, 13(6):575–592, November 2005. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-005-9002-5.

Wang:2006:IMG

[581] Jiaxi Wang, Huiji Liu, Wenqiang Min, Jingo Tong, Min Guan, Yuzhang Han, Luojun Gong, Zhen Huang, Jie Ren, Jinping Zhang, and Hongping Zheng. Induced meiotic gynogenesis in tench, Tinca tinca (L.) using irradiated heterogenic sperm. Aquaculture International, 14(1– 2):??, February 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-9012-3.

Kamler:2006:P

[582] Ewa Kamler, Jacek Wolnicki, and Zdzisław Zakeś. Preface. Aquaculture International, 14(1–2):1–2, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-005-9008-z.

Kohlmann:2006:MLT

[583] Klaus Kohlmann and Petra Kersten. Microsatellite loci in tench: isolation and variability in a test population. Aquaculture International, 14 (1–2):3–7, February 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-9009-y.

Linhart:2006:SSD

[584] Otomar Linhart, Marek Rodina, Martin Flajshans, Nicola Mavrodiev, Jana Nebesarova, David Gela, and Martin Kocour. Studies on sperm of diploid and triploid tench, Tinca tinca (L.). Aquaculture International, 14 (1–2):9–25, February 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-9010-5.

Mamcarz:2006:RHT

[585] Andrzej Mamcarz, Dariusz Kucharczyk, and Roman Kujawa. Reciprocal hybrids of tench Tinca tinca (L.) × bream Abramis brama (L.), and tench × carp Cyprinus carpio L., and some characteristics of their early development. Aquaculture International, 14(1–2):27–33, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-005-9011-4.

Piackova:2006:LTE

[586] Veronika Piačková and Martin Flajšhans. Long-term examination of health conditions in monoculture of communally tested amphimictic diploid, diploid gynogenic and triploid tench, Tinca tinca (L.). Aquaculture International, 14(1–2):43–59, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-005-9013-2.

Linhart:2006:IFG

[587] Otomar Linhart, Marek Rodina, Martin Kocour, and David Gela. Insemination, fertilization and gamete management in tench, Tinca tinca (L.). Aquaculture International, 14(1–2):61–73, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-005-9014-1.

Caille:2006:QMF

[588] Nicolas Caille, Marek Rodina, Martin Kocour, David Gela, Martin Flajšhans, and Otomar Linhart. Quantity, motility and fertility of tench Tinca tinca (L.) sperm in relation to LHRH analogue and carp pituitary treatments. Aquaculture International, 14(1–2):75–87, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-005-9015-0.

Wolnicki:2006:EDD

[589] Jacek Wolnicki, Leszek Myszkowski, MichaŁ Korwin-kossakowski, Rafał Kamiński, and L. Andrzej Stanny. Effects of different diets on juvenile tench, Tinca tinca (L.) reared under controlled conditions. Aquaculture International, 14(1–2):89–98, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-005-9017-y.

Kamler:2006:DOA

[590] Ewa Kamler, Leszek Myszkowski, Rafał Kamiński, Michał Korwin-Kossakowski, and Jacek Wolnickil. Does overfeeding affect tench Tinca tinca (L.) juveniles? Aquaculture International, 14(1–2):99–111, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-005-9018-x.

Ostaszewska:2006:MCD

[591] Teresa Ostaszewska, MichaŁ Korwin-Kossakowski, and Jacek Wolnicki. Morphological changes of digestive structures in starved tench Tinca tinca (L.) juveniles. Aquaculture International, 14(1–2):113–126, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-005-9016-z.

Zakes:2006:EFO

[592] ZdzisŁaw Zakeś, Krystyna Demska-Zakeś, PrzemysŁaw Jarocki, and Konrad Stawecki. The effect of feeding on oxygen consumption and ammonia excretion of juvenile tench Tinca tinca (L.) reared in a water recirculating system. Aquaculture International, 14(1–2):127–140, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-005-9019-9.

Witeska:2006:RHR

[593] Małgorzata Witeska, Barbara Jezierska, and Jacek Wolnicki. Respiratory and hematological response of tench, Tinca tinca (L.) to a shortterm cadmium exposure. Aquaculture International, 14(1–2):141–152, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-005-9020-3.

Siwicki:2006:IHM

[594] Andrzej K. Siwicki, Zdzisław Zakeś, John C. Fuller, Steven Nissen, Krzysztof Kazuń, and Edward Głabski. The influence of β-hydroxy- βmethylbutyrate (HMB) on cell-mediated immunity in tench Tinca tinca (L.): in vitro and in vivo study. Aquaculture International, 14(1–2): 153–161, February 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-9021-2.

Gallardo:2006:SBT

[595] José Martín Gallardo, José Carlos Escudero García, Encarnación García Ceballos-Zúñiga, and Juan José Pérez-Regadera Pérez. Selective behaviour of a tench, Tinca tinca (L.), stock in different light-substrate combined conditions. Aquaculture International, 14(1–2):163–170, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-005-9022-1.

Mamcarz:2006:CCE

[596] Andrzej Mamcarz and Andrzej Skrzypczak. Changes in commercially exploited populations of tench, Tinca tinca (L.), in littoral zones of lakes of northeastern Poland. Aquaculture International, 14(1–2):171– 177, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-005-9023-0.

Skrzypczak:2006:CCE

[597] Andrzej Skrzypczak and Andrzej Mamcarz. Changes in commercially exploited populations of tench, Tinca tinca (L.) in lakes of Northeastern Poland. Aquaculture International, 14(1–2):179–193, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-005-9024-z.

Gela:2006:TTT

[598] David Gela, Martin Flajšhans, Martin Kocour, Marek Rodina, and Otomar Linhart. Tench (Tinca tinca) broodstock management in breeding station under conditions of pond culture: a review. Aquaculture International, 14(1–2):195–203, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-005-9025-y.

Wang:2006:TFG

[599] Jiaxi Wang, Wenqiang Min, Min Guan, Luojun Gong, Jie Ren, Zhen Huang, Hongping Zheng, Jinping Zhang, Huiji Liu, and Yuzhang Han. Tench farming in China: present status and future prospects. Aquaculture International, 14(1–2):205–208, February 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-005-9026-x.

FernandezLoFaso:2006:TTT

[600] Raquel P. Fernández Lo Faso, Ramón Velasco Gemio, José C. Escudero García, Encarnación García Ceballos-Zúñiga, Carmen Bueno, and José Martín Gallardo. Tench, Tinca tinca (L.), fish farms as a tool for environmental education. Aquaculture International, 14(1–2): 209–218, February 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-9027-9.

Mahious:2006:EDI

[601] A. S. Mahious, F. J. Gatesoupe, M. Hervi, R. Metailler, and F. Ollevier. Effect of dietary inulin and oligosaccharides as prebiotics for weaning turbot, Psetta maxima (Linnaeus, c. 1758). Aquaculture International, 14(3):219–229, June 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-9003-4.

Milstein:2006:FLZa

[602] Ana Milstein, Arie Valdenberg, and Sheenan Harpaz. Fish larvae — zooplankton relationships in microcosm simulations of earthen nursery ponds. I. Freshwater system. Aquaculture International, 14(3):231– 246, June 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-005-9006-1.

Du:2006:IFR

[603] Zhen-Yu Du, Yong-Jian Liu, Li-Xia Tian, Jian-Guo He, Jun-Ming Cao, and Gui-Ying Liang. The influence of feeding rate on growth, feed efficiency and body composition of juvenile grass carp (Ctenopharyngodon idella). Aquaculture International, 14(3):247–257, June 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-005-9029-7.

Kucukbay:2006:EDC

[604] F. Z. Küçükbay, H. Yazlak, N. Sahin, and M. N. Cakmak. Effects of dietary chromium picolinate supplementation on serum glucose, cholesterol and minerals of rainbow trout (Oncorhynchus mykiss). Aquaculture International, 14(3):259–266, June 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-005-9030-1.

AvendanoDiaz:2006:EIT

[605] Miguel Avendaño Díaz, Marcela Cantillánez Silva, and Juan Peña Forner. Effect of immersion time of cultch on spatfall of the scallop Argopecten purpuratus (Lamarck 1819) in the Marine Reserve at La Riconada, Antofagasta, Chile. Aquaculture International, 14(3):267–283, June 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-005-9033-y.

Martins:2006:GDN

[606] D. Alves Martins, E. Gomes, P. Rema, J. Dias, R. O. A. Ozório, and L. M. P. Valente. Growth, digestibility and nutrient utilization of rainbow trout (Oncorhynchus mykiss) and European sea bass (Dicentrarchus labrax) juveniles fed different dietary soybean oil levels. Aquaculture International, 14(3):285–295, June 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-005-9034-x.

White:2006:EIM

[607] Hannah White and Helen Glenn. Environmental impact mitigation and bi-culture: a comparative legal analysis of flexibility within European legal regimes — biofilter deployment. Aquaculture International, 14 (3):297–317, June 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-9035-9.

Jobling:2006:BRa

[608] Malcolm Jobling. Book review. Aquaculture International, 14(3):319– 321, June 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-005-9028-8.

Jobling:2006:BRb

[609] Malcolm Jobling. Book review. Aquaculture International, 14(3):323– 325, June 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-005-9031-0.

Marinho-Soriano:2006:SAG

[610] E. Marinho-Soriano, W. S. C. Moreira, and M. A. A. Carneiro. Some aspects of the growth of Gracilaria birdiae (Gracilariales, Rhodophyta) in an estuary in Northeast Brazil. Aquaculture International, 14(4): 327–336, August 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-9032-z.

Viaplana-Marin:2006:EPC

[611] I. Viaplana-Marín, J. Fernández-Borrás, and J. Blasco. Effects of the protein/carbohydrate ratio of extruded diets on protein synthesis, protein growth and body composition in juvenile brown trout (Salmo trutta). Aquaculture International, 14(4):337–353, August 2006. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-005-9036-8.

Wong:2006:RCP

[612] Melisa C. Wong and Myriam A. Barbeau. Rock crab predation of juvenile sea scallops: the functional response and its implications for bottom culture. Aquaculture International, 14(4):355–376, August 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-005-9038-6.

Magnesen:2006:YGS

[613] Thorolf Magnesen, Øivind Bergh, and Gyda Christophersen. Yields of great scallop, Pecten maximus, larvae in a commercial flow-through rearing system in Norway. Aquaculture International, 14(4):377–394, August 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-005-9039-5.

Jones:2006:WWT

[614] J. B. Jones. Why won’t they grow? — inhibitory substances and mollusc hatcheries. Aquaculture International, 14(4):395–403, August 2006. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-005-9040-z.

Alfaro:2006:CCP

[615] Andrea C. Alfaro, Brent R. Copp, David R. Appleton, Shane Kelly, and Andrew G. Jeffs. Chemical cues promote settlement in larvae of the green-lipped mussel, Perna canaliculus. Aquaculture International, 14 (4):405–412, August 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-005-9041-y.

Peruzzi:2006:BR

[616] Stefano Peruzzi. Book review. Aquaculture International, 14(4):413– 414, August 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-005-9037-7.

Losinger:2006:FIV

[617] Willard C. Losinger. Factors influencing the variance in expected yield on catfish farms in the United States. Aquaculture International, 14 (5):415–419, October 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-006-9044-3.

Battaglene:2006:TST

[618] Stephen C. Battaglene and David T. Morehead. Tolerance of striped trumpeter Latris lineata embryos to ozonated seawater. Aquaculture International, 14(5):421–429, October 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-006-9045-2.

Milstein:2006:FLZb

[619] Ana Milstein, Arie Valdenberg, and Sheenan Harpaz. Fish larvae: zooplankton relationships in microcosm simulations of earthen nursery ponds. II. Brackish water system. Aquaculture International, 14(5): 431–442, October 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-006-9046-1.

Samuelsen:2006:MDP

[620] Ole Bent Samuelsen. Multiple dose pharmacokinetic study of oxolinic acid in cod, Gadus morhua L. Aquaculture International, 14(5):443– 450, October 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-006-9047-0.

Kostopoulou:2006:CPS

[621] Venetia Kostopoulou, Helen Miliou, George Katis, and George Verriopoulos. Changes in the population structure of the lineage ‘Nevada’ belonging to the Brachionus plicatilis species complex, batch–cultured under different feeding regimes. Aquaculture International, 14(5):451– 466, October 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-006-9048-z.

Nga:2006:EDR

[622] B. T. Nga, R. Roijackers, T. T. Nghia, V. N. Ut, and M. Scheffer. Effects of decomposing Rhizophora apiculata leaves on larvae of the shrimp Penaeus monodon. Aquaculture International, 14(5):467– 477, October 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-006-9049-y.

Jana:2006:EVD

[623] S. N. Jana, S. K. Garg, U. K. Barman, A. R. T. Arasu, and B. C. Patra. Effect of varying dietary protein levels on growth and production of Chanos chanos (Forsskål) in inland saline groundwater: laboratory and field studies. Aquaculture International, 14(5):479–498, October 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-006-9050-5.

Mente:2006:EFF

[624] Eleni Mente, Graham J. Pierce, Maria Begoña Santos, and Christos Neofitou. Effect of feed and feeding in the culture of salmonids on the marine aquatic environment: a synthesis for European aquaculture. Aquaculture International, 14(5):499–522, October 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-006-9051-4.

McLean:2006:BRB

[625] Ewen McLean. Book review: B. A. Costa–Pierce, A. Desbonnet, P. Edwards and D. Baker (eds). Urban Aquaculture. Aquaculture International, 14(5):523–525, October 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-006-9043-4.

Sarkis:2006:LRC

[626] Samia Sarkis, Michael Helm, and Claudia Hohn. Larval rearing of calico scallops, Argopecten gibbus, in a flow-through system. Aquaculture International, 14(6):527–538, December 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-006-9052-3.

Lennard:2006:CTD

[627] Wilson A. Lennard and Brian V. Leonard. A comparison of three different hydroponic sub-systems (gravel bed, floating and nutrient film technique) in an aquaponic test system. Aquaculture International, 14(6): 539–550, December 2006. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-006-9053-2.

Sykes:2006:EUL

[628] António V. Sykes, Pedro M. Domingues, and José P. Andrade. Effects of using live grass shrimp (Palaemonetes varians) as the only source of food for the culture of cuttlefish, Sepia officinalis (Linnaeus, 1758). Aquaculture International, 14(6):551–568, December 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-006-9054-1.

Bolasina:2006:CHR

[629] Sergio Néstor Bolasina. Cortisol and hematological response in Brazilian codling, Urophycis brasiliensis (Pisces, Phycidae) subjected to anesthetic treatment. Aquaculture International, 14(6):569–575, December 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-006-9055-0.

Schulz:2006:GPB

[630] Carsten Schulz, Steffen Günther, Manfred Wirth, and Bernhard Rennert. Growth performance and body composition of pike perch (Sander lucioperca) fed varying formulated and natural diets. Aquaculture International, 14(6):577–586, December 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-006-9056-z.

Kritsanapuntu:2006:EWE

[631] S. Kritsanapuntu, N. Chaitanawisuti, W. Santhaweesuk, and S. Y. Natsukari. Effects of water exchange regimes on growth, survival and shell normality of the hatchery reared juvenile spotted babylon (Babylonia areolata Link 1807) in a recirculating seawater system. Aquaculture International, 14(6):587–594, December 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-006-9057-y.

Dionne:2006:PSM

[632] Mark Dionne, Jean-Sébastien Lauzon-Guay, Diana J. Hamilton, and Myriam A. Barbeau. Protective socking material for cultivated mussels: a potential non-disruptive deterrent to reduce losses to diving ducks. Aquaculture International, 14(6):595–613, December 2006. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-006-9058-x.

Jha:2006:CGS

[633] Prithwiraj Jha, S. Barat, and C. R. Nayak. A comparison of growth, survival rate and number of marketable koi carp produced under different management regimes in earthen ponds and concrete tanks. Aquaculture International, 14(6):615–626, December 2006. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-006-9059-9.

Patil:2007:FAC

[634] Vishwanath Patil, Torsten Källqvist, Elisabeth Olsen, Gjermund Vogt, and Hans R. Gislerød. Fatty acid composition of 12 microalgae for possible use in aquaculture feed. Aquaculture International, 15(1):1– 9, February 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-006-9060-3.

Bergh:2007:SCW

[635] Øivind Bergh and Ole B. Samuelsen. Susceptibility of corkwing wrasse Symphodus melops, goldsinny wrasse Ctenolabrus rupestis, and Atlantic salmon Salmo salar smolt, to experimental challenge with Vibrio tapetis and Vibrio splendidus isolated from corkwing wrasse. Aquaculture International, 15(1):11–18, February 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-006-9061-2.

Garci:2007:XSL

[636] M. E. Garci, J. E. Trigo, S. Pascual, A. F. González, F. Rocha, and A. Guerra. Xenostrobus securis (Lamarck, 1819) (Mollusca: Bivalvia): first report of an introduced species in Galician waters. Aquaculture International, 15(1):19–24, February 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-006-9062-1.

Roycroft:2007:BIS

[637] Daphne Roycroft, Thomas C. Kelly, and Lesley J. Lewis. Behavioural interactions of seabirds with suspended mussel longlines. Aquaculture International, 15(1):25–36, February 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-006-9065-y.

Celada:2007:EBR

[638] J. D. Celada, J. I. Antolín, J. M. Carral, J. R. Pérez, and M. Sáez-Royuela. Effects of breeder reuse on the reproductive potential of the signal crayfish (Pacifastacus leniusculus Dana. Astacidae) in culture. Aquaculture International, 15(1):37–42, February 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-006-9066-x.

Cooper:2007:BVM

[639] Marie Cooper and Kjell Ø. Midling. Blood vessel melanosis: a physiological detoxification mechanism in Atlantic cod (Gadus morhua). Aquaculture International, 15(1):43–54, February 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-006-9067-9.

Erler:2007:NDS

[640] Dirk Erler, Putth Songsangjinda, Teeyaporn Keawtawee, and Kanit Chaiyakam. Nitrogen dynamics in the settlement ponds of a small-scale recirculating shrimp farm (Penaeus monodon) in rural Thailand. Aquaculture International, 15(1):55–66, February 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-006-9068-8.

Szkudlarek:2007:ESD

[641] Maciej Szkudlarek and Zdzisław Zakeś. Effect of stocking density on survival and growth performance of pikeperch, Sander lucioperca (L.), larvae under controlled conditions. Aquaculture International, 15(1): 67–81, February 2007. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-006-9069-7.

Bodis:2007:EDD

[642] M. Bódis, B. Kucska, and M. Bercsényi. The effect of different diets on the growth and mortality of juvenile pikeperch (Sander lucioperca) in the transition from live food to formulated feed. Aquaculture International, 15(1):83–90, February 2007. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-006-9063-0.

Gandra:2007:PGU

[643] André L. Gandra, Daniel R. Ituassú, Manoel Pereira-Filho, Rodrigo Roubach, Roger Crescêncio, and Bruno A. S. Cavero. Pirarucu growth under different feeding regimes. Aquaculture International, 15(1):91– 96, February 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-006-9064-z.

Rasmussen:2007:FCG

[644] Richard Skøtt Rasmussen, Flemming Hausgaard Larsen, and Stig Jensen. Fin condition and growth among rainbow trout reared at different sizes, densities and feeding frequencies in high-temperature re-circulated water. Aquaculture International, 15(2):97–107, April 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-006-9070-1.

Cek:2007:MCC

[645] Şehriban Çek, Funda Turan, and Esin Atik. Masculinization of convict cichlid (Cichlasoma nigrofasciatum) by immersion in Tribulus terrestris extract. Aquaculture International, 15(2):109–119, April 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-006-9071-0.

Verbeke:2007:CPV

[646] Wim Verbeke, Isabelle Sioen, Karen Brunsø, Stefaan De Henauw, and John Van Camp. Consumer perception versus scientific evidence of farmed and wild fish: exploratory insights from Belgium. Aquaculture International, 15(2):121–136, April 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-007-9072-7.

Immanuel:2007:DHP

[647] G. Immanuel, T. Citarasu, V. Sivaram, M. Michael Babu, and A. Palavesam. Delivery of HUFA, probionts and biomedicine through bioencapsulated Artemia as a means to enhance the growth and survival and reduce the pathogenesity in shrimp Penaeus monodon postlarvae. Aquaculture International, 15(2):137–152, April 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9074-5.

Staykov:2007:EMO

[648] Y. Staykov, P. Spring, S. Denev, and J. Sweetman. Effect of a mannan oligosaccharide on the growth performance and immune status of rainbow trout (Oncorhynchus mykiss). Aquaculture International, 15 (2):153–161, April 2007. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9096-z.

Barreto:2007:MDI

[649] R. E. Barreto, A. M. M. C. Gontijo, R. O. Alves de Lima, V. C. Raymundi, D. Pinhal, V. A. V. Reyes, G. L. Volpato, and D. M. F. Salvadori. MS222 does not induce primary DNA damage in fish. Aquaculture International, 15(2):163–168, April 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-007-9073-6.

Jobling:2007:BRC

[650] Malcolm Jobling. Book review: C. Lim, C. D. Webster (eds), Tilapia — Biology, Culture and Nutrition. Aquaculture International, 15(2): 169–170, April 2007. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9095-0.

Adamek:2007:P

[651] Zdeněk Adámek. Preface. Aquaculture International, 15(3–4):171– 172, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-007-9075-4.

Gal:2007:EOC

[652] Dénes Gál, Ferenc Pekár, Éva Kerepeczki, and László Váradi. Experiments on the operation of a combined aquaculture-algae system. Aquaculture International, 15(3–4):173–180, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9089-y.

Terziyski:2007:EOF

[653] Doychin Terziyski, George Grozev, Roumen Kalchev, and Angelina Stoeva. Effect of organic fertilizer on plankton primary productivity in fish ponds. Aquaculture International, 15(3–4):181–190, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9086-1.

Ritterbusch:2007:GPR

[654] David Ritterbusch. Growth patterns of reed (Phragmites australis): the development of reed stands in carp ponds. Aquaculture International, 15 (3–4):191–199, June 2007. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9091-4.

Potuzak:2007:CFP

[655] Jan Potužák, Jan Hůda, and Libor Pechar. Changes in fish production effectivity in eutrophic fishponds — impact of zooplankton structure. Aquaculture International, 15(3–4):201–210, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9085-2.

Adamek:2007:CAI

[656] Zdeněk Adámek, Jiří Kortan, and Martin Flajšhans. Computer-assisted image analysis in the evaluation of fish wounding by cormorant [ Phalacrocorax carbo sinensis (L.)] attacks. Aquaculture International, 15 (3–4):211–216, June 2007. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9087-0.

Musil:2007:SDF

[657] Jiří Musil, Zdenek Adámek, and Christian Baranyi. Seasonal dynamics of fish assemblage in a pond canal. Aquaculture International, 15 (3–4):217–226, June 2007. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9092-3.

Bekefi:2007:MPF

[658] Emese Bekefi and Laszlo Varadi. Multifunctional pond fish farms in Hungary. Aquaculture International, 15(3–4):227–233, June 2007. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9090-5.

Hubenova:2007:IFF

[659] Tania Hubenova, Angel Zaikov, and Penka Vasileva. Investigation on fecundity, follicles and free embryo size of pond-reared pike (Esox lucius) of different age and size. Aquaculture International, 15(3–4):235– 240, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-007-9077-2.

Fullner:2007:EPH

[660] Gert Füllner, Tim Gottschalk, and Matthias Pfeifer. Experiments for the production of hybrid striped bass in in-pond circulation systems. Aquaculture International, 15(3–4):241–248, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9093-2.

Hubenova:2007:MPF

[661] Tania Hubenova, Angel Zaikov, and Penka Vasileva. Management of paddlefish fry and juveniles in Bulgarian conditions. Aquaculture International, 15(3–4):249–253, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-007-9082-5.

Hamackova:2007:SNI

[662] Jitka Hamáčková, Andrea Lepičová, Miroslav Prokeš, Pavel Lepič, Pavel Kozák, Tomáš Policar, and Leon Andrzej Stanny. Success of nursing ide (Leuciscus idus, L.) fry related to the period of feeding with live food. Aquaculture International, 15(3–4):255–265, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9083-4.

Machova:2007:TDE

[663] J. Máchová, M. Prokeš, Z. Svobodová, V. Žlábek, M. Peňáz, and V. Baruš. Toxicity of Diazinon 60 EC for Cyprinus carpio and Poecilia reticulata. Aquaculture International, 15(3–4):267–276, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9078-1.

Svobodova:2007:AAC

[664] Zdeňka Svobodová, Jana Máchová, Hana Kroupová, Miriam Smutná, and Ladislav Groch. Ammonia autointoxication of common carp: case studies. Aquaculture International, 15(3–4):277–286, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9079-0.

Vallod:2007:AFR

[665] D. Vallod, J. P. Cravedi, A. Hillenweck, and J. Robin. Analysis of the off-flavor risk in carp production in ponds in Dombes and Forez (France). Aquaculture International, 15(3–4):287–298, June 2007. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9080-7.

Marsalek:2007:CTM

[666] Petr Maršálek, Zdeňka Svobodová, and Tomáš Randák. The content of total mercury and methylmercury in common carp from selected Czech ponds. Aquaculture International, 15(3–4):299–304, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9076-3.

Pickova:2007:IFA

[667] Jana Pickova, Eva Brännäs, and Torleif Andersson. Importance of fatty acids in broodstock diets with emphasis on Arctic char (Salvelinus alpinus) eggs. Aquaculture International, 15(3–4):305–311, June 2007. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9084-3.

Steffens:2007:INL

[668] Werner Steffens and Manfred Wirth. Influence of nutrition on the lipid quality of pond fish: common carp (Cyprinus carpio) and tench (Tinca tinca). Aquaculture International, 15(3–4):313–319, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9088-z.

Vacha:2007:ISC

[669] F. Vacha, P. Vejsada, J. Huda, and P. Hartvich. Influence of supplemental cereal feeding on the content and structure of fatty acids during long-lasting storage of common carp (Cyprinus carpio L.). Aquaculture International, 15(3–4):321–329, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9081-6.

Szathmari:2007:IDM

[670] Lászaló Szathmári and Eszter Molnár. Investigations of dry matter and fat content in carp species smoked by hot and cold methods. Aquaculture International, 15(3–4):331–336, June 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9094-1.

Ahmed:2007:DAA

[671] Imtiaz Ahmed. Dietary amino acid l-threonine requirement of fingerling Indian catfish, Heteropneustes fossilis (Bloch) estimated by growth and biochemical parameters. Aquaculture International, 15(5):337–350, October 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-007-9097-y.

Song:2007:ESR

[672] Liang Song, Xiaoxu Li, Steven Clarke, Ting Wang, and Kriston Bott. The effect of size on the response of Pacific oysters (Crassostrea gigas) to changes in water temperature and air exposure. Aquaculture International, 15(5):351–362, October 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-007-9098-x.

Alam:2007:ERD

[673] S. M. Nazmul Alam, Bob Pokrant, Amararatne Yakupitiyage, and Michael J. Phillips. Economic returns of disease-affected extensive shrimp farming in southwest Bangladesh. Aquaculture International, 15(5): 363–370, October 2007. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9100-7.

Xu:2007:EUV

[674] Jian-He Xu, Feng You, Bin-Lun Yan, and Pei-Jun Zhang. Effects of ultra-violet irradiation on sperm motility and diploid gynogenesis induction in large yellow croaker (Pseudosciaena crocea) undergoing cold shock. Aquaculture International, 15(5):371–382, October 2007. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9101-6.

Karlsson:2007:TMS

[675] Sten Karlsson, Liang Ma, Eric Saillant, and John Rush Gold. Tests of Mendelian segregation and linkage-group relationships among 31 microsatellite loci in red drum, Sciaenops ocellatus. Aquaculture International, 15(5):383–391, October 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-007-9102-5.

Lanszki:2007:DFC

[676] József Lanszki, Zsuzsanna S. Pallos, Dénes Nagy, and Grace Yoxon. Diet and fish choice of Eurasian otters (Lutra lutra L.) in fish wintering ponds in Hungary. Aquaculture International, 15(5):393–402, October 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9103-4.

Imsland:2007:IGJ

[677] Albert K. Imsland, Edward Schram, Bjorn Roth, Rian Schelvis-Smit, and Kees Kloet. Improving growth in juvenile turbot (Scophthalmus maximus Rafinesque) by rearing fish in switched temperature regimes. Aquaculture International, 15(5):403–407, October 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9099-9.

Jobling:2007:BRR

[678] Malcolm Jobling. Book review: R. C. Ray (ed), Microbial Biotechnology in Agriculture and Aquaculture, Volume II, Science Publishers, Enfield, 2006, Hardback, 569 pp, £60.20, ISBN-10: 1-57808-443-1. Aquaculture International, 15(5):409–410, October 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9104-3.

Jobling:2007:BRP

[679] Malcolm Jobling. Book review: P. S. Leung and C. Engle (eds): Shrimp culture — Economics, market and trade. Aquaculture International, 15 (5):411–412, October 2007. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9107-0.

Shumway:2007:BRD

[680] Sandra E. Shumway. Book review: D. Hardy, Scallop Farming (2nd edition). Aquaculture International, 15(5):413–414, October 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9108-z.

Dellatorre:2007:FPG

[681] Fernando G. Dellatorre, Marcela S. Pascual, and Pedro J. Barón. Feeding physiology of the Argentine mussel Mytilus edulis platensis (d’Orbigny, 1846): does it feed faster in suspended culture systems? Aquaculture International, 15(6):415–424, December 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9105-2.

Hossain:2007:MCE

[682] M. Shahadat Hossain, Sayedur Rahman Chowdhury, Nani Gopal Das, and M. Moshiur Rahaman. Multi-criteria evaluation approach to GISbased land-suitability classification for tilapia farming in Bangladesh. Aquaculture International, 15(6):425–443, December 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9109-y.

Kaliba:2007:PEA

[683] Aloyce R. Kaliba, Charles C. Ngugi, John M. Mackambo, Kajitanus O. Osewe, Ephraim Senkondo, Berno V. Mnembuka, and Steven Amisah. Potential effect of aquaculture promotion on poverty reduction in sub– saharan Africa. Aquaculture International, 15(6):445–459, December 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-007-9110-5.

Celada:2007:ESD

[684] Jesus D. Celada, Amelia Aguilera, Jose M. Carral, María Sáez-Royuela, Pedro M. Melendre, and Jose R. Pérez. Effects of stocking density on survival and growth of juvenile tench (Tinca tinca L.). Aquaculture International, 15(6):461–465, December 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-007-9111-4.

Wu:2007:EDD

[685] Xiao-Yi Wu, Yong-Jian Liu, Li-Xia Tian, Kang-Sen Mai, Ran Guo, and Sheng-Jie Jin. Effect of different dietary raw to pre-gelatinized starch ratios on growth performance, feed utilization and body composition of juvenile yellowfin seabream (Sparus latus). Aquaculture International, 15 (6):467–477, December 2007. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9113-2.

You:2007:RJR

[686] Kui You, Caihua Ma, Huiwang Gao, Fengqi Li, Meizhao Zhang, Yantao Qiu, and Bo Wang. Research on the jellyfish (Rhopilema esculentum Kishinouye) and associated aquaculture techniques in China: current status. Aquaculture International, 15(6):479–488, December 2007. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9114-1.

Celada:2007:TTT

[687] J. D. Celada, J. M. Carral, R. Rodríguez, M. Sáez-Royuela, A. Aguilera, P. Melendre, and J. Martín. Tench (Tinca tinca L.) larvae rearing under controlled conditions: density and basic supply of Artemia nauplii as the sole food. Aquaculture International, 15(6):489–495, December 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9116-z.

Jobling:2007:BRM

[688] Malcolm Jobling. Book review: M. Fingerman & R. Nagabhushanam (eds): Biomaterials from Aquatic and Terrestrial Organisms. Aquaculture International, 15(6):497–498, December 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9106-1.

Otton:2007:BRC

[689] David Otton. Book review: C. R. Engle & K. Quagrainie, Aquaculture Marketing Handbook. Aquaculture International, 15(6):499–500, December 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-007-9112-3.

Jobling:2007:BRJ

[690] Malcolm Jobling. Book review: J. B. Luten, C. Jacobsen, K. Bekaert, A. Sæbø and J. Oehlenschläger (eds), Seafood Research from Fish to Dish — Quality, Safety and Processing of Wild and Farmed Fish. Aquaculture International, 15(6):501–502, December 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9115-0.

Jobling:2007:BRL

[691] Malcolm Jobling. Book review: O.-I. Lekang, Aquaculture Engineering. Aquaculture International, 15(6):503–504, December 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9119-9.

Jobling:2007:BRV

[692] Malcolm Jobling. Book review: T. V. R. Pillay and M. N. Kutty, Aquaculture: Principles and Practices (2nd edition). Aquaculture International, 15(6):505–506, December 2007. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-007-9122-1.

Avendano-Herrera:2008:EDR

[693] Ruben Avendaño-Herrera, Soledad Núñez, Juan L. Barja, and Alicia E. Toranzo. Evolution of drug resistance and minimum inhibitory concentration to enrofloxacin in Tenacibaculum maritimum strains isolated in fish farms. Aquaculture International, 16(1):1–11, February 2008. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9117-y.

Papadakis:2008:WGA

[694] Ioannis E. Papadakis, Stavros Chatzifotis, Pascal Divanach, and Maroudio Kentouri. Weaning of greater amberjack (Seriola dumerilii Risso 1810) juveniles from moist to dry pellet. Aquaculture International, 16(1):13–25, February 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9118-x.

Alam:2008:DSG

[695] Mohammad Ashraful Alam and Masaru Nakamura. Determination of sex and gonadal maturity in the honeycomb grouper, Epinephelus merra, through biopsy. Aquaculture International, 16(1):27–32, February 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9120-3.

Johnston:2008:EPR

[696] Matthew D. Johnston, Danielle J. Johnston, and Clive M. Jones. Evaluation of partial replacement of live and fresh feeds with a formulated diet and the influence of weaning Panulirus ornatus phyllosomata onto a formulated diet during early ontogeny. Aquaculture International, 16 (1):33–47, February 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9121-2.

Woods:2008:CEU

[697] Chris M. C. Woods, Philip J. James, Graeme A. Moss, Johnny Wright, and Sten Siikavuopio. A comparison of the effect of urchin size and diet on gonad yield and quality in the sea urchin Evechinus chloroticus Valenciennes. Aquaculture International, 16(1):49–68, February 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9124-z.

Herrera:2008:FRS

[698] Marcelino Herrera, Ismael Hachero, Montserrat Rosano, José Francisco Ferrer, José Manuel Márquez, and José Ignacio Navas. First results on spawning, larval rearing and growth of the wedge sole (Dicologoglossa cuneata) in captivity, a candidate species for aquaculture. Aquaculture International, 16(1):69–84, February 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9126-x.

Alvarado:2008:EEC

[699] Daniela Alvarado, Esperanza Buitrago, María Solé, and Khenia Frontado. Experimental evaluation of a composted seaweed extract as microalgal culture media. Aquaculture International, 16(1):85–90, February 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-007-9125-y.

Jobling:2008:BRD

[700] M. Jobling. Book review: A. D. Scarfe, C.-S. Lee & P. J. O’Bryen (eds): Aquaculture Biosecurity: Prevention, Control and Eradication of Aquatic Animal Disease. Aquaculture International, 16(1):91–92, February 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9123-0.

Heinsbroek:2008:GME

[701] Leon T. N. Heinsbroek, Bart J. Goedegebuur, Gertrude Bloemhof, Robert B. Flach, and Gert D. C. de Jong. Gastrointestinal and metabolic effects of feeding schedule on voluntary feed intake and growth of European eel, Anguilla anguilla. Aquaculture International, 16(2):93– 108, April 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-007-9128-8.

Saoud:2008:ESD

[702] I. Patrick Saoud, Joly Ghanawi, and Nada Lebbos. Effects of stocking density on the survival, growth, size variation and condition index of juvenile rabbitfish Siganus rivulatus. Aquaculture International, 16 (2):109–116, April 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9129-7.

Kop:2008:ESN

[703] Aysun Kop and Yaşar Durmaz. The effect of synthetic and natural pigments on the colour of the cichlids (Cichlasoma severum sp., Heckel 1840). Aquaculture International, 16(2):117–122, April 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9130-1.

Seginer:2008:DFD

[704] Ido Seginer. A dynamic fish digestion–assimilation model: oxygen consumption and ammonia excretion in response to feeding. Aquaculture International, 16(2):123–142, April 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-007-9131-0.

Han:2008:ETE

[705] Kyung Nam Han, Soon Woo Lee, and Soon Young Wang. The effect of temperature on the energy budget of the Manila clam, Ruditapes philippinarum. Aquaculture International, 16(2):143–152, April 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9133-y.

Bunting:2008:HIA

[706] Stuart W. Bunting. Horizontally integrated aquaculture development: Exploring consensus on constraints and opportunities with a stakeholder Delphi. Aquaculture International, 16(2):153–169, April 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9134-x.

Christophersen:2008:PTR

[707] Gyda Christophersen, Guillermo Román, Jerry Gallagher, and Thorolf Magnesen. Post-transport recovery of cultured scallop (Pecten maximus) spat, juveniles and adults. Aquaculture International, 16(2):171– 185, April 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-007-9135-9.

Jobling:2008:BRB

[708] Malcolm Jobling. Book review: B. Aarset and G. Rusten (eds): Havbruk: Akvakultur på norsk. Aquaculture International, 16(2):187– 188, April 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-007-9127-9.

Jobling:2008:BRP

[709] Malcolm Jobling. Book review: P. S. Leung, C.-S. Lee & P. J. O’Bryan (eds): Species and system selection for sustainable aquaculture. Aquaculture International, 16(2):189–190, April 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9132-z.

Park:2008:GPD

[710] Jeonghwan Park, Pyong-Kih Kim, and Jae-Yoon Jo. Growth performance of disk abalone Haliotis discus hannai in pilot- and commercialscale recirculating aquaculture systems. Aquaculture International, 16 (3):191–202, June 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9136-8.

You:2008:GSD

[711] Cuihong You, Xiaomu Yu, Deqing Tan, and Jingou Tong. Gynogenesis and sex determination in large-scale loach Paramisgurnus dabryanus (Sauvage). Aquaculture International, 16(3):203–214, June 2008. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9138-6.

Domingues:2008:GAA

[712] Pedro Domingues, Ana Ferreira, Lorenzo Marquez, José P. Andrade, Nelda López, and Carlos Rosas. Growth, absorption and assimilation efficiency by mature cuttlefish (Sepia officinalis) fed with alternative and artificial diets. Aquaculture International, 16(3):215–229, June 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9139-5.

Asha:2008:RBC

[713] P. S. Asha and P. Muthiah. Reproductive biology of the commercial sea cucumber Holothuria spinifera (Echinodermata: Holothuroidea) from Tuticorin, Tamil Nadu, India. Aquaculture International, 16(3):

231–242, June 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-007-9140-z.

Turchini:2008:BEE

[714] Giovanni M. Turchini and Sena S. De Silva. Bio-economical and ethical impacts of alien finfish culture in European inland waters. Aquaculture International, 16(3):243–272, June 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-007-9141-y.

LeFrancois:2008:SCA

[715] N. R. Le François, S. G. Lamarre, H. Tveiten, P. U. Blier, and J. Bailey. Sperm cryoconservation in Anarhichas sp., endangered cold-water aquaculture species with internal fertilization. Aquaculture International, 16(3):273–279, June 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9137-7.

Jobling:2008:BRM

[716] Malcolm Jobling. Book review: M. King, Fisheries biology, assessment and management, 2nd edn. Aquaculture International, 16(3):281– 282, June 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-007-9148-4.

Edding:2008:CCC

[717] Mario Edding, Erika Fonck, Roberto Acuña, and Fadia Tala. Cultivation of Chondrus canaliculatus (C. Agardh) Greville (Gigartinales, Rhodophyta) in controlled environments. Aquaculture International, 16 (4):283–295, August 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9142-x.

Pronker:2008:HBC

[718] Anna Elisabeth Pronker, Nancy Marie Nevejan, Frank Peene, Pieter Geijsen, and Patrick Sorgeloos. Hatchery broodstock conditioning of the blue mussel Mytilus edulis (Linnaeus 1758). Part I. Impact of different microalgae mixtures on broodstock performance. Aquaculture International, 16(4):297–307, August 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9143-9.

Hashimoto:2008:ETP

[719] Tomoki Hashimoto, Katsuya Hyodoh, Takuma Hirose, Satoshi Nishikawa, Toshiya Katano, and Shin ichi Nakano. Evaluation of three phytoplankton species as food for the pearl oyster Pinctada fucata. Aquaculture International, 16(4):309–318, August 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-007-9144-8.

Kang:2008:CPR

[720] Bin Kang and Weiwei Xian. C, N and P regeneration by a detritivorous fish, Liza haematocheila T. & S.: effects of temperature, diet and body size. Aquaculture International, 16(4):319–331, August 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-007-9146-6.

Ambrosio:2008:SDI

[721] Pedro Pablo Ambrosio, Corrado Costa, Pablo Sánchez, and Rosa Flos. Stocking density and its influence on shape of Senegalese sole adults. Aquaculture International, 16(4):333–343, August 2008. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9147-5.

Mazlum:2008:CSG

[722] Yavuz Mazlum and Arnold G. Eversole. Comparison of the survival, growth and yield of red swamp crayfish and white river crayfish in monoculture and polyculture systems. Aquaculture International, 16 (4):345–350, August 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9149-3.

Harlioglu:2008:HFC

[723] Muzaffer Mustafa Harlioğlu. The harvest of the freshwater crayfish Astacus leptodactylus Eschscholtz in Turkey: harvest history, impact of crayfish plague, and present distribution of harvested populations. Aquaculture International, 16(4):351–360, August 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9145-7.

Wang:2008:EDM

[724] Youji Wang, Menghong Hu, Ling Cao, Yi Yang, and Weimin Wang. Effects of daphnia (Moina micrura) plus chlorella (Chlorella pyrenoidosa) or microparticle diets on growth and survival of larval loach (Misgurnus anguillicaudatus). Aquaculture International, 16(4):361– 368, August 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-007-9150-x.

Peruzzi:2008:BRZ

[725] Stefano Peruzzi. Book review: Zhanjiang (John) Liu (ed): Aquaculture Genome Technologies. Aquaculture International, 16(4):369– 371, August 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-007-9157-3.

Imsland:2008:BRR

[726] Albert K. Imsland. Book review: R. N. Gibson (ed), Flatfishes: Biology and Exploitation. Aquaculture International, 16(4):373–374, August 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-008-9166-x.

Jobling:2008:BRJ

[727] Malcolm Jobling. Book review: E. J. Branson (ed): Fish welfare. Aquaculture International, 16(4):375–376, August 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-008-9179-5.

Avendano:2008:EWD

[728] Miguel Avendaño, Marcela Cantillánez, and Gérard Thouzeau. Effects of water depth on survival and growth of Argopecten purpuratus (Lamarck, 1819) spat in northern Chile. Aquaculture International, 16 (5):377–391, October 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9151-9.

Ghosh:2008:BGW

[729] Shubhadeep Ghosh, Archana Sinha, and Chittaranjan Sahu. Bioaugmentation in the growth and water quality of livebearing ornamental fishes. Aquaculture International, 16(5):393–403, October 2008. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9152-8.

Ayub:2008:PWS

[730] Farhana Ayub, Md. Yusuf Sarker, and Md. Samsul Alam. Prevalence of white spot syndrome virus infection detected by one-step and nested PCR in selected tiger shrimp (Penaeus monodon) hatcheries. Aquaculture International, 16(5):405–415, October 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9153-7.

Alaminos:2008:EDN

[731] J. Alaminos and P. Domingues. Effects of different natural or prepared diets on growth and survival of juvenile spider crabs, Maja brachydactyla (Balss, 1922). Aquaculture International, 16(5):417–425, October 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9154-6.

Palma:2008:EBT

[732] Jorge Palma, Dominique P. Bureau, and José Pedro Andrade. Effects of binder type and binder addition on the growth of juvenile Palaemonetes varians and Palaemon elegans (Crustacea: Palaemonidae). Aquaculture International, 16(5):427–436, October 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9155-5.

Supannapong:2008:DEV

[733] Pipadchaya Supannapong, Teerayut Pimsalee, Teerasak A-komol, Arunee Engkagul, Uthaiwan Kovitvadhi, Satit Kovitvadhi, and Krisna Rungruangsak-Torrissen. Digestive enzymes and in-vitro digestibility of different species of phytoplankton for culture of the freshwater pearl mussel, Hyriopsis (Hyriopsis) bialatus. Aquaculture International, 16 (5):437–453, October 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9156-4.

Wei:2008:CGG

[734] Liu-Zhi Wei, Xiu-Mei Zhang, Jian Li, and Guo-Qiang Huang. Compensatory growth of Chinese shrimp, Fenneropenaeus chinensis following hypoxic exposure. Aquaculture International, 16(5):455–470, October 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9158-2.

Yu:2008:EDB

[735] Ming-Chao Yu, Zhuo-Jia Li, Hei-Zhao Lin, Guo-Liang Wen, and Shen Ma. Effects of dietary Bacillus and medicinal herbs on the growth, digestive enzyme activity, and serum biochemical parameters of the shrimp Litopenaeus vannamei. Aquaculture International, 16(5):471– 480, October 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-007-9159-1.

Jobling:2008:BRQ

[736] Malcolm Jobling. Book review: Q. Bone, R. Moore, Biology of Fishes, 3rd edn. Aquaculture International, 16(5):481–482, October 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-008-9186-6.

Nevejan:2008:HBC

[737] Nancy Marie Nevejan, Anna Elisabeth Pronker, and Frank Peene. Hatchery broodstock conditioning of the blue mussel Mytilus edulis (Linnaeus, 1758). Part II. New formulated feeds offer new perspectives to commercial hatcheries. Aquaculture International, 16(6):483–495, December 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-007-9160-8.

Guo:2008:MMI

[738] Wei Guo and Jian-Fang Gui. Microsatellite marker isolation and cultured strain identification in Carassius auratus gibelio. Aquaculture International, 16(6):497–510, December 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-007-9161-7.

Punitha:2008:IIH

[739] S. M. J. Punitha, M. M. Babu, V. Sivaram, V. S. Shankar, S. A. Dhas, T. C. Mahesh, G. Immanuel, and T. Citarasu. Immunostimulating influence of herbal biomedicines on nonspecific immunity in grouper Epinephelus tauvina juvenile against Vibrio harveyi infection. Aquaculture International, 16(6):511–523, December 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-007-9162-6.

Moctezuma-Malagon:2008:MES

[740] A. Moctezuma-Malagón, C. E. González-Esquivel, G. De la Lanza-Espino, and C. González-Rebeles Islas. A methodology for evaluating the sustainability of inland wetland systems. Aquaculture International, 16 (6):525–537, December 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9163-5.

Wang:2008:TFR

[741] Youji Wang, Menghong Hu, Weimin Wang, Ling Cao, Yi Yang, Biping Lü, and Rongrong Yao. Transpositional feeding rhythm of loach Misgurnus anguillicaudatus from larvae to juveniles and its ontogenesis under artificial rearing conditions. Aquaculture International, 16(6): 539–549, December 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-007-9164-4.

Seo:2008:EDM

[742] Joo-Young Seo and Sang-Min Lee. Effects of dietary macronutrient level and feeding frequency on growth and body composition of juvenile rockfish (Sebastes schlegeli). Aquaculture International, 16(6): 551–560, December 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9165-y.

Babiak:2008:SMC

[743] Igor Babiak, Sylvie Bolla, and Oddvar Ottesen. Suitable methods for cryopreservation of semen from Atlantic halibut, Hippoglossus hippoglossus L. Aquaculture International, 16(6):561–572, December 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-008-9167-9.

delRio-Portilla:2008:SCM

[744] Miguel A. del Rio-Portilla and Andy R. Beaumont. Sperm concentration in the mussel Mytilus edulis L.: a spectrophotometric measurement protocol. Aquaculture International, 16(6):573–580, December 2008. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-008-9168-8.

Zheng:2008:ETS

[745] Zhongming Zheng, Chunhua Jin, Mingyun Li, Peifeng Bai, and Shuanglin Dong. Effects of temperature and salinity on oxygen consumption and ammonia excretion of juvenile miiuy croaker, Miichthys miiuy (Basilewsky). Aquaculture International, 16(6):581–589, December 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-008-9169-7.

Zhou:2008:ICA

[746] Xuxia Zhou, Yanbo Wang, Yuanjiang Pan, and Weifen Li. Investigation of the cytotoxicity of apidaecin on intestinal epithelial cells of tilapia (Oreochromis niloticus) in vitro. Aquaculture International, 16 (6):591–600, December 2008. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9170-1.

Mazur:2008:UCP

[747] Nicole A. Mazur and Allan L. Curtis. Understanding community perceptions of aquaculture: lessons from Australia. Aquaculture International, 16(6):601–621, December 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-008-9171-0.

Xu:2008:IDG

[748] Jian-He Xu, Feng You, Wei Sun, Bin-Lun Yan, Pei-Jun Zhang, and Bi-Xiang Jing. Induction of diploid gynogenesis in turbot Scophthalmus maximus with left-eyed flounder Paralichthys olivaceus sperm. Aquaculture International, 16(6):623–634, December 2008. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-008-9173-y.

Bodis:2009:EDD

[749] M. Bódis and M. Bercsényi. The effect of different daily feed rations on the growth, condition, survival and feed conversion of juvenile pikeperch (Sander lucioperca) reared with dry feed in net cages. Aquaculture International, 17(1):1–6, February 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-008-9226-2.

Simontacchi:2009:WBC

[750] Claudia Simontacchi, E. Negrato, M. Pazzaglia, D. Bertotto, C. Poltronieri, and G. Radaelli. Whole-body concentrations of cortisol and sex steroids in white sturgeon (Acipenser transmontanus, Richardson 1836) during early development and stress response. Aquaculture International, 17(1):7–14, February 2009. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9174-x.

Duc:2009:ECF

[751] Nguyen Minh Duc. Economic contribution of fish culture to farm income in Southeast Vietnam. Aquaculture International, 17(1):15–29, February 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-008-9176-8.

Harlioglu:2009:CGS

[752] Muzaffer Mustafa Harlıoğlu. A comparison of the growth and survival of two freshwater crayfish species, Astacus leptodactylus Eschscholtz and Pacifastacus leniusculus (Dana), under different temperature and density regimes. Aquaculture International, 17(1):31–43, February 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-008-9177-7.

Sui:2009:EDS

[753] L. Y. Sui, X. G. Wu, M. Wille, Y. X. Cheng, and P. Sorgeloos. Effect of dietary soybean lecithin on reproductive performance of Chinese mitten crab Eriocheir sinensis (H. Milne–Edwards) broodstock. Aquaculture International, 17(1):45–56, February 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-008-9178-6.

Kovitvadhi:2009:MRW

[754] Satit Kovitvadhi, Uthaiwan Kovitvadhi, Pichan Sawangwong, Premjai Trisaranuwatana, and Jorge Machado. Morphometric relationship of weight and size of cultured freshwater pearl mussel, Hyriopsis (Limnoscapha) myersiana, under laboratory conditions and earthen pond phases. Aquaculture International, 17(1):57–67, February 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-008-9180-z.

Huang:2009:SDD

[755] Zhi-Yong Huang, Qing-Pi Yan, Qiang Zhang, and Ai-Hong Peng. Sample digestion for determining chloramphenicol residues in carp serum and muscle. Aquaculture International, 17(1):69–76, February 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-008-9181-y.

Kritsanapuntu:2009:GWQ

[756] Sirusa Kritsanapuntu, Nilnaj Chaitanawisuti, and Yutaka Natsukari. Growth and water quality for growing-out of juvenile spotted Babylon, Babylonia areolata, at different water-exchange regimes in a largescale operation of earthen ponds. Aquaculture International, 17(1): 77–84, February 2009. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9182-x.

Hipolito-Morales:2009:UMS

[757] Araceli Hipólito-Morales, A. M. Maeda-Martínez, and S. F. Martínez-Díaz. Use of Microbacterium sp. and Exiguobacterium mexicanum to improve the survival and development of Artemia under xenic conditions. Aquaculture International, 17(1):85–90, February 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-008-9175-9.

Okamura:2009:ASM

[758] Akihiro Okamura, Hideo P. Oka, Noriyuki Horie, Tomoko Utoh, Yoshiaki Yamada, Naomi Mikawa, Satoru Tanaka, and Katsumi Tsukamoto. Assessing sexual maturity of feminized Japanese eel Anguilla japonica by measuring eye size. Aquaculture International, 17(1):91–99, February 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-008-9172-z.

Pirozzi:2009:ESD

[759] Igor Pirozzi, Mark A. Booth, and Patricia M. Pankhurst. The effect of stocking density and repeated handling on the growth of juvenile mulloway, Argyrosomus japonicus (Temminck & Schlegel 1843). Aquaculture International, 17(2):??, April 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-008-9190-x.

Badillo-Salas:2009:CGG

[760] Circe E. Badillo-Salas, Enrique Valenzuela-Espinoza, Marco A. González-Gómez, Griselda Pares-Sierra, Francisco Ley-Lou, and Zaul Garcia-Esquivel. Comparative growth of Pacific oyster (Crassostrea gigas) postlarvae with microfeed and microalgal diets. Aquaculture International, 17(2):??, April 2009. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9189-3.

Zhang:2009:SRM

[761] Peidong Zhang, Xiumei Zhang, Jian Li, and Tianxiang Gao. Starvation resistance and metabolic response to food deprivation and recovery feeding in Fenneropenaeus chinensis juveniles. Aquaculture International, 17(2):??, April 2009. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9188-4.

Wang:2009:EPP

[762] Feng Wang, Yu-Hong Yang, Zhi-Zhong Han, Hong-Wei Dong, Cheng-Hui Yang, and Zuo-Yu Zou. Effects of phytase pretreatment of soybean meal and phytase-sprayed in diets on growth, apparent digestibility coefficient and nutrient excretion of rainbow trout (Oncorhynchus mykiss Walbaum). Aquaculture International, 17(2):??, April 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-008-9187-5.

Sanchez-Barajas:2009:DYH

[763] Maximiliano Sánchez-Barajas, Marco Agustín Liñán-Cabello, and Alfredo Mena-Herrera. Detection of yellow-head disease in intensive freshwater production systems of Litopenaeus vannamei. Aquaculture International, 17(2):101–112, April 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-008-9183-9.

Campos:2009:NFE

[764] Susmara Silva Campos, Ugo Lima Silva, Maria Zita Tabosa Lúcio, and Eudes de Souza Correia. Natural food evaluation and water quality in zero water exchange culture of Litopenaeus vannamei fertilized with wheat bran. Aquaculture International, 17(2):113–124, April 2009. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-008-9184-8.

Miao:2009:EEA

[765] Sha Miao, Chen Cheun Jen, Cheng Ting Huang, and Sing-Hwa Hu. Ecological and economic analysis for cobia Rachycentron canadum commercial cage culture in Taiwan. Aquaculture International, 17(2):125– 141, April 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-008-9185-7.

Rasmussen:2009:DSR

[766] Richard Skøtt Rasmussen and Alfred Jokumsen. Digestibility in selected rainbow trout families and relation to growth and feed utilisation. Aquaculture International, 17(2):187–197, April 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-008-9191-9.

Norouzitallab:2009:CED

[767] P. Norouzitallab, M. Farhangi, M. Babapour, R. Rahimi, A. K. Sinha, and K. Baruah. Comparing the efficacy of dietary α-tocopherol with that of dl-α-tocopheryl acetate, both either alone or in combination with ascorbic acid, on growth and stress resistance of angelfish, Pterophylum scalare, juveniles. Aquaculture International, 17(3):207–216, June 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-008-9192-8.

Imsland:2009:BAU

[768] Albert K. Imsland, Mads D. Jenssen, Thor M. Jonassen, and Sigurd O. Stefansson. Best among unequals? Effect of size grading and different social environments on the growth performance of juvenile Atlantic halibut. Aquaculture International, 17(3):217–227, June 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-008-9193-7.

Guo:2009:AEC

[769] Longgen Guo, Zhongjie Li, Ping Xie, and Leyi Ni. Assessment effects of cage culture on nitrogen and phosphorus dynamics in relation to fallowing in a shallow lake in China. Aquaculture International, 17 (3):229–241, June 2009. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9195-5.

Naz:2009:CDEa

[770] Mehmet Naz and Mustafa Türkmen. The changes in digestive enzymes and hormones of gilthead seabream larvae (Sparus aurata, L. 1758) fed on Artemia nauplii enriched with free methionine. Aquaculture International, 17(3):243–256, June 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-008-9196-4.

Matias:2009:EGO

[771] Domitília Matias, Sandra Joaquim, Alexandra Leitão, and Clara Massapina. Effect of geographic origin, temperature and timing of broodstock collection on conditioning, spawning success and larval viability of Ruditapes decussatus (Linné, 1758). Aquaculture International, 17 (3):257–271, June 2009. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9197-3.

Albentosa:2009:DDB

[772] M. Albentosa and F. J. Moyano. Differences in the digestive biochemistry between the intertidal clam, Ruditapes decussatus, and the subtidal clam, Venerupis pullastra. Aquaculture International, 17(3):273– 282, June 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-008-9199-1.

Domingues:2009:EFT

[773] Pedro M. Domingues, Lorenzo Marquez, Nelda López, and Carlos Rosas. Effects of food thermal treatment on growth, absorption, and assimilation efficiency of juvenile cuttlefish, Sepia officinalis. Aquaculture International, 17(3):283–299, June 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-008-9200-z.

Jobling:2009:BRC

[774] Malcolm Jobling. Book review: C. S. Tucker, J. A. Hargreaves (eds): Environmental Best Management Practices for Aquaculture. Aquaculture International, 17(3):301–302, June 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-008-9194-6.

Qi:2009:EFR

[775] Zizhong Qi, Kristof Dierckens, Tom Defoirdt, Patrick Sorgeloos, Nico Boon, Zhenmin Bao, and Peter Bossier. Effects of feeding regime and probionts on the diverting microbial communities in rotifer Brachionus culture. Aquaculture International, 17(4):303–315, August 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-008-9202-x.

Pourasadi:2009:MIS

[776] Mohsen Pourasadi, Bahram Falahatkar, and Ghobad Azari Takami. Minimally invasive surgical technique for egg collection from the Persian sturgeon, Acipenser persicus. Aquaculture International, 17(4): 317–321, August 2009. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9203-9.

Li:2009:GDF

[777] J. L. Li, G. L. Wang, and Z. Y. Bai. Genetic diversity of freshwater pearl mussel (Hyriopsis cumingii) in populations from the five largest lakes in China revealed by inter-simple sequence repeat (ISSR). Aquaculture International, 17(4):323–330, August 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-008-9204-8.

Jones:2009:SST

[778] M. Jones, S. C. Battaglene, and P. M. Pankhurst. Sensitivity of striped trumpeter, Latris lineata, embryos to mechanical shock and simulated transport. Aquaculture International, 17(4):331–340, August 2009. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-008-9205-7.

Espmark:2009:EHB

[779] Åsa Maria Espmark and Grete Baeverfjord. Effects of hyperoxia on behavioural and physiological variables in farmed Atlantic salmon (Salmo salar) parr. Aquaculture International, 17(4):341–353, August 2009. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-008-9206-6.

Ergun:2009:IUM

[780] Sebahattin Ergün, Murat Soyutürk, Betül Güroy, Derya Güroy, and Daniel Merrifield. Influence of Ulva meal on growth, feed utilization, and body composition of juvenile Nile tilapia (Oreochromis niloticus) at two levels of dietary lipid. Aquaculture International, 17(4): 355–361, August 2009. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9207-5.

Hii:2009:FSC

[781] Yii-Siang Hii, Chen-Lin Soo, and Hock-Chark Liew. Feeding of scleractinian coral, Galaxea fascicularis, on Artemia salina nauplii in captivity. Aquaculture International, 17(4):363–376, August 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-008-9208-4.

Yu:2009:EDM

[782] Ming-Chao Yu, Zhuo-Jia Li, Hei-Zhao Lin, Guo-Liang Wen, and Shen Ma. Effects of dietary medicinal herbs and Bacillus on survival, growth, body composition, and digestive enzyme activity of the white shrimp Litopenaeus vannamei. Aquaculture International, 17(4):377–384, August 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-008-9209-3.

Gunnarsson:2009:CSS

[783] Snorri Gunnarsson, Sindri Sigurdsson, Helgi Thorarensen, and Albert K. Imsland. Cryopreservation of sperm from spotted wolffish. Aquaculture International, 17(4):385–389, August 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-008-9210-x.

Durmaz:2009:ETT

[784] Y. Durmaz, M. Donato, M. Monteiro, L. Gouveia, M. L. Nunes, T. Gama Pereira, Ş. Gökpınar, and N. M. Bandarra. Effect of temperature on α-tocopherol, fatty acid profile, and pigments of Diacronema vlkianum (Haptophyceae). Aquaculture International, 17(4):391–399, August 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-008-9211-9.

Jobling:2009:BRR

[785] Malcolm Jobling. Book review: L. Ross and B. Ross, Anaesthetic and Sedative Techniques for Aquatic Animals (Third Edition). Aquaculture International, 17(4):401–402, August 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-008-9201-y.

Enriquez-Diaz:2009:GRI

[786] M. Enríquez-Díaz, S. Pouvreau, J. Chávez-Villalba, and M. Le Pennec. Gametogenesis, reproductive investment, and spawning behavior of the Pacific giant oyster Crassostrea gigas: evidence of an environmentdependent strategy. Aquaculture International, 17(5):??, October 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-008-9219-1.

Soto-Hernandez:2009:WLG

[787] J. Soto-Hernández, A. Velasco-Murillo, and J. M. Grijalva-Chon. When low genetic variability in feral and hatchery-reared tilapia becomes a vicious circle: Oreochromis niloticus from Oaxaca, Mexico. Aquaculture International, 17(5):??, October 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-008-9217-3.

An:2009:HPT

[788] Zhen-Hua An, Yun-Wei Dong, and Shuang-Lin Dong. A highperformance temperature-control scheme: growth of sea cucumber Apostichopus japonicus with different modes of diel temperature fluctuation. Aquaculture International, 17(5):??, October 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-008-9216-4.

Wassef:2009:VOB

[789] Elham A. Wassef, Norhan E. Saleh, and Heyam A. El-Abd El-Hady. Vegetable oil blend as alternative lipid resources in diets for gilthead seabream, Sparus aurata. Aquaculture International, 17(5):??, October 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-008-9213-7.

Radiarta:2009:BMG

[790] I Nyoman Radiarta and Sei-Ichi Saitoh. Biophysical models for Japanese scallop, Mizuhopecten yessoensis, aquaculture site selection in Funka Bay, Hokkaido, Japan, using remotely sensed data and geographic information system. Aquaculture International, 17(5):??, October 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-008-9212-8.

Dias:2009:MSU

[791] P. J. Dias, M. Bland, A. M. Shanks, A. Beaumont, S. B. Piertney, I. M. Davies, and M. Snow. Mytilus species under rope culture in Scotland: implications for management. Aquaculture International, 17(5): 437–448, October 2009. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9214-6.

Cai:2009:CPA

[792] Junpeng Cai, Shuanyu Lin, and Bing Wu. Characterization of Pseudomonas aeruginosa associated with diseased postlarval abalone in Shenzhen, China. Aquaculture International, 17(5):449–458, October 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-008-9215-5.

Kumar:2009:SBS

[793] V. J. Rejish Kumar, Cini Achuthan, N. J. Manju, Rosamma Philip, and I. S. Bright Singh. Stringed bed suspended bioreactors (SBSBR) for in situ nitrification in penaeid and non-penaeid hatchery systems. Aquaculture International, 17(5):479–489, October 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-008-9218-2.

Leef:2009:PIK

[794] Melanie J. Leef and Peter S. Lee. Preliminary investigation into the killing effect of kingfish (Seriola lalandi) serum and mucus against the monogenean parasites Benedenia seriolae and Zeuxapta seriolae. Aquaculture International, 17(6):??, December 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-008-9228-0.

Carvalho:2009:BDW

[795] Susana Carvalho, Manuela Falcão, João Cúrdia, Ana Moura, Dalila Serpa, Miguel B. Gaspar, Maria Teresa Dinis, Pedro Pousão-Ferreira, and Luís Cancela da Fonseca. Benthic dynamics within a land-based semiintensive aquaculture fish farm: the importance of settlement ponds. Aquaculture International, 17(6):??, December 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-008-9227-1.

Celada:2009:RJT

[796] Jesús D. Celada, Amelia Aguilera, Vanesa García, José M. Carral, María Sáez-Royuela, Rocío González, and Álvaro González. Rearing juvenile tench (Tinca tinca L.) under controlled conditions using Artemia nauplii as supplement to a dry diet. Aquaculture International, 17 (6):??, December 2009. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9225-3.

Park:2009:DQC

[797] Kwan Ha Park, Gustavo A. Rodriguez-Montes de Oca, Pierluigi Bonello, Kyeong-Jun Lee, and Konrad Dabrowski. Determination of quercetin concentrations in fish tissues after feeding quercetin-containing diets. Aquaculture International, 17(6):??, December 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-008-9222-6.

Naz:2009:CDEb

[798] Mehmet Naz and Mustafa Türkmen. Changes in the digestive enzymes and hormones of gilthead seabream larvae (Sparus aurata, L. 1758) fed on Artemia nauplii enriched with free lysine. Aquaculture International, 17(6):??, December 2009. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9221-7.

Azaza:2009:GPF

[799] Mohamed Salah Azaza, Wassim Kammoun, Abdelwaheb Abdelmouleh, and Mohamed Mejdeddine Kraïem. Growth performance, feed utilization, and body composition of Nile tilapia (Oreochromis niloticus L.) fed with differently heated soybean-meal-based diets. Aquaculture International, 17(6):507–521, December 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-008-9220-8.

Pantazis:2009:CPP

[800] Panagiotis A. Pantazis. The culture potential of Paracentrotus lividus (Lamarck 1816) in Greece: a preliminary report. Aquaculture International, 17(6):545–552, December 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-008-9223-5.

Cek:2009:EVD

[801] Şehriban Çek and Erdal Yilmaz. The effect of varying dietary energy on gonad development at first sexual maturity of the sharptooth catfish (Clarias gariepinus Burchell, 1822). Aquaculture International, 17(6): 553–563, December 2009. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9224-4.

Gionet:2009:EER

[802] Chantal Gionet, Elise Mayrand, and Thomas Landry. The effect of energy reserves and cryoprotectants on overwintering mortality in Mercenaria mercenaria notata (say 1822) at two tidal levels. Aquaculture International, 17(6):589–605, December 2009. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-008-9229-z.

Smaal:2010:ERB

[803] Aad C. Smaal and Jeroen W. M. Wijsman. Exploitation and restoration of bivalve shellfish resources: case studies presented at the tenth international conference on shellfish restoration. Aquaculture International, 18(1):1–2, February 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9281-3.

Dolmer:2010:IAB

[804] Per Dolmer and Ea Stenalt. The impact of the adult blue mussel (Mytilus edulis) population on settling of conspecific larvae. Aquaculture International, 18(1):3–17, February 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-009-9266-2.

Lebata-Ramos:2010:GSH

[805] M. Junemie Hazel L. Lebata-Ramos, Koichi Okuzawa, Ronald J. Maliao, Jeff Bogart R. Abrogueña, Mark Darwin N. Dimzon, Ellen Flor C. Doyola-Solis, and Terence U. Dacles. Growth and survival of hatcherybred giant clams (Tridacna gigas) in an ocean nursery in Sagay Marine Reserve, Philippines. Aquaculture International, 18(1):19–33, February 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-009-9272-4.

Roncarati:2010:TGP

[806] Alessandra Roncarati, Alberto Felici, Andrea Dees, Forlini Leila, and Melotti Paolo. Trials on Pacific oyster (Crassostrea gigas Thunberg) rearing in the middle Adriatic Sea by means of different trays. Aquaculture International, 18(1):35–43, February 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-009-9267-1.

Fey:2010:DDN

[807] Frouke Fey, Norbert Dankers, Josien Steenbergen, and Kees Goudswaard. Development and distribution of the non-indigenous Pacific oyster (Crassostrea gigas) in the Dutch Wadden Sea. Aquaculture International, 18 (1):45–59, February 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9268-0.

Rossi-Snook:2010:OCV

[808] Kate Rossi-Snook, Gulnihal Ozbay, and Frank Marenghi. Oyster (Crassostrea virginica) gardening program for restoration in Delaware’s Inland Bays, USA. Aquaculture International, 18(1):61–67, February 2010. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9271-5.

Marenghi:2010:CHV

[809] Frank Marenghi, Gulnihal Ozbay, Patrick Erbland, and Kate Rossi-Snook. A comparison of the habitat value of sub-tidal and floating oyster (Crassostrea virginica) aquaculture gear with a created reef in Delaware’s Inland Bays, USA. Aquaculture International, 18(1):69– 81, February 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-009-9273-3.

Marsden:2010:CSC

[810] Islay D. Marsden and Suzanne C. Adkins. Current status of cockle bed restoration in New Zealand. Aquaculture International, 18(1):83– 97, February 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-009-9270-6.

Frechette:2010:HSB

[811] Marcel Fréchette. Hierarchical structure of bivalve culture systems and optimal stocking density. Aquaculture International, 18(1):99–114, February 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-008-9198-2.

Smaal:2010:SIC

[812] Aad C. Smaal and Jeroen W. M. Wijsman. Synthesis of the 10th ICSR conference: innovation in the exploitation and management of shellfish resources. Aquaculture International, 18(1):115–117, February 2010. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9283-1.

Garcia-Ortega:2010:LID

[813] Armando García-Ortega, Johan Verreth, Karen Vermis, Hans J. Nelis, Patrick Sorgeloos, and Martin Verstegen. Laboratory investigation of daily food intake and gut evacuation in larvae of African catfish Clarias gariepinus under different feeding conditions. Aquaculture International, 18(2):119–134, April 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9230-6.

Wang:2010:ETI

[814] Youji Wang, Menghong Hu, Weimin Wang, S. G. Cheung, P. K. S. Shin, and Ling Cao. Effects of the timing of initial feeding on growth and survival of loach (Misgurnus anguillicaudatus) larvae. Aquaculture International, 18(2):135–148, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-008-9231-5.

Brigolin:2010:MBY

[815] Daniele Brigolin, Roberto Pastres, Paolo Tomassetti, and Salvatore Porrello. Modelling the biomass yield and the impact of seabream mariculture in the Adriatic and Tyrrhenian Seas (Italy). Aquaculture International, 18(2):149–163, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-008-9232-4.

Domingues:2010:ETC

[816] Pedro Domingues, Sandra Garcia, and Diego Garrido. Effects of three culture densities on growth and survival of Octopus vulgaris (Cuvier, 1797). Aquaculture International, 18(2):165–174, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-008-9233-3.

Khan:2010:ORL

[817] Mukhtar A. Khan and Shabi F. Abidi. Optimum ration level for better growth, conversion efficiencies and body composition of fingerling Heteropneustes fossilis (Bloch). Aquaculture International, 18(2):175– 188, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-008-9234-2.

Guzman:2010:EMF

[818] Héctor Mendoza Guzmán, Adelina de la Jara Valido, Laura Carmona Duarte, and Karen Freijanes Presmanes. Estimate by means of flow cytometry of variation in composition of fatty acids from Tetraselmis suecica in response to culture conditions. Aquaculture International, 18 (2):189–199, April 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9235-1.

Barrington:2010:SAS

[819] Kelly Barrington, Neil Ridler, Thierry Chopin, Shawn Robinson, and Bryn Robinson. Social aspects of the sustainability of integrated multi-trophic aquaculture. Aquaculture International, 18(2):201–211, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-008-9236-0.

Selcuk:2010:EDC

[820] Zehra Selcuk, Serap Ustaoglu Tiril, Fikret Alagil, Volkan Belen, Mustafa Salman, Sena Cenesiz, Omer Hakan Muglali, and Feraye Berkay Yagci. Effects of dietary l-carnitine and chromium picolinate supplementations on performance and some serum parameters in rainbow trout (Oncorhynchus mykiss). Aquaculture International, 18(2):213– 221, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-008-9237-z.

Jobling:2010:CLO

[821] Malcolm Jobling and Odd Leknes. Cod liver oil: feed oil influences on fatty acid composition. Aquaculture International, 18(2):223–230, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-008-9238-y.

SanchezDeBock:2010:SRG

[822] Mariana Sánchez De Bock and Laura Susana López Greco. Sex reversal and growth performance in juvenile females of the freshwater crayfish Cherax quadricarinatus (Parastacidae): effect of increasing temperature and androgenic gland extract in the diet. Aquaculture International, 18 (2):231–243, April 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-008-9239-x.

Xu:2010:RBC

[823] Lipu Xu, Bangxi Xiong, Yong Pan, Jingbo Wang, Huan Cao, and Wen Zhao. Relationship between concentrations of odorous compounds and biomass of phytoplankton and actinomycetes in freshwater ponds of Beijing, China. Aquaculture International, 18(3):245–254, April 2010. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9240-z.

Huai:2010:EDP

[824] Ming-Yan Huai, Yong-Jian Liu, Li-Xia Tian, Shi-Xi Deng, An-Long Xu, Wen Gao, and Hui-Jun Yang. Effect of dietary protein reduction with synthetic amino acids supplementation on growth performance, digestibility, and body composition of juvenile Pacific white shrimp, Litopenaeus vannamei. Aquaculture International, 18(3):255–269, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9241-y.

Barord:2010:DES

[825] G. J. Barord, K. N. Keister, and P. G. Lee. Determining the effects of stocking density and temperature on growth and food consumption in the pharaoh cuttlefish, Sepia pharaonis, Ehrenberg 1890. Aquaculture International, 18(3):271–283, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-009-9242-x.

Tabinda:2010:EHP

[826] Amtul Bari Tabinda and Mohammad Ayub. Effect of high phosphate fertilization rate on pond phosphate concentrations, chlorophyll a, and fish growth in carp polyculture. Aquaculture International, 18(3):285– 301, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-009-9243-9.

Chaitanawisuti:2010:EDC

[827] Nilnaj Chaitanawisuti, Tosapon Sungsirin, and Somkiat Piyatiratitivorakul. Effects of dietary calcium and phosphorus supplementation on the growth performance of juvenile spotted babylon Babylonia areolata culture in a recirculating culture system. Aquaculture International, 18 (3):303–313, April 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9244-8.

Galley:2010:OLC

[828] Thomas H. Galley, Frederico M. Batista, Richard Braithwaite, Jon King, and Andy R. Beaumont. Optimisation of larval culture of the mussel Mytilus edulis (L.). Aquaculture International, 18(3):315–325, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9245-7.

Moschino:2010:UBA

[829] Vanessa Moschino, Francesca Meneghetti, and Luisa Da Ros. Use of biomarkers to assess the welfare of the edible clam, Ruditapes philippinarum: may it be a tool for proving areas of origin? Aquaculture International, 18(3):327–337, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-009-9246-6.

Huang:2010:ERT

[830] Jianrong Huang, Wenhui Lin, Cunbing Shi, Shuqin Wu, and Runlin Xu. The effects of restoration techniques on protozoan communities in mandarin fish culture ponds, based on an artificial substrate. Aquaculture International, 18(3):339–348, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-009-9247-5.

Zhou:2010:CSS

[831] Xiaoyun Zhou, Khalid Abbas, Mingyun Li, Libao Fang, Su Li, and Weimin Wang. Comparative studies on survival and growth performance among diploid, triploid and tetraploid dojo loach Misgurnus anguillicaudatus. Aquaculture International, 18(3):349–359, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-009-9248-4.

Domingues:2010:ECD

[832] P. Domingues and L. Márquez. Effects of culture density and bottom area on growth and survival of the cuttlefish Sepia officinalis (Linnaeus, 1758). Aquaculture International, 18(3):361–369, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-009-9249-3.

Gonzalez:2010:SDI

[833] R. González, J. D. Celada, A. González, V. García, J. M. Carral, and M. Sáez-Royuela. Stocking density for the intensive rearing of juvenile crayfish, Pacifastacus leniusculus (Astacidae), using Artemia nauplii to supplement a dry diet from the onset of exogenous feeding. Aquaculture International, 18(3):371–378, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-009-9250-x.

Soto-Zarazua:2010:FLB

[834] Genaro M. Soto-Zarazúa, Enrique Rico-García, Rosalía Ocampo, R. G. Guevara-González, and Gilberto Herrera-Ruiz. Fuzzy-logic-based feeder system for intensive tilapia production (Oreochromis niloticus). Aquaculture International, 18(3):379–391, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-009-9251-9.

Quagrainie:2010:AUC

[835] Kwamena K. Quagrainie, Charles C. Ngugi, and Stephen Amisah. Analysis of the use of credit facilities by small-scale fish farmers in Kenya. Aquaculture International, 18(3):393–402, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-009-9252-8.

Citarasu:2010:HBN

[836] Thavasimuthu Citarasu. Herbal biomedicines: a new opportunity for aquaculture industry. Aquaculture International, 18(3):403–414, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-009-9253-7.

Sicuro:2010:PEN

[837] Benedetto Sicuro, Paola Badino, Franco Daprà, Francesco Gai, Marco Galloni, Rosangela Odore, Giovanni Battista Palmegiano, and Elisabetta Macchi. Physiological effects of natural olive oil antioxidants utilization in rainbow trout (Onchorynchus mykiss) feeding. Aquaculture International, 18(3):415–431, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-009-9254-6.

Gustavsson:2010:GBC

[838] Arnthór Gústavsson, Albert K. Imsland, Snorri Gunnarsson, Jón Árnason, Ingólfur Arnarson, Arnar F. Jónsson, Heiddís Smáradóttir, and Helgi Thorarensen. Growth and blood chemistry of Atlantic halibut (Hippoglossus hippoglossus L.) in relation to salinity and continuous light. Aquaculture International, 18(3):433–445, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-009-9255-5.

Li:2010:ESD

[839] Li Li and Qi Li. Effects of stocking density, temperature, and salinity on larval survival and growth of the red race of the sea cucumber Apostichopus japonicus (Selenka). Aquaculture International, 18 (3):447–460, April 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9256-4.

Roth:2010:IAM

[840] Bjorn Roth, Albert Imsland, Lars Helge Stien, Rian Schelvis-Smit, Snorri Gunnarsson, and Atle Foss. The influence of anaerobic muscle activity, maturation and season on the flesh quality of farmed turbot. Aquaculture International, 18(3):461–474, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-009-9257-3.

Herrera:2010:WWS

[841] Marcelino Herrera, Ismael Hachero-Cruzado, Catarina Oliveira, José F. Ferrer, José M. Márquez, Montserrat Rosano, and Jose I. Navas. Weaning of the wedge sole Dicologoglossa cuneata (Moreau): influence of initial size on survival and growth. Aquaculture International, 18(3): 475–485, April 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9258-2.

Domingues:2010:UAP

[842] P. Domingues, S. Garcia, I. Hachero-Cruzado, N. Lopez, and C. Rosas. The use of alternative prey (crayfish, Procambarus clarki, and hake, Merlucius gayi) to culture Octopus vulgaris (Cuvier 1797). Aquaculture International, 18(3):487–499, April 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-009-9259-1.

Jobling:2010:CGC

[843] Malcolm Jobling. Are compensatory growth and catch-up growth two sides of the same coin? Aquaculture International, 18(4):501– 510, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-009-9260-8.

Sicuro:2010:OOP

[844] Benedetto Sicuro, Franco Daprà, Francesco Gai, Giovanni Battista Palmegiano, Roberta Schiavone, Loredana Zilli, and Sebastiano Vilella. Olive oil by-product as a natural antioxidant in gilthead sea bream (Sparus aurata) nutrition. Aquaculture International, 18(4):511–522, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-009-9262-6.

Mohanty:2010:IPH

[845] Rajeeb K. Mohanty. Impact of phased harvesting on population structure, feed intake pattern and growth performance of Macrobrachium rosenbergii DeMan (giant freshwater prawn) in polyculture with carps in concurrent rice–fish culture. Aquaculture International, 18(4):523– 537, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-009-9263-5.

Giri:2010:RCF

[846] Shiba Shankar Giri, Sangram Ketan Sahoo, and Satyendra Nath Mohanty. Replacement of by-catch fishmeal with dried chicken viscera meal in extruded feeds: effect on growth, nutrient utilisation and carcass composition of catfish Clarias batrachus (Linn.) fingerlings. Aquaculture International, 18(4):539–544, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-009-9265-3.

Tsutsui:2010:GGT

[847] Isao Tsutsui, Poonmanee Kanjanaworakul, Prapansak Srisapoome, Dusit Aue-umneoy, and Kaoru Hamano. Growth of giant tiger prawn, Penaeus monodon Fabricius, under co-culture with a discarded filamentous seaweed, Chaetomorpha ligustica (Kützing) Kützing, at an aquarium-scale. Aquaculture International, 18(4):545–553, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-009-9274-2.

Ahmed:2010:RFP

[848] Nesar Ahmed, Edward H. Allison, and James F. Muir. Rice fields to prawn farms: a blue revolution in southwest Bangladesh? Aquaculture International, 18(4):555–574, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-009-9276-0.

Ding:2010:MAG

[849] Wei dong Ding, Zhe ming Cao, and Li ping Cao. Molecular analysis of grass carp (Ctenopharyngodon idella) by SRAP and SCAR molecular markers. Aquaculture International, 18(4):575–587, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-009-9277-z.

Onal:2010:HDD

[850] Umur Önal, İhsan Çelik, and Şükran Cirik. Histological development of digestive tract in discus, Symphysodon spp. larvae. Aquaculture International, 18(4):589–601, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-009-9278-y.

Zuniga:2010:DSA

[851] Sergio Zúñiga. A dynamic simulation analysis of Japanese abalone (Haliotis discus hannai) production in Chile. Aquaculture International, 18 (4):603–620, June 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9279-x.

Kohlmann:2010:SML

[852] Klaus Kohlmann, Gert Füllner, and Matthias Pfeifer. Segregation of microsatellite loci in second generation hybrid striped bass (Morone saxatilis × Morone chrysops). Aquaculture International, 18(4):621– 633, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-009-9282-2.

Leal:2010:USP

[853] Albino Luciani Gonçalves Leal, Patrícia Fernandes de Castro, João Paulo Viana de Lima, Eudes de Souza Correia, and Ranilson de Souza Bezerra. Use of shrimp protein hydrolysate in Nile tilapia (Oreochromis niloticus, L.) feeds. Aquaculture International, 18(4):635– 646, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-009-9284-0.

Sun:2010:EZR

[854] Weiming Sun, Shuanglin Dong, Xidan Zhao, Zilin Jie, Hanwen Zhang, and Lechun Zhang. Effects of zooplankton refuge on the growth of tilapia (Oreochromis niloticus) and plankton dynamics in pond. Aquaculture International, 18(4):647–655, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-009-9286-y.

Leal:2010:OPV

[855] C. A. G. Leal, G. A. Carvalho-Castro, P. S. C. Sacchetin, C. O. Lopes, A. M. Moraes, and H. C. P. Figueiredo. Oral and parenteral vaccines against Flavobacterium columnare: evaluation of humoral immune response by ELISA and in vivo efficiency in Nile tilapia (Oreochromis niloticus). Aquaculture International, 18(4):657–666, June 2010. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9287-x.

Chu:2010:GEP

[856] Wuying Chu, Guihong Fu, Jia Chen, Dinggen Chen, Tao Meng, Ruixue Zhou, Xinjie Xia, and Jianshe Zhang. Gene expression profiling in muscle tissues of the commercially important teleost, Siniperca chuatsi L. Aquaculture International, 18(4):667–678, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-009-9289-8.

Yusufzai:2010:EDM

[857] S. I. Yusufzai, H. Singh, and M. M. Shirdhankar. An evaluation of different methods for transportation of the freshwater mussel Lamellidens corrianus (Lea 1834). Aquaculture International, 18(4):679–692, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9290-2.

Tanyaros:2010:PTR

[858] Suwat Tanyaros, Apirak Songrak, Rattanaporn Anantasuk, Suchat Sangchan, and Bill Templer. Post-tsunami rehabilitation of fish cage farms on the Andaman coast of Thailand. Aquaculture International, 18(4):693–703, June 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9291-1.

Jobling:2010:BRMa

[859] Malcolm Jobling. Book review: M. Cargill and P. O’Connor: Writing scientific research articles: strategy and steps. Aquaculture International, 18(4):705–706, June 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9261-7.

Jobling:2010:BRG

[860] Malcolm Jobling. Book review: G. Helfman, B. B. Collette, D. E. Facey, and B. W. Bowen: The diversity of fishes: biology, evolution and ecology, 2nd edn. Aquaculture International, 18(4):707–708, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-009-9264-4.

Jobling:2010:BRJ

[861] Malcolm Jobling. Book review: J. Morrissey and J. L. Sumich: Introduction to the Biology of Marine Life, 9th edn. Aquaculture International, 18(4):709–710, June 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9275-1.

Jobling:2010:BRR

[862] Malcolm Jobling. Book review: R. R. Stickney: Aquaculture: an introductory text, 2nd edn. Aquaculture International, 18(4):711–712, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-009-9288-9.

Jobling:2010:BRMb

[863] Malcolm Jobling. Book review: I. M. Struthers and D. Rissik (eds): Plankton: a guide to their ecology and monitoring for water quality. Aquaculture International, 18(4):713–714, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-009-9285-z.

Jobling:2010:EBR

[864] Malcolm Jobling. Erratum to: Book Review: I. M. Suthers and D. Rissik (eds): Plankton: a guide to their ecology and monitoring for water quality. Aquaculture International, 18(4):715, June 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-009-9294-y. See [863].

Viau:2010:SSE

[865] Verónica E. Viau and Enrique M. Rodríguez. Substrate selection and effect of different substrates on survival and growth of juveniles of the freshwater crayfish Cherax quadricarinatus (von Martens 1868) (Decapoda, Parastacidae). Aquaculture International, 18(5):717–724, August 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9292-0.

Afero:2010:EAT

[866] Farok Afero, Sha Miao, and Arlenie A. Perez. Economic analysis of tiger grouper Epinephelus fuscoguttatus and humpback grouper Cromileptes altivelis commercial cage culture in Indonesia. Aquaculture International, 18(5):725–739, August 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9295-x.

Damaso-Rodrigues:2010:LEF

[867] Maria Luísa Dâmaso-Rodrigues, Pedro Pousão-Ferreira, Laura Ribeiro, Joana Coutinho, Narcisa M. Bandarra, Paulo J. Gavaia, Luís Narciso, and Sofia Morais. Lack of essential fatty acids in live feed during larval and post-larval rearing: effect on the performance of juvenile Solea senegalensis. Aquaculture International, 18(5):741–757, August 2010. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9296-9.

Naegel:2010:MSR

[868] Ludwig C. A. Naegel. Management strategies to reduce operating costs in a commercial shrimp hatchery in NW Mexico. Aquaculture International, 18(5):759–770, August 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9297-8.

Bastami:2010:SFC

[869] Kazem Darvish Bastami, Mohammad Reza Imanpour, and Seyed Hossein Hoseinifar. Sperm of feral carp Cyprinus carpio: optimization of activation solution. Aquaculture International, 18(5):771–776, August 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9299-6.

Zhao:2010:EYP

[870] Wen Zhao, Miao Liang, and Peng Zhang. Effect of yeast polysaccharide on the immune function of juvenile sea cucumber, Apostichopus japonicus Selenka under pH stress. Aquaculture International, 18(5): 777–786, August 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9300-4.

Drummond:2010:SHS

[871] L. Drummond, M. F. Mulcahy, and S. C. Culloty. A survey of the health status of the Manila clam Ruditapes philippinarum in Ireland with specific reference to brown ring disease. Aquaculture International, 18(5): 787–800, August 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9301-3.

Gionet:2010:EAB

[872] Chantal Gionet, Elise Mayrand, and Thomas Landry. Elimination of animals with best growth potential as a possible effect of the culling of Mercenaria mercenaria notata (L.) larvae in hatchery procedure. Aquaculture International, 18(5):801–812, August 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-009-9298-7.

Yu:2010:SVG

[873] Zonghe Yu, Baozhong Liu, Hongsheng Yang, Yi Zhou, Kun Xing, Qiang Xu, and Libin Zhang. Seasonal variations in growth and clearance rate of the Zhikong scallop Chlamys farreri suspended in the deep water of Haizhou Bay, China. Aquaculture International, 18(5): 813–824, August 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9302-2.

Wang:2010:FGL

[874] Wei-Ji Wang, Han-Ping Wang, Hong Yao, Geoff K. Wallat, Laura G. Tiu, and Qing-Yin Wang. A first genetic linkage map of bluegill sunfish (Lepomis macrochirus) using AFLP markers. Aquaculture International, 18(5):825–835, August 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9303-1.

Nazari:2010:ASH

[875] Rajab Mohammad Nazari, Maryam Modanloo, Mohammad Reza Ghomi, and Mahmoud Reza Ovissipor. Application of synthetic hormone LHRH-A_{2} on the artificial propagation of Persian sturgeon Acipenser persicus. Aquaculture International, 18(5):837–841, August 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-009-9304-0.

Harlioglu:2010:EDM

[876] Muzaffer Mustafa Harlioğlu and Tuba Çakmak Duran. The effect of darkness on mating and pleopodal egg production time in a freshwater crayfish, Astacus leptodactylus Eschscholtz. Aquaculture International, 18(5):843–849, August 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9305-z.

Karami:2010:ETJ

[877] A. Karami, A. Christianus, Z. Ishak, S. C. Courtenay, M. A. Syed, M. Noor Azlina, and H. Noorshinah. Effect of triploidization on juvenile African catfish (Clarias gariepinus). Aquaculture International, 18(5):851–858, August 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9307-x.

Chatterjee:2010:ESD

[878] N. Chatterjee, A. K. Pal, T. Das, R. Dalvi, M. S. Mohammad, K. Sarma, S. C. Mukherjee, and K. Baruah. Effect of stocking density and journey length on the welfare of rohu (Labeo rohita Hamilton) fry. Aquaculture International, 18(5):859–868, August 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-009-9309-8.

Kumar:2010:EIS

[879] Vivek Kumar, Ravi Fotedar, and Ken Dods. Effect of inland saline water ionic profiles on growth, chemical composition and agar characteristics of Gracilaria cliftonii (Withell, Miller and Kraft 1994) under laboratory conditions. Aquaculture International, 18(5):869–881, August 2010. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9306-y.

Pham:2010:SMW

[880] Minh Anh Pham, Gyu-Deok Hwang, Yi-Oh Kim, Joo-Young Seo, and Sang-Min Lee. Soybean meal and wheat flour, proper dietary protein sources for optimal growth of snail (Semisulcospira coreana). Aquaculture International, 18(5):883–895, August 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-009-9308-9.

Luo:2010:DTP

[881] Z. Luo, X.-Y. Tan, X. Liu, and W.-M. Wang. Dietary total phosphorus requirement of juvenile yellow catfish Pelteobagrus fulvidraco. Aquaculture International, 18(5):897–908, August 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-009-9310-2.

Bonaldo:2010:GFU

[882] Alessio Bonaldo, Gloria Isani, Ramon Fontanillas, Luca Parma, Ester Grilli, and Pier Paolo Gatta. Growth and feed utilization of gilthead sea bream (Sparus aurata, L.) fed to satiation and restrictively at increasing dietary energy levels. Aquaculture International, 18(5): 909–919, August 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9312-0.

Guo:2010:ECT

[883] Biao Guo, Fang Wang, Shuanglin Dong, Yunwei Dong, and Xiangli Tian. The effects of cyclical temperature changes on growth and physiological status of Litopenaeus vannamei. Aquaculture International, 18(5): 921–932, August 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9314-y.

Rasmussen:2010:IGV

[884] Richard Skøtt Rasmussen and Torsten Ostenfeld. Intraspecific growth variation among rainbow trout and brook trout: impact of initial body weight and feeding level. Aquaculture International, 18(5):933– 941, August 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-009-9313-z.

El-Barbary:2010:SCM

[885] Manal I. El-Barbary. Some clinical, microbiological and molecular characteristics of Aeromonas hydrophila isolated from various naturally infected fishes. Aquaculture International, 18(5):943–954, August 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9315-x.

Garcia:2010:EMV

[886] García J. García and García B. García. Econometric model of viability/ profitability of ongrowing sharp snout sea bream (Diplodus puntazzo) in sea cages. Aquaculture International, 18(5):955–971, August 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9311-1.

Jobling:2010:MDR

[887] Malcolm Jobling. Montet D and ray RC (eds): Aquaculture microbiology and biotechnology, vol 1. Aquaculture International, 18(5): 973–974, August 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9323-x.

Garcia:2010:EEM

[888] J. García García and B. García García. Erratum to: Econometric model of viability/profitability of ongrowing sharp snout sea bream (Diplodus puntazzo) in sea cages. Aquaculture International, 18(5): 975, August 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-010-9338-3.

Li:2010:PAB

[889] Qiufen Li, Y. Zhang, D. Juck, N. Fortin, Charles W. Greer, and Qisheng Tang. Phylogenetic analysis of bacterial communities in the shrimp and sea cucumber aquaculture environment in northern China by culturing and PCR–DGGE. Aquaculture International, 18(6):977– 990, October 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-009-9316-9.

Bratland:2010:FAU

[890] Silje Bratland, Lars Helge Stien, Victoria A. Braithwaite, Jon-Erik Juell, Ole Folkedal, Jonatan Nilsson, Frode Oppedal, Jan Erik Fosseidengen, and Tore S. Kristiansen. From fright to anticipation: using aversive light stimuli to investigate reward conditioning in large groups of Atlantic salmon (Salmo salar). Aquaculture International, 18(6): 991–1001, October 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-009-9317-8.

Uno:2010:PBO

[891] Kazuaki Uno, Tidaporn Chaweepack, and Lila Ruangpan. Pharmacokinetics and bioavailability of oxytetracycline in vannamei shrimp (Penaeus vannamei) and the effect of processing on the residues in muscle and shell. Aquaculture International, 18(6):1003–1015, October 2010. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-009-9318-7.

Zadeh:2010:AGC

[892] Shahram Shakiba Zadeh, Che Roos Saad, Annie Christianus, Mohd Salleh Kamarudin, Kamaruzaman Sijam, Mariana Nor Shamsudin, and Vasantha Kumari Neela. Assessment of growth condition for a candidate probiotic, Shewanella algae, isolated from digestive system of a healthy juvenile Penaeus monodon. Aquaculture International, 18(6): 1017–1026, October 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9319-6.

Wang:2010:GTK

[893] Jinfeng Wang, Peng Jiang, Yulin Cui, Xiangyuan Deng, Fuchao Li, Jianguo Liu, and Song Qin. Genetic transformation in Kappaphycus alvarezii using micro-particle bombardment: a potential strategy for germplasm improvement. Aquaculture International, 18(6):1027– 1034, October 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-010-9320-0.

Li:2010:EVD

[894] Xiang Li, Qing hui Liu, Lin Hou, and Jie Huang. Effect of VP28 DNA vaccine on white spot syndrome virus in Litopenaeus vannamei. Aquaculture International, 18(6):1035–1044, October 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9321-z.

Chen:2010:MDA

[895] Changsheng Chen, Chaotian Xie, Dehua Ji, Yan Liang, and Lingmin Zhao. Molecular divergence and application of the ITS-5.8S rDNA and RUBISCO spacer in Porphyra haitanensis Chang et Zheng (Bangiales, Rhodophyta). Aquaculture International, 18(6):1045–1060, October 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-010-9322-y.

Nikapitiya:2010:CBE

[896] Chamilani Nikapitiya, Chulhong Oh, Mahanama De Zoysa, Ilson Whang, Do-Hyung Kang, Sun-Ryung Lee, Se-Jae Kim, and Jehee Lee. Characterization of beta-1,4-endoglucanase as a polysaccharide-degrading digestive enzyme from disk abalone, Haliotis discus discus. Aquaculture International, 18(6):1061–1078, October 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-010-9324-9.

Hu:2010:IIP

[897] Xiao-Juan Hu, Zhuo-Jia Li, Yu-Cheng Cao, Jun Zhang, Ying-Xue Gong, and Yu-Feng Yang. Isolation and identification of a phosphatesolubilizing bacterium Pantoea stewartii subsp. stewartii g6, and effects of temperature, salinity, and pH on its growth under indoor culture conditions. Aquaculture International, 18(6):1079–1091, October 2010. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9325-8.

Rigos:2010:EAI

[898] George Rigos, Alexandros Samartzis, Morgane Henry, Eleni Fountoulaki, Efthimia Cotou, John Sweetman, Simon Davies, and Ioannis Nengas. Effects of additive iron on growth, tissue distribution, haematology and immunology of gilthead sea bream, Sparus aurata. Aquaculture International, 18(6):1093–1104, October 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-010-9326-7.

El-Husseiny:2010:EAL

[899] Osama M. El-Husseiny, Galal M. Abdul-Aziz, Ashraf M. A.-S. Goda, and Ashraf Suloma. Effect of altering linoleic acid and linolenic acid dietary levels and ratios on the performance and tissue fatty acid profiles of Nile tilapia Oreochromis niloticus fry. Aquaculture International, 18(6): 1105–1119, October 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9327-6.

Seixas:2010:GFA

[900] Pedro Seixas, Ana Otero, Luísa M. P. Valente, Jorge Dias, and Manuel Rey-Méndez. Growth and fatty acid composition of Octopus vulgaris paralarvae fed with enriched Artemia or co-fed with an inert diet. Aquaculture International, 18(6):1121–1135, October 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9328-5.

Ayala:2010:ETT

[901] M. D. Ayala, A. García-Alcázar, I. Abdel, G. Ramírez-Zarzosa, and O. López-Albors. Effect of thermal treatment on muscle tissue structure and ultrastructure of wild and farmed sea bass, Dicentrarchus labrax L. Aquaculture International, 18(6):1137–1149, October 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-010-9329-4.

Ji:2010:AGA

[902] Xiang-Shan Ji, Song-Lin Chen, Jing-Feng Yang, Hong-Yu Ma, and Yun-Liang Jiang. Artificial gynogenesis and assessment of homozygosity in meiotic gynogens of spotted halibut (Verasper variegatus). Aquaculture International, 18(6):1151–1161, October 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9330-y.

Biandolino:2010:IND

[903] Francesca Biandolino, Giuseppe Portacci, and Ermelinda Prato. Influence of natural diet on growth and biochemical composition of Octopus vulgaris Cuvier, 1797. Aquaculture International, 18(6):1163– 1175, October 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-010-9331-x.

Benavente:2010:CJG

[904] Gonzalo Perez Benavente, Ingebrigt Uglem, Ronan Browne, and Carlos Marino Balsa. Culture of juvenile European lobster (Homarus gammarus L.) in submerged cages. Aquaculture International, 18(6): 1177–1189, October 2010. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9332-9.

Zhang:2010:ERG

[905] Peidong Zhang, Xiumei Zhang, Jian Li, and Tianxiang Gao. Effect of refeeding on the growth and digestive enzyme activities of Fenneropenaeus chinensis juveniles exposed to different periods of food deprivation. Aquaculture International, 18(6):1191–1203, October 2010. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9333-8.

Romana-Eguia:2010:GCG

[906] Maria Rowena R. Romana-Eguia, Minoru Ikeda, Zubaida U. Basiao, and Nobuhiko Taniguchi. Growth comparison of Asian Nile and red tilapia strains in controlled and uncontrolled farm conditions. Aquaculture International, 18(6):1205–1221, October 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9334-7.

Garcia-Garrido:2010:LCM

[907] Sandra García-Garrido, Ismael Hachero-Cruzado, Diego Garrido, Carlos Rosas, and Pedro Domingues. Lipid composition of the mantle and digestive gland of Octopus vulgaris juveniles (Cuvier, 1797) exposed to prolonged starvation. Aquaculture International, 18(6):1223– 1241, October 2010. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-010-9335-6.

Chovatiya:2011:ECS

[908] Sandeep G. Chovatiya, Sujata S. Bhatt, and Amita R. Shah. Evaluation of corn steep liquor as a supplementary feed for Labeo rohita (Ham.) fingerlings. Aquaculture International, 19(1):1–12, February 2011. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9336-5.

Chaitanawisuti:2011:EDP

[909] Nilnaj Chaitanawisuti, Sirusa Kritsanapuntu, and Wannanee Santaweesuk. Effects of dietary protein and lipid levels and protein to energy ratios on growth performance and feed utilization of hatchery-reared juvenile spotted babylon (Babylonia areolata). Aquaculture International, 19 (1):13–21, February 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9337-4.

Chaitanawisuti:2011:DFA

[910] N. Chaitanawisuti, S. Sangsawangchote, and S. Piyatiratitivorakul. Differences in fatty acid composition of egg capsules from broodstock spotted babylon, Babylonia areolata, fed a local trash fish and formulated diet under hatchery conditions. Aquaculture International, 19(1):23– 31, February 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-010-9353-4.

Wetengere:2011:SEF

[911] Kitojo Wetengere. Socio-economic factors critical for intensification of fish farming technology. A case of selected villages in Morogoro and Dar es Salaam regions, Tanzania. Aquaculture International, 19(1): 33–49, February 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9339-2.

Hunt:2011:EFR

[912] Arzu Özlüer Hunt, Ferbal Özkan, Kenan Engin, and Nazmi Tekelioğlu. The effects of freshwater rearing on the whole body and muscle tissue fatty acid profile of the European sea bass (Dicentrarchus labrax). Aquaculture International, 19(1):51–61, February 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9340-9.

Nguyen:2011:EEW

[913] Huy Quang Nguyen, Helge Reinertsen, Per-Arvid Wold, Thien Mai Tran, and Elin Kjørsvik. Effects of early weaning strategies on growth, survival and digestive enzyme activities in cobia (Rachycentron canadum L.) larvae. Aquaculture International, 19(1):63–78, February 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-010-9341-8.

Xu:2011:EDL

[914] Jian-He Xu, Jie Qin, Bin-Lun Yan, Ming Zhu, and Gang Luo. Effects of dietary lipid levels on growth performance, feed utilization and fatty acid composition of juvenile Japanese seabass (Lateolabrax japonicus) reared in seawater. Aquaculture International, 19(1):79–89, February 2011. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9342-7.

Immanuel:2011:AEM

[915] G. Immanuel, M. Sivagnanavelmurugan, and A. Palavesam. Antibacterial effect of medium-chain fatty acid: caprylic acid on gnotobiotic Artemia franciscana nauplii against shrimp pathogens Vibrio harveyi and V. parahaemolyticus. Aquaculture International, 19(1):91– 101, February 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-010-9343-6.

Karagiannis:2011:EAV

[916] Dimitrios Karagiannis, Ioannis N. Vatsos, and Panagiotis Angelidis. Effects of atrazine on the viability and the formation of byssus of the mussel Mytilus galloprovincialis. Aquaculture International, 19(1): 103–110, February 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9344-5.

Mazlum:2011:PSU

[917] Yavuz Mazlum, Erdal Yılmaz, M. A. Genç, and Ozlem Guner. A preliminary study on the use of mannan oligosaccharides (MOS) in freshwater crayfish, Astacus leptodactylus Eschscholtz, 1823 juvenile diets. Aquaculture International, 19(1):111–119, February 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9345-4.

Kortan:2011:BRC

[918] Jiří Kortan and Zdeněk Adámek. Behavioural response of carp (Cyprinus carpio, L.) pond stock upon occurrence of hunting great cormorant (Phalacrocorax carbo sinensis) flocks. Aquaculture International, 19(1): 121–129, February 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9346-3.

Garcia:2011:EIB

[919] Benjamín García García, Jesús Cerezo Valverde, Felipe Aguado-Giménez, José García García, and María D. Hernández. Effect of the interaction between body weight and temperature on growth and maximum daily food intake in sharpsnout sea bream (Diplodus puntazzo). Aquaculture International, 19(1):131–141, February 2011. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9347-2.

Ye:2011:EPF

[920] Jidan Ye, Xianghe Liu, Zijia Wang, and Kun Wang. Effect of partial fish meal replacement by soybean meal on the growth performance and biochemical indices of juvenile Japanese flounder Paralichthys olivaceus. Aquaculture International, 19(1):143–153, February 2011. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9348-1.

Reyes:2011:GSG

[921] Ofelia S. Reyes, Marietta N. Duray, Corazon B. Santiago, and Manuele Ricci. Growth and survival of grouper Epinephelus coioides (Hamilton) larvae fed free-living nematode Panagrellus redivivus at first feeding. Aquaculture International, 19(1):155–164, February 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-010-9349-0.

Rosas:2011:ETD

[922] Carlos Rosas, Ariadna Sánchez, Cristina Pascual, Josué Aguila, Teresita Maldonado, and Pedro Domingues. Effects of two dietary protein levels on energy balance and digestive capacity of Octopus maya. Aquaculture International, 19(1):165–180, February 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9350-7.

Gago:2011:CSI

[923] João Gago and Orlando J. Luís. Comparison of spawning induction techniques on Paracentrotus lividus (Echinodermata: Echinoidea) broodstock. Aquaculture International, 19(1):181–191, February 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-010-9351-6.

Feyjoo:2011:TRA

[924] Paloma Feyjoo, Rodrigo Riera, Beatriz C. Felipe, Alí Skalli, and Eduardo Almansa. Tolerance response to ammonia and nitrite in hatchlings paralarvae of Octopus vulgaris and its toxic effects on prey consumption rate and chromatophores activity. Aquaculture International, 19(1): 193–204, February 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9352-5.

Mousavi:2011:AAN

[925] Seyed Mohammad Mousavi, George Wilson, David Raftos, Seyed Saeed Mirzargar, and Reza Omidbaigi. Antibacterial activities of a new combination of essential oils against marine bacteria. Aquaculture International, 19(1):205–214, February 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-010-9354-3.

LeVay:2011:ISA

[926] Lewis Le Vay, Colin Shelley, Chaoshu Zeng, and Patrick Sorgeloos. International symposium on aquaculture, biology and management of commercially important crabs, Shanghai (China), 8–11 November 2009 (ISABMC-2009). Aquaculture International, 19(2):215–216, April 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9431-2.

Hamasaki:2011:RSP

[927] Katsuyuki Hamasaki, Yasuhiro Obata, Shigeki Dan, and Shuichi Kitada. A review of seed production and stock enhancement for commercially important portunid crabs in Japan. Aquaculture International, 19(2):217–235, April 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9387-7.

Quinitio:2011:DMC

[928] Emilia T. Quinitio, Joana Joy de la Cruz, Maria Rowena R. Eguia, Fe Dolores Parado-Estepa, Gaudioso Pates, and Celia R. Lavilla-Pitogo. Domestication of the mud crab Scylla serrata. Aquaculture International, 19(2):237–250, April 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9381-0.

Li:2011:APR

[929] X. Li, Z. Li, J. Liu, and Sena S. De Silva. Advances in precocity research of the Chinese mitten crab Eriocheir sinensis. Aquaculture International, 19(2):251–267, April 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9400-1.

Sui:2011:EDH

[930] L. Y. Sui, H. X. Sun, X. G. Wu, M. Wille, Y. X. Cheng, and P. Sorgeloos. Effect of dietary HUFA on tissue fatty acid composition and reproductive performance of Chinese mitten crab Eriocheir sinensis (H. Milne–Edwards) broodstock. Aquaculture International, 19 (2):269–282, April 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9379-7.

Qiu:2011:EHI

[931] Renjie Qiu, Yongxu Cheng, Xuxiong Huang, Xugan Wu, Xiaozhen Yang, and Rui Tong. Effect of hypoxia on immunological, physiological response, and hepatopancreatic metabolism of juvenile Chinese mitten crab Eriocheir sinensis. Aquaculture International, 19(2):283–299, April 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9390-z.

Li:2011:ELI

[932] Xiaowu Li, Zhongjie Li, Jiashou Liu, Tanglin Zhang, and Chaowen Zhang. Effects of light intensity on molting, growth, precocity, digestive enzyme activity, and chemical composition of juvenile Chinese mitten crab Eriocheir sinensis. Aquaculture International, 19(2):301– 311, April 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-010-9414-8.

Ye:2011:ENC

[933] Haihui Ye, Yong Tao, Guizhong Wang, Qiongwu Lin, Xuelei Chen, and Shaojing Li. Experimental nursery culture of the mud crab Scylla paramamosain (Estampador) in China. Aquaculture International, 19 (2):313–321, April 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9399-3.

Dan:2011:ESD

[934] Shigeki Dan and Katsuyuki Hamasaki. Effects of salinity and dietary n − 3 highly unsaturated fatty acids on the survival, development, and morphogenesis of the larvae of laboratory-reared mud crab Scylla serrata (Decapoda, Portunidae). Aquaculture International, 19 (2):323–338, April 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9374-z.

Liao:2011:EAN

[935] Yong Yang Liao, Hong Hu Wang, and Zhi Gan Lin. Effect of ammonia and nitrite on vigour, survival rate, moulting rate of the blue swimming crab Portunus pelagicus zoea. Aquaculture International, 19 (2):339–350, April 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9398-4.

Wan:2011:ICV

[936] Xihe Wan, Hui Shen, Libao Wang, and Yongxu Cheng. Isolation and characterization of Vibrio metschnikovii causing infection in farmed Portunus trituberculatus in China. Aquaculture International, 19(2):351– 359, April 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-011-9422-3.

Parkes:2011:PDB

[937] Lee Parkes, Emilia T. Quinitio, and Lewis Le Vay. Phenotypic differences between hatchery-reared and wild mud crabs, Scylla serrata, and the effects of conditioning. Aquaculture International, 19(2):361– 380, April 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-010-9372-1.

Galley:2011:DLC

[938] Thomas H. Galley, Benjamin C. Green, Lloyd Watkins, and Lewis Le Vay. Development of larval culture techniques for the shore crab, Carcinus maenas (L.). Aquaculture International, 19(2):381–394, April 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9380-1.

Zhong:2011:ECG

[939] Guofang Zhong, Xueming Hua, Kun Yuan, and Hongqi Zhou. Effect of CGM on growth performance and digestibility in puffer (Takifugu fasciatus). Aquaculture International, 19(3):395–403, June 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-010-9355-2.

Yang:2011:ERF

[940] Yu-Hong Yang, Yu-Yu Wang, Yang Lu, and Qing-Zhang Li. Effect of replacing fish meal with soybean meal on growth, feed utilization and nitrogen and phosphorus excretion on rainbow trout (Oncorhynchus mykiss). Aquaculture International, 19(3):405–419, June 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9359-y.

Sharma:2011:DBV

[941] S. R. Krupesha Sharma, K. M. Shankar, M. L. Sathyanarayana, Raj Reddy Patil, H. D. Narayana Swamy, and Suguna Rao. Development of biofilm of Vibrio alginolyticus for oral immunostimulation of shrimp. Aquaculture International, 19(3):421–430, June 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9357-0.

Carter:2011:LFD

[942] Chris G. Carter and M. Sajjadi. Low fishmeal diets for Atlantic salmon, Salmo salar L., using soy protein concentrate treated with graded levels of phytase. Aquaculture International, 19(3):431–444, June 2011. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9358-z.

Martinez:2011:PCH

[943] R. Martínez, R. Sántos, A. Álvarez, G. Cuzón, L. Arena, M. Mascaró, C. Pascual, and C. Rosas. Partial characterization of hepatopancreatic and extracellular digestive proteinases of wild and cultivated Octopus maya. Aquaculture International, 19(3):445–457, June 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-010-9360-5.

Hu:2011:EDM

[944] Yi Hu, Beiping Tan, Kangsen Mai, Qinghui Ai, Lu Zhang, and Shixuan Zheng. Effects of dietary menhaden oil, soybean oil and soybean lecithin oil at different ratios on growth, body composition and blood chemistry of juvenile Litopenaeus vannamei. Aquaculture International, 19(3):459–473, June 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9361-4.

Altintzoglou:2011:AHI

[945] Themistoklis Altintzoglou, Filiep Vanhonacker, Wim Verbeke, and Joop Luten. Association of health involvement and attitudes towards eating fish on farmed and wild fish consumption in Belgium, Norway and Spain. Aquaculture International, 19(3):475–488, June 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9363-2.

Chaitanawisuti:2011:EDS

[946] Nilnaj Chaitanawisuti, Chatchaleeya Choeychom, and Somkiat Piyatiratitivorakul. Effect of dietary supplementation of brewers yeast and nucleotide singularly on growth, survival and vibriosis resistance on juveniles of the gastropod spotted babylon (Babylonia areolata). Aquaculture International, 19(3):489–496, June 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-010-9364-1.

Baer:2011:AGT

[947] A. Baer, C. Schulz, I. Traulsen, and J. Krieter. Analysing the growth of turbot (Psetta maxima) in a commercial recirculation system with the use of three different growth models. Aquaculture International, 19 (3):497–511, June 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9365-0.

Kujawa:2011:ASC

[948] Roman Kujawa, Dariusz Kucharczyk, Andrzej Mamcarz, Daniel Żarski, and Katarzyna Targońska. Artificial spawning of common tench Tinca tinca (Linnaeus, 1758), obtained from wild and domestic stocks. Aquaculture International, 19(3):513–521, June 2011. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9366-z.

Garces:2011:ESS

[949] J. P. Garcês and J. Pereira. Effect of salinity on survival and growth of Marphysa sanguinea Montagu (1813) juveniles. Aquaculture International, 19(3):523–530, June 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-010-9368-x.

Vaseeharan:2011:AAA

[950] B. Vaseeharan, G. Sai Prasad, P. Ramasamy, and G. Brennan. Antibacterial activity of Allium sativum against multidrug-resistant Vibrio harveyi isolated from black gill–diseased Fenneropenaeus indicus. Aquaculture International, 19(3):531–539, June 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-010-9369-9.

Pellizzato:2011:RTP

[951] Michele Pellizzato, Thomas Galvan, Raoul Lazzarini, and Pierpaolo Penzo. Recruitment of Tapes philippinarum in the Venice Lagoon (Italy) during 2002–2007. Aquaculture International, 19(3):541–554, June 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9370-3.

Son:2011:PEE

[952] Vu Nam Son, Nguyen Thanh Phuong, Tran Ngoc Hai, and Amararatne Yakupitiyage. Production and economic efficiencies of intensive black tiger prawn (Penaeus monodon) culture during different cropping seasons in the Mekong delta, Vietnam. Aquaculture International, 19 (3):555–566, June 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9371-2.

Ahmed:2011:EDN

[953] Imtiaz Ahmed. Effect of dietary niacin on growth and body composition of two Indian major carps rohu, Labeo rohita, and mrigal, Cirrhinus mrigala (Hamilton), fingerlings based on dose–response study. Aquaculture International, 19(3):567–584, June 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-010-9373-0.

Gwak:2011:GVH

[954] W. S. Gwak and K. Nakayama. Genetic variation of hatchery and wild stocks of the pearl oyster Pinctada fucata martensii (Dunker, 1872), assessed by mitochondrial DNA analysis. Aquaculture International, 19 (3):585–591, June 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9385-9.

Jobling:2011:BRMa

[955] Malcolm Jobling. Book review: M. Speight and P. Henderson: Marine ecology: concepts and applications. Aquaculture International, 19 (3):593–594, June 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9362-3.

Jobling:2011:BRGa

[956] Malcolm Jobling. Book review: G. M. Turchini, W.-K. Ng and D. R. Tocher (eds): Fish oil replacement and alternative lipid sources in aquaculture feeds. Aquaculture International, 19(3):595–596, June 2011. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9367-y.

Jobling:2011:BRDa

[957] Malcolm Jobling. Book review: A. D. McIntyre (ed): Life in the world’s oceans — diversity, distribution and abundance. Aquaculture International, 19(3):597–598, June 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-010-9396-6.

Kiatmetha:2011:ESM

[958] Pauline Kiatmetha, Wipawan Siangdang, Boosya Bunnag, Saengchan Senapin, and Boonsirm Withyachumnarnkul. Enhancement of survival and metamorphosis rates of Penaeus monodon larvae by feeding with the diatom Thalassiosira weissflogii. Aquaculture International, 19(4): 599–609, August 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9375-y.

Sotoudeh:2011:GRB

[959] Ebrahim Sotoudeh, Abdolmohammad Abedian Kenari, and Mehran Habibi Rezaei. Growth response, body composition and fatty acid profile of Caspian brown trout (Salmo trutta Caspius) juvenile fed diets containing different levels of soybean phosphatidylcholine. Aquaculture International, 19(4):611–623, August 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-010-9376-x.

Kristiansen:2011:VRL

[960] Tore S. Kristiansen, Lars H. Stien, Jan Erik Fosseidengen, Espen Strand, and Jon-Erik Juell. Voluntary responses and limits of tolerance to pressure reduction and swimbladder expansion in farmed Atlantic cod. Aquaculture International, 19(4):625–636, August 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9377-9.

Saadiah:2011:RFC

[961] I. Saadiah, A. M. Abol-Munafi, and C. M. Che Utama. Replacement of fishmeal in cobia (Rachycentron canadum) diets using poultry byproduct meal. Aquaculture International, 19(4):637–648, August 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9378-8.

Huang:2011:SRP

[962] Cheng-Ting Huang, Sha Miao, Fan-Hua Nan, and Shi-Mu Jung. Study on regional production and economy of cobia Rachycentron canadum commercial cage culture. Aquaculture International, 19(4):649–664, August 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-010-9391-y.

Rahman:2011:DEE

[963] Sanzidur Rahman, Basanta K. Barmon, and Nesar Ahmed. Diversification economies and efficiencies in a ‘blue-green revolution’ combination: a case study of prawn-carp-rice farming in the ‘gher ’ system in Bangladesh. Aquaculture International, 19(4):665–682, August 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9382-z.

Ouraji:2011:DOT

[964] Hossein Ouraji, Khosro Jani Khalili, Ghaffar Ebrahimi, and Seyed Ali Jafarpour. Determination of the optimum transfer time of kutum (Rutilus frisii kutum) larvae from live food to artificial dry feed. Aquaculture International, 19(4):683–691, August 2011. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9383-y.

Senapin:2011:DPP

[965] Saengchan Senapin, Kornsunee Phiwsaiya, Pauline Kiatmetha, and Boonsirm Withyachumnarnkul. Development of primers and a procedure for specific identification of the diatom Thalassiosira weissflogii. Aquaculture International, 19(4):693–704, August 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9384-x.

Esmaeili:2011:PSG

[966] Abdoulkarim Esmaeili and Mohammad Hassan Tarazkar. Prediction of shrimp growth using an artificial neural network and regression models. Aquaculture International, 19(4):705–713, August 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-010-9386-8.

Tettelbach:2011:UHD

[967] Stephen T. Tettelbach, Debra Barnes, John Aldred, Gregg Rivara, Dennis Bonal, Andrew Weinstock, Chelsea Fitzsimons-Diaz, Josh Thiel, M. Chase Cammarota, Adam Stark, Katherine Wejnert, Richard Ames, and John Carroll. Utility of high-density plantings in bay scallop, Argopecten irradians irradians, restoration. Aquaculture International, 19 (4):715–739, August 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9388-6.

Siikavuopio:2011:LTE

[968] Sten I. Siikavuopio and Philip J. James. Long-term effects of stocking density on somatic growth, gonad index and survival of the juvenile sea urchin, Strongylocentrotus droebachiensis. Aquaculture International, 19 (4):741–747, August 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9389-5.

Sukumaran:2011:IDP

[969] Krishna Sukumaran, Asim K. Pal, Narottam P. Sahu, Shrinivas Jahageerdar, Deepak Khambadakone, and Dipesh Debnath. Influence of dietary phosphorus on mineralization and morphology of Indian major carp, catla (Catla catla) fingerlings. Aquaculture International, 19(4): 749–763, August 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9392-x.

Costa:2011:ANI

[970] Corrado Costa, Stefano D’Andrea, Riccardo Russo, Francesca Antonucci, Federico Pallottino, and Paolo Menesatti. Application of non-invasive techniques to differentiate sea bass (Dicentrarchus labrax, L. 1758) quality cultured under different conditions. Aquaculture International, 19 (4):765–778, August 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9393-9.

Sun:2011:INI

[971] Weiming Sun, Shuanglin Dong, Zilin Jie, Xidan Zhao, Hanzhen Zhang, and Jianxing Li. The impact of net-isolated polyculture of tilapia (Oreochromis niloticus) on plankton community in saline–alkaline pond of shrimp (Penaeus vannamei). Aquaculture International, 19(4):779– 788, August 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-010-9394-8.

Sreedevi:2011:BAP

[972] P. R. Sreedevi and V. Ramasubramanian. Biocontrol of ammonia pollution in the rearing water of fish by inducing a heterotrophic bacterialbased food chain in the medium. Aquaculture International, 19(4): 789–796, August 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9395-7.

Soto-Zarazua:2011:ARA

[973] Genaro M. Soto-Zarazúa, Rocío Peniche-Vera, Enrique Rico-García, Manuel Toledano-Ayala, Rosalía Ocampo-Velázquez, and Gilberto Herrera-Ruiz. An automated recirculation aquaculture system based on fuzzy logic control for aquaculture production of tilapia (Oreochromis niloticus). Aquaculture International, 19(4):797–808, August 2011. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9397-5.

Jobling:2011:BRMb

[974] Malcolm Jobling. Book review: M. A. Pavlidis and C. C. Mylonas (eds): Sparidae: biology and aquaculture of gilthead sea bream and other species. Aquaculture International, 19(4):809–810, August 2011. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9455-7.

Kalinowski:2011:EDA

[975] C. T. Kalinowski, L. E. Robaina, and M. S. Izquierdo. Effect of dietary astaxanthin on the growth performance, lipid composition and post-mortem skin colouration of red porgy Pagrus pagrus. Aquaculture International, 19(5):811–823, October 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9401-0.

Larsen:2011:STR

[976] Sondre V. Larsen, Albert K. Imsland, Petter Lohne, Karin Pittman, and Atle Foss. Stepwise temperature regulation and its effect on growth, feeding and muscle growth patterns of juvenile Atlantic halibut (Hippoglossus hippoglossus L.). Aquaculture International, 19(5): 825–837, October 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9402-z.

Kaggwa:2011:STR

[977] Mary N. Kaggwa, David M. Liti, and Michael Schagerl. Small tropical reservoirs and fish cage culture: a pilot study conducted in Machakos district, Kenya. Aquaculture International, 19(5):839–853, October 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-010-9403-y.

Schenone:2011:FFW

[978] Nahuel Francisco Schenone, Lenka Vackova, and Alicia Fernández Cirelli. Fish-farming water quality and environmental concerns in Argentina: a regional approach. Aquaculture International, 19(5):855–863, October 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-010-9404-x.

Chaitanawisuti:2011:GFE

[979] Nilnaj Chaitanawisuti, Sirusa Kritsanapun, and Wannanee Santhaweesuk. Growth, food efficiency, and biochemical composition of juvenile spotted babylon Babylonia areolata (Link) fed on conventional trash fish and a formulated moist diet. Aquaculture International, 19 (5):865–872, October 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9406-8.

Lai:2011:CCG

[980] Xiao Fang Lai, Huan Gao, Jie Kong, Qing Yin Wang, Wei Ji Wang, and Xian Hong Meng. Cloning and characterization of the glutamine synthetase gene from Chinese shrimp Fenneropenaeus chinensis. Aquaculture International, 19(5):873–889, October 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9407-7.

Emerenciano:2011:EBT

[981] Maurício Emerenciano, Eduardo L. C. Ballester, Ronaldo O. Cavalli, and Wilson Wasielesky. Effect of biofloc technology (BFT) on the early postlarval stage of pink shrimp Farfantepenaeus paulensis: growth performance, floc composition and salinity stress tolerance. Aquaculture International, 19(5):891–901, October 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-010-9408-6.

Ouraji:2011:GSF

[982] H. Ouraji, A. M. Abedian Kenari, B. Shabanpour, A. Shabani, M. Sodagar, S. A. Jafarpour, and G. H. Ebrahimi. Growth, survival, and fatty acid composition of Indian white shrimp Fenneropenaeus indicus (Milne Edwards) fed diets containing different levels of vitamin E and lipid. Aquaculture International, 19(5):903–916, October 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-010-9409-5.

Nogales-Merida:2011:GPH

[983] Silvia Nogales-Mérida, Ana Tomás-Vidal, Miguel Jover Cerdá, and Silvia Martínez-Llorens. Growth performance, histological alterations and fatty acid profile in muscle and liver of sharp snout sea bream (Diplodus puntazzo) with partial replacement of fish oil by pork fat. Aquaculture International, 19(5):917–929, October 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-010-9410-z.

Lee:2011:BNF

[984] Jae Seong Lee, Yong-Tae Kim, Kyung-Hoon Shin, Jung-Ho Hyun, and Sook-Yang Kim. Benthic nutrient fluxes at longline sea squirt and oyster aquaculture farms and their role in coastal ecosystems. Aquaculture International, 19(5):931–944, October 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-010-9411-y.

Wang:2011:DPA

[985] Qing-Kui Wang, Cheng-Xun Chen, Yong-Jun Guo, Hai-Yun Zhao, Jing-Feng Sun, Shen Ma, and Ke-Zhi Xing. Dietary polysaccharide from Angelica sinensis enhanced cellular defence responses and disease resistance of grouper Epinephelus malabaricus. Aquaculture International, 19 (5):945–956, October 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9412-x.

Hayden:2011:EWV

[986] Barbara J. Hayden and Chris M. C. Woods. Effect of water velocity on growth and retention of cultured Greenshell^{T M} mussel spat, Perna canaliculus (Gmelin, 1791). Aquaculture International, 19(5): 957–971, October 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-010-9413-9.

Gopalakannan:2011:IAP

[987] Ayyaru Gopalakannan and Venkatesan Arul. Inhibitory activity of probiotic Enterococcus faecium MC13 against Aeromonas hydrophila confers protection against hemorrhagic septicemia in common carp Cyprinus carpio. Aquaculture International, 19(5):973–985, October 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-011-9415-2.

Jiwyam:2011:ESD

[988] Wirat Jiwyam. The effect of stocking density on yield, growth, and survival of Asian river catfish (Pangasius bocourti Sauvage, 1880) cultured in cages. Aquaculture International, 19(5):987–997, October 2011. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9416-1.

Conejeros:2011:HID

[989] Pablo A. Conejeros, Carlos Calderón, Daniela Gómez, Luis Nilo, and Sergio H. Marshall. High immune diversity in farmed Atlantic salmon (Salmo salar L.). Aquaculture International, 19(5):999–1005, October 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-011-9417-0.

Xu:2011:EDS

[990] Shude Xu, Liang Zhang, Qingyang Wu, Xuebing Liu, Shuqi Wang, Cuihong You, and Yuanyou Li. Evaluation of dried seaweed Gracilaria lemaneiformis as an ingredient in diets for teleost fish Siganus canaliculatus. Aquaculture International, 19(5):1007–1018, October 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-011-9418-z.

Jobling:2011:BRGb

[991] Malcolm Jobling. Book review: G. J. Holt (ed): Larval fish nutrition. Aquaculture International, 19(5):1019–1020, October 2011. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9458-4.

Jobling:2011:BRF

[992] Malcolm Jobling. Book review: F. Asche and T. Bjørndal: The economics of salmon aquaculture, 2nd edn. Aquaculture International, 19(5): 1021–1023, October 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9459-3.

Karami:2011:OTP

[993] Ali Karami, Annie Christianus, Hadi Zokaeifar, Khairul Zamri Saad, Fahmi T. J. Imraan, Shahram Shakibazadeh, Hossien Negarestan, and Simon C. Courtenay. Ovaprim treatment promotes oocyte development and milt fertilization rate in diploid and triploid African catfish (Clarias gariepinus). Aquaculture International, 19(6):1025–1034, December 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-011-9419-y.

Imanpoor:2011:RHU

[994] Mohamad Reza Imanpoor, Masoud Asghari, and Reza Asadi. Requirements for n − 3 highly unsaturated fatty acids in feeding juvenile Iranian sturgeon (Acipenser persicus) and its effects on growth, carcass quality, and fatty acid composition. Aquaculture International, 19(6): 1035–1046, December 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9420-5.

Andrade:2011:DMA

[995] Lorenzo I. Andrade, Daniel A. López, and Boris A. López. Dynamic models applied to giant barnacle culture. Aquaculture International, 19(6): 1047–1060, December 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9421-4.

Deng:2011:EDP

[996] Jun-Ming Deng, Kang-Sen Mai, Qing-Hui Ai, Wen-Bing Zhang, Wei Xu, and Zhi-Guo Liufu. Effects of dietary protein sources on feed intake, growth and plasma thyroid hormones levels of Japanese flounder (Paralichthys olivaceus). Aquaculture International, 19(6):1061– 1074, December 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9423-2.

Worrall:2011:ECP

[997] K. L. Worrall, C. G. Carter, R. J. Wilkinson, and M. J. R. Porter. The effects of continuous photoperiod (24L:0D) on growth of juvenile barramundi (Lates calcarifer). Aquaculture International, 19(6):1075– 1082, December 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9425-0.

Hernandez:2011:APW

[998] José Juan Carbajal Hernández, Luis P. Sánchez Fernández, and Oleksiy Pogrebnyak. Assessment and prediction of water quality in shrimp culture using signal processing techniques. Aquaculture International, 19(6): 1083–1104, December 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9426-z.

Zarski:2011:ISD

[999] Daniel Żarski, Katarzyna Targońska, Sławomir Krejszeff, Maciej Kwiatkowski, Krzysztof Kupren, and Dariusz Kucharczyk. Influence of stocking density and type of feed on the rearing of crucian carp, Carassius carassius (L.), larvae under controlled conditions. Aquaculture International, 19(6):1105–1117, December 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-011-9427-y.

Chakraborty:2011:IAR

[1000] Suman Bhusan Chakraborty, Samir Banerjee, and Saibal Chatterjee. Increased androgen receptor expression in muscle tissue contributing to growth increase in androgen-treated Nile tilapia. Aquaculture International, 19(6):1119–1137, December 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-011-9428-x.

Chaitanawisuti:2011:PSG

[1001] Nilnaj Chaitanawisuti, Wannanee Santhaweesuk, and Sirusa Kritsanapuntu. Performance of the seaweeds Gracilaria salicornia and Caulerpa lentillifera as biofilters in a hatchery scale recirculating aquaculture system for juvenile spotted babylons (Babylonia areolata). Aquaculture International, 19(6):1139–1150, December 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-011-9429-9.

Kloskowski:2011:DEA

[1002] Janusz Kloskowski. Differential effects of age-structured common carp (Cyprinus carpio) stocks on pond invertebrate communities: implications for recreational and wildlife use of farm ponds. Aquaculture International, 19(6):1151–1164, December 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-011-9435-y.

Wang:2011:EHS

[1003] Qing-Lin Wang, Yun-Wei Dong, Shuang-Lin Dong, and Fang Wang. Effects of heat-shock selection during pelagic stages on thermal sensitivity of juvenile sea cucumber, Apostichopus japonicus Selenka. Aquaculture International, 19(6):1165–1175, December 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-011-9436-x.

Garcia:2011:EMV

[1004] José García García and Benjamín García García. Econometric model of viability/profitability of octopus (Octopus vulgaris) ongrowing in sea cages. Aquaculture International, 19(6):1177–1191, December 2011. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9432-1.

Chowdhury:2011:ECS

[1005] Md. Arif Chowdhury, Yahya Khairun, Md. Salequzzaman, and Md. Mizanur Rahman. Effect of combined shrimp and rice farming on water and soil quality in Bangladesh. Aquaculture International, 19(6):1193– 1206, December 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9433-0.

Jancula:2011:OFR

[1006] Daniel Jančula, Eliška Maršálková, and Blahoslav Maršálek. Organic flocculants for the removal of phytoplankton biomass. Aquaculture International, 19(6):1207–1216, December 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-011-9434-z.

Sales:2011:FFF

[1007] J. Sales. First feeding of freshwater fish larvae with live feed versus compound diets: a meta-analysis. Aquaculture International, 19(6): 1217–1228, December 2011. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9424-1.

Jobling:2011:BRDb

[1008] Malcolm Jobling. Book review: D. Whitmarsh: Economic management of marine living resources — a practical introduction. Earthscan, London, 2011, xv + 171 pp., £24.99 (Paperback), ISBN 978-1-84971-259-0. Aquaculture International, 19(6):1229–1230, December 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-011-9444-x.

Jobling:2011:BRC

[1009] Malcolm Jobling. Book review: C. E. Nash: The history of aquaculture. Aquaculture International, 19(6):1231–1232, December 2011. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-011-9465-5.

Shankar:2012:ENG

[1010] R. Shankar, H. Shivananda Murthy, H. R. Sujatha, E. G. Jayaraj, C. S. Tejpal, and V. S. Chinthamani. Effect of nucleotide on growth, immune responses and resistance of Macrobrachium rosenbergii (De Man) to Macrobrachium rosenbergii nodavirus (MrNV) and extra small virus (XSV) and Aeromonas hydrophila infection. Aquaculture International, 20(1):1–12, February 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9430-3.

Corteel:2012:MCL

[1011] M. Corteel, J. J. Dantas-Lima, M. Wille, V. Alday-Sanz, M. B. Pensaert, P. Sorgeloos, and H. J. Nauwynck. Moult cycle of laboratory-raised Penaeus (Litopenaeus) vannamei and P. monodon. Aquaculture International, 20(1):13–18, February 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-011-9437-9.

Xie:2012:EPR

[1012] Quan-Sen Xie, Zhen-Cai Yang, Jun-Wei Li, and Yu-Juan Li. Effect of protein restriction with subsequent re-alimentation on compensatory growth of juvenile soft-shelled turtles (Pelodiscus sinensis). Aquaculture International, 20(1):19–27, February 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-011-9438-8.

Flood:2012:ECF

[1013] Matthew J. Flood, G. John Purser, and Chris G. Carter. The effects of changing feeding frequency simultaneously with seawater transfer in Atlantic salmon Salmo salar L. smolt. Aquaculture International, 20 (1):29–40, February 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9439-7.

Khoeyi:2012:ELI

[1014] Zahra Amini Khoeyi, Jafar Seyfabadi, and Zohreh Ramezanpour. Effect of light intensity and photoperiod on biomass and fatty acid composition of the microalgae, Chlorella vulgaris. Aquaculture International, 20 (1):41–49, February 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9440-1.

Ahmed:2012:PCC

[1015] Nesar Ahmed, James A. Young, Madan M. Dey, and James F. Muir. From production to consumption: a case study of tilapia marketing systems in Bangladesh. Aquaculture International, 20(1):51–70, February 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-011-9441-0.

Valverde:2012:LCM

[1016] Jesús Cerezo Valverde, María D. Hernández, Sandra García-Garrido, Carmen Rodríguez, Juán Estefanell, Joan I. Gairín, Carlos J. Rodríguez, Ana Tomás, and Benjamín García García. Lipid classes from marine species and meals intended for cephalopod feeding. Aquaculture International, 20(1):71–89, February 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-011-9442-z.

Chatzifotis:2012:EPL

[1017] Stavros Chatzifotis, Maria Panagiotidou, and Pascal Divanach. Effect of protein and lipid dietary levels on the growth of juvenile meagre (Argyrosomus regius). Aquaculture International, 20(1):91–98, February 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9443-y.

Booth:2012:IPM

[1018] M. A. Booth, G. L. Allan, and A. J. Anderson. Influence of poultry meal, meat meal or soybean meal inclusion on weight gain and production characteristics of Australian snapper Pagrus auratus. Aquaculture International, 20(1):99–115, February 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-011-9445-9.

Oujifard:2012:FMR

[1019] Amin Oujifard, Jafar Seyfabadi, Abdolmohammad Abedian Kenari, and Masood Rezaei. Fish meal replacement with rice protein concentrate in a practical diet for the Pacific white shrimp, Litopenaeus vannamei Boone, 1931. Aquaculture International, 20(1):117–129, February 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9446-8.

Akpinar:2012:EDL

[1020] Zafer Akpınar, Hüseyin Sevgili, Abdullah Demir, Talip Özgen, Yılmaz Emre, and O. Tufan Eroldoğan. Effects of dietary lipid levels on growth, nutrient utilization, and nitrogen and carbon balances in shi drum (Umbrina cirrosa L.). Aquaculture International, 20(1):131–143, February 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-011-9447-7.

Zhao:2012:ELT

[1021] Hong-Xia Zhao, Jun-Ming Cao, An-Li Wang, Zhen-Yu Du, Chao-Xia Ye, Yan-Hua Huang, Han-Bing Lan, Ting-Ting Zhou, and Guo-Li Li. Effect of long-term administration of dietary β-1,3-glucan on growth, physiological, and immune responses in Litopenaeus vannamei (Boone, 1931). Aquaculture International, 20(1):145–158, February 2012. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9448-6.

Welker:2012:SGN

[1022] Thomas L. Welker, Chhorn Lim, Mediha Yildirim-Aksoy, and Phillip H. Klesius. Susceptibility of Nile tilapia (Oreochromis niloticus) fed with dietary sodium chloride to nitrite toxicity. Aquaculture International, 20 (1):159–176, February 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9449-5.

Hou:2012:ECC

[1023] Chunqiang Hou, Fang Wang, Shuanglin Dong, Yujie Zhu, and Ting Yu. Effects of constant ca^{2+} concentration in salinity fluctuations on growth and energy budget of juvenile Litopenaeus vannamei. Aquaculture International, 20(1):177–188, February 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-011-9450-z.

Magnesen:2012:PPR

[1024] Thorolf Magnesen and Kirsten Jayne Redmond. Potential predation rates by the sea stars Asterias rubens and Marthasterias glacialis, on juvenile scallops, Pecten maximus, ready for sea ranching. Aquaculture International, 20(1):189–199, February 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-011-9451-y.

Tziouveli:2012:ELE

[1025] V. Tziouveli, M. Hall, and G. G. Smith. Evaluation of lipid-enriched Artemia on the reproductive performance of the white-striped cleaner shrimp, Lysmata amboinensis. Aquaculture International, 20(2):201– 211, April 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-011-9496-y.

Overturf:2012:MAD

[1026] Ken Overturf, Roger L. Vallejo, Yniv Palti, Fredrick T. Barrows, and James E. Parsons. Microarray analysis of differential utilization of plantbased diets by rainbow trout. Aquaculture International, 20(2):213– 232, April 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-011-9490-4.

Panda:2012:IES

[1027] Nirupama Panda, A. S. Mahapatra, and Rasmiprava Samal. Impact evaluation of SGSY on socio-economic development of women in aquaculture in Eastern Hills of Orissa. Aquaculture International, 20(2): 233–247, April 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9452-x.

Ghomi:2012:EED

[1028] Mohammad Reza Ghomi, Ramezan Shahriari, Hamid Faghani Langroudi, Mehdi Nikoo, and Eric von Elert. Effects of exogenous dietary enzyme on growth, body composition, and fatty acid profiles of cultured great sturgeon Huso huso fingerlings. Aquaculture International, 20(2):249–254, April 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9453-9.

Wassef:2012:EDV

[1029] Elham A. Wassef, Olfat M. Wahbi, and Shaymaa H. Shalaby. Effects of dietary vegetable oils on liver and gonad fatty acid metabolism and gonad maturation in gilthead seabream (Sparus aurata) males and females. Aquaculture International, 20(2):255–281, April 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-011-9454-8.

Tian:2012:EDD

[1030] Li Xia Tian, Yong Jian Liu, Hui Jun Yang, Gui Ying Liang, and Jin Niu. Effects of different dietary wheat starch levels on growth, feed efficiency and digestibility in grass carp (Ctenopharyngodon idella). Aquaculture International, 20(2):283–293, April 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-011-9456-6.

Oronti:2012:AFS

[1031] Annabelle Oronti, Andy J. Danylchuk, Christina E. Elmore, Rocco Auriemma, and Giusto Pesle. Assessing the feasibility of sponge aquaculture as a sustainable industry in The Bahamas. Aquaculture International, 20(2):295–303, April 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9457-5.

LeMoullac:2012:GGD

[1032] Gilles Le Moullac, Jérôme Tiapari, Hinano Teissier, Elodie Martinez, and Jean-Claude Cochard. Growth and gonad development of the tropical black-lip pearl oyster, Pinctada margaritifera (L.), in the Gambier archipelago (French Polynesia). Aquaculture International, 20(2):305– 315, April 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-011-9460-x.

Abdel-Tawwab:2012:IED

[1033] Mohsen Abdel-Tawwab. Interactive effects of dietary protein and live bakery yeast, Saccharomyces cerevisiae on growth performance of Nile tilapia, Oreochromis niloticus (L.) fry and their challenge against Aeromonas hydrophila infection. Aquaculture International, 20(2):317– 331, April 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-011-9462-8.

Perez:2012:BES

[1034] Eduardo P. Pérez, Andrea Araya, Marcelo Araneda, and Claudio Zúñiga. Bioeconomic effect from the size selection in red abalone intensive culture Haliotis rufescens as a production strategy. Aquaculture International, 20(2):333–345, April 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9461-9.

Yang:2012:EEA

[1035] Xiaozhen Yang, Jinfeng Wang, Liulan Zhao, Peng Fan, Xugan Wu, Yongxu Cheng, and Chaoshu Zeng. Effects of elevated ambient histamine level on survival, growth, sexual maturity and tissue histamine accumulation of the mysis Neomysis awatschensis and Neomysis japonica Nakazawa. Aquaculture International, 20(2):347–356, April 2012. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9463-7.

Sookying:2012:USP

[1036] Daranee Sookying and D. Allen Davis. Use of soy protein concentrate in practical diets for Pacific white shrimp (Litopenaeus vannamei) reared under field conditions. Aquaculture International, 20(2):357– 371, April 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-011-9464-6.

Barbieri:2012:ATA

[1037] Edison Barbieri and Sonia Assami Doi. Acute toxicity of ammonia on juvenile cobia (Rachycentron canadum, Linnaeus, 1766) according to the salinity. Aquaculture International, 20(2):373–382, April 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-011-9467-3.

Zehra:2012:DPR

[1038] Seemab Zehra and Mukhtar A. Khan. Dietary protein requirement for fingerling Channa punctatus (Bloch), based on growth, feed conversion, protein retention and biochemical composition. Aquaculture International, 20(2):383–395, April 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-011-9470-8.

Jobling:2012:BRM

[1039] Malcolm Jobling. Book review: M. B. New, W. C. Valenti, J. Tidwell, L. R. D’Abramo and M. N. Kutty (eds): Freshwater prawns: biology and farming. Aquaculture International, 20(2):397–398, April 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-011-9469-1.

Koprucu:2012:ELE

[1040] Kenan Köprücü and Ebru Sertel. The effects of less-expensive plant protein sources replaced with soybean meal in the juvenile diet of grass carp (Ctenopharyngodon idella): growth, nutrient utilization and body composition. Aquaculture International, 20(3):399–412, June 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-011-9471-7.

Ak:2012:EOF

[1041] İlknur Ak. Effect of an organic fertilizer on growth of blue-green alga Spirulina platensis. Aquaculture International, 20(3):413–422, June 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9473-5.

Sykes:2012:EES

[1042] Peter J. Sykes, Carol A. McClure, Debbie J. Martin-Robichaud, Charles G. Caraguel, and K. Larry Hammell. Economic evaluation of a stratified transport method for Atlantic halibut (Hippoglossus hippoglossus) juveniles. Aquaculture International, 20(3):423–430, June 2012. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9474-4.

Fu:2012:ELL

[1043] Longlong Fu, Zhiyi Bai, Wu Jin, Genfang Zhang, and Jiale Li. Evaluation of laser labeling on growth and survival of the freshwater pearl mussel, Hyriopsis cumingii. Aquaculture International, 20(3):431–441, June 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-011-9475-3.

Slawski:2012:TFM

[1044] H. Slawski, H. Adem, R.-P. Tressel, K. Wysujack, U. Koops, Y. Kotzamanis, S. Wuertz, and C. Schulz. Total fish meal replacement with rapeseed protein concentrate in diets fed to rainbow trout (Oncorhynchus mykiss Walbaum). Aquaculture International, 20(3):443–453, June 2012. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9476-2.

Mahboob:2012:HDG

[1045] Shahid Mahboob, Abdul Rauf, Muhammed Ashraf, Tayyaba Sultana, Salma Sultana, Farhat Jabeen, Muhammad Ibrahim Rajoka, H. F. Alkahem Al-Balawi, and Khalid A. Al-Ghanim. High-density growth and crude protein productivity of a thermotolerant Chlorella vulgaris: production kinetics and thermodynamics. Aquaculture International, 20 (3):455–466, June 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9477-1.

Koedijk:2012:LRE

[1046] Roland Koedijk, Albert Kjartansson Imsland, Arild Folkvord, Sigurd Olav Stefansson, Thor Magne Jonassen, and Atle Foss. Larval rearing environment influences the physiological adaptation in juvenile Atlantic cod, Gadus morhua. Aquaculture International, 20(3):467– 479, June 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-011-9478-0.

Guroy:2012:EFO

[1047] Derya Güroy, Betül Güroy, Daniel Lee Merrifield, Ahmet Adem Tekinay, Simon John Davies, and İzzet Şahin. Effects of fish oil and partial fish meal substitution with oilseed oils and meals on growth performance, nutrient utilization and health of the rainbow trout Oncorhynchus mykiss. Aquaculture International, 20(3):481–497, June 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-011-9479-z.

Keysami:2012:PAB

[1048] Mehran Avakh Keysami, Masoumeh Mohammadpour, and Che Roos Saad. Probiotic activity of Bacillus subtilis in juvenile freshwater prawn, Macrobrachium rosenbergii (de Man) at different methods of administration to the feed. Aquaculture International, 20(3):499–511, June 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9481-5.

Bae:2012:EDP

[1049] Jun-Young Bae, Gun Hyun Park, Jeong-Yeol Lee, Okorie Eme Okorie, and Sungchul C. Bai. Effects of dietary propolis supplementation on growth performance, immune responses, disease resistance and body composition of juvenile eel, Anguilla japonica. Aquaculture International, 20(3):513–523, June 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9482-4.

Skoric:2012:SEI

[1050] Stefan Skorić, Božidar Rašković, Vesna Poleksić, Zoran Gačić, and Mirjana Lenhardt. Scoring of the extent and intensity of carp (Cyprinus carpio) skin changes made by cormorants (Phalacrocorax carbo sinensis): relationship between morphometric and histological indices. Aquaculture International, 20(3):525–535, June 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-011-9483-3.

Woods:2012:BGM

[1051] Chris M. C. Woods, Oliver Floerl, and Barbara J. Hayden. Biofouling on Greenshell^{T M} mussel (Perna canaliculus) farms: a preliminary assessment and potential implications for sustainable aquaculture practices. Aquaculture International, 20(3):537–557, June 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-011-9484-2.

Godoy:2012:EDS

[1052] Leandro Cesar Godoy, Clarisse Odebrecht, Eduardo Ballester, Tatiana Germano Martins, and Wilson Wasielesky. Effect of diatom supplementation during the nursery rearing of Litopenaeus vannamei (Boone, 1931) in a heterotrophic culture system. Aquaculture International, 20 (3):559–569, June 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9485-1.

Soula:2012:EVV

[1053] Mohamed Soula, Ana Navarro, Silvia Hildebrandt, María J. Zamorano, Javier Roo, Carmen M. Hernández-Cruz, and Juan M. Afonso. Evaluation of VIE (visible implant elastomer) and PIT (passive integrated transponder) physical tagging systems for the identification of red porgy fingerlings (Pagrus pagrus). Aquaculture International, 20(3):571–583, June 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-011-9486-0.

Laiz-Carrion:2012:IFD

[1054] Raúl Laiz-Carrión, Inês Rosa Viana, Juana Rosa Cejas, Ignacio Ruiz-Jarabo, Salvador Jerez, Juan Antonio Martos, Almansa Berro Eduardo, and Juan Miguel Mancera. Influence of food deprivation and high stocking density on energetic metabolism and stress response in red porgy, Pagrus pagrus L. Aquaculture International, 20(3):585–599, June 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9488-y.

Jobling:2012:NRC

[1055] Malcolm Jobling. National Research Council (NRC): Nutrient requirements of fish and shrimp. Aquaculture International, 20(3):601– 602, June 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-011-9480-6.

Jobling:2012:BRP

[1056] Malcolm Jobling. Book review: P. K. T. Woo and K. Buchmann (eds): Fish parasites: pathobiology and protection. Aquaculture International, 20(3):603–604, June 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9517-5.

Nguyen:2012:IPM

[1057] Thong Tien Nguyen. Implicit price of mussel characteristics in the auction market. Aquaculture International, 20(4):605–618, August 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9489-x.

Alam:2012:TET

[1058] Md. Ferdous Alam, Md. Akhtaruzzaman Khan, and A. S. M. Anwarul Huq. Technical efficiency in tilapia farming of Bangladesh: a stochastic frontier production approach. Aquaculture International, 20 (4):619–634, August 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9491-3.

Subramannian:2012:PLC

[1059] Selven Subramannian and Rosamma Philip. Pharmacological level of copper induces the immune and antioxidant mechanisms of Fenneropenaeus indicus conferring better protection against white spot syndrome virus infection. Aquaculture International, 20(4):635–647, August 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9492-2.

Wang:2012:AMD

[1060] Hongxia Wang, Zhaoxia Cui, Danhua Wu, Enmian Guo, Yuan Liu, Chunlin Wang, Xiurong Su, and Taiwu Li. Application of microsatellite DNA parentage markers in the swimming crab Portunus trituberculatus. Aquaculture International, 20(4):649–656, August 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-011-9493-1.

Fu:2012:DEF

[1061] Songzhe Fu, Jiazheng Shen, Kang Chen, Junling Tu, and Ying Liu. A dead-end filtration method to remove particle-associated pathogens in aquaculture systems. Aquaculture International, 20(4):657–672, August 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-011-9494-0.

Stefani:2012:ECP

[1062] Gianluca Stefani, Riccardo Scarpa, and Alessio Cavicchi. Exploring consumer’s preferences for farmed sea bream. Aquaculture International, 20 (4):673–691, August 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-011-9495-z.

Motlagh:2012:MGM

[1063] Hamid Reza Ahmadnia Motlagh, Mehrdad Farhangi, Gholamreza Rafiee, and Farzaneh Noori. Modulating gut microbiota and digestive enzyme activities of Artemia urmiana by administration of different levels of Bacillus subtilis and Bacillus licheniformis. Aquaculture International, 20(4):693–705, August 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9497-5.

Francavilla:2012:ECL

[1064] Matteo Francavilla, Pasquale Trotta, Marianna Marangi, Paolo Breber, and Annunziata Giangaspero. Environmental conditions in a lagoon and their possible effects on shellfish contamination by Giardia and Cryptosporidium. Aquaculture International, 20(4):707–724, August 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9498-4.

Oryan:2012:CHI

[1065] Ahmad Oryan, Nasrollah Ahmadi, Mostafa Akhlaghi, and Arsalan Hosseini. A comparative histopathological, immunohistochemical and nested-PCR study for diagnosis of infectious pancreatic necrosis in the naturally infected cultured rainbow trout (Oncorhynchus mykiss). Aquaculture International, 20(4):725–734, August 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9499-3.

Jana:2012:CEF

[1066] B. B. Jana, Sujoy Kumar Bag, and Sukanta Rana. Comparative evaluation of the fertilizer value of human urine, cow manure and their mix for the production of carp fingerlings in small holding tanks. Aquaculture International, 20(4):735–749, August 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9500-1.

Brenner:2012:HGP

[1067] M. Brenner, C. Buchholz, O. Heemken, B. H. Buck, and A. Koehler. Health and growth performance of the blue mussel (Mytilus edulis L.) from two hanging cultivation sites in the German Bight: a nearshore– offshore comparison. Aquaculture International, 20(4):751–778, August 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-012-9501-0.

Iribarren:2012:PEE

[1068] Diego Iribarren, Paula Dagá, María Teresa Moreira, and Gumersindo Feijoo. Potential environmental effects of probiotics used in aquaculture. Aquaculture International, 20(4):779–789, August 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-012-9502-z.

Sun:2012:CIC

[1069] Yun-Zhang Sun, Hong-Ling Yang, Ru-Long Ma, Kun-Peng Huang, and Ji-Dan Ye. Culture-independent characterization of the autochthonous gut microbiota of grouper Epinephelus coioides following the administration of probiotic Enterococcus faecium. Aquaculture International, 20 (4):791–801, August 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9503-y.

Jobling:2012:BRG

[1070] Malcolm Jobling. Book review: G. L. Fletcher and M. L. Rise (eds): Aquaculture biotechnology. Aquaculture International, 20(4):803–805, August 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-012-9526-4.

Jobling:2012:BRJ

[1071] Malcolm Jobling. Book review: J. S. Lucas and P. C. Southgate (eds): Aquaculture — Farming aquatic animals and plants. Aquaculture International, 20(4):807–809, August 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9530-8.

Jobling:2012:BRR

[1072] Malcolm Jobling. Book review: R. J. Roberts (ed): Fish pathology. Aquaculture International, 20(4):811–812, August 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9552-2.

Mitchell:2012:DNH

[1073] Susan O. Mitchell, Emily J. Baxter, Celia Holland, and Hamish D. Rodger. Development of a novel histopathological gill scoring protocol for assessment of gill health during a longitudinal study in marinefarmed Atlantic salmon (Salmo salar). Aquaculture International, 20 (5):813–825, October 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9504-x.

Quoc:2012:FAF

[1074] Nam Cao Quoc, Nico Vromant, Duong Le Thanh, and Frans Ollevier. Factors affecting fish yield and profit in fish pen culture in flooded ricefields. Aquaculture International, 20(5):827–845, October 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-012-9505-9.

Alonso:2012:MCM

[1075] Mercedes Alonso, Fátima C. Lago, Juan M. Vieites, and Montserrat Espiñeira. Molecular characterization of microalgae used in aquaculture with biotechnology potential. Aquaculture International, 20(5):

847–857, October 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9506-8.

Amirkolaie:2012:DFC

[1076] A. Keramat Amirkolaie, S. Mahdavi, and S. A. Hosseini. Dietary fat content and feed supply influence growth and body composition in juvenile beluga sturgeon (Huso huso). Aquaculture International, 20(5): 859–867, October 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9507-7.

Guroy:2012:SNC

[1077] Betül Güroy, İzzet Şahin, Serhan Mantoğlu, and Selin Kayalı. Spirulina as a natural carotenoid source on growth, pigmentation and reproductive performance of yellow tail cichlid Pseudotropheus acei. Aquaculture International, 20(5):869–878, October 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9512-x.

Oliveira:2012:BTB

[1078] J. Oliveira, F. Castilho, A. Cunha, and M. J. Pereira. Bacteriophage therapy as a bacterial control strategy in aquaculture. Aquaculture International, 20(5):879–910, October 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9515-7.

Mahalakshmi:2012:DDM

[1079] P. Mahalakshmi, K. Ganesan, and V. Venkatasubramanian. DMTI-OLA: decision making tool for identification of optimal location for aquaculture farming development. Aquaculture International, 20(5): 911–925, October 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9516-6.

Jin:2012:GAE

[1080] Wu Jin, Zhiyi Bai, Longlong Fu, Genfang Zhang, and Jiale Li. Genetic analysis of early growth traits of the triangle shell mussel, Hyriopsis cumingii, as an insight for potential genetic improvement to pearl quality and yield. Aquaculture International, 20(5):927–933, October 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9518-4.

Farhat:2012:EDA

[1081] Farhat and Mukhtar A. Khan. Effects of dietary arginine levels on growth, feed conversion, protein productive value and carcass composition of stinging catfish fingerling Heteropneustes fossilis (Bloch). Aquaculture International, 20(5):935–950, October 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9519-3.

Parmar:2012:EBY

[1082] Prakash V. Parmar, H. Shivanada Murthy, C. S. Tejpal, and B. T. Naveen Kumar. Effect of brewer’s yeast on immune response of giant freshwater prawn, Macrobrachium rosenbergii, and its resistance to white muscle disease. Aquaculture International, 20(5):951–964, October 2012. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9520-x.

Glaropoulos:2012:ERB

[1083] Alexios Glaropoulos, Vassilis M. Papadakis, Ioannis E. Papadakis, and Maroudio Kentouri. Escape-related behavior and coping ability of sea bream due to food supply. Aquaculture International, 20(5):965– 979, October 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-012-9521-9.

Matsunari:2012:EDA

[1084] Hiroyuki Matsunari, Hiroshi Hashimoto, Kentaro Oda, Yoshitsugu Masuda, Hitoshi Imaizumi, Kazuhisa Teruya, Hirofumi Furuita, Takeshi Yamamoto, Kazuhisa Hamada, and Keiichi Mushiake. Effect of different algae used for enrichment of rotifers on growth, survival, and swim bladder inflation of larval amberjack Seriola dumerili. Aquaculture International, 20(5):981–992, October 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9522-8.

Anguiano-Beltran:2012:ETB

[1085] Casandra Anguiano-Beltrán, Ricardo Searcy-Bernal, Adrián M. García-Ortega, Zaúl García-Esquivel, and Enrique Valenzuela-Espinoza. Effect of three bacterial isolates from a commercial hatchery on early red abalone (Haliotis rufescens) postlarvae. Aquaculture International, 20 (5):993–1001, October 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9523-7.

Foote:2012:TIP

[1086] Andrew R. Foote, Graham C. Mair, Andrew T. Wood, and Melony J. Sellars. Tetraploid inductions of Penaeus monodon using cold shock. Aquaculture International, 20(5):1003–1007, October 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9524-6.

Rajeswari:2012:SIS

[1087] P. Raja Rajeswari, S. Velmurugan, M. Michael Babu, S. Albin Dhas, K. Kesavan, and T. Citarasu. A study on the influence of selected Indian herbal active principles on enhancing the immune system in Fenneropenaeus indicus against Vibrio harveyi infection. Aquaculture International, 20(5):1009–1020, October 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9525-5.

Saroglia:2012:P

[1088] Marco Saroglia and Juan Miguel Mancera. Preface. Aquaculture International, 20(6):1021–1023, December 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9613-6.

Tapia-Paniagua:2012:UPS

[1089] S. T. Tapia-Paniagua, P. Díaz-Rosales, J. M. León-Rubio, I. García de La Banda, C. Lobo, F. J. Alarcón, M. Chabrillón, P. Rosas-Ledesma, J. L. Varela, I. Ruiz-Jarabo, S. Arijo, M. A. Esteban, E. Martínez-Manzanares, J. M. Mancera, M. C. Balebona, and M. A. Moriñigo. Use of the probiotic Shewanella putrefaciens Pdp11 on the culture of Senegalese sole (Solea senegalensis, Kaup 1858) and gilthead seabream (Sparus aurata L.). Aquaculture International, 20(6):1025–1039, December 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9509-5.

Herrera:2012:SDA

[1090] M. Herrera, I. Ruíz-Jarabo, I. Hachero, L. Vargas-Chacoff, A. Amo, and J. M. Mancera. Stocking density affects growth and metabolic parameters in the brill (Scophthalmus rhombus). Aquaculture International, 20(6): 1041–1052, December 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9513-9.

Rodiles:2012:EDP

[1091] Ana Rodiles, Ester Santigosa, Marcelino Herrera, Ismael Hachero-Cruzado, María Luisa Cordero, Silvia Martínez-Llorens, Santosh P. Lall, and Francisco Javier Alarcón. Effect of dietary protein level and source on digestive proteolytic enzyme activity in juvenile Senegalese sole, Solea senegalensis Kaup 1850. Aquaculture International, 20(6): 1053–1070, December 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9508-6.

Rodriguez:2012:IPP

[1092] Rafael Rodríguez, Alicia Felip, Vinicius Cerqueira, Edmond Hala, Silvia Zanuy, and Manuel Carrillo. Identification of a photolabile period for reducing sexual maturation in juvenile male sea bass (Dicentrarchus labrax) by means of a continuous light regime. Aquaculture International, 20(6): 1071–1083, December 2012. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9510-z.

Lopez-Olmeda:2012:DRD

[1093] J. F. López-Olmeda, I. López-García, M. J. Sánchez-Muros, B. Blanco-Vives, R. Aparicio, and F. J. Sánchez-Vázquez. Daily rhythms of digestive physiology, metabolism and behaviour in the European eel (Anguilla anguilla). Aquaculture International, 20(6):1085–1096, December 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9547-z.

Khatooni:2012:ESF

[1094] M. Mardaneh Khatooni, B. Mojazi Amiri, A. Mirvaghefi, V. Jafari, and S. H. Hoseinifar. The effects of salinity on the fertilization rate and rearing of the Persian sturgeon (Acipenser persicus) larvae. Aquaculture International, 20(6):1097–1105, December 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9560-2.

Tomasso:2012:ENA

[1095] J. R. Tomasso. Environmental nitrite and aquaculture: a perspective. Aquaculture International, 20(6):1107–1116, December 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-012-9532-6.

Ribeiro:2012:EPF

[1096] C. S. Ribeiro, A. D. Gomes, V. A. R. O. Vieira, Y. A. Tabata, N. S. Takahashi, and R. G. Moreira. The effect of ploidy on the fatty acid profile during the reproductive cycle of female rainbow trout (Oncorhynchus mykiss). Aquaculture International, 20(6):1117–1137, December 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9468-2.

Mello:2012:FAT

[1097] P. H. Mello, B. C. Araújo, R. L. L. Venturieri, and R. G. Moreira. Fatty acids as a tool to compare cachara (Pseudoplatystoma reticulatum) (Siluriformes: Pimelodidae) and hybrid (Pseudoplatystoma corruscans × Pseudoplatystoma reticulatum) larvae during early development. Aquaculture International, 20(6):1139–1160, December 2012. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-011-9466-4.

Araujo:2012:ICB

[1098] Bruno C. Araújo, Renato M. Honji, Paulo H. de Mello, and Renata Guimarães Moreira. The influence of captive breeding on the fatty acid profiles of Salminus hilarii (Characiformes: Characidae) eggs and larvae. Aquaculture International, 20(6):1161–1181, December 2012. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-011-9472-6.

Adamek:2013:BSB

[1099] Zdeněk Adámek and Blahoslav Maršálek. Bioturbation of sediments by benthic macroinvertebrates and fish and its implication for pond ecosystems: a review. Aquaculture International, 21(1):1–17, February 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9527-3.

Chattopadhyay:2013:ESD

[1100] D. N. Chattopadhyay, B. C. Mohapatra, S. Adhikari, K. C. Pani, J. K. Jena, and A. E. Eknath. Effects of stocking density of Labeo rohita on survival, growth and production in cages. Aquaculture International, 21 (1):19–29, February 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9528-2.

Mamun:2013:RSM

[1101] Shamsuddin Mohammed Mamun, Ulfert Focken, and Klaus Becker. A respirometer system to measure critical and recovery oxygen tensions of fish under simulated diurnal fluctuations in dissolved oxygen. Aquaculture International, 21(1):31–44, February 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9529-1.

Jiang:2013:GFS

[1102] Zengjie Jiang, Guanghua Wang, Jianguang Fang, and Yuze Mao. Growth and food sources of Pacific oyster Crassostrea gigas integrated culture with sea bass Lateolabrax japonicus in Ailian Bay, China. Aquaculture International, 21(1):45–52, February 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9531-7.

Gao:2013:ENF

[1103] Xu Gao, Yukio Agatsuma, and Kazuya Taniguchi. Effect of nitrate fertilization of gametophytes of the kelp Undaria pinnatifida on growth and maturation of the sporophytes cultivated in Matsushima Bay, northern Honshu, Japan. Aquaculture International, 21(1):53–64, February 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9533-5.

Porchas-Cornejo:2013:SZC

[1104] M. A. Porchas-Cornejo, L. R. Martínez-Córdova, M. Martínez-Porchas, R. Barraza-Guardado, and L. Ramos-Trujillo. Study of zooplankton communities in shrimp earthen ponds, with and without organic nutrient-enriched substrates. Aquaculture International, 21(1):65–73, February 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-012-9534-4.

Wang:2013:OAB

[1105] Yi-Da Wang, Hsiao-Yun Chang, Jyh-Yih Chen, and Jian-Chyi Chen. Oral administration of bovine lactoferrin inhibits bacterial infection in tilapia and elevates survival after bacterial infection: an examination of its immune-modulating properties. Aquaculture International, 21(1): 75–96, February 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9537-1.

Yun:2013:ELS

[1106] Biao Yun, Min Xue, Jia Wang, Zhi ying Fan, Xiu-Feng Wu, Yin-Hua Zheng, and Yu-Chang Qin. Effects of lipid sources and lipid peroxidation on feed intake, growth, and tissue fatty acid compositions of largemouth bass (Micropterus salmoides). Aquaculture International, 21 (1):97–110, February 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9538-0.

Shapouri:2013:UPL

[1107] M. Shapouri, P. C. Tavares, C. Martins, P. Pereira, M. Falcão, M. Machado, L. Ribeiro, and L. Cancela da Fonseca. Using pigment level as a primary production indicator to assess organic matter variability in two linked wetland systems with different disturbance levels and its effect on secondary communities. Aquaculture International, 21 (1):111–128, February 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9539-z.

Tong:2013:SED

[1108] X. H. Tong, S. H. Xu, Q. H. Liu, J. Li, Z. Z. Xiao, and D. Y. Ma. Stages of embryonic development and changes in enzyme activities in embryogenesis of turbot (Scophthalmus maximus L.). Aquaculture International, 21 (1):129–142, February 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9540-6.

Hassankiadeh:2013:EDF

[1109] Majid Nikzad Hassankiadeh, Hossein Khara, Mohammad A. Yazdani Sadati, and Hossein Parandavar. Effects of dietary fish oil substitution with mixed vegetable oils on growth and fillet fatty acid composition of juvenile Caspian great sturgeon (Huso huso). Aquaculture International, 21 (1):143–155, February 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9541-5.

George:2013:OOM

[1110] E. Marie George and Christopher C. Parrish. Output of organic material from land-based juvenile Atlantic cod (Gadus morhua) tanks. Aquaculture International, 21(1):157–176, February 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9542-4.

Sabate:2013:CSS

[1111] F. de la S. Sabate, Y. Nakagawa, T. Nasu, W. Sakamoto, and S. Miyashita. Critical swimming speed and maximum sustainable swimming speed of juvenile Pacific bluefin tuna, Thunnus orientalis. Aquaculture International, 21(1):177–181, February 2013. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9543-3.

Sarma:2013:RFA

[1112] Kamal Sarma, K. Prabakaran, P. Krishnan, G. Grinson, and A. Anand Kumar. Response of a freshwater air-breathing fish, Clarias batrachus to salinity stress: an experimental case for their farming in brackishwater areas in Andaman, India. Aquaculture International, 21(1): 183–196, February 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9544-2.

Phinchongsakuldit:2013:PGC

[1113] Jaros Phinchongsakuldit, Puangtip Chaipakdee, James F. Collins, Mullica Jaroensutasinee, and John F. Y. Brookfield. Population genetics of cobia (Rachycentron canadum) in the Gulf of Thailand and Andaman Sea: fisheries management implications. Aquaculture International, 21 (1):197–217, February 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9545-1.

Vaseeharan:2013:MMT

[1114] Baskaralingam Vaseeharan, Perumal Rajakamaran, David Jayaseelan, and Anita Yeshvadha Vincent. Molecular markers and their application in genetic diversity of penaeid shrimp. Aquaculture International, 21 (2):219–241, April 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9582-9.

Hannon:2013:RTC

[1115] Colin Hannon, Rick A. Officer, and Jean Le Dorven. Review of the technical challenges facing aquaculture of the European abalone Haliotis tuberculata in Ireland. Aquaculture International, 21(2):243–254, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9584-7.

Ahmed:2013:DAR

[1116] Imtiaz Ahmed. Dietary arginine requirement of fingerling Indian catfish (Heteropneustes fossilis, Bloch). Aquaculture International, 21(2): 255–271, April 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9548-y.

Harlioglu:2013:EDS

[1117] Muzaffer Mustafa Harlioğlu, Kenan Köprücü, Ayşe Gül Harlioğlu, Serpil Mişe Yonar, Tuba Çakmak Duran, Mehmet Nuri Çakmak, Önder Aksu, Sinan Özcan, Filiz Kutluyer, and Hande Gündoğdu. Effect of dietary n − 3 series fatty acids on sperm production in the freshwater crayfish, Astacus leptodactylus (Eschscholtz) (Astacidae). Aquaculture International, 21(2):273–282, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9549-x.

Guzman-Aguero:2013:FPS

[1118] Juan Eduardo Guzmán-Agüero, Mario Nieves-Soto, Miguel Ángel Hurtado, Pablo Piña-Valdez, and María del Carmen Garza-Aguirre. Feeding physiology and scope for growth of the oyster Crassostrea corteziensis (Hertlein, 1951) acclimated to different conditions of temperature and salinity. Aquaculture International, 21(2):283–297, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-012-9550-4.

Campos-Montes:2013:GPG

[1119] Gabriel Ricardo Campos-Montes, Hugo Horacio Montaldo, Alfonso Martínez-Ortega, Alejandro Martínez Jiménez, and Héctor Castillo-Juárez. Genetic parameters for growth and survival traits in Pacific white shrimp Penaeus (Litopenaeus) vannamei from a nucleus population undergoing a two-stage selection program. Aquaculture International, 21 (2):299–310, April 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9553-1.

Ge:2013:AFN

[1120] Changzi Ge, Fan Zhang, Binduo Xu, Yiping Ren, and Zhenlin Liang. Accumulation flux of nitrogen in mudflats and its implications for benthic shellfish culture. Aquaculture International, 21(2):311–326, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9554-0.

Biswas:2013:ODL

[1121] Biswajit Kumar Biswas, Amal Biswas, Ito Junichi, Yang-Su Kim, and Kenji Takii. The optimal dietary level of ascorbic acid for juvenile Pacific bluefin tuna, Thunnus orientalis. Aquaculture International, 21 (2):327–336, April 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9555-z.

Daga:2013:BPI

[1122] P. Dagá, G. Feijoo, M. T. Moreira, D. Costas, A. G. Villanueva, and J. M. Lema. Bioencapsulated probiotics increased survival, growth and improved gut flora of turbot (Psetta maxima) larvae. Aquaculture International, 21(2):337–345, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9556-y.

Mahanand:2013:OFF

[1123] Sudhansu Shekhar Mahanand, Sanjib Moulick, and P. Srinivasa Rao. Optimum formulation of feed for rohu, Labeo rohita (Hamilton), with biofloc as a component. Aquaculture International, 21(2):347–360, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9557-x.

Ho:2013:DEA

[1124] Adeljean L. F. C. Ho, Stephen K. O’Shea, and Harold F. Pomeroy. Dietary esterified astaxanthin effects on color, carotenoid concentrations, and compositions of clown anemonefish, Amphiprion ocellaris, skin. Aquaculture International, 21(2):361–374, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-012-9558-9.

Suloma:2013:MBM

[1125] Ashraf Suloma, Rania S. Mabroke, and Ehab R. El-Haroun. Meat and bone meal as a potential source of phosphorus in plant-protein-based diets for Nile tilapia (Oreochromis niloticus). Aquaculture International, 21(2):375–385, April 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9559-8.

Kumar:2013:EDS

[1126] N. Ranjit Kumar, Ram Prakash Raman, Sanjay B. Jadhao, Rajive Kumar Brahmchari, Kundan Kumar, and Gyanaranjan Dash. Effect of dietary supplementation of Bacillus licheniformis on gut microbiota, growth and immune response in giant freshwater prawn, Macrobrachium rosenbergii (de Man, 1879). Aquaculture International, 21(2):387–403, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-012-9567-8.

Kekalainen:2013:HSM

[1127] Jukka Kekäläinen, Lars Figenschou, Matti Janhunen, Raine Kortet, Nina Peuhkuri, and Geir Rudolfsen. Hatchery selection may depress the number of motile sperm but intensify selection for their swimming velocity in the Arctic charr. Aquaculture International, 21(2):405–411, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9568-7.

Valverde:2013:AAC

[1128] Jesús Cerezo Valverde, Silvia Martínez-Llorens, Ana Tomás Vidal, Miguel Jover, Carmen Rodríguez, Juán Estefanell, Joan I. Gairín, Pedro Miguel Domingues, Carlos J. Rodríguez, and Benjamín García García. Amino acids composition and protein quality evaluation of marine species and meals for feed formulations in cephalopods. Aquaculture International, 21(2):413–433, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9569-6.

Andersen:2013:ILD

[1129] Sissel Andersen, Gyda Christophersen, and Thorolf Magnesen. Implications of larval diet concentration on post-larval yield in a production scale flow-through system for scallops (Pecten maximus Lamarck) in Norway. Aquaculture International, 21(2):435–452, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9570-0.

Sahu:2013:CNP

[1130] Bharat Chandra Sahu, Subhendu Adhikari, and Lambodar Dey. Carbon, nitrogen and phosphorus budget in shrimp (Penaeus monodon) culture ponds in eastern India. Aquaculture International, 21(2):453– 466, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-012-9573-x.

Naylor:2013:EDS

[1131] Matthew A. Naylor, Horst Kaiser, and Clifford L. W. Jones. The effect of dosing with sodium hydroxide (NaOH^{−}) on water pH and growth of Haliotis midae in an abalone serial-use raceway. Aquaculture International, 21(2):467–479, April 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9574-9.

Touraki:2013:AEL

[1132] Maria Touraki, Gerda Karamanlidou, Mihalis Koziotis, and Ioannis Christidis. Antibacterial effect of Lactococcus lactis subsp. lactis on Artemia franciscana nauplii and Dicentrarchus labrax larvae against the fish pathogen Vibrio anguillarum. Aquaculture International, 21 (2):481–495, April 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9579-4.

Azarm:2013:GRF

[1133] Hamid Mohammadi Azarm, Abdolmohammad Abedian-Kenari, and Mehdi Hedayati. Growth response and fatty acid composition of rainbow trout (Oncorhynchus mykiss) fry fed diets containing different levels of soybean and egg lecithin. Aquaculture International, 21(2):497– 509, April 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-012-9583-8.

Anand:2013:ECR

[1134] P. S. Shyne Anand, Sujeet Kumar, A. Panigrahi, T. K. Ghoshal, J. Syama Dayal, G. Biswas, J. K. Sundaray, D. De, R. Ananda Raja, A. D. Deo, S. M. Pillai, and P. Ravichandran. Effects of C:N ratio and substrate integration on periphyton biomass, microbial dynamics and growth of Penaeus monodon juveniles. Aquaculture International, 21 (2):511–524, April 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9585-6.

Hwang:2013:DGT

[1135] Jae-Ho Hwang, Si-Woo Lee, Sung-Ju Rha, Ho-Seop Yoon, Eun-Sik Park, Kyeong-Ho Han, and Seon-Jae Kim. Dietary green tea extract improves growth performance, body composition, and stress recovery in the juvenile black rockfish, Sebastes schlegeli. Aquaculture International, 21 (3):525–538, June 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9586-5.

Frechette:2013:STD

[1136] Marcel Fréchette, José Manuel Urquiza, Gaétan Daigle, Dominique Maheux, and Jean-François Dumais. Self-thinning dynamics in experimental scallop populations. Aquaculture International, 21(3):539– 551, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-012-9587-4.

Keysami:2013:EBS

[1137] Mehran Avakh Keysami and Masoumeh Mohammadpour. Effect of Bacillus subtilis on Aeromonas hydrophila infection resistance in juvenile freshwater prawn, Macrobrachium rosenbergii (de Man). Aquaculture International, 21(3):553–562, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9588-3.

Santos:2013:DEA

[1138] Juliana Ferreira Santos, Patrícia Fernandes Castro, Albino Luciani Gonçalves Leal, Augusto Cézar Vasconcelos de Freitas Júnior, Daniel Lemos, Luiz Bezerra Carvalho, and Ranilson Souza Bezerra. Digestive enzyme activity in juvenile Nile tilapia (Oreochromis niloticus,

l) submitted to different dietary levels of shrimp protein hydrolysate. Aquaculture International, 21(3):563–577, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9589-2.

Zhang:2013:EDA

[1139] Juanjuan Zhang, Xiaoqin Li, Xiangjun Leng, Chenglong Zhang, Zhiying Han, and Feige Zhang. Effects of dietary astaxanthins on pigmentation of flesh and tissue antioxidation of rainbow trout (Oncorhynchus mykiss). Aquaculture International, 21(3):579–589, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9590-9.

Caipang:2013:EGI

[1140] Christopher Marlowe A. Caipang. Expression of genes involved in the early immune response at the distal segment of the gut in Atlantic cod, Gadus morhua L. after vaccination with a bacterial antigen. Aquaculture International, 21(3):591–603, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9591-8.

Guroy:2013:EAU

[1141] Betül Güroy, Sebahattin Ergün, Daniel L. Merrifield, and Derya Güroy. Effect of autoclaved Ulva meal on growth performance, nutrient utilization and fatty acid profile of rainbow trout, Oncorhynchus mykiss. Aquaculture International, 21(3):605–615, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9592-7.

Jiang:2013:PCS

[1142] Hai-Bo Jiang, Li-Qiao Chen, Jian-Guang Qin, Lu-Jiao Gao, Er-Chao Li, Na Yu, Sheng-Ming Sun, and Xue-Qin Jiang. Partial or complete substitution of fish meal with soybean meal and cottonseed meal in Chinese mitten crab Eriocheir sinensis diets. Aquaculture International, 21 (3):617–628, June 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9594-5.

Nowosad:2013:DAE

[1143] Joanna Nowosad, Daniel Żarski, Maria Biłas, Katarzyna Dryl, Sławomir Krejszeff, and Dariusz Kucharczyk. Dynamics of ammonia excretion in juvenile common tench, Tinca tinca (L.), during intensive rearing under controlled conditions. Aquaculture International, 21(3):629– 637, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-012-9596-3.

Peharda:2013:PAB

[1144] Melita Peharda, Daria Ezgeta-Balić, John Davenport, and Nedo Vrgoč. The potential for aquaculture of the bearded horse mussel (Modiolus barbatus) and Noah’s Ark shell (Arca noae) in southern Croatia. Aquaculture International, 21(3):639–653, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9598-1.

Nimrat:2013:EGP

[1145] Subuntith Nimrat, Phattanunt Tanutpongpalin, Kallaya Sritunyalucksana, Traimat Boonthai, and Verapong Vuthiphandchai. Enhancement of growth performance, digestive enzyme activities and disease resistance in black tiger shrimp (Penaeus monodon) postlarvae by potential probiotics. Aquaculture International, 21(3):655–666, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-012-9600-y.

Xie:2013:RER

[1146] Zhongguo Xie, Furong Wang, Bao Lou, Haiying Liu, and Shidong Guo. Retention efficiency and release of nutrients in the digestive tract of larval shrimp (Penaeus japonicus Bate) using different microencapsulated diets. Aquaculture International, 21(3):667–678, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-012-9601-x.

Marchand:2013:EWH

[1147] Pierre-André Marchand, David L. Straus, Andreas Wienke, Lars-Flemming Pedersen, and Thomas Meinelt. Effect of water hardness on peracetic acid toxicity to zebrafish, Danio rerio, embryos. Aquaculture International, 21(3):679–686, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9602-9.

Xu:2013:EDD

[1148] Wei-Na Xu, Wen-Bin Liu, Mei fang Shen, Gui-Feng Li, Ying Wang, and Wei wei Zhang. Effect of different dietary protein and lipid levels on growth performance, body composition of juvenile red swamp crayfish (Procambarus clarkii). Aquaculture International, 21(3):687– 697, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-012-9603-8.

Zhao:2013:ECD

[1149] Chong Zhao, Weijie Zhang, Yaqing Chang, Haisen Zhou, Jian Song, and Shibin Luo. Effects of continuous and diel intermittent feeding regimes on food consumption, growth and gonad production of the sea urchin Strongylocentrotus intermedius of different size classes. Aquaculture International, 21(3):699–708, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9604-7.

Noseda:2013:DMQ

[1150] Bert Noseda, Anh Ngoc Tong Thi, Lodewijk Rosseel, Frank Devlieghere, and Liesbeth Jacxsens. Dynamics of microbiological quality and safety of Vietnamese Pangasianodon hypophthalmus during processing. Aquaculture International, 21(3):709–727, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9605-6.

Jobling:2013:BRC

[1151] Malcolm Jobling. Book review: C. C. Mischke (ed): Aquaculture pond fertilization: impacts of nutrient input on production. Aquaculture International, 21(3):729–731, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9593-6.

Jobling:2013:BRJ

[1152] Malcolm Jobling. Book review: J. H. Tidwell (ed): Aquaculture production systems. Aquaculture International, 21(3):733–735, June 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9599-0.

Policar:2013:P

[1153] Tomáš Policar and Zdeněk Adámek. Preface. Aquaculture International, 21(4):737–738, August 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9620-7.

Myslowski:2013:GSW

[1154] Bartosz Mysłowski, Remigiusz Panicz, Jacek Sadowski, and Paulina Hofsoe. Genetic structure of the whitefish (Coregonus lavaretus) population inhabiting the Miedwie Lake, Poland, based on partial ND-1 and ITS-1 gene sequences. Aquaculture International, 21(4):739–750, August 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9575-8.

Panicz:2013:MMC

[1155] Remigiusz Panicz, Paulina Hofsoe, Jacek Sadowski, Bartosz Mysłowski, and Mirosław Półgesek. Morphometric and molecular characterisation of Cyprinus carpio × Carassius auratus hybrids. Aquaculture International, 21(4):751–758, August 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9577-6.

Li:2013:IVG

[1156] Ya-Juan Li, Zhuo Yu, Ming-Zhao Zhang, Cong Qian, Syuiti Abe, and Katsutoshi Arai. Induction of viable gynogenetic progeny using eggs and UV-irradiated sperm from the Chinese tetraploid loach, Misgurnus anguillicaudatus. Aquaculture International, 21(4):759–768, August 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9551-3.

Fujimoto:2013:HSI

[1157] Takafumi Fujimoto, Suzu Sakao, Kouzou Oshima, Etsuro Yamaha, and Katsutoshi Arai. Heat-shock-induced tetraploid and diploid/tetraploid mosaic in pond loach, Misgurnus anguillicaudatus. Aquaculture International, 21(4):769–781, August 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9581-x.

Svinger:2013:SOB

[1158] Viktor W. Svinger, Tomas Policar, Christoph Steinbach, Simona Polakova, Antonin Jankovych, and Jan Kouril. Synchronization of ovulation in brook char (Salvelinus fontinalis, Mitchill 1814) using emulsified d-Arg^{6} Pro^{9} NEt sGnRHa. Aquaculture International, 21(4): 783–799, August 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9578-5.

Zakes:2013:EGH

[1159] Zdzisław Zakeś, Mirosław Szczepkowski, Konrad Partyka, and Krzysztof Wunderlich. Effect of gonadotropin hormonal stimulation on out-of-season propagation success of different year classes of indoor-reared pikeperch (Sander lucioperca (L.)). Aquaculture International, 21(4): 801–810, August 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9562-0.

Kristan:2013:HIO

[1160] Jiří Křišt’an, Sayyed Mohammad Hadi Alavi, Vlastimil Stejskal, and Tomáš Policar. Hormonal induction of ovulation in pikeperch (Sander lucioperca L.) using human chorionic gonadotropin (hCG) and mammalian GnRH analogue. Aquaculture International, 21(4):811–818, August 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9572-y.

Zarski:2013:EDC

[1161] Daniel Żarski, Katarzyna Targońska, Rafał Kaszubowski, Patrick Kestemont, Pascal Fontaine, Sławomir Krejszeff, Krzysztof Kupren, and Dariusz Kucharczyk. Effect of different commercial spawning agents and thermal regime on the effectiveness of pikeperch, Sander lucioperca (L.), reproduction under controlled conditions. Aquaculture International, 21 (4):819–828, August 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9597-2.

Kaminski:2013:ETG

[1162] R. Kamiński, J. Wolnicki, J. Sikorska, and V. Garcia. Effects of temperature on growth, survival and body composition in larvae of barbel, Barbus barbus (L.). Aquaculture International, 21(4):829–841, August 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9571-z.

SchillerVestergren:2013:ECL

[1163] AnnaLotta Schiller Vestergren, Sofia Trattner, Jinfeng Pan, Pernilla Johnsson, Afaf Kamal-Eldin, Eva Brännäs, Ali A. Moazzami, and Jana Pickova. The effect of combining linseed oil and sesamin on the fatty acid composition in white muscle and on expression of lipid-related genes in white muscle and liver of rainbow trout (Oncorhynchus mykiss). Aquaculture International, 21(4):843–859, August 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9511-y.

Havasi:2013:PTT

[1164] Máté Havasi, Tamás Oláh, Zoltán Felföldi, Szabolcs Nagy, and Miklós Bercsényi. Passing times of two types of feeds in wels (Silurus glanis) at three different temperatures. Aquaculture International, 21(4): 861–867, August 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-012-9564-y.

Policar:2013:EFS

[1165] Tomas Policar, Vlastimil Stejskal, Jiri Kristan, Peter Podhorec, Viktor Svinger, and Martin Blaha. The effect of fish size and stocking density on the weaning success of pond-cultured pikeperch Sander lucioperca L. juveniles. Aquaculture International, 21(4):869–882, August 2013. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9563-z.

Blaha:2013:NRK

[1166] Martin Bláha, Irena Šetlíkova, Jiří Musil, and Tomáš Policar. No reason for keeping 0+ perch (Perca fluviatilis L.) with the prey fish. Aquaculture International, 21(4):883–896, August 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9576-7.

Vsetickova:2013:ICP

[1167] Lucie Všetičková and Zdeněk Adámek. The impact of carp pond management upon macrozoobenthos assemblages in recipient pond canals. Aquaculture International, 21(4):897–925, August 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9565-x.

Gal:2013:PNR

[1168] Dénes Gál, Ferenc Pekár, Tünde Kosáros, and Éva Kerepeczki. Potential of nutrient reutilisation in combined intensive–extensive pond systems. Aquaculture International, 21(4):927–937, August 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-012-9561-1.

Jarmolowicz:2013:IED

[1169] Sylwia Jarmołowicz, Zdzisław Zakeś, Andrzej Siwicki, Elżbieta Terech-Majewska, Agata Kowalska, Konrad Partyka, and Marek Hopko. Immunomodulatory effect of dietary brewer’s yeast extract in Sander lucioperca juveniles against the challenge of Aeromonas salmonicida. Aquaculture International, 21(4):939–945, August 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9546-0.

Falahatkar:2013:SRG

[1170] Bahram Falahatkar and Samaneh Poursaeid. Stress responses of great sturgeon Huso huso subjected to husbandry stressors. Aquaculture International, 21(4):947–959, August 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9566-9.

Kroupova:2013:WDS

[1171] H. Kroupova, V. Stejskal, J. Kouril, J. Machova, V. Piackova, and E. Zuskova. A wide difference in susceptibility to nitrite between Eurasian perch (Perca fluviatilis L.) and largemouth bass (Micropterus salmoides Lac.). Aquaculture International, 21(4):961–967, August 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9580-y.

Ahmed:2013:EDG

[1172] Faruq Ahmed, Yasuyuki Koike, Carlos Augusto Strüssmann, and Seiichi Watanabe. Effect of density on growth and feed consumption of the abalones Haliotis discus discus, H. gigantea, H. madaka and their hybrids. Aquaculture International, 21(5):969–986, October 2013. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9606-5.

Emerenciano:2013:EST

[1173] Maurício Emerenciano, Gerard Cuzon, Miguel Arévalo, Maite Mascaró Miquelajauregui, and Gabriela Gaxiola. Effect of short-term fresh food supplementation on reproductive performance, biochemical composition, and fatty acid profile of Litopenaeus vannamei (Boone) reared under biofloc conditions. Aquaculture International, 21(5):987–1007, October 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-012-9607-4.

Galimany:2013:SCA

[1174] Eve Galimany, Jennifer H. Alix, Mark S. Dixon, and Gary H. Wikfors. Short communication: adaptability of the feeding behavior of intertidal ribbed mussels (Geukensia demissa) to constant submersion. Aquaculture International, 21(5):1009–1015, October 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9608-3.

Vanhonacker:2013:GEC

[1175] Filiep Vanhonacker, Zuzanna Pieniak, and Wim Verbeke. European consumer image of farmed fish, wild fish, seabass and seabream. Aquaculture International, 21(5):1017–1033, October 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9609-2.

Han:2013:CDF

[1176] Tingting Han, Zengjie Jiang, Jianguang Fang, Jihong Zhang, Yuze Mao, Jian Zou, Yao Huang, and Dongzhe Wang. Carbon dioxide fixation by the seaweed Gracilaria lemaneiformis in integrated multi-trophic aquaculture with the scallop Chlamys farreri in Sanggou Bay, China. Aquaculture International, 21(5):1035–1043, October 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9610-9.

Mordenti:2013:CRW

[1177] Oliviero Mordenti, Andrea Di Biase, Giuseppe Bastone, Rubina Sirri, Annalisa Zaccaroni, and Albamaria Parmeggiani. Controlled reproduction in the wild European eel (Anguilla anguilla): two populations compared. Aquaculture International, 21(5):1045–1063, October 2013. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9611-8.

Aarab:2013:EDI

[1178] Lahoussine Aarab, Alejandro Pérez-Camacho, María del Pino Viera-Toledo, Gercende Courtois de Viçose, Hipólito Fernández-Palacios, and Lucia Molina. Embryonic development and influence of egg density on early veliger larvae and effects of dietary microalgae on growth of brown mussel Perna perna (L. 1758) larvae under laboratory conditions. Aquaculture International, 21(5):1065–1076, October 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9612-7.

Florenciano:2013:PMD

[1179] M. D. Ayala Florenciano, R. Cal, J. Hernández-Urcera, A. Blanco, and O. López-Albors. Post-mortem degradation of the muscle tissue in diploid and triploid turbots (Scophthalmus maximus L.). Aquaculture International, 21(5):1077–1090, October 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9614-5.

Almli:2013:SRI

[1180] Valérie Lengard Almli and Margrethe Hersleth. Salt replacement and injection salting in smoked salmon evaluated from descriptive and hedonic sensory perspectives. Aquaculture International, 21(5):1091–1108, October 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-012-9615-4.

Li:2013:PER

[1181] Yan Li, André Moreira Bordinhon, D. Allen Davis, Wenbing Zhang, and Xuezhi Zhu. Protein: energy ratio in practical diets for Nile tilapia Oreochromis niloticus. Aquaculture International, 21(5):1109– 1119, October 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-012-9616-3.

Vinoj:2013:IEB

[1182] Gopalakrishnan Vinoj, Baskaralingam Vaseeharan, Baranabas David-Jayaseelan, Perumal Rajakumaran, and Cyril Ravi. Inhibitory effects of Bacillus licheniformis (DAB1) and Pseudomonas aeruginosa (DAP1) against Vibrio parahaemolyticus isolated from Fenneropenaeus indicus. Aquaculture International, 21(5):1121–1135, October 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-012-9617-2.

Rekha:2013:AHC

[1183] P. Nila Rekha, P. Ravichandran, R. Gangadharan, J. H. Bhatt, A. Panigrahi, S. M. Pillai, and M. Jayanthi. Assessment of hydrogeochemical characteristics of groundwater in shrimp farming areas in coastal Tamil Nadu, India. Aquaculture International, 21(5):1137–1153, October 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-012-9618-1.

Pan:2013:IED

[1184] Chieh-Yu Pan, Yi-Da Wang, and Jyh-Yih Chen. Immunomodulatory effects of dietary Bacillus coagulans in grouper (Epinephelus coioides) and zebrafish (Danio rerio) infected with Vibrio vulnificus. Aquaculture International, 21(5):1155–1168, October 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-012-9619-0.

Subramanian:2013:IHM

[1185] Selven Subramanian and Rosamma Philip. Identification of haematological markers in shrimp health assessment from the immune profile of Fenneropenaeus indicus on β-1,3-glucan administration and white spot syndrome virus challenge. Aquaculture International, 21(5):1169– 1184, October 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-013-9621-1.

Deng:2013:ESD

[1186] Yuewen Deng, Shao Fu, Feilong Liang, and Shaohe Xie. Effects of stocking density, diet, and water exchange on growth and survival of pearl oyster Pinctada maxima larvae. Aquaculture International, 21(6): 1185–1194, December 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9622-0.

Fu:2013:MMP

[1187] Jianjun Fu, Yubang Shen, Xiaoyan Xu, Yong Chen, Da Li, and Jiale Li. Multiplex microsatellite PCR sets for parentage assignment of grass carp (Ctenopharyngodon idella). Aquaculture International, 21(6):1195– 1207, December 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9623-z.

Han:2013:ETR

[1188] C.-Y. Han, Q.-M. Zheng, and L.-N. Feng. Effects of total replacement of dietary fish oil on growth performance and fatty acid compositions of hybrid tilapia (Oreochromis niloticus × O. aureus). Aquaculture International, 21(6):1209–1217, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9624-y.

Wang:2013:CLL

[1189] Aimin Wang, Guangming Han, Zhitao Qi, Fu Lv, Yebing Yu, Jintian Huang, Tian Wang, and Pao Xu. Cloning of lipoprotein lipase (LPL) and the effects of dietary lipid levels on LPL expression in GIFT tilapia (Oreochromis niloticus). Aquaculture International, 21(6):1219– 1232, December 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9625-x.

Woolley:2013:SIA

[1190] Lindsey D. Woolley, Stewart D. Fielder, and Jian G. Qin. Swimbladder inflation associated with body density change and larval survival in southern bluefin tuna Thunnus maccoyii. Aquaculture International, 21(6):1233–1242, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-013-9626-9.

Zehra:2013:DIR

[1191] Seemab Zehra and Mukhtar A. Khan. Dietary isoleucine requirement of fingerling catla, Catla catla (Hamilton), based on growth, protein productive value, isoleucine retention efficiency and carcass composition. Aquaculture International, 21(6):1243–1259, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9627-8.

Wang:2013:LES

[1192] Yiyan Wang, Haina Wang, Dongchun Yan, Lei Wang, Zhenxing Sun, and Hushan Sun. Lentinan extracted from shiitake mushrooms (Lentinus edodes) improves the non-specific immunity of sea cucumber (Apostichopus japonicus). Aquaculture International, 21(6):1261–1277, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9628-7.

MacDonald:2013:CRG

[1193] Camilla L. E. MacDonald, Selina M. Stead, and Matthew J. Slater. Consumption and remediation of European seabass (Dicentrarchus labrax) waste by the sea cucumber Holothuria forskali. Aquaculture International, 21(6):1279–1290, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9629-6.

Almeida:2013:RPF

[1194] Diones Bender Almeida, Marco André Paldês da Costa, Liane Ney Bassini, Cecilia Irene Pérez Calabuig, Carla Giovane Avila Moreira, Marília Danyelle Nunes Rodrigues, Harold Julian Pérez, Rafael Aldrighi Tavares, Antonio Sergio Varela, and Heden Luiz Marques Moreira. Reproductive performance in female strains of Nile tilapia, Oreochromis niloticus. Aquaculture International, 21(6):1291–1300, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9630-0.

Zhao:2013:CGP

[1195] Yongchao Zhao, Zhiyi Bai, Longlong Fu, Yue Liu, Guiling Wang, and Jiale Li. Comparison of growth and pearl production in males and females of the freshwater mussel, Hyriopsis cumingii, in China. Aquaculture International, 21(6):1301–1310, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9632-y.

Adams:2013:ESW

[1196] Amanda L. Adams, Evan W. Needham, and Jens Knauer. The effect of shade on water quality parameters and survival and growth of juvenile fluted giant clams, Tridacna squamosa, cultured in a land-based growth trial. Aquaculture International, 21(6):1311–1324, December 2013. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9634-9.

Melo:2013:MSS

[1197] Mauro André Damasceno Melo, Adam Rick Bessa da Silva, Colin Robert Beasley, and Claudia Helena Tagliaro. Multiplex species-specific PCR identification of native and non-native oysters (Crassostrea) in Brazil: a useful tool for application in oyster culture and stock management. Aquaculture International, 21(6):1325–1332, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9635-8.

Huang:2013:AAR

[1198] Zhitao Huang, Rong Wan, Xiefa Song, and Eric Hallerman. Assessment of AquaMats for removing ammonia in intensive commercial Pacific white shrimp Litopenaeus vannamei aquaculture systems. Aquaculture International, 21(6):1333–1342, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9636-7.

Sui:2013:ICN

[1199] L. Y. Sui, J. Wang, V. H. Nguyen, P. Sorgeloos, P. Bossier, and G. Van Stappen. Increased carbon and nitrogen supplementation in Artemia culture ponds results in higher cyst yields. Aquaculture International, 21(6): 1343–1354, December 2013. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9637-6.

Rios-Duran:2013:NEA

[1200] M. G. Ríos-Durán, I. R. Valencia, L. G. Ross, and C. A. Martínez-Palacios. Nutritional evaluation of autoclaved Salicornia bigelovii Torr. seed meal supplemented with varying levels of cholesterol on growth, nutrient utilization and survival of the Nile tilapia (Oreochromis niloticus). Aquaculture International, 21(6):1355–1371, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9638-5.

Laghari:2013:QTL

[1201] Muhammad Younis Laghari, Yan Zhang, Punhal Lashari, Xiaofeng Zhang, Peng Xu, Baoping Xin, and Xiaowen Sun. Quantitative trait loci (QTL) associated with growth rate trait in common carp (Cyprinus carpio). Aquaculture International, 21(6):1373–1379, December 2013. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9639-4.

Emerenciano:2013:EBT

[1202] Maurício Emerenciano, Gerard Cuzon, Adriana Paredes, and Gabriela Gaxiola. Evaluation of biofloc technology in pink shrimp Farfantepenaeus duorarum culture: growth performance, water quality, microorganisms profile and proximate analysis of biofloc. Aquaculture International, 21(6):1381–1394, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9640-y.

Wezel:2013:UBV

[1203] Alexander Wezel, Carole Chazoule, and Dominique Vallod. Using biodiversity to valorise local food products: the case of fish ponds in a cultural landscape, their biodiversity, and carp production. Aquaculture International, 21(6):1395–1408, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9641-x.

Jobling:2013:BRW

[1204] Malcolm Jobling. Book review: S. W. Bunting: Principles of sustainable aquaculture. Aquaculture International, 21(6):1409–1411, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-013-9631-z.

Jobling:2013:SCM

[1205] Malcolm Jobling. Suresh Chandra and Mohit Kumar Sharma: Research methodology. Aquaculture International, 21(6):1413–1414, December 2013. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-013-9633-x.

Kucharczyk:2014:P

[1206] Dariusz Kucharczyk, Daniel Żarski, and Andrzej Mamcarz. Preface. Aquaculture International, 22(1):1–3, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9740-8.

Horvath:2014:EDB

[1207] Ákos Horváth, György Hoitsy, Balázs Kovács, Dóra Kánainé Sipos, Ágnes Ősz, Klavdija Bogataj, and Béla Urbányi. The effect of domestication on a brown trout (Salmo trutta m fario) broodstock in Hungary. Aquaculture International, 22(1):5–11, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9665-2.

Palinska-Zarska:2014:DYS

[1208] K. Palińska-Żarska, D. Żarski, S. Krejszeff, J. Nowosad, M. Biłas, K. Trejchel, and D. Kucharczyk. Dynamics of yolk sac and oil droplet utilization and behavioural aspects of swim bladder inflation in burbot, Lota lota L., larvae during the first days of life, under laboratory conditions. Aquaculture International, 22(1):13–27, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9663-4.

Kupren:2014:EDA

[1209] Krzysztof Kupren, Izabela Trabska, Daniel Żarski, Sławomir Krejszeff, Katarzyna Palińska-Żarska, and Dariusz Kucharczyk. Early development and allometric growth patterns in burbot Lota lota L. Aquaculture International, 22(1):29–39, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9680-3.

Furgala-Selezniow:2014:FSB

[1210] Grażyna Furgała-Selezniow, Andrzej Skrzypczak, Dariusz Kucharczyk, Roman Kujawa, Andrzej Mamcarz, Daniel Żarski, and Katarzyna Targońska. Food selection of burbot (Lota lota L.) larvae reared in illuminated net cages in mesotrophic Lake Maróz (north-eastern Poland). Aquaculture International, 22(1):41–52, February 2014. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9707-9.

Krol:2014:EDE

[1211] Jarosław Król, Wojciech Pobłocki, Tomasz Bockenheimer, and Piotr Hliwa. Effect of diethylstilbestrol (DES) and 17 β-estradiol (E2) on growth, survival and histological structure of the internal organs in juvenile European catfish Silurus glanis (L.). Aquaculture International, 22(1):53–62, February 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9664-3.

Chenais:2014:FCC

[1212] Nathalie Chenais, Alexandra Depince, Pierre-Yves Le Bail, and Catherine Labbe. Fin cell cryopreservation and fish reconstruction by nuclear transfer stand as promising technologies for preservation of finfish genetic resources. Aquaculture International, 22(1):63–76, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9671-4.

Ciesla:2014:CIL

[1213] Mirosław Cieśla, Robert Jończyk, Dariusz Gozdowski, Jerzy Śliwiński, Jacek Rechulicz, and Wojciech Andrzejewski. Changes in ide Leuciscus idus (L.) females’ reproductive parameters after stimulation with carp pituitary homogenate (CPH) and ovopel: the effect of domestication? Aquaculture International, 22(1):77–88, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9668-z.

Hliwa:2014:UEG

[1214] Piotr Hliwa, Mamadou Bah, Henryk Kuźmiński, Stefan Dobosz, and Andrzej Ciereszko. Ultrasound evaluation of the gonadal structure in sex-reversed rainbow trout females. Aquaculture International, 22(1): 89–96, February 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9646-5.

Cejko:2014:ETC

[1215] Beata Irena Cejko, Daniel Żarski, Sylwia Judycka, Dariusz Kucharczyk, Beata Sarosiek, and Radosław Kajetan Kowalski. Effect of two commercial preparations containing different GnRH analogues with dopamine antagonists on barbel Barbus barbus (L.) sperm quantity and quality. Aquaculture International, 22(1):97–109, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9657-2.

Cejko:2014:SQS

[1216] Beata Irena Cejko, Sławomir Krejszeff, Sylwia Judycka, Beata Sarosiek, Mariola Dietrich, Dariusz Kucharczyk, and Radosław Kajetan Kowalski. Sperm quality and selected biochemical markers of seminal plasma at the beginning of the reproductive period of common carp, Cyprinus carpio L. Aquaculture International, 22(1):111–122, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9672-3.

Kowalski:2014:EAC

[1217] Radosław Kajetan Kowalski, Beata Irena Cejko, Sławomir Krejszeff, Beata Sarosiek, Sylwia Judycka, Katarzyna Targońska, Dariusz Kucharczyk, and Jan Glogowski. Effect of albumin and casein supplementation on the common carp Cyprinus carpio L. sperm motility parameters measured by CASA. Aquaculture International, 22(1):123–129, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-013-9673-2.

Bozkurt:2014:EDA

[1218] Yusuf Bozkurt, İlker Yavaş, and Cengiz Yıldız. Effect of different avian egg yolk types on fertilization ability of cryopreserved common carp (Cyprinus carpio) spermatozoa. Aquaculture International, 22(1): 131–139, February 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9728-4.

Yavas:2014:CSC

[1219] İlker Yavaş, Yusuf Bozkurt, and Cengiz Yıldız. Cryopreservation of scaly carp (Cyprinus carpio) sperm: effect of different cryoprotectant concentrations on post-thaw motility, fertilization and hatching success of embryos. Aquaculture International, 22(1):141–148, February 2014. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9698-6.

Brzuska:2014:CRE

[1220] Elżbieta Brzuska. Characteristics of the reproduction effectiveness of four Hungarian breeding lines of carp Cyprinus carpio (L.). Aquaculture International, 22(1):149–158, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9675-0.

Sarosiek:2014:MPPa

[1221] Beata Sarosiek, Katarzyna Dryl, Dariusz Kucharczyk, Daniel Żarski, and Radosław K. Kowalski. Motility parameters of perch spermatozoa (Perca fluviatilis L.) during short-term storage with antioxidants addition. Aquaculture International, 22(1):159–165, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9679-9.

Sarosiek:2014:MPPb

[1222] Beata Sarosiek, Sylwia Judycka, Dariusz Kucharczyk, Daniel Żarski, and Radosław K. Kowalski. Motility parameters of perch spermatozoa (Perca fluviatilis L.) with cryoprotectors addition. Aquaculture International, 22 (1):167–172, February 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9687-9.

Zarski:2014:ADA

[1223] Daniel Żarski, Ákos Horváth, Gergely Bernáth, Katarzyna Palińska-Żarska, Sławomir Krejszeff, Tamás Müller, and Dariusz Kucharczyk. Application of different activating solutions to in vitro fertilization of crucian carp, Carassius carassius (L.), eggs. Aquaculture International, 22 (1):173–184, February 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9692-z.

Kowalska:2014:ECC

[1224] Agata Kowalska and Radosław Kajetan Kowalski. The effect of cyclooxygenase (COX) inhibitors on Japanese medaka (Oryzias latipes) reproduction parameters fed with high level of arachidonic acid (20:4n−6). Aquaculture International, 22(1):185–193, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9684-z.

Trejchel:2014:DOF

[1225] Karol Trejchel, Daniel Żarski, Katarzyna Palińska-Żarska, Sławomir Krejszeff, Bartłomiej Dryl, Krzysztof Dakowski, and Dariusz Kucharczyk. Determination of the optimal feeding rate and light regime conditions in juvenile burbot, Lota lota (L.), under intensive aquaculture. Aquaculture International, 22(1):195–203, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9670-5.

Florczyk:2014:GPF

[1226] Krzysztof Florczyk, Jan Mazurkiewicz, Katarzyna Przybylska, Dariusz Ulikowski, Mirosław Szczepkowski, Wojciech Andrzejewski, and Janusz Golski. Growth performance, feed intake and morphology of juvenile European catfish, Silurus glanis (L.) fed diets containing different protein and lipid levels. Aquaculture International, 22(1):205–214, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-013-9667-0.

Falahatkar:2014:ECR

[1227] Bahram Falahatkar, Sobhan R. Akhavan, and Gholamreza Ghaedi. Egg cortisol response to stress at early stages of development in Persian sturgeon Acipenser persicus. Aquaculture International, 22(1):215–223, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-013-9669-y.

Gomulka:2014:AAT

[1228] Piotr Gomułka, Daniel Żarski, Krzysztof Kupren, Sławomir Krejszeff, Katarzyna Targońska, and Dariusz Kucharczyk. Acute ammonia toxicity during early ontogeny of ide Leuciscus idus (Cyprinidae). Aquaculture International, 22(1):225–233, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9677-y.

Falahatkar:2014:EHM

[1229] Bahram Falahatkar and Samaneh Poursaeid. Effects of hormonal manipulation on stress responses in male and female broodstocks of pikeperch Sander lucioperca. Aquaculture International, 22(1):235–244, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-013-9678-x.

Kamaszewski:2014:EFM

[1230] Maciej Kamaszewski and Teresa Ostaszewska. The effect of feeding on morphological changes in intestine of pike-perch (Sander lucioperca L.). Aquaculture International, 22(1):245–258, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9693-y.

Kamaszewski:2014:IFD

[1231] Maciej Kamaszewski, Marta Prasek, Teresa Ostaszewska, and Konrad Dabrowski. The influence of feeding diets containing wheat gluten supplemented with dipeptides or free amino acids on structure and development of the skeletal muscle of carp (Cyprinus carpio). Aquaculture International, 22(1):259–271, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9683-0.

Saber:2014:SSD

[1232] Mohammad Hassanzadeh Saber and Ali Hallajian. Study of sex determination system in ship sturgeon, Acipenser nudiventris using meiotic gynogenesis. Aquaculture International, 22(1):273–279, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9676-z.

Ocalewicz:2014:CHS

[1233] Konrad Ocalewicz, Piotr Hliwa, Konrad Pomianowski, Rodrigo Lisboa, and Malgorzata Jankun. Cytogenetic and histological studies of the brook trout, Salvelinus fontinalis (Mitchill), and the Arctic char, S. alpinus (L.) hybrids. Aquaculture International, 22(1):281– 288, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-013-9655-4.

Poleksic:2014:MPE

[1234] Vesna Poleksić, Marko Stanković, Zoran Marković, Renata Relić, Nada Lakić, Zorka Dulić, and Božidar Rašković. Morphological and physiological evaluation of common carp (Cyprinus carpio L., 1758) fed extruded compound feeds containing different fat levels. Aquaculture International, 22(1):289–298, February 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9654-5.

Hlavac:2014:ESF

[1235] David Hlaváč, Zdeněk Adámek, Pavel Hartman, and Jan Másílko. Effects of supplementary feeding in carp ponds on discharge water quality: a review. Aquaculture International, 22(1):299–320, February 2014. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9718-6.

Saetae:2014:EPC

[1236] Panthida Saetae, Sakshin Bunthawin, and Raymond J. Ritchie. Environmental persistence of chlorine from prawn farm discharge monitored by measuring the light reactions of photosynthesis of phytoplankton. Aquaculture International, 22(2):321–338, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9642-9.

Zuniga-Jara:2014:BMR

[1237] Sergio Zuniga-Jara and Mauricio Goycolea-Homann. A bioeconomic model for red tilapia culture on the coast of Ecuador. Aquaculture International, 22(2):339–359, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9643-8.

Choi:2014:ETG

[1238] W. M. Choi, W. Y. Mo, S. C. Wu, N. K. Mak, Z. X. Bian, X. P. Nie, and M. H. Wong. Effects of traditional Chinese medicines (TCM) on the immune response of grass carp (Ctenopharyngodon idellus). Aquaculture International, 22(2):361–377, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9644-7.

Guo:2014:CEI

[1239] Biao Guo, Chunguang Gong, Peng Shao, and Lei Jia. The changes of enzymes involved in metabolism and immunity of Apostichopus japonicus (Selenka) after short-term transportations. Aquaculture International, 22(2):379–389, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9645-6.

Drillet:2014:DPC

[1240] Guillaume Drillet and Jörg Dutz. Dealing with the presence of the ciliate Euplotes sp. in cultures of the copepod Acartia tonsa. Aquaculture International, 22(2):391–398, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9647-4.

Honnens:2014:ENP

[1241] Hilke Honnens, Thomas Assheuer, and Ralf-Udo Ehlers. Enrichment of the nematode Panagrolaimus sp., a potential live food for marine aquaculture, with essential n − 3 fatty acids. Aquaculture International, 22(2):399–409, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9648-3.

Dash:2014:CSI

[1242] Supriya Dash, Juhi Samal, and Hrudaynath Thatoi. A comparative study on innate immunity parameters in the epidermal mucus of Indian major carps. Aquaculture International, 22(2):411–421, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9649-2.

Qiu:2014:EMS

[1243] Tianlong Qiu, Libin Zhang, Tao Zhang, and Hongsheng Yang. Effects of mud substrate and water current on the behavioral characteristics and growth of the sea cucumber Apostichopus japonicus in the Yuehu lagoon of northern China. Aquaculture International, 22(2):423–433, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9650-9.

Le:2014:EVD

[1244] Ky T. Le, Thuy T. T. Dao, Ravi Fotedar, and Gavin J. Partrigde. Effects of variation in dietary contents of selenium and vitamin E on growth and physiological and haematological responses of yellowtail kingfish, Seriola lalandi. Aquaculture International, 22(2):435–446, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9651-8.

Aguado-Gimenez:2014:DFO

[1245] F. Aguado-Giménez, M. D. Hernández, J. Cerezo-Valverde, M. A. Piedecausa, and B. García-García. Does flat oyster (Ostrea edulis) rearing improve under open-sea integrated multi-trophic conditions? Aquaculture International, 22(2):447–467, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9653-6.

Trieu:2014:END

[1246] Tran Thi Ngoc Trieu and Minjiao Lu. Estimates of nutrient discharge from striped catfish farming in the Mekong River, Vietnam, by using a 3D numerical model. Aquaculture International, 22(2):469–483, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9656-3.

Lebeda:2014:OSI

[1247] I. Lebeda, B. Dzyuba, M. Rodina, and M. Flajshans. Optimization of sperm irradiation protocol for induced gynogenesis in Siberian sturgeon, Acipenser baerii. Aquaculture International, 22(2):485–495, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9658-1.

Brito:2014:WQG

[1248] Luis Otavio Brito, Rafael Arantes, Caio Magnotti, Rafael Derner, Francisco Pchara, Alfredo Olivera, and Luis Vinatea. Water quality and growth of Pacific white shrimp Litopenaeus vannamei (Boone) in coculture with green seaweed Ulva lactuca (Linnaeus) in intensive system. Aquaculture International, 22(2):497–508, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9659-0.

Jancula:2014:SMC

[1249] Daniel Jančula, Přemysl Mikula, Blahoslav Maršálek, Pavel Rudolf, and František Pochylý. Selective method for cyanobacterial bloom removal: hydraulic jet cavitation experience. Aquaculture International, 22(2):509–521, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9660-7.

Selim:2014:ETM

[1250] Khaled M. Selim, Hana El-hofy, and Riad H. Khalil. The efficacy of three mycotoxin adsorbents to alleviate aflatoxin B_{1}-induced toxicity in Oreochromis niloticus. Aquaculture International, 22(2):523–540, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9661-6.

Bordon:2014:IDG

[1251] Isabella C. A. C. Bordon, Helcio L. A. Marques, José L. Alves, André Rossi, and Iracy L. Pecora. Influence of densities on the growth of brown mussel Perna perna L. cultivated in suspended socks at Caraguatatuba, southeastern Brazil. Aquaculture International, 22(2): 541–549, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9662-5.

Radhakrishnan:2014:IMH

[1252] S. Radhakrishnan, P. Saravana Bhavan, C. Seenivasan, R. Shanthi, and R. Poongodi. Influence of medicinal herbs (Alteranthera sessilis, Eclipta alba and Cissus quadrangularis) on growth and biochemical parameters of the freshwater prawn Macrobrachium rosenbergii. Aquaculture International, 22(2):551–572, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9666-1.

Carton-Kawagoshi:2014:LWE

[1253] Rhea Joy Carton-Kawagoshi, Bessie Joy Elle, Valeriano Corre, Shuichi Satoh, Masahiro Notoya, and Daisuke Fujita. Low water exchange Gracilariopsis bailiniae Zhang & B. M. Xia culture in intensive milkfish culture effluents: effects of seaweed density on seaweed production and effluent treatment. Aquaculture International, 22(2):573–584, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9674-1.

Nugroho:2014:CED

[1254] Rudy Agung Nugroho and Ravi Fotedar. Comparing the effects of dietary selenium and mannan oligosaccharide supplementation on the growth, immune function, and antioxidant enzyme activity in the cultured marron Cherax cainii (Austin, 2002). Aquaculture International, 22(2):585–596, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9682-1.

Altundag:2014:ESO

[1255] Mehmet Sukru Altundag, Serap Ustaoglu Tiril, and Atilla Ozdemir. Effects of safflower oil supplementation in diet on growth performance and body fatty acid composition of turbot (Psetta maxima). Aquaculture International, 22(2):597–605, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9686-x.

Dong:2014:DSB

[1256] Hong biao Dong, Yong quan Su, Yong Mao, Xin xin You, Shao xiong Ding, and Jun Wang. Dietary supplementation with Bacillus can improve the growth and survival of the kuruma shrimp Marsupenaeus japonicus in high-temperature environments. Aquaculture International, 22 (2):607–617, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9688-8.

Hung:2014:MMY

[1257] Dinh Hung and Nguyen Hong Nguyen. Modeling meat yield based on measurements of body traits in genetically improved giant freshwater prawn (GFP) Macrobrachium rosenbergii. Aquaculture International, 22 (2):619–631, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9690-1.

Roy:2014:INV

[1258] Tanami Roy, Goutam Banerjee, Suhas Kumar Dan, Pinki Ghosh, and Arun Kumar Ray. Improvement of nutritive value of sesame oilseed meal in formulated diets for rohu, Labeo rohita (Hamilton), fingerlings after fermentation with two phytase-producing bacterial strains isolated from fish gut. Aquaculture International, 22(2):633–652, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9691-0.

Furtado:2014:SCA

[1259] Plínio S. Furtado, Fabiane P. Serra, Luis H. Poersch, and Wilson Wasielesky. Short communication: Acute toxicity of hydrogen peroxide in juvenile white shrimp Litopenaeus vannamei reared in biofloc technology systems. Aquaculture International, 22(2):653–659, April 2014. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9694-x.

Zhang:2014:VHC

[1260] Xiao-Jun Zhang, Xue-Song Bai, Bin-Lun Yan, Ke-Ran Bi, and Lei Qin. Vibrio harveyi as a causative agent of mass mortalities of megalopa in the seed production of swimming crab Portunus trituberculatus. Aquaculture International, 22(2):661–672, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9695-9.

Behera:2014:AEA

[1261] Truptimayee Behera and Priyabrat Swain. Antigen encapsulated alginate-coated chitosan microspheres stimulate both innate and adaptive immune responses in fish through oral immunization. Aquaculture International, 22(2):673–688, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9696-8.

Charan:2014:EAI

[1262] Ravi Charan, Padinhate Purayil Suresh Babu, Gundaboena Venugopal, Narinder Kumar Chadha, and Kommula Bhaskara Sreeramamurty. Effect of aromatase inhibitors on the ovarian development of stunted yearlings of rohu (Labeo rohita): a preliminary study. Aquaculture International, 22(2):689–697, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9697-7.

Lamari:2014:SLA

[1263] Faouzi Lamari, Khouadja Sadok, Amina Bakhrouf, and François-Joël Gatesoupe. Selection of lactic acid bacteria as candidate probiotics and in vivo test on Artemia nauplii. Aquaculture International, 22 (2):699–709, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9699-5.

Makinen:2014:TIF

[1264] Timo Mäkinen, Pekka Salmi, and Leena Forsman. Towards interactive fish farming governance? A comparison of Finland and Sweden. Aquaculture International, 22(2):711–721, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9700-3.

Kucuk:2014:EFF

[1265] Ercan Küçük, İlhan Aydın, Hamza Polat, Orhan Tufan Eroldoğan, and Temel Şahin. Effect of feeding frequency on growth, feed efficiency and nutrient utilization of juvenile flounder (Platichthys flesus luscus). Aquaculture International, 22(2):723–732, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9701-2.

Harlioglu:2014:EDT

[1266] Muzaffer Mustafa Harlıoğlu, Ayşe Gül Harlıoğlu, Serpil Mişe Yonar, and Tuba Çakmak Duran. Effects of dietary l-tryptophan on the agonistic behavior, growth, and survival of freshwater crayfish Astacus leptodactylus Eschscholtz. Aquaculture International, 22(2):733–748, April 2014. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9702-1.

Ma:2014:EPD

[1267] Xiao Ma, Yi Hu, Xiao-Qing Wang, Qing-Hui Ai, Zhi-Gang He, Fu-Xian Feng, and Xiang-Yang Lu. Effects of practical dietary protein to lipid levels on growth, digestive enzyme activities and body composition of juvenile rice field eel (Monopterus albus). Aquaculture International, 22 (2):749–760, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9703-0.

Li:2014:ETD

[1268] Bin Li, Jian-An Xian, Hui Guo, An-Li Wang, Yu-Tao Miao, Jian-Min Ye, Chao-Xia Ye, and Shao-An Liao. Effect of temperature decrease on hemocyte apoptosis of the white shrimp Litopenaeus vannamei. Aquaculture International, 22(2):761–774, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9704-z.

Allu:2014:PBS

[1269] Prasanna K. R. Allu, B. Chakraborty, Madhusudan Das, Nitish R. Mahapatra, and Sajalendu Ghosh. PCR-based segregation of one hybrid variety of Labeo rohita and Catla catla from their wild-types. Aquaculture International, 22(2):775–782, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9705-y.

Musson:2014:DDS

[1270] Maryke Musson and Horst Kaiser. Development of the digestive system in dusky kob, Argyrosomus japonicus, larvae. Aquaculture International, 22(2):783–796, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9706-x.

Sevgili:2014:HDL

[1271] Hüseyin Sevgili, Adem Kurtoğlu, Masahiko Oikawa, Erol Öztürk, Nurgül Dedebali, Nesrin Emre, and Faruk Pak. High dietary lipids elevate carbon loss without sparing protein in adequate protein-fed juvenile turbot (Psetta maxima). Aquaculture International, 22(2):797–810, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9708-8.

Ren:2014:ISB

[1272] Lihua Ren, Jihong Zhang, Jianguang Fang, Qisheng Tang, Mingliang Zhang, and Meirong Du. Impact of shellfish biodeposits and rotten seaweed on the sediments of Ailian Bay, China. Aquaculture International, 22(2):811–819, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9709-7.

Song:2014:VAS

[1273] Chao Song, Liqiang Zhong, Xiaohui Chen, Wenji Bian, Liping Qiu, Junchao Ming, and Jiazhang Chen. Variation analysis and sample size estimation for growth indicators during PIT-tag-assisted family construction of channel catfish (Ictalurus punctatus). Aquaculture International, 22 (2):821–831, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9710-1.

Luo:2014:BPP

[1274] Shibin Luo, Chong Zhao, Yaqing Chang, Wenping Feng, and Xiaofei Tian. Banana peel provides a new insight into improving gonad flavor in the sea urchin Strongylocentrotus intermedius. Aquaculture International, 22(2):833–841, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9711-0.

Neofitou:2014:MFI

[1275] Nikos Neofitou, Nikos Charizopoulos, Dimitris Vafidis, Konstantinos Skordas, Lamprini Tziantziou, and Christos Neofitou. Mussel farming impacts on trophic status and benthic community structure in Maliakos Gulf (Eastern Mediterranean). Aquaculture International, 22(2):843– 857, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-013-9712-z.

Riisgaard:2014:PPM

[1276] Hans Ulrik Riisgård, Kim Lundgreen, and Poul S. Larsen. Potential for production of ‘mini-mussels’ in Great Belt (Denmark) evaluated on basis of actual and modeled growth of young mussels Mytilus edulis. Aquaculture International, 22(2):859–885, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9713-y.

Byadgi:2014:PIT

[1277] Omkar V. Byadgi, K. M. Shankar, B. T. Naveen Kumar, Rajreddy Patil, and Iqlas Ahmed. Passive immunity in tiger shrimp (Penaeus monodon) fed with monoclonal antibody to white spot syndrome virus. Aquaculture International, 22(2):887–900, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9714-x.

Kumar:2014:ECS

[1278] Sujeet Kumar, P. S. Shyne Anand, D. De, J. K. Sundaray, R. Ananda Raja, G. Biswas, A. G. Ponniah, T. K. Ghoshal, A. D. Deo, A. Panigrahi, and M. Muralidhar. Effects of carbohydrate supplementation on water quality, microbial dynamics and growth performance of giant tiger prawn (Penaeus monodon). Aquaculture International, 22 (2):901–912, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9715-9.

Nicholaus:2014:EBV

[1279] Regan Nicholaus and Zhongming Zheng. The effects of bioturbation by the venus clam Cyclina sinensis on the fluxes of nutrients across the sediment–water interface in aquaculture ponds. Aquaculture International, 22(2):913–924, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9716-8.

Zhang:2014:EBM

[1280] Weijie Zhang, Yaqing Chang, Shibin Luo, Haisen Zhou, Xiaofei Tian, Jun Ding, and Xukai Chen. Effects of biofilms as the main and as a supplementary food on the survival, somatic growth and gonad enhancement of sea urchin Strongylocentrotus intermedius. Aquaculture International, 22(2):925–936, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9717-7.

Blay:2014:IND

[1281] Carole Blay, Manaarii Sham-Koua, Vincent Vonau, Roger Tetumu, Philippe Cabral, and Chin-Long Ky. Influence of nacre deposition rate on cultured pearl grade and colour in the black-lipped pearl oyster Pinctada margaritifera using farmed donor families. Aquaculture International, 22 (2):937–953, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9719-5.

Jobling:2014:BRW

[1282] Malcolm Jobling. Book review: W. McLarney: Freshwater aquaculture. Aquaculture International, 22(2):955–957, April 2014. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9652-7.

Jobling:2014:BRP

[1283] Malcolm Jobling. Book review: A. Petrie and P. Watson: Statistics for veterinary and animal science (third edition). Aquaculture International, 22(2):959–960, April 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9681-2.

Jobling:2014:BRG

[1284] Malcolm Jobling. Book review: G. Syda Rao, Imelda–Joseph, K. K. Philipose and M. Suresh Kumar: Cage aquaculture in India. Aquaculture International, 22(2):961–962, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9685-y.

Jobling:2014:BRM

[1285] Malcolm Jobling. Book review: M. Cargill and P. O’Connor: Writing scientific research articles: strategy and steps (2nd edition). Aquaculture International, 22(2):963–964, April 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9689-7.

Bulut:2014:EDP

[1286] Musa Bulut, Murat Yiğit, Sebahattin Ergün, Osman Sabri Kesbiç, Ümit Acar, Nejdet Gültepe, Mustafa Karga, Sevdan Yılmaz, and Derya Güroy. Evaluation of dietary protein and lipid requirements of two-banded seabream (Diplodus vulgaris) cultured in a recirculating aquaculture system. Aquaculture International, 22(3):965–973, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9720-z.

Qiang:2014:OCC

[1287] J. Qiang, J. He, P. Xu, M. D. Kpundeh, and Z. X. Zhu. Optimization of culture conditions for larval GIFT tilapia Oreochromis niloticus using response surface methodology and effects of HAMP-1 and ctype lysozyme mRNA expression in liver. Aquaculture International, 22 (3):975–991, June 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9721-y.

Luan:2014:SRS

[1288] Sheng Luan, Guoliang Yang, Junyi Wang, Kun Luo, Xuefeng Chen, Qiang Gao, Honglang Hu, and Jie Kong. Selection responses in survival of Macrobrachium rosenbergii after performing five generations of multi-trait selection for growth and survival. Aquaculture International, 22(3):993–1007, June 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9722-x.

Furtado:2014:ADD

[1289] Plínio S. Furtado, Carlos A. P. Gaona, Luis H. Poersch, and Wilson Wasielesky. Application of different doses of calcium hydroxide in the farming shrimp Litopenaeus vannamei with the biofloc technology (BFT). Aquaculture International, 22(3):1009–1023, June 2014. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9723-9.

Bayot:2014:EIC

[1290] Bonny Bayot, Jenny Rodríguez, Wilfrido Arguello, María Herminia Cornejo-Rodríguez, and Stanislaus Sonnenholzner. An evaluation of intraspecific competition for Pacific white shrimp Penaeus vannamei (Boone) in extensive/semi-intensive ponds. Aquaculture International, 22(3):1025–1039, June 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9724-8.

Iglesias:2014:EUC

[1291] J. Iglesias, G. Pazos, J. Fernández, F. J. Sánchez, J. J. Otero, P. Domingues, M. J. Lago, and F. Linares. The effects of using crab zoeae (Maja brachydactyla) on growth and biochemical composition of Octopus vulgaris (Cuvier 1797) paralarvae. Aquaculture International, 22(3):1041–1051, June 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9725-7.

Lalrinsanga:2014:YCM

[1292] P. L. Lalrinsanga, Bindu R. Pillai, Gunamaya Patra, Swagatika Mohanty, Namita Kumari Naik, Rashmi Ranjan Das, Sovan Sahu, and Rajesh Nelliyoura. Yield characteristics and morphometric relationships of giant freshwater prawn, Macrobrachium rosenbergii (de Man, 1879). Aquaculture International, 22(3):1053–1066, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9726-6.

Cui:2014:EEE

[1293] Yanting Cui, Bo Liu, Jun Xie, Pao Xu, Yuanyuan Zhang, and Jianhua Ming. Effect of enrofloxacin and emodin on heat-shock proteins’ expression in hepatic cells of grass carp (Ctenopharyngodon idellus). Aquaculture International, 22(3):1067–1077, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9727-5.

Vaseeharan:2014:MPD

[1294] Baskaralingam Vaseeharan and Rajagopalan Thaya. Medicinal plant derivatives as immunostimulants: an alternative to chemotherapeutics and antibiotics in aquaculture. Aquaculture International, 22(3):1079– 1091, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-013-9729-3.

Amirkolaie:2014:PSO

[1295] A. Keramat Amirkolaie, M. Darvishy Shahkolaie, S. Karimzadeh, and M. K. Khalesi. The potential of soya oil industry products as oil alternatives in rainbow trout (Oncorhynchus mykiss) diet. Aquaculture International, 22(3):1093–1103, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9730-x.

Rupia:2014:EHI

[1296] Emmanuel Joseph Rupia, Jiaojiao Shen, Jiamin Wu, Wenyin Chen, Liping Liu, Kristof Dierckens, Patrick Sorgeloos, and Weiqun Lu. Effect of hormone injection frequency on the lipid content and fatty acid compositions in gonad, muscle and liver of Anguilla japonica during artificial maturation. Aquaculture International, 22(3):1105–1120, June 2014. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9731-9.

Mizanur:2014:EFR

[1297] Rahman M. Mizanur, Gunhyun Park, Hyeonho Yun, Seunghan Lee, Sera Choi, and Sungchul C. Bai. The effects of feeding rates in juvenile Korean rockfish, (Sebastes schlegeli) reared at 17^{o}C and 20^{o}C water temperatures. Aquaculture International, 22(3):1121–1130, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-013-9732-8.

Huang:2014:PEP

[1298] Jung-Fu Huang and Jie-Min Lee. Production economics and profitability analysis of horizontal rack culture and horizontal rack culture coupled with raft-string culture methods: a case study of Pacific oyster (Crassostrea gigas) farming in Chiayi and Yunlin Counties, Taiwan. Aquaculture International, 22(3):1131–1147, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9733-7.

Wang:2014:CEA

[1299] Qing lin Wang, Shan shan Yu, Chuan xin Qin, Shuang lin Dong, and Yun wei Dong. Combined effects of acute thermal and hypo-osmotic stresses on osmolality and hsp70, hsp90 and sod expression in the sea cucumber Apostichopus japonicus Selenka. Aquaculture International, 22(3):1149–1161, June 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9734-6.

Pradeep:2014:OCC

[1300] Padmaja Jayaprasad Pradeep, Thekkeparambil Chandrabose Srijaya, Anuar Hassan, Anil Kumar Chatterji, Boonsirm Withyachumnarnkul, and Andrew Jeffs. Optimal conditions for cold-shock induction of triploidy in red tilapia. Aquaculture International, 22(3):1163–1174, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-013-9736-4.

Kim:2014:CVH

[1301] Wi-Sik Kim, Jong-Oh Kim, Jae-Kwon Cho, and Myung-Joo Oh. Change of viral hemorrhagic septicemia virus (VHSV) titer in olive flounder (Paralichthys olivaceus) following Poly(I:C) administration. Aquaculture International, 22(3):1175–1179, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9737-3.

Sombatjinda:2014:WQC

[1302] Siriphorn Sombatjinda, Chalermraj Wantawin, Somkiet Techkarnjanaruk, Boonsirm Withyachumnarnkul, and Marasri Ruengjitchatchawalya. Water quality control in a closed re-circulating system of Pacific white shrimp (Penaeus vannamei) postlarvae co-cultured with immobilized Spirulina mat. Aquaculture International, 22(3):1181–1195, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9738-2.

Aramli:2014:UMI

[1303] Mohammad Sadegh Aramli, Karim Golshahi, and Rajab Mohammad Nazari. Use of minimally invasive surgical technique for egg removal from the beluga, Huso huso. Aquaculture International, 22(3):1197– 1201, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-013-9739-1.

Glaropoulos:2014:SBI

[1304] Alexios Glaropoulos, Vassilis M. Papadakis, Ioannis E. Papadakis, Antonia Georgara, and Maroudio Kentouri. Sea bream interactions toward the aquaculture net due to the presence of micro-fouling. Aquaculture International, 22(3):1203–1214, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-013-9741-7.

Comeau:2014:NVG

[1305] Luc A. Comeau and Jose M. F. Babarro. Narrow valve gaping in the invasive mussel Limnoperna securis: implications for competition with the indigenous mussel Mytilus galloprovincialis in NW Spain. Aquaculture International, 22(3):1215–1227, June 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9742-6.

Avendano:2014:RCA

[1306] Miguel Avendaño and Marcela Cantillánez. Reproductive cycle of Aulacomya ater [Bivalvia: Mytilidae (Molina 1782)] in Punta Arenas Cove (Antofagasta Region, Chile). Aquaculture International, 22(4):1229– 1244, August 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-013-9743-5.

Hlophe:2014:CSU

[1307] S. N. Hlophe and N. A. G. Moyo. A comparative study on the use of Pennisetum clandestinum and Moringa oleifera as protein sources in the diet of the herbivorous Tilapia rendalli. Aquaculture International, 22 (4):1245–1262, August 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-013-9744-4.

Li:2014:EOR

[1308] Xian Li, Jean-Paul Blancheton, Ying Liu, Sebastien Triplet, and Luigi Michaud. Effect of oxidation–reduction potential on performance of European sea bass (Dicentrarchus labrax) in recirculating aquaculture systems. Aquaculture International, 22(4):1263–1282, August 2014. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-013-9745-3.

Caipang:2014:STH

[1309] Christopher Marlowe A. Caipang, Effrosyni Fatira, Carlo C. Lazado, and Michail Pavlidis. Short-term handling stress affects the humoral immune responses of juvenile Atlantic cod, Gadus morhua. Aquaculture International, 22(4):1283–1293, August 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9746-2.

Drillet:2014:ECA

[1310] Guillaume Drillet, Rémi Maguet, Mohamed-Sofiane Mahjoub, François Roullier, and Mark James Fielding. Egg cannibalism in Acartia tonsa: effects of stocking density, algal concentration, and egg availability. Aquaculture International, 22(4):1295–1306, August 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9747-1.

Jafari:2014:EDS

[1311] Seyed Mohammad Ali Jafari, Mohammad Rabbani, Mozhgan Emtyazjoo, and Fatemeh Piryaei. Effect of dietary Spirulina platensis on fatty acid composition of rainbow trout (Oncorhynchus mykiss) fillet. Aquaculture International, 22(4):1307–1315, August 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-013-9748-0.

Ma:2014:FIP

[1312] Zhenhua Ma. Food ingestion, prey selectivity, feeding incidence, and performance of yellowtail kingfish Seriola lalandi larvae under constant and varying temperatures. Aquaculture International, 22(4):1317– 1330, August 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-013-9749-z.

Imsland:2014:PTE

[1313] Albert Kjartansson Imsland, Sigurd Olav Handeland, and Sigurd Olav Stefansson. Photoperiod and temperature effects on growth and maturation of pre- and post-smolt Atlantic salmon. Aquaculture International, 22(4):1331–1345, August 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9750-1.

Fraser:2014:ETE

[1314] T. W. K. Fraser, I. Mayer, J. E. Skjæraasen, T. Hansen, and P. G. Fjelldal. The effect of triploidy on the efficacy and physiological response to anesthesia with MS 222 and isoeugenol in Atlantic salmon post-smolts. Aquaculture International, 22(4):1347–1359, August 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9751-0.

Peaydee:2014:IAH

[1315] Phacharakamon Peaydee, Sirawut Klinbunga, Piamsak Menasveta, Pikul Jiravanichpaisal, and Narongsak Puanglarp. An involvement of aquaporin in heat acclimation and cross-tolerance against ammonia stress in black tiger shrimp, Penaeus monodon. Aquaculture International, 22(4): 1361–1375, August 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9752-z.

Suloma:2014:CRB

[1316] A. Suloma, O. M. El-Husseiny, M. I. Hassane, R. S. Mabroke, and E. R. El-Haroun. Complementary responses between hydrolyzed feather meal, fish meal and soybean meal without amino acid supplementation in Nile tilapia Oreochromis niloticus diets. Aquaculture International, 22(4): 1377–1390, August 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9753-y.

deCarvalho:2014:FGJ

[1317] Cristina Vaz Avelar de Carvalho, Gabriel Passini, Wanessa de Melo Costa, and Vinicius Ronzani Cerqueira. Feminization and growth of juvenile fat snook Centropomus parallelus fed diets with different concentrations of estradiol-17β. Aquaculture International, 22(4):1391–1401, August 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9754-x.

Zhang:2014:NSB

[1318] Libin Zhang, Yan Gao, Tao Zhang, Hongsheng Yang, Qiang Xu, Lina Sun, and Zonghe Yu. A new system for bottom co-culture of the scallop, Patinopecten yessoensis, with the sea cucumber, Apostichopus japonicus, and the sea urchin, Anthocidaris crassispina, in shallow water in China. Aquaculture International, 22(4):1403–1415, August 2014. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9755-9.

Boonanuntanasarn:2014:ESA

[1319] Surintorn Boonanuntanasarn, Pranorm Khaomek, Taratip Pitaksong, and Yanling Hua. The effects of the supplementation of activated charcoal on the growth, health status and fillet composition-odor of Nile tilapia (Oreochromis niloticus) before harvesting. Aquaculture International, 22(4):1417–1436, August 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9756-8.

Hu:2014:DSR

[1320] Qiaomu Hu, Song-Lin Chen, Fengtao Gao, Yangzhen Li, Shanshan Liu, Feng Liu, Jingfeng Yang, and Min Du. Differences in sex reversion and growth between normal and neomale stock in half-smooth tongue sole, Cynoglossus semilaevis. Aquaculture International, 22(4): 1437–1449, August 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9757-7.

Babu:2014:MAB

[1321] Padinhate Purayil Suresh Babu, Kalkuli Mariappa Shankar, Badami Ramalingappa Honnananda, and Purandara Ballyaya Abhiman. Monoclonal antibody-based immunodot for evaluating serum immunoglobulin levels in rohu (Labeo rohita). Aquaculture International, 22(4): 1451–1456, August 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9758-6.

Yue:2014:DTR

[1322] Yirong Yue, Zhiying Zou, Jinglin Zhu, Dayu Li, Wei Xiao, Jue Han, and Hong Yang. Dietary threonine requirement of juvenile Nile tilapia, Oreochromis niloticus. Aquaculture International, 22(4):1457–1467, August 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-014-9759-5.

Ghasemi:2014:PCL

[1323] Seyed Mahdi Ghasemi, Majid Bouzari, and Giti Emtiazi. Preliminary characterization of Lactococcus garvieae bacteriophage isolated from wastewater as a potential agent for biological control of lactococcosis in aquaculture. Aquaculture International, 22(4):1469–1480, August 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9760-z.

Liu:2014:QDI

[1324] Fu-Jia Liu, Yong-Jian Liu, Li-Xia Tian, Xue-Fei Li, Zhi-Hao Zhang, Hui-Jun Yang, and Zhen-Yu Du. Quantitative dietary isoleucine requirement of juvenile Pacific white shrimp, Litopenaeus vannamei (Boone) reared in low-salinity water. Aquaculture International, 22(4):1481– 1497, August 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9761-y.

Mirera:2014:SEI

[1325] David Oersted Mirera, Jacob Ochiewo, and Fridah Munyi. Social and economic implications of small-scale mud crab (Scylla serrata) aquaculture: the case of organised community groups. Aquaculture International, 22 (4):1499–1514, August 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9762-x.

El-Magsodi:2014:HNQ

[1326] Mohamed Omar El-Magsodi, Peter Bossier, Patrick Sorgeloos, and Gilbert Van Stappen. Hatching and nutritional quality of Artemia cysts progressively deteriorates as a function of increased exposure to hydration/dehydration cycles. Aquaculture International, 22(5):1515– 1532, October 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9763-9.

Zuniga-Jara:2014:BMG

[1327] Sergio Zuniga-Jara and María C. Marín-Riffo. A bioeconomic model of a genetic improvement program of abalone. Aquaculture International, 22 (5):1533–1562, October 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9764-8.

Ghiasi:2014:ETI

[1328] Sareh Ghiasi, Bahram Falahatkar, Konrad Dabrowski, Alireza Abasalizadeh, and Murat Arslan. Effect of thiamine injection on growth performance, hematology and germinal vesicle migration in sterlet sturgeon Acipenser ruthenus L. Aquaculture International, 22(5):1563– 1576, October 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9765-7.

Li:2014:CAT

[1329] Xilei Li, Zhiyi Bai, Hongrui Luo, Guiling Wang, and Jiale Li. Comparative analysis of total carotenoid content in tissues of purple and white inner-shell color pearl mussel, Hyriopsis cumingii. Aquaculture International, 22(5):1577–1585, October 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9766-6.

Yoshimatsu:2014:RAH

[1330] Takao Yoshimatsu and M. Amzad Hossain. Recent advances in the high-density rotifer culture in Japan. Aquaculture International, 22(5): 1587–1603, October 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9767-5.

Valentim-Neto:2014:DEP

[1331] Pedro A. Valentim-Neto, Ana P. M. Fraga, and Maria R. F. Marques. Differential expression of proteins in the gills of Litopenaeus vannamei infected with white spot syndrome virus. Aquaculture International, 22 (5):1605–1620, October 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9768-4.

Pantazis:2014:EWD

[1332] P. A. Pantazis, G. Benekos, and G. Papadomichelakis. Early-weaning diets for gilthead sea bream (Sparus aurata L.) and their potential use in Hellenic marine fish hatcheries. Aquaculture International, 22(5): 1621–1636, October 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9769-3.

Sen:2014:IRP

[1333] Shib Sankar Sen, Sib Sankar Giri, and V. Sukumaran. Immune responses and protection in rohu vaccinated against Aeromonas hydrophila infection. Aquaculture International, 22(5):1637–1648, October 2014. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9770-x.

Brito:2014:WQP

[1334] Luis Otavio Brito, Luis Alejandro Vinatea Arana, Roberta Borda Soares, William Severi, Rayzza Helena Miranda, Suzianny Maria Bezerra Cabral da Silva, Maria Raquel Moura Coimbra, and Alfredo Olivera Gálvez. Water quality, phytoplankton composition and growth of Litopenaeus vannamei (Boone) in an integrated biofloc system with Gracilaria birdiae (Greville) and Gracilaria domingensis (Kützing). Aquaculture International, 22(5):1649–1664, October 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9771-9.

Liu:2014:BET

[1335] Fuli Liu, Xiutao Sun, Feijiu Wang, Wenjun Wang, Zhourui Liang, Zhelong Lin, and Zhian Dong. Breeding, economic traits evaluation, and commercial cultivation of a new Saccharina variety “Huangguan No. 1”. Aquaculture International, 22(5):1665–1675, October 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-014-9772-8.

Giangrande:2014:REP

[1336] Adriana Giangrande, Cataldo Pierri, Giovanni Fanelli, Roberto Schirosi, Margherita Licciano, and Loredana Stabili. Rearing experiences of the polychaete Sabella spallanzanii in the Gulf of Taranto (Mediterranean Sea, Italy). Aquaculture International, 22(5):1677–1688, October 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9773-7.

Ky:2014:ISG

[1337] Chin-Long Ky, Nicolas Molinari, Elisabeth Moe, and Steve Pommier. Impact of season and grafter skill on nucleus retention and pearl oyster mortality rate in Pinctada margaritifera aquaculture. Aquaculture International, 22(5):1689–1701, October 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9774-6.

Srijila:2014:RRC

[1338] C. K. Srijila, A. M. Babitha Rani, P. Girish Babu, and V. K. Tiwari. Ration restriction, compensatory growth and pituitary growth hormone gene expression in Labeo rohita. Aquaculture International, 22(5): 1703–1710, October 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9775-5.

Schamber:2014:GPB

[1339] Christiano Rodrigues Schamber, Wilson Rogério Boscolo, Maria Raquel Marçal Natali, Mariana Michelato, Valéria Rossetto Barriviera Furuya, and Wilson Massamitu Furuya. Growth performance and bone mineralization of large Nile tilapia (Oreochromis niloticus) fed graded levels of available phosphorus. Aquaculture International, 22(5):1711– 1721, October 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9776-4.

Gonzalez-Rodriguez:2014:EVP

[1340] Á. González-Rodríguez, J. D. Celada, J. M. Carral, M. Sáez-Royuela, and J. B. Fuertes. Effects of varying protein level in practical diets on survival, growth, feed utilization and body composition of juvenile tench (Tinca tinca L.). Aquaculture International, 22(5):1723–1735, October 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9777-3.

Zhao:2014:HPC

[1341] Chong Zhao, Ping Sun, Haisen Zhou, Xiaofei Tian, Wenping Feng, and Yaqing Chang. Heritability and phenotypic correlations of gonad sweetness in the sea urchin Strongylocentrotus intermedius. Aquaculture International, 22(6):1737–1742, December 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9778-2.

Chambel:2014:HPI

[1342] João Chambel, Ricardo Costa, Mónica Gomes, Susana Mendes, Teresa Baptista, and Rui Pedrosa. Hydrogen peroxide, iodine solution and methylene solution highly enhance the hatching rate of freshwater ornamental fish species. Aquaculture International, 22(6):1743–1751, December 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-014-9779-1.

Zhao:2014:EYP

[1343] Wen Zhao, Miao Liang, Qing Liu, Xunwang Yin, and Jie Wei. The effect of yeast polysaccharide (YSP) on the immune function of Apostichopus japonicus Selenka under salinity stress. Aquaculture International, 22(6): 1753–1766, December 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9780-8.

Degremont:2014:MAO

[1344] Lionel Degremont and Abdellah Benabdelmouna. Mortality associated with OsHV-1 in spat Crassostrea gigas: role of wild-caught spat in the horizontal transmission of the disease. Aquaculture International, 22(6): 1767–1781, December 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9781-7.

Guo:2014:SFW

[1345] Kai Guo, Wen Zhao, Shan Wang, and Shuang Lin Dong. Study of food web structure and trophic level in the sea ponds of an optimized culture model (jellyfish–shellfish–fish–prawn). Aquaculture International, 22(6): 1783–1791, December 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9782-6.

Das:2014:MSD

[1346] Sweta Das and P. K. Sahoo. Markers for selection of disease resistance in fish: a review. Aquaculture International, 22(6):1793–1812, December 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9783-5.

Hannon:2014:CML

[1347] Colin Hannon, Rick A. Officer, Jean Le Dorven, and John Chamberlain. Culture methods of live algal feeds for European aquaculture: optimising culture conditions for Ulvella lens. Aquaculture International, 22(6): 1813–1822, December 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9784-4.

Su:2014:IGD

[1348] Xionggao Su, Xiaoqin Li, Xiangjun Leng, Chonggui Tan, Bo Liu, Xianqi Chai, and Ting Guo. The improvement of growth, digestive enzyme activity and disease resistance of white shrimp by the dietary citric acid. Aquaculture International, 22(6):1823–1835, December 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-014-9785-3.

Llorente:2014:EOS

[1349] Ignacio Llorente and Ladislao Luna. Economic optimisation in seabream (Sparus aurata) aquaculture production using a particle swarm optimisation algorithm. Aquaculture International, 22(6):1837–1849, December 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-014-9786-2.

Xie:2014:EDT

[1350] Zhongguo Xie, Furong Wang, Haiying Liu, Shidong Guo, Huilai Shi, Wei Zhan, and Bao Lou. Effect of dietary taurine levels on growth performance and taurine content of Nibea albiflora larvae. Aquaculture International, 22(6):1851–1862, December 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9787-1.

Hao:2014:EHT

[1351] Zhen-Lin Hao, Xue-Jiao Tang, Jun Ding, Yue Ben, and Ya-Qing Chang. Effect of high temperature on survival, oxygen consumption, behavior, ammonia-n excretion, and related immune indicators of the Japanese scallop Mizuhopecten yessoensis. Aquaculture International, 22(6):1863– 1876, December 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9788-0.

Zhong:2014:QMG

[1352] Xiaoxiao Zhong, Qi Li, Xiang Guo, Hong Yu, and Lingfeng Kong. QTL mapping for glycogen content and shell pigmentation in the Pacific oyster Crassostrea gigas using microsatellites and SNPs. Aquaculture International, 22(6):1877–1889, December 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9789-z.

Chavez-Sanchez:2014:AAR

[1353] María Cristina Chávez-Sánchez, Miguel Angel Olvera-Novoa, Briseida Osuna-Durán, Isabel Abdo de la Parra, Selene María Abad-Rosales, and Irma Martínez-Rodríguez. Ascorbic acid requirement and histopathological changes due to its deficiency in juvenile spotted rose snapper Lutjanus guttatus (Steindachner, 1869). Aquaculture International, 22(6): 1891–1909, December 2014. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9790-6.

Anzueto-Sanchez:2014:EFC

[1354] M. A. Anzueto-Sánchez, B. Barón-Sevilla, B. Cordero-Esquivel, and A. Celaya-Ortega. Effects of food concentration and temperature on development, growth, reproduction and survival of the copepod Pseudodiaptomus euryhalinus. Aquaculture International, 22(6):1911–1923, December 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9791-5.

deBarros:2014:EVP

[1355] Leonardo Castilho de Barros, Oscar José Sallée Barreto, and Marcelo Barbosa Henriques. The economic viability for the production of live baits of white shrimp (Litopenaeus schmitti) in recirculation culture system. Aquaculture International, 22(6):1925–1935, December 2014. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-014-9792-4.

Ak:2015:ESA

[1356] İlknur Ak, Cenk Öztaşkent, Yeliz Özüdoğru, and Tolga Göksan. Effect of sodium acetate and sodium nitrate on biochemical composition of green algae Ulva rigida. Aquaculture International, 23(1):1–11, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9793-3.

Nowosad:2015:CGE

[1357] Joanna Nowosad, Dariusz Kucharczyk, Joanna Łuczyńska, Katarzyna Targońska, Tomasz Kajetan Czarkowski, Maria Biłas, Sławomir Krejszeff, László Horváth, and Tamás Müller. Changes in European eel ovary development and body and ovary chemistry during stimulated maturation under controlled conditions: preliminary data. Aquaculture International, 23(1):13–27, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9794-2.

Buen-Ursua:2015:IRP

[1358] Shelah Mae A. Buen-Ursua, Teruo Azuma, Katsutoshi Arai, and Relicardo M. Coloso. Improved reproductive performance of tiger tail seahorse, Hippocampus comes, by mysid shrimp fed singly or in combination with other natural food. Aquaculture International, 23(1): 29–43, February 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9795-1.

Nowosad:2015:CTS

[1359] Joanna Nowosad, Dariusz Kucharczyk, Tomasz Liszewski, Katarzyna Targońska, and Roman Kujawa. Comparison of temperature shock timing to induced artificial mitotic gynogenesis and androgenesis in common tench. Aquaculture International, 23(1):45–53, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-014-9796-0.

Yarahmadi:2015:ESD

[1360] Peyman Yarahmadi, Hamed Kolangi Miandare, Seyed Hossein Hoseinifar, Nahid Gheysvandi, and Arash Akbarzadeh. The effects of stocking density on hemato-immunological and serum biochemical parameters of rainbow trout (Oncorhynchus mykiss). Aquaculture International, 23 (1):55–63, February 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9797-z.

Velmurugan:2015:SCA

[1361] S. Velmurugan, N. Jerin, M. Michael Babu, F. Bindhu, S. Albin Dhas, and T. Citarasu. Screening and characterization of antiviral compounds from Enteromorpha flexuosa against white spot syndrome virus (WSSV) and its in vivo influence on Indian white shrimp Fenneropenaeus indicus. Aquaculture International, 23(1):65–80, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-014-9798-y.

Li:2015:SGC

[1362] Ya-Juan Li, Bo Liu, He Zhou, Yang-Chun Gao, Min Gao, Yu-Sheng Wang, Ding-Chen Cao, and Xiao-Wen Sun. Study of the genomic constitution of the German mirror carp. Aquaculture International, 23(1): 81–91, February 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9799-x.

Oh:2015:QCR

[1363] So-Young Oh, Wi-Sik Kim, Myung-Joo Oh, and Toyohiko Nishizawa. Quantitative change of red seabream iridovirus (RSIV) in rock bream Oplegnathus fasciatus, following Poly(I:C) administration. Aquaculture International, 23(1):93–98, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9800-8.

Jiang:2015:EDL

[1364] Yu dong Jiang, Ji teng Wang, Tao Han, Xin yu Li, and Shui xin Hu. Effect of dietary lipid level on growth performance, feed utilization and body composition by juvenile red spotted grouper (Epinephelus akaara). Aquaculture International, 23(1):99–110, February 2015. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9801-7.

Sartori:2015:ERP

[1365] D. Sartori, A. Scuderi, G. Sansone, and A. Gaion. Echinoculture: the rearing of Paracentrotus lividus in a recirculating aquaculture system — experiments of artificial diets for the maintenance of sexual maturation. Aquaculture International, 23(1):111–125, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-014-9802-6.

Vizzini:2015:UFV

[1366] Salvatrice Vizzini, Luca Miccichè, Antonino Vaccaro, and Antonio Mazzola. Use of fresh vegetable discards as sea urchin diet: effect on gonad index and quality. Aquaculture International, 23(1):127–139, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9803-5.

Tahil:2015:ERP

[1367] Abduraji S. Tahil and Danilo T. Dy. Effects of reduced pH on the growth and survival of postlarvae of the donkey’s ear abalone, Haliotis asinina (L.). Aquaculture International, 23(1):141–153, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-014-9804-4.

Mirera:2015:CPW

[1368] David Oersted Mirera and Per-Olav Moksnes. Comparative performance of wild juvenile mud crab (Scylla serrata) in different culture systems in East Africa: effect of shelter, crab size and stocking density. Aquaculture International, 23(1):155–173, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9805-3.

Oh:2015:FFI

[1369] Sung-Yong Oh and B. A. Venmathi Maran. Feeding frequency influences growth, feed consumption and body composition of juvenile rock bream (Oplegnathus fasciatus). Aquaculture International, 23(1): 175–184, February 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9806-2.

Kaldre:2015:ETD

[1370] Katrin Kaldre, Kerli Haugjärv, Mari Liiva, and Riho Gross. The effect of two different feeds on growth, carapace colour, maturation and mortality in marbled crayfish (Procambarus fallax f. virginalis). Aquaculture International, 23(1):185–194, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9807-1.

Ravi:2015:ETM

[1371] M. Ravi and A. S. Sahul Hameed. Experimental transmission of Macrobrachium rosenbergii nodavirus (Mr NV ) and extra small virus (XSV) in Macrobrachium malcolmsonii and Macrobrachium rude. Aquaculture International, 23(1):195–201, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9808-0.

Reda:2015:EBA

[1372] Rasha M. Reda and Khaled M. Selim. Evaluation of Bacillus amyloliquefaciens on the growth performance, intestinal morphology, hematology and body composition of Nile tilapia, Oreochromis niloticus. Aquaculture International, 23(1):203–217, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9809-z.

Zhang:2015:IGF

[1373] Sicong Zhang, Hongsheng Yang, Shilin Liu, Libin Zhang, Xiaoshang Ru, and Dongxue Xu. The influence of genetics factor on key growth traits and quantitative genetic analysis of sea cucumber Apostichopus japonicus (Selenka) heat-resistant and fast-growing strain. Aquaculture International, 23(1):219–233, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9810-6.

Sun:2015:ETS

[1374] Yu-Ping Sun, Li-Zeng Guan, Jie-Hua Xiong, Qian-Yun Xi, and Yong-Liang Zhang. Effects of l-tryptophan-supplemented dietary on growth performance and 5-HT and GABA levels in juvenile Litopenaeus vannamei. Aquaculture International, 23(1):235–251, February 2015. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9811-5.

Cassiano:2015:UPA

[1375] Eric J. Cassiano, Matthew L. Wittenrich, Thomas B. Waltzek, Natalie K. Steckler, Kevin P. Barden, and Craig A. Watson. Utilizing public aquariums and molecular identification techniques to address the larviculture potential of Pacific blue tangs (Paracanthurus hepatus), semicircle angelfish (Pomacanthus semicirculatus), and bannerfish (Heniochus sp.). Aquaculture International, 23(1):253–265, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-014-9813-3.

Rohmana:2015:WQP

[1376] Dasu Rohmana, Enang Harris Surawidjaja, Sukenda Sukenda, and Julie Ekasari. Water quality and production performance of catfish–prawn co-culture with organic carbon source addition. Aquaculture International, 23(1):267–276, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9814-2.

Dan:2015:EEP

[1377] Shigeki Dan and Katsuyuki Hamasaki. Evaluation of the effects of probiotics in controlling bacterial necrosis symptoms in larvae of the mud crab Scylla serrata during mass seed production. Aquaculture International, 23(1):277–296, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9815-1.

Pucher:2015:PMS

[1378] Johannes Pucher, Richard Mayrhofer, Mansour El-Matbouli, and Ulfert Focken. Pond management strategies for small-scale aquaculture in northern Vietnam: fish production and economic performance. Aquaculture International, 23(1):297–314, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9816-0.

Furtado:2015:ENT

[1379] Plínio S. Furtado, Bruno R. Campos, Fabiane P. Serra, Marta Klosterhoff, Luis A. Romano, and Wilson Wasielesky. Effects of nitrate toxicity in the Pacific white shrimp, Litopenaeus vannamei, reared with biofloc technology (BFT). Aquaculture International, 23(1):315–327, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9817-z.

Hadipour:2015:LSE

[1380] Abouzar Hadipour, Freydoon Vafaie, and Vahid Hadipour. Land suitability evaluation for brackish water aquaculture development in coastal area of Hormozgan, Iran. Aquaculture International, 23(1):329–343, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9818-y.

Furtado:2015:EDA

[1381] Plínio S. Furtado, Luis H. Poersch, and Wilson Wasielesky, Jr. The effect of different alkalinity levels on Litopenaeus vannamei reared with biofloc technology (BFT). Aquaculture International, 23(1):345–358, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9819-x.

Janson:2015:PGS

[1382] S. Janson, J. Wouters, M. Bonow, I. Svanberg, and K. H. Olsén. Population genetic structure of crucian carp (Carassius carassius) in man-made ponds and wild populations in Sweden. Aquaculture International, 23 (1):359–368, February 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9820-4.

Hussain:2015:EWF

[1383] Tanveer Hussain, A. K. Verma, V. K. Tiwari, Chandra Prakash, G. Rathore, A. P. Shete, and Neelam Saharan. Effect of water flow rates on growth of Cyprinus carpio var. koi (Cyprinus carpio L., 1758) and spinach plant in aquaponic system. Aquaculture International, 23 (1):369–384, February 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9821-3.

Chaweepack:2015:EGA

[1384] Tidaporn Chaweepack, Surachart Chaweepack, Boonyee Muenthaisong, Lila Ruangpan, Kei Nagata, and Kaeko Kamei. Effect of galangal (Alpinia galanga Linn.) extract on the expression of immune-related genes and Vibrio harveyi resistance in Pacific white shrimp (Litopenaeus vannamei). Aquaculture International, 23(1):385–399, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9822-2.

Lalrinsanga:2015:RAL

[1385] P. L. Lalrinsanga, Bindu R. Pillai, Kanta D. Mahapatra, Lopamudra Sahoo, Raul W. Ponzoni, Nguyen Hong Nguyen, Swagatika Mohanty, Swagatika Sahu, Sovan Sahu, Vijay Kumar, Gunamaya Patra, and Shivani Patnaik. Retracted article: Length–weight relationship and condition factor of nine possible crosses of three stocks of giant freshwater prawn, Macrobrachium rosenbergii from different agro-ecological regions of India. Aquaculture International, 23(1):401, February 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-012-9595-4.

Obirikorang:2015:LAI

[1386] Kwasi Adu Obirikorang, Stephen Amisah, Simon Cudjoe Fialor, and Peter Vilhelm Skov. Local agro-industrial by-products with potential use in Ghanaian aquaculture: a review. Aquaculture International, 23 (2):403–425, April 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9831-1.

Winkelbach:2015:COR

[1387] Anja Winkelbach, Rüdiger Schade, Carsten Schulz, and Sven Wuertz. Comparison of oral, rectal and intraperitoneal administration of IgY antibodies in passive immunization of rainbow trout (Oncorhynchus mykiss). Aquaculture International, 23(2):427–438, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9823-1.

Meshram:2015:EDG

[1388] S. J. Meshram, H. Shivananda Murthy, H. Ali, H. S. Swain, and Abhiman Ballyaya. Effect of dietary β-glucan on immune response and disease resistance against Aeromonas hydrophila in giant freshwater prawn, Macrobrachium rosenbergii (de Man. 1879). Aquaculture International, 23(2):439–447, April 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9824-0.

Suita:2015:DHQ

[1389] Sabrina Medeiros Suita, Alessandro Pereira Cardozo, Luis Alberto Romano, Paulo Cesar Abreu, and Wilson Wasielesky. Development of the hepatopancreas and quality analysis of post-larvae Pacific white shrimp Litopenaeus vannamei produced in a BFT system. Aquaculture International, 23(2):449–463, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9825-z.

Smith:2015:HMS

[1390] Peter Smith and Göran Kronvall. How many strains are required to set an epidemiological cut-off value for MIC values determined for bacteria isolated from aquatic animals? Aquaculture International, 23 (2):465–470, April 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9827-x.

Ju:2015:NIM

[1391] Bao Ju, Lihong Chen, Ronglian Xing, and Aili Jiang. A new integrated multi-trophic aquaculture system consisting of Styela clava, microalgae, and Stichopus japonicus. Aquaculture International, 23 (2):471–497, April 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9829-8.

Yang:2015:PDG

[1392] Min Yang, Wenjie Guo, Xu-Fang Liang, Changxu Tian, Liyuan Lv, Kecheng Zhu, and Hanping Wang. Parentage determination in golden mandarin fish (Siniperca scherzeri) based on microsatellite DNA markers. Aquaculture International, 23(2):499–507, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-014-9830-2.

Fugimura:2015:BRS

[1393] Michelle Midori Sena Fugimura, Helaine dos Reis Flor, Emanuela Paula de Melo, Tiago Viana da Costa, Wilson Wasielesky, and Lidia Miyako Yoshii Oshiro. Brewery residues as a source of organic carbon in Litopenaeus schmitti white shrimp farms with BFT systems. Aquaculture International, 23(2):509–522, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9832-0.

Trieu:2015:ICC

[1394] Tran Thi Ngoc Trieu and Nguyen Thanh Phong. The impact of climate change on salinity intrusion and Pangasius (Pangasianodon hypophthalmus) farming in the Mekong Delta, Vietnam. Aquaculture International, 23(2):523–534, April 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9833-z.

Lalonde:2015:CPC

[1395] Benoit A. Lalonde, William Ernst, and Christine Garron. Chemical and physical characterisation of effluents from land-based fish farms in Atlantic Canada. Aquaculture International, 23(2):535–546, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9834-y.

Flores-Miranda:2015:EDF

[1396] Ma. del Carmen Flores-Miranda, Antonio Luna-González, Diana Verónica Cortés-Espinosa, Píndaro Álvarez-Ruiz, Edilmar Cortés-Jacinto, Francisco Javier Valdez-González, Ruth Escamilla-Montes, and Héctor Abelardo González-Ocampo. Effects of diets with fermented duckweed (Lemna sp.) on growth performance and gene expression in the Pacific white shrimp, Litopenaeus vannamei. Aquaculture International, 23(2):547– 561, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9835-x.

daSilva:2015:EII

[1397] Suzianny Maria Bezerra Cabral da Silva, João Luís Rocha, Pedro Carlos Cunha Martins, Alfredo Olivera Gálvez, Fernando Leandro dos Santos, Humber Agrelli Andrade, and Maria Raquel Moura Coimbra. Experimental infection of infectious myonecrosis virus (IMNV) in the Pacific white shrimp Litopenaeus vannamei (Boone, 1931). Aquaculture International, 23(2):563–576, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9836-9.

Zehra:2015:DLR

[1398] Seemab Zehra and Mukhtar A. Khan. Dietary leucine requirement of fingerling Catla catla (Hamilton) based on growth, feed conversion ratio, RNA/DNA ratio, leucine gain, blood indices and carcass composition. Aquaculture International, 23(2):577–595, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9837-8.

Hassantabar:2015:CGD

[1399] Fatemeh Hassantabar, Abolghasem Esmaeili Fereidouni, Hossein Ouraji, Sedigheh Babaei, and Ali Jafarpour. Comparison of growth and digestive enzymes activity of kutum (Rutilus kutum) during ontogeny fed with live prey and artificial feed. Aquaculture International, 23 (2):597–612, April 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9838-7.

Panda:2015:SMW

[1400] Nirupama Panda, A. S. Mahapatra, and D. P. Rath. Statistical modeling to women self-help groups of Odisha based on socio-economic developmental parameters in aquaculture. Aquaculture International, 23 (2):613–637, April 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9839-6.

Imsland:2015:ASS

[1401] Albert K. Imsland, Patrick Reynolds, Gerhard Eliassen, Thor A. Hangstad, Ane V. Nytrø, Atle Foss, Erik Vikingstad, and Tor Anders Elvegård. Assessment of suitable substrates for lumpfish in sea pens. Aquaculture International, 23(2):639–645, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9840-0.

Darmody:2015:TTG

[1402] G. Darmody, A. P. Maloy, S. A. Lynch, M. Prado-Alvarez, J. Cotterill, T. Wontner-Smith, and S. C. Culloty. Tissue targeting of the European flat oyster, Ostrea edulis, using microencapsulated microbeads as a biological proxy. Aquaculture International, 23(2):647–659, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9842-y.

Wei:2015:PEP

[1403] Jianteng Wei, Shuxian Wang, Dong Pei, Yongfeng Liu, Yi Liu, and Duolong Di. Polysaccharide from Enteromorpha prolifera enhances non-specific immune responses and protection against Vibrio splendidus infection of sea cucumber. Aquaculture International, 23(2):661– 670, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9844-9.

Sutthinon:2015:JHG

[1404] Parinya Sutthinon, Karun Thongprajukaew, Suktianchai Saekhow, and Rinpawee Ketmanee. Juvenile hybrid grouper (Epinephelus coioides × E. lanceolatus) are euryhaline and can grow in a wide range of salinities. Aquaculture International, 23(2):671–682, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9845-8.

Zhang:2015:EDG

[1405] Yao bei-ping Zhang, Yi bin Zhou, Ba yi Sang, Xiao chun Wan, Yan ou Yang, Jian li Zhang, Thomas L. Welker, and Keshun Liu. Effect of dietary Chinese tea on growth performance, disease resistance and muscle fatty acid profile of channel catfish (Ictalurus punctatus). Aquaculture International, 23(2):683–698, April 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9846-7.

Sorgeloos:2015:F

[1406] Patrick Sorgeloos and Margaret Eleftheriou. Foreword. Aquaculture International, 23(3):699–701, June 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-015-9901-z.

Eleftheriou:2015:ATT

[1407] Margaret Eleftheriou, Marieke Reuver, John Bostock, Patrick Sorgeloos, and Jean Dhont. AQUA-TNET thematic network: an 18-year chronicle of development and achievement in European aquaculture education. Aquaculture International, 23(3):703–713, June 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9857-4.

Lekang:2015:ATO

[1408] Odd-Ivar Lekang, Michael Moulton, and Margaret Eleftheriou. AQUA-TNET: an overview of the development of European master programmes in the aquaculture sector. Aquaculture International, 23(3):715–726, June 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-014-9858-3.

Bossier:2015:DBD

[1409] Peter Bossier and Margaret Eleftheriou. Designing and building dedicated Ph.D. courses contributing to international EU mobility at doctoral level. Aquaculture International, 23(3):727–749, June 2015. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9841-z.

Eleftheriou:2015:PLLa

[1410] Margaret Eleftheriou and Sónia Seixas. Positioning lifelong learning in aquaculture: challenges and opportunities. Aquaculture International, 23(3):751–766, June 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9826-y.

Pita:2015:GSN

[1411] Cristina Pita, Margaret Eleftheriou, Jaime Fernández-Borrás, Susana Gonçalves, Eleni Mente, M. Begoña Santos, Sónia Seixas, and Graham J. Pierce. Generic skills needs for graduate employment in the aquaculture, fisheries and related sectors in Europe. Aquaculture International, 23 (3):767–786, June 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9843-x.

Tort:2015:SMM

[1412] Lluis Tort and Margaret Eleftheriou. Student mobility measures in the aquatic sciences: the development of the AQUA–TNET education gate. Aquaculture International, 23(3):787–803, June 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9896-5.

Eleftheriou:2015:PLLb

[1413] Margaret Eleftheriou and Annie Heral. The promotion of language learning and linguistic diversity through the aquatic sciences ERAS-MUS network AQUA–TNET. Aquaculture International, 23(3):805– 823, June 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9812-4.

Seixas:2015:EUL

[1414] Sónia Seixas, Clive Dove, Bernd Ueberschär, and John Bostock. Evaluation on the use of e-learning tools to support teaching and learning in aquaculture and aquatic sciences education. Aquaculture International, 23(3):825–841, June 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9828-9.

Seixas:2015:IEC

[1415] Sónia Seixas, Sathappan Saravanan, and Susana Gonçalves. Innovation and educational changes: two e-learning cases in aquaculture. Aquaculture International, 23(3):843–859, June 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9850-y.

Bostock:2015:IHC

[1416] John Bostock and Sónia Seixas. Investing in the human capital of the aquatic food sector: AQUA–TNET and the road ahead. Aquaculture International, 23(3):861–881, June 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-015-9915-6.

Jia:2015:HAE

[1417] Rui Jia, Jin-Liang Du, Li-Ping Cao, Ying-Juan Liu, Pao Xu, and Guo-Jun Yin. Hepatoprotective and antioxidant effects of phyllanthin against carbon tetrachloride-induced liver injury in Cyprinus carpio. Aquaculture International, 23(4):883–893, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9847-6.

Butts:2015:RWF

[1418] Ian A. E. Butts, Ines Ben Khemis, and Matthew K. Litvak. Receptivity of winter flounder larvae to artificial diet from the yolk-sac stage to metamorphosis. Aquaculture International, 23(4):895–901, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9848-5.

Feng:2015:EIG

[1419] Wenping Feng, Yaqing Chang, Chong Zhao, Ping Sun, and Jing Wei. Effects of inbreeding on growth, gametogenesis, gonad production, quality and MYP expression in the sea urchin Strongylocentrotus intermedius. Aquaculture International, 23(4):903–912, August 2015. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9849-4.

Li:2015:PRP

[1420] Zhimin Li, Li Xu, Wenshu Liu, Yuchun Liu, Einar Ringø, Zhenyu Du, and Zhigang Zhou. Protein replacement in practical diets altered gut allochthonous bacteria of cultured cyprinid species with different food habits. Aquaculture International, 23(4):913–928, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-014-9851-x.

Hache:2015:SNP

[1421] Rémy Haché, Claude J. Pelletier, and André Dumas. Selected nutrient profiles in first larvae and postlarvae of American lobster (Homarus americanus). Aquaculture International, 23(4):929–941, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9852-9.

Karakatsouli:2015:ERD

[1422] Nafsika Karakatsouli, Nikolaos Kassianos, and Sofronios E. Papoutsoglou. Effects of rearing density and tank colour on juvenile sharpsnout seabream (Diplodus puntazzo) growth performance. Aquaculture International, 23(4):943–953, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9853-8.

Li:2015:EDC

[1423] Junwei Li, Shuanglin Dong, Xiangli Tian, Ce Shi, Fang Wang, Qinfeng Gao, and Changbo Zhu. Effects of the diatom Cylindrotheca fusiformis on the growth of the sea cucumber Apostichopus japonicus and water quality in ponds. Aquaculture International, 23(4):955–965, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9854-7.

Zhong:2015:ILV

[1424] Dasen Zhong, Fang Wang, Shuanglin Dong, and Li Li. Impact of Litopenaeus vannamei bioturbation on nitrogen dynamics and benthic fluxes at the sediment–water interface in pond aquaculture. Aquaculture International, 23(4):967–980, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9855-6.

Yang:2015:DLC

[1425] Wenping Yang, Aimin Wang, Feng Gao, Yebing Yu, Linlan Lv, and Fu Lv. Dietary lipid concentrations influence growth and chemical and fatty acid compositions of juvenile redlip mullet, Liza haematocheila. Aquaculture International, 23(4):981–996, August 2015. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9856-5.

Seemann:2015:GPN

[1426] Uli B. Seemann, Kai Lorkowski, Matthew J. Slater, Friedrich Buchholz, and Bela H. Buck. Growth performance of noble crayfish Astacus astacus in recirculating aquaculture systems. Aquaculture International, 23(4): 997–1012, August 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9859-2.

Iliyasu:2015:PGT

[1427] Abdullahi Iliyasu, Zainal Abidin Mohamed, and Mazuki Hashim. Productivity growth, technical change and efficiency change of the Malaysian cage fish farming: an application of Malmquist Productivity Index approach. Aquaculture International, 23(4):1013–1024, August 2015. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9860-9.

Piccolo:2015:ELP

[1428] Giovanni Piccolo, Fulvia Bovera, Pietro Lombardi, Vincenzo Mastellone, Sandra Nizza, Carmelo Di Meo, Stefania Marono, and Antonino Nizza. Effect of Lactobacillus plantarum on growth performance and hematological traits of European sea bass (Dicentrarchus labrax). Aquaculture International, 23(4):1025–1032, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9861-8.

Mechri:2015:IGP

[1429] Badreddine Mechri, Imen Ben Salem, Amel Medhioub, Mohamed Nejib Medhioub, and Mahjoub Aouni. Isolation and genotyping of potentially pathogenic Vibrio alginolyticus associated with Ruditapes decussatus larva and juvenile mass mortalities. Aquaculture International, 23 (4):1033–1047, August 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9862-7.

Yang:2015:MIR

[1430] Lishi Yang, Qibin Yang, Shigui Jiang, Yong Li, Falin Zhou, Tao Li, and Jianhua Huang. Metabolic, immune responses in prawn (Penaeus monodon) exposed to ambient ammonia. Aquaculture International, 23 (4):1049–1062, August 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9863-6.

Perez-Casanova:2015:EBD

[1431] Juan Carlos Pérez-Casanova, Matthew Webb, Lynn Lush, Christopher Parrish, Isabel A. S. F. Costa, and Dounia Hamoutene. Effects of broodstock diets on growth of larval Atlantic cod (Gadus morhua L.). Aquaculture International, 23(4):1063–1070, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9864-5.

Rolland:2015:EPP

[1432] Marine Rolland, Bodil K. Larsen, Jørgen Holm, Johanne Dalsgaard, and Peter V. Skov. Effect of plant proteins and crystalline amino acid supplementation on postprandial plasma amino acid profiles and metabolic response in rainbow trout (Oncorhynchus mykiss). Aquaculture International, 23(4):1071–1087, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9865-4.

Methe:2015:SGP

[1433] D. Méthé, L. A. Comeau, H. Stryhn, T. Guyondet, J. F. Burka, T. Landry, and J. Davidson. Survival and growth performance of Crassostrea virginica along an estuarine gradient. Aquaculture International, 23(4):1089–1103, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-014-9866-3.

Peruzzi:2015:GMD

[1434] Stefano Peruzzi, Ørjan Hagen, and Malcolm Jobling. Gut morphology of diploid and triploid Atlantic salmon (Salmo salar L.). Aquaculture International, 23(4):1105–1108, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9867-2.

Mazumder:2015:LOW

[1435] Sabuj Kanti Mazumder, Mazlan Abd. Ghaffar, Simon Kumar Das, and Md. Samsul Alam. Low occurrence of WSSV in Penaeus monodon nauplii and post-larvae produced from PCR-negative broodstocks. Aquaculture International, 23(4):1109–1123, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9868-1.

Jobling:2015:BRD

[1436] Malcolm Jobling. Book review: D. Merrifield and E. Ringø (eds): Aquaculture nutrition: gut health, probiotics and prebiotics. Aquaculture International, 23(4):1125–1126, August 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-014-9869-0.

Verma:2015:EDR

[1437] Vipin Kumar Verma, Kumari Vandana Rani, Neeta Sehgal, and Om Prakash. Enhanced disease resistance in the Indian snakehead, Channa punctata against Aeromonas hydrophila, through 5 feed supplementation with F. benghalensis (aerial root) and L. leucocephala (pod seed). Aquaculture International, 23(5):1127–1140, October 2015. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-014-9870-7.

Wang:2015:CMS

[1438] Xingqiang Wang, Lingling Wang, Huan Zhang, Rui Liu, and Linsheng Song. The carbohydrate metabolism of scallop Chlamys farreri in the immune response against acute challenge of Vibrio anguillarum. Aquaculture International, 23(5):1141–1155, October 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9871-6.

Hansen:2015:DMA

[1439] Øyvind J. Hansen, Velmurugu Puvanendran, and Rama Bangera. Do maternal age and experience contribute to better growth, survival and disease resistance of offspring in Atlantic cod (Gadus morhua)? Aquaculture International, 23(5):1157–1164, October 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9872-5.

Goosen:2015:CHP

[1440] N. J. Goosen, L. F. De Wet, and J. F. Görgens. Comparison of hydrolysed proteins from different raw materials in diets for Mozambique tilapia Oreochromis mossambicus. Aquaculture International, 23(5): 1165–1178, October 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9873-4.

Kongrueng:2015:LAD

[1441] Jetnaphang Kongrueng, Natta Tansila, Pimonsri Mitraparp-arthorn, Mitsuaki Nishibuchi, Gary J. Vora, and Varaporn Vuddhakul. LAMP assay to detect Vibrio parahaemolyticus causing acute hepatopancreatic necrosis disease in shrimp. Aquaculture International, 23(5):1179– 1188, October 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9874-3.

Kpundeh:2015:EDP

[1442] Mathew Didlyn Kpundeh, Jun Qiang, Jie He, Hong Yang, and Pao Xu. Effects of dietary protein levels on growth performance and haematoimmunological parameters of juvenile genetically improved farmed tilapia (GIFT), Oreochromis niloticus. Aquaculture International, 23(5):1189– 1201, October 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-014-9876-1.

Wu:2015:VCR

[1443] Fan Wu, Feng Huang, Hua Wen, Ming Jiang, Wei Liu, Juan Tian, and Chang Geng Yang. Vitamin C requirement of adult genetically improved farmed tilapia, Oreochromis niloticus. Aquaculture International, 23(5): 1203–1215, October 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-014-9877-0.

Zou:2015:APA

[1444] Guiwei Zou, Yuanyuan Zhu, Hongwei Liang, and Zhong Li. Association of pituitary adenylate cyclase-activating polypeptide and myogenic factor 6 genes with growth traits in Nile tilapia (Oreochromis niloticus). Aquaculture International, 23(5):1217–1225, October 2015. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9878-7.

Barnes:2015:CPT

[1445] Michael E. Barnes, Michael L. Brown, and Regg Neiger. Comparative performance of two rainbow trout strains fed fermented soybean meal. Aquaculture International, 23(5):1227–1238, October 2015. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9879-6.

Alishahi:2015:EAD

[1446] M. Alishahi, M. Karamifar, and M. Mesbah. Effects of astaxanthin and Dunaliella salina on skin carotenoids, growth performance and immune response of Astronotus ocellatus. Aquaculture International, 23(5): 1239–1248, October 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9880-0.

Obirikorang:2015:DPA

[1447] Kwasi Adu Obirikorang, Stephen Amisah, Simon Cudjoe Fialor, and Peter Vilhelm Skov. Digestibility and postprandial ammonia excretion in Nile tilapia (Oreochromis niloticus) fed diets containing different oilseed by-products. Aquaculture International, 23(5):1249–1260, October 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9881-z.

Abdel-Tawwab:2015:EDO

[1448] Mohsen Abdel-Tawwab, Ahmed E. Hagras, Heba Allah M. Elbaghdady, and Mohamed N. Monier. Effects of dissolved oxygen and fish size on Nile tilapia, Oreochromis niloticus (L.): growth performance, whole-body composition, and innate immunity. Aquaculture International, 23(5): 1261–1274, October 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9882-y.

Kumar:2015:EDS

[1449] Rajesh Kumar, Subhas C. Mukherjee, Ritesh Ranjan, T. Vani, Rajeev K. Brahmachari, and Sukanta K. Nayak. Effect of dietary supplementation of Bacillus subtilis on haematological and immunological parameters of Catla catla (Hamilton). Aquaculture International, 23(5): 1275–1292, October 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9883-x.

Nalder:2015:CLC

[1450] Tim D. Nalder, Matthew R. Miller, and Michael A. Packer. Changes in lipid class content and composition of Isochrysis sp. (T-Iso) grown in batch culture. Aquaculture International, 23(5):1293–1312, October 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-015-9884-9.

Zhou:2015:TSE

[1451] Bo Zhou, Shuanglin Dong, and Fang Wang. Trophic structure and energy fluxes in a grass carp (Ctenopharyngodon idellus) cultured pond ecosystem. Aquaculture International, 23(5):1313–1324, October 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9886-7.

Serra:2015:UDC

[1452] Fabiane P. Serra, Carlos A. P. Gaona, Plínio S. Furtado, Luis H. Poersch, and Wilson Wasielesky. Use of different carbon sources for the biofloc system adopted during the nursery and grow-out culture of Litopenaeus vannamei. Aquaculture International, 23(6):1325–1339, December 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9887-6.

Ma:2015:EEP

[1453] Fei Ma, Xiaoqin Li, Baian Li, and Xiangjun Leng. Effects of extruded and pelleted diets with differing protein levels on growth and nutrient retention of tilapia, Oreochromis niloticus × O. aureus. Aquaculture International, 23(6):1341–1356, December 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9888-5.

Portilla:2015:PIE

[1454] Sixto E. Portilla, Brett F. Branco, and John T. Tanacredi. Preliminary investigation into the effects of two dietary fatty acids, 20:5n − 3 and 22:6n − 3, on mortality of juvenile Mercenaria mercenaria during the approach to winter. Aquaculture International, 23(6):1357– 1376, December 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9889-4.

Zhang:2015:EPC

[1455] Jiasong Zhang, Huayang Guo, Zhenhua Ma, Shigui Jiang, Kaichang Wu, Youning Li, and Jian G. Qin. Effects of prey color, wall color and water color on food ingestion of larval orange-spotted grouper Epinephelus coioides (Hamilton, 1822). Aquaculture International, 23(6): 1377–1386, December 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9890-y.

Guillen:2015:EEP

[1456] Jordi Guillen, Fabrizio Natale, and José M. Fernández Polanco. Estimating the economic performance of the EU aquaculture sector. Aquaculture International, 23(6):1387–1400, December 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9891-x.

Gao:2015:EIG

[1457] Baoquan Gao, Ping Liu, Jian Li, Qingyin Wang, and Zhike Han. Effect of inbreeding on growth and genetic diversity of Portunus trituberculatus based on the full-sibling inbreeding families. Aquaculture International, 23(6):1401–1410, December 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-015-9892-9.

Lee:2015:CFD

[1458] Soxi Lee, Neil D. Hartstein, and Andrew Jeffs. Characteristics of faecal and dissolved nitrogen production from tropical spiny lobster, Panulirus ornatus. Aquaculture International, 23(6):1411–1425, December 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9893-8.

Fu:2015:GPE

[1459] Jianjun Fu, Yubang Shen, Xiaoyan Xu, Chengchu Liu, and Jiale Li. Genetic parameter estimates and genotype by environment interaction analyses for early growth traits in grass carp (Ctenopharyngodon idella). Aquaculture International, 23(6):1427–1441, December 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-015-9894-7.

Lukwambe:2015:ECM

[1460] Betina Lukwambe, Linlin Qiuqian, Jinfeng Wu, Demin Zhang, Kai Wang, and Zhongming Zheng. The effects of commercial microbial agents (probiotics) on phytoplankton community structure in intensive white shrimp (Litopenaeus vannamei) aquaculture ponds. Aquaculture International, 23(6):1443–1455, December 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-015-9895-6.

Jeong:2015:AET

[1461] Min Ho Jeong, Ji Hyun Yun, Kwangmo Yang, and Wol Soon Jo. Application of electrochemically treated seawater as basic culture media for mass cultivation of Tetraselmis suecica. Aquaculture International, 23(6): 1457–1471, December 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9897-4.

Waller:2015:IMT

[1462] Uwe Waller, Anne K. Buhmann, Anneliese Ernst, Verena Hanke, Andreas Kulakowski, Bert Wecker, Jaime Orellana, and Jutta Papenbrock. Integrated multi-trophic aquaculture in a zero-exchange recirculation aquaculture system for marine fish and hydroponic halophyte production. Aquaculture International, 23(6):1473–1489, December 2015. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9898-3.

Cardenas:2015:ADL

[1463] Jessie Vargas Cárdenas, Alfredo Olivera Gálvez, Luis Otavio Brito, Elsa Vega Galarza, David Cano Pitta, and Victor Vergara Rubin. Assessment of different levels of green and brown seaweed meal in experimental diets for whiteleg shrimp (Litopenaeus vannamei, Boone) in recirculating aquaculture system. Aquaculture International, 23(6): 1491–1504, December 2015. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9899-2.

Liu:2015:SCM

[1464] Yibing Liu, Xiaoxu Li, Nicholas Robinson, and Jianguang Qin. Sperm cryopreservation in marine mollusk: a review. Aquaculture International, 23(6):1505–1524, December 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-015-9900-0.

Ferrara:2015:HMM

[1465] Ermete Ferrara, Andrea Gustinelli, Maria Letizia Fioravanti, Brunella Restucci, Francesco Quaglio, Stefania Marono, and Giovanni Piccolo. Histological and micro-/macro-morphological evaluation of intestine in sharpsnout seabream (Diplodus puntazzo) fed soybean meal-based diets added with MOS and inulin as prebiotics. Aquaculture International, 23(6):1525–1537, December 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-015-9902-y.

VanHoestenberghe:2015:EWA

[1466] Stijn Van Hoestenberghe, Mathieu Wille, Evelien De Swaef, Bruno M. Goddeeris, and Nancy Nevejan. Effect of weaning age and the use of different sized Artemia nauplii as first feed for jade perch Scortum barcoo. Aquaculture International, 23(6):1539–1552, December 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-015-9903-x.

Harikrishnan:2015:RAE

[1467] Ramasamy Harikrishnan, Ju-Sang Kim, Chellam Balasundaram, Sundaram Jawahar, and Moon-Soo Heo. Retracted article: Effect of Magnifera indica kernel–enriched feed on immune response of Penaeus indicus against white spot syndrome virus (WSSV). Aquaculture International, 23(6):1553, December 2015. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-012-9535-3.

Li:2016:EGP

[1468] Fei Li, Zhi min Gu, Xi lian Li, Jian lin Guo, Yong yi Jia, Yu fei Zhang, Xian ming Huang, and Wen min Zhang. Estimates of genetic parameters for growth-related traits of the red swamp crayfish, Procambarus clarkii. Aquaculture International, 24(1):1–10, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9904-9.

Lakeh:2016:GER

[1469] Amir Abbas Bazyar Lakeh, Hamid Farahmand, Werner Kloas, Alireza Mirvaghefi, Achim Trubiroha, Brian C. Peterson, and Sven Wuertz. Growth enhancement of rainbow trout (Oncorhynchus mykiss) by passive immunization against somatostatin-14. Aquaculture International, 24(1):11–21, February 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9905-8.

Johari:2016:ANC

[1470] Seyed Ali Johari, Mohammad Reza Kalbassi, Mehdi Soltani, and Il Je Yu. Application of nanosilver-coated zeolite as water filter media for fungal disinfection of rainbow trout (Oncorhynchus mykiss) eggs. Aquaculture International, 24(1):23–38, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9906-7.

Manriquez-Hernandez:2016:EIB

[1471] Juan A. Manríquez-Hernández, Jim Duston, and David J. Garbary. Effect of irradiance on bioremediation capacity of the red algae Chondrus crispus and Palmaria palmata. Aquaculture International, 24(1): 39–55, February 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9907-6.

Hagemann:2016:ELS

[1472] Andreas Hagemann, Gunvor Øie, Jan Ove Evjemo, and Yngvar Olsen. Effects of light and short-term temperature elevation on the 48-h hatching success of cold-stored Acartia tonsa Dana eggs. Aquaculture International, 24(1):57–68, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-015-9908-5.

Khanzadeh:2016:EPR

[1473] Mostafa Khanzadeh, Abolghasem Esmaeili Fereidouni, and Saleh Seifi Berenjestanaki. Effects of partial replacement of fish meal with Spirulina platensis meal in practical diets on growth, survival, body composition, and reproductive performance of three-spot gourami (Trichopodus trichopterus) (Pallas, 1770). Aquaculture International, 24(1):69– 84, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-015-9909-4.

Giri:2016:DEA

[1474] Sib Sankar Giri, S. Jai Suda, V. Sukumaran, and Se Chang Park. Dietary emodin affects the growth performance, immune responses, and disease resistance of Labeo rohita against Aeromonas hydrophila. Aquaculture International, 24(1):85–99, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9910-y.

Niu:2016:EDF

[1475] Jin Niu, Xu Chen, Yun-Qiang Zhang, Li-Xia Tian, Hei-Zhao Lin, Jun Wang, Yun Wang, and Yong-Jian Liu. The effect of different feeding rates on growth, feed efficiency and immunity of juvenile Penaeus monodon. Aquaculture International, 24(1):101–114, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-015-9911-x.

Pinheiro:2016:DMR

[1476] Ana C. A. S. Pinheiro, Enrico Volpe, Donatella Principi, Santino Prosperi, and Sara Ciulli. Development of a multiplex RT–PCR assay for simultaneous detection of the major viruses that affect rainbow trout (Oncorhynchus mykiss). Aquaculture International, 24(1): 115–125, February 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9912-9.

Akter:2016:DSM

[1477] Mst. Nahid Akter, Amalia Sutriana, Allah Dad Talpur, and Roshada Hashim. Dietary supplementation with mannan oligosaccharide influences growth, digestive enzymes, gut morphology, and microbiota in juvenile striped catfish, Pangasianodon hypophthalmus. Aquaculture International, 24(1):127–144, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-015-9913-8.

Peng:2016:ETR

[1478] Xianghe Peng, Fajian Li, Shimei Lin, and Yongjun Chen. Effects of total replacement of fish oil on growth performance, lipid metabolism and antioxidant capacity in tilapia (Oreochromis niloticus). Aquaculture International, 24(1):145–156, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9914-7.

Ahmed:2016:SSD

[1479] Raju Ahmed, S. M. Rafiquzaman, Muhammad Tofazzal Hossain, Jong-Min Lee, and In-Soo Kong. Species-specific detection of Vibrio alginolyticus in shellfish and shrimp by real-time PCR using the groEL gene. Aquaculture International, 24(1):157–170, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9916-5.

Ky:2016:HSF

[1480] Chin-Long Ky, Carole Blay, and Cédrik Lo. Half-sib family effects on cultured pearl quality traits in the black-lipped pearl oyster Pinctada margaritifera: testing for indirect benefits of polyandry and polygyny. Aquaculture International, 24(1):171–182, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9917-4.

Kotani:2016:EFF

[1481] Tomonari Kotani, Hiroyuki Imari, Aki Miyashima, and Hiroshi Fushimi. Effects of feeding with frozen freshwater cladoceran Moina macrocopa on the performance of red sea bream Pagrus major larviculture. Aquaculture International, 24(1):183–197, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9918-3.

Serracca:2016:MOH

[1482] Laura Serracca, Irene Rossini, Roberta Battistini, and Carlo Ercolini. Mortality and ostreid herpesvirus 1 infection in the Pacific oyster Crassostrea gigas in the Gulf of La Spezia, Italy. Aquaculture International, 24 (1):199–209, February 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9919-2.

Xie:2016:ASM

[1483] Xi Xie, Libin Zhang, Shilin Liu, Tao Zhang, and Hongsheng Yang. Aerated sea mud is beneficial for post-nursery culture of early juvenile sea cucumber Apostichopus japonicus (Selenka). Aquaculture International, 24 (1):211–224, February 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9920-9.

Mohammadian:2016:PEL

[1484] Takavar Mohammadian, Mojtaba Alishahi, Mohammad Reza Tabandeh, Masoud Ghorbanpoor, Dariush Gharibi, Morteza Tollabi, and Siamak Rohanizade. Probiotic effects of Lactobacillus plantarum and L. delbrueckii ssp. bulguricus on some immune-related parameters in Barbus grypus. Aquaculture International, 24(1):225–242, February 2016. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9921-8.

Martinez-Porchas:2016:PSC

[1485] Marcel Martínez-Porchas, Susana Maria Scheuren-Acevedo, Luis Rafael Martínez-Córdova, Teresa Gollas-Galvan, Ramon H. Barraza-Guardado, Fernando Enríquez-Ocaña, Edilmar Cortés-Jacinto, and Marco A. Porchas-Cornejo. Physiological and sanitary condition of the white clam Dosinia ponderosa collected from a coastal area impacted by shrimp farm effluent. Aquaculture International, 24(1):243–256, February 2016. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9922-7.

Kokou:2016:EDM

[1486] Fotini Kokou, George Rigos, Maroudio Kentouri, and Maria Alexis. Effects of DL-methionine-supplemented dietary soy protein concentrate on growth performance and intestinal enzyme activity of gilthead sea bream (Sparus aurata L.). Aquaculture International, 24(1):257–271, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-015-9923-6.

Harris:2016:DFB

[1487] L. J. Harris, M. J. Sellars, and F. Perez. Differential family-based survival of Pacific white shrimp during communal larval culture and implications for selective breeding programs. Aquaculture International, 24 (1):273–279, February 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9924-5.

Wang:2016:MPM

[1488] Chunfang Wang, Qin Tang, Wei Luo, and Yingming Zhao. Modeling phosphorus metabolism in fish species: an example for juvenile yellow catfish, Pelteobagrus fulvidraco. Aquaculture International, 24(1): 281–294, February 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9925-4.

Tallar:2016:AWQ

[1489] Robby Yussac Tallar and Jian-Ping Suen. Aquaculture water quality index: a low-cost index to accelerate aquaculture development in Indonesia. Aquaculture International, 24(1):295–312, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-015-9926-3.

Cheirsilp:2016:EOC

[1490] Benjamas Cheirsilp, Yohanis Irenius Mandik, and Poonsuk Prasertsan. Evaluation of optimal conditions for cultivation of marine Chlorella sp. as potential sources of lipids, exopolymeric substances and pigments. Aquaculture International, 24(1):313–326, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9927-2.

Guo:2016:BAO

[1491] Lin Guo, Fei Xu, Zhigang Feng, and Guofan Zhang. A bibliometric analysis of oyster research from 1991 to 2014. Aquaculture International, 24 (1):327–344, February 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9928-1.

Ma:2016:RGM

[1492] Niannian Ma, Guozhi Luo, Hongxin Tan, Li Li, and Xiaoyong Wang. Removal of geosmin and 2-methylisoborneol by bioflocs produced with aquaculture waste. Aquaculture International, 24(1):345–356, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9929-0.

Shan:2016:ANN

[1493] H. W. Shan, W. Y. Bao, S. Ma, D. P. Wei, and L. Gao. Ammonia and nitrite nitrogen removal in shrimp culture by Vibrio alginolyticus VZ5 immobilized in SA beads. Aquaculture International, 24(1): 357–372, February 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9930-7.

Zhang:2016:SIO

[1494] Dongsheng Zhang, Hua Li, Yang Liu, Guo Qiao, Shuang Chi, and Jian Song. Screening and identification of organics-degrading bacteria from the sediment of sea cucumber Apostichopus japonicus ponds. Aquaculture International, 24(1):373–384, February 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9931-6.

Nuwansi:2016:EWF

[1495] K. K. T. Nuwansi, A. K. Verma, Chandra Prakash, V. K. Tiwari, M. H. Chandrakant, A. P. Shete, and G. P. W. A. Prabhath. Effect of water flow rate on polyculture of koi carp (Cyprinus carpio var. koi) and goldfish (Carassius auratus) with water spinach (Ipomoea aquatica) in recirculating aquaponic system. Aquaculture International, 24(1): 385–393, February 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9932-5.

Soltanzadeh:2016:GPB

[1496] Somayeh Soltanzadeh, Abolghasem Esmaeili Fereidouni, Hossein Ouraji, and Khosrow Jani Khalili. Growth performance, body composition, hematological, and serum biochemical responses of beluga (Huso huso) juveniles to different dietary inclusion levels of faba bean (Vicia faba) meal. Aquaculture International, 24(1):395–413, February 2016. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9933-4.

Medina:2016:APG

[1497] Miles Medina, Krishnaswamy Jayachandran, Mahadev G. Bhat, and Alok Deoraj. Assessing plant growth, water quality and economic effects from application of a plant-based aquafeed in a recirculating aquaponic system. Aquaculture International, 24(1):415–427, February 2016. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9934-3.

El-Magsodi:2016:HDC

[1498] Mohamed Omar El-Magsodi, Kartik Baruah, Parisa Norouzitallab, Peter Bossier, Patrick Sorgeloos, and Gilbert Van Stappen. Hydration/ dehydration cycles imposed on Artemia cysts influence the tolerance limit of nauplii against abiotic and biotic stressors. Aquaculture International, 24(2):429–439, April 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9935-2.

Krol:2016:EDT

[1499] Jarosław Król and Zdzisław Zakeś. Effect of dietary l-tryptophan on cannibalism, survival and growth in pikeperch Sander lucioperca (L.) post-larvae. Aquaculture International, 24(2):441–451, April 2016. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9936-1.

Zhou:2016:GPI

[1500] Qun lan Zhou, Jun Xie, Xian ping Ge, H. Michael Habte-Tsion, Bo Liu, and Mingchun Ren. Growth performance and immune responses of gibel carp, Carassius auratus gibelio, fed with graded level of rare earthchitosan chelate. Aquaculture International, 24(2):453–463, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9937-0.

Isnansetyo:2016:NSI

[1501] Alim Isnansetyo, Amiqatul Fikriyah, Noer Kasanah, and Murwantoko. Non-specific immune potentiating activity of fucoidan from a tropical brown algae (Phaeophyceae), Sargassum cristaefolium in tilapia (Oreochromis niloticus). Aquaculture International, 24(2):465– 477, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-015-9938-z.

Rezaee:2016:GDS

[1502] Saiwan Rezaee, Hamid Farahmand, and Mohammad Ali Nematollahi. Genetic diversity status of Pacific white shrimp (Litopenaeus vannamei) using SSR markers in Iran. Aquaculture International, 24(2):479–489, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-015-9939-y.

Zhu:2016:EEN

[1503] Dashi Zhu, Xiaobo Wen, Shengkang Li, Xiongzhi Xuan, and Yuanyou Li. Effects of exogenous non-starch polysaccharide-degrading enzymes in diets containing Gracilaria lemaneiformis on white-spotted snapper Lutjanus stellatus Akazaki. Aquaculture International, 24(2):491– 502, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-015-9940-5.

Sheedy:2016:HNM

[1504] John R. Sheedy, Sébastien Lachambre, David K. Gardner, and Robert W. Day. ^{1} H-NMR metabolite profiling of abalone digestive gland in response to short-term starvation. Aquaculture International, 24(2):503– 521, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-015-9941-4.

Shen:2016:BSL

[1505] Ya nan Shen, Dong ling Zhang, Bo Zhang, Song Jiang, Bo suo Liu, Gui ju Huang, and Da hui Yu. A baseline study on lipid and fatty acid composition in the pearl oyster, Pinctada fucata. Aquaculture International, 24(2):523–536, April 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9942-3.

Miyake:2016:EGC

[1506] T. Miyake, K. Isowa, T. Ishikawa, A. Komaru, and K. Kawamura. Evaluation of genetic characteristics of wild and cultured populations of the Japanese pearl oyster Pinctada fucata martensii by using AFLP markers. Aquaculture International, 24(2):537–548, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9943-2.

Okamura:2016:ESO

[1507] Akihiro Okamura, Yoshiaki Yamada, Naomi Mikawa, Noriyuki Horie, and Katsumi Tsukamoto. Effect of salinity on occurrence of notochord deformities in Japanese eel Anguilla japonica larvae. Aquaculture International, 24(2):549–555, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-015-9944-1.

Mbiru:2016:CPM

[1508] Moses Mbiru, Samwel Mchele Limbu, Sebastian W. Chenyambuga, Hieromin Amon Lamtane, Rashid Tamatamah, Nazael Amos Madalla, and Augustine Warioba Mwandya. Comparative performance of mixed-sex and hormonal-sex-reversed Nile tilapia Oreochromis niloticus and hybrids (Oreochromis niloticus × Oreochromis urolepis hornorum) cultured in concrete tanks. Aquaculture International, 24(2):557–566, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-015-9946-z.

Zhang:2016:PDY

[1509] Jin Zhang, Wenge Ma, Weimin Wang, Jian-Fang Gui, and Jie Mei. Parentage determination of yellow catfish (Pelteobagrus fulvidraco) based on microsatellite DNA markers. Aquaculture International, 24(2):567– 576, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-015-9947-y.

Hong:2016:PRA

[1510] Xupeng Hong, Liqun Lu, and Dan Xu. Progress in research on acute hepatopancreatic necrosis disease (AHPND). Aquaculture International, 24(2):577–593, April 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9948-x.

Byadgi:2016:IEN

[1511] Omkar Byadgi, Chin-I Chang, Jai-Wei Lee, Chih-Chun Jean Huang, and Ta-Chih Cheng. Immunostimulatory effects of non-CpG oligodeoxynucleotides on cobia (Rachycentron canadum). Aquaculture International, 24(2):595–608, April 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9950-3.

deBezerra:2016:EAC

[1512] Thales Ramon Q. de Bezerra, Ernesto C. Domingues, Luiz Flávio A. Maia Filho, Artur N. Rombenso, Santiago Hamilton, and Ronaldo O. Cavalli. Economic analysis of cobia (Rachycentron canadum) cage culture in large- and small-scale production systems in Brazil. Aquaculture International, 24(2):609–622, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-015-9951-2.

Ho:2016:EPC

[1513] Li-Ming Ho, Jung-Fu Huang, and Jie-Min Lee. The effects of price changes on oyster farmers in Changhua County, Taiwan, who cultivate oysters at low and high densities. Aquaculture International, 24 (2):623–635, April 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9952-1.

Castillo-Castellanos:2016:IEN

[1514] D. Castillo-Castellanos, I. Zavala-Leal, J. M. J. Ruiz-Velazco, A. Radilla-García, J. T. Nieto-Navarro, C. A. Romero-Bañuelos, and J. González-Hernández. Implementation of an experimental nutrient film techniquetype aquaponic system. Aquaculture International, 24(2):637–646, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-015-9954-z.

Grassi:2016:BPG

[1515] Thiago Luís Magnani Grassi, Edson Francisco do Espírito Santo, Marcelo Tacconi de Siqueira Marcos, Jefferson Felipe Cavazzana, Dayse Lícia Oliveira, Iderlipes Luiz Carvalho Bossolani, and Elisa Helena Giglio Ponsano. Bacterial pigment for Nile tilapia feeding. Aquaculture International, 24(2):647–660, April 2016. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9955-y.

Shilta:2016:EBW

[1516] M. T. Shilta, Narinder Kumar Chadha, P. K. Pandey, and Paramita Banerjee Sawant. Effect of biofilm on water quality and growth of Etroplus suratensis (Bloch, 1790). Aquaculture International, 24(2):661– 674, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-015-9956-x.

Yang:2016:EDR

[1517] Gang Yang, Xiangli Tian, Shuanglin Dong, Mo Peng, Dongdong Wang, and Kai Zhang. Effects of dietary rhubarb, Bacillus cereus, yeast polysaccharide, and florfenicol supplementation on growth, intestinal morphology, and immune responses of sea cucumber (Apostichopus japonicus). Aquaculture International, 24(2):675–690, April 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9957-9.

Burnell:2016:NEA

[1518] Gavin Burnell. Note of the Editor on the 20th anniversary of aquaculture international. Aquaculture International, 24(3):691–692, June 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0002-4.

Mente:2016:ISI

[1519] E. Mente and A. C. Smaal. Introduction to the special issue on “European aquaculture development since 1993: the benefits of aquaculture to Europe and the perspectives of European aquaculture production”. Aquaculture International, 24(3):693–698, June 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0003-3.

Bostock:2016:AEC

[1520] John Bostock, Alistair Lane, Courtney Hough, and Koji Yamamoto. An assessment of the economic contribution of EU aquaculture production and the influence of policies for its sustainable development. Aquaculture International, 24(3):699–733, June 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-9992-1.

Jansen:2016:FOA

[1521] Henrice M. Jansen, Sander Van Den Burg, Bas Bolman, Robbert G. Jak, Pauline Kamermans, Marnix Poelman, and Marian Stuiver. The feasibility of offshore aquaculture and its potential for multi-use in the North Sea. Aquaculture International, 24(3):735–756, June 2016. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-9987-y.

Lekang:2016:CET

[1522] O. I. Lekang, C. Salas-Bringas, and J. C. Bostock. Challenges and emerging technical solutions in on-growing salmon farming. Aquaculture International, 24(3):757–766, June 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-9994-z.

Jobling:2016:FNR

[1523] Malcolm Jobling. Fish nutrition research: past, present and future. Aquaculture International, 24(3):767–786, June 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-014-9875-2.

Steffens:2016:APW

[1524] Werner Steffens. Aquaculture produces wholesome food: cultured fish as a valuable source of n − 3 fatty acids. Aquaculture International, 24(3):787–802, June 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9885-8.

Ferreira:2016:GSE

[1525] J. G. Ferreira and S. B. Bricker. Goods and services of extensive aquaculture: shellfish culture and nutrient trading. Aquaculture International, 24(3):803–825, June 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9949-9.

Kamermans:2016:RNS

[1526] Pauline Kamermans, Ainhoa Blanco, Sandra Joaquim, Domitília Matias, Thorolf Magnesen, Jean Louis Nicolas, Bruno Petton, and Rene Robert. Recirculation nursery systems for bivalves. Aquaculture International, 24(3):827–842, June 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-9990-3.

Young:2016:MPM

[1527] Tim Young, Andrea C. Alfaro, and Silas G. Villas-Bôas. Metabolic profiling of mussel larvae: effect of handling and culture conditions. Aquaculture International, 24(3):843–856, June 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-015-9945-0.

Petersen:2016:USE

[1528] Jens Kjerulf Petersen, Camille Saurel, Pernille Nielsen, and Karen Timmermann. The use of shellfish for eutrophication control. Aquaculture International, 24(3):857–878, June 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-015-9953-0.

Chen:2016:PSP

[1529] Li hong Chen, Rong lian Xing, Ai li Jiang, Li Teng, and Chang hai Wang. A preliminary study on the potential value of a novel integrated aquaculture system on water purification. Aquaculture International, 24 (4):879–892, August 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9958-8.

Zhu:2016:GGE

[1530] Lin Zhu, Xuan Che, Huang Liu, Xingguo Liu, Chong Liu, Xiaolong Chen, and Xu Shi. Greenhouse gas emissions and comprehensive greenhouse effect potential of Megalobrama amblycephala culture pond ecosystems in a 3-month growing season. Aquaculture International, 24(4): 893–902, August 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9959-7.

Anane-Taabeah:2016:AFT

[1531] Gifty Anane-Taabeah, Kwamena Quagrainie, and Steve Amisah. Assessment of farmed tilapia value chain in Ghana. Aquaculture International, 24(4):903–919, August 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9960-1.

Ferreira:2016:BVB

[1532] Lise M. H. Ferreira, Gabriele Lara, Wilson Wasielesky, and Paulo Cesar Abreu. Biofilm versus biofloc: Are artificial substrates for biofilm production necessary in the BFT system? Aquaculture International, 24 (4):921–930, August 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9961-0.

Llorente:2016:BMA

[1533] Ignacio Llorente and Ladislao Luna. Bioeconomic modelling in aquaculture: an overview of the literature. Aquaculture International, 24 (4):931–948, August 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9962-z.

Hansen:2016:ORF

[1534] Benni Winding Hansen, Elisa Blanda, Guillaume Drillet, Jacob Kring Højgaard, Mohamed-Sofiane Mahjoub, and Thomas Allan Rayner. Outdoor rearing facilities of free spawning calanoid copepods for turbot larva can host a bank of resting eggs in the sediment. Aquaculture International, 24(4):949–964, August 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-015-9963-y.

Rico:2016:IDI

[1535] R. M. Rico, M. T. Tejedor-Junco, S. T. Tapia-Paniagua, F. J. Alarcón, J. M. Mancera, F. López-Figueroa, M. C. Balebona, R. T. Abdala-Díaz, and M. A. Moriñigo. Influence of the dietary inclusion of Gracilaria cornea and Ulva rigida on the biodiversity of the intestinal microbiota of Sparus aurata juveniles. Aquaculture International, 24(4): 965–984, August 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9964-x.

Medeiros:2016:ANT

[1536] Rafael Soriani Medeiros, Bianca Azevedo Lopez, Luís André Sampaio, Luis Alberto Romano, and Ricardo Vieira Rodrigues. Ammonia and nitrite toxicity to false clownfish Amphiprion ocellaris. Aquaculture International, 24(4):985–993, August 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-015-9965-9.

Wang:2016:GPB

[1537] Qingheng Wang, Chuangye Yang, Xiaodong Du, Xingwang Liu, Ruijiao Sun, and Yuewen Deng. Growth performance and biochemical composition of juvenile pearl oyster Pinctada martensii fed on artificial diets. Aquaculture International, 24(4):995–1005, August 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-015-9966-8.

Guo:2016:ENG

[1538] Jingping Guo, Benyue Guo, Hanle Zhang, Wei Xu, Wenbing Zhang, and Kangsen Mai. Effects of nucleotides on growth performance, immune response, disease resistance and intestinal morphology in shrimp Litopenaeus vannamei fed with a low fish meal diet. Aquaculture International, 24(4):1007–1023, August 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-015-9967-7.

Sotelano:2016:CAL

[1539] M. Paula Sotelano, Gustavo A. Lovrich, and Federico Tapella. Cannibalism among Lithodes santolla (Molina 1782) juveniles: effect of stocking density, stage and molt condition. Aquaculture International, 24(4): 1025–1037, August 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-015-9968-6.

Zhang:2016:EHP

[1540] Bin Zhang, Zhou rong Shi, Xiao ling Wang, and Shang gui Deng. The effects of hairtail protein hydrolysate–fe^{2+} complexes on growth and non-specific immune response of red swamp crayfish (Procambarus clarkii). Aquaculture International, 24(4):1039–1048, August 2016. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-9969-0.

Hisano:2016:CGM

[1541] Hamilton Hisano, José Luiz Pilecco, and Jorge Antônio Ferreira de Lara. Corn gluten meal in pacu Piaractus mesopotamicus diets: effects on growth, haematology, and meat quality. Aquaculture International, 24 (4):1049–1060, August 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-9970-7.

Lim:2016:EEH

[1542] Han Kyu Lim. Effect of exogenous hormones on ovulation and gonadal steroid plasma levels in starry flounder, Platichthys stellatus. Aquaculture International, 24(4):1061–1071, August 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-9971-6.

Wassef:2016:RGE

[1543] Elham A. Wassef, Olfat M. Wahbi, Eman M. Saqr, and Norhan E. Saleh. Response of European seabass (Dicentrarchus labrax) to canola oil diets: effect on growth performance, fish health and liver and intestine histomorphology. Aquaculture International, 24(4):1073–1088, August 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-9972-5.

Sundaray:2016:SSR

[1544] Jitendra Kumar Sundaray, Kiran Dashrath Rasal, Vemulawada Chakrapani, Pranati Swain, Dinesh Kumar, Arun S. Ninawe, Samiran Nandi, and Pallipuram Jayasankar. Simple sequence repeats (SSRs) markers in fish genomic research and their acceleration via next-generation sequencing and computational approaches. Aquaculture International, 24 (4):1089–1102, August 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-9973-4.

delaPena:2016:UTS

[1545] Milagros R. de la Peña, Myrna B. Teruel, Jose M. Oclarit, Mary Jane A. Amar, and Ellen Grace T. Ledesma. Use of thraustochytrid Schizochytrium sp. as source of lipid and fatty acid in a formulated diet for abalone Haliotis asinina (Linnaeus) juveniles. Aquaculture International, 24(4):1103–1118, August 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-9974-3.

Lin:2016:MCM

[1546] Jian Lin, Bao Ju, Yanan Yao, Xiaoming Lin, Ronglian Xing, Li Teng, and Aili Jiang. Microbial community in a multi-trophic aquaculture system of Apostichopus japonicus, Styela clava and microalgae. Aquaculture International, 24(4):1119–1140, August 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-9975-2.

Qinzeng:2016:RSS

[1547] Xu Qinzeng, Zhang Libin, Zhang Xuelei, Zhou Yi, and Yang Hongsheng. Release size and stocking density for grow-out of Apostichopus japonicus in the sea with raft-cultured macroalgae. Aquaculture International, 24 (4):1141–1152, August 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-9976-1.

Lim:2016:ECV

[1548] Yun-Jin Lim, Do-Hyung Kim, Heyong Jin Roh, Myoung-Ae Park, Chan-Il Park, and Peter Smith. Epidemiological cut-off values for disc diffusion data generated by standard test protocols from Edwardsiella tarda and Vibrio harveyi. Aquaculture International, 24(4):1153–1161, August 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-016-9977-0.

Islam:2016:ISS

[1549] S. M. Didar-Ul Islam and Mohammad Amir Hossain Bhuiyan. Impact scenarios of shrimp farming in coastal region of Bangladesh: an approach of an ecological model for sustainable management. Aquaculture International, 24(4):1163–1190, August 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-9978-z.

Liu:2016:ORM

[1550] Jinhui Liu, Yongqin Zhang, Saiyu Gui, Wenbin Liu, Jun Xiao, and Yamei Xiao. Observation and regression models on body colour inheritance and development in crucian carp and carp. Aquaculture International, 24(4): 1191–1199, August 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-9979-y.

Soliman:2016:AGE

[1551] Naglaa F. Soliman and Dalia M. M. Yacout. Aquaculture in Egypt: status, constraints and potentials. Aquaculture International, 24(5): 1201–1227, October 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-9989-9.

Huang:2016:EDL

[1552] Yisheng Huang, Xiaobo Wen, Shengkang Li, Wenjia Li, and Dashi Zhu. Effects of dietary lipid levels on growth, feed utilization, body composition, fatty acid profiles and antioxidant parameters of juvenile chu’s croaker Nibea coibor. Aquaculture International, 24(5):1229– 1245, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-016-9980-5.

Martinez-Pita:2016:NLM

[1553] Inés Martínez-Pita, Clara Sánchez-Lazo, and Marcelino Herrera. A nonlethal method for establishing sexual maturation in mussels (Mytilus galloprovincialis (Lamarck, 1819)) during broodstock conditioning in hatcheries. Aquaculture International, 24(5):1247–1254, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-9981-4.

Nass:2016:ELF

[1554] Danilo Henrique Nass, Eduardo Luiz Tavares Gonçalves, and Mônica Yumi Tsuzuki. Effect of live food transition time on survival, growth and metamorphosis of yellowtail clownfish, Amphiprion clarkii, larvae. Aquaculture International, 24(5):1255–1261, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-9982-3.

Gaona:2016:BMD

[1555] Carlos Augusto Prata Gaona, Fabiane da Paz Serra, Plínio Schmidt Furtado, Luis Henrique Poersch, and Wilson Wasielesky. Biofloc management with different flow rates for solids removal in the Litopenaeus vannamei BFT culture system. Aquaculture International, 24(5):1263– 1275, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-016-9983-2.

Midhun:2016:MDE

[1556] Sebastian J. Midhun, Damodaran Arun, Lincy Edatt, M. V. Sruthi, V. V. Thushara, Oommen V. Oommen, V. B. Sameer Kumar, and Lekha Divya. Modulation of digestive enzymes, GH, IGF-1 and IGF-2 genes in the teleost, tilapia (Oreochromis mossambicus) by dietary curcumin. Aquaculture International, 24(5):1277–1286, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-9984-1.

Chavanne:2016:CSS

[1557] Hervé Chavanne, Kasper Janssen, Johann Hofherr, Franca Contini, Pierrick Haffray, Hans Komen, Einar Eg Nielsen, and Luca Bargelloni. A comprehensive survey on selective breeding programs and seed market in the European aquaculture fish industry. Aquaculture International, 24 (5):1287–1307, October 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-9985-0.

Day:2016:GCA

[1558] Scott B. Day, Khalid Salie, and Henk B. Stander. A growth comparison among three commercial tilapia species in a biofloc system. Aquaculture International, 24(5):1309–1322, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-9986-z.

Qi:2016:LTE

[1559] Shoubing Qi, Weijie Zhang, Chenchen Jing, Haifeng Wang, Shuai Zhao, Mi Zhou, and Yaqing Chang. Long-term effects of stocking density on survival, growth performance and marketable production of the sea urchin Strongylocentrotus intermedius. Aquaculture International, 24(5): 1323–1339, October 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-9988-x.

Loh:2016:EPP

[1560] Jiun Yan Loh and Adeline Su Yien Ting. Effects of potential probiotic Lactococcus lactis subsp. lactis on digestive enzymatic activities of live feed Artemia franciscana. Aquaculture International, 24(5): 1341–1351, October 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-9991-2.

Zhao:2016:DLC

[1561] Peng-Fei Zhao, Fa-Jian Li, Xiao-Ru Chen, Yong-Jun Chen, Shi-Mei Lin, Lu Zhang, and Yun Li. Dietary lipid concentrations influence growth, liver oxidative stress, and serum metabolites of juvenile hybrid snakehead (Channa argus × Channa maculata). Aquaculture International, 24(5): 1353–1364, October 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-9993-0.

Qiang:2016:CCE

[1562] J. Qiang, J. He, H. Yang, P. Xu, H. Michael Habte-Tsion, X. Y. Ma, and Z. X. Zhu. The changes in cortisol and expression of immune genes of GIFT tilapia Oreochromis niloticus (L.) at different rearing densities under Streptococcus iniae infection. Aquaculture International, 24(5): 1365–1378, October 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-9995-y.

Gupta:2016:GPF

[1563] Akhil Gupta, Geetika Verma, and Paromita Gupta. Growth performance, feed utilization, digestive enzyme activity, innate immunity and protection against Vibrio harveyi of freshwater prawn, Macrobrachium rosenbergii fed diets supplemented with Bacillus coagulans. Aquaculture International, 24(5):1379–1392, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-9996-x.

Huang:2016:MAS

[1564] Zhitao Huang, Rong Wan, Xiefa Song, Ying Liu, Eric Hallerman, Dengpan Dong, Jieming Zhai, Hesen Zhang, and Liyuan Sun. Metagenomic analysis shows diverse, distinct bacterial communities in biofilters among different marine recirculating aquaculture systems. Aquaculture International, 24(5):1393–1408, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-9997-9.

Garcia-Celdran:2016:GAT

[1565] M. García-Celdrán, G. Ramis, E. María-Dolores, J. Peñalver, Y. J. Borrell, M. Manchado, A. Estévez, J. M. Afonso, and E. Armero. Genetic assessment of three gilthead sea bream (Sparus aurata L.) populations along the Spanish coast and of three broodstocks managements. Aquaculture International, 24(5):1409–1420, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-9998-8.

Doskeland:2016:ELT

[1566] Inge Døskeland, Albert Kjartan D. Imsland, Per Gunnar Fjelldal, Sigurd O. Stefansson, Bjorn Roth, Katrine Eikeland Eriksen, Bjørn Mikalsen, and Sigurd Handeland. The effect of low temperatures and photoperiods on growth and vertebra morphometry in Atlantic salmon. Aquaculture International, 24(5):1421–1434, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-016-9999-7.

Esparza-Leal:2016:PLV

[1567] Héctor M. Esparza-Leal, João A. Amaral Xavier, and Wilson Wasielesky, Jr. Performance of Litopenaeus vannamei postlarvae reared in indoor nursery tanks under biofloc conditions at different salinities and zero-water exchange. Aquaculture International, 24(5):1435–1447, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-016-0001-5.

Delfosse:2016:UPZ

[1568] Cyril Delfosse, Cécile Bienboire-Frosini, Camille Chabaud, Céline Lafont-Lecuelle, Alessandro Cozzi, and Patrick Pageat. Using physiological and zootechnical profiles to evaluate welfare in farmed rainbow trout Oncorhynchus mykiss (Walbaum) under stressful conditions. Aquaculture International, 24(5):1449–1457, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0004-2.

Capelle:2016:RSC

[1569] Jacob J. Capelle, Gerard Scheiberlich, Jeroen W. M. Wijsman, and Aad C. Smaal. The role of shore crabs and mussel density in mussel losses at a commercial intertidal mussel plot after seeding. Aquaculture International, 24(5):1459–1472, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0005-1.

Suita:2016:CBC

[1570] Sabrina M. Suita, André Braga, Eduardo Ballester, Alessandro P. Cardozo, Paulo Cesar Abreu, and Wilson Wasielesky. Contribution of bioflocs to the culture of Litopenaeus vannamei post-larvae determined using stable isotopes. Aquaculture International, 24(5):1473–1487, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-016-0006-0.

Mancuso:2016:ESC

[1571] T. Mancuso, L. Baldi, and L. Gasco. An empirical study on consumer acceptance of farmed fish fed on insect meals: the Italian case. Aquaculture International, 24(5):1489–1507, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0007-z.

Imsland:2016:IBI

[1572] Albert K. Imsland, Patrick Reynolds, Gerhard Eliassen, Thor A. Hangstad, Ólöf D. B. Jónsdóttir, Tor Anders Elvegård, Sebastiaan C. A. Lemmens, Randi Rydland, and Ane V. Nytrø. Investigation of behavioural interactions between lumpfish (Cyclopterus lumpus) and goldsinny wrasse (Ctenolabrus rupestris) under controlled conditions. Aquaculture International, 24(5):1509–1521, October 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-016-0008-y.

Linhartova:2016:P

[1573] Zuzana Linhartová and Jan Mráz. Preface. Aquaculture International, 24(6):1523–1525, December 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-016-0072-3.

Piria:2016:CCC

[1574] Marina Piria, Tea Tomljanović, Tomislav Treer, Roman Safner, Ivica Aničić, Daniel Matulić, and Lorenzo Vilizzi. The common carp Cyprinus carpio in Croatia (Danube and Adriatic basins): a historical review. Aquaculture International, 24(6):1527–1541, December 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-016-0029-6.

Gal:2016:SEI

[1575] Dénes Gál, Ferenc Pekár, and Éva Kerepeczki. A survey on the environmental impact of pond aquaculture in Hungary. Aquaculture International, 24(6):1543–1554, December 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0034-9.

Lunda:2016:EDT

[1576] Roman Lunda, Zuzana Linhartová, Jan Másílko, Petr Dvořák, Sonja Smole Možina, and Jan Mráz. Effect of different types of descaling methods on shelf life of air-/vacuum-packaged common carp (Cyprinus carpio L.) fillets under refrigerated storage conditions. Aquaculture International, 24(6):1555–1568, December 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0040-y.

Sternisa:2016:MAC

[1577] Meta Sterniša, Jan Mraz, and Sonja Smole Možina. Microbiological aspects of common carp (Cyprinus carpio) and its processing — relevance for final product quality: a review. Aquaculture International, 24(6): 1569–1590, December 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0051-8.

Adamek:2016:EQN

[1578] Z. Adámek, M. Mrkvová, J. Zukal, K. Roche, L. Mikl, L. Šlapanský, M. Janáč, and P. Jurajda. Environmental quality and natural food performance at feeding sites in a carp (Cyprinus carpio) pond. Aquaculture International, 24(6):1591–1606, December 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0048-3.

Policar:2016:CPE

[1579] Tomáš Policar, Miroslav Blecha, Jiří Křišťan, Jan Mráz, Josef Velíšek, Alžběta Stará, Vlastimil Stejskal, Oleksandr Malinovskyi, Petr Svačina, and Azin Mohagheghi Samarin. Comparison of production efficiency and quality of differently cultured pikeperch (Sander lucioperca L.) juveniles as a valuable product for ongrowing culture. Aquaculture International, 24(6):1607–1626, December 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-016-0050-9.

Jurajda:2016:CFA

[1580] P. Jurajda, Z. Adámek, K. Roche, M. Mrkvová, D. Štarhová, V. Prášek, and J. Zukal. Carp feeding activity and habitat utilisation in relation to supplementary feeding in a semi-intensive aquaculture pond. Aquaculture International, 24(6):1627–1640, December 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0061-6.

Zajic:2016:FFS

[1581] Tomas Zajic, Jan Mraz, Sabine Sampels, and Jana Pickova. Finishing feeding strategy as an instrument for modification of fatty acid composition of brook char (Salvelinus fontinalis). Aquaculture International, 24(6):1641–1656, December 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-016-0067-0.

Nebesky:2016:TIE

[1582] Václav Nebeský, Tomáš Policar, Miroslav Blecha, Jiří Křišťan, and Petr Svačina. Trends in import and export of fishery products in the Czech Republic during 2010–2015. Aquaculture International, 24(6): 1657–1668, December 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0066-1.

Masilko:2016:RMF

[1583] Jan Másílko, Tomáš Zajíc, David Hlaváč, Sabine Sampels, Jan Mráz, and Martin Oberle. Rapid measurements of fat content in live and slaughtered common carp (Cyprinus carpio L.). Aquaculture International, 24(6): 1669–1679, December 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0058-1.

Hlavac:2016:SFT

[1584] David Hlaváč, Maria Anton-Pardo, Jan Másílko, Pavel Hartman, Ján Regenda, Pavel Vejsada, Marek Baxa, Libor Pechar, Olga Valentová, Lucie Všetičková, Bořek Drozd, and Zdeněk Adámek. Supplementary feeding with thermally treated cereals in common carp (Cyprinus carpio L.) pond farming and its effects on water quality, nutrient budget and zooplankton and zoobenthos assemblages. Aquaculture International, 24(6): 1681–1697, December 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0059-0.

Markovic:2016:CAU

[1585] Zoran Marković, Marko Stanković, Božidar Rašković, Zorka Dulić, Ivana Živić, and Vesna Poleksić. Comparative analysis of using cereal grains and compound feed in semi-intensive common carp pond production. Aquaculture International, 24(6):1699–1723, December 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-016-0076-z.

Potuzak:2016:MBF

[1586] Jan Potužák, Jindřich Duras, and Bořek Drozd. Mass balance of fishponds: are they sources or sinks of phosphorus? Aquaculture International, 24(6):1725–1745, December 2016. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0071-4.

Hartman:2016:CCP

[1587] P. Hartman, M. Kutý, D. Hlaváč, J. Regenda, and P. Vejsada. Calcium content in pond sediment and its effect on neutralizing capacity of water and fish production. Aquaculture International, 24(6):1747– 1754, December 2016. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0087-9.

Hermelink:2017:PTM

[1588] B. Hermelink, W. Kleiner, C. Schulz, W. Kloas, and S. Wuertz. Photothermal manipulation for the reproductive management of pikeperch Sander lucioperca. Aquaculture International, 25(1):1–20, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0009-x.

Kaiser:2017:BRF

[1589] Horst Kaiser, Bernard Erasmus, and Matthew Naylor. Behavioural responses of farmed South African abalone Haliotis midae L. to disturbances caused by husbandry procedures. Aquaculture International, 25 (1):21–29, February 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0010-4.

Najafi:2017:CEF

[1590] Morteza Najafi, Bahram Falahatkar, Ali Safarpour Amlashi, and Mohammad Hossein Tolouei Gilani. The combined effects of feeding time and dietary lipid levels on growth performance in juvenile beluga sturgeon Huso huso. Aquaculture International, 25(1):31–45, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0011-3.

Shi:2017:TRS

[1591] Ce Shi, Ying Liu, MengMeng Yi, Jimeng Zheng, Huiqin Tian, and Yishuai Du. Time-restricted self-feeding causes fin damage of Atlantic salmon. Aquaculture International, 25(1):47–55, February 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0012-2.

Zhang:2017:MED

[1592] Ning Zhang, Yuehua Pu, Lijun Sun, Yaling Wang, Qi Deng, Defeng Xu, Ying Liu, Malik Hussain, and Ravi Gooneratne. Modeling the effects of different conditions on the inhibitory activity of antimicrobial lipopeptide (AMPNT-6) against Staphylococcus aureus growth and enterotoxin production in shrimp meat. Aquaculture International, 25 (1):57–70, February 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0013-1.

Mahmood:2017:SDS

[1593] Tariq Mahmood, Jianguang Fang, Zengjie Jiang, Wu Ying, and Jing Zhang. Seasonal distribution, sources and sink of dissolved organic carbon in integrated aquaculture system in coastal waters. Aquaculture International, 25(1):71–85, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0014-0.

Pucher:2017:UNN

[1594] Johannes Pucher and Ulfert Focken. Uptake of nitrogen from natural food into fish in differently managed polyculture ponds using^{15} N as tracer. Aquaculture International, 25(1):87–105, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-016-0015-z.

Tian:2017:EDP

[1595] Jingjing Tian, Ke Yang, Haojie Chen, Hong Ji, Jie Li, Wenqiang Nie, and Jishu Zhou. Effect of dietary prickly ash (Zanthoxylum bungeanum) seeds (PAS) on growth, body composition, and health of juvenile Jian carp (Cyprinus carpio var. Jian). Aquaculture International, 25(1): 107–120, February 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0016-y.

Karadal:2017:EFF

[1596] Onur Karadal, Derya Güroy, and Gürel Türkmen. Effects of feeding frequency and Spirulina on growth performance, skin coloration and seed production on kenyi cichlids (Maylandia lombardoi). Aquaculture International, 25(1):121–134, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0017-x.

Rathman:2017:IWD

[1597] Mara Rathman, Jakša Bolotin, Nikša Glavić, and Josip Barišić. Influence of water depth on growth and mortality of Chlamys varia (Linnaeus, 1758): implications for cage culture in Mali Ston Bay, Croatia. Aquaculture International, 25(1):135–146, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0018-9.

Lara:2017:UDA

[1598] Gabriele Lara, Dariano Krummenauer, Paulo C. Abreu, Luís H. Poersch, and Wilson Wasielesky. The use of different aerators on Litopenaeus vannamei biofloc culture system: effects on water quality, shrimp growth and biofloc composition. Aquaculture International, 25(1):147– 162, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-016-0019-8.

Pattanasiri:2017:ATA

[1599] Thanapat Pattanasiri, Wara Taparhudee, and Panuwat Suppakul. Acute toxicity and anaesthetic effect of clove oil and eugenol on Siamese fighting fish, Betta splendens. Aquaculture International, 25(1):163–175, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-016-0020-2.

Feucht:2017:AGG

[1600] Yvonne Feucht and Katrin Zander. Aquaculture in the German print media. Aquaculture International, 25(1):177–195, February 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0021-1.

Pattanasiri:2017:AEC

[1601] Thanapat Pattanasiri, Wara Taparhudee, and Panuwat Suppakul. Anaesthetic efficacy of clove oil-coated LDPE bag on improving water quality and survival in the Siamese fighting fish, Betta splendens, during transportation. Aquaculture International, 25(1):197–209, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-016-0022-0.

Kovacic:2017:MGL

[1602] Ines Kovačić, Dijana Pavičić-Hamer, Martin Pfannkuchen, and Matteo Usich. Mytilus galloprovincialis (Lamarck, 1819) as host of Mytilicola orientalis (Mori, 1935) in the northern Adriatic Sea: presence and effect. Aquaculture International, 25(1):211–221, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-016-0023-z.

Cecchini:2017:PUB

[1603] Stefano Cecchini and Anna Rocchina Caputo. Potential use of bovine lactoferrin as adjuvant for the modulation of the specific immune response in rainbow trout (Oncorhynchus mykiss W.). Aquaculture International, 25(1):223–232, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0024-y.

Zhang:2017:EOK

[1604] Liao Zhang, Yinglong Wu, Huailiang Xu, and Yongfang Yao. Effects of oxidized konjac glucomannan on the intestinal microbial flora and intestinal morphology of Schizothorax prenanti. Aquaculture International, 25 (1):233–250, February 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0025-x.

Ninawe:2017:ADC

[1605] A. S. Ninawe, A. S. Sahul Hameed, and J. Selvin. Advancements in diagnosis and control measures of viral pathogens in aquaculture: an Indian perspective. Aquaculture International, 25(1):251–264, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0026-9.

Ju:2017:OCI

[1606] Bao Ju, Ai li Jiang, Rong lian Xing, Li hong Chen, and Li Teng. Optimization of conditions for an integrated multi-trophic aquaculture system consisting of sea cucumber Apostichopus japonicus and ascidian Styela clava. Aquaculture International, 25(1):265–286, February 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0027-8.

Xie:2017:BEB

[1607] Xi Xie, Wen Zhao, Miao Yang, Shuxuan Zhao, and Jie Wei. Beneficial effects of benthic diatoms on growth and physiological performance in juvenile sea cucumber Apostichopus japonicus (Selenka). Aquaculture International, 25(1):287–302, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0028-7.

Kumar:2017:PEI

[1608] Saurav Kumar, T. Sathish Kumar, R. Vidya, and P. K. Pandey. A prospective of epidemiological intervention in investigation and management of argulosis in aquaculture. Aquaculture International, 25(1): 303–325, February 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0030-0.

Huang:2017:EHT

[1609] Lili Huang, Jilin Xu, Chunguang Zong, Si Zhu, Mengwei Ye, Chengxu Zhou, Haimin Chen, and Xiaojun Yan. Effect of high temperature on the lipid composition of Isochrysis galbana Parke in logarithmic phase. Aquaculture International, 25(1):327–339, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0031-z.

Cipriano-Maack:2017:IED

[1610] Ashlie N. Cipriano-Maack, Camila T. Wood, and Sarah C. Culloty. Immunostimulatory effect of diet (Laminaria digitata and Mytilus edulis) in the edible sea urchin, Paracentrotus lividus. Aquaculture International, 25(1):341–354, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0032-y.

Carrier:2017:SDK

[1611] Tyler J. Carrier, Stephen D. Eddy, and Sarah Redmond. Solar-dried kelp as potential feed in sea urchin aquaculture. Aquaculture International, 25 (1):355–366, February 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0033-x.

Ding:2017:EDM

[1612] Jing Jing Ding, Bo Huang, Ya Qiang Hu, and Xiao Bing Wang. The effects of different monospecific benthic diatoms on larval settlement, metamorphosis, survival, and growth of Haliotis asinina Linnaeus in the South China Sea. Aquaculture International, 25(1):367–377, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0035-8.

Zhang:2017:ELC

[1613] Dan Zhang, Biao Guo, Fang Wang, and Lei Jia. Effects of light color change on carbohydrate-related enzymes in Litopenaeus vannamei. Aquaculture International, 25(1):379–391, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-016-0036-7.

Kumar:2017:ESW

[1614] S. Dinesh Kumar, P. Santhanam, S. Ananth, M. Kaviyarasan, P. Nithya, B. Dhanalakshmi, Min S. Park, and Mi-Kyung Kim. Evaluation of suitability of wastewater-grown microalgae (Picochlorum maculatum) and copepod (Oithona rigida) as live feed for white leg shrimp Litopenaeus vannamei post-larvae. Aquaculture International, 25(1):393–411, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-016-0037-6.

Liu:2017:EVC

[1615] Shaoyang Liu, Tao Liao, Stephen T. McCrummen, Terrill R. Hanson, and Yifen Wang. Exploration of volatile compounds causing off-flavor in farm-raised channel catfish (Ictalurus punctatus) fillet. Aquaculture International, 25(1):413–422, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0038-5.

Shekhawat:2017:VPM

[1616] Mahipal S. Shekhawat, M. Manokari, and J. Revathi. In vitro propagation, micromorphological studies and ex vitro rooting of Alternanthera philoxeroides (Mart.) Griseb.: an important aquatic plant. Aquaculture International, 25(1):423–435, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0039-4.

Wang:2017:ETC

[1617] Chang’an Wang, Qiyou Xu, Jinnan Li, Liansheng Wang, Zhigang Zhao, Xue Du, Liang Luo, and Jiasheng Yin. Effects of tank colour on growth and survival of taimen Hucho taimen (Pallas, 1773) larvae. Aquaculture International, 25(1):437–446, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0041-x.

Wu:2017:ESS

[1618] Xugan Wu, Chaoshu Zeng, and Paul C. Southgate. Effects of starvation on survival, biomass, and lipid composition of newly hatched larvae of the blue swimmer crab, Portunus pelagicus (Linnaeus, 1758). Aquaculture International, 25(1):447–461, February 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0042-9.

Bahabadi:2017:AAA

[1619] Mahmoud Nafisi Bahabadi, Fatemeh Hosseinpour Delavar, Maryam Mirbakhsh, Khodabakhsh Niknam, and Seyed Ali Johari. Assessment of antibacterial activity of two different sizes of colloidal silver nanoparticle (cAgNPs) against Vibrio harveyi isolated from shrimp Litopenaeus vannamei. Aquaculture International, 25(1):463–472, February 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0043-8.

Rego:2017:FVI

[1620] Marcelo Augusto Soares Rego, Omar Jorge Sabbag, Roberta Soares, and Silvio Peixoto. Financial viability of inserting the biofloc technology in a marine shrimp Litopenaeus vannamei farm: a case study in the state of Pernambuco, Brazil. Aquaculture International, 25(1): 473–483, February 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0044-7.

Amin:2017:VSL

[1621] Muhamad Amin, Mark Adams, Christopher J. S. Bolch, and Christopher M. Burke. In vitro screening of lactic acid bacteria isolated from gastrointestinal tract of Atlantic salmon (Salmo salar) as probiont candidates. Aquaculture International, 25(1):485–498, February 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0045-6.

Fernandes:2017:EGP

[1622] Tiago Fernandes, Marine Herlin, María Dolores López Belluga, Guillermo Ballón, Paulino Martinez, Miguel A. Toro, and Jesús Fernández. Estimation of genetic parameters for growth traits in a hatchery population of gilthead sea bream (Sparus aurata L.). Aquaculture International, 25 (1):499–514, February 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0046-5.

Pereira:2017:UMP

[1623] T. S. B. Pereira, C. N. P. Boscolo, R. G. Moreira, and S. R. Batlouni. The use of mGnRHa provokes ovulation but not viable embryos in Leporinus macrocephalus. Aquaculture International, 25(2):515–529, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0049-2.

Hagemann:2017:EDW

[1624] Andreas Hagemann, Sigbjørn Hjetland Vorren, Yngve Attramadal, Jan Ove Evjemo, and Yngvar Olsen. Effects of different wavelengths and intensities of visible light on the hatching success of Acartia tonsa Dana eggs. Aquaculture International, 25(2):531–541, April 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0047-4.

Wan:2017:ESM

[1625] Jin juan Wan, Mei fang Shen, Jian qing Tang, Hai Lin, Wei hui Yan, Jia jia Li, and Lin Zhu. Effects of soybean meal processing treatments on growth performance, nutrient digestibility, nitrogen and phosphorus excretion in red swamp crayfish, Procambarus clarkii. Aquaculture International, 25(2):543–554, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0052-7.

Ensibi:2017:AEA

[1626] Cherif Ensibi, Charaf M’rabet, Mohamed Chalghaf, and Mohamed Nejib Daly-Yahia. Antioxidant enzyme activity and lipid peroxidation in gills of fish (Sparus aurata) upon exposure to swarms of Pelagia noctiluca. Aquaculture International, 25(2):555–565, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0056-3.

Biradar:2017:DSM

[1627] Sharanappa Biradar, H. Shivananda Murthy, Prakash Patil, E. G. Jayaraj, and Naveen Kumar B. Tammegowda. Dietary supplementation of microbial phytase improves growth and protein efficiency ratio of freshwater prawn (Macrobrachium rosenbergii). Aquaculture International, 25(2):567–575, April 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0057-2.

Boelen:2017:GPS

[1628] Peter Boelen, Audrey van Mastrigt, Henk H. van de Bovenkamp, Hero J. Heeres, and Anita G. J. Buma. Growth phase significantly decreases the DHA-to-EPA ratio in marine microalgae. Aquaculture International, 25 (2):577–587, April 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0053-6.

Jeong:2017:ECE

[1629] Dong Hyeok Jeong, Min Ho Jeong, Soo Kyung Jeong, Kwangmo Yang, and Wol Soon Jo. Effect of continuous exposure to low-dose-rate gamma irradiation on cell growth and lipid accumulation of marine microalgae. Aquaculture International, 25(2):589–601, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0054-5.

Medhioub:2017:IRT

[1630] Amel Medhioub, Badreddine Mechri, Sondes Bchir, Younes Limeyem, Wissem Slimani, Mahjoub Aouni, and Mohamed Nejib Medhioub. Impacts of rearing techniques on growth, survival and bacterial microbiota of carpet shell clam (Ruditapes decussatus) larvae. Aquaculture International, 25(2):603–617, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0055-4.

Nakphet:2017:APB

[1631] Siriporn Nakphet, Raymond J. Ritchie, and Suphada Kiriratnikom. Aquatic plants for bioremediation in red hybrid tilapia (Oreochromis niloticus × Oreochromis mossambicus) recirculating aquaculture. Aquaculture International, 25(2):619–633, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0060-7.

Jepsen:2017:RFR

[1632] Per M. Jepsen, Niels S. Bjørbæk, Thomas A. Rayner, Minh T. T. Vu, and Benni W. Hansen. Recommended feeding regime and light climate in live feed cultures of the calanoid copepod Acartia tonsa Dana. Aquaculture International, 25(2):635–654, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0063-4.

Zhan:2017:ESA

[1633] Yaoyao Zhan, Wanbin Hu, Lizhu Duan, Minbo Liu, Weijie Zhang, Yaqing Chang, and Cong Li. Effects of seawater acidification on early development of the sea urchin Hemicentrotus pulcherrimus. Aquaculture International, 25(2):655–678, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0064-3.

Gansel:2017:QBN

[1634] L. C. Gansel, N. Bloecher, O. Floerl, and J. Guenther. Quantification of biofouling on nets: a comparison of wet weight measurements and optical (image analysis) methods. Aquaculture International, 25 (2):679–692, April 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0062-5.

Arrokhman:2017:SOJ

[1635] Salim Arrokhman, Nastiti Wijayanti, and Agoes Soegianto. Survival and osmoregulation of juvenile of hybrid grouper (Epinephelus fuscoguttatus × Epinephelus lanceolatus) during acclimation in calcium-supplemented freshwater. Aquaculture International, 25(2):693–704, April 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0069-y.

Li:2017:EDC

[1636] M. Li, M. M. Rahman, B. Wu, and Y.-C. Lin. Effects of dietary canthaxanthin on growth and body colour of blood parrot cichlid Amphilophus citrinellus × Paraneetroplus synspilus. Aquaculture International, 25(2):705–713, April 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0068-z.

Golubev:2017:SMA

[1637] A. P. Golubev, A. V. Alekhnovich, O. A. Bodilovskaya, and Anilkumar Gopinathan. Some methodological approaches to the definition of limiting density for aquaculture of freshwater crayfish fingerlings. Aquaculture International, 25(2):715–725, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0065-2.

Abdy:2017:CEA

[1638] Esmail Abdy, Mojtaba Alishahi, Morteza Tollabi, Masoud Ghorbanpour, and Takavar Mohammadian. Comparative effects of Aloe vera gel and Freund’s adjuvant in vaccination of common carp (Cyprinus carpio L.) against Aeromonas hydrophila. Aquaculture International, 25 (2):727–742, April 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0074-1.

Priyaja:2017:EPT

[1639] P. Priyaja, P. Jayesh, M. Haseeb, Blessy Jose, Rosamma Philip, and I. S. Bright Singh. Evaluation of pyocyanin toxicity in various life stages of Penaeus monodon and in nitrifying bacterial consortia for their safe application in recirculating aquaculture systems (RAS) to abrogate pathogenic vibrios. Aquaculture International, 25(2):743–753, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0075-0.

Ha:2017:DCP

[1640] Natalia Ha, André Fernando Nascimento Gonçalves, Luana Camargo Sousa, Jaqueline Dalbello Biller-Takahashi, and Leonardo Susumu Takahashi. Dietary carbohydrates and protein of yeast modulate the early stages of innate immune response in tilapia (Oreochromis niloticus) primarily after LPS inoculation. Aquaculture International, 25(2):755– 776, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-016-0073-2.

He:2017:ESC

[1641] Yan fu He, Lan mei Wang, Wen bin Zhu, Zai jie Dong, and Nian Liu. Effects of salinity on cold tolerance of Malaysian red tilapia. Aquaculture International, 25(2):777–792, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0077-y.

Imanpoor:2017:EDV

[1642] Mehdi Imanpoor, Mohammad Reza Imanpoor, and Zahra Roohi. Effects of dietary vitamin C on skeleton abnormalities, blood biochemical factors, haematocrit, growth, survival and stress response of Cyprinus carpio fry. Aquaculture International, 25(2):793–803, April 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0080-3.

Josip:2017:PMG

[1643] Barišić Josip, Rozelindra Čož-Rakovac, Ivančica Delaš, Natalija Topić Popović, Ana Gavrilović, Jurica Jug-Dujaković, Marina Brailo, Roberta Sauerborn-Klobučar, Sanja Babić, and Ivančica Strunjak-Perović. Predictive modeling of European flat oyster (Ostrea edulis L.) fatty acid composition. Aquaculture International, 25(2):805–825, April 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0079-9.

Zhu:2017:PDB

[1644] Kecheng Zhu, Wenbo Yu, Jianhua Huang, Falin Zhou, Huayang Guo, Nan Zhang, Shigui Jiang, and Dianchang Zhang. Parentage determination in black tiger shrimp (Penaeus monodon) based on microsatellite DNA markers. Aquaculture International, 25(2):827–836, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0082-1.

Ali:2017:EDF

[1645] Sajjad Syed Raffic Ali, Kondusamy Ambasankar, Peter Ezhil Praveena, Sambasivam Nandakumar, and Jagabatula Syamadayal. Effect of dietary fructooligosaccharide supplementation on growth, body composition, hematological and immunological parameters of Asian seabass (Lates calcarifer). Aquaculture International, 25(2):837–848, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0081-2.

Talib:2017:BEM

[1646] Anita Talib, Kwong Kok Onn, Md Arif Chowdury, Wan Mustaffa Wan Din, and Khairun Yahya. The beneficial effects of multispecies Bacillus as probiotics in enhancing culture performance for mud crab Scylla paramamosain larval culture. Aquaculture International, 25(2):849– 866, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-016-0070-5.

Karim:2017:CMP

[1647] Manjurul Karim, Hadayet Ullah, Sarah Castine, Mohammad Mahmudul Islam, Hendrik Jan Keus, Mrityunjoy Kunda, Shakuntala Haraksingh Thilsted, and Michael Phillips. Carp–mola productivity and fish consumption in small-scale homestead aquaculture in Bangladesh. Aquaculture International, 25(2):867–879, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0078-x.

Hanh:2017:DAN

[1648] Le Minh Hanh, Van Thi Phan, Nguyen Huu Nghia, and Martin Rudbeck Jepsen. Dependency on aquaculture in northern Vietnam. Aquaculture International, 25(2):881–891, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0083-0.

Zhao:2017:ESM

[1649] Daxian Zhao, Zhenfang Mao, Qichen Lin, Chuanjie Qin, and Yijiang Hong. Effect of supplementary manganese on growth performance, tissue manganese content, antioxidant activities, and cytosolic manganese superoxide dismutase gene mRNA expression in juvenile Chinese mitten crabs (Eriocheir sinensis). Aquaculture International, 25 (2):893–903, April 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0086-x.

Zhang:2017:GSH

[1650] Lisheng Zhang, Chong Zhao, Dongtao Shi, Wanbin Hu, Jing Wei, and Yaqing Chang. Gulfweed Sargassum horneri is an alternative diet for aquaculture of juvenile sea urchins Strongylocentrotus intermedius in summer. Aquaculture International, 25(2):905–914, April 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0088-8.

Korwin-Kossakowski:2017:SMD

[1651] Michał Korwin-Kossakowski, Leszek Myszkowski, and Rafał Kamiński. A simple method to detect body morphological abnormalities in juvenile cyprinid fish — a case study on ide Leuciscus idus. Aquaculture International, 25(2):915–925, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0084-z.

Bernal:2017:PES

[1652] Milagro García Bernal, Ricardo Medina Marrero, Ángel Isidro Campa-Córdova, and José Manuel Mazón-Suástegui. Probiotic effect of Streptomyces strains alone or in combination with Bacillus and Lactobacillus in juveniles of the white shrimp Litopenaeus vannamei. Aquaculture International, 25(2):927–939, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0085-y.

Sivaramakrishnan:2017:ODL

[1653] T. Sivaramakrishnan, N. P. Sahu, K. K. Jain, A. P. Muralidhar, K. Saravanan, S. Ferosekhan, J. Praveenraj, and N. Artheeswaran. Optimum dietary lipid requirement of Pangasianodon hypophthalmus juveniles in relation to growth, fatty acid profile, body indices and digestive enzyme activity. Aquaculture International, 25(2):941–954, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-016-0090-1.

Jin:2017:DTR

[1654] Yan Jin, Fu-Jia Liu, Yong-Jian Liu, Li-Xia Tian, and Zhi-Hao Zhang. Dietary tryptophan requirements of juvenile Pacific white shrimp, Litopenaeus vannamei (Boone) reared in low-salinity water. Aquaculture International, 25(2):955–968, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0098-6.

Eysturskareth:2017:ART

[1655] Jonhard Eysturskard, Ása Johannesen, and Kirstin Eliasen. Application of real-time PCR for specific detection of Lepeophtheirus salmonis in fluid samples from lumpfish (Cyclopterus lumpus) stomachs. Aquaculture International, 25(2):969–973, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-016-0093-y.

Unger:2017:PRM

[1656] Patrick Unger and Harry Wilhelm Palm. Parasite risk of maricultured rainbow trout (Oncorhynchus mykiss Walbaum, 1792) in the Western Baltic Sea, Germany. Aquaculture International, 25(2):975– 989, April 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-016-0096-8.

Wu:2017:DVE

[1657] Fan Wu, Ming Jiang, Hua Wen, Wei Liu, Juan Tian, Chang geng Yang, and Feng Huang. Dietary vitamin E effects on growth, fillet textural parameters, and antioxidant capacity of genetically improved farmed tilapia (GIFT), Oreochromis niloticus. Aquaculture International, 25 (2):991–1003, April 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0089-7.

Pinedo-Gil:2017:EQR

[1658] Julia Pinedo-Gil, Ana Tomás-Vidal, Ana María Larrán-García, Cristina Tomás-Almenar, Miguel Jover-Cerdá, Miguel Ángel Sanz-Calvo, and Ana Belén Martín-Diana. Enhancement of quality of rainbow trout (Oncorhynchus mykiss) flesh incorporating barley on diet without negative effect on rearing parameters. Aquaculture International, 25(3): 1005–1023, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0091-0. See correction [1783].

Theodoridis:2017:TEM

[1659] Alexandros Theodoridis, Christos Batzios, Athanasios Ragkos, and Panagiotis Angelidis. Technical efficiency measurement of mussel aquaculture in Greece. Aquaculture International, 25(3):1025–1037, June 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-016-0092-z.

Yeh:2017:EAH

[1660] Chun-Yuan Yeh, Jung-Fu Huang, Jie-Min Lee, and Christian Schafferer. An economic analysis of hard clam (Meretrix meretrix) farmer polyculture with milkfish (Chanos chanos), silver sea bream (Rhabdosargus sarba), and shrimps at different hard clam stocking densities: a case study of Yunlin County, Taiwan. Aquaculture International, 25 (3):1039–1055, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0094-x.

Phuc:2017:ESS

[1661] Nguyen Trong Hong Phuc, Peter B. Mather, and David A. Hurwood. Effects of sublethal salinity and temperature levels and their interaction on growth performance and hematological and hormonal levels in tra catfish (Pangasianodon hypophthalmus). Aquaculture International, 25 (3):1057–1071, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0097-7.

Welker:2017:EDG

[1662] Thomas L. Welker, Xiao chun Wan, Yi bin Zhou, Yan ou Yang, Ken Overturf, Frederic Barrows, and Keshun Liu. Effect of dietary green tea supplementation on growth, fat content, and muscle fatty acid profile of rainbow trout (Oncorhynchus mykiss). Aquaculture International, 25 (3):1073–1094, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0099-5.

Stratev:2017:OMA

[1663] Deyan Stratev and Olumide A. Odeyemi. An overview of motile Aeromonas septicaemia management. Aquaculture International, 25(3): 1095–1105, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0100-3.

Junior:2017:ECD

[1664] O. Tomazelli Junior, F. Kuhn, P. J. Mendonça Padilha, L. R. Mota Vicente, S. Winckler da Costa, B. Corrêa da Silva, D. Dias Schleder, A. A. Boligon, J. Scapinello, C. Nunes Nesi, J. Dal Magro, and S. De Lamo Castellví. Effect of Cynodon dactylon extract on white spot virus-infected Litopenaeus vannamei. Aquaculture International, 25(3): 1107–1122, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0101-2.

Xu:2017:LPJ

[1665] Panpan Xu, Yajun Wang, Juanjuan Chen, Rui Yang, and Qicun Zhou. Lipidomic profiling of juvenile yellow head catfish (Pelteobagrus fulvidraco) in response to fucoidan diet. Aquaculture International, 25 (3):1123–1143, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0102-1.

Sun:2017:EEE

[1666] Yiyi Sun, Minghai Chen, Chongmin Kong, Huijuan Tang, Lian Gan, and Min Zhang. Enclosure experiment of effects of dietary phosphorus level on water quality, phosphorus budget, and plankton composition in intensive culture of crucian carp. Aquaculture International, 25 (3):1145–1158, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0103-0.

Sahoo:2017:MTP

[1667] L. Sahoo, P. K. Meher, K. Das Mahapatra, J. N. Saha, P. Jayasankar, and P. Das. A molecular tool for parentage analysis in Indian major carp, Labeo rohita (Hamilton, 1822). Aquaculture International, 25 (3):1159–1166, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0104-z.

Barman:2017:BBC

[1668] Prasenjit Barman, Ahmet Kati, Amit Kumar Mandal, Partha Bandyopadhyay, and Pradeep Kumar Das Mohapatra. Biopotentiality of Bacillus cereus PB45 for nitrogenous waste detoxification in ex situ model. Aquaculture International, 25(3):1167–1183, June 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0105-y.

Dai:2017:MIS

[1669] Ping Dai, Hongxia Wang, and Baozhong Liu. Mining and identification of SNP markers associated with growth traits in the clam Meretrix meretrix. Aquaculture International, 25(3):1185–1196, June 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0106-x.

Dale-Kuys:2017:GAA

[1670] Ruth Dale-Kuys, Jessica Vervalle, Rouvay Roodt-Wilding, and Clint Rhode. Genetic association analysis of candidate loci under selection with size in the South African abalone. Aquaculture International, 25 (3):1197–1214, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-016-0107-9.

Ahmad:2017:BTE

[1671] Irshad Ahmad, A. M. Babitha Rani, A. K. Verma, and Mudasir Maqsood. Biofloc technology: an emerging avenue in aquatic animal healthcare and nutrition. Aquaculture International, 25(3):1215–1226, June 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-016-0108-8.

Cerozi:2017:EDP

[1672] Brunno da Silva Cerozi and Kevin Fitzsimmons. Effect of dietary phytase on phosphorus use efficiency and dynamics in aquaponics. Aquaculture International, 25(3):1227–1238, June 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-016-0109-7.

Bautista-Teruel:2017:UAB

[1673] M. N. Bautista-Teruel, J. R. H. Maquirang, M. R. dela Peña, and V. T. Balinas. Use of agar-bound microparticulate diet as alternative food for tropical abalone, Haliotis asinina (Linnaeus 1758) post-larvae in largescale cultures. Aquaculture International, 25(3):1239–1252, June 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0110-9.

Yuan:2017:EPT

[1674] Yuan Yuan, Yongming Yuan, Yunyun Dai, and Yunchong Gong. Economic profitability of tilapia farming in China. Aquaculture International, 25(3):1253–1264, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0111-8.

Meng:2017:ILM

[1675] Xiao Meng, Jiting Wang, Wenju Wan, Mengmeng Xu, and Tingting Wang. Influence of low molecular weight chitooligosaccharides on growth performance and non-specific immune response in Nile tilapia Oreochromis niloticus. Aquaculture International, 25(3):1265–1277, June 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-017-0112-7.

Ge:2017:CGA

[1676] Hongxing Ge, Jian Li, Ping Chen, Zhiqiang Chang, Mingming Shen, and Fazhen Zhao. Cultivation of green algae Platymonas helgolandica in rearing water enhances the growth performance and resistance of Litopenaeus vannamei against Vibrio parahaemolyticus infection. Aquaculture International, 25(3):1279–1290, June 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0113-6.

Shete:2017:EDH

[1677] A. P. Shete, A. K. Verma, N. K. Chadha, Chandra Prakash, M. H. Chandrakant, and K. K. T. Nuwansi. Evaluation of different hydroponic media for mint (Mentha arvensis) with common carp (Cyprinus carpio) juveniles in an aquaponic system. Aquaculture International, 25 (3):1291–1301, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0114-5.

Chen:2017:SSP

[1678] Jianan Chen, Xiaoqin Li, Huaibing Xu, Wentong Sun, and Xiangjun Leng. Substitute of soy protein concentrate for fish meal in diets of white shrimp (Litopenaeus vannamei Boone). Aquaculture International, 25(3):1303–1315, June 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0115-4.

Basavaraja:2017:HSR

[1679] N. Basavaraja and C. H. Raghavendra. Hormonal sex reversal in red tilapia (Oreochromis niloticus and Oreochromis mossambicus) and inheritance of body colour in O. mossambicus and red tilapia: implications for commercial farming. Aquaculture International, 25(3):1317– 1331, June 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-017-0116-3.

Fu:2017:EAP

[1680] Longlong Fu, Gang Zhou, Jianlin Pan, Yuehua Li, Quanping Lu, Jun Zhou, and Xuguang Li. Effects of Astragalus polysaccharides on antioxidant abilities and non-specific immune responses of Chinese mitten crab, Eriocheir sinensis. Aquaculture International, 25(3):1333– 1343, June 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-017-0117-2.

Lin:2017:AMV

[1681] Siheng Lin, Fucun Wu, and Guofan Zhang. Analysis of monthly variation of biological conditions, metabolic parameters and antioxidative capacities in sea-based farmed Pacific abalone during summer days. Aquaculture International, 25(3):1345–1359, June 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0118-1.

Xu:2017:EDF

[1682] Huaibing Xu, Xiaoqin Li, Wentong Sun, Jianan Chen, Qipin Gao, Ke Shuai, and Xiangjun Leng. Effects of different feeding rates of extruded and pelleted feeds on growth and nutrient retention in channel catfish (Ictalurus punctatus). Aquaculture International, 25(3):1361– 1372, June 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-017-0119-0.

Stevant:2017:SAG

[1683] Pierrick Stévant, Céline Rebours, and Annelise Chapman. Seaweed aquaculture in Norway: recent industrial developments and future perspectives. Aquaculture International, 25(4):1373–1390, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0120-7.

Maji:2017:IMD

[1684] Usha Jyoti Maji, Sriprakash Mohanty, Avinash Pradhan, and Nikhil Kumar Maiti. Immune modulation, disease resistance and growth performance of Indian farmed carp, Labeo rohita (Hamilton), in response to dietary consortium of putative lactic acid bacteria. Aquaculture International, 25(4):1391–1407, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0122-5.

Miao:2017:EMA

[1685] Shuyan Miao, Longsheng Sun, Hongyi Bu, Jinyu Zhu, and Guohong Chen. Effect of molasses addition at C:N ratio of 20:1 on the water quality and growth performance of giant freshwater prawn (Macrobrachium rosenbergii). Aquaculture International, 25(4):1409–1425, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0124-3.

Ojanen:2017:CBB

[1686] Suvi Ojanen, Esa Tyystjärvi, Henrik Holmberg, Mikko Kouhia, and Pekka Ahtila. Can bacterial biofiltration be replaced by autotrophic organisms in recirculating fresh water aquaculture? Aquaculture International, 25(4):1427–1440, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0126-1.

Mohamed:2017:EDS

[1687] Khaled Mohamed, Mohamed E. Megahed, and Mervat A. M. Ali. Effect of dietary supplementation of Agrimos(R) on growth performance, feed utilization and immunological parameters of Macrobrachium rosenbergii juveniles. Aquaculture International, 25(4):1441–1452, August 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0123-4.

Harlioglu:2017:FSC

[1688] Muzaffer Mustafa Harlıoğlu and Ardavan Farhadi. Feminization strategies in crustacean aquaculture. Aquaculture International, 25(4):1453– 1468, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-017-0128-z.

Zhou:2017:EOS

[1689] Xia Zhou and Terry Hanson. Economic optimization of super-intensive biosecure recirculating shrimp production systems. Aquaculture International, 25(4):1469–1483, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0129-y.

Can:2017:ETN

[1690] Safak Seyhaneyildiz Can, Edis Koru, and Semra Cirik. Effect of temperature and nitrogen concentration on the growth and lipid content of Spirulina platensis and biodiesel production. Aquaculture International, 25 (4):1485–1493, August 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0121-6.

Nandakumar:2017:RFM

[1691] Sambasivam Nandakumar, Kondusamy Ambasankar, Sajjad Syed Raffic Ali, Jagabatula Syamadayal, and Kumaraguru Vasagam. Replacement of fish meal with corn gluten meal in feeds for Asian seabass (Lates calcarifer). Aquaculture International, 25(4):1495–1505, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0133-2.

Liu:2017:CMS

[1692] Ting Liu, Qi Li, Lingfeng Kong, and Hong Yu. Comparison of microsatellites and SNPs for pedigree analysis in the Pacific oyster Crassostrea gigas. Aquaculture International, 25(4):1507–1519, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0127-0.

Wang:2017:RGC

[1693] Pingping Wang and Jianyue Ji. Research on China’s mariculture efficiency evaluation and influencing factors with undesirable outputs — an empirical analysis of China’s ten coastal regions. Aquaculture International, 25(4):1521–1530, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0131-4.

Tziouveli:2017:CFA

[1694] Vasiliki Tziouveli and Saichiro Yokoyama. A comparison of the fatty acid profiles of newly hatched, fed, and starved juveniles of Amphioctopus fangsiao (d’Orbigny 1839). Aquaculture International, 25(4): 1531–1542, August 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0130-5.

NaveenKumar:2017:BIR

[1695] Singaiah NaveenKumar, Mohamed A. Hassan, Mahmoud A. Mahmoud, Adly Al-Ansari, and Waleed K. Al-Shwared. Betanodavirus infection in reared marine fishes along the Arabian Gulf. Aquaculture International, 25(4):1543–1554, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0134-1.

Amarouayache:2017:FAC

[1696] Mounia Amarouayache, Yavuz Selim Cakmak, Meltem Asan-Ozusaglam, and Abdelkader Amorouayeche. Fatty acid composition of five Algerian bisexual and parthenogenetic strains of Artemia (Anostraca, Crustacea) and their antimicrobial activity. Aquaculture International, 25(4): 1555–1568, August 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0136-z.

Rajabzadeh:2017:MPA

[1697] Nadia Rajabzadeh, Mohsen Naeemipour, and Mohsen Seyedabadi. Multiplex PCR assay for the simultaneous detection of bacterial pathogens in rainbow trout. Aquaculture International, 25(4):1569–1575, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0135-0.

Zhang:2017:ETS

[1698] Tao Zhang, Hao Song, Yu-Cen Bai, Jing-Chun Sun, Xiao-Fang Zhang, Shao-Jun Ban, Zheng-Lin Yu, Mei-Jie Yang, and Hai-Yan Wang. Effects of temperature, salinity, diet and stocking density on development of the veined rapa whelk, Rapana venosa (Valenciennes, 1846) larvae. Aquaculture International, 25(4):1577–1590, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0140-3.

Li:2017:IED

[1699] Xiaoxia Li, Liliu Zhou, Huangyan Mo, Qing Pan, and Lian Gan. Interaction effects of dietary lipid and lysine on growth feed utilization and body composition of juvenile grass carp (Ctenopharyngodon idella). Aquaculture International, 25(4):1591–1606, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0132-3.

Zhang:2017:MGC

[1700] Jing Zhang, Fen Wang, Ye-Lin Jiang, Guan-Jun Hou, Yun-Sheng Cheng, Hong-Lian Chen, and Xiang Li. Modern greenhouse culture of juvenile soft-shelled turtle, Pelodiscus sinensis. Aquaculture International, 25(4): 1607–1624, August 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0137-y.

Guo:2017:BLO

[1701] Yanbiao Guo, Qi Pan, Shiqi Yan, Yuhong Chen, MingJia Li, Dan Chen, Hongcao Han, Bing Wu, and Junpeng Cai. Bdellovibrio and like organisms promoted growth and survival of juvenile abalone Haliotis discus hannai Ino and modulated bacterial community structures in its gut. Aquaculture International, 25(4):1625–1643, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0138-x.

Wang:2017:ESB

[1702] Xiaozhu Wang, Shikai Liu, Rex Dunham, and Zhanjiang Liu. Effects of strain and body weight on low-oxygen tolerance of channel catfish (Ictalurus punctatus). Aquaculture International, 25(4):1645– 1652, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-017-0125-2.

Freites:2017:GSW

[1703] L. Freites, M. Rojas, A. W. Dale, A. Márquez, J. Revilla, L. Troccoli, and C. Lodeiros. Growth and survival of the winged oyster Pteria colymbus in suspended culture: influence of environmental factors associated to upwelling periods. Aquaculture International, 25(4): 1653–1666, August 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0141-2.

Ngambi:2017:IES

[1704] Jonas Wiza Ng’ambi, Ronghua Li, Changkao Mu, Weiwei Song, and Chunlin Wang. The immunostimulatory effect of saponin immersion against Vibrio alginolyticus in swimming crab Portunus trituberculatus. Aquaculture International, 25(4):1667–1678, August 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0139-9.

Thammasorn:2017:PBL

[1705] Thitiporn Thammasorn, Sarocha Jitrakorn, Patai Charoonnart, Suwitchaya Sirimanakul, Triwit Rattanarojpong, Soraya Chaturongakul, and Vanvimon Saksmerprome. Probiotic bacteria (Lactobacillus plantarum) expressing specific double-stranded RNA and its potential for controlling shrimp viral and bacterial diseases. Aquaculture International, 25(5): 1679–1692, October 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0144-z.

Jannathulla:2017:EPS

[1706] Rajabdeen Jannathulla, Jagabattula Syama Dayal, Kondusamy Ambasankar, Hajah Imran Khan, Eda Purudhvi Madhubabu, and Moturi Muralidhar. Effect of protein solubility of soybean meal on growth, digestibility and nutrient utilization in Penaeus vannamei. Aquaculture International, 25(5):1693–1706, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0147-9.

Wang:2017:CNE

[1707] Jiulong Wang and Qi Li. Characterization of novel EST–SNP markers and their association analysis with growth-related traits in the Pacific oyster Crassostrea gigas. Aquaculture International, 25(5):1707– 1719, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-017-0142-1.

Bao:2017:AGI

[1708] Ning Bao, Tongjun Ren, Yuzhe Han, Fuqiang Wang, Fei Chen, and Zhiqiang Jiang. Alteration of growth, intestinal Lactobacillus, selected immune and digestive enzyme activities in juvenile sea cucumber Apostichopus japonicus, fed dietary multiple probiotics. Aquaculture International, 25(5):1721–1731, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0148-8.

David:2017:NBI

[1709] Fernanda S. David, Danilo C. Proença, and Wagner C. Valenti. Nitrogen budget in integrated aquaculture systems with Nile tilapia and Amazon River prawn. Aquaculture International, 25(5):1733–1746, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0145-y.

Treasurer:2017:ODA

[1710] Jim Treasurer. Ovary development and annual egg production of hake, M. merluccius: a promising aquaculture species. Aquaculture International, 25(5):1747–1759, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0146-x.

Taju:2017:DRV

[1711] G. Taju, N. Madan, S. Abdul Majeed, T. Rajkumar, M. A. Farook, S. Vimal, A. Nazeer Basha, and A. S. Sahul Hameed. Distribution of recombinant VP28 protein in tissues and its immunomodulatory effect against white spot syndrome virus in whiteleg shrimp, Litopenaeus vannamei (Boone, 1931). Aquaculture International, 25(5):1761– 1776, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-017-0155-9.

Roncarati:2017:GSC

[1712] Alessandra Roncarati, Alberto Felici, Gian Enrico Magi, Nina Bilandžić, and Paolo Melotti. Growth and survival of cupped oysters (Crassostrea gigas) during nursery and pregrowing stages in open sea facilities using different stocking densities. Aquaculture International, 25(5): 1777–1785, October 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0152-z.

Pascual:2017:DSG

[1713] Mariano M. Pascual, Juan P. Hualde, Virginia A. Bianchi, Juan M. Castro, and Carlos M. Luquet. Diet supplemented with Grifola gargal mushroom enhances growth, lipid content, and nutrient retention of juvenile rainbow trout (Oncorhynchus mykiss). Aquaculture International, 25(5): 1787–1797, October 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0154-x.

Omondi:2017:GPM

[1714] George Alfred Ajuoga Omondi, Elick O. Otachi, and Jonathan M. Munguti. Growth performance of mixed sex Nile tilapia in cage monoculture and polyculture with African catfish and African carp. Aquaculture International, 25(5):1799–1812, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0143-0.

Eysturskareth:2017:FVE

[1715] Jonhard Eysturskard, Sunnvør í Kongsstovu, Daisy Færø, Ása Jacobsen, and Hóraldur Joensen. Fucus vesiculosus extract inhibits the proteolytic activity and gene expression of matrix metalloproteinases in Atlantic salmon (Salmo salar L.). Aquaculture International, 25(5): 1813–1819, October 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0157-7.

Bu:2017:EFR

[1716] Xianyong Bu, Xuqiu Lian, Ying Zhang, Chenghui Yang, Cunhe Cui, Jianfang Che, Baibing Tang, Baohui Su, Qicun Zhou, and Yuhong Yang. Effects of feeding rates on growth, feed utilization, and body composition of juvenile Pseudobagrus ussuriensis. Aquaculture International, 25(5): 1821–1831, October 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0156-8.

Porchas-Cornejo:2017:HRD

[1717] Marco A. Porchas-Cornejo, Marcel Martínez-Porchas, Francisco Vargas-Albores, Teresa Gollas-Galvan, Luis Rafael Martínez-Córdova, Roberto Vazquez-Euan, and Emilio Peña-Messina. High-resolution detection of bacterial profile of ocean water, before and after being used by shrimp farms. Aquaculture International, 25(5):1833–1843, October 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0160-z.

Wang:2017:NOE

[1718] Zi-Chao Wang, Mou-Ming Zhao, Zhongxiang Fang, and De-Wei Chen. Nitric oxide euthanasia: a potential procedure for improving animal welfare and fillet color of tilapia (Oreochromis niloticus). Aquaculture International, 25(5):1845–1856, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0162-x.

Sotoudeh:2017:EDS

[1719] Ebrahim Sotoudeh and Marziyeh Jafari. Effects of dietary supplementation with red seaweed, Gracilaria pygmaea, on growth, carcass composition and hematology of juvenile rainbow trout, Oncorhynchus mykiss. Aquaculture International, 25(5):1857–1867, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0158-6.

Aranda:2017:ENF

[1720] Dalila Aldana Aranda and Nancy Brito Manzano. Effects of near-future-predicted ocean temperatures on early development and calcification of the queen conch Strombus gigas. Aquaculture International, 25(5): 1869–1881, October 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0153-y.

Akber:2017:CSF

[1721] Md. Ali Akber, Md. Atikul Islam, Munir Ahmed, Md. Munsur Rahman, and Mohammad Rezaur Rahman. Changes of shrimp farming in southwest coastal Bangladesh. Aquaculture International, 25(5):1883– 1899, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-017-0159-5.

Zhang:2017:GPG

[1722] Jiachen Zhang, Fujun Cao, Jianyong Liu, and Ruipeng Yuan. Genetic parameters for growth and survival traits in Litopenaeus vannamei at different ages. Aquaculture International, 25(5):1901–1911, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0163-9.

Soltani:2017:GPI

[1723] Mehdi Soltani, Esmail Abdy, Mojtaba Alishahi, Ali Taheri Mirghaed, and Pezhman Hosseini-Shekarabi. Growth performance, immunephysiological variables and disease resistance of common carp (Cyprinus carpio) orally subjected to different concentrations of Lactobacillus plantarum. Aquaculture International, 25(5):1913–1933, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0164-8.

Ghosh:2017:SEP

[1724] S. Ghosh, S. Manna, N. C. Sahu, A. Dutta, and R. Goswami. Social, economic and production characteristics of freshwater prawn, Macrobrachium rosenbergii (De Man, 1879) culture in West Bengal, India. Aquaculture International, 25(5):1935–1957, October 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0165-7.

Lara:2017:UBD

[1725] Gabriele Lara, Marcelo Honda, Luís Poersch, and Wilson Wasielesky. The use of biofilm and different feeding rates in biofloc culture system: the effects in shrimp growth parameters. Aquaculture International, 25 (5):1959–1970, October 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0151-0.

Mercaldo-Allen:2017:ECD

[1726] Renee Mercaldo-Allen, Shannon Meseck, Ronald Goldberg, Paul Clark, Catherine Kuropat, and Julie M. Rose. Effects of clam dredging on benthic ecology of two cultivated northern quahog beds with different harvest histories and sediment grain sizes. Aquaculture International, 25 (5):1971–1985, October 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0167-5. See correction [1749].

Zaefarian:2017:DEG

[1727] Azadeh Zaefarian, Sakineh Yeganeh, and Batoul Adhami. Dietary effects of garlic powder (Allium sativum) on growth, blood indices, carcass composition, and lysozyme activity in brown trout (Salmo caspius) and resistance against Yersinia ruckeri infection. Aquaculture International, 25(6):1987–1996, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0169-3.

Zhu:2017:HGD

[1728] W. B. Zhu, H. Yang, X. H. Yuan, Z. J. Dong, J. J. Fu, L. M. Wang, S. Y. Su, N. Liu, F. B. Song, and X. T. Chen. High genetic diversity and differentiation in three red tilapia stocks revealed by microsatellite DNA marker analysis. Aquaculture International, 25(6): 1997–2006, December 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0150-1.

Zhang:2017:ELM

[1729] Liao Zhang, Yinglong Wu, Ping Si, Yongfeng Yan, Huailiang Xu, and Yongfang Yao. Effects on lipid metabolism and expression of PPARα and FABP of Schizothorax prenanti by oxidized konjac glucomannan. Aquaculture International, 25(6):2007–2025, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0168-4.

Barreiro:2017:CIC

[1730] L. Barreiro, R. Caamaño, S. Cabaleiro, M. V. Ruiz de Ocenda, and F. Villoch. Ceratomyxosis infection in cultured striped red mullet (Mullus surmuletus Linnaeus 1786) broodstock. Aquaculture International, 25(6): 2027–2034, December 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0166-6.

daPaixao:2017:ECP

[1731] Alison Eduardo Melo da Paixão, Jéssica Cerqueira dos Santos, Mariana Sampaio Pinto, Denise Soledade Peixoto Pereira, Carlos Eduardo Crispim de Oliveira Ramos, Robson Bahia Cerqueira, Rodrigo Diana Navarro, and Rodrigo Fortes da Silva. Effect of commercial probiotics (Bacillus subtilis and Saccharomyces cerevisiae) on growth performance, body composition, hematology parameters, and disease resistance against Streptococcus agalactiae in tambaqui (Colossoma macropomum). Aquaculture International, 25(6):2035–2045, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0173-7.

Nehlah:2017:RVP

[1732] R. Nehlah, M. Firdaus-Nawi, N. Y. Nik-Haiha, M. Karim, M. Zamri-Saad, and M. Y. Ina-Salwany. Recombinant vaccine protects juvenile hybrid grouper, Epinephelus fuscoguttatus × Epinephelus lanceolatus, against infection by Vibrio alginolyticus. Aquaculture International, 25 (6):2047–2059, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-017-0172-8.

Wang:2017:SRM

[1733] Chunling Wang, Lingli Jiang, Guoying Qian, and Youling Gao. Supplying rapeseed meal to the diets with or without potassium iodide for yellow catfish (Tachysurus fulvidraco). Aquaculture International, 25(6): 2061–2078, December 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0171-9.

Bai:2017:EGP

[1734] Zhiyi Bai, Qingqing Li, Xuekai Han, and Jiale Li. Estimates of genetic parameters and genotype by environment interactions for shell nacre color and growth traits in the purple freshwater pearl mussel Hyriopsis cumingii. Aquaculture International, 25(6):2079–2090, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0170-x.

Wang:2017:EDC

[1735] Longtao Wang, Chenxia Ge, Jianchao Wang, Jing Dai, Peijun Zhang, and Yuehong Li. Effects of different combinations of Bacillus on immunity and antioxidant activities in common carp. Aquaculture International, 25(6):2091–2099, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0175-5.

Kumar:2017:DSA

[1736] Pankaj Kumar, K. K. Jain, P. Sardar, N. P. Sahu, and S. Gupta. Dietary supplementation of acidifier: effect on growth performance and haemato-biochemical parameters in the diet of Cirrhinus mrigala juvenile. Aquaculture International, 25(6):2101–2116, December 2017. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0176-4.

Li:2017:SGV

[1737] Pengfei Li, Hui Zhang, Xiumei Zhang, Tianxiang Gao, and Zhiqiang Han. Study on the genetic variability of the hatchery-released and wild populations of Chinese white shrimp Fenneropenaeus chinensis in the Yellow Sea and Bohai Sea. Aquaculture International, 25(6):2117– 2126, December 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0174-6.

Kurita:2017:PES

[1738] Yoshihisa Kurita, Ikuo Chiba, and Akihiro Kijima. Physical eradication of small planktonic crustaceans from aquaculture tanks with cavitation treatment. Aquaculture International, 25(6):2127–2133, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0179-1.

Wang:2017:EDP

[1739] Chang’an Wang, Liansheng Wang, Jinnan Li, Dongli Qin, Zhigang Zhao, Liang Luo, Xue Du, Qiyou Xu, and Jiasheng Yin. Effects of dietary phosphorus on growth, body composition, and blood chemistry of juvenile taimen Hucho taimen. Aquaculture International, 25(6):2135– 2148, December 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0178-2.

Deb:2017:ESE

[1740] Saptashish Deb, Md. T. Noori, and P. Srinivasa Rao. Experimental study to evaluate the efficacy of locally available waste carbon sources on aquaculture water quality management using biofloc technology. Aquaculture International, 25(6):2149–2159, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0180-8.

Mohanty:2017:EVS

[1741] Rajeeb K. Mohanty, Atmaram Mishra, Sunil K. Ambast, Krishna Gopal Mandal, Dileep K. Panda, and Pravukalyan Panigrahi. Effects of various shrimp (Penaeus monodon) densities on their growth, water and sediment quality, and water budget. Aquaculture International, 25(6): 2161–2176, December 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0181-7.

Luo:2017:CGS

[1742] Kun Luo, Jie Kong, Xianhong Meng, Sheng Luan, Baolong Chen, and Baoxiang Cao. Comparison of growth and survival among selected population, inbreeding population, and wild population in Fenneropenaeus chinensis. Aquaculture International, 25(6):2177–2188, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0177-3.

Dai:2017:GEF

[1743] Ping Dai, Sheng Luan, Xia Lu, Kun Luo, Baoxiang Cao, Xianhong Meng, and Jie Kong. Genetic evaluation of feed efficiency in the breeding population of Fenneropenaeus chinensis “Huanghai No. 2” using phenotypic, pedigree and genomic information. Aquaculture International, 25(6): 2189–2200, December 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0182-6.

Wang:2017:NSI

[1744] Fang Wang, Xuguang Wang, Chun Liu, Ouqin Chang, Yongyong Feng, Lang Jiang, and Kaibin Li. Nocardia seriolae infection in cultured jade perch, Scortum barcoo. Aquaculture International, 25(6):2201– 2212, December 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0184-4.

Hungria:2017:GSP

[1745] Diogo Barbalho Hungria, Camila Prestes dos Santos Tavares, Leandro Ângelo Pereira, Ubiratã de Assis Teixeira da Silva, and Antonio Ostrensky. Global status of production and commercialization of softshell crabs. Aquaculture International, 25(6):2213–2226, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0183-5.

Wang:2017:SDA

[1746] Chen Wang, Yan Zhao, Yong ping Wei, and Guo qing Shi. Systems dynamics approach to understanding the impacts of aquaculture closure policies on environmental pressures and fishermen households’ incomes in Danjiangkou Reservoir, China. Aquaculture International, 25(6):2227– 2257, December 2017. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0185-3.

Kuradomi:2017:ESI

[1747] Rafael Yutaka Kuradomi, Fausto Foresti, and Sergio Ricardo Batlouni. The effects of sGnRHa implants on Piaractus mesopotamicus female breeders. an approach addressed to aquaculture. Aquaculture International, 25(6):2259–2273, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0186-2.

Challouf:2017:EAI

[1748] Rafika Challouf, Asma Hamza, Mabrouka Mahfoudhi, Khemissa Ghozzi, and Mohamed Nejmeddine Bradai. Environmental assessment of the impact of cage fish farming on water quality and phytoplankton status in Monastir Bay (eastern coast of Tunisia). Aquaculture International, 25(6):2275–2292, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-017-0187-1.

Mercaldo-Allen:2017:CEC

[1749] Renee Mercaldo-Allen, Shannon Meseck, Ronald Goldberg, Paul Clark, Catherine Kuropat, and Julie M. Rose. Correction to: Effects of clam dredging on benthic ecology of two cultivated northern quahog beds with different harvest histories and sediment grain sizes. Aquaculture International, 25(6):2293–2294, December 2017. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0208-0. See [1726].

AndradeRodrigues:2018:DAA

[1750] Robson Andrade Rodrigues, Cleujosí da Silva Nunes, Leticia Emiliani Fantini, Rodrigo Yutaka Dichoff Kasai, Carlos Antonio Lopes Oliveira, Hamilton Hisano, and Cristiane Meldau de Campos. Dietary ascorbic acid influences the intestinal morphology and hematology of hybrid sorubim catfish (Pseudoplatystoma reticulatum × P. corruscans). Aquaculture International, 26(1):1–11, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0188-0. See [1751].

Rodrigues:2018:EDA

[1751] Robson Andrade Rodrigues, Cleujosí da Silva Nunes, Leticia Emiliani Fantini, Rodrigo Yutaka Dichoff Kasai, Carlos Antonio Lopes Oliveira, Hamilton Hisano, and Cristiane Meldau de Campos. Erratum to: Dietary ascorbic acid influences the intestinal morphology and hematology of hybrid sorubim catfish (Pseudoplatystoma reticulatum × P. corruscans). Aquaculture International, 26(1):13–14, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-017-0195-1. See [1750].

Zhou:2018:AET

[1752] Fan Zhou, Ya-Qin Wang, Yi-Jiang Bei, Wing-Keong Ng, Ding-Nan Wang, Shi-Yan Li, Qing-Hui Meng, and Xue-Yan Ding. Assessing the efficacy of three methionine sources in low protein and low fish meal diet for Chinese soft-shelled turtle, Pelodiscus sinensis. Aquaculture International, 26(1):15–26, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0193-3.

Kujawa:2018:RRL

[1753] Roman Kujawa, Dorota Fopp-Bayat, Beata I. Cejko, Dariusz Kucharczyk, Katarzyna Glińska-Lewczuk, Krystian Obolewski, and Mateusz Biegaj. Rearing river lamprey Lampetra fluviatilis (L.) larvae under controlled conditions as a tool for restitution of endangered populations. Aquaculture International, 26(1):27–36, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0190-6.

Chaoruangrit:2018:AMD

[1754] Lalida Chaoruangrit, Paveena Tapaneeyaworawong, Sorawit Powtongsook, and La orsri Sanoamuang. Alternative microalgal diets for cultivation of the fairy shrimp Branchinella thailandensis (Branchiopoda: Anostraca). Aquaculture International, 26(1):37–47, February 2018. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0191-5.

Jonsdottir:2018:PGS

[1755] Ólöf Dóra Bartels Jónsdóttir, Julia Schregel, Snorre B. Hagen, Camilla Tobiassen, Siv Grethe Aarnes, and Albert K. D. Imsland. Population genetic structure of lumpfish along the Norwegian coast: aquaculture implications. Aquaculture International, 26(1):49–60, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:/

/link.springer.com/article/10.1007/s10499-017-0194-2. See cor-

rection [1756].

Jonsdottir:2018:CPG

[1756] Ólöf Dóra Bartels Jónsdóttir, Julia Schregel, Snorre B. Hagen, Camilla Tobiassen, Siv Grethe Aarnes, and Albert K. D. Imsland. Correction to: Population genetic structure of lumpfish along the Norwegian coast: aquaculture implications. Aquaculture International, 26 (1):61, February 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0197-z. See [1755].

Liu:2018:DSP

[1757] Chun-Hung Liu, Kuanchih Wu, Tah-Wei Chu, and Tsung-Meng Wu. Dietary supplementation of probiotic, Bacillus subtilis E20, enhances the growth performance and disease resistance against Vibrio alginolyticus in parrot fish (Oplegnathus fasciatus). Aquaculture International, 26 (1):63–74, February 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0189-z.

Sudagar:2018:EDP

[1758] Mohammad Sudagar, Saeide Keivanloo, and Abasali Hajibeglou. Effect of different permeable and non-permeable cryoprotectants on the hatching rate of rainbow trout (Oncorhynchus mykiss) embryos. Aquaculture International, 26(1):75–84, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0192-4.

deAraujo:2018:PPS

[1759] Ednara Ronise Lima de Araújo, Luis André Luz Barbas, Carlos Massatoshi Ishikawa, Danielle de Carla Dias, Fábio Rosa Sussel, Hélcio Luis de Almeida Marques, and Leonardo Tachibana. Prebiotic, probiotic, and synbiotic in the diet of Nile tilapia post-larvae during the sex reversal phase. Aquaculture International, 26(1):85–97, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0201-7.

Pinho:2018:IPF

[1760] Sara Mello Pinho, Giovanni Lemos de Mello, Kevin M. Fitzsimmons, and Maurício Gustavo Coelho Emerenciano. Integrated production of fish (pacu Piaractus mesopotamicus and red tilapia Oreochromis sp.) with two varieties of garnish (scallion and parsley) in aquaponics system. Aquaculture International, 26(1):99–112, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0198-y.

Zhang:2018:ISM

[1761] Jingxiao Zhang, Qi Li, Qingzhi Wang, Rihao Cong, Jianlong Ge, and Lingfeng Kong. The impact of successive mass selection on population genetic structure in the Pacific oyster (Crassostrea gigas) revealed by microsatellite markers. Aquaculture International, 26(1):113–125, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-017-0196-0.

Rodrigues:2018:MFN

[1762] Marianna Vaz Rodrigues, Claire Juliana Francisco, Gianmarco S. David, Reinaldo José da Silva, Maria Fernanda Falcone-Dias, and João Pessoa Araújo Júnior. Monitoring of Francisella noatunensis subsp. orientalis in farmed Nile tilapia (Oreochromis niloticus) in Brazil. Aquaculture International, 26(1):127–138, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0204-4.

Ali:2018:EDB

[1763] Tamer El-Sayed Ali, Abdelfattah Mohamed El-Sayed, Mohamed Abdel-Razek Eissa, and Hebatollah Moustafa Hanafi. Effects of dietary Biogen and sodium butyrate on hematological parameters, immune response, and histological characteristics of Nile tilapia (Oreochromis niloticus) fingerlings. Aquaculture International, 26(1):139–150, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/

s10499-017-0205-3. See correction [1764].

El-Sayed:2018:CED

[1764] Abdel-Fattah Mohamed El-Sayed, Tamer El-Sayed Ali, Mohamed Abdel-Razek Eissa, and Hebatollah Moustafa Almisherfi. Correction to: Effects of dietary Biogen and sodium butyrate on hematological parameters, immune response, and histological characteristics of Nile tilapia (Oreochromis niloticus) fingerlings. Aquaculture International, 26(1): 151, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-017-0221-3. See [1763].

Liu:2018:EDP

[1765] Wei Liu, Hua Wen, and Zhi Luo. Effect of dietary protein levels and feeding rates on the growth and health status of juvenile genetically improved farmed tilapia (Oreochromis niloticus). Aquaculture International, 26(1):

153–167, February 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0202-6.

Goda:2018:EDS

[1766] A. M. Goda, E. A. Omar, T. M. Srour, A. M. Kotiet, E. El-Haroun, and Simon J. Davies. Effect of diets supplemented with feed additives on growth, feed utilization, survival, body composition and intestinal bacterial load of early weaning European seabass, Dicentrarchus labrax post-larvae. Aquaculture International, 26(1):169–183, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0200-8.

Prato:2018:EFD

[1767] Ermelinda Prato, Mariachiara Chiantore, Maeve S. Kelly, Adam D. Hughes, Philip James, Maria Paola Ferranti, Francesca Biandolino, Isabella Parlapiano, Benedetto Sicuro, and Giovanni Fanelli. Effect of formulated diets on the proximate composition and fatty acid profiles of sea urchin Paracentrotus lividus gonad. Aquaculture International, 26 (1):185–202, February 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0203-5.

Maciel:2018:CGF

[1768] Julio Cesar Maciel, Claire Juliana Francisco, and Kleber Campos Miranda-Filho. Compensatory growth and feed restriction in marine shrimp production, with emphasis on biofloc technology. Aquaculture International, 26(1):203–212, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0209-z.

Hisano:2018:DGM

[1769] Hamilton Hisano, Michelly Pereira Soares, Fabiana Golin Luiggi, and Arielle Cristina Arena. Dietary β-glucans and mannanoligosaccharides improve growth performance and intestinal morphology of juvenile pacu Piaractus mesopotamicus (Holmberg, 1887). Aquaculture International, 26(1):213–223, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-017-0210-6.

Bruno:2018:ISB

[1770] Eleonora Bruno, Jacob Kring Højgaard, Benni Winding Hansen, Peter Munk, and Josianne Gatt Støttrup. Influence of swimming behavior of copepod nauplii on feeding of larval turbot (Scophthalmus maximus). Aquaculture International, 26(1):225–236, February 2018. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0199-x.

Jakobsen:2018:CCP

[1771] H. H. Jakobsen, C. Thoisen, and B. W. Hansen. Crypthecodinium cohnii: a promising prey toward large-scale intensive rearing of the live feed copepod Acartia tonsa (Dana). Aquaculture International, 26(1): 237–251, February 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0207-1.

Ajjabi:2018:RMG

[1772] Leila Chebil Ajjabi, Mouna Abaab, and Raafa Segni. The red macroalga Gracilaria verrucosa in co-culture with the Mediterranean mussels Mytilus galloprovincialis: productivity and nutrient removal performance. Aquaculture International, 26(1):253–266, February 2018. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0206-2.

Ojerio:2018:AAI

[1773] Vianney T. Ojerio, Valeriano L. Corre, Nieves A. Toledo, Karen Grace S. Andrino-Felarca, Lovelyn Marie Nievales, and Rex Ferdinand M. Traifalgar. Alginic acid as immunostimulant: effects of dose and frequency on growth performance, immune responses, and white spot syndrome virus resistance in tiger shrimp Penaeus monodon (Fabricius, 1798). Aquaculture International, 26(1):267–278, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0212-4.

Guerreiro:2018:SCG

[1774] Inês Guerreiro, Cláudia R. Serra, Aires Oliva-Teles, and Paula Enes. Short communication: gut microbiota of European sea bass (Dicentrarchus labrax) is modulated by short-chain fructooligosaccharides and xylooligosaccharides. Aquaculture International, 26(1):279–288, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-017-0220-4.

Stratev:2018:BEM

[1775] Deyan Stratev, Georgi Zhelyazkov, Xavier Siwe Noundou, and Rui W. M. Krause. Beneficial effects of medicinal plants in fish diseases. Aquaculture International, 26(1):289–308, February 2018. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0219-x.

Huang:2018:SCM

[1776] Yu-Juan Huang, Nan-Nan Zhang, Wu-Jiang Fan, Yan-Yan Cui, Samwel Mchele Limbu, Fang Qiao, Yun-Long Zhao, Li-Qiao Chen, Zhen-Yu Du, and Dong-Liang Li. Soybean and cottonseed meals are good candidates for fishmeal replacement in the diet of juvenile Macrobrachium nipponense. Aquaculture International, 26(1):309–324, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0215-1.

Okutsu:2018:UAS

[1777] Tomoyuki Okutsu, Phonenaphet Chanthasone, Phutsamone Phommachan, Aloun Kounthongbang, Oulaytham Lasasimma, Koji Hamada, Shinsuke Morioka, and Sayaka Ito. Use of artificial seawater in the rearing of the fluvial prawn Macrobrachium yui larvae. Aquaculture International, 26(1):325–335, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0218-y.

Lodeiros:2018:SCG

[1778] C. Lodeiros, D. Rodríguez-Pesantes, A. Márquez, J. Revilla, J. Chávez-Villalba, and S. Sonnenholzner. Suspended cultivation of the Pacific oyster Crassostrea gigas in the Eastern Tropical Pacific. Aquaculture International, 26(1):337–347, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0217-z.

Goncalves:2018:DPC

[1779] André Fernando Nascimento Gonçalves, Natalia Ha, Jaqueline Dalbello Biller-Takahashi, Rodrigo Yukihiro Gimbo, Elisabeth Criscuolo Urbinati, and Leonardo Susumu Takahashi. Dietary protein-to-carbohydrate ratios affect metabolism and growth of juvenile surubim cachara (Pseudoplatystoma reticulatum). Aquaculture International, 26(1):349–362, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-017-0213-3.

Das:2018:NEA

[1780] Mousumi Das, Xuan Li, Nguyen Viet Dung, Peter Bossier, and Nguyen Thi Ngoc Tinh. No effect of N-acyl homoserine lactones disruption by lactonase enzyme on the virulence of Vibrio anguillarum towards sea bass (Dicentrarchus labrax) and brine shrimp (Artemia franciscana). Aquaculture International, 26(1):363–374, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0214-2.

Jayant:2018:PRJ

[1781] Manish Jayant, Ande Penchala Muralidhar, Narrottam Prasad Sahu, Kamal Kant Jain, Asim Kumar Pal, and Prem Prakash Srivastava. Protein requirement of juvenile striped catfish, Pangasianodon hypophthalmus. Aquaculture International, 26(1):375–389, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0216-0.

Falahatkar:2018:PSL

[1782] Bahram Falahatkar, Iraj Efatpanah, and Patrick Kestemont. Pikeperch Sander lucioperca production in the south part of the Caspian Sea: technical notes. Aquaculture International, 26(1):391–401, February 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0222-2.

Pinedo-Gil:2018:CEQ

[1783] Julia Pinedo-Gil, Ana Tomás-Vidal, Ana María Larrán-García, Cristina Tomás-Almenar, Miguel Jover-Cerdá, Miguel Ángel Sanz-Calvo, and Ana Belén Martín-Diana. Correction to: Enhancement of quality of rainbow trout (Oncorhynchus mykiss) flesh incorporating barley on diet without negative effect on rearing parameters. Aquaculture International, 26(1):403, February 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0211-5. See [1658].

Ghosh:2018:EFL

[1784] Alokesh Kumar Ghosh. Effect of feeding level on growth, body composition, fatty acid profile, and nutrient accumulation in shrimp (Litopenaeus vannamei). Aquaculture International, 26(2):405–417, April 2018. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0225-z.

Sahrir:2018:EEE

[1785] Muhamad Syahmin Aiman Sahrir, Mushrifah Idris, Abdullah Samat, and Suhairi Alimon. Effects of electrolyte enhanced water on culturing giant freshwater prawn Macrobrachium rosenbergii. Aquaculture International, 26(2):419–431, April 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0223-1.

Kritsanapuntu:2018:EEB

[1786] Sirusa Kritsanapuntu and Nilnaj Chaitanawisuti. Effects of extensive bottom cultivation of tropical oyster Crassostrea belcheri on benthic invertebrate community structure in Ban Don Bay, Suratthani Province, Southern Thailand. Aquaculture International, 26(2):433– 449, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-017-0227-x.

Ngo:2018:AHR

[1787] Thi Thu Hang Ngo, Hossein Azadi, Huu Cuong Tran, and Philippe Lebailly. Assessment of household risk management strategies for coastal aquaculture: the case of clam farming in Thaibinh Province, Vietnam. Aquaculture International, 26(2):451–468, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0226-y.

Suzuki:2018:BMP

[1788] Aya Suzuki and Vu Hoang Nam. Better management practices and their outcomes in shrimp farming: evidence from small-scale shrimp farmers in Southern Vietnam. Aquaculture International, 26(2):469–486, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-017-0228-9.

AndersonPereira:2018:CPG

[1789] Felipe Anderson Pereira, Natalia Ha, André Fernando Nascimento Gonçalves, Hélio Antunes, Wagner C. Valenti, and Thiago El Hadi Perez Fabregat. Can the polyculture with South American catfish improve the feeding efficiency of rainbow trout culture? Aquaculture International, 26(2):487–493, April 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0234-y.

Abreu:2018:EFV

[1790] Ruan E. F. Abreu, Thaís C. Magalhães, Renilde C. Souza, Samira T. L. Oliveira, Adriana M. G. Ibelli, Fábio N. Demarqui, João J. S. Gouveia, Mateus M. Costa, and Gisele V. Gouveia. Environmental factors on virulence of Aeromonas hydrophila. Aquaculture International, 26 (2):495–507, April 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-017-0230-2.

Khan:2018:EDV

[1791] Hajah Imran Khan, Jagabattuala Syama Dayal, Kondusamy Ambasankar, Eda Purdhvi Madhubabu, Rajabdeen Jannathulla, and Vanjiappan Rajaram. Enhancing the dietary value of palm oil in the presence of lysolecithin in tiger shrimp, Penaeus monodon. Aquaculture International, 26(2):509–522, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0235-x.

Haque:2018:ESD

[1792] M. Rezoanul Haque, M. Ashraful Islam, Zohura Khatun, Md Afzal Hossain, and Md Abdul Wahab. Effects of stocking densities of tilapia Oreochromis niloticus (Linnaeus, 1758) with the inclusion of silver carp Hypophthalmichthys molitrix (Valenciennes, 1844) in C/N-CP prawn Macrobrachium rosenbergii (De Man, 1879) culture pond. Aquaculture International, 26(2):523–541, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0229-8.

Forchino:2018:ATB

[1793] Andrea Alberto Forchino, Fabio Brambilla, Simona Rimoldi, Marco Saroglia, and Genciana Terova. The application of two benthic indices to investigate the effects of land-based fish farms in coastal transitional ecosystems: two case studies in Tuscany region (Italy). Aquaculture International, 26(2):543–555, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-017-0224-0.

Wu:2018:ECA

[1794] Dong-Lei Wu, Zhi-Quan Liu, You-Hui Huang, Wei-Wei Lv, Ming-Hai Chen, Yi-Ming Li, and Yun-Long Zhao. Effects of cold acclimation on the survival, feeding rate, and non-specific immune responses of the freshwater red claw crayfish (Cherax quadricarinatus). Aquaculture International, 26(2):557–567, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0236-4.

Vahdat:2018:LTE

[1795] Saeid Vahdat, Abolghasem Esmaeili Fereidouni, and Mohammad Kazem Khalesi. Long-term effects of vermicompost manure leachate (powder) inclusions on growth and survival, biochemical composition, total carotenoids, and broodstock reproductive performance of Artemia franciscana (Kellogg, 1906). Aquaculture International, 26(2):569–588, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-017-0232-0.

Megahed:2018:GNP

[1796] Mohamed E. Megahed, Gamal Elmesiry, Ahmed Ellithy, and Khaled Mohamed. Genetic, nutritional and pathological investigations on the effect of feeding low protein diet and biofloc on growth performance, survival and disease prevention of Indian white shrimp Fenneropenaeus indicus. Aquaculture International, 26(2):589–615, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-017-0231-1.

Praveenraj:2018:EIB

[1797] J. Praveenraj, P. Ezhil Praveena, T. Bhuvaneswari, A. Navaneeth Krishnan, and K. P. Jithendran. Experimental infection of Betanodavirus in freshwater fish Gambusia affinis (Baird and Girard, 1853) — a potential infection model for viral encephalopathy and retinopathy. Aquaculture International, 26(2):617–627, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0241-7.

Rebolledo:2018:DCS

[1798] Uriel Arreguin Rebolledo, S. Nandini, S. S. S. Sarma, José Cristóbal Román Reyes, and Gustavo Alejandro Rodríguez Montes de Oca. Demographic and competition studies on Brachionus ibericus and Proales similis in relation to salinity and algal (Nannochloropsis oculata) density. Aquaculture International, 26(2):629–644, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-017-0233-z.

Yang:2018:IGP

[1799] Zhigang Yang, Banghong Wei, Qibin Liu, Yongxu Cheng, and Junyu Zhou. Individual growth pattern of juvenile stages of the Chinese mitten crab (Eriocheir sinensis) reared under laboratory conditions. Aquaculture International, 26(2):645–657, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0239-1.

Ninwichian:2018:ETC

[1800] Parichart Ninwichian, Nirandon Phuwan, Kesara Jakpim, and Panya Sae-Lim. Effects of tank color on the growth, stress responses, and skin color of snakeskin gourami (Trichogaster pectoralis). Aquaculture International, 26(2):659–672, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0242-6.

Ajadi:2018:GEP

[1801] Abdullateef Ajadi, M. Y. Sabri, A. B. Dauda, M. Y. Ina-Salwany, and A. H. Hasliza. Growth enhancement and protective potential of feed-based outer membrane proteins against vibriosis in Macrobrachium rosenbergii. Aquaculture International, 26(2):673–684, April 2018. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0244-4.

McCrummen:2018:CEO

[1802] Stephen T. McCrummen, Yifen Wang, Terrill R. Hanson, Lisa Bott, and Shaoyang Liu. Culture environment and the odorous volatile compounds present in pond-raised channel catfish (Ictalurus punctatus). Aquaculture International, 26(2):685–694, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0247-1.

Linhartova:2018:PFA

[1803] Zuzana Linhartová, Jakub Krejsa, Tomáš Zajíc, Jan Másílko, Sabine Sampels, and Jan Mráz. Proximate and fatty acid composition of 13 important freshwater fish species in central Europe. Aquaculture International, 26(2):695–711, April 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0243-5.

Ballester:2018:IBP

[1804] Eduardo Luis Cupertino Ballester, Fabrício Martins Dutra, Tito Luís Pisseti, Ronaldo Olivera Cavalli, Paulo César Abreu, and Wilson Wasielesky. Influence of biofilm on the production of Farfantepenaeus paulensis in pens in the Patos Lagoon estuary. Aquaculture International, 26 (3):713–726, June 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0246-2.

Shahi:2018:IEH

[1805] Neetu Shahi, László Ardó, Gyöngyvér Fazekas, Elen Gócza, Shivendra Kumar, Norbert Rèvèsz, Zsuzsanna Jakabné Sándor, Zsuzsanna Molnár, Galina Jeney, and Zsigmond Jeney. Immunogene expression in head kidney and spleen of common carp (Cyprinus carpio L.) following thermal stress and challenge with Gram-negative bacterium, Aeromonas hydrophila. Aquaculture International, 26(3):727–741, June 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0250-6.

Lee:2018:EGI

[1806] Ji-Hoon Lee, Ju-Wan Kim, Yue-Jai Kang, Dae-Woong Ko, Jong-Man Kim, Sang-Hoon Choi, and Kwan Ha Park. Effects of β-1,3-glucan on innate immunity responses and mortality induced by Vibrio harveyi, hemorrhagic septicemia virus, or Miamiensis avidus in the olive flounder Paralichthys olivaceus. Aquaculture International, 26(3):743–756, June 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0248-0.

Yao:2018:CDY

[1807] Zhi Lei Yao, Yan Zhao, Hong Yu Ma, Hong Jun Liu, Hui Wang, and Xiang Shan Ji. Combined diet of yeast, fermented soybean meal, and microparticulate as larval feed in extensive rearing systems for seed production of the oriental river prawn Macrobrachium nipponense. Aquaculture International, 26(3):757–772, June 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0240-8.

Luo:2018:NRE

[1808] Guozhi Luo, Zefeng Liu, Jinfang Gao, Zhiwei Hou, and Hongxin Tan. Nitrate removal efficiency and bacterial community of polycaprolactonepacked bioreactors treating water from a recirculating aquaculture system. Aquaculture International, 26(3):773–784, June 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0251-5.

Sukenda:2018:ESA

[1809] Sukenda Sukenda, Rahman Rahman, Khairun Nisaa, Dendi Hidayatullah, and Apriana Vinasyiam. The efficacy of Streptococcus agalactiae vaccine preparations, administered to tilapia broodstock, in preventing streptococcosis in their offspring, via transfer of maternal immunity. Aquaculture International, 26(3):785–798, June 2018. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0252-4.

Huang:2018:FUP

[1810] Hui Huang, Luqing Pan, Shanshan Pan, and Mengsi Song. The feasibility of using primary shrimp hemocyte culture to screen herbal immunostimulants. Aquaculture International, 26(3):799–811, June 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0238-2.

Palm:2018:TCA

[1811] Harry W. Palm, Ulrich Knaus, Samuel Appelbaum, Simon Goddek, Sebastian M. Strauch, Tycho Vermeulen, M. Haïssam Jijakli, and Benz Kotzen. Towards commercial aquaponics: a review of systems, designs, scales and nomenclature. Aquaculture International, 26(3):813– 842, June 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-018-0249-z.

Abass:2018:DSY

[1812] David Attim Abass, Kwasi Adu Obirikorang, Benjamin Betey Campion, Regina Esi Edziyie, and Peter Vilhelm Skov. Dietary supplementation of yeast (Saccharomyces cerevisiae) improves growth, stress tolerance, and disease resistance in juvenile Nile tilapia (Oreochromis niloticus). Aquaculture International, 26(3):843–855, June 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0255-1.

Goosen:2018:EFA

[1813] Neill Jurgens Goosen, Lourens Francois De Wet, and Johann Ferdinand Görgens. Effects of formic acid in abalone diets that contain ingredients derived from fish processing by-products. Aquaculture International, 26 (3):857–868, June 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0257-z.

Mejias:2018:PRB

[1814] Cristian Mejias, Carlos Riquelme, Camila Sayes, Julián Plaza, and Fernando Silva-Aciares. Production of the rotifer Brachionus plicatilis (Müller 1786) in closed outdoor systems fed with the microalgae Nannochloropsis gaditana and supplemented with probiotic bacteria Pseudoalteromonas sp. (SLP1). Aquaculture International, 26(3):869–884, June 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-018-0253-3.

Zhang:2018:SAT

[1815] Defeng Zhang, Xiaoli Ke, Lihui Liu, Maixin Lu, Cunbin Shi, and Zhigang Liu. Streptococcus agalactiae from tilapia (Oreochromis sp.) transmitted to a new host, bighead carp (Aristichthys nobilis), in China. Aquaculture International, 26(3):885–897, June 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0254-2.

Lin:2018:ENE

[1816] Yan Lin, Ling-Hong Miao, Wu-Xiao Zhang, Wen-Jing Pan, Hua-Liang Liang, Xian-Ping Ge, Yan-Shun Xu, Bo Liu, Ming-Chun Ren, Qun-Lan Zhou, and Si-Lei Xia. Effect of nitrite exposure on oxygen-carrying capacity and gene expression of nf-κ b /hif-1α pathway in gill of bighead carp (Aristichthys nobilis). Aquaculture International, 26(3):899– 911, June 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-018-0256-0.

Castelo-Branco:2018:SVL

[1817] Thaís Castelo-Branco, Maria Madalena Pessoa Guerra, Roberta Soares, and Silvio Peixoto. Sperm vitrification of Litopenaeus vannamei: effect of cryoprotectant solutions on sperm viability and spawning quality after artificial insemination. Aquaculture International, 26(3):913– 920, June 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-018-0258-y.

Goncalves:2018:ABA

[1818] Rui A. Gonçalves, Tu Do Cam, Nguyen Nhu Tri, Gonçalo A. Santos, Pedro Encarnação, and Le Thanh Hung. Aflatoxin B_{1} (AFB_{1}) reduces growth performance, physiological response, and disease resistance in tra catfish (Pangasius hypophthalmus). Aquaculture International, 26 (3):921–936, June 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0259-x.

Le:2018:EGR

[1819] Bao Le, Jong-Am Shin, Man-Gu Kang, Sangmi Sun, Seung Hwan Yang, and Gyuhwa Chung. Enhanced growth rate and ulvan yield of Ulva pertusa using light-emitting diodes (LEDs). Aquaculture International, 26(4):937–946, August 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0260-4.

Wang:2018:ESS

[1820] Tao Wang, Qi Li, Jingxiao Zhang, and Ruihai Yu. Effects of salinity, stocking density, and algal density on growth and survival of Iwagaki oyster Crassostrea nippona larvae. Aquaculture International, 26 (4):947–958, August 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0261-3.

Girijakumari:2018:VVE

[1821] Nisha Rajagopalan Girijakumari, Kannapiran Ethiraja, and Prabhu Narayanasamy Marimuthu. In vitro and in vivo evaluation of probiotic properties of Enterobacter cloacae in kenyi cichlid, Maylandia lombardoi. Aquaculture International, 26(4):959–980, August 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0262-2.

Liu:2018:INC

[1822] Gang Liu, Zhangying Ye, Dezhao Liu, and Songming Zhu. Inorganic nitrogen control, growth, and immunophysiological response of Litopenaeus vannamei (Boone, 1931) in a biofloc system and in clear water with or without commercial probiotic. Aquaculture International, 26 (4):981–999, August 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0263-1.

Mendez-Martinez:2018:NEA

[1823] Yuniel Méndez-Martínez, Marcelo Ulises García-Guerrero, María Concepción Lora-Vilchis, Luis Rafael Martínez-Córdova, Fabiola Guadalupe Arcos-Ortega, Juan José Alpuche, and Edilmar Cortés-Jacinto. Nutritional effect of Artemia nauplii enriched with Tetraselmis suecica and Chaetoceros calcitrans microalgae on growth and survival on the river prawn Macrobrachium americanum larvae. Aquaculture International, 26 (4):1001–1015, August 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0264-0.

Song:2018:DIH

[1824] Zhidong Song, Peiyu Li, Jiying Wang, Yongzhi Sun, and Chengqiang Wang. Dietary inclusion of hydrolyzed soybean and cottonseed meals influence digestion, metabolic enzymes, and growth-related hormones and growth of juvenile turbot (Scophthalmus maximus). Aquaculture International, 26(4):1017–1033, August 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0265-z.

PachecoBoaventura:2018:IPL

[1825] Túlio Pacheco Boaventura, Kleber Campos Miranda-Filho, Rodrigo Lambert Oréfice, and Ronald Kennedy Luz. Influence of porosity of low-density polyethylene media on the maturation process of biofilters used in recirculating aquaculture systems. Aquaculture International, 26 (4):1035–1049, August 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0266-y.

Cyr:2018:PMR

[1826] Bernard-Antonin Dupont Cyr, Helge Tveiten, Domynick Maltais, Grant W. Vandenberg, and Nathalie R. Le François. Photoperiod manipulation for the reproductive management of captive wolffish populations: Anarhichas minor and A. lupus. Aquaculture International, 26 (4):1051–1065, August 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0267-x.

Jia:2018:EME

[1827] Xu-Ying Jia, Da-Sen Zhong, Dan Zhang, Fang Wang, and Wen-Li Zhou. Energy metabolic enzyme responses of Litopenaeus vannamei to thermal stress: a comparative study in freshwater and seawater conditions. Aquaculture International, 26(4):1067–1081, August 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0268-9.

Kuradomi:2018:PGS

[1828] Rafael Yutaka Kuradomi and Sergio Ricardo Batlouni. PGF_{2α} and gonadal steroid plasma levels of successful and unsuccessful spawning Piaractus mesopotamicus (Teleostei, Characiformes) females. Aquaculture International, 26(4):1083–1094, August 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0269-8.

Doroteo:2018:EPP

[1829] Alvin M. Doroteo, Fiona L. Pedroso, James David M. Lopez, and Mary Jane S. Apines-Amar. Evaluation of potential probiotics isolated from saline tilapia in shrimp aquaculture. Aquaculture International, 26(4):

1095–1107, August 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0270-2.

Santiago:2018:EDR

[1830] Cabaleiro Santiago, Lucía Barreiro, and Rubén Caamaño. Effect of diet and rearing conditions on growth and survival throughout developmental stages of larvae and juveniles of dusky grouper Epinephelus marginatus (Lowe 1834). Aquaculture International, 26(4):1109–1118, August 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0271-1.

delaPena:2018:MPP

[1831] Milagros R. de la Peña, Annie Villa Franco, Hermoso P. Igcasan, Mary Dianne Grace N. Arnaldo, Ramil M. Piloton, Soledad S. Garibay, and Vicente T. Balinas. Microalgal paste production of the diatom Chaetoceros calcitrans using electrolytic flocculation method at optimum culture conditions. Aquaculture International, 26(4):1119–1134, August 2018. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0272-0.

Muller-Belecke:2018:DPT

[1832] Andreas Müller-Belecke, Sebastian von Plessen, Gregor Schmidt, Carsten Kühn, and Andreas Spranger. Design and practical test of a compact phosphorus elimination module for freshwater RAS discharge water. Aquaculture International, 26(4):1135–1145, August 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0273-z.

Mabrok:2018:IME

[1833] Mahmoud Abd Elaziz Mabrok and Ali Wahdan. The immune modulatory effect of oregano (Origanum vulgare L.) essential oil on Tilapia zillii following intraperitoneal infection with Vibrio anguillarum. Aquaculture International, 26(4):1147–1160, August 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0274-y.

Sanz-Lazaro:2018:DMB

[1834] Carlos Sanz-Lazaro, Victoria Fernandez-Gonzalez, Pablo Arechavala-Lopez, David Izquierdo-Gomez, Elena Martinez-Garcia, and Pablo Sanchez-Jerez. Depth matters for bivalve culture in integrated multitrophic aquaculture (IMTA) and other polyculture strategies under non-eutrophic conditions. Aquaculture International, 26(5):1161–1170, October 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-018-0276-9.

Ballyaya:2018:DSR

[1835] Abhiman Purandara Ballyaya, Moumita Mondal, Shankar Mariappa Kalkuli, and Suresh Babu Padinhate Purayil. Development of a simple and rapid monoclonal antibody-based flow through immunogold assay (FIA) for detection of Aeromonas hydrophila. Aquaculture International, 26(5):1171–1186, October 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0278-7.

Castilho-Barros:2018:EEC

[1836] Leonardo Castilho-Barros, Fábio H. Almeida, Marcelo B. Henriques, and Walter Q. Seiffert. Economic evaluation of the commercial production between Brazilian samphire and whiteleg shrimp in an aquaponics system. Aquaculture International, 26(5):1187–1206, October 2018. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0277-8.

Jobling:2018:BRJ

[1837] Malcolm Jobling. Book review: J. Treasurer (editor): Cleaner fish biology and aquaculture applications. Aquaculture International, 26(5): 1207–1209, October 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0279-6.

Azpeitia:2018:ACP

[1838] K. Azpeitia, M. B. Urrutia, and D. Mendiola. Analysis of culture production and growth rate patterns of mussel (Mytilus galloprovincialis Lmk.) cultured in the open ocean of the SE Bay of Biscay for the commercial product development. Aquaculture International, 26(5): 1211–1228, October 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0275-x.

Bijoy:2018:FMR

[1839] Valiyaparambil Mohanan Bijoy, Sarasan Sabu, and Mahadevan Harikrishnan. Fish meal replacement with squilla (Oratosquilla nepa, Latreille) silage in a practical diet for the juvenile giant freshwater prawn, Macrobrachium rosenbergii de Man, 1879. Aquaculture International, 26(5): 1229–1245, October 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0280-0.

Wu:2018:ESD

[1840] Fan Wu, Hua Wen, Juan Tian, Ming Jiang, Wei Liu, Changgeng Yang, Lijuan Yu, and Xing Lu. Effect of stocking density on growth performance, serum biochemical parameters, and muscle texture properties of genetically improved farm tilapia, Oreochromis niloticus. Aquaculture International, 26(5):1247–1259, October 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0281-z.

He:2018:MIG

[1841] Yongjin He, Gang Lin, Xiaozhen Rao, Langjun Chen, Huang Jian, Mingzi Wang, Zheng Guo, and Bilian Chen. Microalga Isochrysis galbana in feed for Trachinotus ovatus: effect on growth performance and fatty acid composition of fish fillet and liver. Aquaculture International, 26(5): 1261–1280, October 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0282-y.

Rasdi:2018:IFT

[1842] Nadiah W. Rasdi and Jian G. Qin. Impact of food type on growth, survival and reproduction of the cyclopoid copepod Cyclopina kasignete as a potential live food in aquaculture. Aquaculture International, 26(5): 1281–1295, October 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0283-x.

Abbas:2018:IDM

[1843] Ahmed S. A. Abbas, Eman El-Wazzan, Amal R. Khafage, Abdel-Fattah M. El-Sayed, and Fatma A. Abdel Razek. Influence of different microalgal diets on gonadal development of the carpet shell clam Ruditapes decussatus broodstock. Aquaculture International, 26(5):1297– 1309, October 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-018-0284-9.

Goldstein:2018:BSD

[1844] Jason S. Goldstein and Jeffrey D. Shields. Bait-subsidized diets and their effects on ovigerous North American lobsters (Homarus americanus). Aquaculture International, 26(6):1311–1326, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0285-8.

Harlioglu:2018:IGA

[1845] Muzaffer Mustafa Harlıoğlu and Ardavan Farhadi. Importance of Gammarus in aquaculture. Aquaculture International, 26(6):1327–1338, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-018-0287-6.

Pascual:2018:WGF

[1846] Mariano M. Pascual, Juan P. Hualde, Virginia A. Bianchi, Pablo Moreno, Juan M. Castro, and Carlos M. Luquet. Wheat grains fermented by fungal mycelia (Pleurotus ostreatus or Lentinus edodes) as alternative feed ingredients for juvenile rainbow trout (Oncorhynchus mykiss). Aquaculture International, 26(6):1339–1352, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0286-7.

Schroeder:2018:ECW

[1847] Sarah Schroeder, John M. Grigor, Constantinos E. Stathopoulos, Anne Savage, Philip Cassidy, Stefan Toepfl, and Jonathan D. Wilkin. The effect of collagenase, water and calcium chloride on the removal of Salmo salar (salmon) and Oncorhynchus mykiss (trout) pin bones. Aquaculture International, 26(6):1353–1363, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0288-5.

Kohlmann:2018:VSS

[1848] K. Kohlmann, P. Kersten, J. Geßner, O. Eroglu, S. Firidin, M. Ciorpac, E. Taflan, and R. Suciu. Validation of 12 species-specific, tetrasomic microsatellite loci from the Russian sturgeon, Acipenser gueldenstaedtii, for genetic broodstock management. Aquaculture International, 26(6): 1365–1376, December 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0290-y.

Goddek:2018:CLS

[1849] Simon Goddek and Tycho Vermeulen. Comparison of Lactuca sativa growth performance in conventional and RAS-based hydroponic systems. Aquaculture International, 26(6):1377–1386, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0293-8.

Mensah:2018:ESE

[1850] Emmanuel T. D. Mensah, Hederick R. Dankwa, Lauridsen L. Torben, Ruby Asmah, Benjamin B. Campion, and Regina Edziyie. Effects of seasonal and environmental changes on aquaculture production in tropical Lake Volta, Ghana. Aquaculture International, 26(6): 1387–1400, December 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0294-7.

Bedekar:2018:EIG

[1851] Megha Kadam Bedekar, Praveena Soman, Sajal Kole, Deepika Anand, Gayatri Tripathi, M. Makesh, and K. V. Rajendran. Evaluation of interferon gamma (IFN-γ) of Labeo rohita as an immunomodulator: in vitro expression model. Aquaculture International, 26(6):1401–1413, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-018-0292-9.

Khosravi:2018:CBC

[1852] Mohammad Khosravi, Takavar Mohammadian, Mina Tahmasebifard, and Mehdi Pourmehdi Boroujeni. Correlation between c-reactive protein level, immunology, and hematology of a Oncorhynchus mykiss infected with Lactococcus garvieae. Aquaculture International, 26(6):1415– 1425, December 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0295-6.

Cubillo:2018:ESG

[1853] Alhambra Martínez Cubillo, João Gomes Ferreira, Christopher Michael Pearce, Robert Marshall, Dan Cheney, and Bobbi Hudson. Ecosystem services of geoduck farming in South Puget Sound, USA: a modeling analysis. Aquaculture International, 26(6):1427–1443, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0291-x.

Guo:2018:CAG

[1854] Liang Guo, Nan Zhang, Jing-Wen Yang, Hua-Yang Guo, Ke-Cheng Zhu, Bao-Suo Liu, Tian-Tian Liu, and Dian-Chang Zhang. Comprehensive assessment of the genetic diversity and population structure of cultured populations of golden pompano, Trachinotus ovatus (Linnaeus, 1758), by microsatellites. Aquaculture International, 26(6):1445– 1457, December 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0289-4.

Junior:2018:EMT

[1855] O. Tomazelli Júnior, F. Kuhn, P. J. Mendonça Padilha, C. Nunes Nesi, M. Mestres, J. Dal Magro, and S. De Lamo Castellví. Effect of microencapsulated thyme essential oil on white spot virus-infected Litopenaeus vannamei. Aquaculture International, 26(6):1459–1468, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0296-5.

Imsland:2018:AAS

[1856] Albert K. Imsland, Patrick Reynolds, Gerhard Eliassen, Thor A. Hangstad, Tor Anders Elvegård, Tonje Cecilie Urskog, and Bjørn Mikalsen. Assessment of artificial substrates for lumpfish: effect of material thickness and water current speed. Aquaculture International, 26(6): 1469–1479, December 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0298-3.

Wang:2018:EDE

[1857] Chang’an Wang, Hongbai Liu, Jinnan Li, Liansheng Wang, Zhigang Zhao, Liang Luo, and Qiyou Xu. Effects of dietary myo-inositol on growth, chemical composition and plasma chemistry of Amur sturgeon Acipenser schrenckii. Aquaculture International, 26(6):1481–1492, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-018-0299-2.

Srinuanpan:2018:PCO

[1858] Sirasit Srinuanpan, Benjamas Cheirsilp, Poonsuk Prasertsan, Yasuo Kato, and Yasuhisa Asano. Photoautotrophic cultivation of oleaginous microalgae and co-pelletization with filamentous fungi for cost-effective harvesting process and improved lipid yield. Aquaculture International, 26(6):1493–1509, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-018-0300-0.

Castilla-Gavilan:2018:OMP

[1859] Marta Castilla-Gavilán, Vincent Turpin, Florence Buzin, Bruno Cognie, and Priscilla Decottignies. Optimizing metamorphosis in Paracentrotus lividus aquaculture using alternative macroalgae species to Corallina sp. Aquaculture International, 26(6):1511–1518, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0305-8.

Gao:2018:ELF

[1860] Yang Gao, Jingting Zhu, Huajiang Bao, Vector Hector, Bo Zhao, and Zhangjie Chu. Effect of lignite fulvic acid on growth, antioxidant ability, and HSP70 of Pacific white shrimp, Litopenaeus vannamei. Aquaculture International, 26(6):1519–1530, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0302-y.

Huyben:2018:EMF

[1861] David Huyben, David Bevan, Roselynn Stevenson, Hongde Zhou, and Richard Moccia. Evaluation of membrane filtration and UV irradiation to control bacterial loads in recirculation aquaculture systems. Aquaculture International, 26(6):1531–1540, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0301-z.

Gichana:2018:WMR

[1862] Zipporah Moraa Gichana, David Liti, Herwig Waidbacher, Werner Zollitsch, Silke Drexler, and Joseph Waikibia. Waste management in recirculating aquaculture system through bacteria dissimilation and plant assimilation. Aquaculture International, 26(6):1541–1572, December 2018. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0303-x.

Potki:2018:GHB

[1863] Najmeh Potki, Bahram Falahatkar, and Alireza Alizadeh. Growth, hematological and biochemical indices of common carp Cyprinus carpio fed diets containing corn gluten meal. Aquaculture International, 26(6): 1573–1586, December 2018. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0304-9.

Lawrence:2019:LSP

[1864] John M. Lawrence, Chong Zhao, and Ya-Qing Chang. Large-scale production of sea urchin (Strongylocentrotus intermedius) seed in a hatchery in China. Aquaculture International, 27(1):1–7, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0319-2.

Flaherty:2019:PAT

[1865] Mark Flaherty, Gregor Reid, Thierry Chopin, and Erin Latham. Public attitudes towards marine aquaculture in Canada: insights from the Pacific and Atlantic coasts. Aquaculture International, 27(1):9– 32, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-018-0312-9.

Selim:2019:VEP

[1866] Khaled M. Selim, Hassanin M. El-Sayed, M. A. El-Hady, and Rasha M. Reda. In vitro evaluation of the probiotic candidates isolated from the gut of Clarias gariepinus with special reference to the in vivo assessment of live and heat-inactivated Leuconostoc mesenteroides and Edwardsiella sp. Aquaculture International, 27(1):33–51, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0297-4.

Villamil:2019:EDA

[1867] L. Villamil, S. Infante Villamil, G. Rozo, and J. Rojas. Effect of dietary administration of kappa carrageenan extracted from Hypnea musciformis on innate immune response, growth, and survival of Nile tilapia (Oreochromis niloticus). Aquaculture International, 27(1):53– 62, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-018-0306-7.

Tarrab:2019:IGE

[1868] Koby Tarrab, Shay Ravid-Peretz, and Michal Ucko. Immunoserology of European seabass (Dicentrarchus labrax) and white grouper (Epinephelus aeneus) as a non-lethal diagnostic tool for viral nervous necrosis. Aquaculture International, 27(1):63–77, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0307-6.

Cieplinski:2019:EDE

[1869] Mateusz Ciepliński, Mariusz Kasprzak, Monika Grandtke, Aleksandra Steliga, Piotr Kamiński, and Leszek Jerzak. The effect of dipotassium EDTA and lithium heparin on hematologic values of farmed brown trout Salmo trutta (L.) spawners. Aquaculture International, 27(1): 79–87, February 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-0308-5.

Bergstrom:2019:BME

[1870] Per Bergström, Niklas Hällmark, Karl-Johan Larsson, and Mats Lindegarth. Biodeposits from Mytilus edulis: a potentially high-quality food source for the polychaete, Hediste diversicolor. Aquaculture International, 27(1):89–104, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0309-4.

You:2019:EST

[1871] Weiwei You, Bo Wang, Xuan Luo, and Caihuan Ke. Environmental stress tolerance and immune response for the small abalone hybrids. Aquaculture International, 27(1):105–123, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0310-y.

Mohd-Aris:2019:VHP

[1872] Aslizah Mohd-Aris, Mohd-Zamri Saad, Hassan Mohd Daud, Mohd Termizi Yusof, and Md Yasin Ina-Salwany. Vibrio harveyi protease deletion mutant as a live attenuated vaccine candidate against vibriosis and transcriptome profiling following vaccination for Epinephelus fuscoguttatus. Aquaculture International, 27(1):125–140, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0311-x.

Sahraei:2019:EBS

[1873] Fereshte Sahraei, Hamed Ahari, and Shapoor Kakoolaki. Effect of Bacillus subtilis as a probiotic on protein, lipid content, and trypsin and chymotrypsin enzymes in rainbow trout biometry (Oncorhynchus mykiss). Aquaculture International, 27(1):141–153, February 2019. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0313-8.

Falk-Petersen:2019:EDI

[1874] Inger-Britt Falk-Petersen, Stefano Peruzzi, Hege Lysne, Hilde Toften, Helge Tveiten, Bjørn Steinar Sæther, and Velmurugu Puvanendran. Effects of different incubation and start-feeding temperature regimes on growth, survival, and histomorphology of cod larvae. Aquaculture International, 27(1):155–166, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0314-7.

Wu:2019:FSE

[1875] Yu Wu, Horst Kaiser, and Clifford L. W. Jones. A first study on the effect of dietary soya levels and crystalline isoflavones on growth, gonad development and gonad histology of farmed abalone, Haliotis midae. Aquaculture International, 27(1):167–193, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0315-6.

Unuma:2019:MIM

[1876] Tatsuya Unuma, Takashi Ichikawa, Sayumi Sawaguchi, and Natsuki Hasegawa. A microplate incubation method for assessing egg quality of the barfin flounder: effects of well size and rearing medium on larval viability. Aquaculture International, 27(1):195–208, February 2019. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0316-5.

Soto-Alcala:2019:TRI

[1877] Jorge Soto-Alcalá, Píndaro Álvarez-Ruiz, J. M. Audelo-Naranjo, H. M. Esparza-Leal, I. E. Luis-Villaseñor, J. A. Estrada-Godínez, A. Luna-González, C. Gámez-Jiménez, and G. Diarte-Plata. Transcriptional response of immune-related genes in Litopenaeus vannamei post-larvae cultured in recirculating aquaculture systems with and without biofloc. Aquaculture International, 27(1):209–225, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0317-4.

Hindu:2019:RIS

[1878] S. Vidhya Hindu, Natarajan Chandrasekaran, Amitava Mukherjee, and John Thomas. A review on the impact of seaweed polysaccharide on the growth of probiotic bacteria and its application in aquaculture. Aquaculture International, 27(1):227–238, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0318-3.

Mohammadian:2019:EAP

[1879] Takavar Mohammadian, Mojtaba Alishahi, Mohammad Reza Tabandeh, Abdolhossein Jangaran Nejad, Esmaeil Karami, and Mojtaba Zarea. Effects of autochthonous probiotics, isolated from Tor grypus (Karaman, 1971) intestine and Lactobacillus casei (PTCC 1608) on expression of immune-related genes. Aquaculture International, 27(1):239– 260, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-018-0320-9.

Ray:2019:ECF

[1880] Andrew J. Ray, John W. Leffler, and Craig L. Browdy. The effects of a conventional feed versus a fish-free feed and biofloc management on the nutritional and human sensory characteristics of shrimp (Litopenaeus vannamei). Aquaculture International, 27(1):261–277, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0321-8.

Sharma:2019:BAM

[1881] S. R. Krupesha Sharma, M. A. Pradeep, P. N. Dube, T. V. Arun Kumar, Raj Kumar, and T. Raja Swaminathan. Betanodavirus-associated mortality in Asian seabass (Lates calcarifer, Bloch) cultured in indoor tanks and sea cages. Aquaculture International, 27(1):279–286, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0322-7.

Kubiriza:2019:GHA

[1882] Godfrey Kawooya Kubiriza, Jón Árnarson, Ólafur Sigurgeirsson, Patricia Hamaguchi, Sigurdur Snorrason, Tumi Tómasson, and Helgi Thorarensen. Growth and hepatic antioxidant enzyme activity of juvenile Arctic charr (Salvelinus alpinus) fed on diets supplemented with ethoxyquin, rosemary (Rosmarinus officinalis), or bladder wrack (Fucus vesiculosus). Aquaculture International, 27(1):287–301, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0324-5.

Zhang:2019:BCB

[1883] Xiaoping Zhang, Yang Wang, Zheke Zhong, Qingjun Shao, Yibing Wang, and Weifen Li. Bacterial complexes of Bacillus subtilis and Pseudomonas stutzeri alter the microbial composition in grass carp water. Aquaculture International, 27(1):303–312, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0325-4.

PerezRibeiro:2019:EBA

[1884] Paula Adriane Perez Ribeiro, Daniela Chemim de Melo Hoyos, Camila Gomes de Oliveira, Marco Aurélio Lopes Della Flora, and Ronald Kennedy Luz. Eugenol and benzocaine as anesthetics for Lophiosilurus alexandri juvenile, a freshwater carnivorous catfish. Aquaculture International, 27(1):313–321, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0326-3.

Marquez:2019:MDJ

[1885] A. Marquez, C. Lodeiros, A. Loor, J. Revilla, F. Da Costa, and S. Sonnenholzner. Microalgae diet for juveniles of Spondylus limbatu s. Aquaculture International, 27(1):323–335, February 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0327-2.

Nunes:2019:MFE

[1886] Alberto J. P. Nunes, Hassan Sabry-Neto, Francisco Hélio Pires da Silva, Adhemar Rodrigues de Oliveira-Neto, and Karthik Masagounder. Multiple feedings enhance the growth performance and feed efficiency of juvenile Litopenaeus vannamei when fed a low-fish meal amino acidsupplemented diet. Aquaculture International, 27(2):337–347, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0330-7.

Cantelli:2019:DSS

[1887] Liege Cantelli, Priscila Goncalves, Cristhiane Guertler, Mirian Kayser, Mariana Rangel Pilotto, Margherita Anna Barracco, and Luciane Maria Perazzolo. Dietary supplementation with sulfated polysaccharides from Gracilaria birdiae promotes a delayed immunostimulation in marine shrimp challenged by the white spot syndrome virus. Aquaculture International, 27(2):349–367, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0328-1.

Abdel-Tawwab:2019:DAB

[1888] Mohsen Abdel-Tawwab, Abdel-Rahman A. Khattaby, and Mohamed N. Monier. Dietary acidifiers blend enhanced the production of Nile tilapia (Oreochromis niloticus), striped mullet (Mugil cephalus), and African catfish (Clarias gariepinus) polycultured in earthen ponds. Aquaculture International, 27(2):369–379, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0329-0.

Gallardo-Colli:2019:MCN

[1889] Alfredo Gallardo-Collí, Carlos Iván Pérez-Rostro, Martha Patricia Hernández-Vergara, and Ignacio Alejandro Pérez-Legaspi. Microeukaryote community and the nutritional composition of the biofloc during Nile tilapia culture in water-reusing biofloc systems. Aquaculture International, 27(2):381–398, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0335-2.

Valladao:2019:EDT

[1890] Gustavo Moraes Ramos Valladão, Sílvia Umeda Gallani, Suzana Kotzent, Inácio Mateus Assane, and Fabiana Pilarski. Effects of dietary thyme essential oil on hemato-immunological indices, intestinal morphology, and microbiota of Nile tilapia. Aquaculture International, 27(2):399– 411, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-018-0332-5.

Francova:2019:EFF

[1891] Kateřina Francová, Kateřina Šumberová, Georg A. Janauer, and Zdeněk Adámek. Effects of fish farming on macrophytes in temperate carp ponds. Aquaculture International, 27(2):413–436, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-018-0331-6.

Sanchez:2019:FUN

[1892] Fabiola Chong Sánchez, Martha Enriquez Díaz, Eric Murillo Rodríguez, and Dalila Aldana Aranda. First use of a non-invasive technique for determination of sex hormones in the queen conch Lobatus gigas, Mollusca Gastropoda. Aquaculture International, 27(2):437–448, April 2019. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-018-0336-1.

Obirikorang:2019:EWF

[1893] Kwasi Adu Obirikorang, Nelson Winston Agbo, Christian Obirikorang, Daniel Adjei-Boateng, Sefakor Esinam Ahiave, and Peter Vilhelm Skov. Effects of water flow rates on growth and welfare of Nile tilapia (Oreochromis niloticus) reared in a recirculating aquaculture system. Aquaculture International, 27(2):449–462, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00342-0.

Dias:2019:EFP

[1894] Joel Artur Rodrigues Dias, Higo Andrade Abe, Natalino Costa Sousa, Raimundo Darley Figueredo Silva, Carlos Alberto Martins Cordeiro, Grazielle Fernanda Evangelista Gomes, Jonathan Stuart Ready, José Luiz Pedreira Mouriño, Maurício Laterça Martins, Paulo César Falanghe Carneiro, Alexandre Nízio Maria, and Rodrigo Yudi Fujimoto. Enterococcus faecium as potential probiotic for ornamental neotropical cichlid fish, Pterophyllum scalare (Schultze, 1823). Aquaculture International, 27(2):463–474, April 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00339-9.

Kononova:2019:IMS

[1895] Svetlana V. Kononova, Dmitry V. Zinchenko, Tatyana A. Muranova, Nataliya A. Belova, and Anatoly I. Miroshnikov. Intestinal microbiota of salmonids and its changes upon introduction of soy proteins to fish feed. Aquaculture International, 27(2):475–496, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00341-1.

Haldar:2019:SNP

[1896] Chandan Haldar, S. P. Das, Bindu R. Pillai, Annam Pavan-Kumar, P. Gireesh-Babu, P. Das, and Aparna Chaudhari. Single-nucleotide polymorphisms linked to body weight revealed in growth selected Macrobrachium rosenbergii. Aquaculture International, 27(2):497–508, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-018-0334-3.

Katayama:2019:ILR

[1897] Tomoyo Katayama, Masatoshi Kishi, Kazutaka Takahashi, Ken Furuya, Mohd Effendy Abd Wahid, Helena Khatoon, and Nor Azman Kasan. Isolation of lipid-rich marine microalgae by flow cytometric screening with Nile red staining. Aquaculture International, 27(2):509– 518, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00344-y.

Rahman:2019:WGE

[1898] Sanzidur Rahman and Basanta Kumar Barmon. Women’s gainful employment in ‘gher’ farming system (prawn–carp–rice integrated culture) in Bangladesh: trends and determinants. Aquaculture International, 27 (2):519–537, April 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00347-9.

Cozer:2019:CPI

[1899] Nathieli Cozer, Aline Horodesky, Vitor Gomes Rossi, Giorgi Dal Pont, and Antonio Ostrensky. Challenges and potentialities of the integrated production regime implementation in the Brazilian marine shrimp farming: a systematic review. Aquaculture International, 27(2):539– 553, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00348-8.

Su:2019:ERS

[1900] Jianmei Su, Hongli Hou, Chunhong Wang, and Yuliang Luo. Effects of replacing soybean meal with cottonseed meal on growth, muscle amino acids, and hematology of juvenile common carp, Cyprinus carpio. Aquaculture International, 27(2):555–566, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00340-2.

Hamsah:2019:IRR

[1901] Hamsah Hamsah, Widanarni Widanarni, Alimuddin Alimuddin, Munti Yuhana, Muhammad Zairin Junior, and Dendi Hidayatullah. Immune response and resistance of Pacific white shrimp larvae administered probiotic, prebiotic, and synbiotic through the bio-encapsulation of Artemia sp. Aquaculture International, 27(2):567–580, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00346-w.

Ortiz-Estrada:2019:DBS

[1902] Ángel Martín Ortiz-Estrada, Teresa Gollas-Galván, Luis Rafael Martínez-Córdova, Armando Burgos-Hernández, Susana María Scheuren-Acevedo, Mauricio Emerenciano, and Marcel Martínez-Porchas. Diversity and bacterial succession of a phototrophic biofilm used as complementary food for shrimp raised in a super-intensive culture. Aquaculture International, 27(2):581–596, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00345-x.

Hallajian:2019:HEE

[1903] Ali Hallajian, Hossien Ali Abdolhay, Abdol Ahad Shadparvar, Mahtab Yarmohammadi, and Mohammad Ali Yazdani Sadati. Heritability estimation of eggs and the newly hatched larvae of wild Persian sturgeon (Acipenser persicus). Aquaculture International, 27(2):597–608, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00338-w.

Behera:2019:PMP

[1904] B. K. Behera, A. Das, P. Paria, A. K. Sahoo, P. K. Parida, T. Abdulla, and B. K. Das. Prevalence of microsporidian parasite, Enterocytozoon hepatopenaei in cultured Pacific white shrimp, Litopenaeus vannamei (Boone, 1931) in West Bengal, East Coast of India. Aquaculture International, 27(2):609–620, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00350-0.

Jobling:2019:BRJa

[1905] Malcolm Jobling. Book review: J. T. Trushenski: Understanding aquaculture. Aquaculture International, 27(2):621–623, April 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00349-7.

Chary:2019:MSC

[1906] Killian Chary, Annie Fiandrino, Denis Covès, Joël Aubin, Jean-Claude Falguière, and Myriam D. Callier. Modeling sea cage outputs for datascarce areas: application to red drum (Sciaenops ocellatus) aquaculture in Mayotte, Indian Ocean. Aquaculture International, 27(3):625– 646, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00351-z.

Akbarzadeh:2019:UDS

[1907] Arash Akbarzadeh, Somayeh Pakravan, Kimiya Karimi, Kobra Babanejad Abkenar, Masoumeh Eshagh Nimvari, Mohammad Niroomand, Seyed Alireza Sobhani, and Eisa Ebrahimi Dorcheh. Utilization of date seed meal in the diet of Pacific white shrimp (Penaeus vannamei): growth performance, body and fatty acid composition, biochemical parameters, and tolerance of salinity stress. Aquaculture International, 27 (3):647–661, June 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00352-y.

Liu:2019:IHG

[1908] Feng Liu, Yangyang Liu, Tianqi Chu, Bao Lou, Wei Zhan, and Ruiyi Chen. Interspecific hybridization and genetic characterization of Larimichthys polyactis ([female sign]) and L. crocea ([male sign]). Aquaculture International, 27(3):663–674, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00353-x.

Shen:2019:PSM

[1909] Hui Shen, Yi Qiao, Xihe Wan, Ge Jiang, Xianping Fan, Hui Li, Wenjun Shi, Libao Wang, and Xiaoran Zhen. Prevalence of shrimp microsporidian parasite Enterocytozoon hepatopenaei in Jiangsu Province, China. Aquaculture International, 27(3):675–683, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00358-6.

Tancredo:2019:EAD

[1910] Karen Roberta Tancredo, José Victor Ferrarezi, Natália da Costa Marchiori, and Maurício Laterça Martins. Ecotoxicological assays to determine the median lethal concentration (LC_{50}) of formalin for fish. Aquaculture International, 27(3):685–694, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00354-w.

Pounds:2019:OER

[1911] Alexandra Pounds and John Bostock. Open educational resources (OER) in higher education courses in aquaculture and fisheries: opportunities, barriers, and future perspectives. Aquaculture International, 27 (3):695–710, June 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00355-9.

Gallani:2019:DIR

[1912] S. U. Gallani, G. M. R. Valladão, S. Kotzent, E. F. E. Santo, E. H. G. Ponsano, and F. Pilarski. Dietary intake of Rubrivivax gelatinosus biomass enhances phagocytic cells in tropical fish Piaractus mesopotamicus infected with Aeromonas hydrophila. Aquaculture International, 27 (3):711–720, June 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00359-5.

Chen:2019:CNA

[1913] Binbin Chen, Lidong Lin, Zengling Ma, Tiantian Zhang, Weizhou Chen, and Dinghui Zou. Carbon and nitrogen accumulation and interspecific competition in two algae species, Pyropia haitanensis and Ulva lactuca, under ocean acidification conditions. Aquaculture International, 27(3):721–733, June 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00360-y.

Li:2019:AEH

[1914] Mu-Yang Li, Xin-Yu Liu, Chang-Ge Xia, Gui-Qin Wang, and Dong-Ming Zhang. Astaxanthin enhances hematology, antioxidant and immunological parameters, immune-related gene expression, and disease resistance against in Channa argus. Aquaculture International, 27(3): 735–746, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00362-w.

Alaliyat:2019:OFD

[1915] Saleh Alaliyat, Harald Yndestad, and Pål I. Davidsen. Optimal fish densities and farm locations in Norwegian fjords: a framework to use a PSO algorithm to optimize an agent-based model to simulate fish disease dynamics. Aquaculture International, 27(3):747–770, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00366-6.

Kucukgul:2019:IAA

[1916] Azime Küçükgül, Altuğ Küçükgül, Ramazan Gönenci, Şule Yurdagül Özsoy, Banu Kutlu, and Mehmet Mustafa İşgör. Investigation of the anti-apoptotic activity of ozone therapy in rainbow trout macrophages infected with Yersinia ruckeri. Aquaculture International, 27(3):771– 783, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00364-8.

Aheto:2019:SSF

[1917] Denis Worlanyo Aheto, Esther Acheampong, and Justice Odoiquaye Odoi. Are small-scale freshwater aquaculture farms in coastal areas of Ghana economically profitable? Aquaculture International, 27(3): 785–805, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00363-9.

Peng:2019:ELI

[1918] Ruibing Peng, Xia-Min Jiang, Maowang Jiang, and Sihan Chen. Effect of light intensity on embryonic development of the cuttlefish Sepia lycidas. Aquaculture International, 27(3):807–816, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00367-5.

Kahieshesfandiari:2019:SOS

[1919] M. Kahieshesfandiari, M. Y. Sabri, M. Y. Ina-salwany, M. D. Hassan, O. Noraini, A. A. Ajadi, and A. I. Isiaku. Streptococcosis in Oreochromis sp.: is feed-based biofilm vaccine of Streptococcus agalactiae effective? Aquaculture International, 27(3):817–832, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00372-8.

Linhartova:2019:IRE

[1920] Zuzana Linhartová, Roman Lunda, Petr Dvořák, Jan Bárta, Veronika Bártová, Jaromír Kadlec, Eva Samková, Jan Bedrníček, Milan Pešek, Ivana Laknerová, Sonja Smole Možina, Pavel Smetana, and Jan Mráz. Influence of rosemary extract (Rosmarinus officinalis) inolens to extend the shelf life of vacuum-packed rainbow trout (Oncorhynchus mykiss) fillets stored under refrigerated conditions. Aquaculture International, 27(3):833–847, June 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00369-3.

Estrada-SanAgustin:2019:PTD

[1921] E. Estrada-San Agustín, L. Gutiérrez, M. Bernad, H. Castillo-Juárez, S. Sánchez, and H. Sumano. Pharmacokinetics of two dosing forms of a recrystallized enrofloxacin as hydrochloride dihydrate in tilapia (Oreochromis niloticus × Oreochromis mossambicus). Aquaculture International, 27(3):849–857, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-019-00371-9.

Zhang:2019:CSI

[1922] Jiajia Zhang, Xueying Pei, Zhaowen Wu, Jie Li, Tao Wang, and Shaowu Yin. A comparative study of immune response between hybrid yellow catfish “Huangyou-1” and its parental populations after challenge with Aeromonas hydrophila or Edwardsiella ictaluri. Aquaculture International, 27(3):859–873, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-019-00370-w.

Eryalcin:2019:NVP

[1923] Kamil Mert Eryalçın. Nutritional value and production performance of the rotifer Brachionus plicatilis Müller, 1786 cultured with different feeds at commercial scale. Aquaculture International, 27(3):875– 890, June 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00375-5.

Saekhow:2019:BAB

[1924] Suktianchai Saekhow, Karun Thongprajukaew, and Wutiporn Phromkunthong. Blue aquarium background is appropriate for rearing male Siamese fighting fish (Betta splendens). Aquaculture International, 27 (3):891–903, June 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00374-6.

Li:2019:PEC

[1925] Tao Li, Fei Ke, Jian-Fang Gui, Li Zhou, Xiao-Juan Zhang, and Qi-Ya Zhang. Protective effect of Clostridium butyricum against Carassius auratus herpesvirus in gibel carp. Aquaculture International, 27 (3):905–914, June 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00377-3.

Zarski:2019:P

[1926] Daniel Żarski, Damien Toner, Tomáš Policar, Stefan Meyer, and Stefan Teerlinck. Preface. Aquaculture International, 27(4):915–916, August 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00424-z.

Ljubobratovic:2019:AET

[1927] Uroš Ljubobratović, Géza Péter, Rene Alvestad, Zoltán Horváth, and András Rónyai. Alcalase enzyme treatment affects egg incubation and larval quality in pikeperch (Sander lucioperca). Aquaculture International, 27(4):917–929, August 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-018-0323-6.

Palinska-Zarska:2019:EWH

[1928] Katarzyna Palińska-Żarska, Sławomir Krejszeff, Michał Łopata, and Daniel Żarski. Effect of water hardness, temperature, and tank wall color, on the effectiveness of swim bladder inflation and survival of Eurasian perch (Perca fluviatilis, L.) larvae reared under controlled conditions. Aquaculture International, 27(4):931–943, August 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-018-0333-4.

Krol:2019:ESS

[1929] Jarosław Król, Artur Długoński, Maciej Błażejewski, and Piotr Hliwa. Effect of size sorting on growth, cannibalism, and survival in Eurasian perch Perca fluviatilis L.. post-larvae. Aquaculture International, 27(4): 945–955, August 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-018-00337-3.

Schaefer:2019:FIB

[1930] Fabian J. Schaefer, Moritz Tielmann, Julia L. Overton, Angela Krüger, Sven Wuertz, Werner Kloas, Carsten Schulz, and Stefan Meyer. Fate or independency: is batch-specific larval performance determined by egg traits? A case study in farmed pikeperch (Sander lucioperca). Aquaculture International, 27(4):957–969, August 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00356-8.

Zakes:2019:EPT

[1931] Zdzisław Zakeś, Maciej Rożyński, and Krystyna Demska-Zakeś. Effect of PIT tagging on hematology and plasma composition of juvenile pikeperch (Sander lucioperca (L.)). Aquaculture International, 27 (4):971–981, August 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00357-7.

Blecha:2019:SBI

[1932] Miroslav Blecha, Oleksandr Malinovskyi, Lukáš Veselý, Jiří Křišťan, and Tomáš Policar. Swim bladder inflation failure in pikeperch (Sander lucioperca) larvae in pond culture. Aquaculture International, 27(4): 983–989, August 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00361-x.

Sipos:2019:CGA

[1933] Dóra Kánainé Sipos, Gyula Kovács, Eszter Buza, Katalin Csenki-Bakos, Ágnes Ősz, Uroš Ljubobratović, Réka Cserveni-Szücs, Miklós Bercsényi, István Lehoczky, Béla Urbányi, and Balázs Kovács. Comparative genetic analysis of natural and farmed populations of pike-perch (Sander lucioperca). Aquaculture International, 27(4):991–1007, August 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00365-7.

Ljubobratovic:2019:EHP

[1934] Uroš Ljubobratović, Géza Péter, Zsuzsana Sándor, Nándor Kugyela, and András Rónyai. The effect of hormonal preparation (gonadotropins vs. gonadoliberins) on pre-seasonally obtained eggs and larvae quality in pikeperch (Sander lucioperca L.). Aquaculture International, 27(4): 1009–1024, August 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00368-4.

Steinberg:2019:IGR

[1935] Kathrin Steinberg, Jan Zimmermann, Stefan Meyer, and Carsten Schulz. Individual growth rates of pikeperch (Sander lucioperca) depending on water exchange rates in recirculating aquaculture systems. Aquaculture International, 27(4):1025–1035, August 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00373-7.

Samarin:2019:VSP

[1936] Azin Mohagheghi Samarin, Azadeh Mohagheghi Samarin, Miroslav Blecha, Jiri Kristan, and Tomas Policar. In vitro storage of pikeperch (Sander lucioperca) eggs. Aquaculture International, 27(4):1037–1044, August 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00380-8.

Rupp:2019:IIP

[1937] M. Rupp, R. Knüsel, P-D. Sindilariu, and H. Schmidt-Posthaus. Identification of important pathogens in European perch (Perca fluviatilis) culture in recirculating aquaculture systems. Aquaculture International, 27 (4):1045–1053, August 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00382-6.

Vanina:2019:SIP

[1938] Tatyana Vanina, Radek Gebauer, Lola Toomey, Vlastimil Stejskal, Bořek Drozd, Martin Bláha, Jan Kouřil, and Thomas Lecocq. Seeking for the inner potential: comparison of larval growth rate between seven populations of Perca fluviatilis. Aquaculture International, 27(4):1055– 1064, August 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00384-4.

Policar:2019:PSB

[1939] Tomas Policar, Oleksandr Malinovskyi, Jiri Kristan, Vlastimil Stejskal, and Azin Mohagheghi Samarin. Post-spawning bath treatments to reduce morbidity and mortality of pond-cultured pikeperch (Sander lucioperca L.) broodstock. Aquaculture International, 27(4):1065–1078, August 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00392-4.

Ljubobratovic:2019:GHI

[1940] Uroš Ljubobratović, Božidar Rašković, Endre Balogh, Svetlana Lengyel, Gyula Kovàcs, Endre Janurik, Vesna Poleksić, and András Rónyai. Gill histopathological indicators in pikeperch Sander lucioperca larvae reared in a flow-through system: effect of clay-turbid water. Aquaculture International, 27(4):1079–1091, August 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00393-3.

Malinovskyi:2019:BPS

[1941] Oleksandr Malinovskyi, Jitka Kolářová, Miroslav Blecha, Alžběta Stará, Josef Velíšek, Jiří Křišťan, and Tomáš Policar. Behavior and physiological status of pond-cultured pikeperch (Sander lucioperca) broodstock effected by sexual interactions throughout semi-artificial reproduction. Aquaculture International, 27(4):1093–1107, August 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00401-6. See

correction [2057].

Zimmermann:2019:MRU

[1942] J. Zimmermann, K. Steinberg, S. Meyer, and C. Schulz. Maintenance requirement and utilisation efficiency of energy and protein for ongrowing pikeperch (Sander lucioperca (L.)). Aquaculture International, 27(4): 1109–1123, August 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00404-3. See correction [2456].

Pirhonen:2019:SOV

[1943] Juhani Pirhonen, Liisa Muuri, Saara M. Kalliokoski, Marko M. Puranen, and Timo J. Marjomäki. Seasonal and ontogenetic variability in stomach size of Eurasian perch (Perca fluviatilis L.). Aquaculture International, 27 (4):1125–1135, August 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00415-0.

Imentai:2019:OAR

[1944] Aiman Imentai, Carlos Yanes-Roca, Christoph Steinbach, and Tomáš Policar. Optimized application of rotifers Brachionus plicatilis for rearing pikeperch Sander lucioperca L. larvae. Aquaculture International, 27(4): 1137–1149, August 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00420-3.

Policar:2019:RPG

[1945] Tomas Policar, Fabian J. Schaefer, Edson Panana, Stefan Meyer, Stefan Teerlinck, Damien Toner, and Daniel Żarski. Recent progress in European percid fish culture production technology — tackling bottlenecks. Aquaculture International, 27(5):1151–1174, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00433-y. See

correction [1946].

Policar:2019:CRP

[1946] Tomas Policar, Fabian J. Schaefer, Edson Panana, Stefan Meyer, Stefan Teerlinck, Damien Toner, and Daniel Żarski. Correction to: Recent progress in European percid fish culture production technology — tackling bottlenecks. Aquaculture International, 27(5):1175, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00457-4. See [1945].

Toomey:2019:WBG

[1947] Lola Toomey, Martin Bláha, Emilie Mauduit, Tatyana Vanina, Margot Baratçabal, Yannick Ledoré, Sami Vesala, Pascal Fontaine, Alain Pasquet, and Thomas Lecocq. When behavioural geographic differentiation matters: inter-populational comparison of aggressiveness and group structure in the European perch. Aquaculture International, 27(5): 1177–1191, October 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00343-z. See correction [1947].

Toomey:2019:CWB

[1948] Lola Toomey, Martin Bláha, Emilie Mauduit, Tatyana Vanina, Margot Baratçabal, Yannick Ledoré, Sami Vesala, Pascal Fontaine, Alain Pasquet, and Thomas Lecocq. Correction to: When behavioural geographic differentiation matters: inter-populational comparison of aggressiveness and group structure in the European perch. Aquaculture International, 27(5):1193, October 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00458-3. See [1947].

Wan:2019:GWA

[1949] Liang Wan, Weijia Wang, Guijia Liu, Linsong Dong, Wanbo Li, Zhaofang Han, Kun Ye, and Zhiyong Wang. A genome-wide association study of resistance to Pseudomonas plecoglossicida infection in the large yellow croaker (Larimichthys crocea). Aquaculture International, 27(5): 1195–1208, October 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00376-4.

Dong:2019:PSR

[1950] Khuu Thi Phuong Dong, Yoko Saito, Nguyen Thi Ngoc Hoa, Tong Yen Dan, and Takashi Matsuishi. Pressure–state–response of traceability implementation in seafood-exporting countries: evidence from Vietnamese shrimp products. Aquaculture International, 27(5):1209–1229, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00378-2.

vanGorcum:2019:GMI

[1951] Bart van Gorcum, Simon Goddek, and Karel J. Keesman. Gaining market insights for aquaponically produced vegetables in Kenya. Aquaculture International, 27(5):1231–1237, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00379-1.

Nowland:2019:WTS

[1952] Samantha J. Nowland, Wayne A. O’Connor, Shane S. Penny, Matthew W. J. Osborne, and Paul C. Southgate. Water temperature and salinity synergistically affect embryonic and larval development of the tropical black-lip rock oyster Saccostrea echinata. Aquaculture International, 27 (5):1239–1250, October 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00381-7.

Lopardo:2019:PMD

[1953] Cristina R. Lopardo and Hidetoshi Urakawa. Performance and microbial diversity of bioreactors using polycaprolactone and polyhydroxyalkanoate as carbon source and biofilm carrier in a closed recirculating aquaculture system. Aquaculture International, 27(5):1251–1268, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00383-5.

Chong:2019:TSC

[1954] Jia Fong Chong, Widad Fadhullah, Vuanghao Lim, and Chee Keong Lee. Two-stage cultivation of the marine microalga Chlorella salina for starch and carbohydrate production. Aquaculture International, 27(5): 1269–1288, October 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00385-3.

Velasco:2019:NCW

[1955] L. A. Velasco, Y. Villarruel, and B. Toro. Nursery culture of wild and hatchery-produced juveniles of the West Indian top shell, Cittarium pica, fed microalgal and artificial biofilms. Aquaculture International, 27(5): 1289–1299, October 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00386-2.

Fu:2019:SAC

[1956] Longlong Fu, Tianheng Gao, Hucheng Jiang, Faqi Qiang, Yan Zhang, and Jianlin Pan. Staphylococcus aureus causes hepatopancreas browned disease and hepatopancreatic necrosis complications in Chinese mitten crab, Eriocheir sinensis. Aquaculture International, 27(5):1301– 1314, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00387-1.

Abdel-Razek:2019:AAC

[1957] Nashwa Abdel-Razek. Antimicrobial activities of chitosan nanoparticles against pathogenic microorganisms in Nile tilapia, Oreochromis niloticus. Aquaculture International, 27(5):1315–1330, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00388-0.

Huo:2019:LDJ

[1958] Zhongming Huo, Md. Golam Rbbani, Hao Cui, Longqiang Xu, Xiwu Yan, Lei Fang, Ye Wang, and Feng Yang. Larval development, juvenile survival, and burrowing rate of geoduck clams (Panopea japonica) under different pH conditions. Aquaculture International, 27(5): 1331–1342, October 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00389-z.

Kim:2019:DGP

[1959] Junseong Kim, Eun-A Kim, Hyun-Sung Yang, Do-Hyung Kang, Hyun-Soo Kim, Seo-Young Kim, You-Jin Jeon, and Soo-Jin Heo. Development of growth-promoting substances for diatoms (Navicula sp.). Aquaculture International, 27(5):1343–1352, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00391-5.

Cheng:2019:MRO

[1960] Xiangju Cheng, Yuning Xie, Dantong Zhu, and Jun Xie. Modeling re-oxygenation performance of fine-bubble–diffusing aeration system in aquaculture ponds. Aquaculture International, 27(5):1353–1368, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00390-6.

S:2019:VWS

[1961] Satheesh Kumar S, R. Ananda Bharathi, J. J. S. Rajan, V. Chitra, M. Muralidhar, and S. V. Alavandi. Viability of white spot syndrome virus (WSSV) in shrimp pond sediments with reference to physicochemical properties. Aquaculture International, 27(5):1369–1382, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00394-2.

Tsai:2019:GLG

[1962] Ching-Yi Tsai, Harvey M. Santos, Shao-Yang Hu, Cheng Yu Sang, Ciara Alyssa S. Yanuaria, Ernest Nicolo G. Lola, Lemmuel L. Tayo, Karmella Marie A. Nitura, Chun Hung Liu, and Kuo Pin Chuang. LpxD gene knockout elicits protection to Litopenaeus vannamei, white shrimp, against Vibrio parahaemolyticus infection. Aquaculture International, 27 (5):1383–1393, October 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00398-y.

Dijkgraaf:2019:MIA

[1963] Klaas Hielke Dijkgraaf, Simon Goddek, and Karel J. Keesman. Modeling innovative aquaponics farming in Kenya. Aquaculture International, 27 (5):1395–1422, October 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00397-z.

Oliveira:2019:GDN

[1964] Claudivane de Sá Teles Oliveira, Ricardo Franco Cunha Moreira, Aldeney Andrade Soares Filho, Soraia Barreto Aguiar Fonteles, and Norma Suely Evangelista-Barreto. Genetic diversity in natural populations of Colossomamacropomum in the Brazilian Amazon region and in populations farmed in Northeast Brazil based on ISSR markers. Aquaculture International, 27(5):1423–1434, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00395-1.

Reynaga-Franco:2019:MMI

[1965] F. J. Reynaga-Franco, E. A. Aragón-Noriega, J. M. Grijalva-Chon, R. Castro-Longoria, J. A. Arreola-Lizárraga, R. H. Barraza-Guardado, and J. Chávez-Villalba. Multi-model inference as criterion to determine differences in growth patterns of distinct Crassostrea gigas stocks. Aquaculture International, 27(5):1435–1450, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00396-0.

Vogeley:2019:GIG

[1966] Joana Lyra Vogeley, Juliana Aguiar Interaminense, Diego Souza Buarque, Suzianny Maria Bezerra Cabral da Silva, Maria Raquel Moura Coimbra, Sílvio Maurano Peixoto, and Roberta Borda Soares. Growth and immune gene expression of Litopenaeus vannamei fed Bacillus subtilis and Bacillus circulans supplemented diets and challenged with Vibrio parahaemolyticus. Aquaculture International, 27(5):1451–1464, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00399-x.

Poblete:2019:ITM

[1967] Exequiel Gonzalez Poblete, Benjamin M. Drakeford, Felipe Hurtado Ferreira, Makarena Garrido Barraza, and Pierre Failler. The impact of trade and markets on Chilean Atlantic salmon farming. Aquaculture International, 27(5):1465–1483, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00400-7.

Stavrakidis-Zachou:2019:DDS

[1968] Orestis Stavrakidis-Zachou, Anneliese Ernst, Christian Steinbach, Kai Wagner, and Uwe Waller. Development of denitrification in semiautomated moving bed biofilm reactors operated in a marine recirculating aquaculture system. Aquaculture International, 27(5):1485– 1501, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00402-5.

Nguyen:2019:RDA

[1969] Vuong Viet Nguyen, Channarong Rodkhum, Saengchan Senapin, and Ha Thanh Dong. Retrospective diagnosis of archived marine fish experienced unexplained mortality reveals dual infections of Nocardia seriolae and Streptococcus iniae. Aquaculture International, 27(5):1503– 1512, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00403-4.

Jaffer:2019:ICH

[1970] Y. D. Jaffer, H. Sanath Kumar, R. Vinothkumar, A. B. Irfan, N. M. Ishfaq, Parvaiz Ahmad Ganie, Raja Adil Hussain Bhat, and A. Vennila. Isolation and characterization of heterotrophic nitrification–aerobic denitrification and sulphur-oxidizing bacterium Paracoccus saliphilus strain SPUM from coastal shrimp ponds. Aquaculture International, 27(5): 1513–1524, October 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00407-0.

Bastos:2019:EEL

[1971] Argemiro Midonês Bastos, Jô Farias Lima, and Marcos Tavares-Dias. Effects of environmental light colors on the larviculture of the Amazon River prawn Macrobrachium amazonicum. Aquaculture International, 27 (5):1525–1534, October 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00409-y.

Akter:2019:IMO

[1972] Mst. Nahid Akter, Roshada Hashim, Amalia Sutriana, and Siti Azizah Mohd Nor. Influence of mannan oligosaccharide supplementation on haematological and immunological responses and disease resistance of striped catfish (Pangasianodon hypophthalmus Sauvage, 1878) juveniles. Aquaculture International, 27(5):1535–1551, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00408-z.

Pounds:2019:SEM

[1973] Alexandra Pounds and John Bostock. Suitable economic models for open educational resources initiative in aquaculture higher education. Aquaculture International, 27(5):1553–1563, October 2019. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00406-1.

M:2019:RBV

[1974] Hazreen-Nita M., Azila A., Mukai Y., Firdaus-Nawi M., and Nur-Nazifah M. A review of betanodavirus vaccination as preventive strategy to viral nervous necrosis (VNN) disease in grouper. Aquaculture International, 27 (5):1565–1577, October 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00410-5.

Jobling:2019:BRJb

[1975] Malcolm Jobling. Book review: J. Tidwell, S. Coyle and L. A. Bright (eds): Largemouth bass aquaculture. 5M Publishing, Sheffield, 2019, vi + 274 pp, £135 (Hardback), ISBN: 978-1-78918-048-0. Aquaculture International, 27(5):1579–1581, October 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00405-2.

Zhang:2019:SRD

[1976] Wei Zhang, Qihui Yang, Beiping Tan, Fengmei Wang, Xiaohui Dong, Shuyan Chi, Hongyu Liu, Shuang Zhang, and Hualang Wang. Study of the requirements of dietary cholesterol at two different growth stages of Pacific white shrimps, Litopenaeus vannamei. Aquaculture International, 27(6):1583–1597, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00411-4.

Liu:2019:ESD

[1977] Baoliang Liu, Fan Fei, Xiaotian Li, Xinyi Wang, and Bin Huang. Effects of stocking density on stress response, innate immune parameters, and welfare of turbot (Scophthalmus maximus). Aquaculture International, 27(6):1599–1612, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00413-2.

Sheikhzadeh:2019:SPR

[1978] Najmeh Sheikhzadeh, Shalaleh Mousavi, Ali Khani Oushani, Masoumeh Firouzamandi, and Karim Mardani. Spirulina platensis in rainbow trout (Oncorhynchus mykiss) feed: effects on growth, fillet composition, and tissue antioxidant mechanisms. Aquaculture International, 27(6):1613– 1623, December 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00412-3.

Stojanovic:2019:HCW

[1979] Katarina Stojanović, Miroslav Živić, Zoran Marković, Jelena – Dor– dević, Jelena Jovanović, and Ivana Živić. How changes in water quality under the influence of land-based trout farms shape chemism of the recipient streams — case study from Serbia. Aquaculture International, 27(6): 1625–1641, December 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00414-1.

Baldissera:2019:SPI

[1980] Matheus D. Baldissera, Carine de Freitas Souza, Lorenzo B. Abbad, Maria Izabel U. M. da Rocha, Marcelo L. da Veiga, Aleksandro S. da Silva, and Bernardo Baldisserotto. Saprolegnia parasitica impairs branchial phosphoryl transfer network in naturally infected grass carp (Ctenopharyngodon idella): prejudice on bioenergetic homeostasis. Aquaculture International, 27(6):1643–1654, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00416-z.

Zhou:2019:YCW

[1981] Jianjun Zhou, Zhen Gao, Wenbiao Wang, Feng Huang, Junpeng Hu, Aqiong Gong, Rui Wang, Wumin Yang, Jie Li, Xianqin Hu, and Xuedong Wang. Yeast cell walls stimulate viability, respiratory burst, and phagocytosis in channel catfish (Ictalurus punctatus) head-kidney macrophages. Aquaculture International, 27(6):1655–1665, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00417-y.

Menconi:2019:RLF

[1982] Vasco Menconi, Paolo Pastorino, Giulia Cavazza, Morena Santi, Davide Mugetti, Gaetano Zuccaro, and Marino Prearo. The role of live fish trade in the translocation of parasites: the case of Cystidicola farionis in farmed rainbow trout (Oncorhynchus mykiss). Aquaculture International, 27(6):1667–1671, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00422-1.

Ong:2019:IDR

[1983] Quy Moc Ong, Ravi Fotedar, and Thy Thi Truong Ho. Impact of different rearing systems on survival, growth and quality of mud crab (Scylla paramamosain) megalopae reared from early zoeae. Aquaculture International, 27(6):1673–1687, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00421-2.

Obiero:2019:PUA

[1984] Kevin Odhiambo Obiero, Herwig Waidbacher, Bryan Otieno Nyawanda, Jonathan Mbonge Munguti, Julius Otieno Manyala, and Boaz Kaunda-Arara. Predicting uptake of aquaculture technologies among smallholder fish farmers in Kenya. Aquaculture International, 27(6):1689– 1707, December 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00423-0.

Hynes:2019:DCE

[1985] Stephen Hynes, Elisa Ravagnan, and Brita Gjerstad. Do concerns for the environmental credentials of salmon aquaculture translate into WTP a price premium for sustainably farmed fish? A contingent valuation study in Ireland and Norway. Aquaculture International, 27(6):1709– 1723, December 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00425-y.

AbdElLatif:2019:HMC

[1986] A. M. Abd El Latif, Hiam Elabd, Aziza Amin, A. I. Noor Eldeen, and A. A. Shaheen. High mortalities caused by Aeromonas veronii: identification, pathogenicity, and histopathologicalstudies in Oreochromis niloticus. Aquaculture International, 27(6):1725–1737, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00429-8.

Sakaguchi:2019:MIW

[1987] Sakiko Orui Sakaguchi, Gen Ogawa, Hiroaki Kasai, Yuichi Shimizu, Hiroshi Kitazato, Katsunori Fujikura, and Kiyotaka Takishita. Molecular identification of water molds (oomycetes) associated with chum salmon eggs from hatcheries in Japan and possible sources of their infection. Aquaculture International, 27(6):1739–1749, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00427-w.

Zibiene:2019:ICP

[1988] Grazina Zibiene and Alvydas Zibas. Impact of commercial probiotics on growth parameters of European catfish (Silurus glanis) and water quality in recirculating aquaculture systems. Aquaculture International, 27(6): 1751–1766, December 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00428-9.

Gichana:2019:EPC

[1989] Zipporah Gichana, David Liti, Joseph Wakibia, Erick Ogello, Silke Drexler, Paul Meulenbroek, Robert Ondiba, Werner Zollitsch, and Herwig Waidbacher. Efficiency of pumpkin (Cucurbita pepo), sweet wormwood (Artemisia annua) and amaranth (Amaranthus dubius) in removing nutrients from a smallscale recirculating aquaponic system. Aquaculture International, 27(6):1767–1786, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00442-x.

Diab:2019:ABC

[1990] Amany M. Diab, R. H. Khalil, and M. Khallaf. Autogenous bacterins cross-protection as a trial for streptococcosis control in Oreochromis niloticus. Aquaculture International, 27(6):1787–1800, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00432-z.

Vanegas-Espinoza:2019:MAR

[1991] Pablo Emilio Vanegas-Espinoza, Verónica Pérez-Escalante, Gabriel Aguirre-Guzman, Javier Darío Hoyos-Leyva, and Alma Angélica Del Villar-Martínez. Microencapsulation of anthocyanins from roselle (Hibiscus sabdariffa) and its application on a pigment supplied diet to fantail goldfish (Carassius auratus). Aquaculture International, 27(6): 1801–1811, December 2019. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00430-1.

Forte:2019:MCC

[1992] Jamille Martins Forte, Luiz Fagner Ferreira Nogueira, Rafael dos Santos Rocha, Rodrigo Maggioni, and Oscarina Viana de Sousa. Multienzymatic capacity of cultivable intestinal bacteria from captive Litopenaeus vannamei (Boone, 1931) shrimp reared in green water. Aquaculture International, 27(6):1813–1824, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00431-0.

Filho:2019:DGI

[1993] Fernando de O. Roberti Filho, João Fernando Albers Koch, Christian Wallace, and Miguel Costa Leal. Dietary β-1,3/1,6-glucans improve the effect of a multivalent vaccine in Atlantic salmon infected with Moritella viscosa or infectious salmon anemia virus. Aquaculture International, 27(6):1825–1834, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-019-00436-9.

Panigrahi:2019:CEB

[1994] Akshaya Panigrahi, Palanichamy Esakkiraj, Sundaresan Jayashree, Chakrapani Saranya, Rashmi Ranjan Das, and Mani Sundaram. Colonization of enzymatic bacterial flora in biofloc grown shrimp Penaeus vannamei and evaluation of their beneficial effect. Aquaculture International, 27(6):1835–1846, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00434-x.

Abdollahi:2019:CEA

[1995] Yaser Abdollahi, Nasrollah Ahmadifard, Naser Agh, Kaveh Rahmanifarah, and Mohammad Amin Hejazi. β-carotene-enriched Artemia as a natural carotenoid improved skin pigmentation and enhanced the mucus immune responses of platyfish Xiphophorus maculatus. Aquaculture International, 27(6):1847–1858, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00437-8.

Han:2019:CIF

[1996] Ming Ming Han, Fu Hong Ding, and Jun Li. Characterization of the immunological factors in turbot (Scophthalmus maximus) seminal plasma and their correlation with sperm quality. Aquaculture International, 27(6):1859–1868, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-019-00435-w.

Crichigno:2019:RTO

[1997] Sonia Alejandra Crichigno and Víctor Enrique Cussac. Rainbow trout (Oncorhynchus mykiss) adaptation to a warmer climate: the performance of an improved strain under farm conditions. Aquaculture International, 27(6):1869–1882, December 2019. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00438-7.

Nafiqoh:2020:APA

[1998] Nunak Nafiqoh, Sukenda, M. Zairin, Alimuddin, Angela Lusiastuti, Samira Sarter, Domenico Caruso, and Jean-Christophe Avarre. Antimicrobial properties against Aeromonas hydrophila and immunostimulant effect on Clarias gariepinus of Piper betle, Psidium guajava, and Tithonia diversifolia plants. Aquaculture International, 28(1):1– 13, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00439-6.

Han:2020:AAE

[1999] Qing Han, Hongbo Fan, Jin Peng, Liliu Zhou, and Lian Gan. Ascorbic acid enhanced the growth performance, oxidative status, and resistance to Aeromonas hydrophila challenge of juvenile grass carp (Ctenopharyngodon idella). Aquaculture International, 28(1):15–30, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00440-z.

Shen:2020:ARG

[2000] Min Shen, Yanting Cui, Renjie Wang, Tiantian Dong, Haibin Ye, Shusheng Wang, Ranghui Fu, and Yuquan Li. Acute response of Pacific white shrimp Litopenaeus vannamei to high-salinity reductions in osmosis-, metabolism-, and immune-related enzyme activities. Aquaculture International, 28(1):31–39, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00441-y.

Antonucci:2020:PAS

[2001] Francesca Antonucci and Corrado Costa. Precision aquaculture: a short review on engineering innovations. Aquaculture International, 28(1): 41–57, February 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00443-w.

Jiang:2020:SGY

[2002] Weiwei Jiang, Fan Lin, Meirong Du, Jianguang Fang, Jinghui Fang, Yaping Gao, Xiaoqin Wang, Fengxue Li, Shipeng Dong, Xing Hou, and Zengjie Jiang. Simulation of Yesso scallop, Patinopecten yessoensis, growth with a dynamic energy budget (DEB) model in the mariculture area of Zhangzidao Island. Aquaculture International, 28(1): 59–71, February 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00447-6.

Bilen:2020:ECM

[2003] Soner Bilen, Abdelsalam M. O. Filogh, Abobaker Barka Ali, Osman Nezih Kenanoğlu, and Mehmet Arif Zoral. Effect of common mallow (Malva sylvestris) dietary supplementation on growth performance, digestive enzyme activities, haemotological and immune responses of common carp (Cyprinus carpio). Aquaculture International, 28(1):73– 84, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00444-9.

Zhang:2020:MIR

[2004] Weijia Zhang, Ronghua Li, Xingqiang Chen, Chunlin Wang, Zhongqi Gu, Changkao Mu, Weiwei Song, Pingping Zhan, and Ji Huang. Molecular identification reveals hybrids of Mytilus coruscus × Mytilus galloprovincialis in mussel hatcheries of China. Aquaculture International, 28 (1):85–93, February 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00445-8.

Wassef:2020:SPD

[2005] Elham A. Wassef, Norhan E. Saleh, Nabila E. Abdel-Meguid, Khouloud M. Barakat, Heba H. Abdel-Mohsen, and Nagi M. El-bermawy. Sodium propionate as a dietary acidifier for European seabass (Dicentrarchus labrax) fry: immune competence, gut microbiome, and intestinal histology benefits. Aquaculture International, 28(1):95–111, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00446-7.

Palanikumar:2020:UPN

[2006] Pandi Palanikumar, Dinamella Wahjuningrum, Paramachandran Abinaya, Mariavincent Michael Babu, and Thavasimuthu Citarasu. Usage of plant natural products for prevention and control of white feces syndrome (WFS) in Pacific whiteleg shrimp Litopenaeus vannamei farming in India. Aquaculture International, 28(1):113–125, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00448-5.

Bertolini:2020:SAC

[2007] Rafaela Manchin Bertolini, Lucia Suarez Lopez, Nivaldo Ferreira do Nascimento, Dilberto Ribeiro Arashiro, Diógenes Henrique de Siqueira Silva, Silvio Carlos Alves dos Santos, José Augusto Senhorini, and George Shigueki Yasui. Strategies for aquaculture and conservation of neotropical catfishes based on the production of triploid Pimelodus maculatus. Aquaculture International, 28(1):127–137, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00449-4.

Lu:2020:ARF

[2008] Yongzhong Lu, Qian Qiu, Chen Li, Linyue Cheng, and Jie Liu. Antioxidant responses of Fenneropenaeus chinensis to white spot syndrome virus challenge. Aquaculture International, 28(1):139–151, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00450-x.

Korni:2020:RMO

[2009] Fatma M. M. Korni, Fatma I. Abo El-Ela, and Usama K. Moawad. Role of Moringa oleifera leaves and aqueous extract in prevention of Motile Aeromonas Septicemia in common carp, Cyprinus carpio fingerlings with a reference to histopathological alterations. Aquaculture International, 28 (1):153–168, February 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00452-9.

Santos:2020:DPT

[2010] Harvey M. Santos, Ching-Yi Tsai, Kenth Roger A. Maquiling, Lemmuel L. Tayo, Abdul R. Mariatulqabtiah, Chi-Wen Lee, and Kuo Pin Chuang. Diagnosis and potential treatments for acute hepatopancreatic necrosis disease (AHPND): a review. Aquaculture International, 28(1): 169–185, February 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00451-w.

Yavuzer:2020:IIP

[2011] Emre Yavuzer, Fatih Özogul, and Yesim Özogul. Impact of icing with potato, sweet potato, sugar beet, and red beet peel extract on the sensory, chemical, and microbiological changes of rainbow trout (Oncorhynchus mykiss) fillets stored at (3 ± 1^{o} C). Aquaculture International, 28(1): 187–197, February 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00454-7.

Vavrecka:2020:IEA

[2012] Antonín Vavrečka, Petra Šánová, and Lukáš Kalous. Insight into the economy of aquaculture production in Czechia: assessment of aquaculture enterprises. Aquaculture International, 28(1):199–209, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00453-8.

Mukherjee:2020:IEN

[2013] Priyanka Mukherjee, Prakash Chandra Gorain, Ishita Paul, Rahul Bose, P. B. S. Bhadoria, and Ruma Pal. Investigation on the effects of nitrate and salinity stress on the antioxidant properties of green algae with special reference to the use of processed biomass as potent fish feed ingredient. Aquaculture International, 28(1):211–234, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00455-6.

Das:2020:DMV

[2014] Soumya Das, R. Aswani, B. Jasim, K. S. Sebastian, E. K. Radhakrishnan, and Jyothis Mathew. Distribution of multi-virulence factors among Aeromonas spp. isolated from diseased Xiphophorus hellerii. Aquaculture International, 28(1):235–248, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00456-5.

Capelle:2020:ESD

[2015] Jacob J. Capelle, Eva Hartog, Jorik Creemers, Jouke Heringa, and Pauline Kamermans. Effects of stocking density and immersion time on the performance of oysters in intertidal off-bottom culture. Aquaculture International, 28(1):249–264, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00460-9.

Li:2020:ICP

[2016] Wenjuan Li, Jiamin Liu, Xingrong Xuan, Qian Li, Zhiyi Shi, Ye Zhang, Zhiyi Bai, and Yuanshuai Fu. Induced cell proliferation in pearl oyster. Aquaculture International, 28(1):265–276, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00459-2.

Kuebutornye:2020:MNU

[2017] Felix K. A. Kuebutornye, Emmanuel Delwin Abarike, Michael Essien Sakyi, Yishan Lu, and Zhiwen Wang. Modulation of nutrient utilization, growth, and immunity of Nile tilapia, Oreochromis niloticus: the role of probiotics. Aquaculture International, 28(1):277–291, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00463-6.

Valente:2020:TDR

[2018] Cecília de Souza Valente, Karla Oliveira Ortiz, Rachel Depperschmidt, Ana Paula de Medeiros Fraga, Felipe do Nascimento Vieira, and Maria Risoleta Freire Marques. Transcription of defense related genes in Pacific white shrimp, Litopenaeus vannamei, kept in biofloc and in clear seawater and challenged with the white spot syndrome virus. Aquaculture International, 28(1):293–307, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00461-8.

Farzadnia:2020:MTD

[2019] Akram Farzadnia and Mohsen Naeemipour. Molecular techniques for the detection of bacterial zoonotic pathogens in fish and humans. Aquaculture International, 28(1):309–320, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00462-7.

Saleh:2020:TSN

[2020] Norhan E. Saleh, Elham A. Wassef, and Ahmed M. Ashry. Is a taurine supplement necessary in fishmeal-based feeds for juvenile European sea bass (Dicentrarchus labrax)? Aquaculture International, 28 (1):321–333, February 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00464-5.

Xu:2020:CET

[2021] Chengxun Xu, Qi Li, and Jindou Chong. Combined effect of temperature, salinity, and rearing density on the larval growth of the black shell strain and wild population of the Pacific oyster Crassostrea gigas. Aquaculture International, 28(1):335–347, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00465-4.

Megarajan:2020:MDB

[2022] Sekar Megarajan, Ritesh Ranjan, Biji Xavier, Shubhadeep Ghosh, Narasimhulu Sadu, Chinnibabu Bathina, and Achamveettil Gopalakrishnan. Molecular detection of betanodavirus in orange-spotted grouper (Epinephelus coioides) broodstock maintained in recirculating aquaculture systems and sea cages. Aquaculture International, 28(1):349–362, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00466-3.

Eissa:2020:MEI

[2023] Alaa Eldin Eissa, Said K. Abolghait, Nehal A. Younis, Amina A. Dessouki, Maather M. El-Lamie, Abdelsalam A. Abu Mhara, and Mohamed Abdelsalam. Myxobolus episquamalis infection in farmed flathead grey mullet Mugil cephalus L. and thin-lipped mullet Liza ramada. Aquaculture International, 28(1):363–376, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00467-2.

Hassani:2020:ECS

[2024] Mir Hamed Sayed Hassani, Ayoub Yousefi Jourdehi, Ali Hosseinpour Zelti, Alireza Shenavar Masouleh, and Frozan Bagherzadeh Lakani. Effects of commercial superzist probiotic on growth performance and hematological and immune indices in fingerlings Acipenser baerii. Aquaculture International, 28(1):377–387, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00468-1.

El-Gawad:2020:EDM

[2025] Eman A. Abd El-Gawad, Amel M. El Asely, Eman I. Soror, Amany A. Abbass, and Brian Austin. Effect of dietary Moringa oleifera leaf on the immune response and control of Aeromonas hydrophila infection in Nile tilapia (Oreochromis niloticus) fry. Aquaculture International, 28 (1):389–402, February 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00469-0.

Campos:2020:PDS

[2026] Clarissa Vilela Figueiredo da Silva Campos, Renata da Silva Farias, Suzianny Maria Bezerra Cabral da Silva, William Severi, Luis Otavio Brito, and Alfredo Olivera Gálvez. Production of Daphnia similis Claus, 1876 using wastewater from tilapia cultivation in a biofloc system. Aquaculture International, 28(1):403–419, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00470-7.

Hoai:2020:RSP

[2027] Truong Dinh Hoai. Reproductive strategies of parasitic flatworms (Platyhelminthes, Monogenea): the impact on parasite management in aquaculture. Aquaculture International, 28(1):421–447, February 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00471-6.

Rivadeneyra:2020:HAI

[2028] Norma L. S. Rivadeneyra, Omar Mertins, Ruben C. Cuadros, Jose C. O. Malta, Lorena V. de Matos, and Patrick D. Mathews. Histopathology associated with infection by Procamallanus (Spirocamallanus) inopinatus (Nematoda) in farmed Brycon cephalus (Characiformes) from Peru: a potential fish health problem. Aquaculture International, 28 (2):449–461, April 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00474-3.

Amin:2020:GET

[2029] Muhamad Amin, Christopher J. S. Bolch, Mark B. Adams, and Christopher M. Burke. Growth enhancement of tropical abalone, Haliotis asinina L., through probiotic supplementation. Aquaculture International, 28(2):463–475, April 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00473-4.

Moses:2020:CGS

[2030] Mbiru Moses, Matern S. P. Mtolera, Leonard J. Chauka, Fernando A. Lopes, Dirk Jan de Koning, Ross D. Houston, and Christos Palaiokostas. Characterizing the genetic structure of introduced Nile tilapia (Oreochromis niloticus) strains in Tanzania using double digest RAD sequencing. Aquaculture International, 28(2):477–492, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00472-5.

Abdel-Tawwab:2020:DGC

[2031] Mohsen Abdel-Tawwab, Eman Khalifa, Amany M. Diab, Mohamed A. Khallaf, Nashwa Abdel-Razek, and Riad H. Khalil. Dietary garlic and chitosan alleviated zearalenone toxic effects on performance, immunity, and challenge of European sea bass, Dicentrarchus labrax, to Vibrio alginolyticus infection. Aquaculture International, 28(2):493– 510, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00477-0.

Lin:2020:EDK

[2032] Xue Lin, Baitao Jin, Hongquan Wang, and Yurong Zhao. Effects of diet α-ketoglutarate (AKG) supplementation on the growth performance, antioxidant defense system, intestinal digestive enzymes, and immune response of grass carp (Ctenopharyngodon idellus). Aquaculture International, 28(2):511–524, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00475-2.

Ouyang:2020:OCE

[2033] Ping Ouyang, Yaojia Zhou, Ruixue Yang, Zhuangzhi Yang, Kaiyu Wang, Yi Geng, Weiming Lai, Xiaoli Huang, Defang Chen, Jing Fang, Zhengli Chen, Li Tang, Chao Huang, and Lizi Yin. Outbreak of carp edema virus disease in cultured ornamental koi in a lower temperature in China. Aquaculture International, 28(2):525–537, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00476-1.

Mbokane:2020:EDA

[2034] Esau Matthews Mbokane and Ngonidzashe Adreck Gukuta Moyo. Effect of dietary Artemisia afra on growth, some innate immunological parameters in Clarias gariepinus challenged with Aeromonas hydrophila. Aquaculture International, 28(2):539–553, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00479-y.

Mokrani:2020:ETR

[2035] Ahmed Mokrani, Mingchun Ren, Hualiang Liang, Qiang Yang, Ke Ji, Hopeson Chisomo Kasiya, and Xianping Ge. Effect of the total replacement of fishmeal with plant proteins and supplemental essential amino acids in the extruded diet on antioxidants genes, enzyme activities, and immune response in juvenile blunt snout bream. Aquaculture International, 28(2):555–568, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00480-5.

Armstrong:2020:TCS

[2036] Ethan G. Armstrong, Joseph Mersereau, Flora Salvo, Dounia Hamoutene, and Suzanne C. Dufour. Temporal change in the spatial distribution of visual organic enrichment indicators at aquaculture sites in Newfoundland, Canada. Aquaculture International, 28(2):569–586, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00478-z.

Samayanpaulraj:2020:ERR

[2037] Vignesh Samayanpaulraj, Vijay Velu, Muthukumar Sivaramapillai, Krishnaveni Govindaraj, and Ramesh Uthandakalaipandiyan. Extraction and recovery response of Penaeus indicus chitosan against Aeromonas hydrophila Ah17 infected snakehead murrel Channa striata. Aquaculture International, 28(2):587–602, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00481-4. See correction [2058].

Yang:2020:GEG

[2038] Tianyan Yang, Shuang Yang, Wei Meng, Tianxiang Gao, and Xiumei Zhang. Genetic evaluation of Chinese shrimp (Fenneropenaeus chinensis) stock enhancement in the Yellow Sea and Bohai Sea based on mitochondrial DNA control region. Aquaculture International, 28 (2):603–614, April 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00482-3.

Yan:2020:EPT

[2039] Mao-Cang Yan, Yao-Hua Wang, De-Wei Ji, Fang Li, Li-Hua Hu, and Min Zhang. Expression profiles of three crustin-like genes in Litopenaeus vannamei under Vibrio harveyi infection. Aquaculture International, 28 (2):615–624, April 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00484-1.

Wen:2020:EPA

[2040] Xiaobin Wen, Zhongjie Wang, Yi Ding, Yahong Geng, and Yeguang Li. Enhancing the production of astaxanthin by mixotrophic cultivation of Haematococcus pluvialis in open raceway ponds. Aquaculture International, 28(2):625–638, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00483-2.

Nuwansi:2020:EHL

[2041] K. K. T. Nuwansi, A. K. Verma, G. Rathore, M. H. Chandrakant, G. P. W. A. Prabhath, and R. M. Peter. Effect of hydraulic loading rate on the growth of koi carp (Cyprinus carpio var. koi) and Gotukola (Centella asiatica (L.)) using phytoremediated aquaculture wastewater in aquaponics.

Aquaculture International, 28(2):639–652, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00485-0.

Jiang:2020:SCR

[2042] Fangyan Jiang, Ning Yang, Hai Huang, Huimin Feng, Yu Li, and Bingbing Han. Short communication: recovery of Vibrio vulnificus from head ulceration in seahorse (Hippocampus kuda). Aquaculture International, 28(2):653–660, April 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00486-z.

Campos-Garcia:2020:FDH

[2043] Julio C. Campos-García, Roberto Rodríguez-Ramírez, Luz A. Avila-Villa, Oscar Gómez-Aldama, Joel Arias-Martínez, and Teresa Gollas-Galván. Fractal dimension of hepatopancreas of white shrimp Litopenaeus vannamei infected with Hepatobacter penaei bacteria (NHPB). Aquaculture International, 28(2):661–673, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00487-y.

Theodoridis:2020:RPE

[2044] Alexandros Theodoridis, Athanasios Ragkos, and Georgia Koutouzidou. Revealing the profile of economically efficient mussel farms: a restricted data envelopment analysis application. Aquaculture International, 28 (2):675–689, April 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00490-3.

Gupta:2020:CEM

[2045] Shobha Gupta, Z. P. Bhathena, Sarvendra Kumar, P. M. Nuzaiba, P. P. Srivastava, Subodh Gupta, and Sanjay Balkrishna Jadhao. Comparative efficacy of mannan-oligosaccharides from two yeast species fed alone or in combination with probiotic Bacillus subtilis ATCC 6633 to catla (Catla catla) juveniles. Aquaculture International, 28(2):691– 710, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00488-x.

Cardoso:2020:LCS

[2046] C. Cardoso, H. Pereira, J. Franca, J. Matos, I. Monteiro, P. Pousão-Ferreira, A. Gomes, L. Barreira, J. Varela, N. Neng, J. M. Nogueira, C. Afonso, and N. M. Bandarra. Lipid composition and some bioactivities of 3 newly isolated microalgae (Tetraselmis sp. IMP3, Tetraselmis sp. CTP4, and Skeletonema sp.). Aquaculture International, 28(2):711– 727, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00489-w.

Loh:2020:OIE

[2047] Jiun Yan Loh, Hor Kuan Chan, Hok Chai Yam, Lionel Lian Aun In, and Crystale Siew Ying Lim. An overview of the immunomodulatory effects exerted by probiotics and prebiotics in grouper fish. Aquaculture International, 28(2):729–750, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00491-2.

Wachirachaikarn:2020:DNM

[2048] Anyalak Wachirachaikarn, Onprang Sutthakiet, Wansuk Senanan, and Uthairat Na-Nakorn. Development of the new microsatellite multiplex PCR panel and genetic variation of farmed snakeskin gourami, Trichopodus pectoralis. Aquaculture International, 28(2):751–765, April 2020. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-019-00492-
1.

Pan:2020:STG

[2049] Tingshuang Pan and Maocang Yan. The screening of traditional Chinese herbs on nonspecific immune response and protection of Pacific white shrimp (Litopenaeus vannamei) from Vibrio harveyi infection. Aquaculture International, 28(2):767–776, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00493-0.

Emmanuele:2020:ARA

[2050] Pietro Emmanuele, Antonio Casalini, Daniela Pisati, Rebecca Andreini, Niccolò Guercilena, Albamaria Parmeggiani, Annalisa Zaccaroni, and Oliviero Mordenti. Artificial reproduction of Anguilla anguilla: evaluation of biometrics characteristics of a population from Valle Campo Lagoon, Comacchio (Italy). Aquaculture International, 28(2):777–790, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00494-z.

Dai:2020:CGC

[2051] Yun-Yun Dai, Yong-Ming Yuan, Yuan Yuan, Zhen Zhou, and Hong-Yan Zhang. Competitiveness of Chinese and Indonesian tilapia exports in the US market. Aquaculture International, 28(2):791–804, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00496-x.

Kumaran:2020:ABP

[2052] M. Kumaran, T. K. Ghoshal, Debasis De, Gouranga Biswas, R. Ananda Raja, P. S. Anand, A. Panigrahi, and K. K. Vijayan. Aquaculture-based production systems for the livelihood security of coastal farm families in the risk-prone agro-ecosystem of India: an appraisal. Aquaculture International, 28(2):805–814, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00495-y.

Duncan:2020:ATP

[2053] Wallice Paxiúba Duncan, Jhonatan Junior Silva Idalino, André Gentil da Silva, Rebeca Fontenele Moda, Hallana Cristina Menezes da Silva, Daniele Aparecida Matoso, and Ana Lúcia Silva Gomes. Acute toxicity of the pesticide trichlorfon and inhibition of acetylcholinesterase in Colossoma macropomum (Characiformes: Serrasalmidae). Aquaculture International, 28(2):815–830, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-019-00497-w.

Cai:2020:ODP

[2054] Zuo-Nan Cai, Xue-Qiao Qian, and Shou-Qi Xie. Optimal dietary protein concentrations for largemouth bass (Micropterus salmoides) of different sizes (10–500 g). Aquaculture International, 28(2):831–840, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-019-00498-9.

Mantilla-Aldana:2020:PIN

[2055] Leonardo Mantilla-Aldana, Pedro Campoy-López, Estefanía Pereira-Pinto, and Ricardo Beiras. Pixelar index: a new quantitative approach for the assessment of reproductive condition in the purple sea urchin, Paracentrotus lividus, by image analysis. Aquaculture International, 28 (2):841–850, April 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00499-8.

Ucar:2020:ENS

[2056] Yilmaz Ucar, Yesim Özogul, Fatih Özogul, Mustafa Durmuş, and Ali Riza Köşker. Effect of nisin on the shelf life of sea bass (Dicentrarchus labrax L.) fillets stored at chilled temperature (4 ± 2^{o} C). Aquaculture International, 28(2):851–863, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00512-5.

Malinovskyi:2020:CBP

[2057] Oleksandr Malinovskyi, Jitka Kolářová, Miroslav Blecha, Alžběta Stará, Josef Velíšek, Jiří Křišťan, and Tomáš Policar. Correction to: Behavior and physiological status of pond-cultured pikeperch (Sander lucioperca) broodstock effected by sexual interactions throughout semi-artificial reproduction. Aquaculture International, 28(2):865, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00418-x. See

[1941].

Samayanpaulraj:2020:CER

[2058] Vignesh Samayanpaulraj, Vijay Velu, Muthukumar Sivaramapillai, Krishnaveni Govindaraj, and Ramesh Uthandakalaipandiyan. Correction to: Extraction and recovery response of Penaeus indicus chitosan against Aeromonas hydrophila Ah17 infected snakehead murrel Channa striata. Aquaculture International, 28(2):867, April 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-019-00500-4. See [2037].

Ahmed:2020:IEE

[2059] Imtiaz Ahmed, Quseen Mushtaq Reshi, and Francesco Fazio. The influence of the endogenous and exogenous factors on hematological parameters in different fish species: a review. Aquaculture International, 28(3):869–899, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00501-3.

El-Jakee:2020:ICM

[2060] Jakeen El-Jakee, Sarah Elshamy, Abdel-Wahed Hassan, Mohamed Abdelsalam, Nehal Younis, Maha A. El-Hady, and Alaa Eldin Eissa. Isolation and characterization of Mycoplasmas from some moribund Egyptian fishes. Aquaculture International, 28(3):901–912, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00502-2.

deOliveira:2020:NPF

[2061] Samira Teixeira Leal de Oliveira, Riani Ananda Nunes Soares, Sília Maria de Negreiros Sousa, Antônio Wilton Cavalcante Fernandes, Gisele Veneroni Gouveia, and Mateus Matiuzzi da Costa. Natural products as functional food ingredients for Nile tilapia challenged with Aeromonas hydrophila. Aquaculture International, 28(3):913–926, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-019-00503-1.

Araujo:2020:DEG

[2062] J. Araújo, F. Soares, A. Medeiros, N. M. Bandarra, M. Freire, M. Falcão, and P. Pousão-Ferreira. Depth effect on growth and fatty acid profile of Mediterranean mussel (Mytilus galloprovincialis) produced on a longline off south Portugal. Aquaculture International, 28(3):927– 946, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-019-00504-0.

Abdel-Aziz:2020:RDS

[2063] Mohamed Abdel-Aziz, Mohamed Bessat, Amr Fadel, and Samar Elblehi. Responses of dietary supplementation of probiotic effective microorganisms (EMs) in Oreochromis niloticus on growth, hematological, intestinal histopathological, and antiparasitic activities. Aquaculture International, 28(3):947–963, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-019-00505-z.

Kuebutornye:2020:CMP

[2064] Felix K. A. Kuebutornye and Emmanuel Delwin Abarike. The contribution of medicinal plants to tilapia aquaculture: a review. Aquaculture International, 28(3):965–983, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00506-3.

Zhang:2020:DON

[2065] Chenglin Zhang, Benben Song, Jianjun Shan, Qi Ni, Fan Wu, and Shiming Wang. Design and optimization of a new tube aeration device. Aquaculture International, 28(3):985–999, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00507-2.

Rossi:2020:IMM

[2066] L. T. Rossi, A. Romero Sharpen, J. A. Zimmermann, C. R. Olivero, M. V. Zbrun, L. S. Frizzo, M. L. Signorini, C. Bacchetta, R. E. Cian, J. Cazenave, L. P. Soto, and S. R. Drago. Intestinal microbiota modulation in juvenile Pacú (Piaractus mesopotamicus) by supplementation with Pyropia columbina and β-carotene. Aquaculture International, 28 (3):1001–1016, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00508-1.

Amenyogbe:2020:EPP

[2067] Eric Amenyogbe, Gang Chen, Zhongliang Wang, JianSheng Huang, Baosong Huang, and Hongjuan Li. The exploitation of probiotics, prebiotics and synbiotics in aquaculture: present study, limitations and future directions: a review. Aquaculture International, 28(3):1017– 1041, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00509-0.

Rivero-Martinez:2020:APE

[2068] Psique Victoria Rivero-Martínez, Héctor Castillo-Juárez, Eugenia Guadalupe Cienfuegos Rivas, José Luis Pablos-Hash, David Alberto Martínez-Espinosa, and Gabriel Ricardo Campos-Montes. Analysis of the preferences of the esthetic traits and their morphotypes candidates to selection criteria in tiger barb (Puntius tetrazona) in the actors of the productive chain. Aquaculture International, 28(3):1043–1055, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00510-7.

Yadav:2020:ESE

[2069] Anamika Yadav, Avinash Kumar, and Sudipto Sarkar. An experimental study to evaluate the efficacy of air entrainment holes on the throat of a venturi aeration system. Aquaculture International, 28(3): 1057–1068, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00511-6.

Sukenda:2020:IGG

[2070] Sukenda Sukenda, Lila Gardenia, M. Zairin, Angela Lusiastuti, and Alimuddin Alimudin. Identification of giant gourami iridovirus (GGIV): a new infectious spleen and kidney necrosis virus (ISKNV) from natural outbreak in cultured Osphronemus goramy. Aquaculture International, 28(3):1069–1082, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00513-4.

Hayatgheib:2020:RFF

[2071] Niki Hayatgheib, Emmanuelle Moreau, Ségolène Calvez, Didier Lepelletier, and Hervé Pouliquen. A review of functional feeds and the control of Aeromonas infections in freshwater fish. Aquaculture International, 28 (3):1083–1123, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00514-3.

Musharraf:2020:DCR

[2072] Mohammad Musharraf and Mukhtar A. Khan. Dietary calcium requirement of fingerling Indian major carp, Labeo rohita (Hamilton) based on growth performance, tissue mineralization, whole body, and serum biochemical composition. Aquaculture International, 28(3):1125– 1139, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00515-2.

Souza:2020:MAR

[2073] José Gilmar da Silva Souza, Lais Teodoro Libeck, Bárbara do Carmo Rodrigues Virote, Renata Catão Egger, Gabriel Cândido Ribeiro de Sá, Gilmara Junqueira Machado, and Luis David Solis Murgas. A method to analyze the relationship between locomotor activity and feeding behaviour in larvae of Betta splendens. Aquaculture International, 28(3): 1141–1152, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00516-1.

Xie:2020:FPC

[2074] Dizhi Xie, Meng Wang, Shuqi Wang, Cuihong You, Mengmeng Li, Yongcai Ma, Lihua Sun, Yong Wang, and Yuanyou Li. Fat powder can be a feasible lipid source in aquafeed for the carnivorous marine teleost golden pompano, Trachinotus ovatus. Aquaculture International, 28(3): 1153–1168, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00517-0.

Kaur:2020:AET

[2075] Amandeep Kaur, Shanthanagouda A. Holeyappa, Neelam Bansal, Vaneet Inder Kaur, and Anuj Tyagi. Ameliorative effect of turmeric supplementation in feed of Labeo rohita (Linn.) challenged with pathogenic Aeromonas veronii. Aquaculture International, 28(3):1169– 1182, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00518-z.

Veiga:2020:SCB

[2076] Pamela Thainara do Nascimento Veiga, Marco Shizuo Owatari, André Luiz Nunes, Robson Andrade Rodrigues, Rodrigo Y. Dichoff Kasai, Carlos Eurico Fernandes, and Cristiane Meldau de Campos. Short communication: Bacillus subtilis C-3102 improves biomass gain, innate defense, and intestinal absorption surface of native Brazilian hybrid surubim (Pseudoplatystoma corruscans × P. reticulatum). Aquaculture International, 28(3):1183–1193, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00519-y.

Yu:2020:SNA

[2077] Jin-Kai Yu and Jian-Qiu Ma. Social network analysis as a tool for the analysis of the international trade network of aquatic products. Aquaculture International, 28(3):1195–1211, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00520-5.

Changpasert:2020:OPP

[2078] Waleewan Changpasert and Saou-Lien Wong. Optimized process parameters for production of Chlorella pyrenoidosa biomass and lipids using response surface methodology. Aquaculture International, 28(3): 1213–1226, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00522-3.

Olier:2020:DVS

[2079] Bruno Silva Olier, Jiovani Sergio Bee Tubin, Giovanni Lemos de Mello, Marcel Martínez-Porchas, and Maurício Gustavo Coelho Emerenciano. Does vertical substrate could influence the dietary protein level and zootechnical performance of the Pacific white shrimp Litopenaeus vannamei reared in a biofloc system? Aquaculture International, 28(3): 1227–1241, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00521-4.

Rahman:2020:MCF

[2080] M. Mojibar Rahman, Shahroz Mahean Haque, Shams M. Galib, M. Ashraful Islam, Md. Taskin Parvez, Md. Nazmul Hoque, M. Abdul Wahab, Hillary Egna, and Christopher Brown. Mud crab fishery in climate vulnerable coastal Bangladesh: an analysis towards sustainable development. Aquaculture International, 28(3):1243–1268, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00523-2.

Domingues:2020:EPF

[2081] Pedro Domingues, Carlos Rosas, Ariadna Sánchez, Maria Eugenia Chimal, and Pedro Gallardo. Effect of the processing of four marine species on their in vitro digestibility of Octopus vulgaris adults. Aquaculture International, 28(3):1269–1277, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00524-1.

Zhang:2020:HLE

[2082] Tao Zhang, Wenju Xu, Shude Xu, Jude Juventus Aweya, and Wenhua Liu. High-level expression and characterization of a lipase from Trichosporon fermentans Y3 and its application as an aquafeed additive for grouper (Epinephelus coioides). Aquaculture International, 28 (3):1279–1292, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00525-0.

Azhar:2020:BBC

[2083] Muhammad H. Azhar, Suciyono Suciyono, Darmawan S. Budi, Mohammad F. Ulkhaq, Mai Anugrahwati, and Julie Ekasari. Bioflocbased co-culture systems of Nile tilapia (Oreochromis niloticus) and redclaw crayfish (Cherax quadricarinatus) with different carbon–nitrogen ratios. Aquaculture International, 28(3):1293–1304, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00526-z.

Grassi:2020:GPF

[2084] Thiago Luís Magnani Grassi, Natália Mingues Paiva, Dayse Lícia Oliveira, Fábio Taniwaki, Jefferson Felipe Cavazzana, Gisele Cristina Rocha da Costa Camargo, Juliana Campos Pereira Diniz, Rubén Bermejo-Poza, Ricardo Borghesi, Morris Villarroel, and Elisa Helena Giglio Ponsano. Growth performance and flesh quality of tilapia (Oreochromis niloticus) fed low concentrations of Rubrivivax gelatinosus, Saccharomyces cerevisiae and Spirulina platensis. Aquaculture International, 28(3): 1305–1317, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00527-y.

Du:2020:PGR

[2085] Huanmin Du, Jiali Ren, Zhe Li, Haonan Zhang, Kang Wang, Bin Lin, Shanmin Zheng, Changyu Zhao, Chunxiao Meng, and Zhengquan Gao. Plant growth regulators affect biomass, protein, carotenoid, and lipid production in Botryococcus braunii. Aquaculture International, 28(3): 1319–1340, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00528-x.

Geetha:2020:TCI

[2086] R. Geetha, T. Ravisankar, P. K. Patil, Satheesha Avunje, S. Vinoth, C. V. Sairam, and K. K. Vijayan. Trends, causes, and indices of import rejections in international shrimp trade with special reference to India: a 15-year longitudinal analysis. Aquaculture International, 28 (3):1341–1369, June 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00529-w.

Yalcin:2020:GLC

[2087] Dilek Yalcin. Growth, lipid content, and fatty acid profile of freshwater cyanobacteria Dolichospermum affine (Lemmermann) Wacklin, Hoffmann, & Komárek by using modified nutrient media. Aquaculture International, 28(3):1371–1388, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00531-2.

Asuman:2020:ESL

[2088] Yilmaz Hatice Asuman. The effects of soy lecithin–enriched artemia on growth and survival of the early stages of green tiger shrimp (Penaeus semisulcatus). Aquaculture International, 28(3):1389–1402, June 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00536-x.

Yu:2020:IIP

[2089] Peng Yu, Teng Wang, Haibin Ye, Hongwei Shan, and Shen Ma. Isolation and identification of pathogenic Vibrio spp. retrieved from diseased Litopenaeus vannamei and beneficial role of some functional probiotic bacteria for control. Aquaculture International, 28(4):1403– 1420, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00530-3.

Cicala:2020:GMS

[2090] Francesco Cicala, Asunción Lago-Lestón, Bruno Gomez-Gil, Teresa Gollas-Galván, Jennyfers Chong-Robles, Edilmar Cortés-Jacinto, and Marcel Martínez-Porchas. Gut microbiota shifts in the giant tiger shrimp, Penaeus monodon, during the postlarvae, juvenile, and adult stages. Aquaculture International, 28(4):1421–1433, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00532-1.

deOliveira:2020:ETM

[2091] Carlos Yure Barbosa de Oliveira, Thayna Lie Viegas, Maria Fernanda Oliveira da Silva, Debora Machado Fracalossi, Rafael Garcia Lopes, and Roberto Bianchini Derner. Effect of trace metals on growth performance and accumulation of lipids, proteins, and carbohydrates on the green microalga Scenedesmus obliquus. Aquaculture International, 28 (4):1435–1444, August 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00533-0.

Liu:2020:SUD

[2092] Mei Liu, Xiang-Hu Huang, Chang-Ling Li, and Binhe Gu. Study on the uptake of dissolved nitrogen by Oocystis borgei in prawn (Litopenaeus vannamei) aquaculture ponds and establishment of uptake model. Aquaculture International, 28(4):1445–1458, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00534-z.

Aly:2020:TMI

[2093] S. Aly, M. Fathi, E. M. Youssef, and M. Mabrok. Trichodinids and monogeneans infestation among Nile tilapia hatcheries in Egypt: prevalence, therapeutic and prophylactic treatments. Aquaculture International, 28 (4):1459–1471, August 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00537-w.

Esparza-Leal:2020:ESD

[2094] Héctor M. Esparza-Leal, Jesús T. Ponce-Palafox, Píndaro Álvarez-Ruiz, Ely S. López-Álvarez, Nadia Vázquez-Montoya, Mariel López-Espinoza, Magnolia Montoya Mejia, Rosa L. Gómez-Peraza, and Eusebio Nava-Perez. Effect of stocking density and water exchange on performance and stress tolerance to low and high salinity by Litopenaeus vannamei postlarvae reared with biofloc in intensive nursery phase. Aquaculture International, 28(4):1473–1483, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00535-y.

Dey:2020:EBC

[2095] Bipul Kumar Dey, Md. Mer Mosharraf Hossain, and Md. Eftakher Alam. Effect of black cumin seed oil on growth, innate immunity and resistance against Pseudomonas fluorescens infection in Nile tilapia Oreochromis niloticus. Aquaculture International, 28(4):1485–1499, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00539-8.

Yang:2020:CSB

[2096] Sen Yang, Bin Fan, Xinghan Chen, and Zining Meng. Cryopreservation of sperm in brown-marbled grouper (Epinephelus fuscoguttatus). Aquaculture International, 28(4):1501–1516, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00538-9.

Ma:2020:CFS

[2097] Hangke Ma, Huan Gao, Wanyuan Xu, Jinqiu Sun, Jixuan Zhu, Qin Dai, Guangwei Hu, and Binlun Yan. Cloning and functional study of fatty acid-binding protein-like gene of the ridgetail white prawn, Exopalaemon carinicauda. Aquaculture International, 28(4):1517–1530, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00540-1.

Sun:2020:EDF

[2098] Yunfei Sun, Shaicheng Zhu, Xiaowen Long, Xugan Wu, and Yongxu Cheng. Effects of dietary fish meal replacement with protein mixtures on growth performance, biochemical composition, and physiological metabolism of juvenile swimming crab, Portunus trituberculatus. Aquaculture International, 28(4):1531–1545, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00541-0.

Teng:2020:CEG

[2099] Teng Teng, Xuqian Zhao, Caijuan Li, Jinqiang Guo, Yunfeng Wang, Chenglong Pan, Enguang Liu, and Qufei Ling. Cloning and expression of IGF-I, IGF-II, and GHR genes and the role of their single-nucleotide polymorphisms in the growth of pikeperch (Sander lucioperca). Aquaculture International, 28(4):1547–1561, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00542-z.

Silva:2020:ZPB

[2100] Shadai M. Silva, Patricia B. Ramos, Juan R. Buitrago, Thamyres V. N. da Silva, Cleber S. Simião, Grecica M. Colombo, Marcos Schmitz, Marcelo B. Tesser, Carlos Prentice, Wilson Wasielesky, and José M. Monserrat. Zootechnical performance, biochemical response, and chromaticity in Pacific white shrimp (Litopenaeus vannamei) (Boone, 1931) after the inclusion of lyophilized açaí (Euterpe oleracea) in the diet. Aquaculture International, 28(4):1563–1577, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00544-x.

McDougall:2020:VET

[2101] Daniel R. McDougall, Julien Vignier, Norman L. C. Ragg, Bridget Finnie, Andrew Jeffs, and Serean Adams. The value of EDTA treatment of hatchery water to rear Greenshell^{T M} mussel (Perna canaliculus) larvae. Aquaculture International, 28(4):1579–1592, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00543-y.

Singaravel:2020:PDS

[2102] V. Singaravel, A. Gopalakrishnan, N. K. Dewangan, D. Kannan, N. Shettu, and Gary G. Martin. Photobacterium damselae subsp. damselae associated with bacterial myonecrosis and hepatopancreatic necrosis in broodstock Pacific white leg shrimp, Litopenaeus vannamei (Boone, 1931). Aquaculture International, 28(4):1593–1608, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00545-w.

Sales:2020:PCI

[2103] Rafael Sales and Lília Pereira de Souza-Santos. Production of concentrated inocula from the microalgae Nannochloropsis oculata. Aquaculture International, 28(4):1609–1620, August 2020. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00546-9.

Anton-Pardo:2020:EBV

[2104] Maria Anton-Pardo, David Hlaváč, Christian Bauer, and Zdeněk Adámek. Environmental and biotic variables in carp (Cyprinus carpio) ponds: organic vs. conventional management. Aquaculture International, 28(4):1621–1637, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00547-8.

Sarojini:2020:SAA

[2105] Kuttalampillai Sarojini, Sethunathan Soja Ajitha, Ramasamy Ramasubburayan, Arunachalam Palavesam, and Grasian Immanuel. Studies on the adhesion, aggregative properties and the probiotic efficiency of a potent bacteriocin-producing shrimp gut isolate Bacillus subtilis subsp. inaquosorum v_{1} against Carassius auratus. Aquaculture International, 28 (4):1639–1656, August 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00548-7.

Nakayama:2020:DTA

[2106] Cintia Labussière Nakayama, Diogo Lopes, André Braga, Wilson Wasielesky, and Ronaldo O. Cavalli. Does temperature affect the occurrence of penaeid spermatophore melanization related to successive manual extrusions? A case study with the closed thelycum Farfantepenaeus paulensis. Aquaculture International, 28(4):1657–1663, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00549-6.

Fazio:2020:IVB

[2107] Francesco Fazio, Giovanni Lanteri, Concetta Saoca, Carmelo Iaria, Giuseppe Piccione, Tiziana Orefice, Emilia Calabrese, and Irene Vazzana. Individual variability of blood parameters in striped bass Morone saxatilis: possible differences related to weight and length. Aquaculture International, 28(4):1665–1673, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00550-z.

daCosta:2020:HCG

[2108] Fiz da Costa, Ana Cerviño-Otero, Óscar Iglesias, Andreia Cruz, and Eric Guévélou. Hatchery culture of European clam species (family Veneridae). Aquaculture International, 28(4):1675–1708, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00552-x.

Ahmmed:2020:HGD

[2109] Fatema Ahmmed, Mirja Kaizer Ahmmed, Sharmin S. Khushi, Mohammad Saifuddin Sumon, Sree Soundarya Karamcheti, and Md. Golam Sarower. Host gut-derived probiotic Lactobacillus sp. improves resistance of giant freshwater prawn Macrobrachium rosenbergii against Vibrio harveyi. Aquaculture International, 28(4):1709–1724, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00551-y.

Howlader:2020:AAC

[2110] Pallabi Howlader, Alokesh Kumar Ghosh, Shikder Saiful Islam, Joyanta Bir, and Ghausiatur Reza Banu. Antiviral activity of Cynodon dactylon on white spot syndrome virus (WSSV)-infected shrimp: an attempt to mitigate risk in shrimp farming. Aquaculture International, 28(4): 1725–1738, August 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00553-w.

Hayat:2020:LAG

[2111] M. Hayat, M. Y. Sabri, A. R. Intan-Shameha, M. Y. Ina-Salwany, and Kim D. Thompson. Localisation of antigens in the gut postchallenge with Streptococcus iniae in vaccinated and non-vaccinated red hybrid tilapia (Oreochromis sp.). Aquaculture International, 28(4): 1739–1752, August 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00554-9.

vanderMeeren:2020:TGA

[2112] Terje van der Meeren and Anders Mangor-Jensen. Tolerance of Atlantic cod (Gadus morhua L.) larvae to acute ammonia exposure. Aquaculture International, 28(4):1753–1769, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00555-8.

Zhou:2020:FLA

[2113] Qun-Lan Zhou, Su-Fei Jiang, Yi-Wei Xiong, Bo Liu, Cunxin Sun, Zhenting Jiang, and Hongtuo Fu. Fishmeal level affects growth performance of Macrobrachium nipponense via regulating protein and lipid metabolism. Aquaculture International, 28(4):1771–1785, August 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00556-7.

Cui:2020:EAP

[2114] Qingman Cui, Ziyue Zhao, and Chunying Yuan. Effects of astragalus polysaccharides on hemocytes phagocytosis and gene expression of immune-related factors in Eriocheir sinensis. Aquaculture International, 28(5):1787–1796, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00557-6.

Zhang:2020:EDL

[2115] Hai peng Zhang, Wen long Dong, Long Chen, Yi ming Wang, Inam Muhammad, An qi Ju, Xiao feng Shan, Hong xia Ma, and Ling cong Kong. Effects of dietary Lactobacillus plantarum C20015 on growth, immunity, and disease resistance in koi carp. Aquaculture International, 28 (5):1797–1809, October 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00558-5.

Taha:2020:EVN

[2116] Engy Taha, Mohamed Shawky, Basem Ahmed, Mohamed Moustafa, Ausama Yousif, and Mohamed Abdelaziz. Emergence of viral nervous necrosis is associated with mass mortality in hatchery-reared tilapia (Oreochromis niloticus) in Egypt. Aquaculture International, 28(5): 1811–1823, October 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00559-4.

Seixas:2020:ATL

[2117] Sónia Seixas and Margaret Eleftheriou. AQUA-TNET lifelong learning strategy changes E & T attitudes towards ECVET implementation and the EQF. Aquaculture International, 28(5):1825–1836, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00560-x.

LimaFerreira:2020:BMA

[2118] Andre Lima Ferreira, Walisson de Souza e Silva, Luanna do Carmo Neves, Nathália Soares Ferreira, Rodrigo Takata, and Ronald Kennedy Luz. Benzocaine and menthol as anesthetics for the African cichlid Aulonocara nyassae. Aquaculture International, 28(5):1837–1846, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00561-w.

Song:2020:CWS

[2119] Mengsi Song, Luqing Pan, Mengyu Zhang, Fei Huang, Shuo Gao, and Changcheng Tian. Changes of water, sediment, and intestinal bacterial communities in Penaeus japonicus cultivation and their impacts on shrimp physiological health. Aquaculture International, 28(5):1847– 1865, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00562-9.

Zhou:2020:DMM

[2120] Ying Zhou, Jingou Tong, Junru Wang, and Xiaomu Yu. Development of microsatellite markers and genetic diversity in wild and cultured populations of black carp (Mylopharyngodon piceus) along the Yangtze River. Aquaculture International, 28(5):1867–1882, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00563-8.

Minabi:2020:EDC

[2121] Khalil Minabi, Iman Sourinejad, Morteza Alizadeh, Ebrahim Rajabzadeh Ghatrami, and Mohammad Hossein Khanjani. Effects of different carbon to nitrogen ratios in the biofloc system on water quality, growth, and body composition of common carp (Cyprinus carpio L.) fingerlings. Aquaculture International, 28(5):1883–1898, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00564-7.

Battazza:2020:ICS

[2122] Alexandre Battazza, Felipe César da Silva Brasileiro, Eduardo Ferreira Machado, Magna Gomes de Matos, Cauê Bastos Tertuliano dos Santos, Marianna Vaz Rodrigues, Diana do Nascimento, and Noeme Sousa Rocha. Identification and characterization of Sinuolinea niloticus from Nile tilapia (Oreochromis niloticus) farmed in Botucatu, Brazil. Aquaculture International, 28(5):1899–1906, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00565-6.

Rodrigues:2020:HBH

[2123] Robson Andrade Rodrigues, André Luiz do Nascimento Silva, Mayara Schueroff Siqueira, Fabiana Pilarski, Cassia Rejane Brito Leal, Karin Virgínia Kuibida, Cristiane Meldau de Campos, and Carlos Eurico Fernandes. Hematological, biochemical, and histopathological responses in sorubim Pseudoplatystoma spp. experimentally infected with Lactococcus garvieae. Aquaculture International, 28(5):1907–1923, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00566-5.

Ninawe:2020:BAD

[2124] A. S. Ninawe, S. Sivasankari, P. Ramasamy, G. Seghal Kiran, and Joseph Selvin. Bacteriophages for aquaculture disease control. Aquaculture International, 28(5):1925–1938, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00567-4.

Arulmoorthy:2020:MVD

[2125] M. P. Arulmoorthy, E. Anandajothi, S. Vasudevan, and E. Suresh. Major viral diseases in culturable penaeid shrimps: a review. Aquaculture International, 28(5):1939–1967, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00568-3.

Guarizo:2020:EDD

[2126] Matheus Guarizo, Tânia Marcia Costa, and Murilo Zanetti Marochi. Effect of diet during larval development of Menippe nodifrons Stimpson, 1859 and Callinectes danae Smith, 1869. Aquaculture International, 28 (5):1969–1980, October 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00569-2.

Pontin:2020:PEN

[2127] Mariana C. F. Pontin, Wiolene M. Nordi, Jéssica Pampolini, Raul Machado-Neto, and Débora B. Moretti. Protective effect of nutraceutical food on the intestinal mucosa of juvenile pacu Piaractus mesopotamicus under high stocking density. Aquaculture International, 28(5): 1981–1995, October 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00570-9.

Durmus:2020:IOO

[2128] Mustafa Durmus, Abdelkader Bensid, Yeşim Ozogul, Ali Riza Kosker, Yilmaz Ucar, Esmeray Kuley Boga, Murat Durmus, and Fatih Ozogul. Influence of olive oil–based nanoemulsion on the fatty acid profiles of rainbow trout fillets. Aquaculture International, 28(5):1997–2014, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00574-5.

Rehman:2020:CFC

[2129] Tayyab Rehman, Lizi Yin, Muhammad Bilal Latif, Yaojia Zhou, Kaiyu Wang, Yi Geng, Xiaoli Huang, Defang Chen, Jing Fang, Zhengli Chen, Hongrui Guo, Weiming Lai, and Ping Ouyang. Current findings on carp edema virus, control challenges, and future outlook. Aquaculture International, 28(5):2015–2026, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00573-6.

Zhu:2020:SSO

[2130] Wenliang Zhu, Peng Tan, Bao Lou, Ruiyi Chen, Ligai Wang, and Dongdong Xu. Supplementation of a soybean oil-based diet with tributyrin influences growth, muscle composition, intestinal morphology, and expression of immune-related genes of juvenile yellow drum (Nibea albiflora Richardson, 1846). Aquaculture International, 28(5):2027–2043, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00572-7.

Angkha:2020:MMB

[2131] B. Angkha, A. K. Verma, Sanath H. Kumar, Chandra Prakash, and R. M. Thomas. Mobilization of mica by Bacillus sp. and its effect on Nile tilapia (Oreochromis niloticus) cum holy basil (Ocimum tenuiflorum)–based aquaponic system. Aquaculture International, 28(5):2045– 2058, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00575-4.

Gao:2020:SOE

[2132] Yaping Gao, Jianguang Fang, Fan Lin, Fengxue Li, Wenhao Li, Xiaoqin Wang, Meirong Du, Jinghui Fang, Weiwei Jiang, Hui Liu, Yitao Zhang, Junwei Wang, and Zengjie Jiang. Simulation of oyster ecological carrying capacity in Sanggou Bay in the ecosystem context. Aquaculture International, 28(5):2059–2079, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00576-3.

Eissa:2020:CFG

[2133] Alaa Eldin Eissa, Mohamed Abdelsalam, Mahmoud A. Mahmoud, Nehal A. Younis, Abdelsalam A. Abu Mhara, and Rabia A. El Zlitne. Cutaneous fibropapilloma in Egyptian-farmed gilthead seabream (Sparus aurata; Linnaeus, 1758). Aquaculture International, 28(5):2081–2091, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00579-0.

Harohau:2020:UAT

[2134] Daykin Harohau, Jessica Blythe, Marcus Sheaves, and Amy Diedrich. Uneven adoption of tilapia aquaculture in rural Solomon Islands. Aquaculture International, 28(5):2093–2109, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00577-2.

Barbosa:2020:EAF

[2135] Andressa Steffen Barbosa, Raquel Grande Pereira, Laurindo André Rodrigues, Jorge de Matos Casaca, Wagner Cotroni Valenti, and Thiago El Hadi Perez Fabregat. Economic analysis of family trout farming in Southern Brazil. Aquaculture International, 28(5):2111–2120, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00580-7.

Lam:2020:DNS

[2136] Chelsea Lam, Isnita Khairunissa, Lily Damayanti, Tomofumi Kurobe, Swee J. Teh, Heather Pfahl, Saipul Rapi, Noel Janetski, and Dolores V. Baxa. Detection of a new strain of lymphocystis disease virus (LCDV) in captive-bred clownfish Amphiprion percula in South Sulawesi, Indonesia. Aquaculture International, 28(5):2121–2137, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00581-6.

Banerjee:2020:RTN

[2137] Rajanya Banerjee, Mrinal Samanta, and Surajit Das. Regulation of tumor necrosis superfamily molecules, BAFF and APRIL, in response to pathogenic exposure and ligand stimulation in freshwater carp, Catla catla. Aquaculture International, 28(5):2139–2159, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00582-5.

Chakrobortty:2020:VCT

[2138] Dipankar Chakrobortty, Md. Rayhan Ali, Bipul Kumar Dey, Nipa Gupta, Shikder Saiful Islam, and Liying Sui. Viral contamination of tiger shrimp Penaeus monodon broodstock in Bangladesh. Aquaculture International, 28(5):2161–2172, October 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-020-00583-4.

Pasookhush:2020:MIC

[2139] Phongthana Pasookhush, Akapon Vaniksampanna, Paisarn Sithigorngul, Siwaporn Longyant, and Parin Chaivisuthangkura. Molecular isolation and characterization of translationally controlled tumor protein (TCTP) gene from Macrobrachium rosenbergii. Aquaculture International, 28(6): 2173–2190, December 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00571-8.

Haghparast:2020:EHI

[2140] Mohammad Mahdi Haghparast, Mojtaba Alishahi, Masoud Ghorbanpour, and Ali Shahriari. Evaluation of hemato-immunological parameters and stress indicators of common carp (Cyprinus carpio) in different C/ N ratio of biofloc system. Aquaculture International, 28(6): 2191–2206, December 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00578-1.

Rahman:2020:AII

[2141] Md. Sadique Rahman, Mohammad Mizanul Haque Kazal, and Shah Johir Rayhan. Adoption and impacts of improved mud crab fattening practices on the productivity and wellbeing of coastal farmers in Bangladesh. Aquaculture International, 28(6):2207–2219, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00584-3.

Amoah:2020:CSH

[2142] Kwaku Amoah, Patricia Adu-Asiamah, Xiao hui Dong, Akwasi Ampofo-Yeboah, and Emmanuel D. Abarike. A comparative study on the hatchability and survival rate of African catfish, Clarias gariepinus (Burchell, 1822), induced with catfish’s pituitary gland hormone from farmed and wild sources. Aquaculture International, 28(6):2221–2234, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00591-4.

Zhang:2020:AES

[2143] Lin Zhang, Shuping Yang, Jilin Xu, Tong Liu, Dongjie Yang, Zuyao Wu, and Mengjie Shao. Application of exogenous salicylic acid on improving high temperature resistance of Nannochloropsis oceanica. Aquaculture International, 28(6):2235–2246, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00592-3.

Younis:2020:ISH

[2144] Nehal A. Younis, Samah E. Laban, Asmaa K. Al-Mokaddem, and Marwa M. Attia. Immunological status and histopathological appraisal of farmed Oreochromis niloticus exposed to parasitic infections and heavy metal toxicity. Aquaculture International, 28(6):2247–2262, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00589-y.

Workagegn:2020:GPG

[2145] Kassaye Balkew Workagegn, P. Natarajan, and Andargachew Gedebo. Genetic parameters and genotype by environment interaction of the Nile tilapia (Oreochromis niloticus) reared in two test environments. Aquaculture International, 28(6):2263–2273, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00593-2.

Yu:2020:PMQ

[2146] Chengchen Yu, Xiaoyan Xu, Jiale Li, and Yubang Shen. Primary mapping of QTL for growth-related traits in grass carp (Ctenopharyngodon idellus). Aquaculture International, 28(6):2275–2285, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00594-1.

Chen:2020:DLC

[2147] Shi-Jun Chen, Peng Liu, You-Ming Jia, Hong-Ping Liao, Shao-Xuan Zhu, Li-Liu Zhou, Xue-Ming Dan, Li Liu, Jia-Hao Li, Su-Bi Zheng, Jun-Jiang Yang, and Lian Gan. Dietary lipid concentrations influence growth, body composition, morphology of the liver and mid-intestine, and antioxidant status of marbled eel (Anguilla marmorata). Aquaculture International, 28(6):2287–2302, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00590-5.

Garcia-Medel:2020:BLB

[2148] Darío Israel García-Medel, Carlos Angulo, Ruth Escamilla-Montes, Jesús Arturo Fierro-Coronado, Genaro Diarte-Plata, Carina Gámez-Jiménez, and Antonio Luna-González. Bacillus licheniformis BCR 4-3 increases immune response and survival of Litopenaeus vannamei challenged with Vibrio parahaemolyticus IPNGS16. Aquaculture International, 28(6):2303–2318, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00585-2.

Xiong:2020:VCT

[2149] Jing Xiong, Shan gong Wu, Ying Liang, Yue lian Zou, Xiao mei Xie, Wen shu Huang, and Yi bing Zhang. Visual closed-tube loopmediated isothermal amplification (LAMP) assay targeting aerolysin gene: a practical screening method for virulent Aeromonas species affecting cultured eels in China. Aquaculture International, 28(6):2319– 2332, December 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00588-z.

Raguraman:2020:SPP

[2150] Vasantharaja Raguraman, Namitha Ravindran, Koushalya Selvaraju, and Govindaraju Kasivelu. Seaweed polysaccharides as potential therapeutic agents against white spot syndrome virus (WSSV): a mini review. Aquaculture International, 28(6):2333–2343, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00587-0.

Nunes:2020:EBS

[2151] André Luiz Nunes, Marco Shizuo Owatari, Robson Andrade Rodrigues, Letícia Emiliani Fantini, Rodrigo Yutaka Dichoff Kasai, Maurício Laterça Martins, José Luiz Pedreira Mouriño, and Cristiane Meldau de Campos. Effects of Bacillus subtilis C-3102-supplemented diet on growth, non-specific immunity, intestinal morphometry and resistance of hybrid juvenile Pseudoplatystoma sp. challenged with Aeromonas hydrophila. Aquaculture International, 28(6):2345–2361, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00586-1.

Younus:2020:CED

[2152] Naima Younus, Amina Zuberi, Tariq Mahmoood, Wasim Akram, and Muhammad Ahmad. Comparative effects of dietary micro- and nanoscale chitosan on the growth performance, non-specific immunity, and resistance of silver carp Hypophthalmichthys molitrix against Staphylococcus aureus infection. Aquaculture International, 28(6):2363–2378, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00595-0.

Meneses:2020:ATH

[2153] Juliana Oliveira Meneses, Fernanda dos Santos Cunha, Joel Artur Rodrigues Dias, Ana Flávia Santos da Cunha, Francisco José dos Santos, Natalino da Costa Sousa, Márcia Valéria Silva do Couto, Peterson Emmanuel Guimarães Paixão, Higo Andrade Abe, Bruno dos Santos Lima, Antônio Guilherme de Carvalho Neto, Adriano Antunes de Souza Araújo, Luiz Pereira da Costa, Juliana Cordeiro Cardoso, and Rodrigo Yudi Fujimoto. Acute toxicity of hot aqueous extract from leaves of the Terminalia catappa in juvenile fish Colossoma macropomum. Aquaculture International, 28(6):2379–2396, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00596-z.

Fox:2020:PMF

[2154] M. Fox, R. Christley, C. Lupo, H. Moore, M. Service, and K. Campbell. Preventing and mitigating farmed bivalve disease: a Northern Ireland case study. Aquaculture International, 28(6):2397–2417, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00597-y.

Chahinez:2020:SSF

[2155] Laama Chahinez, Hassani Abderrahim, and Bachari Nour El Islem. Site selection for finfish cage farming using spatial multi-criteria evaluation and their validation at field in the Bay of Souahlia (Algeria). Aquaculture International, 28(6):2419–2436, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00598-x.

Lulijwa:2020:PPA

[2156] Ronald Lulijwa, Andrea C. Alfaro, Fabrice Merien, Mark Burdass, Jill Meyer, Leonie Venter, and Tim Young. Polyinosinic:polycytidylic acid in vivo enhances Chinook salmon (Oncorhynchus tshawytscha) immunity and alters the fish metabolome. Aquaculture International, 28(6): 2437–2463, December 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00599-w.

Przybyla:2020:GES

[2157] C. Przybyla, G. Dutto, M. Bernard, B. Rollin, X. Laurand, J. Averseng, E. Gasset, I. Tadadjeu, and L. Dusseau. European sea bass (Dicentrarchus labrax) and meagre (Argyrosomus regius) fertilized egg resistance to a spacecraft launcher vibration qualifying test. Aquaculture International, 28(6):2465–2479, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00601-5.

Goddek:2020:INW

[2158] Simon Goddek and Karel J. Keesman. Improving nutrient and water use efficiencies in multi-loop aquaponics systems. Aquaculture International, 28(6):2481–2490, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00600-6.

Mahdy:2020:MCH

[2159] Olfat A. Mahdy, Mahmoud A. Mahmoud, and Mohamed Abdelsalam. Morphological characterization and histopathological alterations of homologs heterophyid metacercarial coinfection in farmed mullets and experimental infected pigeons. Aquaculture International, 28(6):2491– 2504, December 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00602-4.

Suljevic:2020:FRB

[2160] Damir Suljević and Maja Mitrašinović-Brulić. The first record of brook trout (Salvelinus fontinalis, salmonidae) blood cell characteristics and hematological profile: the influence of fish sex on leukocyte count. Aquaculture International, 28(6):2505–2516, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00603-3.

Ghazy:2020:CPA

[2161] Haneen A. Ghazy, Abeer F. El-Nahas, Shawky A. Mahmoud, Hanan A. Fahmy, Ramadan A. El-Domany, Hoda E. Mahmoud, and Amira A. Omar. Characterization of Pseudomonas aeruginosa ghost and evaluation of its immune proficiency in Nile tilapia (Oreochromis niloticus). Aquaculture International, 28(6):2517–2529, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00604-2.

Hossain:2020:DGV

[2162] Md. Mer Mosharraf Hossain, Md. Imtiaz Uddin, Habiba Islam, Jannatul Fardoush, Md. Ariful Haque Rupom, Md. Monjur Hossain, Nawshin Farjana, Rukaiya Afroz, Hasan-Uj-Jaman, Hironmoy Shovon Roy, Md. Asif Shahriar Shehab, and Md. Anisur Rahman. Diagnosis, genetic variations, virulence, and toxicity of AHPND-positive Vibrio parahaemolyticus in Penaeus monodon. Aquaculture International, 28(6): 2531–2546, December 2020. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00607-z.

Yu:2020:DAC

[2163] Lean Huat Yu, Cindy Shuan Ju Teh, Kien Pong Yap, and Kwai Lin Thong. Diagnostic approaches and contribution of next-generation sequencing technologies in genomic investigation of Vibrio parahaemolyticus that caused acute hepatopancreatic necrosis disease (AHPND). Aquaculture International, 28(6):2547–2559, December 2020. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00610-4.

Wang:2021:CPS

[2164] Ya-Yu Wang, Sheng-Hua Duan, Sai-Sai Dong, Xiao-Yu Cui, Gui-Ling Wang, and Jia-Le Li. Comparative proteomic study on fem-1b in female and male gonads in Hyriopsis cumingii. Aquaculture International, 29 (1):1–18, February 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00605-1.

Lima:2021:HAF

[2165] Adriano Lima, Bárbara Marinho, and Tiago Morais. Hydrodynamic analysis of flow in raceway ponds for algae cultivation under versatile conditions. Aquaculture International, 29(1):19–35, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00606-0.

Pinho:2021:IPG

[2166] Sara M. Pinho, Luiz Henrique C. David, Simon Goddek, Maurício G. C. Emerenciano, and Maria Célia Portella. Integrated production of Nile tilapia juveniles and lettuce using biofloc technology. Aquaculture International, 29(1):37–56, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00608-y.

Zhu:2021:OAT

[2167] Ze Zhu, Uri Yogev, Karel J. Keesman, and Amit Gross. Onsite anaerobic treatment of aquaponics lettuce waste: digestion efficiency and nutrient recovery. Aquaculture International, 29(1):57–73, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00609-x.

Xie:2021:CEA

[2168] Rui tao Xie, Eric Amenyogbe, Wei zheng Wang, Zhi xiong Guo, Gang Chen, and Jian sheng Huang. Cloning and expression analysis of hypoxia-related gene HO in cobia. Aquaculture International, 29(1): 75–89, February 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00611-3.

Guimaraes:2021:OAB

[2169] Mateus Cardoso Guimarães, Amanda Iracy Cavalcante da Silva Guimarães, Mariene Miyoko Natori, Miguel Frederico Fernandez Alarcon, Danielle de Carla Dias, Carlos Massatoshi Ishikawa, Silvana Tapia-Paniagua, Miguel Ángel Moriñigo, Francisco Javier Moyano, and Leonardo Tachibana. Oral administration of Bacillus subtilis and Lactobacillus plantarum modulates the gut microbiota and increases the amylase activity of Nile tilapia (Oreochromis niloticus). Aquaculture International, 29(1):91–104, February 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00612-2.

El-Tohamy:2021:SAS

[2170] Wael El-Tohamy, Jianguang Qin, Nagwa Abdel-Aziz, Ahmed El-Ghobashy, and Mohamed Dorgham. Suitable algal species and density for the culture of copepod Gladioferens imparipes as a potential live food for fish larvae. Aquaculture International, 29(1):105–125, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00614-0.

Kim:2021:AAM

[2171] Kwang Il Kim, Eun Young Min, Tae-Ho Kim, Hye Sung Choi, and Hyun-Ja Han. Application of alginate microparticles incorporating formalininactivated Streptococcus parauberis for oral vaccination in olive flounder. Aquaculture International, 29(1):127–138, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00613-1.

Jimenez-Prada:2021:AWF

[2172] P. Jiménez-Prada, I. Hachero-Cruzado, and J. M. Guerra-García. Aquaculture waste as food for amphipods: the case of Gammarus insensibilis in marsh ponds from southern Spain. Aquaculture International, 29(1): 139–153, February 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00615-z.

Akbary:2021:SFA

[2173] Paria Akbary, Lawrence M. Liao, Zahra Aminikhoei, Kamran Rezaeie Tavabe, Mehrdad Hobbi, and Elahe Erfanifar. Sterol and fatty acid profiles of three macroalgal species collected from the Chabahar coasts, southeastern Iran. Aquaculture International, 29(1):155–165, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00616-y.

Xu:2021:SAA

[2174] Min Xu, Xin yuan Yang, Xiao jing Song, Kai da Xu, and Lin lin Yang. Seasonal analysis of artificial oyster reef ecosystems: implications for sustainable fisheries management. Aquaculture International, 29(1): 167–192, February 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00617-x.

Oztaskent:2021:ELL

[2175] Cenk Öztaşkent and İlknur Ak. Effect of LED light sources on the growth and chemical composition of brown seaweed Treptacantha barbata. Aquaculture International, 29(1):193–205, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-020-00619-9.

dosSantos:2021:FTC

[2176] Driely Kathriny Monteiro dos Santos, Juliana Tomomi Kojima, Thiago Macedo Santana, Diogo Pereira de Castro, Paula Taquita Serra, Naiara Silva Menezes Dantas, Flávio Augusto Leão da Fonseca, Luís André Morais Mariúba, and Ligia Uribe Gonçalves. Farming tambaqui (Colossoma macropomum) in static clear water versus a biofloc system with or without Bacillus subtilis supplementation. Aquaculture International, 29(1):207–218, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00618-w.

Naiel:2021:ULA

[2177] Mohammed A. E. Naiel, Mayada R. Farag, Ahmed G. A. Gewida, Mahmoud A. Elnakeeb, Mahmoud S. Amer, and Mahmoud Alagawany. Using lactic acid bacteria as an immunostimulants in cultured shrimp with special reference to Lactobacillus spp. Aquaculture International, 29(1): 219–231, February 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00620-2.

Sarkar:2021:BEI

[2178] Soumyabrata Sarkar, P. Nila Rekha, K. Ambasankar, and K. K. Vijayan. Bioremediation efficiency of indigenous seaweeds of Chennai coast in brackishwater system. Aquaculture International, 29(1):233– 251, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00621-1.

Wang:2021:DRP

[2179] Hongquan Wang, Hong Yang, Rui Li, Yaoguo Li, and Tiaoyi Xiao. Different resistance potential to reovirus in grass carp (Ctenopharyngodon idella) populations and their immune characteristics. Aquaculture International, 29(1):253–260, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00623-z.

Sangari:2021:GBC

[2180] Mohammad Sangari, Ebrahim Sotoudeh, Dara Bagheri, Salim Morammazi, and Mansour Torfi Mozanzadeh. Growth, body composition, and hematology of yellowfin seabream (Acanthopagrus latus) given feeds supplemented with organic acid salts (sodium acetate and sodium propionate). Aquaculture International, 29(1):261–273, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00625-x.

Sari:2021:EHI

[2181] Diana Purnama Sari, Sukenda Sukenda, Munti Yuhana, and Sri Nuryati. Effect of the hyperosmotic infiltration method on immune response in tilapia vaccinated with Streptococcus agalactiae. Aquaculture International, 29(1):275–288, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00624-y.

delPazo:2021:GDS

[2182] Felipe del Pazo, Sebastián Sánchez, Victoria Posner, Andrés A. Sciara, Silvia E. Arranz, and Gabriela V. Villanova. Genetic diversity and structure of the commercially important native fish pacu (Piaractus mesopotamicus) from cultured and wild fish populations: relevance for broodstock management. Aquaculture International, 29(1):289–305, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00626-w.

Khanjani:2021:EDC

[2183] Mohammad Hossein Khanjani, Morteza Alizadeh, and Moslem Sharifinia. Effects of different carbon sources on water quality, biofloc quality, and growth performance of Nile tilapia (Oreochromis niloticus) fingerlings in a heterotrophic culture system. Aquaculture International, 29 (1):307–321, February 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00627-9.

James:2021:BAF

[2184] Greeshma James, Bini C. Das, Seena Jose, and Rejish Kumar V. J. Bacillus as an aquaculture friendly microbe. Aquaculture International, 29 (1):323–353, February 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00630-0.

Notaro:2021:PSI

[2185] Dean Anthony Notaro, Sarah C. Culloty, and Sharon A. Lynch. A pilot study investigating the potential of antimicrobial photodynamic therapy (aPDT) to control Vibrio spp. development in microalgae and seawater. Aquaculture International, 29(1):355–372, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00631-z.

Zhao:2021:EAW

[2186] Qiong Zhao, Ke Zhou, Fengfeng Zhang, Yujie Zhao, Haibo Sun, and Fengxing Xie. Effect of aquaculture water eutrophication on color development in Biolog EcoPlates. Aquaculture International, 29(1):373– 386, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00632-y.

Ding:2021:LST

[2187] Hongchang Ding, Linbin Huang, and Xinghong Yan. Low-salt tolerance of the thin-blade strain in Pyropia haitanensis. Aquaculture International, 29(1):387–397, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00637-7.

Soares:2021:GPA

[2188] Roberta Soares, Silvio Peixoto, Harsha S. C. Galkanda-Arachchige, and D. Allen Davis. Growth performance and acoustic feeding behavior of two size classes of Litopenaeus vannamei fed pelleted and extruded diets. Aquaculture International, 29(1):399–415, February 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00636-8.

Satpathy:2021:FBP

[2189] Ark Ansuman Satpathy, Supriya Dash, Swagat Kumar Das, Shyamalini Shyamasuta, and Suvendu Pradhan. Functional and bioactive properties of chitosan from Indian major carp scale. Aquaculture International, 29 (2):417–430, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00622-0.

Buwono:2021:GHT

[2190] Ibnu Dwi Buwono, Iskandar Iskandar, and Roffi Grandiosa. Growth hormone transgenesis and feed composition influence growth and protein and amino acid content in transgenic g_{3} mutiara catfish (Clarias gariepinus). Aquaculture International, 29(2):431–451, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00628-8.

Smith:2021:UMC

[2191] J. K. Penhaul Smith, C. Beveridge, V. A. Laudicella, A. D. Hughes, L. McEvoy, and J. G. Day. Utilising mixotrophically cultured “designer algae” as blue mussel larval feed. Aquaculture International, 29 (2):453–475, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00629-7.

Abidin:2021:DTS

[2192] Aisamuddin Ardi Zainal Abidin, Nur Adila Othman, Fatimah Md. Yusoff, and Zetty Norhana Balia Yusof. Determination of transgene stability in Nannochloropsis sp. transformed with immunogenic peptide for oral vaccination against vibriosis. Aquaculture International, 29 (2):477–486, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00634-w.

Lahteenmaki-Uutela:2021:GEU

[2193] Anu Lähteenmäki-Uutela, Moona Rahikainen, María Teresa Camarena-Gómez, Jonna Piiparinen, Kristian Spilling, and Baoru Yang. European Union legislation on macroalgae products. Aquaculture International, 29 (2):487–509, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00633-x.

Rao:2021:MIT

[2194] Madhusudhana Rao, Sanath H. Kumar, Saurav Kumar, Megha Kadam Bedekar, Gayatri Tripathi, and Rajendran Kooloth Valappil. Microbiological investigation of tilapia lake virus–associated mortalities in cagefarmed Oreochromis niloticus in India. Aquaculture International, 29 (2):511–526, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00635-9.

Rungrueng:2021:ASI

[2195] Naruporn Rungrueng, Watcharachai Meemetta, Kornsunee Phiwsaiya, Ha Thanh Dong, Wattana Panphut, and Saengchan Senapin. Ammonium sulfate improves sensitivity and avoids false negatives of polymerase chain reaction (PCR) for scale drop disease virus (SDDV) detection. Aquaculture International, 29(2):527–538, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-020-00639-5.

Panana:2021:ATA

[2196] Edson Panana, Boris Delaide, Stefan Teerlinck, and Peter Bleyaert. Aerobic treatment and acidification of pikeperch (Sander lucioperca L.) sludge for nutrient recovery. Aquaculture International, 29(2):539– 552, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-020-00640-y.

Kaplan:2021:EDP

[2197] Çağlayan Kaplan and Fatime Erdoğan. Effect of dietary propolis on growth, body composition, and serum biochemistry of juvenile sea bream (Sparus aurata). Aquaculture International, 29(2):553–563, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-020-00642-w.

Haji:2021:CHS

[2198] Bereket Haji and Kassaye Balkew Workagegn. Constraints hindering small scale aquaculture production in southern Ethiopia. Aquaculture International, 29(2):565–574, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-020-00641-x.

AftabUddin:2021:PCM

[2199] Sheikh AftabUddin, M. Gulam Hussain, Mamun Abdullah Al, Pierre Failler, and Benjamin M. Drakeford. On the potential and constraints of mariculture development in Bangladesh. Aquaculture International, 29(2):575–593, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00643-9.

Nasrullah:2021:EIG

[2200] Hasan Nasrullah, Dwi Hany Yanti, Nurly Faridah, Dian Hardiantho, Yanti Inneke Nababan, Sukenda Sukenda, and Alimuddin Alimuddin. Early immune gene development and expression in African catfish Clarias gariepinus after challenged with Aeromonas hydrophila. Aquaculture International, 29(2):595–607, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00645-1.

deSouzaeSilva:2021:ESD

[2201] Walisson de Souza e Silva, Andre Lima Ferreira, Luanna do Carmo Neves, Nathália Soares Ferreira, Glauber David Almeida Palheta, Rodrigo Takata, and Ronald Kennedy Luz. Effects of stocking density on survival, growth and stress resistance of juvenile tambaqui (Colossoma macropomum) reared in a recirculating aquaculture system (RAS). Aquaculture International, 29(2):609–621, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00647-z.

Jannathulla:2021:MPT

[2202] Rajabdeen Jannathulla, Oduru Sravanthi, Syed Moomeen, Gopalapillai Gopikrishna, and Jagabattula Syama Dayal. Microbial products in terms of isolates, whole-cell biomass, and live organisms as aquafeed ingredients: production, nutritional values, and market potential — a review.

Aquaculture International, 29(2):623–650, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00644-2.

Mani:2021:ENA

[2203] Sundaram Mani, Sivakumar Mullaivanam Ramasamy, Saranya Chakrapani, Ambiganandam Krishna, Shyne Anand P. S, Christina Lalramchhani, Jose Antony, and Akshaya Panigrahi. The effect of natural and artificial periphytic substrates with biofloc system on shrimp Penaeus vannamei (Boone 1931) culture: growth and immune response. Aquaculture International, 29(2):651–668, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00646-0.

ZeinEddine:2021:PAR

[2204] Razan ZeinEddine, Bridget Ireland, Samer Monzer, and I. Patrick Saoud. Preliminary assessment of restaurant food waste as a feed ingredient for small juvenile rainbow trout (Oncorhynchus mykiss). Aquaculture International, 29(2):669–679, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00650-4.

Lemos:2021:PDI

[2205] Daniel Lemos, Rafael Coelho, Sjo Zwart, and Albert G. J. Tacon. Performance and digestibility of inorganic phosphates in diets for juvenile shrimp (Litopenaeus vannamei): dicalcium phosphate, monocalcium phosphate, and monoammonium phosphate. Aquaculture International, 29(2):681–695, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00651-3.

Jithendran:2021:SOC

[2206] K. P. Jithendran, A. Navaneeth Krishnan, P. T. Aneesh, P. Ezhil Praveena, and T. Bhuvaneswari. Susceptibility of orange chromide, Etroplus maculatus (Bloch, 1795) to experimental infection of Betanodavirus. Aquaculture International, 29(2):697–710, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00648-y.

Ashaf-Ud-Doulah:2021:IWT

[2207] Mohammad Ashaf-Ud-Doulah, S. M. Majharul Islam, Md Mahiuddin Zahangir, Md Sadiqul Islam, Christopher Brown, and Md Shahjahan.

Increased water temperature interrupts embryonic and larval development of Indian major carp rohu Labeo rohita. Aquaculture International, 29(2):711–722, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00649-x.

deAlmeida:2021:BAP

[2208] Marcos Souza de Almeida, Juliana Rosa Carrijo-Mauad, Régio Marcio Toesca Gimenes, Carlos Augusto Prata Gaona, Plínio Schmidt Furtado, Luís Henrique Poersch, Wilson Wasielesky, and Geraldo Kipper Fóes. Bioeconomic analysis of the production of marine shrimp in greenhouses using the biofloc technology system. Aquaculture International, 29(2):723–741, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00653-1.

Lopez-Cervantes:2021:TSM

[2209] Gabriela López-Cervantes, Píndaro Álvarez-Ruiz, Silvia Luna-Suárez, Antonio Luna-González, Héctor M. Esparza-Leal, Claudia Castro-Martínez, Carina Gámez-Jiménez, and Jorge Soto-Alcalá. Temperature and salinity modulate virulence and PirA gene expression of Vibrio parahaemolyticus, the causative agent of AHPND. Aquaculture International, 29(2):743–756, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00654-0.

Marinho:2021:EID

[2210] Yllana Ferreira Marinho, Carolina Barbosa Malafaia, Katarynna Santos de Araújo, Túlio Diego da Silva, Ana Paula Felipe dos Santos, Laenne Barbara de Moraes, and Alfredo Olivera Gálvez. Evaluation of the influence of different culture media on growth, life cycle, biochemical composition, and astaxanthin production in Haematococcus pluvialis. Aquaculture International, 29(2):757–778, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00655-z.

deOrtiz:2021:FRP

[2211] Danielle Ortiz de Ortiz, Ivan Luiz Gavioli, José Guilherme Filho Bersano, and Erica Alves Gonzalez Vidal. Feeding rates and prey preference in Octopus americanus paralarvae fed with different prey densities and types, Artemia, copepods, and zoeae. Aquaculture International, 29 (2):779–800, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00657-x.

Lebeda:2021:GHC

[2212] Ievgen Lebeda, Marek Rodina, David Gela, Sidika Sakali, Sahana Shivaramu, and Martin Flajšhans. Gonadal histology and concentration of 11-ketotesterone of meiotic gynogens confirm female heterogametic sex determination in sterlet (Acipenser ruthenus). Aquaculture International, 29(2):801–811, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00658-w.

Osei:2021:ESB

[2213] Isaac Kofi Osei, Kobina Yankson, and Edward Adzesiwor Obodai. Effect of sedentary biofoulers on the growth and survival of cultured oysters (Crassostrea tulipa) towards its mass culture in the Densu Delta, Ghana. Aquaculture International, 29(2):813–826, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00659-9.

Folorunso:2021:CFS

[2214] Ewumi Azeez Folorunso, Muhammad Arifur Rahman, Isaac Sarfo, George Darko, and Olumide Samuel Olowe. Catfish farming: a sustainability study at eriwe fish farming village in southwest Nigeria. Aquaculture International, 29(2):827–843, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00662-0.

Agarwal:2021:AAA

[2215] Pranay Agarwal, Pooja Kayala, Natrajan Chandrasekaran, Amitava Mukherjee, Shreshtha Shah, and John Thomas. Antioxidant and antibacterial activity of Gelidium pusillum (Stackhouse) against Aeromonas caviae and its applications in aquaculture. Aquaculture International, 29(2):845–858, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00661-1.

S:2021:PAI

[2216] Ranjani S, Pradeep Parthasarathy, Rameshkumar P, VimalKumar U, and Hemalatha S. Pungent anti-infective nanocolloids manipulate growth, biofilm formation, and CTX-M-15 gene expression in pathogens causing vibriosis. Aquaculture International, 29(2):859–869, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-021-00660-2.

Young:2021:SCA

[2217] Benjamin C. Young, Riyadh Hussain Alfaggeh, and Ibrahim AlMoutiri. Status and cost analysis of Sabaki tilapia farming in Saudi Arabia. Aquaculture International, 29(2):871–878, April 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00663-z.

Kersante:2021:FAA

[2218] Pierrick Kersanté, Guillaume Le Reste, Benoitr Diringer, Juan Quimi, Renaud Sergheraert, and Joël Duperray. Free amino acids mix made of poultry keratin improves survival of whiteleg shrimp post larvae (Litopenaeus vannamei) challenged with acute hepatopancreatic necrosis disease and white spot syndrome virus. Aquaculture International, 29 (2):879–890, April 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00665-x.

Zare:2021:IDA

[2219] Rasool Zare, Abdolmohammad Abedian Kenari, and Mohammadali Yazdani Sadati. Influence of dietary acetic acid, protexin (probiotic), and their combination on growth performance, intestinal microbiota, digestive enzymes, immunological parameters, and fatty acids composition in Siberian sturgeon (Acipenser baerii, Brandt, 1869). Aquaculture International, 29(3):891–910, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00652-2.

Ma:2021:EDC

[2220] Shuoli Ma, Xinxin Li, Dong Huang, Yanlin Guo, Junming Deng, Wanyou Zhou, Wenbing Zhang, and Kangsen Mai. Effects of dietary chromium yeast and astaxanthin on the growth performance, anti-oxidative capacity, and resistance to heat stress of abalone Haliotis discus hannai. Aquaculture International, 29(3):911–924, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00656-y.

Fan:2021:TAR

[2221] Wenqiao Fan, Mengjun Huang, Jingming Xu, Meixia Zhang, Yusong Jiang, Mingcheng Cai, Hanchang Sun, and Xiaoying Li. Transcriptome analysis reveals the effects of temperature on growth in tadpoles of spiny-bellied frog (Quasipaa boulengeri). Aquaculture International, 29(3):925–939, June 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00664-y.

Wu:2021:CSA

[2222] Shenji Wu, Jinqiang Huang, Yongjuan Li, Zhe Liu, Qian Zhang, Yucai Pan, and Xiaolan Wang. Cloning, sequence analysis, and expression of tyrp1a and tyrp2 genes related to body colour in different developmental stages and tissues of rainbow trout Oncorhynchus mykiss. Aquaculture International, 29(3):941–961, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00666-w.

Sati:2021:SBB

[2223] Himanshu Sati, Kaumeel Chokshi, Rajapitamahuni Soundarya, Arup Ghosh, and Sandhya Mishra. Seaweed-based biostimulant improves photosynthesis and effectively enhances growth and biofuel potential of a green microalga Chlorella variabilis. Aquaculture International, 29 (3):963–975, June 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00667-9.

Sousa:2021:CSH

[2224] Andria Gama Sousa, Acácio H. B. Pacheco, Gilson A. Siqueira-Pinto, Gleika T. J. dos Reis, Michelle M. S. Fugimura, Luciano J. Vaz, Paulo F. Marcusso, Fernando C. Ramos-Espinoza, and Gustavo da Silva Claudiano. Comparative study of hematological parameters of Colossoma macropomum anesthetized with benzocaine and eugenol by using different anticoagulants. Aquaculture International, 29(3):977–988, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-021-00668-8.

Cuadros:2021:IHA

[2225] Ruben C. Cuadros, Norma L. S. Rivadeneyra, Anai Flores-Gonzales, Omar Mertins, Jose C. O. Malta, M. Enrique Serrano-Martínez, and Patrick D. Mathews. Intestinal histological alterations in farmed redbellied pacu Piaractus brachypomus (Characiformes: Serrasalmidae) heavily infected by roundworms. Aquaculture International, 29(3):989– 998, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00670-0.

Just:2021:VSM

[2226] Philip Niclas Just, Bernd Köllner, and Matthew James Slater. Video surveillance methods to evaluate individual feeding response in rainbow trout (Oncorhynchus mykiss, Walbaum) — implications for feeding regime optimisation. Aquaculture International, 29(3):999–1013, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-021-00671-z.

Fu:2021:IMM

[2227] Longlong Fu, Tianheng Gao, Hucheng Jiang, Yan Zhang, and Jianlin Pan. Integrated miRNA-mRNA transcriptomic analysis of hepatopancreas reveals molecular strategies in Chinese mitten crab (Eriocheir sinensis) under acute nitrite stress. Aquaculture International, 29(3): 1015–1030, June 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00672-y.

Zhou:2021:RBE

[2228] Xuanyu Zhou, Shun Zhang, Xiaofang Zhu, Chunpu Zhao, Shun Cheng, Shanliang Xu, and Danli Wang. Relationship between expression of matrix metalloproteinase-9 and feeding on jellyfish in Pampus argenteus. Aquaculture International, 29(3):1031–1047, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00673-x.

Jyotsna:2021:AAS

[2229] Jyotsna, Parameswaran Vijayakumar, Tharmathass Stalin Dhas, Ravi Mani, and Vasantharaja Raguraman. Antiviral activity of sulfated polysaccharides from Sargassum ilicifolium against fish betanodavirus infection. Aquaculture International, 29(3):1049–1067, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00675-9.

Gomes:2021:ESM

[2230] Ana Gomes, Sílvia Lourenço, Pedro M. Santos, Andreia Raposo, Susana Mendes, Sílvia C. Gonçalves, Susana M. F. Ferreira, and Ana Pombo. Effects of single and mixed-diatom diets on growth, condition, and survival of larvae of the sea urchin Paracentrotus lividus (Lamarck, 1816). Aquaculture International, 29(3):1069–1090, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00676-8.

Iyapparaj:2021:EPF

[2231] Palanisamy Iyapparaj, Peranandam Revathi, Sesuraj Divakar, Thangapandi Marudhupandi, and Shanmuganathan Niroshan. Evaluation of polyherbal feed supplement (ban-v) on the enhancement of growth and non-specific immune responses in shrimp Penaeus vannamei. Aquaculture International, 29(3):1091–1101, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00680-y.

Elumalai:2021:SLC

[2232] Preetham Elumalai, Abdul Salam Rubeena, Sreeja Lakshmi, Veerappan Anbazhagan, Jesu Arockiaraj, Mani Divya, Sekar Vijayakumar, and Baskaralingam Vaseeharan. Shrimp lectin–conjugated copper sulfide nanoparticles enhance immune response and gene expression in Etroplus suratensis infected with Aeromonas hydrophila. Aquaculture International, 29(3):1103–1120, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00679-5.

Minaz:2021:OPA

[2233] Mert Minaz and Aysegul Kubilay. Operating parameters affecting biofloc technology: carbon source, carbon/nitrogen ratio, feeding regime, stocking density, salinity, aeration, and microbial community manipulation. Aquaculture International, 29(3):1121–1140, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00681-x.

Zhu:2021:CSP

[2234] Jiajie Zhu, Kai Yu, Qiuwei Ao, Yun Tan, Qiang Fu, and Hesheng Jiang. Comparative splenic proteomic analysis of susceptible and resistant GIFT tilapia following challenge with Streptococcus agalactiae. Aquaculture International, 29(3):1141–1159, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00683-9.

Adamek:2021:IAP

[2235] Zdeněk Adámek, Zorka Dulić, David Hlaváč, and Maria Anton-Pardo. Insect attractants and plant biomass as natural food complements in pond culture of stock chub (Squalius cephalus). Aquaculture International, 29 (3):1161–1179, June 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00682-w.

Roy:2021:DAF

[2236] Subha M. Roy, Jayraj P, Rajendra Machavaram, C. M. Pareek, and B. C. Mal. Diversified aeration facilities for effective aquaculture systems — a comprehensive review. Aquaculture International, 29(3): 1181–1217, June 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00685-7.

Beninger:2021:EBF

[2237] Peter G. Beninger, Daphné Chérel, and Lucie Kessler. Examining bivalve fecundity: oocyte viability revealed by neutral red vital staining. Aquaculture International, 29(3):1219–1231, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00686-6.

Ye:2021:FMF

[2238] Tingqi Ye, Jiayan Wu, Cuimin Mu, Guangchen Pan, Jiakang Chen, and Xuepeng Wang. FucP mediates fucose to regulate T3SS in Edwardsiella piscicida and promotes intestinal colonization in tilapia. Aquaculture International, 29(3):1233–1243, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00688-4.

Cammies:2021:EEL

[2239] Christopher Cammies, David Mytton, and Rosemary Crichton. Exploring economic and legal barriers to commercial aquaponics in the EU through the lens of the UK and policy proposals to address them. Aquaculture International, 29(3):1245–1263, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00690-w.

Dong:2021:EAI

[2240] Khuu Thi Phuong Dong, Nguyen Minh Duc, Nguyen Thi Ngoc Hoa, Bui Le Thai Hanh, and Le Thi Thu Trang. Effects of application for international quality assurance certifications on the farm’s profit: evidence from Vietnamese shrimp. Aquaculture International, 29(3): 1265–1277, June 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00691-9.

Aranda:2021:LCG

[2241] Dalila Aldana Aranda, Martha Enríquez-Díaz, Wendy González-López, Jean-Louis Mansot, and Olivier Gros. Larval calcification and growth of veligers to early pediveliger of the queen conch Strombus gigas in mesocosm and laboratory conditions. Aquaculture International, 29(3): 1279–1294, June 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00696-4.

Abakari:2021:EBM

[2242] Godwin Abakari, Guozhi Luo, Lina Shao, Yusuf Abdullateef, and Samuel Jerry Cobbina. Effects of biochar on microbial community in bioflocs and gut of Oreochromis niloticus reared in a biofloc system. Aquaculture International, 29(3):1295–1315, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00697-3.

Ruiz-Velazco:2021:EPH

[2243] Javier M. J. Ruiz-Velazco, Miguel A. González-Romero, Nallely Estrada-Perez, and Alfredo Hernandez-Llamas. Evaluating partial harvesting strategies for whiteleg shrimp Litopenaeus (Penaeus) vannamei semiintensive commercial production: profitability, uncertainty, and economic risk. Aquaculture International, 29(3):1317–1329, June 2021. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-021-00695-
5.

Luo:2021:AHF

[2244] Wei Luo, Ning Zhang, Dongjie Wang, Zhou Xu, Tianzhu Wang, Xiaoyang Zhang, Tong Gao, Min Liao, Yuejin Long, and Zongjun Du. Assessment of the heterosis of F_{1} hybrids of Misgurnus anguillicaudatus and Paramisgurnus dabryanus using microsatellite-based parentage assignment. Aquaculture International, 29(3):1331–1341, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00700-x.

Zhang:2021:EDB

[2245] Jianxin Zhang, Mengyuan Huang, Junchang Feng, Yongyan Chen, Meng Li, and Xulu Chang. Effects of dietary Bacillus licheniformis on growth performance, intestinal morphology, intestinal microbiome, and disease resistance in common carp (Cyprinus carpio L.). Aquaculture International, 29(3):1343–1358, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00701-w.

Chen:2021:EAN

[2246] Jian Chen, Hui Luo, Ziqin Zhai, Hongchen Wang, Baosuo Liu, Lirong Bai, and Dahui Yu. Estimates of additive and non-additive genetic effects on growth traits in a diallel cross of three strains of pearl oyster (Pinctada fucata). Aquaculture International, 29(3):1359–1371, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00702-9.

Xue:2021:RSL

[2247] Hai-Bo Xue, Can Liu, Yuan Liu, Wei-Na Wang, and Bin Xu. Roles of surface layer proteins in the regulation of Pediococcus pentosaceus on growth performance, intestinal microbiota, and resistance to Aeromonas hydrophila in the freshwater prawn Macrobrachium rosenbergii. Aquaculture International, 29(3):1373–1391, June 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00704-7.

Attia:2021:MMI

[2248] Marwa M. Attia, Mamdouh Y. Elgendy, Mohamed Abdelsalam, Azza Hassan, Abdelbary Prince, Nagla M. K. Salaeh, and Nehal A. Younis. Morpho-molecular identification of Heterophyes heterophyes encysted metacercariae and its immunological and histopathological effects on farmed Mugil cephalus in Egypt. Aquaculture International, 29(3):1393–1407, June 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00708-3.

Koussovi:2021:OST

[2249] Guillaume Koussovi, Farokh Niass, Diane Kpoguè, Arsène M. Houssou, Clément A. Bonou, and Elie Montchowui. Optimal salinity and temperature requirements for the early developmental stages and survival of freshwater prawn Macrobrachium macrobrachion (Herklots, 1851) in a controlled environment. Aquaculture International, 29(4):1409– 1425, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00669-7.

Pinho:2021:GSA

[2250] Sara M. Pinho, Luiz H. David, Fabiana Garcia, Karel J. Keesman, Maria Célia Portella, and Simon Goddek. South American fish species suitable for aquaponics: a review. Aquaculture International, 29(4): 1427–1449, August 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00674-w.

G:2021:EGD

[2251] Vidya Sagar Reddy G. and Vijaya Ch. Efficacy of β-glucan from Debaryomyces hansenii as an immunostimulant in Litopenaeus vannamei culture. Aquaculture International, 29(4):1451–1458, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00678-6.

Nasr-Eldahan:2021:RAN

[2252] Sameh Nasr-Eldahan, Asmaa Nabil-Adam, Mohamed Attia Shreadah, Adham M. Maher, and Tamer El-Sayed Ali. A review article on nanotechnology in aquaculture sustainability as a novel tool in fish disease control. Aquaculture International, 29(4):1459–1480, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00677-7.

Bekkozhayeva:2021:AIN

[2253] Dinara Bekkozhayeva, Mohammadmehdi Saberioon, and Petr Cisar. Automatic individual non-invasive photo-identification of fish (Sumatra barb Puntigrus tetrazona) using visible patterns on a body. Aquaculture International, 29(4):1481–1493, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00684-8.

Farias:2021:APC

[2254] Caio Francisco Santana Farias, Franmir Rodrigues Brandão, Fernanda de Alexandre Sebastião, Damy Caroline de Melo Souza, Patrícia Castro Monteiro, Cláudia Majolo, and Edsandra Campos Chagas. Albendazole and praziquantel for the control of Neoechinorhynchus buttnerae in tambaqui (Colossoma macropomum). Aquaculture International, 29(4): 1495–1505, August 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00687-5.

Sathyaruban:2021:PNC

[2255] Sutharshiny Sathyaruban, Deepthi Inoka Uluwaduge, Shivatharsiny Yohi, and Sivashanthini Kuganathan. Potential natural carotenoid sources for the colouration of ornamental fish: a review. Aquaculture International, 29(4):1507–1528, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00689-3.

Hossain:2021:GEE

[2256] Md. Shahdat Hossain, Jingcheng Dai, and Dongru Qiu. European eel (Anguilla anguilla) GI tract conserves a unique metagenomics profile in the recirculation aquaculture system (RAS). Aquaculture International, 29(4):1529–1544, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00692-8.

Purbosari:2021:EAS

[2257] Ninik Purbosari, Endang Warsiki, Khaswar Syamsu, Joko Santoso, and Irzal Effendi. Evaluation of the application of seaweed (Eucheuma cottonii) extract as fish anesthetic agent. Aquaculture International, 29(4): 1545–1560, August 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00693-7.

Cubillo:2021:DEC

[2258] Alhambra Martínez Cubillo, João G. Ferreira, João Lencart-Silva, Nick G. H. Taylor, Adam Kennerley, James Guilder, Susan Kay, and Pauline Kamermans. Direct effects of climate change on productivity of European aquaculture. Aquaculture International, 29(4):1561–1590, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-021-00694-6.

Seethalakshmi:2021:SDM

[2259] P. S. Seethalakshmi, Riya Rajeev, George Seghal Kiran, and Joseph Selvin. Shrimp disease management for sustainable aquaculture: innovations from nanotechnology and biotechnology. Aquaculture International, 29(4):1591–1620, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00698-2.

Chen:2021:EMP

[2260] Yung-Sheng Chen, Ming-Sheng Ko, Yu-Cheng Wang, Ping Chen, and Ta-Kang Liu. Enrichment of marine productivity utilizing steelmaking slag: managing conflicting stakeholders’ interests in Taiwan. Aquaculture International, 29(4):1621–1638, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00699-1.

Ouyang:2021:SIC

[2261] Ping Ouyang, Jiehao Chen, Lizi Yin, Yi Geng, Defang Chen, Kaiyu Wang, Weiming Lai, Hongrui Guo, Jing Fang, Zhengli Chen, Li Tang, Chao Huang, and Xiaoli Huang. The sub-inhibitory concentration of cinnamaldehyde resists Aeromonas hydrophila pathogenicity via inhibition of W-pili production. Aquaculture International, 29(4):1639– 1655, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00705-6.

Rodrigues:2021:RVD

[2262] Ricardo Vieira Rodrigues, Lucas C. Maltez, Cristiano C. Ferreira, Thiago P. A. P. Oliveira, Luís A. Sampaio, and José María Monserrat. ROS in vivo determination and antioxidant responses in rotifers Brachionus plicatilis fed with commercial yeast Saccharomyces cerevisiae or microalgae Nannochloropsis oculata. Aquaculture International, 29(4): 1657–1667, August 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00707-4.

Mishra:2021:IHD

[2263] Induja Mishra, Namita Joshi, Pashupati Nath, and B. D. Joshi. The influences of haematozoans on different haematological and morphological parameters in a freshwater catfish Heteropneustes fossilis (Bloch). Aquaculture International, 29(4):1669–1681, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00706-5.

Skelton:2021:AUM

[2264] Bradley M. Skelton and Andrew G. Jeffs. An assessment of the use of macroalgae to improve the retention of Greenshell^{T M} mussel (Perna canaliculus) spat in longline culture. Aquaculture International, 29(4): 1683–1695, August 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00710-9.

Zhang:2021:DMR

[2265] Yuanyuan Zhang, Ze Fan, Di Wu, Jinnan Li, Qiyou Xu, Hongbai Liu, and Liansheng Wang. Dietary magnesium requirement on dietary minerals and physiological function of juvenile hybrid sturgeon (Acipenser schrenckii [female sign] × Acipenser baerii [male sign]). Aquaculture International, 29(4):1697–1709, August 2021. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00712-7.

Acharjee:2021:PHF

[2266] Debasish Chandra Acharjee, Mohammad Ismail Hossain, and G. M. Monirul Alam. Post-harvest fish loss in the fish value chain and the determinants: empirical evidence from Bangladesh. Aquaculture International, 29(4): 1711–1720, August 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00711-8.

Delaide:2021:SSA

[2267] Boris Delaide, Edson Panana, Stefan Teerlinck, and Peter Bleyaert. Suitability of supernatant of aerobic and anaerobic pikeperch (Sander lucioperca L.) sludge treatments as a water source for hydroponic production of lettuce (Lactuca sativa L. var. capitata). Aquaculture International, 29 (4):1721–1735, August 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00709-2.

Capelle:2021:ORM

[2268] Jacob J. Capelle, Ainhoa Blanco Garcia, Pauline Kamermans, Marc Y. Engelsma, and Henrice M. Jansen. Observations on recent mass mortality events of marine mussels in the Oosterschelde, the Netherlands. Aquaculture International, 29(4):1737–1751, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00713-6.

Babatunde:2021:QSA

[2269] Adeleke Babatunde, Robertson-Andersson Deborah, Moodley Gan, and Taylor Simon. A quantitative SWOT analyses of key aquaculture players in Africa. Aquaculture International, 29(4):1753–1770, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-021-00715-4.

Pieroni:2021:CUS

[2270] Soraya Pieroni, Bruno Silva Olier, Isabela Ramos Lima, Isadora Marini Sanches, Vanessa Villanova Kuhnen, and Eduardo Gomes Sanches. Can use of substrates affect water quality in aquatic organism culture? Aquaculture International, 29(4):1771–1783, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00718-1.

Shi:2021:TAR

[2271] Wenjun Shi, Pan Wang, Runhao Hu, Xihe Wan, Hui Shen, Hui Li, Libao Wang, Yi Qiao, Ge Jiang, Jie Cheng, and Zeyu Yang. Transcriptome analysis reveals hub genes in the hepatopancreas of Exopalaemon carinicauda in response to hypoxia and reoxygenation. Aquaculture International, 29(4):1785–1811, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00717-2.

Parappurathu:2021:GMP

[2272] Shinoj Parappurathu, N. K. Sanil, P. K. Asokan, S. R. Krupesha Sharma, M. A. Pradeep, Shelton Padua, Suja Gangadharan, Gishnu Mohan, and Prasanna Kumar Patil. Green mussel (Perna viridis L.) farming in India: an analysis of major growth milestones, recent decline due to disease incidence, and prospects for revival. Aquaculture International, 29(4): 1813–1828, August 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00716-3.

Seth:2021:PET

[2273] Mrinal Seth, Natrajan Chandrasekaran, Amitava Mukherjee, and John Thomas. Pathogenicity of Edwardsiella tarda in Oreochromis mossambicus and treatment by Tamarindus indica seed extract. Aquaculture International, 29(4):1829–1841, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00719-0.

De:2021:YGC

[2274] Himansu Kumar De, Simantini Shasani, and Manoj Kumar Das. Yield gaps in composite carp culture in eastern India. Aquaculture International, 29(4):1843–1851, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00720-7.

Wang:2021:IEA

[2275] Huizheng Wang, Delong Li, Minghong Fan, Junke Zhu, and Xin Han. Identification and expression analysis of Ricin B-like lectin genes in the fish pathogen Saprolegnia parasitica. Aquaculture International, 29(4): 1853–1868, August 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00722-5.

Jiang:2021:MCS

[2276] Xiaona Jiang, Jiaxin Sun, Chitao Li, Xuesong Hu, Yanlong Ge, Bao Li, Lianyu Shi, and Zhiying Jia. Molecular cloning and sequence characterization of common carp (Cyprinus carpio) integrin β1 (ITGβ1) and its temporal expression in response to CyHV-3. Aquaculture International, 29(4):1869–1884, August 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00723-4.

Chaiw-Yee:2021:UFS

[2277] Teoh Chaiw-Yee and Wong Yin-Ying. Use of fish and shrimp hydrolysates as dietary supplements to increase feeding and growth of juvenile striped catfish (Pangasius hypophthalmus). Aquaculture International, 29(4): 1885–1894, August 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00725-2.

Xue:2021:ASI

[2278] Shuqun Xue, Jiawen Lin, Yue Han, and Ying Han. Ammonia stress– induced apoptosis by p53-BAX/ BCL-2 signal pathway in hepatopancreas of common carp (Cyprinus carpio). Aquaculture International, 29 (4):1895–1907, August 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00724-3.

Bergamo:2021:EFM

[2279] Gabriela Claudia Arato Bergamo, Bruno Silva Olier, Otávio Mesquita de Sousa, Vanessa Villanova Kuhnen, Maria Fernanda Guedes Pessoa, and Eduardo Gomes Sanches. Economic feasibility of mussel (Perna perna) and cobia (Rachycentron canadum) produced in a multitrophic system. Aquaculture International, 29(5):1909–1924, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-021-00762-x.

Das:2021:IGE

[2280] Sweta Das, Arathi Dharmaratnam, Charan Ravi, Raj Kumar, and Thangaraj Raja Swaminathan. Immune gene expression in cyprinid herpesvirus-2 (CyHV-2)-sensitized peripheral blood leukocytes (PBLs) co-cultured with CyHV-2-infected goldfish fin cell line. Aquaculture International, 29(5):1925–1934, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00721-6.

Siagian:2021:DLA

[2281] Desi Rahmadani Siagian, Dedi Jusadi, Julie Ekasari, and Mia Setiawati. Dietary α-lipoic acid supplementation to improve growth, blood chemistry, and liver antioxidant status of African catfish Clarias gariepinus. Aquaculture International, 29(5):1935–1947, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00726-1.

Lahnsteiner:2021:EDN

[2282] Franz Lahnsteiner. Effect of disinfection of non-hardened Salmo trutta eggs with chloramine T(R), Wofasteril(R), and hydrogen peroxide on embryo and larvae viability, microorganism load, lipid peroxidation, and protein carbonylation. Aquaculture International, 29(5):1949–1962, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00727-0.

Weston:2021:FAS

[2283] Kim Weston, Sam Jahangard, Brett A. Ingram, Adam D. Miller, Geordie Jennings, and Craig D. H. Sherman. Factors affecting settlement, growth and metamorphosis of hatchery-produced Australian blue mussel larvae. Aquaculture International, 29(5):1963–1977, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00728-z.

Agriandini:2021:FTD

[2284] Maulina Agriandini, Sukenda Sukenda, Widanarni Widanarni, and Angela Mariana Lusiastuti. Fate and tissue distribution of Mycobacterium fortuitum through immersion challenge as a model of natural infection in Osphronemus goramy. Aquaculture International, 29(5): 1979–1989, October 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00729-y.

Rioja:2021:AES

[2285] Rose Angeli Rioja, Nadia Palomar-Abesamis, and Marie Antonette Juinio-Meñez. Associated effects of shading on the behavior, growth, and survival of Stichopus cf. horrens juveniles. Aquaculture International, 29 (5):1991–2007, October 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00732-3. See correction [2393].

Arahou:2021:ICC

[2286] Fadia Arahou, Rachida Hassikou, Moustapha Arahou, Laila Rhazi, and Imane Wahby. Influence of culture conditions on Arthrospira platensis growth and valorization of biomass as input for sustainable agriculture. Aquaculture International, 29(5):2009–2020, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00730-5.

Bokor:2021:SSM

[2287] Z. Bokor, D. Żarski, K. Palińska-Żarska, S. Krejszeff, J. Król, J. Ifj. Radóczi, Á. Horváth, L. Várkonyi, B. Urbányi, and G. Bernáth. Standardization of sperm management for laboratory assessment of sperm quality and in vitro fertilization in Eurasian perch (Perca fluviatilis). Aquaculture International, 29(5):2021–2033, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00731-4.

Xu:2021:CFL

[2288] Baohong Xu, Ligang Lv, Tiaoyi Xiao, Qiaolin Liu, Hang Su, Yi Liu, and Jiajia Ni. Cloning of the full-length cDNA of the gene encoding complement C5 from grass carp (Ctenopharyngodon idella) and its expression in different tissues by following grass carp reovirus infection. Aquaculture International, 29(5):2035–2048, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00733-2.

Spelta:2021:DWQ

[2289] Anna Carolina Ferreira Spelta, João Paulo Silva Lorenzini, Luanna do Carmo Neves, Luciano dos Santos Rodrigues, Ângela Maria Quintão Lana, Cintia Labussière Nakayama, Rebeca Valentim Marques, Marco Yuri Rodrigues Napoli, Gustavo Moreira Alves, and Kleber Campos Miranda Miranda-Filho. Dynamics of the water quality parameters in the super-intensive culture of Litopenaeus vannamei in BFT system on artificial brackish water. Aquaculture International, 29(5):2049– 2063, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00734-1.

Salam:2021:PMA

[2290] Hala Sayed Hassan Salam, Walaa M. S. Mohamed, Sahar Abdel Aleem Abdel Aziz, Asmaa N. Mohammed, and Fatma M. M. Korni. Prevention of motile Aeromonas septicemia in Nile tilapia, Oreochromis niloticus, using thyme essential oil and its nano-emlusion. Aquaculture International, 29(5):2065–2084, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00735-0.

Khatua:2021:DPL

[2291] Rooprekha Khatua, Kedar Nath Mohanta, Nitish Kumar Chandan, Rojalin Pattanayak, Choudhury Suryakant Mishra, and Prem Kumar. Dietary protein and lipid concentrations affect the growth, nutritional indices, and whole-body composition of long-whisker catfish, Mystus gulio, fry. Aquaculture International, 29(5):2085–2099, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00738-x.

Paolacci:2021:EPL

[2292] Simona Paolacci, Vlastimil Stejskal, and Marcel A. K. Jansen. Estimation of the potential of Lemna minor for effluent remediation in integrated multi-trophic aquaculture using newly developed synthetic aquaculture wastewater. Aquaculture International, 29(5):2101–2118, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00736-z.

Chakraborty:2021:EMP

[2293] Prasun Chakraborty, Md. Rabiul Islam, Md. Amzad Hossain, Umme Kaniz Fatema, Dinesh Chandra Shaha, Md. Shah Alam Sarker, and Taslima Akter. Earthworm meal (Perionyx excavatus) as an alternative protein source to fish meal in feed for juvenile butter catfish (Ompok pabda). Aquaculture International, 29(5):2119–2129, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00737-y.

Chen:2021:ILI

[2294] Xiurong Chen, Yingying Yang, Quanling Lu, Xiaoli Sun, Shanshan Wang, Qiuyue Li, Xiao Wei, and Yuan Wang. The influence of light intensity and organic content on cultivation of Chlorella vulgaris in sludge extracts diluted with BG11. Aquaculture International, 29(5): 2131–2144, October 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00740-3.

Sarkar:2021:IBF

[2295] Soumyabrata Sarkar, P. Nila Rekha, A. Panigrahi, R. R. Das, S. Rajamanickam, and C. P. Balasubramanian. Integrated brackishwater farming of red seaweed Agarophyton tenuistipitatum and Pacific white leg shrimp Litopenaeus vannamei (Boone) in biofloc system: a production and bioremediation way out. Aquaculture International, 29(5): 2145–2159, October 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00739-w.

Rahman:2021:IMM

[2296] Mohammad R. Rahman, Md. R. I. Sarder, Ananaya A. Nishat, Rafiul Islam, and A. H. M. Kohinoor. Induction of meiotic and mitotic gynogenesis in silver barb (Barbonymus gonionotus) through cold shock treatment. Aquaculture International, 29(5):2161–2179, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00744-z.

Dagoudo:2021:EAH

[2297] Missinhoun Dagoudo, Jun Qiang, Jin-Wen Bao, Yi-Fan Tao, Hao-Jun Zhu, Ezra Mutebi Tumukunde, Tlou Kevin Ngoepe, and Pao Xu. Effects of acute hypoxia stress on hemato-biochemical parameters, oxidative resistance ability, and immune responses of hybrid yellow catfish (Pelteobagrus fulvidraco × P. vachelli) juveniles. Aquaculture International, 29(5): 2181–2196, October 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00742-1.

Cipriani:2021:DID

[2298] Luiz Augusto Cipriani, Natalia Ha, Nandara Soares de Oliveira, and Thiago El Hadi Perez Fabregat. Does ingestion of duckweed (Lemna minor) improve the growth and productive performance of juvenile Nile tilapia (Oreochromis niloticus) given formulated feeds in a recirculation system? Aquaculture International, 29(5):2197–2205, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00743-0.

Gheytasi:2021:FRT

[2299] Abbas Gheytasi, Seyed Pezhman Hosseini Shekarabi, Houman Rajabi Islami, and Mehdi Shamsaie Mehrgan. Feeding rainbow trout, Oncorhynchus mykiss, with lemon essential oil loaded in chitosan nanoparticles: effect on growth performance, serum hemato-immunological parameters, and body composition. Aquaculture International, 29(5): 2207–2221, October 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00741-2.

Turcihan:2021:EFD

[2300] Gülşah Turcihan, Emre Turgay, Remziye Eda Yardımcı, and Kamil Mert Eryalçın. The effect of feeding with different microalgae on survival, growth, and fatty acid composition of Artemia franciscana metanauplii and on predominant bacterial species of the rearing water. Aquaculture International, 29(5):2223–2241, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00745-y.

Tartila:2021:EDS

[2301] Shobrina Silmi Qori Tartila, Dedi Jusadi, Mia Setiawati, and Ichsan Achmad Fauzi. Evaluation of dietary supplementation with cinnamon products on growth, blood composition, liver structure, and meat quality of striped catfish (Pangasianodon hypophthalmus). Aquaculture International, 29(5):2243–2257, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00746-x.

DAgaro:2021:GSS

[2302] Edo D’Agaro, Andea Favaro, Stefano Matiussi, Pier Paolo Gibertoni, and Stefano Esposito. Genomic selection in salmonids: new discoveries and future perspectives. Aquaculture International, 29(5):2259– 2289, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00747-w.

Du:2021:DSR

[2303] Lian cai Du, Hai rui Yu, Ling yao Li, Qin Zhang, Qi Tian, Jin qian Liu, and Ling ling Shan. Dietary selenium requirement of Coho salmon (Oncorhynchus kisutch W.) alevins. Aquaculture International, 29(5): 2291–2304, October 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00749-8.

Ekasari:2021:IBQ

[2304] Julie Ekasari, Utomo Adi Nugroho, Nurul Fatimah, Deasy Angela, Yuni Puji Hastuti, Gde Sasmita Julyantoro Pande, and F. M. I. Natrah. Improvement of biofloc quality and growth of Macrobrachium rosenbergii in biofloc systems by Chlorella addition. Aquaculture International, 29 (5):2305–2317, October 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00750-1.

Sgnaulin:2021:CPP

[2305] Tayna Sgnaulin, Sara M. Pinho, Emerson Giuliani Durigon, Micheli Cristina Thomas, Giovanni Lemos de Mello, and Maurício Gustavo Coelho Emerenciano. Culture of pacu Piaractus mesopotamicus in biofloc technology (BFT): insights on dietary protein sparing and stomach content. Aquaculture International, 29(5):2319–2335, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00748-9.

Oviedo-Bolanos:2021:MIS

[2306] Karen Oviedo-Bolaños, Jorengeth Abad Rodríguez-Rodríguez, Carolina Sancho-Blanco, Juan Esteban Barquero-Chanto, Nelson Peña-Navarro, Cesar Marcial Escobedo-Bonilla, and Rodolfo Umaña-Castro. Molecular identification of Streptococcus sp. and antibiotic resistance genes present in tilapia farms (Oreochromis niloticus) from the Northern Pacific region, Costa Rica. Aquaculture International, 29(5):2337–2355, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00751-0.

Yuan:2021:IIR

[2307] Yongzhu Yuan, Yonggang Zhang, Guoshan Qi, Hai Ren, Guisheng Gao, Xiaomin Jin, and Hai Fang. Isolation, identification, and resistance gene detection of Vibrio harveyi from Scophthalmus maximus. Aquaculture International, 29(5):2357–2368, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00752-z.

Maciel:2021:PAM

[2308] Patricia Oliveira Maciel and Elizabeth Gusmão Affonso. Praziquantel against monogeneans of tambaqui (Colossoma macropomum). Aquaculture International, 29(5):2369–2386, October 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00753-y.

Sushila:2021:CVH

[2309] Ngairangbam Sushila, Basanta Kumar Das, K. Pani Prasad, and Gayatri Tripathi. Characterisation and validation of housekeeping genes for qRT-PCR expression analysis in Pterophyllum scalare. Aquaculture International, 29(6):2387–2402, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00754-x.

Broom:2021:EDD

[2310] Mohammed Broom, Mohamed Hosny Gabr, Mamdouh Al-Harbi, and Sathianeson Satheesh. Effect of different diets on growth of the subadult sea cucumber Holothuria scabra. Aquaculture International, 29(6): 2403–2414, December 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00756-9.

Makhutova:2021:FAC

[2311] Olesia N. Makhutova and Kirill N. Stoyanov. Fatty acid content and composition in tissues of baikal grayling (Thymallus baicalensis), with a special focus on DHA synthesis. Aquaculture International, 29(6): 2415–2433, December 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00755-w.

Pandolfi:2021:DMM

[2312] Victor César Freitas Pandolfi, Andrei Lincoln Yamachita, Felipe Pinheiro de Souza, Sara Mataroli de Godoy, Ed Christian Suzuki de Lima, Daniele Cassiano Feliciano, Ulisses de Pádua Pereira, Jayme Aparecido Povh, Denise Rocha Ayres, Annaiza Braga Bignardi, Jefferson Murici Penafort, Claudete de Fátima Ruas, and Nelson Mauricio Lopera-Barrero. Development of microsatellite markers and evaluation of genetic diversity of the Amazonian ornamental fish Pterophyllum scalare. Aquaculture International, 29(6):2435–2449, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00757-8.

Cui:2021:EDG

[2313] Qingman Cui, Ruibin Sun, and Chunying Yuan. Effects of dietary glutamine dipeptide on growth performance, intestinal digestive enzyme activity, and gene expression of related factors in Penaeus vannamei. Aquaculture International, 29(6):2451–2461, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00759-6.

Lee:2021:DSB

[2314] Chorong Lee, Soohwan Kim, Jaehyeong Shin, Min-Gi Kim, Buddhi E. Gunathilaka, Sung Hun Kim, Ji Eun Kim, Seung-Cheol Ji, Jee Eun Han, and Kyeong-Jun Lee. Dietary supplementations of Bacillus probiotic improve digestibility, growth performance, innate immunity, and water ammonia level for Pacific white shrimp, Litopenaeus vannamei.

Aquaculture International, 29(6):2463–2475, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00760-z.

Castro:2021:AED

[2315] Liliane Araújo Castro, Sanny Maria Andrade-Porto, Rebeca Guimarães Oliveira, Yasmin Luzeiro Batista, Felipe Moura Araújo Silva, Christiane Patrícia Feitosa Oliveira, Elizabeth Gusmão Affonso, Claudinei Cruz, and Marcos Tavares-Dias. Antiparasitic efficacy of dietary administration of trichlorfon (Masoten(R)) in the control of Neoechinorhynchus buttnerae (Neochinorhynchidae) in Colossoma macropomum (Serrasalmidae). Aquaculture International, 29(6):2477–2488, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-021-00763-w.

Mirbakhsh:2021:EBS

[2316] Maryam Mirbakhsh, Masoud Mahjoub, Mohammad Afsharnasab, Shapoor Kakoolaki, Mostafa Sayyadi, and Saeid Hosseinzadeh. Effects of Bacillus subtilis on the water quality, stress tolerance, digestive enzymes, growth performance, immune gene expression, and disease resistance of white shrimp (Litopenaeus vannamei) during the early hatchery period. Aquaculture International, 29(6):2489–2506, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00758-7.

Miao:2021:FAF

[2317] Miao Miao, Huang Liu, and Jun Chen. Factors affecting fluctuations in China’s aquatic product exports to Japan, the USA, South Korea, Southeast Asia, and the EU. Aquaculture International, 29(6):2507– 2533, December 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00761-y.

Zhu:2021:ERF

[2318] Zhan-Hao Zhu, Qi hui Yang, Bei ping Tan, Xiao-Qiu Zhou, Xiao hui Dong, Shu yan Chi, Hong yu Liu, and Shuang Zhang. Effects of replacing fishmeal with soybean protein concentrate (SPC) on growth, blood biochemical indexes, non-specific immune enzyme activity, and nutrient apparent digestibility for juvenile Litopenaeus vannamei. Aquaculture International, 29(6):2535–2554, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00765-8.

Fan:2021:EDB

[2319] Ying Fan, Xiaolu Wang, Youhong Wang, Haibin Ye, Xiaoqing Yu, Shuxian Wang, Jing Diao, La Xu, Chunlei Gai, Hongjun Liu, and Danping Ma. Effect of dietary Bacillus licheniformis on growth, intestinal health, and resistance to nitrite stress in Pacific white shrimp Litopenaeus vannamei. Aquaculture International, 29(6):2555–2573, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00764-9.

Meng:2021:OHP

[2320] Zhen Meng, Xinfu Liu, Yudong Jia, Bin Liu, Zhi Yang, and Hesen Zhang. Optimization of hydrostatic pressure, timing, and duration parameters for the induction of tetraploidy in turbot, Scophthalmus maximus. Aquaculture International, 29(6):2575–2589, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00766-7.

Mirzargar:2021:AEB

[2321] Seyed Saeed Mirzargar, Reza Habibnejad Roshan, Ali Taheri Mirghaed, Hooman Rahmati-Holasoo, and Melika Ghelichpour. Anesthetic efficacy and biochemical effects of 1,8-cineole on Caspian trout, Salmo caspius. Aquaculture International, 29(6):2591–2603, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00767-6.

Lv:2021:EAE

[2322] Huirong Lv, Mrope Peter, Jun wook Hur, Yang Gao, and Zhangjie Chu. Effects of ammonia exposure on oxidative stress, immune enzyme activities, and intestinal microbiota of Pacific white shrimp Litopenaeus vannamei. Aquaculture International, 29(6):2605–2618, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-021-00770-x.

Zhang:2021:RNR

[2323] Yanpeng Zhang, Zhen Zhang, Xu-Fang Liang, Shan He, and Jing Xu. Role of NPY receptor 8 in regulating of food intake in Chinese perch (Siniperca chuatsi). Aquaculture International, 29(6):2619– 2634, December 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00771-w.

Wang:2021:HPA

[2324] Zhen Wang, Zhongdian Dong, Dawei Zhang, Liqun Liang, and Weijie Mu. Histopathological parameters, antioxidant enzyme levels, transcriptome, and hematology parameters of Amur minnow (Phoxinus lagowskii) infection with Tetrahymena pyriformis. Aquaculture International, 29(6): 2635–2659, December 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00772-9.

Fan:2021:CCE

[2325] Rui Fan, Yundong Li, Shigui Jiang, Song Jiang, Qibin Yang, Lishi Yang, Jianhua Huang, and Falin Zhou. cDNA cloning and expression analysis of glutaredoxin 3 in black tiger shrimp Penaeus monodon. Aquaculture International, 29(6):2661–2679, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00774-7.

Wang:2021:IFF

[2326] Cong Wang, Zhen Li, Tan Wang, Xianbao Xu, Xiaoshuan Zhang, and Daoliang Li. Intelligent fish farm — the future of aquaculture. Aquaculture International, 29(6):2681–2711, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00773-8.

Kumaran:2021:APF

[2327] Kumaran M, P. R. Anand, J. Ashok Kumar, M. Muralidhar, K. P. Kumaraguru Vasagam, and K. K. Vijayan. Assessment of perceived farming risks, communication of risk management practices, and evaluation of their efficiency in Pacific white shrimp (Penaeus vannamei) farming — a survey-based cross-sectional study. Aquaculture International, 29(6): 2713–2730, December 2021. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00775-6.

Kazemi:2021:EDD

[2328] Esmaeil Kazemi, Sajad Nazari, Iman Sourinejad, Mohammad Pourkazemi, Hamed Paknejad, and Khalil Eslamloo. Effect of different dietary zinc sources on seminal plasma enzymatic activity, antioxidant, and immunerelated gene expression in rainbow trout (Oncorhynchus mykiss). Aquaculture International, 29(6):2731–2750, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00778-3.

Mohammadi:2021:GPB

[2329] Yasaman Mohammadi, Mehrzad Mesbah, Mojdeh Chelemal Dezfoulnejad, Mehdi Shamsaie Mehrgan, and Houman Rajabi Islami. Growth performance, blood biochemical parameters, immune response, and antioxidant defense of Asian seabass (Lates calcarifer) fingerlings exposed to monovalent and bivalent vaccines against Streptococcus iniae and Vibrio harveyi. Aquaculture International, 29(6):2751–2767, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-021-00776-5.

Sutriana:2021:ESC

[2330] Amalia Sutriana, Mst. Nahid Akter, Roshada Hashim, and Siti Azizah Mohd Nor. Effectiveness of single and combined use of selected dietary probiotic and prebiotics on growth and intestinal conditions of striped catfish (Pangasianodon hypophthalmus, Sauvage, 1978) juvenile. Aquaculture International, 29(6):2769–2791, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00777-4.

Fasakhodi:2021:CEB

[2331] Maliheh Taheri Fasakhodi, Amirreza Abed-Elmdoust, Alireza Mirvaghefi, Seyed Vali Hosseini, and Kamran Rezaei Tavabe. Changes of extracted bioactive compounds from brown algae (Cystoseira indica) after conversion to mill and tablet using a quantitative metabolomics approach. Aquaculture International, 29(6):2793–2804, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-021-00779-
2.

Ren:2021:RSF

[2332] Jindong Ren, Rong Na, Honglin Chen, Bao Lou, and Baolong Niu. RNA sequencing and functional analysis of adult gonadal tissue to identify candidate key genes in Macrobrachium rosenbergii sex development. Aquaculture International, 29(6):2805–2821, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00780-9.

Sun:2021:IRP

[2333] Wei Sun, Na Rong, Sijie Jian, Chao Kang, Chunlin Chen, Rui Chen, Rui Ding, Chen Chen, and Xiang Liu. Immune responses and protective efficacy of outer membrane protein ExbB of Pseudomonas fluorescens against Aeromonas hydrophila and Pseudomonas fluorescens affecting Carassius auratus. Aquaculture International, 29(6):2823–2840, December 2021. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00784-5.

Ma:2022:GHG

[2334] Fang Ma, Lin Tong Luo, and Qianning Wang. Hsp60/10 and sHsp families of heat shock protein genes in rainbow trout (Oncorhynchus mykiss) and their expression under heat stress. Aquaculture International, 30 (1):1–18, February 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-020-00638-6.

Huang:2022:DSA

[2335] Zhangfan Huang, Youling Ye, Anle Xu, Zhongbao Li, and Zhen Wang. Dietary supplementation with an acidifier blend (citric, lactic, and phosphoric acids) influences growth, digestive enzymes, and blood chemistry of juvenile Japanese sea-bass (Lateolabrax japonicus). Aquaculture International, 30(1):19–32, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00703-8.

Chen:2022:IIB

[2336] Jianjun Chen, Yidi Zhao, Dandan Sun, Chenyang Rao, Lulu Li, Suqi Guo, Shuai Yang, and Xianglin Cao. Improvement of intestinal barrier, immunity, and meat quality in common carp infected by Aeromonas hydrophila using probiotics. Aquaculture International, 30(1):33–49, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00769-4.

Zafar:2022:BTE

[2337] Md Abu Zafar and Md Masud Rana. Biofloc technology: an eco-friendly “green approach” to boost up aquaculture production. Aquaculture International, 30(1):51–72, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00781-8.

Falahatkar:2022:EVH

[2338] Bahram Falahatkar, Kamyar Javid Rahmdel, Eshagh Rasouli Kargar, and Shapoor Gholami. Evaluation of various hCG treatment strategies applied to domesticated pikeperch (Sander lucioperca) broodstock on nest-spawning performance. Aquaculture International, 30(1):73– 85, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00785-4.

Wang:2022:DAM

[2339] Yude Wang, Kejie Chen, Dishan Hong, and Duansheng Wu. Development and application of microsatellite DNA molecular markers for the laboratory red crucian carp (Carassius auratus) strain C1HD. Aquaculture International, 30(1):87–97, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00786-3.

Ranjan:2022:SMU

[2340] Amit Ranjan, Sarvendra Kumar, Narottam Prasad Sahu, Kamal Kant Jain, and Ashutosh Dharmendra Deo. Strategies for maximizing utilization of de-oiled rice bran (DORB) in the fish feed. Aquaculture International, 30(1):99–114, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00791-6.

Liang:2022:SAM

[2341] Chao Liang, Shuanghui Huang, Yi Geng, Xiaoli Huang, Defang Chen, Weiming Lai, Hongrui Guo, Huidan Deng, Jing Fang, Lizi Yin, and Ping Ouyang. A study on the antibacterial mechanism of thymol against Aeromonas hydrophila in vitro. Aquaculture International, 30 (1):115–129, February 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00789-0.

Dehkordi:2022:PPG

[2342] Seyed Majid Hashemi Dehkordi, Seyed Amirali Anvar, Ebrahim Rahimi, Hamed Ahari, and Maryam Ataee. Prevalence, phenotypic and genotypic diversity, antibiotic resistance, and frequency of virulence genes in Pseudomonas aeruginosa isolated from shrimps. Aquaculture International, 30(1):131–156, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00798-z.

Kumaran:2022:TEE

[2343] M. Kumaran, K. P. Kumaraguru Vasagam, R. Subburaj, P. R. Anand, K. Ramachandran, R. Geetha, D. Deboral Vimala, R. Anandaraja, M. Jayanthi, and C. V. Sairam. Techno-economic evaluation of Asian seabass (Lates calcarifer) nursery rearing in small net cages (hapas) under different coastal salinities. Aquaculture International, 30(1):157– 172, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00788-1.

Boamah:2022:FHR

[2344] Grace Afumwaa Boamah, Feng Yu, Yawei Shen, Weiwei You, Changan Xu, Xuan Luo, and Caihuan Ke. Fluctuations in the heart rate of abalone in response to low salinity stress. Aquaculture International, 30 (1):173–186, February 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00790-7.

Maitland:2022:PGD

[2345] Danielle M. Maitland, Joe Baker, Greg Chambers, Neil W. Ross, and Stefanie M. Colombo. Population growth dynamics and their implications for fish welfare in mixed-size cohorts of Cyprinus carpio var. koi grown in a commercial-scale aquaponics system. Aquaculture International, 30 (1):187–210, February 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00792-5.

Nugraha:2022:FEF

[2346] Tito Aria Nugraha, Alim Isnansetyo, Triyanto, and Mukhlisna Djalil. Fermented earthworms as a feed additive enhances non-specific immune response in catfish (Clarias gariepinus). Aquaculture International, 30 (1):211–226, February 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00794-3.

Mondal:2022:VIC

[2347] Haimanti Mondal, Natarajan Chandrasekaran, Amitava Mukherjee, and John Thomas. Viral infections in cultured fish and shrimps: current status and treatment methods. Aquaculture International, 30(1): 227–262, February 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00795-2.

Gonzales-Flores:2022:EHE

[2348] Anai P. Gonzales-Flores, Felix I. Mejia Perez, Katty A. Huanuiri Quinteros, Ingrid Y. Sanchez Callejas, Jorge L. Vasquez Rojas, and Christian Fernandez-Mendez. Effect of heparin and EDTA as anticoagulants on hematological values in farmed juvenile of Arapaima gigas. Aquaculture International, 30(1):263–271, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00796-1.

Shah:2022:SDP

[2349] Yashesh Shah, Manali Rathod, Monica G. Kavale, Santlal Jaiswar, and Vaibhav A. Mantri. Socio-demographic profiling and asset indicators of Gracilaria dura farmers from northern west coast of India useful for longitudinal analysis. Aquaculture International, 30(1): 273–287, February 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00797-0.

Rahman:2022:PPI

[2350] Md Sadique Rahman, Mohammad Mizanul Haque Kazal, Shah Johir Rayhan, and Md Hayder Khan Sujan. Potential for productivity improvements by implementing improved management practices in pond polyculture of Indian major carps in Bangladesh. Aquaculture International, 30(1):289–303, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00799-y.

Greenfeld:2022:AGE

[2351] Asael Greenfeld, Nir Becker, Janet F. Bornman, Sabrina Spatari, and Dror L. Angel. Is aquaponics good for the environment? — evaluation of environmental impact through life cycle assessment studies on aquaponics systems. Aquaculture International, 30(1):305–322, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00800-8.

Guilherme:2022:FGB

[2352] Helder O. Guilherme, Fabio A. C. Santos, Leandro S. Costa, Verônica G. L. Prado, Glauber D. A. Palheta, Nuno F. A. C. de Melo, Ronald K. Luz, and Paula A. P. Ribeiro. Feeding, growth, and blood chemistry of the tambaqui (Colossoma macropomum) held under self-feeding and time-restricted automatic feeding conditions. Aquaculture International, 30(1):323–339, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-021-00801-7.

Rathinam:2022:SMR

[2353] R. Bharathi Rathinam, S. Abuthagir Iburahim, S. Suresh Ramanan, and Gayatri Tripathi. A scientometric mapping of research on Aeromonas infection in fish across the world (1998–2020). Aquaculture International, 30(1):341–363, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00802-6.

Ha:2022:PPS

[2354] Natalia Ha, Luiz Augusto Cipriani, Nandara Soares de Oliveira, Juliano Uczay, Marcos Luiz Pessatti, and Thiago El Hadi Perez Fabregat. Peptide profile of the sardine protein hydrolysate affects food utilization and intestinal microbiota of Nile tilapia. Aquaculture International, 30(1): 365–382, February 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00804-4.

Khanjani:2022:BTA

[2355] Mohammad Hossein Khanjani and Moslem Sharifinia. Biofloc technology with addition molasses as carbon sources applied to Litopenaeus vannamei juvenile production under the effects of different C/ N ratios. Aquaculture International, 30(1):383–397, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00803-5.

Keivanloo:2022:PID

[2356] Saeide Keivanloo, Mohammad Sudagar, Abasali Hajibeglou, and Moradmohammad Shakiba. Preliminary investigation to determine the toxicity of various cryoprotectants on striped gourami (Trichogaster fasciata) embryos. Aquaculture International, 30(1):399–408, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-021-00809-z.

Raknak:2022:TMA

[2357] Kantinan Raknak, Panyawut Rattanarom, and Nalena Praphairaksit. Transformation of mugwort (Artemisia vulgaris) oil into nanoemulsion as an ethanol-free herbal anesthetic for surgery on koi carp (Cyprinus carpio). Aquaculture International, 30(1):409–418, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00808-0.

Cagatay:2022:FCT

[2358] Ifakat Tülay Çağatay. FTA(R) card tool for sampling and rapid diagnosis of bacterial diseases from rainbow trout (Oncorhynchus mykiss) tissue. Aquaculture International, 30(1):419–428, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00810-6.

Miyoshi:2022:PRM

[2359] Koji Miyoshi and Susumu Chiba. Predation risk management of sea stars (Asterias amurensis and Distolasterias nipon) by adjusting the density and size of seeded scallops (Mizuhopecten yessoensis): an improvement to local mariculture. Aquaculture International, 30(1):429– 443, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00805-3.

Arifa:2022:EFP

[2360] Arifa, Mst Khadiza Begum, Raad Mozib Lalon, A. B. M. Shamsul Alam, and Mohammad Shamsur Rahman. Economic feasibility of Pabda and stinging catfish culture in recirculating aquaculture systems (RAS) in Bangladesh. Aquaculture International, 30(1):445–465, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00807-1.

Wan:2022:CGP

[2361] Jin-Juan Wan, Jian lin Pan, Mei-Fang Shen, Hui Xue, Meng lin Sun, Mei-Qin Zhang, Xi-He Zhu, and Xing kong Ma. Changes in the growth performance, antioxidant enzymes and stress resistance caused by dietary administration of synbiotic (fructooligosaccharide and probiotics) in juvenile Chinese mitten crab, Eriocheir sinensis. Aquaculture International, 30(1):467–481, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00811-5.

Li:2022:CCT

[2362] Yun Li, Qifei Zhu, Yong Huang, Qi Xu, Xilin Dai, and Chenxi Ju. Cloning and characterization of two types of growth hormone receptors in tomato clownfish (Amphiprion frenatus), and their expression under different light spectra and photoperiods. Aquaculture International, 30 (1):483–500, February 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00814-2.

Furlan-Murari:2022:IGO

[2363] Pâmela Juliana Furlan-Murari, Ed Christian Suzuki de Lima, Felipe Pinheiro de Souza, Angela Maria Urrea-Rojas, Andreia Carla Eugenio Pupim, Eduardo José de Almeida Araújo, Paulo César Meletti, Cindy Namie Seino Leal, Letícia Lima Fernandes, and Nelson Mauricio Lopera-Barrero. Inclusion of β-1,3/1,6-glucan in the ornamental fish, jewel tetra (Hyphessobrycon eques), and its effects on growth, blood glucose, and intestinal histology. Aquaculture International, 30(1):501– 515, February 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00815-1.

Suarez-Caceres:2022:CAP

[2364] Gina Patricia Suárez-Cáceres, Víctor M. Fernández-Cabanás, José Lobillo-Eguíbar, and Luis Pérez-Urrestarazu. Characterisation of aquaponic producers and small-scale facilities in Spain and Latin America. Aquaculture International, 30(2):517–532, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00793-4.

Luo:2022:EAE

[2365] Yaqi Luo, Chenchen Shen, Dan Tang, Yuze Bai, Lv Wu, Yiping Zhang, Ya Wu, and Zhengfei Wang. The effects of ammonia exposure on immune response, oxidative stress, and apoptosis in Procambarus clarkii. Aquaculture International, 30(2):533–546, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00828-w.

Dantas:2022:EEP

[2366] Daniela P. Dantas, Dallas L. Flickinger, Gelcirene A. Costa, Patricia Moraes-Valenti, and Wagner C. Valenti. Economic effects of production scale, use of agricultural greenhouses, and integration of tropical aquaculture species when farming in a subtropical climate. Aquaculture International, 30(2):547–579, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00822-2.

Godoy-Olmos:2022:IDF

[2367] Sergio Godoy-Olmos, Ignacio Jauralde, Raquel Monge-Ortiz, María C. Milián-Sorribes, Miguel Jover-Cerdá, Ana Tomás-Vidal, and Silvia Martínez-Llorens. Influence of diet and feeding strategy on the performance of nitrifying trickling filter, oxygen consumption and ammonia excretion of gilthead sea bream (Sparus aurata) raised in recirculating aquaculture systems. Aquaculture International, 30(2):581– 606, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-021-00821-3.

Yang:2022:MCE

[2368] Yuting Yang, Zhongdian Dong, Xi Chen, Zhen Wang, Dawei Zhang, Liqun Liang, and Weijie Mu. Molecular characterization and expression analysis of hypoxia-inducible factor-1α, factor-2α, and factor-3α and physiological response to hypoxia exposure in Amur minnow (Phoxinus lagowskii). Aquaculture International, 30(2):607–632, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-021-00826-y.

Rahmati-Holasoo:2022:PFF

[2369] Hooman Rahmati-Holasoo, Amin Marandi, Hosseinali Ebrahimzadeh Mousavi, and Ali Taheri Mirghaed. Parasitic fauna of farmed freshwater ornamental fish in the northwest of Iran. Aquaculture International, 30(2):633–652, April 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00832-0.

Abbaszadeh:2022:EAC

[2370] Akbar Abbaszadeh, Mansour Torfi Mozanzadeh, Ahmad Qasemi, Amin Oujifard, and Mahmod Nafisi Bahabadi. Effects of the addition of Calanopia elliptica, Artemia franciscana, and Brachionus rotundiformis in a nursery biofloc system on water quality, growth, gut morphology, health indices, and transcriptional response of immune and antioxidantrelated genes in Penaeus vannamei. Aquaculture International, 30(2): 653–676, April 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00823-1.

Khieokhajonkhet:2022:SIM

[2371] Anurak Khieokhajonkhet, Niran Aeksiri, Jiraporn Rojtinnakorn, Hien Van Doan, and Gen Kaneko. Sacha inchi meal as a fish-meal replacer in red hybrid tilapia (Oreochromis niloticus × O. mossambicus) feeds: effects on dietary digestibility, growth metrics, hematology, and liver and intestinal histology. Aquaculture International, 30 (2):677–698, April 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00833-7.

Shehata:2022:ERA

[2372] Akram Ismael Shehata, Mayada Alhoshy, Tao Wang, Jianfeng Wang, Ruoxuan Wang, Mahmoud A. O. Dawood, Mohamed A. A. Zaki, Yilei Wang, and Ziping Zhang. Expression of reproduction and antioxidantrelated genes in crayfish Cherax quadricarinatus fed with dietary feed additives. Aquaculture International, 30(2):699–720, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00816-0.

Sun:2022:IFR

[2373] Lishuang Sun, Anni Ni, Yulian Tang, Shu Li, and Genliang Li. Involvement of the fusome-related protein add1 in spermatogenesis of the Chinese mitten crab (Eriocheir sinensis) by organization of cytoskeleton. Aquaculture International, 30(2):721–733, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00825-z.

daSilvaCerozi:2022:NRC

[2374] Brunno da Silva Cerozi, Josué Wenceslau de Oliveira Neto, Roger Abraão Ribeiro Cardeal Dias Leonel, and Angélica Priscila do Carmo Alves. Nutrient release coefficient: a proposed batch reactor assay to determine the elemental composition of aquaponic nutrient solutions. Aquaculture International, 30(2):735–746, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00817-z.

Zhang:2022:EPP

[2375] Yurou Zhang, Zhenchao Wu, Mengnan Yu, Dongming Zhang, Qiuju Wang, Lili Lin, Guiqin Wang, Mahmoud M. Elsadek, Qi Yao, Yuke Chen, and Zhixin Guo. Evaluating the probiotic potential and adhesion characteristics of Bacillus spp. isolated from the intestine of Rhynchocypris lagowskii Dybowski. Aquaculture International, 30(2):747– 772, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00840-8.

Heal:2022:UEF

[2376] Richard Heal, Mohammad Mahfujul Haque, Neaz A. Hasan, Joseph Nagoli, Syed Arifuzzaman, Charles R. Tyler, and David Bass. Understanding the economic and farming practices driving species selection in aquaculture within the Mymensingh division of Bangladesh. Aquaculture International, 30(2):773–789, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00818-y.

Martelli:2022:EDA

[2377] Antonela Martelli, Melania Fernandez, Cynthia Sequeiros, and Pedro J. Barón. Effects of dietary administration of two Patagonian probiotics on the zoeae I from the swimming crab Ovalipes trimaculatus. Aquaculture International, 30(2):791–802, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00836-4.

Huang:2022:CEA

[2378] Jian sheng Huang, Eric Amenyogbe, Lin tong Yang, Zhong liang Wang, Gang Chen, Wei zheng Wang, and Jian dong Zhang. Cloning and expression analysis of hif-1α and downstream genes during hypoxic stress in cobia (Rachycentron canadum). Aquaculture International, 30 (2):803–824, April 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00820-4.

Younis:2022:EHA

[2379] Abuelhassan Elshazly Younis, Awatef Hamed Hamouda, and Eman Moustafa Moustafa. Euclinostomum heterostomum and E. ardeolae in tilapia species of Aswan Governorate, Egypt: morphological, molecular, and histopathological characterization. Aquaculture International, 30(2):825– 844, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00839-1.

Elgendy:2022:MCV

[2380] Mamdouh Y. Elgendy, Mohamed Abdelsalam, Samaa Ahmed Mohamed, and Shimaa E. Ali. Molecular characterization, virulence profiling, antibiotic susceptibility, and scanning electron microscopy of Flavobacterium columnare isolates retrieved from Nile tilapia (Oreochromis niloticus). Aquaculture International, 30(2):845–862, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00819-x.

Lai:2022:CLS

[2381] Quan T. Lai, Vu Anh Tuan, Nguyen Thi Bich Thuy, Le Dinh Huynh, and Nguyen Minh Duc. A closer look into shrimp yields and mangrove coverage ratio in integrated mangrove-shrimp farming systems in Ca Mau, Vietnam. Aquaculture International, 30(2):863–882, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00831-1.

Wang:2022:EDS

[2382] Xin Wang, Qi Yao, Xin yu Lei, Dong ming Zhang, Sen Wang, Ji wu Wan, Hong jian Liu, Yu ke Chen, Yun long Zhao, Gui qin Wang, Qiu ju Wang, and Zhi xin Guo. Effects of different stocking densities on the growth performance and antioxidant capacity of Chinese mitten crab (Eriocheir sinensis) in rice crab culture system. Aquaculture International, 30 (2):883–898, April 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00829-9.

Syazwina:2022:BPB

[2383] S. Siti Syazwina, E. Mohammad Basri, and W. O. Wan Maznah. Biomass, proximate biochemical composition and fatty acid profiles associated with the growth phase of Chlorella salina Butcher and Isochrysis maritima Billard and Gayral isolated from the coastal waters of Penang, Malaysia. Aquaculture International, 30(2):899–918, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00843-5.

StLouis:2022:CFD

[2384] Tracy Jeanel St. Louis, Manoel Xavier Pedroza Filho, and Roberto Manolio Valladao Flores. Consumption frequencies, determinants, and habits of aquaculture species in Brazil. Aquaculture International, 30(2):919– 936, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00838-2.

Cole:2022:EOB

[2385] Anthony J. Cole, Smita S. Tulsankar, Benjamin J. Saunders, and Ravi Fotedar. Effects of an oil-based substrate (The Water Cleanser^{T M}) and bacterial additives on nitrogen and phosphorous dynamics in freshwater crayfish (Cherax cainii, Austin and Ryan 2002) aquaculture. Aquaculture International, 30(2):937–954, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00842-6.

Knoll:2022:SBP

[2386] Lauren Knoll and Mazz Marry. Structural and biophysical properties of whole leaf and root tissue and isolated cell walls of common green bean and tomato seedlings grown in an aquaponics system relative to soil-grown counterparts. Aquaculture International, 30(2):955– 988, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00846-2.

Karthikeyan:2022:AVA

[2387] K. Karthikeyan, A. T. Manish Kumar, D. Harshini, G. Pooja, J. Arul Daniel, Pavani Yasam, G. Gracevictoria, M. Priayanka, P. Krishnaraj, R. Vidya, S. Asha Devi, Sabbasani Rajasekhara Reddy, and Raja Sudhakaran. Anti-viral activity of methyl 1-chloro-7-methyl-2-propyl-1h-benzo[d] imidazole-5-carboxylate against white spot syndrome virus in freshwater crab (Paratelphusa hydrodromous). Aquaculture International, 30(2):989–998, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-021-00714-5.

Bahrioglu:2022:RDF

[2388] Ergi Bahrioğlu and Ertan Ercan. Replacement of dried freshwater alga Arthrospira maxima with marine diatom Schizochytrium limacinum in a diet of freshwater mussel Unio crassus (Philipsson, 1788). Aquaculture International, 30(2):999–1010, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00852-4.

Baliarsingh:2022:MCS

[2389] Snigdha Baliarsingh, Sonalina Sahoo, Yong Hun Jo, Yeon Soo Han, Arup Sarkar, Yong Seok Lee, Jyotirmaya Mohanty, and Bharat Bhusan Patnaik. Molecular cloning, sequence characterization, and expression analysis of C-type lectin (CTL) and ER–Golgi intermediate compartment 53-kDa protein (ERGIC-53) homologs from the freshwater prawn, Macrobrachium rosenbergii. Aquaculture International, 30(2): 1011–1035, April 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00845-3.

Mirmazloomi:2022:CCV

[2390] Sina Mirmazloomi, Maryam Ghiasi, and Ali Reza Khosravi. Chemical composition and in vitro antifungal activity of Sambucus ebulus and Actinidia deliciosa on the fish pathogenic fungus, Saprolegnia parasitica. Aquaculture International, 30(2):1037–1046, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00835-5.

Shamsuddin:2022:ABT

[2391] Md Shamsuddin, M. Belal Hossain, Moshiur Rahman, Mst. Salamun Kawla, Md. Badiul Alam Shufol, Muhammad Mamunur Rashid, Md. Asadujjaman, and Md. Refat Jahan Rakib. Application of biofloc technology for the culture of Heteropneustes fossilis (Bloch) in Bangladesh: stocking density, floc volume, growth performance, and profitability. Aquaculture International, 30(2):1047–1070, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00849-z.

Jeyavani:2022:PCN

[2392] Jeyaraj Jeyavani, Ashokkumar Sibiya, Jeyachandran Sivakamavalli, Mani Divya, Elumalai Preetham, Baskaralingam Vaseeharan, and Caterina Faggio. Phytotherapy and combined nanoformulations as a promising disease management in aquaculture: a review. Aquaculture International, 30(2):1071–1086, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00848-0.

Rioja:2022:CAE

[2393] Rose Angeli Rioja, Nadia Palomar-Abesamis, and Marie Antonette Juinio-Meñez. Correction to: Associated effects of shading on the behavior, growth, and survival of Stichopus cf. horrens juveniles. Aquaculture International, 30(2):1087–1088, April 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00768-5. See [2285].

Debnath:2022:ESD

[2394] Sudip Debnath, Moin Uddin Ahmed, Md. Shahin Parvez, Anup Kumar Karmokar, and Md. Nazmul Ahsan. Effect of stocking density on growth performance and body composition of climbing perch (Anabas testudineus) in biofloc system. Aquaculture International, 30(3): 1089–1100, June 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00812-4.

Long:2022:ETF

[2395] Xiaowen Long, Jie Pan, Nicholas M. Wade, Yunfei Sun, Yuming Liu, Yongxu Cheng, and Xugan Wu. Effects of three feeding modes on the metabolism, antioxidant capacity, and metabolome of the adult male Chinese mitten crab Eriocheir sinensis. Aquaculture International, 30 (3):1101–1119, June 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00782-7.

Sridhar:2022:CAO

[2396] Arun Sridhar, Francisco A. Guardiola, Rajkumar Krishnasamy Sekar, Sathiya Deepika Murugesan, Sivagaami Palaniyappan, Dinesh Babu Manikandan, Manikandan Arumugam, and Thirumurugan Ramasamy. Comparative assessment of organic solvent extraction on non-specific immune defences of skin mucus from freshwater fish. Aquaculture International, 30(3):1121–1138, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00847-1.

Wang:2022:CTA

[2397] Teng Wang, Lei Gao, Wenheng Li, Yun Li, and Hongwei Shan. Comparative transcriptome analysis reveals a strategy involving dietary manipulation for reducing the mortality of Litopenaeus vannamei exposed to sublethal ammonia through the energy metabolism pathway. Aquaculture International, 30(3):1139–1154, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00855-1.

Dagoudo:2022:SST

[2398] Missinhoun Dagoudo, Jun Qiang, and Maurice P. Solevo. Status in science and technology developments in Benin’s aquaculture industry: a review. Aquaculture International, 30(3):1155–1169, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00853-3.

Defaee:2022:EPF

[2399] Somayeh Defaee, Bahram Falahatkar, Fatemeh Lavajoo, and Iraj Efatpanah. The effect of a phytogenic feed additive (Digestrom P.E.P) on growth performance, proximate composition, hematological and immunological indices of juvenile beluga sturgeon Huso huso. Aquaculture International, 30(3):1171–1183, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-021-00824-0. See correction

[2455].

Moawad:2022:RAG

[2400] Madelyn N. Moawad, Mary Ghobrial, and Soha Shabaka. The red alga Grateloupia gibbesii, as a valuable source of lipids: lipids quality indices, spectroscopic characterization, and potential industrial applications. Aquaculture International, 30(3):1185–1209, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00813-3.

Preena:2022:QHF

[2401] Prasannan Geetha Preena, Thaliyil Veetil Arun Kumar, Tina Kollannoor Johny, Arathi Dharmaratnam, and Thangaraj Raja Swaminathan. Quick hassle-free detection of cyprinid herpesvirus 2 (CyHV-2) in goldfish using recombinase polymerase amplification-lateral flow dipstick (RPA-LFD) assay. Aquaculture International, 30(3):1211–1220, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-021-00806-2.

Vignesh:2022:ECF

[2402] Samayanpaulraj Vignesh, Govindaraj Krishnaveni, J. C. Walter Devaa, Sivaramapillai Muthukumar, and Ramesh Uthandakalaipandian. Experimental challenge of the freshwater fish pathogen Aeromonas hydrophila Ah17 and its effect on snakehead murrel Channa striata. Aquaculture International, 30(3):1221–1238, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00856-0.

Li:2022:MHP

[2403] Mengmeng Li, Dongdong Wei, Shuaishuai Huang, Lin Huang, Fengqiao Xu, Qing Yu, Mingzhu Liu, and Pengfei Li. Medicinal herbs and phytochemicals to combat pathogens in aquaculture. Aquaculture International, 30(3):1239–1259, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00841-7.

Armuelles-Bernal:2022:EGE

[2404] Candy Armuelles-Bernal, Lidia Cucala-García, Benjamín Barón-Sevilla, Lourdes Díaz-Jiménez, Miguel Rubio-Godoy, and Mónica Hernández-Rodríguez. Effect of garlic extract and diallyl sulfide immersion baths in the life stages of Zeuxapta seriolae, a parasite of California yellowtail (Seriola dorsalis). Aquaculture International, 30(3):1261–1277, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00844-4.

Meusel:2022:GDF

[2405] Esther Meusel, Simon Menanteau-Ledouble, Matthew Naylor, Horst Kaiser, and Mansour El-Matbouli. Gonad development in farmed male and female South African abalone, Haliotis midae, fed artificial and natural diets under a range of husbandry conditions. Aquaculture International, 30(3):1279–1293, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00850-6.

deMoraes:2022:MFP

[2406] Laenne Barbara Silva de Moraes, Rudã Fernandes Brandão Santos, Genes Fernando Gonçalves Junior, Géssica Cavalcanti Pereira Mota, Danielli Matias de Macêdo Dantas, Ranilson de Souza Bezerra, and Alfredo Olivera Gálvez. Microalgae for feeding of penaeid shrimp larvae: an overview. Aquaculture International, 30(3):1295–1313, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00857-z.

Nogueira:2022:PAB

[2407] Wesclen Vilar Nogueira, Francisco Javier Moyano, María Jesús Aznar García, Marcelo Borges Tesser, and Jaqueline Garda Buffon. Preliminary assessment of bioaccessibility of aflatoxin B_{1} in fish. Aquaculture International, 30(3):1315–1325, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00860-4.

Liu:2022:EIE

[2408] Jian Liu, Yunfei Sun, Wenfeng Han, Jinghao Li, Shihui Wang, Zhigang Yang, and Yongxu Cheng. Evaluation of the inhibitory effects of four different microecological preparations on Cladophora. Aquaculture International, 30(3):1327–1340, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-021-00783-6.

Sebastiao:2022:ECS

[2409] Fernanda de Alexandre Sebastião, Maria Juliete Souza Rocha, Franmir Rodrigues Brandão, Maria Inês Braga de Oliveira, Damy Caroline de Melo Souza, Bruna Carlos do Nascimento Barbosa, Patrícia Castro Monteiro, Cláudia Majolo, Roger Crescêncio, Marcos Tavares-Dias, and Edsandra Campos Chagas. Evaluation of the clinical safety and efficacy of fenbendazole and levamisole in the control of Neoechinorhynchus buttnerae in Colossoma macropomum. Aquaculture International, 30(3): 1341–1351, June 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00854-2.

Angwin:2022:GEP

[2410] Renee E. Angwin, Brian T. Hentschel, and Todd W. Anderson. Gonad enhancement of the purple sea urchin, Strongylocentrotus purpuratus, collected from barren grounds and fed prepared diets and kelp. Aquaculture International, 30(3):1353–1367, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00863-1.

Liu:2022:SPD

[2411] Qian Liu, Junnan Li, Hongwei Shan, Yicheng Xie, and Dongxu Zhang. Specific patterns and drivers of the bacterial communities in the sediment of two typical integrated multitrophic aquaculture systems. Aquaculture International, 30(3):1369–1388, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00862-2.

Sequeiros:2022:ZIC

[2412] Cynthia Sequeiros, Marisa E. Garcés, Melania Fernández, Magali Marcos, Cecilia Castaños, Mariano Moris, and Nelda L. Olivera. Zebrafish intestinal colonization by three lactic acid bacteria isolated from Patagonian fish provides evidence for their possible application as candidate probiotic in aquaculture. Aquaculture International, 30(3):1389– 1405, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00864-0.

Ahsan:2022:AHM

[2413] Md. Emranul Ahsan, Seikh Razibul Islam, Muhammad Abdur Razzak, Md. Lokman Ali, and Mohammad Mahfujul Haque. Assessment of heavy metals from pangasius and tilapia aquaculture in Bangladesh and human consumption risk. Aquaculture International, 30(3):1407– 1434, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00903-w.

Amoussou:2022:CFA

[2414] Nellya Amoussou, Michel Marengo, Ogouyôm Herbert Iko Afé, Pierre Lejeune, Éric Dominique Henri Durieux, Caroline Douny, Marie-Louise Scippo, and Sylvie Gobert. Comparison of fatty acid profiles of two cultivated and wild marine fish from Mediterranean Sea. Aquaculture International, 30(3):1435–1452, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00861-3.

Gouife:2022:PDS

[2415] Moussa Gouife, Suyang Chen, Kejing Huang, Mateen Nawaz, Shan Jin, Rongrong Ma, Yajun Wang, Liangyi Xue, and Jiasong Xie. Photobacterium damselae subsp. damselae in mariculture. Aquaculture International, 30(3):1453–1480, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-022-00867-x.

Hoerterer:2022:SFF

[2416] Christina Hoerterer, Jessica Petereit, Gisela Lannig, Johan Johansen, Gabriella V. Pereira, Luis E. C. Conceição, Roberto Pastres, and Bela H. Buck. Sustainable fish feeds: potential of emerging protein sources in diets for juvenile turbot (Scophthalmus maximus) in RAS. Aquaculture International, 30(3):1481–1504, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00859-x.

Mantri:2022:HPA

[2417] Vaibhav A. Mantri, Ramalingam Dineshkumar, Anshul Yadav, V. Veeragurunathan, M. Ganesan, K. Eswaran, and S. Thiruppathi. How profitability assessment parameters score under large-scale commercial cultivation of different agarophyte seaweeds along south-eastern coast of India. Aquaculture International, 30(3):1505–1525, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00866-y.

Du:2022:REG

[2418] Jinliang Du, Rui Jia, Liping Cao, Zhengyan Gu, Qin He, Pao Xu, Guojun Yin, and Yuzhong Ma. Regulatory effects of Glycyrrhiza total flavones on fatty liver injury induced by a high-fat diet in tilapia (Oreochromis niloticus) via the Nrf2 and TLR signaling pathways. Aquaculture International, 30(3):1527–1548, June 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00787-2.

Ferreira:2022:EHR

[2419] Andre Lima Ferreira, Fábio Aremil Costa dos Santos, André de Sena Souza, Gisele Cristina Favero, Bernardo Baldisserotto, Carlos Garrido Pinheiro, Berta Maria Heinzmann, and Ronald Kennedy Luz. Efficacy of Hesperozygis ringens essential oil as an anesthetic and for sedation of juvenile tambaqui (Colossoma macropomum) during simulated transport. Aquaculture International, 30(3):1549–1561, June 2022. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00868-w.

Xu:2022:EDP

[2420] Cong mei Xu, Hai rui Yu, Qin Zhang, Ling yao Li, Jing jing Fan, Meng jie Guo, Fang hui Li, Xiang yi Qiu, and Ling ling Shan. Effects of dietary protein concentration on growth and anti-oxidant capacity of Coho salmon Oncorhynchus kisutch (Walbaum, 1792) alevins. Aquaculture International, 30(3):1563–1574, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00837-3.

Adeleke:2022:PLC

[2421] Babatunde Adeleke, Shahida Cassim, and Simon Taylor. Pathways to low-cost aquaponic systems for sustainable livelihoods and economic development in poor communities: defining critical success factors. Aquaculture International, 30(3):1575–1591, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00865-z.

Roy:2022:AGA

[2422] Subha M. Roy, Mohammad Tanveer, and Rajendra Machavaram. Applications of gravity aeration system in aquaculture — a systematic review. Aquaculture International, 30(3):1593–1621, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00851-5.

Boateng:2022:DAA

[2423] Charles Narteh Boateng, Austin Mtethiwa, and Seth Koranteng Agyakwah. Drivers of aquaculture adoption and disadoption: the case of pond aquaculture in Ghana. Aquaculture International, 30(3):1623– 1643, June 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00858-y.

delValle:2022:MCP

[2424] Juana Cristina del Valle, Aldo Nahuel Zanazzi, Yamila Eliana Rodriguez, Nora Selma Haran, María Victoria Laitano, Juan Carlos Mallo, and Analía Verónica Fernández-Gimenez. Morphological changes, peptidase activity, and effects of exogenous enzymes in the early ontogeny of Nile tilapia, Oreochromis niloticus. Aquaculture International, 30(4): 1645–1658, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00932-5.

Sui:2022:EIR

[2425] Juan Sui, Sheng Luan, Guoliang Yang, Zhenglong Xia, Qiongying Tang, Kun Luo, Xianhong Meng, and Jie Kong. Effects of the individual rearing stage on the growth traits of candidate giant freshwater prawns (Macrobrachium rosenbergii). Aquaculture International, 30(4): 1659–1673, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-021-00827-x.

Li:2022:DPR

[2426] Meifeng Li, Li Zhang, Bing Hu, Lei Liu, Feng Huang, Juan Tian, Xianqin Hu, and Yuanyuan Wang. Dietary phosphorus requirement for juvenile bighead carp (Aristichthys nobilis). Aquaculture International, 30(4): 1675–1692, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00834-6.

Chen:2022:EMM

[2427] Kai Chen, Gladstone Sagada, Bingying Xu, Yuechong Liu, Lu Zheng, Arnaud Fabrice Tegomo, Yifei Yang, Yuxiao Sun, Jia Wang, Matt Longshaw, and Qingjun Shao. Evaluation of methanotroph (Methylococcus capsulatus, bath) bacteria meal as an alternative protein source for growth performance, digestive enzymes, and health status of Pacific white shrimp (Litopenaeus vannamei). Aquaculture International, 30(4): 1693–1710, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00869-9.

Kumar:2022:TAE

[2428] Munish Kumar, Gyandeep Gupta, Nuzaiba P. Muhammed, Karthik R, Tincy Varghese, Prem Prakash Srivastava, Shashi Bhushan, Satya Prakash Shukla, Gopal Krishna, and Subodh Gupta. Toxicity ameliorative effect of vitamin E against super-paramagnetic iron oxide nanoparticles on haemato-immunological responses, antioxidant capacity, oxidative stress, and metabolic enzymes activity during exposure and recovery in Labeo rohita fingerlings. Aquaculture International, 30 (4):1711–1739, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00870-2.

Mazon-Suastegui:2022:SDE

[2429] José Manuel Mazón-Suástegui, Guadalupe Fabiola Arcos-Ortega, César Lodeiros, Dariel Tovar-Ramírez, and Jesús Antonio López-Carvallo. Stocking density evaluation on Catarina scallop (Argopecten ventricosus, Sowerby II, 1842) larvae to improve hatchery production. Aquaculture International, 30(4):1741–1754, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00873-z.

Chen:2022:DNS

[2430] Xiaochun Chen, Xiaoquan Huang, Yiwen Tang, Lei Zhang, and Feng Lin. Dietary nucleotide supplementation enhances the growth, immune response, and disease resistance of the juvenile giant freshwater prawn, Macrobrachium rosenbergii. Aquaculture International, 30(4): 1755–1768, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00871-1.

Xu:2022:PAS

[2431] Hanying Xu, Juan Dou, Qingyang Wu, Yangfang Ye, Chunlin Wang, Changbin Song, Changkao Mu, Zhiming Ren, and Ce Shi. Photoperiod affects the survival rate but not the development of larval swimming crab Portunus trituberculatus. Aquaculture International, 30(4): 1769–1778, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00875-x.

Wang:2022:EHP

[2432] Yang Wang, Mingzhu Sun, Yizhen Tang, Ailing Xu, Jingchao Tang, and Zhiwen Song. Effects of Haematococcus pluvialis on the water quality and performance of Litopenaeus vannamei using artificial substrates and water exchange systems. Aquaculture International, 30(4): 1779–1797, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00872-0.

Alias:2022:SSA

[2433] Sharifah Lia Farliana Wan Alias, Mohammad Bodrul Munir, Roslianah Asdari, Lein En Yao, and Chua Sing Ying. Selection of suitable aquaponics system for empurau (Tor tambroides) fries nursery in polyculture method. Aquaculture International, 30(4):1799–1816, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00874-y.

Estrada-Mata:2022:DEP

[2434] Fernando Estrada-Mata, Juan M. Pacheco-Vega, O. Iram Zavala-Leal, Daniel E. Godínez-Siordia, Viridiana Peraza-Gómez, José Ángel Hinojosa-Larios, Erika Torres-Ochoa, and Julián Gamboa-Delgado. Demand and effect of potassium, magnesium, and calcium chlorides on hemolymph parameters, immune-related gene expression, and growth of Litopenaeus vannamei, Boone (1931) under biofloc technology. Aquaculture International, 30(4):1817–1833, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00880-0.

Wang:2022:EDF

[2435] Rui-Fang Wang, Yuan Wang, Jun Zhang, Mei-Qi Weng, Yu-Hui Liu, Qiu-Yan Cheng, Min Song, Yan-Ping Yang, Xiao-Ping An, and Jing-Wei Qi. The effects of dietary fermented wheat bran polysaccharides on mucosal and serum immune parameters, hepatopancreas antioxidant indicators, and immune-related gene expression of common carp (Cyprinus carpio) juveniles. Aquaculture International, 30(4): 1835–1853, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00877-9.

Guroy:2022:CIG

[2436] Derya Güroy, Onur Karadal, Serhan Mantoğlu, Halit Kuşku, and Betül Güroy. Color intensity and growth performance of common clownfish (Amphiprion ocellaris) enhanced by dietary spirulina (Arthrospira platensis). Aquaculture International, 30(4):1855–1868, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00878-8.

Eldessouki:2022:DAM

[2437] Elsayed A. A. Eldessouki, Amany M. Diab, Talal A. M. Abo Selema, Nader M. Sabry, Mahmoud M. Abotaleb, Riad H. Khalil, Nasser El-Sabbagh, Nehal A. Younis, and Mohsen Abdel-Tawwab. Dietary astaxanthin modulated the performance, gastrointestinal histology, and antioxidant and immune responses and enhanced the resistance of Litopenaeus vannamei against Vibrio harveyi infection. Aquaculture International, 30 (4):1869–1887, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00876-w.

daSilveira:2022:EMP

[2438] Lucas Genésio Pereira da Silveira, Victor Torres Rosas, Dariano Krummenauer, Charles Fróes, Adriana da Silva, Luís Henrique Poersch, Geraldo Fóes, and Wilson Wasielesky. Establishing the most productive stocking densities for each stage of a multi-phase shrimp culture in BFT system. Aquaculture International, 30(4):1889–1903, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00879-7.

Alves:2022:RPI

[2439] Angélica Priscila do Carmo Alves, Ana Paula Peconick, Brunno da Silva Cerozi, and José Eurico Possebon Cyrino. Role of probiotics on the immunity of Nile tilapia Oreochromis niloticus: a review. Aquaculture International, 30(4):1905–1929, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00881-z.

Wang:2022:MCE

[2440] Zhen Wang, Jing Wang, Weijie Mu, and Liqun Liang. Molecular characterization, expression following cold stress, and functional characterization of YB-1 gene in the spotted sea bass (Lateolabrax maculatus). Aquaculture International, 30(4):1931–1950, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00882-y. See correction

[2441].

Wang:2022:CMC

[2441] Zhen Wang, Jing Wang, Weijie Mu, and Liqun Liang. Correction to: Molecular characterization, expression following cold stress, and functional characterization of YB-1 gene in the spotted sea bass (Lateolabrax maculatus). Aquaculture International, 30(4):1951, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00914-7. See [2440].

Krishnapriya:2022:MCE

[2442] R. S. Krishnapriya, Avinash Premraj, and T. P. Sajeevan. Molecular characterization and expression analysis of two RING-between-RING (RBR) ubiquitin ligase orthologues from the Asian seabass (Lates calcarifer). Aquaculture International, 30(4):1953–1970, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00883-x.

Mondal:2022:RRA

[2443] Haimanti Mondal and John Thomas. A review on the recent advances and application of vaccines against fish pathogens in aquaculture. Aquaculture International, 30(4):1971–2000, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00884-w.

Veeragurunathan:2022:IOS

[2444] V. Veeragurunathan, Kamalesh Prasad, Rosy Alphons Sequeira, Ramavatar Meena, Monica Gajanan Kavale, and P. Gwen Grace. Identifying other suitable and potential indigenous carrageenophytes for commercial cultivation in India. Aquaculture International, 30(4):2001– 2015, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00886-8.

Shil:2022:DAB

[2445] Bikash Shil, Biswajit Lahiri, Prasenjit Pal, Amitava Ghosh, Pradyut Biswas, and Yumlembam Jackie Singh. Determinants of adoption behaviour of the fish farmers of Pabda fish culture (Ompok bimaculatus Bloch, 1794) in Tripura, Northeast India. Aquaculture International, 30 (4):2017–2041, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00885-9.

Gebauer:2022:NRD

[2446] Radek Gebauer, Liliana Lehman, Hendrik Monsees, Bernhard Rennert, Jan Mráz, and Werner Kloas. Nitrogen recovery in a decoupled aquaponic system with lamellar settler and trickling biofilter: implications for system management. Aquaculture International, 30(4): 2043–2058, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00888-6.

Jin:2022:GRM

[2447] Shubo Jin, Wenyi Zhang, Yiwei Xiong, Sufei Jiang, Hui Qiao, Yongsheng Gong, Yan Wu, and Hongtuo Fu. Genetic regulation of male sexual development in the oriental river prawn Macrobrachium nipponense during reproductive vs. non-reproductive season. Aquaculture International, 30 (4):2059–2079, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00887-7.

Lelin:2022:IPC

[2448] Chinnadurai Lelin, Eswaramoorthy Thirumalaikumar, Ganapathi Uma, Mariavincent Michael Babu, Chellappan Ajan, Sugumar Vimal, and Thavasimuthu Citarasu. Isolation and partial characterization of bacteriophages infecting Vibrio harveyi from shrimp farm effluent water. Aquaculture International, 30(4):2081–2094, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00891-x.

Lin:2022:EDC

[2449] Tingting Lin, Yishan Cai, Xin Liu, Siping Li, Dong Zhang, and Lianjun Xia. Effects of different carbon sources and carbon–nitrogen ratios on the survival, growth, intestinal biochemical parameters, and water quality of seahorse juveniles cultured under zero-water exchange conditions. Aquaculture International, 30(4):2095–2112, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00889-5.

Dey:2022:HCT

[2450] Biswadeep Dey, Thangapalam Jawahar Abraham, Jasmine Singha, Anwesha Roy, Sutanu Karmakar, Prasanna Kumar Patil, and Utsa Roy. Histopathological changes and tissue residue concentrations of monosex Nile tilapia (Oreochromis niloticus, l) fries exposed to oxytetracycline. Aquaculture International, 30(4):2113–2128, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00892-w.

Balakrishnan:2022:PSA

[2451] Soumya Balakrishnan, Ahna Ameer, Sowmya Pazhur Mohandas, Ambadi Kannan Maliyekkal Sajeevan, Anoop Bhaskaran Sathyabhama, and Bright Singh. Pyocyanin as a safe aquaculture drug for the control of vibriosis in shrimp recirculating aquaculture system (RAS). Aquaculture International, 30(4):2129–2144, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00890-y.

Javanmardi:2022:AVC

[2452] Sina Javanmardi, Kamran Rezaei Tavabe, Gholamreza Rafiee, Saeed Moradi, and Divya Sivaji. Assessment of vitamin C different levels in aquaponic system on Nile tilapia (Oreochromis niloticus) and saffron (Crocus sativus): growth performances, hematology, serum biochemistry, and digestive enzyme activity. Aquaculture International, 30(4):

2145–2163, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00894-8.

deOliveiraRamos:2022:EWR

[2453] Cassio de Oliveira Ramos, Francisco Carlos da Silva, Matthew Gray, Carlos Henrique Araujo de Miranda Gomes, and Claudio Manoel Rodrigues De Melo. Effect of water recirculation rate and initial stocking densities on competent larvae and survival of the Pacific oyster Crassostrea gigas in a recirculation aquaculture system. Aquaculture International, 30(4): 2165–2178, August 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00896-6.

Tanjung:2022:PAI

[2454] Livia Rossila Tanjung, Anggoro Prihutomo, Fitria Nawir, Tjandra Chrismadha, and Tri Widiyanto. Productivity assessment of an intensive whiteleg shrimp Penaeus vannamei farm based on powersimsimulated growth rates. Aquaculture International, 30(4):2179–2196, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00895-7.

Defaee:2022:CEP

[2455] Somayeh Defaee, Bahram Falahatkar, Fatemeh Lavajoo, and Iraj Efatpanah. Correction to: The effect of a phytogenic feed additive (Digestrom P.E.P) on growth performance, proximate composition, hematological and immunological indices of juvenile beluga sturgeon Huso huso. Aquaculture International, 30(4):2197, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00893-9. See [2399].

Zimmermann:2022:CMR

[2456] J. Zimmermann, K. Steinberg, C. Hund, S. Meyer, and C. Schulz. Correction to: Maintenance requirement and utilisation efficiency of energy and protein for ongrowing pikeperch (Sander lucioperca (L.)). Aquaculture International, 30(4):2199, August 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00933-4. See [1942].

Ozil:2022:AAS

[2457] Öznur Özil, Öznur Diler, and Mevlüt Nazıroğlu. Antifungal activity of some essential oil nanoemulsions against saprolegniasis in rainbow trout (Oncorhynchus mykiss) eggs. Aquaculture International, 30(5): 2201–2212, October 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00897-5.

Casimiro:2022:HFS

[2458] Armando Cesar Rodrigues Casimiro, Ana Carolina Vizintim Marques, Alexander Claro-Garcia, Diego Azevedo Zoccal Garcia, Fernanda Simões de Almeida, and Mário Luís Orsi. Hatchery fish stocking: case study, current Brazilian state, and suggestions for improvement. Aquaculture International, 30(5):2213–2230, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00898-4.

Kozlowski:2022:ESG

[2459] Michał Kozłowski and Iwona Piotrowska. Effect of size grading on growth, survival, and cannibalism in larval and juvenile pike, Esox lucius (L.), reared in recirculating systems. Aquaculture International, 30(5): 2231–2244, October 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00899-3.

Gonzales:2022:VET

[2460] Anai Flores Gonzales, Vanessa Mamani, Manuel Pereyra, Edwin Aguilar, Patrick Delgado Mathews, Marcos Tavares-Dias, and Christian Fernández-Méndez. In vitro efficacy and tolerance of the essential oils of three species of the lamiaceae family against monogeneans from the gills of Piaractus brachypomus from the Peruvian Amazon. Aquaculture International, 30(5):2245–2261, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00900-z.

Elhag:2022:FAI

[2461] Abdelrahman Ibrahim Elhag, Sharifah Rahmah, Rasina Abdul Rasid, Sharif Shahin, Gusti Afiz Gusti Ruslan Noor, Mohd Sabri Muda, Mohd Zaidi Ibrahim, Sairatul Dahlianis Ishak, Jiun-Yan Loh, Thumronk Amornsakun, Hua Thai Nhan, Young-Mao Chen, and Hon Jung Liew. Fatty acids in the inedible parts of jade perch Scortum barcoo. Aquaculture International, 30(5):2263–2277, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00901-y.

Sahandi:2022:CAF

[2462] Javad Sahandi, Patrick Sorgeloos, and Wenbing Zhang. Culture of Artemia franciscana nauplii with selected microbes suppressed Vibrio loading and enhanced survival, population stability, enzyme activity, and chemical composition. Aquaculture International, 30(5):2279– 2293, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00905-8.

Lu:2022:EDS

[2463] Yuting Lu, Yilin Zhang, Peijun Zhang, Jia Liu, Bo Wang, Xianglu Bu, Qianyan Wei, Shaojun Liu, and Yuehong Li. Effects of dietary supplementation with Bacillus subtilis on immune, antioxidant, and histopathological parameters of Carassius auratus gibelio juveniles exposed to acute saline-alkaline conditions. Aquaculture International, 30 (5):2295–2310, October 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00902-x.

Zhang:2022:SMO

[2464] Lu Zhang, Wensheng Li, Chunhong Liu, and Qingling Duan. Segmentation method for overlapping fish images based on point detection and pairing. Aquaculture International, 30(5):2311–2341, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00904-9.

Villa-Franco:2022:GPG

[2465] Annie U. Villa-Franco, Milagros R. dela Peña, and Marie Frances J. Nievales. Grazing periodicity, grazing rate, feeding preference, and gut examination of early juveniles of abalone Haliotis asinina-fed five benthic diatom species. Aquaculture International, 30(5):2343–2364, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00906-7.

Murugan:2022:LAA

[2466] Raghul Murugan, Ajay Guru, B. Haridevamuthu, Gokul Sudhakaran, Aziz Arshad, and Jesu Arockiaraj. Lantibiotics: an antimicrobial asset in combating aquaculture diseases. Aquaculture International, 30(5): 2365–2387, October 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00908-5.

Choi:2022:OPR

[2467] Seo Yeol Choi, Eun Hye Lee, Sang Su Shin, Young Hui Lim, and Ho Young Soh. Optimal photoperiod for the reproduction of Eurytemora pacifica: potential live feed for fish larvae. Aquaculture International, 30 (5):2389–2401, October 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00907-6.

Flores:2022:ETS

[2468] Roberto Manolio Valladão Flores, Paul V. Preckel, Kwamena Quagrainie, Nicole Olynk Widmar, Laura Silva, Jesaias Ismael da Costa, Sara M. Pinho, Maria Célia Portella, Thais Castelo Branco, and Manoel Xavier Pedroza Filho. Efficiency tests for screening production strategies in a lettuce-juvenile tilapia aquaponics system in Brazil. Aquaculture International, 30(5):2403–2424, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00912-9.

Zhou:2022:MRG

[2469] Yan Zhou, Yan Li, Luo Lei, Xingxing Deng, Yuting Duan, Suxing Fu, Jingsen Zhang, Dengyue Yuan, Chaowei Zhou, and Wenping He. The melanocortin-4 receptor (MC4R) gene in the gibel carp Carassius auratus gibelio: cloning, tissue distribution, and fasting effects. Aquaculture International, 30(5):2425–2438, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00911-w.

T:2022:EDC

[2470] Naveen Kumar B. T., Nishchal Thakur, Chetna Sharma, A. H. Shanthanagouda, Anuj Taygi, and Amandeep Singh. Effect of dietary chitosan nanoparticles on immune response and disease resistance against Aeromonas hydrophila infection in tropical herbivore fish (rohu, Labeo rohita). Aquaculture International, 30(5):2439–2452, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00910-x.

Fazeli:2022:VVA

[2471] Neda Fazeli, Seyed Amir Hossein Jalali, Akram Sadat Naeemi, and Hojjatolah Zamani. In vitro and in vivo antibacterial activity of sea anemoneisolated Vibrio parahaemolyticus against Yersinia ruckeri. Aquaculture International, 30(5):2453–2475, October 2022. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00913-8.

Liu:2022:EDS

[2472] Liangfang Liu, Yuanxiao Wang, Jiping Ren, Huafeng Zou, and Chun Wang. Effect of dietary supplementation with sodium butyrate and tributyrin on the growth performance and intestinal microbiota of Pacific white shrimp (Litopenaeus vannamei). Aquaculture International, 30(5): 2477–2489, October 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00909-4.

Mbokane:2022:EFR

[2473] Esau Matthews Mbokane, Lucia Matlale Mbokane, and Chris Henri Fouche. The effect of fishmeal replacement with acid-fermented chicken silage on growth, digestive enzyme activity and histology of the intestine and liver of juvenile Mozambique tilapia (Oreochromis mossambicus). Aquaculture International, 30(5):2491–2512, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00916-5.

Chattaraj:2022:RRP

[2474] Sourav Chattaraj, Arindam Ganguly, Asish Mandal, and Pradeep K. Das Mohapatra. A review of the role of probiotics for the control of viral diseases in aquaculture. Aquaculture International, 30(5):2513–2539, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00915-6.

Adamek:2022:GCP

[2475] Zdeněk Adámek and Lenka Kajgrová. Great cormorant (Phalacrocorax carbo sinensis) occurrence in carp aquacultural ponds: a case study from the South Bohemia (Czech Republic) pond region. Aquaculture International, 30(5):2541–2556, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00917-4.

Kabir:2022:ODR

[2476] Khondokar H. Kabir, Mohammed Nasir Uddin, Saifur Rahman, Dietrich Darr, and Md. A. N. Zaman Siddiqi Drubo. Opportunities and determinants for rural youth engagement in catfish farming: empirical evidence from north-central Bangladesh. Aquaculture International, 30(5): 2557–2578, October 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00918-3.

Wang:2022:PDB

[2477] Ruixuan Wang, Xiaojing Liu, Jiangyong Wang, Huizhu, Xiaozhi Lin, Jingfeng Sun, Hongli Mou, Ting Zhang, and Xilan Ma. Proteomic differences between Vibrio tubiashii strains with high- or low-virulence levels isolated from diseased ivory snail Babylonia areolata. Aquaculture International, 30(5):2579–2591, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00919-2.

Rahmati-Holasoo:2022:ICL

[2478] Hooman Rahmati-Holasoo, Sohrab Ahmadivand, Amin Marandi, Sara Shokrpoor, Dušan Palić, and Adib Jahangard. Identification and characterization of lymphocystis disease virus (LCDV) from Indian glassy fish (Parambassis ranga Hamilton, 1822) in Iran. Aquaculture International, 30(5):2593–2602, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00922-7.

Romano:2022:ASI

[2479] Nicholas Romano, Shaun Francis, Shahidul Islam, Austin Powell, and Hayden Fischer. Aquaponics substantially improved sweetpotato (Ipomoea batatas) slip production compared to soil but decreased phenol and antioxidant capacity. Aquaculture International, 30(5):2603– 2610, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00923-6.

Wahjuningrum:2022:LDN

[2480] Dinamella Wahjuningrum, Sri Hariati, Munti Yuhana, Irzal Effendi, Thavasimuthu Citarasu, Diah Ayu Satyari Utami, and Kustiariyah Tarman. Low dose of Nodulisporium sp. KT29 metabolite promotes production performance and innate immunity of Pacific white leg shrimp (Litopenaeus vannamei) against co-infection of white spot syndrome virus and Vibrio harveyi. Aquaculture International, 30(5):2611–2628, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00924-5.

Paula:2022:CSS

[2481] Daniella Aparecida de Jesus Paula, Naiara Melo, William Franco Carneiro, Pedro Leite de Castro Uzeda, Bruna Resende Chaves, Isaac Filipe Moreira Konig, and Luis David Solis Murgas. Cryopreservation of semen in streaked prochilod (Prochilodus lineatus): use of ultra-freezer as alternative technique to liquid nitrogen. Aquaculture International, 30(5): 2629–2639, October 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00926-3.

Monir:2022:FBB

[2482] Md Shirajum Monir, Sabri Mohd Yusoff, Zarirah Mohamed Zulperi, Hasliza Abu Hassim, Mohd Zamri-Saad, Mohammad Noor Azmai Amal, Annas Salleh, Aslah Mohamad, Mohamad Azzam-Sayuti, and Yasin Ina-Salwany. Feed-based bivalent vaccine upregulates expressions of immune-related genes in systemic and mucosal tissues of red hybrid tilapia (Oreochromis niloticus × O. mossambicus) against Streptococcus iniae and Aeromonas hydrophila. Aquaculture International, 30(5): 2641–2659, October 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00925-4.

Elgendy:2022:VOF

[2483] Mamdouh Y. Elgendy, Mohamed Abdelsalam, Amany M. Kenawy, and Shimaa E. Ali. Vibriosis outbreaks in farmed Nile tilapia (Oreochromis niloticus) caused by Vibrio mimicus and V. cholerae. Aquaculture International, 30(5):2661–2677, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00921-8.

Ma:2022:VQC

[2484] Chao Ma, Zhuo Tian, Lili Yang, and Jijuan Cao. Validation of qPCR from a crude extract for the rapid detection of white spot syndrome virus. Aquaculture International, 30(5):2679–2690, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00920-9.

Kaewtapee:2022:ODD

[2485] Chantana Kaewtapee, Karun Thongprajukaew, Patcharee Nualsrithong, Nongphanga Katchoo, Waraporn Hahor, Suktianchai Saekhow, and Nattawipa Thongsawai. Ontogenic development of digestive enzymes in veliger larvae of dog conch (Laevistrombus canarium). Aquaculture International, 30(5):2691–2703, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00927-2.

Nethaji:2022:BMI

[2486] M. Nethaji, B. Ahilan, A. Kathirvelpandiyan, N. Felix, A. Uma, T. L. S. Samuel Mosses, and R. Somu Sunder Lingam. Biofloc meal incorporated diet improves the growth and physiological responses of Penaeus vannamei. Aquaculture International, 30(5):2705–2724, October 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00929-0.

Kumaran:2022:DMA

[2487] Kumaran M, Anand P. R., Ashok Kumar J., Muralidhar M., Ambasankar K., Panigrahi A., and Otta S. K. Development of a mobile application for Pacific white shrimp (Penaeus vannamei) farming and evaluation of its efficiency in technology communication and feedback. Aquaculture International, 30(6):2725–2739, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00930-7.

Borkens:2022:MO

[2488] Yannick Borkens and Paul Koppe. Mytilicola orientalis. Aquaculture International, 30(6):2741–2749, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00928-1.

Yadav:2022:ECV

[2489] Anamika Yadav, Avinash Kumar, and Sudipto Sarkar. Economic comparison of venturi aeration system. Aquaculture International, 30(6): 2751–2774, December 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00931-6.

Kartika:2022:MEM

[2490] Nurmayuni Kartika, Sukenda Sukenda, Sri Nuryati, Angela Mariana Lusiastuti, and Dendi Hidayatullah. Monovalent Mycobacterium fortuitum oral vaccination on giant gourami (Osphronemus goramy) to prevent mycobacteriosis. Aquaculture International, 30(6):2775–2791, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00943-2.

Ou:2022:RNK

[2491] Haixin Ou and Jianyong Liu. Role of Nrf2-Keap1 signaling in the antioxidant defense response induced by low salinity in the kuruma shrimp (Marsupenaeus japonicus). Aquaculture International, 30(6): 2793–2811, December 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00941-4.

Khaudtong:2022:EGM

[2492] Chotiga Khaudtong, Nutt Nuntapong, Karun Thongprajukaew, Waraporn Hahor, and Suktianchai Saekhow. Effects on growth metrics, fillet composition, and health biomarkers of juvenile striped catfish (Pangasianodon hypophthalmus) when commercial feed is replaced with graded levels of restaurant food waste. Aquaculture International, 30(6): 2813–2832, December 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00934-3.

Jaiswar:2022:EPE

[2493] Santlal Jaiswar, Urvashi Mungalapara, Mudassar Anisoddin Kazi, and Nikunj Balar. Evidence from preliminary experiments revealed drifted Ulva biomass has seedling and aquaculture potential. Aquaculture International, 30(6):2833–2846, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00940-5.

El-Saadony:2022:RSA

[2494] Mohamed T. El-Saadony, Abdelrazeq M. Shehata, Mahmoud Alagawany, Abdel-Moneim E. Abdel-Moneim, Dina A. Selim, Mohamed Abdo, Asmaa F. Khafaga, Khaled A. El-Tarabily, Nahed A. El-Shall, and Mohamed E. Abd El-Hack. A review of shrimp aquaculture and factors affecting the gut microbiome. Aquaculture International, 30(6):2847– 2869, December 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00936-1.

Wang:2022:EDG

[2495] Chang’an Wang, Shaoxia Lu, Ze Fan, Shicheng Han, Yang Liu, Haibo Jiang, and Hongbai Liu. Effects of dietary glutathione on the growth performance, skin mucus antioxidant capacity, and immune responses of juvenile taimen Hucho taimen. Aquaculture International, 30(6): 2871–2884, December 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00939-y.

Saleh:2022:DPH

[2496] Reda Saleh, Monica Betancor, Asaad Hassan Mohamed, Tamer El-Sayed Ali, and Marisol Izquierdo. Different phosphatidylcholine and n-3 HUFA contents in microdiets for gilthead seabream (Sparus aurata) larvae: effects on histological changes in intestine and liver. Aquaculture International, 30(6):2885–2900, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00935-2.

Pinto:2022:BPO

[2497] Daniel de Sá Britto Pinto, Lucas Pellegrin, Lilian Fiori Nitz, José Maria Monserrat, and Luciano Garcia. Blood parameters and oxidative stress responses in pacu (Piaractus mesopotamicus) reared at different temperatures. Aquaculture International, 30(6):2901–2918, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00937-0.

Musa:2022:SBA

[2498] Babatunde Olaseni Musa, Alvaro Hernández-Flores, Oludare Akanni Adeogun, José A. Duarte, and Raúl Villanueva-Poot. Stochastic bioeconomic analysis of intensive African catfish cultivation with three sources of uncertainty. Aquaculture International, 30(6):2919–2935, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00938-z.

Falola:2022:EAS

[2499] Abraham Falola, Ridwan Mukaila, and Joshua Oluwasegun Emmanuel. Economic analysis of small-scale fish farms and fund security in North-Central Nigeria. Aquaculture International, 30(6):2937–2952, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00944-1.

Yu:2022:PIM

[2500] Xiaowen Yu, Ningbo Sun, Yongxu Cheng, and Xiaozhen Yang. Povidoneiodine modulates the antioxidant capacity, immunity, and resistance to Aeromonas hydrophila of the Chinese mitten crab, Eriocheir sinensis. Aquaculture International, 30(6):2953–2967, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00942-3.

Demmer:2022:LDE

[2501] Jonathan Demmer, Simon P. Neill, Olga Andres, Shelagh K. Malham, Trevor Jones, and Peter Robins. Larval dispersal from an energetic tidal channel and implications for blue mussel (Mytilus edulis) shellfisheries. Aquaculture International, 30(6):2969–2995, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00948-x.

Alishahi:2022:IEP

[2502] Mojtaba Alishahi, Tahereh Shirali, Mohammad Reza Tabandeh, and Masoud Ghorbanpour. Influence of p-coumaric acid, as a medicinal plant phenolic compound, on expression of virulence genes and pathogenicity of Aeromonas hydrophila in common carp. Aquaculture International, 30(6):2997–3016, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-022-00947-y.

Pilnick:2022:DLP

[2503] Aaron R. Pilnick, Keri L. O’Neil, Matthew A. DiMaggio, and Joshua T. Patterson. Development of larviculture protocols for the long-spined sea urchin (Diadema antillarum) and enhanced performance with diets containing the cryptophyte Rhodomonas lens. Aquaculture International, 30(6):3017–3034, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-022-00945-0.

Thaiparambil:2022:CAE

[2504] Naveen Arakkal Thaiparambil and Vidya Radhakrishnan. Challenges in achieving an economically sustainable aquaponic system: a review. Aquaculture International, 30(6):3035–3066, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00946-z.

Laramore:2022:SVN

[2505] Susan E. Laramore, Paul S. Wills, and M. Dennis Hanisak. Seasonal variation in the nutritional profile of Ulva lactuca produced in a land-based IMTA system. Aquaculture International, 30(6):3067– 3079, December 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00950-3.

Prabhath:2022:DPB

[2506] Gardiyawasam Pussewalage Wineetha A Prabhath, Satya Prakash Shukla, Prem Prakash Srivastava, Kundan Kumar, Paramita Banerjee Sawant, Ajit Kumar Verma, Mithlesh Kumar Chouksey, and Kithulampitiya Koralege Tharaka Nuwansi. Downstream processing of biomass produced in aquaculture wastewater for valuable pigments from the cyanobacterium Spirulina (Arthrospira) platensis: a green and sustainable approach. Aquaculture International, 30(6):3081–3106, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00949-w.

Ruiz-Velazco:2022:EIC

[2507] Javier M. J. Ruiz-Velazco, Margarita Estrada-Perez, Nallely Estrada-Perez, and Alfredo Hernández-Llamas. Estimating the impact of COVID-19 pandemic on alternative semi-intensive shrimp (Penaeus vannamei) production schedules in Mexico: a stochastic bioeconomic approach. Aquaculture International, 30(6):3107–3121, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00951-2.

Oliveira:2022:GPP

[2508] Caio Rubens do Rêgo Oliveira, Valdemir Queiroz de Oliveira, Otávio Augusto Lacerda Ferreira Pimentel, Elizabeth Pereira dos Santos, Paulo Roberto Campagnoli de Oliveira Filho, Alfredo Olivera Gálvez, and Luis Otavio Brito. Growth performance and proximate composition of Penaeus vannamei reared in low-salinity water with different ionic compositions in a synbiotic system. Aquaculture International, 30(6): 3123–3141, December 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00952-1.

DOrbcastel:2022:MEA

[2509] Emmanuelle Roque D’Orbcastel, Mathieu Lutier, Emilie Le Floc’h, François Ruelle, Sébastien Triplet, Patrik Le Gall, Clarisse Hubert, Martine Fortune, Thierry Laugier, Thibault Geoffroy, Anaïs Crottier, Angélique Gobet, and Eric Fouilland. Marine ecological aquaculture: a successful Mediterranean integrated multi-trophic aquaculture case study of a fish, oyster and algae assemblage. Aquaculture International, 30(6): 3143–3157, December 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00953-0.

dosSantos:2022:EDS

[2510] Robson Batista dos Santos, Petrônio Alves Coelho-Filho, Chaiane Santos Assunção, Tais Nunes dos Santos, Josefa Honorio da Silva, Gênison Carneiro Silva, and Luis Otavio Brito. The effect of different synbiotic preparation strategies on water fertilization and zootechnical performance of Macrobrachium rosenbergii reared in the nursery stage. Aquaculture International, 30(6):3159–3178, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00955-y.

Mehrgan:2022:SES

[2511] Mehdi Shamsaie Mehrgan, Seyed Pezhman Hosseini Shekarabi, Azin Azari, Sevdan Yilmaz, Christian Lückstädt, and Houman Rajabi Islami. Synergistic effects of sodium butyrate and sodium propionate on the growth performance, blood biochemistry, immunity, and immune-related gene expression of goldfish (Carassius auratus). Aquaculture International, 30(6):3179–3193, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00954-z.

Danso:2022:MPC

[2512] George Danso, Albert Boaitey, and Miriam Otoo. Market potential and challenges for wastewater aquaculture in Peru. Aquaculture International, 30(6):3195–3212, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00956-x.

Barlaya:2022:GDE

[2513] Gangadhar Barlaya, Umalatha Harish, and Ganesh Hegde. Growth, digestive enzyme activities, and carcass composition of fringe-lipped carp (Labeo fimbriatus) and catla (Catla catla) held in polyculture in tanks with sugarcane bagasse as periphyton substrate. Aquaculture International, 30(6):3213–3228, December 2022. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00957-w.

Nageswari:2022:EDS

[2514] Ponpandy Nageswari, Ajit Kumar Verma, Subodh Gupta, Annamalai Jeyakumari, and Chandrakant Mallikarjun Hittinahalli. Effects of different stocking densities on haematological, non-specific immune, and antioxidant defence parameters of striped catfish (Pangasianodon hypophthalmus) fingerlings reared in finger millet-based biofloc system. Aquaculture International, 30(6):3229–3245, December 2022. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00958-9.

Hosseini:2022:MEL

[2515] Seyed Samad Hosseini, Mojtaba Alishahi, Kourosh Amini, Masoud Ghorbanpour, and Takavar Mohammadian. Microencapsulation of Lactobacillus bulgaricus with alginate-chitosan improves probiotic potency in great sturgeon (Huso huso). Aquaculture International, 30(6):3247– 3268, December 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00959-8.

Malawa:2022:ADC

[2516] Saowalak Malawa, Nutt Nuntapong, Naraid Suanyuk, and Karun Thongprajukaew. Addition of different concentrations of Indian almond (Terminalia catappa) leaf extract to aquarium water resulted in improved water quality and increased bubble nest formation by male Siamese fighting fish (Betta splendens) without having any consistent negative effects on growth metrics and blood chemistry. Aquaculture International, 30(6): 3269–3288, December 2022. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00960-1.

Alam:2023:AQH

[2517] S. M. Nazmul Alam. Advancing quality and health management practices in extensive shrimp (Penaeus monodon) farming in Bangladesh. Aquaculture International, 31(1):1–13, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00961-0.

Moghadam:2023:DEA

[2518] Houriyeh Moghadam, Iman Sourinejad, and Seyed Ali Johari. Digestive enzyme activities, intestinal histology, and gut microbiota of Pacific white shrimp Litopenaeus vannamei fed with turmeric, curcumin, and nanomicelle curcumin. Aquaculture International, 31(1): 15–30, February 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00962-z.

Chen:2023:LDM

[2519] Yingyi Chen, Huihui Liu, Ling Yang, Huihui Yu, Daoliang Li, Siyuan Mei, and Yeqi Liu. A lightweight detection method for the spatial distribution of underwater fish school quantification in intensive aquaculture. Aquaculture International, 31(1):31–52, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00963-y.

Utomo:2023:IRG

[2520] Agil Setya Utomo, Munti Yuhana, Widanarni Widanarni, and Usamah Afiff. Immune response, gene expression, and intestinal microbial composition of Pacific white shrimp fed with multispecies synbiotic for the prevention of coinfection disease. Aquaculture International, 31(1): 53–64, February 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00966-9.

Langi:2023:RFU

[2521] Sandra Langi, Edson Panana, Ceder Alloo, Gilbert Van Stappen, and Wouter Meeus. Replacement of fishmeal using poultry-based protein sources in feeds for pikeperch (Sander lucioperca, Linnaeus, 1758) during grow out phase. Aquaculture International, 31(1):65–80, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00964-x.

Hajar-Azira:2023:PIE

[2522] Zulkifli Hajar-Azira, Mohd Amran Aaqillah-Amr, Nadiah W. Rasdi, Hongyu Ma, and Mhd Ikhwanuddin. Preliminary investigation on the effect of fiddlehead fern, Diplazium esculentum, extract to the growth performance of giant freshwater prawn, Macrobrachium rosenbergii, postlarvae. Aquaculture International, 31(1):81–101, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00965-w.

Chelyadina:2023:MCS

[2523] Natalya S. Chelyadina, Mark A. Popov, and Sergey V. Kapranov. Morphometric characteristics, sex structure, and gonadal ripening of Mytilus galloprovincialis Lam. cultivated in Lake Donuzlav (northwestern Crimea, Black Sea). Aquaculture International, 31(1):103–116, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00967-8.

Millot:2023:ELD

[2524] Rémi Millot, Mikaël Demolliens, Salomé Ducos, Sally Pugliese, Lucie Vanalderweireldt, Alban Delmas, Alizée Boussard, Antoine Aiello, and Eric D. H. Durieux. Embryonic and larval development of Corsican brown meagre, Sciaena umbra (Linnaeus 1758), rearing in captivity from the Mediterranean Sea. Aquaculture International, 31(1): 117–140, February 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00968-7.

Peiro-Alcantar:2023:MBH

[2525] Cecilia Isabel Peiro-Alcantar, Anselmo Miranda-Baeza, Estefanía Garibay-Valdez, Luis Rafael Martínez-Córdova, Francisco Vargas-Albores, Francesco Cicala, Ricardo Gómez-Reyes, and Marcel Martínez-Porchas. Mature biofloc harbor similar bacterial communities regardless of the vegetal floating substrates (oat, amaranth, or wheat) used as promoters. Aquaculture International, 31(1):141–155, February 2023. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-022-00970-
z.

Li:2023:ESC

[2526] Ruoming Li, Teng Chi, Qing Xu, Juntong Liu, Xiaofeng Shan, Rui Zhou, Jiayun Yao, Wuwen Sun, and Guiqin Wang. Effects of single or conjoint administration of lactic acid bacteria as potential probiotics on the growth, immune responses, and disease resistance of Carassius auratus. Aquaculture International, 31(1):157–177, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00969-6.

Osmond:2023:EAC

[2527] Angelisa T. Y. Osmond, Sylvain Charlebois, and Stefanie M. Colombo. Exploratory analysis on Canadian consumer perceptions, habits, and opinions on salmon consumption and production in Canada. Aquaculture International, 31(1):179–193, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00972-x.

Sato-Okoshi:2023:PSA

[2528] Waka Sato-Okoshi, Kenji Okoshi, Hirokazu Abe, and Jean-Claude Dauvin. Polydorid species (Annelida: Spionidae) associated with commercially important oyster shells and their shell infestation along the coast of Normandy, in the English Channel, France. Aquaculture International, 31 (1):195–230, February 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00971-y.

Slette:2023:EUS

[2529] Hans Tobias Slette, Bjørn Egil Asbjørnslett, Kjetil Fagerholt, Ingeborg Margrete Lianes, and Maren Theisen Noreng. Effective utilization of service vessels in fish farming: fleet design considering the characteristics of the locations. Aquaculture International, 31(1):231–247, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00974-9.

Ibrahim:2023:SEC

[2530] Mustafa M. Ibrahim and Marwa M. Attia. Studies on endoparasitic copepods: Sarcotaces arcticus (Copepoda: Phylichthyid) infested dotted grouper (Epinephelus epilistictus) from the Arabian Gulf water, Saudi Arabia. Aquaculture International, 31(1):249–260, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00978-5.

Ramiro:2023:EPV

[2531] Bianca de Oliveira Ramiro, Alda Lúcia de Lima Amâncio, Júlio César da Silva Cacho, Telma de Sousa Lima, Márcia Dantas dos Santos, Otávio Augusto Lacerda Ferreira Pimentel, Luciana Diniz Rola, and Ricardo Romão Guerra. Penaeus vannamei post-larvae growth and economic aspects after hatchery with artemia replacement by commercial feeds. Aquaculture International, 31(1):261–272, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00973-w.

Das:2023:GSI

[2532] Rashmi Ranjan Das, Akshaya Panigrahi, Soumyabrata Sarkar, A. Saravanan, I. F. Biju, K. Ambikanandham, M. Jayanthi, and S. Kannappan. Growth, survival, and immune potential of post larvae of Indian white shrimp, Penaeus indicus (H. Milne Edwards, 1937) in different salinities with biofloc system (BFT) during nursery phase. Aquaculture International, 31(1):273–293, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00979-4.

Leite:2023:DCF

[2533] Jordana Sampaio Leite, Alexandre Firmino Diógenes, and Alberto Jorge Pinto Nunes. Dietary contribution of fermented grain pellets to the growth of juvenile Litopenaeus vannamei raised in an intensive biofloc-based rearing system. Aquaculture International, 31(1):

295–316, February 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00975-8.

Ren:2023:DNA

[2534] Yuting Ren, Yi Mu, Bo Zhao, Yang Gao, Xiaoxin Dai, and Zhangjie Chu. dmrt3, nom1, abce1, and pkmyt1 play key roles in gonadal sex determination in Acrossocheilus fasciatus. Aquaculture International, 31 (1):317–332, February 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00976-7.

Hamilton:2023:GCB

[2535] Matthew G. Hamilton, Mohammed Yeasin, Vishnumurthy Mohan Chadag, Jérôme Delamare-Deboutteville, Partho Pratim Debnath, and John A. H. Benzie. Genetic correlations between harvest weight and secondary traits in a silver carp (Hypophthalmichthys molitrix) genetic improvement program. Aquaculture International, 31(1):333–348, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00977-6.

Yang:2023:ETG

[2536] Ying Yang, Jiangtao Tian, Xinglin Du, Yingying Huang, Yiming Li, Yinying Huang, Qichen Jiang, and Yunlong Zhao. Effects of temperature on the growth parameters, hepatopancreas structures, antioxidant ability, and non-specific immunity of the crayfish, Cherax destructor. Aquaculture International, 31(1):349–365, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00980-x.

Ramli:2023:EDC

[2537] Siti Sarah Ramli, Iswadi Jauhari, Hidayah Manan, Mhd Ikhwanuddin, and Nor Azman Kasan. Effect of different C/N ratio, carbon sources, and aeration flow rates on ammonia fluctuations during startup period of biofloc-based system. Aquaculture International, 31(1): 367–380, February 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00981-w.

Gao:2023:HVH

[2538] Yingli Gao, Gaoshang Han, Lu Qiang, Liyuan Zhang, Ruiming Tan, and Yongxiang Yu. Hematological varieties, histological changes, and immune responses in the early stage of infection with Vibrio parahaemolyticus in black rockfish Sebastes schlegelii. Aquaculture International, 31(1): 381–399, February 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00982-9.

Sherif:2023:PRI

[2539] Ahmed H. Sherif, Elsayed A. Eldessouki, Nader M. Sabry, and Nadia G. Ali. The protective role of iodine and MS-222 against stress response and bacterial infections during Nile tilapia (Oreochromis niloticus) transportation. Aquaculture International, 31(1):401–416, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/

s10499-022-00984-7. See correction [2577].

Abdelsalam:2023:RMD

[2540] Mohamed Abdelsalam, Mamdouh Y. Elgendy, Medhat R. Elfadadny, Sameh Samir Ali, Ahmed H. Sherif, and Said Kamal Abolghait. A review of molecular diagnoses of bacterial fish diseases. Aquaculture International, 31(1):417–434, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-00983-8.

Mukoma:2023:SFF

[2541] Marvin Gaye Mukoma, Steve Omondi Odour, and Elick Onyango Otachi. Sustainable fish feeds: optimization of levels of inorganic fertilizers for mass production of Oocystis sp. for climate smart aquaculture. Aquaculture International, 31(1):435–445, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00986-5.

Tan:2023:CSF

[2542] Hui Teng Tan, Fatimah Md. Yusoff, Yam Sim Khaw, Muhammad Farhan Nazarudin, Nur Amirah Izyan Noor Mazli, Siti Aqlima Ahmad, Noor Azmi Shaharuddin, and Tatsuki Toda. Characterisation and selection of freshwater cyanobacteria for phycobiliprotein contents. Aquaculture International, 31(1):447–477, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-00985-6.

Ibanez:2023:ESS

[2543] Glaiza Ibañez, Rona Cabanayan-Soy, Jerwin Baure, and Marie Antonette Juinio-Meñez. Sargassum sp. juice as an early juvenile supplemental feed for Stichopus cf. horrens. Aquaculture International, 31(1): 479–492, February 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00987-4.

Janhunen:2023:EVS

[2544] Matti Janhunen, Juha-Pekka Turkka, and Jukka Kekäläinen. Extended in vitro storage of eggs and milt increases maternal but not paternal variation in embryo viability of landlocked Atlantic salmon (Salmo salar m. sebago). Aquaculture International, 31(1):493–507, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00989-2.

GusmaoAffonso:2023:SCA

[2545] Elizabeth Gusmão Affonso, Marieta Nascimento de Queiroz, Eduardo Akifumi Ono, Sanny Maria de Andrade Porto, and Gustavo da Silva Claudiano. Sodium chloride against Dawestrema cycloancistrium in juvenile Arapaima gigas. Aquaculture International, 31(1):509–521, February 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00988-3.

Wang:2023:WGS

[2546] Libao Wang, Hui Li, Wenjun Shi, Yi Qiao, Pan Wang, Zhijun Yu, Ran Zhao, Runhao Hu, Hui Shen, Jie Cheng, Ge Jiang, Xintong Huang, and Xihe Wan. Whole-genome sequencing and comparative genomic analysis of a pathogenic Enterocytozoon hepatopenaei strain isolated from Litopenaeus vannamei. Aquaculture International, 31(1): 523–546, February 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00990-9.

Muznebin:2023:EPE

[2547] Farhana Muznebin, Andrea C. Alfaro, and Stephen C. Webb. Perkinsus olseni and other parasites and abnormal tissue structures in New Zealand Greenshell^{T M} mussels (Perna canaliculus) across different seasons. Aquaculture International, 31(2):547–582, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00991-8.

Wang:2023:OPM

[2548] Xin Wang, Zhi xin Guo, Xin yu Lei, Sen Wang, Ji wu Wan, Hong jian Liu, Yu ke Chen, Yun long Zhao, Gui qin Wang, Qiu ju Wang, and Dong ming Zhang. Osmoregulation, physiological metabolism, and oxidative stress responses to water salinity in adult males of Chinese mitten crabs (Eriocheir sinensis). Aquaculture International, 31 (2):583–601, April 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00998-1.

Wu:2023:DLF

[2549] Meili Wu, Meifeng Li, Hua Wen, Lijuan Yu, Ming Jiang, Xing Lu, Juan Tian, and Feng Huang. Dietary lysine facilitates muscle growth and mediates flesh quality of Pacific white shrimp (Litopenaeus vannamei) reared in low-salinity water. Aquaculture International, 31 (2):603–625, April 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-00997-2.

Tavares:2023:RVD

[2550] Camila Prestes dos Santos Tavares, Ubiratan Assis Teixeira Silva, Marcio Pie, and Antonio Ostrensky. A review of viral diseases in cultured brachyuran crustaceans. Aquaculture International, 31(2):627– 655, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00993-6.

Panigrahi:2023:BBF

[2551] A. Panigrahi, Rashmi R. Das, Soumyabrata Sarkar, R. Aravind, P. N. Rekha, and K. P. Sandeep. Biofloc-based farming of Indian white shrimp, Penaeus indicus, in recirculating aquaculture system (RAS) enriched with rotifers as feed supplement. Aquaculture International, 31 (2):657–680, April 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-01000-8.

Akbary:2023:GSN

[2552] Paria Akbary, Ashkan Ajdari, and Bijan Ajang. Growth, survival, nutritional value and phytochemical, and antioxidant state of Litopenaeus vannamei shrimp fed with premix extract of brown Sargassum ilicifolium, Nizimuddinia zanardini, Cystoseira indica, and Padina australis macroalgae. Aquaculture International, 31(2):681–701, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00994-5.

Hamdan:2023:PEE

[2553] Nur Amanina Hamdan, Mohd Ihwan Zakariah, Nor Asma Husna Yusoff, Nor Azri-Shah Norhan, Anur Melad Nagi, Wahidah Wahab, and Marina Hassan. Physiological effects of Melaleuca cajuputi extract on Macrobrachium rosenbergii, and its sensitivity against Probopyrus buitendijki. Aquaculture International, 31(2):703–718, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-00996-3.

Macedo:2023:EEC

[2554] Juliano dos Santos Macedo, Carlos Eduardo Copatti, Emmanoel Vilaça Costa, Felipe Moura Araújo da Silva, Lívia Macedo Dutra, Victória Laysna dos Anjos Santos, Jackson Roberto Guedes da Silva Almeida, Marcos Tavares-Dias, and José Fernando Bibiano Melo. Effects of Citrus limon extract on growth performance and immunity in striped catfish (Pangasius hypophthalmus). Aquaculture International, 31(2):719–738, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-00995-4.

Ganesan:2023:ROM

[2555] Sathiyaraj Ganesan, Babu Baskaran, Mithun Raj, Saravanan Marimuthu, Velmurugan Krishnasamy, Ruban Lamech, Anup Mandal, Kandan Shanmuganathan, and Prabhu Narayanasamy Marimuthu. Reovirus occurrence in mud crab farming systems and wild-caught brooders located in eastern coastal area of India. Aquaculture International, 31(2):739– 758, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-00992-7.

Wang:2023:EOP

[2556] Xiaoqing Wang, Maosheng Wei, Jing Wang, Jing Liu, and Quanbin Zhang. Effects of oligo-porphyran on immunological parameters related to immunoregulation and growth in RAW264.7 macrophages and zebrafish model. Aquaculture International, 31(2):759–775, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-022-00999-
0.

Nokhwal:2023:BTE

[2557] Alka Nokhwal, Taruna Anand, Ravikant, and Rajesh Kumar Vaid. Bacteriophage therapy: an emerging paradigm in fish disease management. Aquaculture International, 31(2):777–805, April 2023. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01001-7.

Fabrini:2023:OEH

[2558] Sabrina Aparecida Fabrini, Raquel Stroher, Fabiano Bisinella Scheufele, Marlise Teresinha Mauerwerk, and Eduardo Luis Cupertino Ballester. Optimization of the enzymatic hydrolysis process of shrimp viscera (Macrobrachium rosenbergii) with two commercial enzymes, aiming to produce an alternative protein source for aquaculture feed formulation. Aquaculture International, 31(2):807–825, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01010-6.

Buwono:2023:ECE

[2559] Ibnu Dwi Buwono, Iskandar Iskandar, and Roffi Grandiosa. CgGH and igf-1 expression level and growth response of G_{4} transgenic mutiara strain catfish (Clarias gariepinus) reared at different stocking densities. Aquaculture International, 31(2):827–846, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01002-6.

Afshari:2023:DMS

[2560] Fatemeh Afshari, Ebrahim Sotoudeh, Mansour Torfi Mozanzadeh, and Ahmad Ghasemi. Dietary mixture of short-chain fatty acids, a phytogenic agent, and a permeabilizer improved growth, antioxidant enzymes, and immunocompetence in whiteleg shrimp juveniles (Penaeus vannamei). Aquaculture International, 31(2):847–866, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-01003-5.

Radhakrishnan:2023:OVS

[2561] Akshaya Radhakrishnan, Baskaralingam Vaseeharan, Pasiyappazham Ramasamy, and Sivakamavalli Jeyachandran. Oral vaccination for sustainable disease prevention in aquaculture — an encapsulation approach. Aquaculture International, 31(2):867–891, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01004-4.

Siddiqua:2023:SPO

[2562] Kazi Sabnam Siddiqua and Mukhtar A. Khan. Substituting palm oil for fish oil in feeds for juvenile rohu, Labeo rohita: effects on growth performance, fillet fatty acid composition, and antioxidant capacity. Aquaculture International, 31(2):893–913, April 2023. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01005-3.

Pan:2023:MCC

[2563] Chenglong Pan, Lingling Li, Hao Zhu, Wenjia Mao, Tao Han, Xuqian Zhao, Caijuan Li, and Qufei Ling. Molecular cloning, characterization, and expression patterns of the hatching enzyme genes during embryonic development of pikeperch (Sander lucioperca). Aquaculture International, 31(2):915–930, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-01006-2.

Huang:2023:TAJ

[2564] Jian-Sheng Huang, Zhi-Xiong Guo, Jian-Dong Zhang, Wei-Zheng Wang, Zhong-Liang Wang, Rui-Tao Xie, Eric Amenyogbe, and Gang Chen. Transcriptomic analysis of juvenile cobia in response to hypoxic stress. Aquaculture International, 31(2):931–955, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01007-1.

Yang:2023:AAM

[2565] Hu Yang, Jia Li, Huizhong Xu, Chunfeng Peng, Jun Cui, Shengbiao Hu, Liqiu Xia, and Youming Zhang. Antioomycete activity and mechanism of acidic electrolyzed water: a novel sanitizer to prevent saprolegniasis in grass carp. Aquaculture International, 31(2):957–973, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-01008-0.

Moser:2023:SEN

[2566] Juliana Righetto Moser and Maria Risoleta Freire Marques. Susceptibility of Neohelice granulata (Decapoda, Varunidae) to white spot syndrome virus (WSSV). Aquaculture International, 31(2):975–996, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-01009-z.

Adhikary:2023:IIC

[2567] Anamika Adhikary, Gayatree Hazarika, Rajdeep Das, Kangkan Sarma, S. Vimal, S. Abdul Majeed, A. S. Sahul Hameed, and Dandadhar Sarma. Isolation, identification and characterisation of Aeromonas veronii from diseased Indian major carp, Cirrhinus mrigala. Aquaculture International, 31(2):997–1010, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-022-01012-4.

dosSantos:2023:GBC

[2568] Fabio Aremil Costa dos Santos, Leandro Santos Costa, Helder de Oliveira Guilherme, Pedro Gomes Gamarano, Jose Fernando López-Olmeda, Verônica Guimarães Landa Prado, Débora de Almeida Freitas, Luiz Felipe da Silveira Silva, Ronald Kennedy Luz, and Paula Adriane Perez Ribeiro. Growth and blood chemistry of juvenile neotropical catfish (Lophiosilurus alexandri) self-feeding on diets that differ in protein-to-energy (P:E) ratio. Aquaculture International, 31(2):1011– 1029, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-01013-3.

daSilva:2023:EDI

[2569] Thamyres Vanessa N. da Silva, Camila F. dos Santos, Jessica M. L. dos Santos, Marcos J. Schmitz, Juan R. B. Ramírez, Marcelo F. Torres, Luis André L. Barbas, Luís A. Sampaio, Pablo E. Verde, Marcelo B. Tesser, and José M. Monserrat. Effects of dietary inclusion of lyophilized açai berries (Euterpe oleracea) on growth metrics, metabolic and antioxidant biomarkers, and skin color of juvenile tambaqui (Colossoma macropomum). Aquaculture International, 31(2):1031–1056, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-01014-2. See

correction [2753].

Martinez-Cordova:2023:ATM

[2570] Luis R. Martinez-Cordova, Mauricio G. C. Emerenciano, Anselmo Miranda-Baeza, Sara M. Pinho, Estefanía Garibay-Valdez, and Marcel Martínez-Porchas. Advancing toward a more integrated aquaculture with polyculture > aquaponics > biofloc technology > FLOCponics. Aquaculture International, 31(2):1057–1076, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01016-0.

Zoral:2023:MPT

[2571] Mehmet Arif Zoral. Medicinal plants: are they safe enough for fish health? Aquaculture International, 31(2):1077–1096, April 2023. CO-DEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-022-01015-

1. See correction [2578].

Gao:2023:MHR

[2572] Yingli Gao, Lu Qiang, Liyuan Zhang, Ruiming Tan, Xingqiang Wang, and Yongxiang Yu. Mucosal and humoral responses of javelin goby, Synechogobius hasta, after immersion vaccination using killed Vibrio vulnificus bacterin. Aquaculture International, 31(2):1097–1113, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-01017-z.

Basumatary:2023:GRT

[2573] Bwsrang Basumatary, A. K. Verma, and Manoj Kumar Verma. Global research trends on aquaponics: a systematic review based on computational mapping. Aquaculture International, 31(2):1115–1141, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-01018-y.

Lahiri:2023:RFS

[2574] Taushik Lahiri, Muhammad Arifur Rahman, and Abdullah-Al Mamun. Reassessing the food security implications of export-oriented aquaculture in Bangladesh. Aquaculture International, 31(2):1143–1162, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-01019-x.

Xu:2023:SBC

[2575] Liting Xu, Yilong Ruan, Maowang Jiang, Ruibing Peng, Xiamin Jiang, Weiwei Zhang, and Qingxi Han. Succession of bacterial communities during a disease progress in cuttlefish Sepia pharaonis. Aquaculture International, 31(2):1163–1175, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-01022-2.

Lozano-Olvera:2023:PSS

[2576] Rodolfo Lozano-Olvera and Selene M. Abad-Rosales. Practical scoring system for diagnosis presumptive of acute necrosis of the hepatopancreas (AHPND) in shrimp Penaeus vannamei. Aquaculture International, 31 (2):1177–1189, April 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-01020-4.

Sherif:2023:CPR

[2577] Ahmed H. Sherif, Elsayed A. Eldessouki, Nader M. Sabry, and Nadia G. Ali. Correction to: The protective role of iodine and MS-222 against stress response and bacterial infections during Nile tilapia (Oreochromis niloticus) transportation. Aquaculture International, 31(2): 1191, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-01011-5. See [2539].

Zoral:2023:CMP

[2578] Mehmet Arif Zoral. Correction to: Medicinal plants: are they safe enough for fish health? Aquaculture International, 31(2):1193, April 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-01034-y. See [2571].

Chen:2023:LTA

[2579] Xiumei Chen, Jiasong Zhang, Muyang Li, Jiaxin Tian, Xiaotian Niu, Xiaofeng Shan, Sha Luo, Guiqin Wang, and Aidong Qian. Liver transcriptome analysis and identification of differentially expressed immune gene response to Aeromonas veronii infection in Channa argus. Aquaculture International, 31(3):1195–1211, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01079-7.

Mai:2023:AED

[2580] Qianyi Mai, Yuqi Jin, Yanfeng Chen, Hanxu Dong, Yali Wu, Dongli Sun, Weiqiang Liu, Yingying Yu, Xinxian Wei, Ying Yang, Hui Yu, Mingjian Lai, and Weiwei Zeng. Assessing the effects of dietary live prey versus an artificial compound feed on growth performance, immune response, and intestinal microflora of largemouth bass Micropterus salmoides. Aquaculture International, 31(3):1213–1230, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01021-3.

Dong:2023:RLQ

[2581] Miaoping Dong, Yi Zhang, Qiuyan Yu, Qing Liu, Xiaojian Zhou, and Cuili Jin. Regulation of light quality on lipid production, biodiesel quality, and nutritional quality of Phaeodactylum tricornutum. Aquaculture International, 31(3):1231–1251, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01024-0.

Youssef:2023:VGC

[2582] Hadeer A. Youssef, Hala F. Ayoub, Eman I. Soror, and Aya F. Matter. Virulence genes contributing to Aeromonas veronii pathogenicity in Nile tilapia (Oreochromis niloticus): approaching the development of live and inactivated vaccines. Aquaculture International, 31 (3):1253–1267, June 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-01023-1.

Kim:2023:EAC

[2583] Bong-Tae Kim and Sangchoul Yi. Exploring Asian consumers’ seafood purchasing behavior: a case of farmed Atlantic salmon products in South Korea. Aquaculture International, 31(3):1269–1286, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-01027-x.

Royer:2023:DAK

[2584] Edouard Royer and Roberto Pastres. Data assimilation as a key step towards the implementation of an efficient management of dissolved oxygen in land-based aquaculture. Aquaculture International, 31(3): 1287–1301, June 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-01028-w.

Silva:2023:EDF

[2585] Gênison Carneiro Silva, Agatha Catharina Limeira, Gisely Karla de Almeida Costa, Suzianny Maria Bezerra Cabral da Silva, Paulo Roberto Campagnoli de Oliveira Filho, and Luis Otavio Brito. Effects of different forms of artificially salinized in low-salinity water of Penaeus vannamei in the grow-out phase in a synbiotic system. Aquaculture International, 31(3):1303–1324, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-022-01025-z.

Peter:2023:DFE

[2586] Nila Rekha Peter, Albin Sunny, Nishan Raja Raja, Nijan Josephraj, Hemanth Kumar, and Ambasankar Kondusamy. Design, fabrication, and evaluation of a low-cost portable solar dryer for shrimp feeds. Aquaculture International, 31(3):1325–1348, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01026-y.

Kong:2023:MCT

[2587] Xianghui Kong, Hairong Tang, Yachen Zhu, Jie Zhang, Chunjing Li, Xianliang Zhao, Chao Pei, Yong Zhou, and Lingbing Zeng. Molecular characterizations of TLR1 and TLR2 in Qihe crucian carp (Carassius auratus) and responses to stimulations of Aeromonas hydrophila and TLR ligands. Aquaculture International, 31(3):1349–1374, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-01030-2.

Simtoe:2023:PRF

[2588] Ambakisye Poland Simtoe, Siwema Amran Luvanga, and Blandina Robert Lugendo. Partial replacement of fishmeal with Chaetomorpha algae improves feed utilization, survival, biochemical composition, and fatty acids profile of farmed shrimp Penaeus monodon post larvae. Aquaculture International, 31(3):1375–1388, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01029-9.

Jaiswar:2023:ESF

[2589] Santlal Jaiswar, Pankaj S. Dawange, Chetna M. Zala, and Nikunj Balar. Effect of shrimp farm effluent (SFE) on growth, survival, regeneration, and biochemical composition in indigenous red seaweed Gracilaria corticata var. cylindrica. Aquaculture International, 31(3): 1389–1400, June 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01055-1.

Gilson:2023:EFD

[2590] F. Gilson, M. B. New, L. A. Rodrigues, and W. C. Valenti. Effect of fish downstream supply chain on wealth creation: the case of tambatinga in the Brazilian Midnorth. Aquaculture International, 31(3): 1401–1421, June 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01056-0.

Phan:2023:PBT

[2591] Hong-Nhung T. Phan, Hong-Loan T. Nguyen, Tuan-Hung Dinh, Ngoc T. Le, Ha-Giang Vu, and Tuan-Nghia Phan. Phenoloxidases from black tiger shrimp (Penaeus monodon): gene expression and activity distribution in different tissues. Aquaculture International, 31(3):1423– 1437, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-01033-z.

Preenanka:2023:ETP

[2592] Rajan Preenanka, Muhammed P. Safeena, and Rahul Krishnan. Evaluation of therapeutic potential of Streptococcus phage-1A against Streptococcus agalactiae infection in Nile tilapia (Oreochromis niloticus). Aquaculture International, 31(3):1439–1456, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01036-w.

Elgendy:2023:OEA

[2593] Mamdouh Y. Elgendy, Shimaa E. Ali, Mohamed Abdelsalam, Tamer H. Abd El-Aziz, Faten Abo-Aziza, Hussien A. Osman, Mohammad M. N. Authman, and Wafaa T. Abbas. Onion (Allium cepa) improves Nile tilapia (Oreochromis niloticus) resistance to saprolegniasis (Saprolegnia parasitica) and reduces immunosuppressive effects of cadmium. Aquaculture International, 31(3):1457–1481, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01035-x.

Shin:2023:EMS

[2594] Jaebeom Shin, Julie Feyaerts, Mirasha Hasanthi, Chorong Lee, Danbi Shin, Yeonji Lee, Jongho Lim, Han-Se Kim, and Kyeong-Jun Lee. Evaluation of a mixture of short-chain and medium-chain fatty acid glycerides as a dietary supplement in diets for Pacific white shrimp (Litopenaeus vannamei). Aquaculture International, 31(3):1483–1498, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-01037-9.

Prabu:2023:EWT

[2595] Dhanasekaran Linga Prabu, C. Kalidas, L. Ranjith, Sanal Ebeneezar, M. Kavitha, P. U. Zacharia, P. Vijayagopal, A. Mathan Babu, and B. Ramar Muniswaran. Effect of water temperature on growth, blood biochemistry, digestive, metabolic enzymology, and antioxidant defences of Trachinotus blochii juveniles. Aquaculture International, 31(3):1499– 1522, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-01038-8.

Panigrahi:2023:GPI

[2596] Akshaya Panigrahi, Rashmi Ranjan Das, I. F. Biju, A. Saravanan, Soumyabrata Sarkar, P. S. S. Shyne Anand, L. Christina, J. Antony, K. P. Kumaraguru Vasagam, K. Ambasankar, M. Jayanthi, M. Kumaran, P. Mahalaxmi, S. Rajamanickam, and C. P. Balasubramanian. Growth potential and immunity of the Indian white shrimp, Penaeus indicus (H. Milne-Edwards, 1837), cultured in grow-out ponds at varying densities and salinities. Aquaculture International, 31(3): 1523–1549, June 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-01039-7.

Mramba:2023:SCA

[2597] Rosemary Peter Mramba and Emmanuel Jacob Kahindi. The status and challenges of aquaculture development in Dodoma, a semiarid region in Tanzania. Aquaculture International, 31(3):1551–1568, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-022-01041-z.

Li:2023:EOS

[2598] Jinghao Li, Chen Qian, Chenlu Li, Zhengyou Li, Yewen Xi, Yongxu Cheng, and Jiayao Li. Exploration of the optimal stocking density of red swamp crayfish (Procambarus clarkii) larvae by using the biofloc technology. Aquaculture International, 31(3):1569–1582, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-01042-y.

Bhendarkar:2023:ICI

[2599] Mukesh Parasram Bhendarkar, Bhaskar Bharat Gaikwad, Amol Kamalakar Bhalerao, Ankush Lala Kamble, K. Viswanatha Reddy, Sreekanth Giri Bhavan, R. Sendhil, P. Ramasundaram, and Sonal Rajendra Kalbande. Impacts of COVID-19-induced lockdown and key reforms in the Indian fisheries sector — a stakeholders’ perspective. Aquaculture International, 31(3):1583–1605, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-022-01040-0.

Thakur:2023:IIB

[2600] Kushal Thakur, Ankita Sharma, Dixit Sharma, Bhavna Brar, Kanika Choudhary, Amit Kumar Sharma, Danish Mahajan, Ranjit Kumar, Sunil Kumar, and Rakesh Kumar. An insight into the interaction between Argulus siamensis and Labeo rohita offers future therapeutic strategy to combat argulosis. Aquaculture International, 31(3):1607– 1621, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-022-01043-x.

Alavijeh:2023:EME

[2601] Maryam Karimi Alavijeh, Sepideh Safi, and Abdolkarim Zarei. An efficient method for economic micropropagation of three aquatic plant species (Lobelia cardinalis, Staurogyne repens, and Alternanthera reineckii). Aquaculture International, 31(3):1623–1636, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-022-01044-w.

Jamal:2023:CVS

[2602] Md Roushon Jamal. Can vannamei shrimp (Litopenaeus vannamei) revitalise Bangladesh’s dying shrimp industry? Aquaculture International, 31(3):1637–1641, June 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-022-01045-9.

Reis:2023:PAF

[2603] João Reis, Aya S. Hussain, Alexis Weldon, Samuel Walsh, William Stites, Melanie Rhodes, and D. Allen Davis. Passive acoustic feeders as a tool to assess feed response and growth in shrimp pond production. Aquaculture International, 31(3):1643–1657, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01053-3.

Azhar:2023:OSE

[2604] Nuril Azhar and Ervia Yudiati. Outbreak simulation of Litopenaeus vannamei recovery rate with oral alginate and spirulina diet supplementation against Vibrio parahaemolyticus AHPND. Aquaculture International, 31 (3):1659–1676, June 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01050-6.

Duong:2023:MSP

[2605] Ngoc-Diem Duong, Thuy-Dung Mai-Hoang, Khai-Hoan Nguyen-Phuoc, Kim-Yen Thi Do, Nguyet-Thu Thi Nguyen, Thuoc Linh Tran, and Hieu Tran-Van. Monitoring the secreted profile of PirA^{vp} and PirB^{vp} toxins from Vibrio parahaemolyticus causing acute hepatopancreatic necrosis disease. Aquaculture International, 31(3):1677–1684, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01048-0.

dosSantos:2023:DPR

[2606] Raphael Brito dos Santos, Judá Izel-Silva, Paulo Adelino de Medeiros, Michelle Midori Sena Fugimura, Thiago Mendes de Freitas, Eduardo Akifumi Ono, Gustavo da Silva Claudiano, and Elizabeth Gusmão Affonso. Dietary protein requirement for tambaqui cultivated in biofloc and clear water systems. Aquaculture International, 31(3):1685–1704, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01047-1.

Gao:2023:EDM

[2607] Xin Gao, Haojie Zhai, Libo Wei, Lidong Shi, Lin Yan, Zuxiang Peng, Wei Wang, Tongjun Ren, and Yuzhe Han. Effects of dietary mannan oligosaccharides on growth, non-specific immunity, and intestinal health in juveniles of the Japanese sea cucumber (Apostichopus japonicus). Aquaculture International, 31(3):1705–1727, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01054-2.

Medeiros:2023:DAR

[2608] Luiza Medeiros, Bruna Nornberg, Raíza Azevedo, Arthur Cardoso, Victor Torres Rosas, Marcelo Borges Tesser, Virgínia F. Pedrosa, Luis A. Romano, Wilson Wasielesky, and Luis Fernando Marins. Dietary addition of recombinant Bacillus subtilis expressing a fungal phytase increases phosphorus fixation in muscle of Pacific white shrimp Litopenaeus vannamei. Aquaculture International, 31(3):1729–1742, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01049-z.

Ghosh:2023:IAP

[2609] Alokesh Kumar Ghosh, Sujogya Kumar Panda, and Walter Luyten. Immunomodulatory activity of plants against white spot syndrome virus (WSSV) in shrimp culture: a review. Aquaculture International, 31 (3):1743–1774, June 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01051-5.

Braga:2023:IBW

[2610] Ítalo Felipe Mascena Braga, Marcele Trajano Araújo, Luis Otavio Brito, and Eudes de Souza Correia. Influence of BFT and water exchange systems on growth, ammonia tolerance, and water footprint in Macrobrachium rosenbergii nursery in intensive systems. Aquaculture International, 31(3):1775–1788, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01046-2.

Saleem:2023:UDB

[2611] Rimsha Saleem, Syed Zakir Hussain Shah, Mahroze Fatima, Maryam, and Hafsa Nadeem. Use of dietary β-mannanase supplementation to increase nutrient digestibility and improve growth of juvenile rohu (Labeo rohita) given a feed based on plant ingredients. Aquaculture International, 31(3):1789–1804, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01052-4.

Andriawan:2023:RGF

[2612] Soni Andriawan, Hung Tran Bao, Wahyu Purbiantoro, Ating Yuniarti, Hso Chi Chaung, Tsair-Bor Yen, and Ta-Chih Cheng. Reversegavage feeding as a novel administration to investigate the immunomodulatory effects of Radix scutellaria water extract on Macrobrachium rosenbergii immunity. Aquaculture International, 31(3):1805–1820, June 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01058-y.

Nunes:2023:TCE

[2613] Bruna Rafaela Caetano Nunes, Yuri Vinicius de Andrade Lopes, Ricardo Henrique B. de Souza, Raul Dirceu Pazdiora, Thiago Tetsuo Ushizima, Simone Paiva Medeiros, Lorrayne Tavares Corrêa, Igor Izidoro Cardoso, Matheus Vitor Vieira Andrade, Patricia Oliveira Maciel-Honda, Edsandra Campos Chagas, and Fernanda de Alexandre Sebastião. Treatments for the control of Neoechinorhynchus buttnerae (Acanthocephala) in tambaqui Colossoma macropomum. Aquaculture International, 31 (3):1821–1835, June 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01057-z.

Hittinahalli:2023:DPE

[2614] Chandrakant Mallikarjun Hittinahalli, Bimal Chandra Mal, Appidi Krishna Reddy, Ajit Kumar Verma, and Abinaya Pattusamy. Design and performance evaluation of rotating biological contactors for recirculating freshwater prawn (Macrobrachium rosenbergii) hatchery using artificial seawater. Aquaculture International, 31(4):1837–1854, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01060-4.

Olutumise:2023:IRF

[2615] Adewale Isaac Olutumise. Impact of relaxing flood policy interventions on fish production: lessons from earthen pond-based farmers in Southwest Nigeria. Aquaculture International, 31(4):1855–1878, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01062-2.

Ullah:2023:IPG

[2616] Hafiz Ullah, Bakhtiar Gul, Haroon Khan, Khushnood ur Rehman, Ishfaq Hameed, Umar Zeb, Sohaib Roomi, and Zill-E-Huma. Impact of pH on the growth and nutritional profile of Lemna minor L. as a sustainable alternative for Pakistan’s feed sector. Aquaculture International, 31(4): 1879–1891, August 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01063-1.

Amin:2023:EPS

[2617] Muhamad Amin, Bunga Intan, Mashielda Arbias Ridwan Putri, Akhmad Taufiq Mukti, and Mochammad Amin Alamsjah. Effect of protein sources in formulated diets on growth performance, feed utilization, survival rate, and reproductive performance of Artemia franciscana. Aquaculture International, 31(4):1893–1910, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01059-x.

Chen:2023:ESM

[2618] Defang Chen, Hao Zhu, Daiyu Chen, Lei Yang, Fei Yang, Qiaolin Pei, Yue He, Xin Zhang, Lianshi Qu, Tianqiang Liu, Yi Geng, and Zhiqiong Li. Salvia miltiorrhiza polysaccharides enhance the antioxidant capacity and immune resistance of hybrid sturgeon (Acipenser baerii [female sign] × Acipenser schrenckii [male sign]) against Streptococcus iniae infection. Aquaculture International, 31(4):1911–1923, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01061-3.

Natarajan:2023:HRA

[2619] Annamalai Natarajan, Radhakrishnan Prabakarakrishnan, Kasivelu Govindaraju, Vimal Sugumar, Kuppusamy Sathiskumar, Jayaraman Narenkumar, Aadhikesavan Ramanan, and Balasubramanian Senthil Kumar. Hepato and renoprotective activity of Kappaphycus alvarezii ethanolic extract in cisplatin causes hepatic and kidney harm in albino Wistar rats. Aquaculture International, 31(4):1925–1940, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01064-0.

Wang:2023:EEA

[2620] Yuzhen Wang, Zhao Chen, Zhiqiang Chang, Gao Meng, and Jian Li. Ecological and economic analysis for different shrimp farming models in China. Aquaculture International, 31(4):1941–1958, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01066-y.

Mansour:2023:VVA

[2621] Abdallah Tageldein Mansour, Elsayed A. Eldessouki, Riad H. Khalil, Amany M. Diab, Talal A. M. Abo Selema, Nehal A. Younis, and Nashwa Abdel-Razek. In vitro and in vivo antifungal and immune stimulant activities of oregano and orange peel essential oils on Fusarium solani infection in whiteleg shrimp. Aquaculture International, 31(4): 1959–1977, August 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01065-z.

PereiraBesen:2023:BTB

[2622] Kayane Pereira Besen, Larissa da Cunha, Nandara Soares de Oliveira, Luiz Augusto Cipriani, Mariana Bender, Rafaela Gomes, Everton Skoronski, and Thiago El Hadi Perez Fabregat. Biofloc technology (BFT) system improves survival and intestinal health of Carassius auratus larvae subjected to different food management. Aquaculture International, 31(4):1979–1994, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-023-01068-w.

Lebata-Ramos:2023:SEG

[2623] Ma. Junemie Hazel L. Lebata-Ramos, Ellen Flor D. Solis, and Joseph B. Biñas. Seaweed Gracilariopsis heteroclada, formulated flaked diet, and a combination of both as feed for the abalone Haliotis asinina: effect on growth and survival. Aquaculture International, 31(4):1995– 2009, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01069-9.

Haobijam:2023:IFF

[2624] J. W. Haobijam and Souvik Ghosh. Integrated fish farming and its influence on farm livelihoods in Manipur, India. Aquaculture International, 31 (4):2011–2034, August 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01067-x.

Lopez-Mas:2023:ECB

[2625] Laura López-Mas, Anna Claret, Gaëlle Arvisenet, Roser Romero del Castillo, Zein Kallas, Massimo Zuccaro, and Luis Guerrero. European consumers’ beliefs about the main pillars of the sustainability: a comparison between wild and farmed fish. Aquaculture International, 31(4): 2035–2055, August 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01070-2.

Eyayu:2023:RPS

[2626] Alamrew Eyayu, Abebe Getahun, and James Last Keyombe. A review of the production status, constraints, and opportunities in East African freshwater capture and culture fisheries. Aquaculture International, 31 (4):2057–2078, August 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01071-1. See [2752].

Hernandez-Arzaba:2023:CSP

[2627] Juan Cristóbal Hernández-Arzaba, Pierre Failler, Alberto Asiain-Hoyos, Diego Esteban Platas-Rosado, Andy Forse, Benjamin M. Drakeford, Alba Rocío Muñoz-Madrid, and Rogelio Limón-Rivera. Cluster strategy as a public policy option for aquaculture development in Mexico: the case for tilapia. Aquaculture International, 31(4):2079–2098, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01073-z.

Wang:2023:ITM

[2628] Qishuai Wang, Qian Hu, Siqi Yang, and Yanhe Li. Isolation of tetrameric microsatellite markers and its application on parentage identification in Procambarus clarkii. Aquaculture International, 31(4):2099–2111, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01072-0.

dosSantos:2023:SRA

[2629] Nayara Netto dos Santos, Adriana Xavier Alves, Gean Paulo Andrade Reis, Victor Ramos Pawlowski, Núbia Gomes Braga, Jéssica Almeida da Silva, Marcos Ferreira Brabo, Daniel Abreu Vasconcelos Campelo, and Galileu Crovatto Veras. Stress and recovery of American bullfrog after biometry management: biochemistry and erythrogram responses. Aquaculture International, 31(4):2113–2125, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01074-y.

Reyden:2023:ISD

[2630] Carrie A. R. Reyden, Natalí J. Delorme, Paul M. South, and J. David Aguirre. Impacts of seeding density on the oxidative stress response of the Greenshell^{T M} mussel, Perna canaliculus. Aquaculture International, 31 (4):2127–2143, August 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01078-8.

Jateen:2023:SBB

[2631] Sinha Jateen, Vidya Shree Bharti, Satya Prakash, Sreedharan Krishnan, Tapas Paul, and Saurav Kumar. Sugarcane bagasse biocharamended sediment improves growth, survival, and physiological profiles of white-leg shrimp, Litopenaeus vannamei (Boone, 1931) reared in inland saline water. Aquaculture International, 31(4):2145–2164, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01077-9.

Raja:2023:AAE

[2632] Bharath Raja, Naveen Arakkal Thaiparambil, Kariyanna B, Vidya Radhakrishnan, and Sudhakaran Raja. Antiviral activity of Turbinaria ornata against white spot syndrome virus in freshwater crab (Paratelphusa hydrodromous). Aquaculture International, 31(4):2165–2177, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01075-x.

Korni:2023:USN

[2633] Fatma M. M. Korni, Asmaa N. Mohammed, and Usama K. Moawad. Using some natural essential oils and their nano-emulsions for ammonia management, anti-stress and prevention of streptococcosis in Nile tilapia, Oreochromis niloticus. Aquaculture International, 31(4):2179– 2198, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01076-w.

Wu:2023:EDC

[2634] Lin Wu, Lei Wang, Shuchang Cui, Zuxiang Peng, Zexin Liu, Muzi Li, Yuzhe Han, and Tongjun Ren. Effects of dietary compound probiotics and heat-killed compound probiotics on antioxidative capacity, plasma biochemical parameters, intestinal morphology, and microbiota of Cyprinus carpio haematopterus. Aquaculture International, 31(4): 2199–2219, August 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01080-0.

Bakthavachalam:2023:ICB

[2635] Dhivya Bakthavachalam and Sivakumar Arumugam. Identification and characterization of bioactive peptides from marine crustacean crabs: a possible drug candidate for Alzheimer’s disease. Aquaculture International, 31(4):2221–2234, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01083-x.

Nisa:2023:JEA

[2636] S. Amreen Nisa, K. Govindaraju, R. Vasantharaja, M. Kannan, and K. Raja. Jellyfish Acromitus flagellatus (Maas) nematocyst venom-mediated biogenic synthesis of gold nanoparticles and its antiproliferative effects. Aquaculture International, 31(4):2235–2244, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01081-z.

Hoem:2023:STN

[2637] Kine Samset Hoem and Ann-Kristin Tveten. Sea transfer and net pen cleaning induce changes in stress-related gene expression in commercial Atlantic salmon (Salmo salar) gill tissue. Aquaculture International, 31 (4):2245–2262, August 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01084-w.

Gao:2023:CNQ

[2638] Xin Gao, Haojie Zhai, Zuxiang Peng, Jinxi Yu, Lin Yan, Wei Wang, Tongjun Ren, and Yuzhe Han. Comparison of nutritional quality, flesh quality, muscle cellularity, and expression of muscle growth-related genes between wild and recirculating aquaculture system (RAS)-farmed black rockfish (Sebastes schlegelii). Aquaculture International, 31(4): 2263–2280, August 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01087-7.

Zhou:2023:CCH

[2639] Huihua Zhou, Chunlei Gai, Jie Liu, La Xu, Haipeng Cao, and Jian An. Compound Chinese herbal extract (RMCP) supplementation improves defense against Aeromonas veronii infection in Chinese mitten crab Eriocheir sinensis. Aquaculture International, 31(4):2281– 2295, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01086-8.

Tawfik:2023:PIN

[2640] M. A. Tawfik, M. A. Salem, and R. I. Zaki. Performance investigation of a novel design of vertical micro-screen drum filter for a recirculating aquaculture system (RAS). Aquaculture International, 31(4): 2297–2322, August 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01085-9.

Huynh-Phuoc:2023:IAA

[2641] Vinh Huynh-Phuoc, Wahyu Purbiantoro, Thuyen Quyen Ly, Fernando Afonso, Ngoc Ut Vu, and Ta-Chih Cheng. The inhibitory ability and adhesion properties of a novel probiotic Bacillus safensis strain VQV8 against Vibrio spp. Aquaculture International, 31(4):2323–2337, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01088-6.

Ni:2023:ECA

[2642] Qian Ni, Jialing Liu, Xiao Huang, Hongxing Ge, Zhiguo Dong, and Yongxing Peng. Effects of chronic ammonia nitrogen stress on hydrolases and interleukin 17-3 (IL-17-3 ) in clam Cyclina sinensis. Aquaculture International, 31(4):2339–2354, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01090-y.

Elias:2023:PLB

[2643] Nurul Ashikin Elias, Mohamad Sofi Abu Hassan, Nor Asma Husna Yusoff, Okomoda Victor Tosin, Noor Aniza Harun, Sharifah Rahmah, and Marina Hassan. Potential and limitation of biocontrol methods against vibriosis: a review. Aquaculture International, 31(4):2355– 2398, August 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01091-x.

Mallik:2023:CPA

[2644] Sumanta Kumar Mallik, Shivam Singh, Neetu Shahi, Richa Pathak, Krishna Kala, Partha Das, Bhupendra Singh, Abhay Kumar Giri, Suresh Chandra, Debajit Sarma, and Pramod Kumar Pandey. Characterization and pathological analysis of Flavobacterium tructae recovered from farmed rainbow trout, Oncorhynchus mykiss (Walbaum, 1792), in the Indian Himalayan Region. Aquaculture International, 31(4): 2399–2420, August 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01089-5.

Radwan:2023:SBS

[2645] Mahmoud Radwan, Mahmoud A. El-Sharkawy, Ahmed N. Alabssawy, Sara F. Ghanem, Amaal Mohammadein, Jamila S. Al Malki, Asma W. Al-Thomali, Eman A. Manaa, Ragab A. Soliman, Shahd Yassir, Alsayed E. Mekky, Mansour A. E. Bashar, and Kareem F. Darweesh. The synergy between serious parasitic pathogens and bacterial infestation in the cultured Nile tilapia (Oreochromis niloticus): a severe threat to fish immunity, causing mass mortality and significant economic losses. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01093-9.

Zaki:2023:AZP

[2646] Mohamed A. A. Zaki, Hala Saber Khalil, Belal W. Allam, Riad H. Khalil, Mohammed F. El Basuini, Abd El-Aziz M. Nour, Eman M. H. Labib, Islam S. E. Elkholy, Marc Verdegem, and Hany M. R. Abdel-Latif. Assessment of zootechnical parameters, intestinal digestive enzymes, haemato-immune responses, and hepatic antioxidant status of Pangasianodon hypophthalmus fingerlings reared under different stocking densities. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01092-w.

Radwan:2023:RRP

[2647] Mahmoud Radwan, Kareem F. Darweesh, Sara F. Ghanem, Yasser Abdelhadi, Zana H. Kareem, Annie Christianus, Murni Karim, Rania M. Waheed, and Mahmoud A. El-Sharkawy. Regulatory roles of pawpaw (Carica papaya) seed extract on growth performance, sexual maturity, and health status with resistance against bacteria and parasites in Nile tilapia (Oreochromis niloticus). Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01094-8.

Parra:2023:CFF

[2648] Alini Beloto Parra, João Henrique Pinheiro Dias, Hugo Marques, Juan Antonio Balbuena, and Igor Paiva Ramos. Cage fish farming as a driver of changes in the functional diversity and structure of ichthyofauna in a neotropical reservoir. Aquaculture International, 31(5): ??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01096-6.

Li:2023:FDM

[2649] Lei Li, Guosheng Shi, and Tao Jiang. Fish detection method based on improved YOLOv5. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01095-7.

Rizkita:2023:SSA

[2650] Meidina Rizkita, Mia Rosmiati, Magdalena Lenny Situmorang, Made Dendy Pratama, Syafira Rosefa, and Gede Suantika. Sustainability status analysis and strategy development for common carp (Cyprinus carpio L.) hatchery industry in Ciparay District, West Java, Indonesia. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-023-01097-5.

Davis:2023:ELC

[2651] Barbara A Davis and Malcolm D Devine. Evaluation of long-chain omega-3 canola oil on Atlantic salmon growth, performance, and essential fatty acid tissue accretion across the life cycle: a review. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01099-3.

Wang:2023:EDP

[2652] Linyue Wang, Yanyu Yao, Zhen Zhao, Lu Li, Hanling Gan, Hesham E. Desouky, Xuexi Wang, and Ziping Zhang. Effect of dietary protein, vitamin E, and vitamin C levels on the growth and gonad development of hybrid male abalone (Haliotis fulgens [male sign] × H. discus hannai [female sign]) by the orthogonal array design. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01100-z.

Bustamam:2023:PED

[2653] Muhammad Safwan Ahamad Bustamam, Khozirah Shaari, Chong Chou Min, Md Shirajum Monir, Hamza Ahmed Pantami, and Intan Safinar Ismail. Prophylactic effects of dietary Isochrysis galbana on immersionchallenged red hybrid tilapia (Oreochromis spp.) with Streptococcus agalactiae by spleen-NMR-metabolomics. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01098-4.

Antunes:2023:BMA

[2654] Mariana Antunes, Samara de Paiva Barros-Alves, Douglas Fernandes Rodrigues Alves, Rafael Augusto Gregati, and Maria Lucia Negreiros-Fransozo. Broodstock management of the arrow crab Stenorhynchus seticornis and effects of temperature and salinity on larvae survival. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01104-9.

Wang:2023:IAC

[2655] Zhenlu Wang, Guowei Liao, Bing Chen, and Lanfen Fan. Impacts of acute cold-stress in Pacific white shrimp Litopenaeus vannamei: investigating the tissue-specific stress resistance response. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01106-7.

Cardoso:2023:ELS

[2656] Lucas Cardoso, Marco Shizuo Owatari, Francisco Célio Maia Chaves, William Eduardo Furtado, Luciana Aparecida Honorato, Jonathan Paulo Agnes, Daniela Coelho dos Santos, Rozangela Curi Pedrosa, Silvia Terra Fontes, José Luiz Pedreira Mouriño, and Maurício Laterça Martins. Lippia sidoides essential oil (Verbenaceae) improves inflammatory response and histological condition in Danio rerio. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01103-w.

Machado:2023:RCM

[2657] Patrícia C. Machado, Bruno Pinto, and Natacha Nogueira. Regional communication and media analysis of aquaculture in Atlantic islands. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01101-y.

Zheng:2023:SAC

[2658] Xinxin Zheng, Changsheng Shao, Jianxia Zhu, Leisheng Zhang, and Qing Huang. Study of the antioxidant capacity of astaxanthin in cells against radiation-induced strong oxidative stress. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01105-8.

Perveen:2023:DEP

[2659] Summia Perveen, Noor Khan, Fei Yin, Fayyaz Rasool, Chunlin Wang, Mahrukh Butt, and Misbah Irm. Development of Pangasius hypophthalmus polyculture production in Pakistan when cultured with freshwater and Chinese carps. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01102-x.

Beg:2023:QEO

[2660] Mirza Masum Beg, Subha M. Roy, Sanjib Moulick, and Basudev Mandal. Quality evaluation of organically farmed fish fillet of Indian major carps. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01107-6.

deRezende:2023:UUI

[2661] Cícero Eduardo de Rezende, Caio Augusto Perazza, Rilke Tadeu Fonseca de Freitas, Eric Hallerman, and Alexandre Wagner Silva Hilsdorf. Use of ultrasonographic imaging for non-invasive carcass yield prediction in Nile tilapia (Oreochromis niloticus). Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01109-4.

Zhou:2023:OSA

[2662] Yan Zhou, Jie Zheng, Jun Zhao, Shuang Li, Jie Xing, Chunqing Ai, Chenxu Yu, Sheng Yang, and Jingfeng Yang. Oxygenated storage alleviates autolysis of the sea cucumber Apostichopus japonicus during transport. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01108-5.

Rajasekaran:2023:ABA

[2663] Jayaseelan Rajasekaran and Pragasam Viswanathan. Anti-bacterial and antibiofilm properties of seaweed polysaccharide-based nanoparticles. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01111-w.

Banerjee:2023:EAM

[2664] Risav Banerjee, Rajdeep Das, Amrutha VM, Asha Devi S, and Sudhakaran R. Effect of algal metabolites in the inhibition of angiogenic channels for breast tumour. Aquaculture International, 31(5): ??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01110-x.

Wang:2023:RIA

[2665] Pengchao Wang, Wenyi Zhang, Yiwei Xiong, Tianyong Chen, Sufei Jiang, Hui Qiao, Yongsheng Gong, Yan Wu, Shubo Jin, and Hongtuo Fu. RNA interference analysis of the potential functions of cyclin-dependent kinase 2 in sexual reproduction of male oriental river prawns (Macrobrachium nipponense). Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01112-9.

Aly:2023:AVU

[2666] Salah M. Aly, Mohamed E. Abou-El-Atta, Heba S. El-Mahallawy, Ahmed Elaswad, Fatma A. ElAbyad, and Noha I. ElBanna. Aeromonas veronii and ulcerative syndrome in cultured Nile tilapia (Oreochromis niloticus) and their associated factors. Aquaculture International, 31(5): ??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01113-8.

Sandeep:2023:IDM

[2667] K. P. Sandeep, T. Sivaramakrishnan, S. Sudhin, J. A. J. Raymond, N. S. Sudheer, R. Ananda Raja, Sujeet Kumar, J. Syama Dayal, C. P. Balasubramanian, Paramita Banerjee Sawant, N. K. Chadha, and K. Ambasankar. Influence of dietary microalgal concentrates on growth, survival and health status of Penaeus vannamei. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01114-7.

Dhande:2023:PRU

[2668] Kranthi Kumar Dhande, Rama Sharma, Ananthan P. S., and Vinay A. Profitability and resource use efficiency of polyculture system in Andhra Pradesh. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01115-6.

Wang:2023:GDS

[2669] Xinyao Wang, Yubing Xu, and Luyao Wang. Growth dynamics and sustainable development of aquatic products export trade of China and Vietnam. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01119-2.

Debnath:2023:TAE

[2670] Sanjit Chandra Debnath, Jamie McMurtrie, Ben Temperton, Jérôme Delamare-Deboutteville, Chadag Vishnumurthy Mohan, and Charles R. Tyler. Tilapia aquaculture, emerging diseases, and the roles of the skin microbiomes in health and disease. Aquaculture International, 31 (5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01117-4.

Lebata-Ramos:2023:MAP

[2671] Ma. Junemie Hazel L. Lebata-Ramos. Molluscan aquaculture in the Philippines: a review. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-023-01120-
9.

Li:2023:FCM

[2672] Ruojing Li, Mantang Xiong, Wang Li, Wei Li, Jiashou Liu, and Tanglin Zhang. The feasibility of Chinese mitten crab (Eriocheir sinensis), mandarin fish (Siniperca chuatsi), and mud carp (Cirrhinus molitorella) polyculture and their effects on growth performance, water quality, antioxidant, and nonspecific immunological capacity. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01118-3.

Diken:2023:WDF

[2673] Gürkan Diken, Hayati Koknaroglu, and Ergi Bahrioğlu. Would dependency on fossil fuels affect food security and sustainable production in aquaculture? Cultural energy use and energy use efficiency for earthen pond European seabass (Dicentrarchus labrax L., 1758) production. Aquaculture International, 31(5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01242-0.

Singh:2023:DGW

[2674] Rajeev Kumar Singh, Raj Kumar, Anamika Bhordia, Sangeeta Mandal, G. Kantharajan, Rejani Chandran, T. T. Ajithkumar, Vindhya Mohindra, and Kuldeep Kumar Lal. Development of genome-wide novel microsatellite markers and genetic variability analysis of Asian seabass, Lates calcarifer (Bloch, 1790). Aquaculture International, 31 (5):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01116-5.

Mondal:2023:AAE

[2675] Haimanti Mondal, Natarajan Chandrasekaran, Amitava Mukherjee, and John Thomas. Antibacterial activity of Bacillus licheniformis isolated from marine sediments and its effect in treating Aeromonas hydrophila infection in freshwater prawn, Macrobrachium rosenbergii. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01121-8.

Elangovan:2023:PCB

[2676] Sujatha Elangovan and Sivakumar Arumugam. Purification, characterization, and biological activities of melanin pigment isolated from Indian squid Uroteuthis duvaucelii. Aquaculture International, 31(6): ??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01158-9.

Sowmiya:2023:AAN

[2677] Prasad Sowmiya, Tharmathass Stalin Dhas, Dhinakarasamy Inbakandan, Ravi Mani, Anandakumar Natarajan, Gopal Dharani, Kasivelu Govindaraju, Malaichamy Kannan, Karthick Velu, and Chirayil Meethalepurayil Vineeth Kumar. Antagonistic activity of a novel chitosan-selenium nanoflower against common aquaculture pathogen Aeromonas caviae. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01146-z.

Velmurugan:2023:HHE

[2678] Aswini Velmurugan and Gothandam Kodiveri Muthukaliannan. Homologous and heterologous expression of phytoene synthase gene in marine microalgae Dunaliella salina and its potential as aquaculture feed. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01195-4.

Parra-Riofrio:2023:ETS

[2679] Geovanna Parra-Riofrio, Patricia Moreno, Esther García-Rosado, M. Carmen Alonso, Eduardo Uribe-Tapia, Roberto Teófilo Abdala-Diaz, and Julia Bejar. Tetraselmis suecica and Porphyridium cruentum exopolysaccharides show anti-VHSV activity on RTG-2 cells. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.com/

article/10.1007/s10499-023-01202-8.

Gayathri:2023:VIE

[2680] G. A. Gayathri, P. Kavya, D. Ashwini, Eshika Chakraborty, Idris Adewale Ahmed, and Gayathri Mahalingam. Vitexin isolated from Acanthus ilicifolius L. leaf enhances GLUT-4 translocation in experimental diabetic rats. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01235-z.

Karthik:2023:EBP

[2681] T. Karthik and M. A. Jayasri. Exploration of biostimulant property of seaweed liquid extract (SLE) on the growth development and yield of Solanum lycopersicum (tomato). Aquaculture International, 31(6): ??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01252-y.

Ramya:2023:OAL

[2682] P. Ramya, K. Selvaraj, K. Suthendran, K. Sundar, and B. Vanavil. Optimization of alginate lyase production using Enterobacter tabaci RAU2C isolated from marine environment by RSM and ANFIS modelling. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01302-5.

Unnikrishnan:2023:EMR

[2683] P. S. Unnikrishnan, K. Suthindhiran, and M. A. Jayasri. Effect of marine red algal (Portieria hornemannii) extracts on starch digestion rate and its possible role in diabetes management. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01311-4.

Hastuti:2023:ANB

[2684] Yuni Puji Hastuti, Anggi Siregar, Yuli Siti Fatma, and Eddy Supriyono. Application of a nitrifying bacterium Pseudomonas sp. HIB D to reduce nitrogen waste in the Litopenaeus vannamei cultivation environment. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01123-6.

Zhao:2023:SIR

[2685] Lu Zhao, Jinqiang Huang, Yongjuan Li, Shenji Wu, and Yujun Kang. Skin immune response of rainbow trout (Oncorhynchus mykiss) infected with infectious hematopoietic necrosis virus. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01122-7.

deFaria:2023:ELW

[2686] Camila de Fátima Pereira de Faria, Allan Emilio Piedade, and Elisabeth Criscuolo Urbinati. Effect of low water temperature on the stress, innate immune, and antioxidant responses of pacu, Piaractus mesopotamicus, a sub-tropical fish. Aquaculture International, 31(6): ??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01125-4.

Attia:2023:HIO

[2687] Marwa M. Attia, Mustafa M. Ibrahim, and Mahmoud A. Mahmoud. Heavy infection of the orange-spotted grouper (Epinephelus coioides) with Huffmanela japonica: morphological, ultrastructural identification, tissue reactions and immunological analysis. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01124-5. See correction [2754].

Wang:2023:THA

[2688] Zhengfei Wang, Yayun Guan, Yue Wang, Shang Zhu, Chong Cui, and Xinyu Wang. Transcriptome and histopathology analyses of the gills of Eriocheir sinensis provide novel insights into the molecular mechanism of Pb stress. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01126-3.

Purbiantoro:2023:IHH

[2689] Wahyu Purbiantoro, Vinh Huynh-Phuoc, Khanh Van Nguyen, Omkar Vijay Byadgi, and Ta-Chih Cheng. The immunocompetence of the hemocytes, hepatopancreas, and midgut of the giant freshwater prawn Macrobrachium rosenbergii in response to reverse gavage stimulation with CpGoligodeoxynucleotides. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL

https://link.springer.com/article/10.1007/s10499-023-01127-
2.

Intriago:2023:AME

[2690] Pablo Intriago, Andres Medina, Jorge Espinoza, Xavier Enriquez, Kelly Arteaga, Luis Fernando Aranguren, Andrew P. Shinn, and Xavier Romero. Acute mortality of Penaeus vannamei larvae in farm hatcheries associated with the presence of Vibrio sp. carrying the Vp PirAB toxin genes. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01129-0.

Jiang:2023:IRL

[2691] Shuiqing Jiang, Xiaohong Huang, Ting Li, Yinan Zhang, and Jingwei Zhang. Immune response of large yellow croaker Larimichthys crocea towards a recombinant vaccine candidate targeting the parasitic ciliate Cryptocaryon irritans. Aquaculture International, 31(6): ??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01131-6.

Yu:2023:SIA

[2692] Shu Yu, Xiaomei Hou, Yingxue Zhang, and Yongtong Mu. Status investigation and analysis of oyster aquaculture industry in China: evidence from Zhanjiang. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01132-5.

Kim:2023:EDM

[2693] Han-Se Kim, Sehee Jo, Kwan-Sik Yun, and Kyeong-Jun Lee. Effects of dietary micelle silymarin on the growth performance, feed utilization and health of olive flounder (Paralichthys olivaceus). Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01135-2.

Monier:2023:EEB

[2694] Mohamed N. Monier, Hoda Kabary, Amal Elfeky, Saadea Saadony, Nadia N. B. Abd El-Hamed, Moaheda E. H. Eissa, and El-Sayed Hemdan Eissa. The effects of Bacillus species probiotics (Bacillus subtilis and B. licheniformis) on the water quality, immune responses, and resistance of whiteleg shrimp (Litopenaeus vannamei) against Fusarium solani infection. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01136-1.

Puvanendran:2023:OIC

[2695] Velmurugu Puvanendran, Trilochan Swain, Helge Tveiten, Øyvind J. Hansen, and Atle Mortensen. Optimizing intensive culture protocols for Atlantic cod (Gadus morhua) larvae. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01133-4.

Mota:2023:UAY

[2696] Lucas dos Santos Mota, Otávio Mesquita de Sousa, Danilo Araujo Soares Pereira, Maria Fernanda Guedes Pessoa, Marcelo Barbosa Henriques, and Eduardo Gomes Sanches. Urban aquaculture of yellow neon goby (Elacatinus figaro): a vulnerable marine ornamental fish. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01134-3.

Quintero-Martinez:2023:OFO

[2697] Gabriel Ángel Quintero-Martínez, Crisantema Hernández, Elena Palacios, María Cristina Chávez-Sánchez, Leonardo Ibarra-Castro, and Miguel Ángel Hurtado-Oliva. Oxidized fish oil in the diet negatively affect rearing performance, health, and tissue fatty acid composition of juvenile spotted rose snapper Lutjanus guttatus (Steindachner, 1869). Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01137-0.

Wahyudi:2023:EDS

[2698] Imam Tri Wahyudi, Dedi Jusadi, Mia Setiawati, and Julie Ekasari. Effects of dietary supplementation with cinnamon powder and lysine on blood chemistry, liver histology, growth performance, and fillet quality of striped catfish Pangasianodon hypophthalmus. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01141-4.

Abgoun:2023:IIL

[2699] Kianoush Abgoun, Iman Sourinejad, and Arash Akbarzadeh. Influence of induced low-, moderate-, and severe-intensity starvation on compensatory growth, body composition, and economic efficiency of Pacific white shrimp Litopenaeus vannamei. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01140-5.

Soudah:2023:OFN

[2700] Boma Soudah, N’Feidé Toï, Lombo Yao, Batimsoga Bougra, Kossoga Assota, Bataka Koutera, Imorou Toko Ibrahim, and Simion Kipkemboi Omasaki. Optimization of feed for Nile tilapia (Oreochromis niloticus) using an Excel programming model in small-scale feed-mixing operations in Togo. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01138-z.

Bodenstein:2023:SAH

[2701] Sarah Bodenstein, Isabelina Nahmens, Brian R. Callam, and Terrence R. Tiersch. Simulation analysis of high-throughput oyster cryopreservation at three scales of production. Aquaculture International, 31(6): ??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01139-y.

Bivareh:2023:LCT

[2702] Mohammadreza Bivareh. Lysozyme chicken type (Lys-c) expression in Persian sturgeon (Acipenser persicus) using real-time PCR. Aquaculture International, 31(6):??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01142-3.

Ran:2023:DSW

[2703] Xun Ran, Beibei Li, Daoliang Li, Jianping Wang, and Qingling Duan. Detection of surfacing white shrimp under hypoxia based on improved lightweight YOLOv5 model. Aquaculture International, 31(6): ??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01149-w.

Li:2023:DPO

[2704] Mingzhi Li, Hongbo Lu, Gongzhi Yu, Ying Liu, Jiequn Wang, and Zhibao Wu. Design and parameter optimization of a submerged water jet scallop skirt cleaning device. Aquaculture International, 31(6): ??, ???? 2023. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01147-y.

Kaimakoudi:2024:PIT

[2705] Eleni Kaimakoudi. Policy initiatives towards enhancing consumer knowledge and tackling consumer confusion in aquaculture sector. Aquaculture International, 32(1):1–9, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01143-2.

Saputra:2024:EDB

[2706] Ishaaq Saputra and Ravi Fotedar. The effect of defatted black soldier fly meal (Hermetia illucens) inclusion in the formulated diet on the growth, gene expression, and histopathology of juvenile lobster (Panulirus ornatus Fabricius, 1798). Aquaculture International, 32(1): 11–29, February 2024. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01151-2.

Liu:2024:EAE

[2707] Guoqing Liu, Meng Zhou, Xiangjie Mao, Dianchao Gu, Wangwang Chen, Xianmei Long, Shouqi Xie, and Qingsong Tan. Evaluation of the appropriate Clostridium autoethanogenum protein level in grass carp (Ctenopharyngodon idellus) diets by growth performance, health status, and intestinal microbiota. Aquaculture International, 32(1):31– 59, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01163-y.

Juarez:2024:PTE

[2708] Oscar E. Juárez, Clara E. Galindo-Sánchez, Fabiola Lafarga-De la Cruz, Sara Enciso, Edgar A. López-Landavery, Camilo Muñoz, Felipe Aguilera, and Juan Pablo Lazo. Physiological and transcriptomic effects of formulated diets including the prebiotics inulin, β-glucan, and chitosan on juveniles of Totoaba macdonaldi. Aquaculture International, 32(1): 61–85, February 2024. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01144-1.

Pellegrin:2024:GPH

[2709] Lucas Pellegrin, Carlos Eduardo Copatti, Lilian Fiori Nitz, Daniel de Sá Britto Pinto, Wilson Wasielesky, and Luciano Garcia. Growth performance and hematological parameters of pacu (Piaractus mesopotamicus) juveniles in different concentrations of total suspended solids in the BFT system. Aquaculture International, 32(1):87–100, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01161-0.

Yigit:2024:AEP

[2710] Ümüt Yigit, Murat Yigit, Sebahattin Ergün, Halit Kusku, Hüseyin Ek, and Masashi Maita. Analysis of the economic performance of salmon farming in submerged and surface cages in the Black Sea. Aquaculture International, 32(1):101–118, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01155-y.

deJesus:2024:EFC

[2711] Diana Pérez de Jesús, Martha Patricia Hernández-Vergara, Carlos Iván Pérez-Rostro, and Carlos Alfonso Frías-Quintana. Effect of fasting on compensatory growth and digestive enzymatic activity of freshwater prawn post larvae (Macrobrachium rosenbergii) during its culture in biofloc. Aquaculture International, 32(1):119–135, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01154-z.

He:2024:SEA

[2712] Li jun He, Zhen Yang, Zhao wei Zhong, Yan Xu, Yan Feng, Zhen Ding, Hong wei Liu, and Yong hua Jiang. Screening and expression analysis of genes related to skin coloration in pearlscale angelfish (Centropyge vrolikii). Aquaculture International, 32(1):137–159, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01145-0.

Akimov:2024:EPS

[2713] Arkady I. Akimov, Ekaterina S. Solomonova, and Natalia Yu. Shoman. Estimation physiological state and carotenoid content of Dunaliella salina (Teod.) using flow cytometry and variable fluorescence methods. Aquaculture International, 32(1):161–174, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01153-0.

Chi:2024:GPG

[2714] Yong Chi, Qi Li, and Chengxun Xu. Genetic parameters and genotype by environment interactions for harvest traits in the Pacific oyster (Crassostrea giga s). Aquaculture International, 32(1):175–195, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01159-8.

Shen:2024:ASP

[2715] Yaqin Shen, Chen Chen, Pengquan Li, Xianfeng Huang, and Yubao Li. Application of a smart pilot electrochemical system for recycling aquaculture seawater. Aquaculture International, 32(1):197–213, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01150-3.

Basumatary:2024:GRT

[2716] Bwsrang Basumatary, A. K. Verma, Saket Kushwaha, and Manoj Kumar Verma. Global research trends and performance measurement on biofloc technology (BFT): a systematic review based on computational techniques. Aquaculture International, 32(1):215–240, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01162-z.

Liu:2024:EDE

[2717] Jing Liu, Sudong Xia, Keyong Jiang, Lihua Geng, Yang Yue, Ning Wu, Quanbin Zhang, and Jing Wang. Effects of dietary Saccharina japonica and Pyropia yezoensis enzymatic hydrolysate on growth, nonspecific immunity, and resistance to Vibrio harveyi in hybrid grouper (Epinephelus fuscoguttatus [female sign] × Epinephelus lanceolatu [male sign]). Aquaculture International, 32(1):241–262, February 2024. CODEN AQINFS.

ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01148-x.

daCosta:2024:USI

[2718] Jesaías Ismael da Costa, Dalton José Carneiro, Ana Cristina Oliveira, Rosângela Kiyoko Jomori, and Maria Inez Martins. Using stable isotopes (C and N) to determine plankton contribution in tambaqui juvenile growth in earthen ponds. Aquaculture International, 32(1):263– 273, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01152-1.

Lemos:2024:AAA

[2719] Daniel Lemos, Rafael Coelho, and Rodrigo Carvalho. Apparent amino acid digestibility of feed ingredients for juvenile shrimp (Litopenaeus vannamei): a new method of determination using soybean meal as an example. Aquaculture International, 32(1):275–297, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01157-w.

Wu:2024:FEE

[2720] Leiming Wu, Guangming Han, Jiahong Zhang, Shouhong Wang, Baoli Qin, Xiangming Kou, Linjie Ma, Jianhua Bi, Qin Yuan, and Hejun Tang. Feeding effect of Alternanthera philoxeroides and Elodea nuttallii on crayfish (Procambarus clarkii) growth, hepatopancreas enzyme activity, and muscle quality. Aquaculture International, 32(1): 299–314, February 2024. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01156-x.

Mbokane:2024:SRM

[2721] Esau Matthews Mbokane and Ngonidzashe Adreck Gukuta Moyo. A systematic review and meta-analysis of the potential effect of medicinal plants on innate immunity of selected freshwater fish species: its implications for fish farming in Southern Africa. Aquaculture International, 32 (1):315–335, February 2024. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01160-1.

Ahmadmoradi:2024:EEE

[2722] Maryam Ahmadmoradi, Mojtaba Alishahi, Siavash Soltanian, Ali Shahriari, and Azadeh Yektaseresht. Effects of encapsulation of Lactobacillus plantarum on probiotic potential and reducing lead toxicity in rainbow trout (Oncorhynchus mykiss Walbaum). Aquaculture International, 32(1):337–359, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01164-x.

Vega-Carranza:2024:SIR

[2723] Ana S. Vega-Carranza, Ruth Escamilla-Montes, Antonio Luna-González, Jesús A. Fierro-Coronado, Genaro Diarte-Plata, and Cipriano García-Gutiérrez. Survival, immune response, and gut microbiota in Litopenaeus vannamei fed with synbiotics and postbiotics and challenged with Vibrio parahaemolyticus. Aquaculture International, 32(1):361– 381, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01165-w.

Thompson:2024:CNM

[2724] Cameron R. S. Thompson, Angelico Madaro, Jonatan Nilsson, Lars Helge Stien, Frode Oppedal, Øyvind Øverli, Wayne J. Korzan, and Samantha Bui. Comparison of non-medicinal delousing strategies for parasite (Lepeophtheirus salmonis) removal efficacy and welfare impact on Atlantic salmon (Salmo salar) hosts. Aquaculture International, 32(1): 383–411, February 2024. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01167-8.

Espinoza-Ortega:2024:FFE

[2725] Manuel Espinoza-Ortega, César Molina-Poveda, Miguel Jover-Cerdá, and Roberto Civera-Cerecedo. Feeding frequency effect on water quality and growth of Litopenaeus vannamei fed extruded and pelleted diets. Aquaculture International, 32(1):413–429, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01166-9.

Sicuro:2024:FMM

[2726] B. Sicuro, B. Castelar, C. Bergamino, S. Mioletti, S. Squadrone, A. Griglione, M. Falzone, E. Colombino, and M. T. Capucchio. Freshwater mussel meal as new alternative ingredient for rainbow trout (Oncorhynchus mykiss) feeds: growth performance and histomorphological analyses. Aquaculture International, 32(1):431–445, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01168-7.

Lu:2024:SCN

[2727] Fei-Fei Lu, Hong-Chang Ding, and Xing-Hong Yan. Selection and characterization of a new strain with thin blades and high-temperature resistance in Neoporphyra haitanensis. Aquaculture International, 32(1): 447–460, February 2024. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01170-z.

Vargas:2024:ATB

[2728] Luana Búrigo Vargas, Thais Agda R. da C. Primo, Rafaela Dexcheimer Alves, Fabio de Farias Neves, Viviane Trevisan, and Everton Skoronski. Application of a tannin-based coagulant for harvesting Parachlorella kessleri microalgae cultivated in chicken meat thermal processing wastewater. Aquaculture International, 32(1):461–476, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01171-y.

Abdelhamid:2024:IDL

[2729] Ahmed F. Abdelhamid, Ahmed G. A. Gewida, Abdel-Fattah M. El-Sayed, and Mohamed F. Badran. Impacts of different levels of vitamin K on the growth performance, hematological parameters, and immunological response of juvenile Nile tilapia (Oreochromis niloticus). Aquaculture International, 32(1):477–488, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01169-6.

Tan:2024:COS

[2730] Sin-Ying Tan, Sumathi Sethupathi, Kah-Hon Leong, and Tanveer Ahmad. Challenges and opportunities in sustaining aquaculture industry in Malaysia. Aquaculture International, 32(1):489–519, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01173-w.

Dominguez-May:2024:OSH

[2731] Roger Domínguez-May, Patricia Borrego-Kim, and Iván Velázquez-Abunader. Optimization of stocking and harvesting strategies in intensive culture of tilapia (Oreochromis niloticus), considering minimum marketable sizes. Aquaculture International, 32(1):521–544, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01172-x.

Elbialy:2024:DTR

[2732] Zizy I. Elbialy, Abdullah S. Salah, Ahmed Elsheshtawy, Nasema M. Elkatatny, Alamira Marzouk Fouad, and Haitham G. Abo-Al-Ela. Differential tissue regulation of nrf2 /keap1 crosstalk in response to Aeromonas infection in Nile tilapia: a comparative study. Aquaculture International, 32(1):545–562, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01175-8.

Gao:2024:TAC

[2733] Jinwei Gao, Xudong Lyu, Wenyu Lu, Shao Peng, Xuying Jia, Wenli Zhou, and Jun Kang. Transcriptional analysis combined with intestinal microbiota sequencing to unveil the effects of astaxanthin from Haematococcus pluvialis on pearl gentian grouper (Epinephelus fuscoguttatus [female sign] × Epinephelus lanceolatu s [male sign]). Aquaculture International, 32(1):563–579, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01176-7.

Eissa:2024:PET

[2734] Alaa Eldin Eissa, Marwa M. Attia, Rabia A. El Zlitne, Ayad A. Magdy, Asmaa Edrees, Mahmoud S. Sharaf, Abeer E. Mahmoud, Awad A. Abdelbaky, Rehab R. Abd ElMaged, Elshaimaa Ismael, Rehab A. Qorany, Doaa F. El Moghazi, Abdelbary Prince, Emad A. Afiffy, Said El Behiri, and Nehal A. Younis. The puzzling etiologies of transient black discoloration in Nile tilapia (Oreochromis niloticus) intensively cultured under RAS system. Aquaculture International, 32(1): 581–592, February 2024. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01328-9.

Liang:2024:CST

[2735] Yuanxin Liang, Geng Cheng, Xianchao Bai, Jianmin Zhou, Haining Zhang, Yong Chi, Gaowei Jiang, Chengxun Xu, and Qi Li. Comparative study on tetraploid induction of the Fujian oyster Crassostrea angulata utilizing three typical methods. Aquaculture International, 32(1): 593–612, February 2024. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01174-9.

Ghosh:2024:OAE

[2736] Alokesh Kumar Ghosh, Sheikh Shaon Ahmmed, H. M. Rakibul Islam, Md. Abir Hasan, Ghausiatur Reza Banu, Sujogya Kumar Panda, Liliane Schoofs, and Walter Luyten. Oral administration of Zingiber officinale and Aegle marmelos extracts enhances growth and immune functions of the shrimp Penaeus monodon against the white spot syndrome virus (WSSV). Aquaculture International, 32(1):613–632, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01177-6.

Marchao:2024:EDT

[2737] Rafael Silva Marchão, Carlos Eduardo Copatti, Felipe Barbosa Ribeiro, Marcos Antonio Delmondes Bomfim, Maylanne Sousa de Lima, Vanessa Ferreira Batista, Aline da Silva Rocha, Gilmar Amaro Pereira, Thaisa Sales Costa, David Ramos da Rocha, Luiz Vítor Oliveira Vidal, and José F. Bibiano Melo. Evaluation of dietary tryptophan requirement on growth, whole-body composition, and hematobiochemical parameters of tambaqui (Colossoma macropomum) in the fattening phase. Aquaculture International, 32(1):633–652, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01179-4.

He:2024:EDB

[2738] Gui-Lun He, Meng-Lin Shi, Yu-Cheng Liu, Liu-Tong Chen, Xin Chen, Wen-Bo Zhu, Zheng-Bang Chen, Bei-Ping Tan, and Shi-Wei Xie. Effects of dietary betaine supplementation on growth performance, feed intake, intestinal histology, lipid metabolism, and immune response of black tiger shrimp (Penaeus monodon) fed diets containing two levels of raw feed attractants. Aquaculture International, 32(1):653–673, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-

143X (electronic). URL https://link.springer.com/article/10.
1007/s10499-023-01180-x.

Kumar:2024:SSD

[2739] Manmohan Kumar, Narinder Kumar Chadha, Satya Prakash, Annam Pavan-Kumar, Vungarala Harikrishna, P. Gireesh-Babu, and Gopal Krishna. Salinity, stocking density, and their interactive effects on growth performance and physiological parameters of white-leg shrimp, Penaeus vannamei (Boone, 1931), reared in inland ground saline water. Aquaculture International, 32(1):675–690, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01181-w.

Akter:2024:EEM

[2740] Urmi Akter, Shaharior Hashem, Farhabun Binte Farhad, K. M. Shakil Rana, Mohammad Mahfujul Haque, and M. A. Salam. Effects of Epsom (magnesium sulfate) salt on growth performance of water spinach (Ipomoea aquatica) in nutrient film technique and media-based aquaponics systems. Aquaculture International, 32(1):691–707, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X

(electronic). URL https://link.springer.com/article/10.1007/
s10499-023-01183-8.

Chuphal:2024:EPR

[2741] Bhawna Chuphal, Umesh Rai, and Brototi Roy. Exploring the possible role of estradiol in the regulation of reproductive phase-dependent expression of NOD-like receptors in female Channa punctata. Aquaculture International, 32(1):709–724, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01184-7.

Wang:2024:EDG

[2742] Yuehan Wang, Zuxiang Peng, Lin Yan, Xin Gao, Lin Wu, Shuchang Cui, Mengyuan Zhou, Tongjun Ren, Wei Wang, Shuhao Sun, Yuzhe Han, and Hang Sun. Effects of dietary glutamine supplementation on growth performance, intestinal digestive ability, antioxidant status and hepatic lipid accumulation in Xenocypris davidi (Bleeker,1871). Aquaculture International, 32(1):725–743, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01187-4.

Gopalraaj:2024:EDS

[2743] Jhanani Gopalraaj, Krishnakumar Velayudhannair, John Paul Arockiasamy, and Divya Kandathil Radhakrishnan. The effect of dietary supplementation of proteases on growth, digestive enzymes, oxidative stress, and intestinal morphology in fishes – a review. Aquaculture International, 32(1):745–765, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01191-8.

Hattori:2024:MMA

[2744] Jahina Fagundes de Assis Hattori, Márcia Regina Piovesan, Denis Rogério Sanches Alves, Suzana Raquel de Oliveira, Ricácio Luan Marques Gomes, Fábio Bittencourt, and Wilson Rogério Boscolo. Mathematical modeling applied to fish feeding behavior. Aquaculture International, 32(1):767–774, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01186-5.

Wang:2024:PID

[2745] Qian Wang, Yanbin Lin, Heng Zhang, Wenhao Fan, Shengxuan Li, Guoliang Ruan, and Liu Fang. Positive impacts of dietary prebiotic inulin on growth performance, antioxidant capacity, immunity, and intestinal microbiota of red swamp crayfish (Procambarus clarkii). Aquaculture International, 32(1):775–794, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.springer.

com/article/10.1007/s10499-023-01188-3.

Shahi:2024:FRC

[2746] Neetu Shahi, Mohan Singh, Sumanta Kumar Mallik, Bhupendra Singh, Krishna Kala, Monalisa Sahoo, Debajit Sarma, and Pramod Kumar Pandey. First report of characterization and pathogenicity of Basidiobolus sp. Ind SN1 recovered from gastrointestinal basidiobolomycosis as an outbreak in a coldwater fish species rainbow trout, Oncorhynchus mykiss (Walbaum, 1792) in India. Aquaculture International, 32(1): 795–815, February 2024. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01190-9.

Nargesi:2024:ESC

[2747] Erfan Akbari Nargesi, Danial Gorouhi, and Bahram Falahatkar. Efficacy of sGnRHa in combination with domperidone on the sperm quality, seminal fluid metabolites, and fertilization ability of spermatozoa in Caspian kutum, Rutilus frisii. Aquaculture International, 32(1): 817–832, February 2024. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01189-2.

Gyan:2024:EDC

[2748] Watson Ray Gyan, Mpwaga Alatwinusa Yohana, Qihui Yang, Beiping Tan, Shuyan Chi, and Yuanming Yi. The effects of dietary L-carnitine supplementation in fish diet containing high corn gluten meal on immunity, lipid metabolism, and metabolomics in juvenile hybrid grouper ([female sign]Epinephelus fuscoguttatus × [male sign]Epinephelus lanceolatus). Aquaculture International, 32(1):833–869, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01193-6.

Wang:2024:EAH

[2749] Fei Wang, Zi-Le Qin, Wei-Sheng Luo, Ning-Xia Xiong, and Sheng-Wei Luo. Aeromonas hydrophila can modulate synchronization of immune response in gut-liver axis of red crucian carp via the breach of gut barrier. Aquaculture International, 32(1):871–885, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https:

//link.springer.com/article/10.1007/s10499-023-01185-6.

Mandal:2024:ADS

[2750] Arghya Mandal and Apurba Ratan Ghosh. AI-driven surveillance of the health and disease status of ocean organisms: a review. Aquaculture International, 32(1):887–898, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01192-7.

Islam:2024:OPC

[2751] Sk Injamamul Islam, Channarong Rodkhum, and Piyanan Taweethavonsawat. An overview of parasitic co-infections in tilapia culture. Aquaculture International, 32(1):899–927, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01198-1.

Eyayu:2024:CRP

[2752] Alamrew Eyayu, Abebe Getahun, and James Last Keyombe. Correction to: A review of the production status, constraints, and opportunities in East African freshwater capture and culture fisheries. Aquaculture International, 32(1):929–930, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01178-5. See [2626].

daSilva:2024:CED

[2753] Thamyres Vanessa N. da Silva, Camila F. dos Santos, Jessica M. L. dos Santos, Marcos J. Schmitz, Juan R. B. Ramírez, Marcelo F. Torres, Luis André L. Barbas, Luís A. Sampaio, Pablo E. Verde, Marcelo B. Tesser, and José M. Monserrat. Correction to: Effects of dietary inclusion of lyophilized açai berries (Euterpe oleracea) on growth metrics, metabolic and antioxidant biomarkers, and skin color of juvenile tambaqui (Colossoma macropomum). Aquaculture International, 32(1): 931–935, February 2024. CODEN AQINFS. ISSN 0967-6120 (print),

1573-143X (electronic). URL https://link.springer.com/article/
10.1007/s10499-023-01128-1. See [2569].

Attia:2024:CHI

[2754] Marwa M. Attia, Mustafa M. Ibrahim, and Mahmoud A. Mahmoud. Correction to: Heavy infection of the orange-spotted grouper (Epinephelus coioides) with Hufmanela japonica: morphological, ultrastructural identification, tissue reactions and immunological analysis. Aquaculture International, 32(1):937–938, February 2024. CODEN AQINFS. ISSN 0967-6120 (print), 1573-143X (electronic). URL https://link.

springer.com/article/10.1007/s10499-023-01182-9. See [2687].